### v8

# Consumption Smoothing in Metropolis: Evidence from the Working-class Households in Prewar Tokyo

Abstract
I analyze the risk-coping behaviors among factory worker households in early 20th-century Tokyo. I have digitized a unique daily longitudinal survey dataset on household budgets to analyze how consumption is impacted by idiosyncratic shocks. I find that while the households were so vulnerable that the shocks impacted their consumption levels, the income elasticity for food consumption is relatively low in the short-run perspective. The result from mechanism analysis suggests that credit purchases with local retailers smoothed short-run food consumption. The event-study analysis using the adverse health shock as the idiosyncratic income shock confirms the robustness of the results. I also find evidence that the misassignment of payday in data aggregation results in a systematic attenuation bias due to measurement error in the standard consumption smoothing regressions.
Keywords: consumption smoothing; credit purchases; daily budget data; local retailers; measurement error; risk-coping strategy; risk-sharing
JEL Codes: E21; N35
1 Introduction
Economic theory illustrates the mechanism behind the smoothing of consumption in complete markets (Arrow and Debreu 1954). A series of empirical studies have tested the smoothness of consumption among households in various economies under systematic empirical specification. Evidence shows that markets in contemporary developing countries are often imperfect.111 The early representative literature includes, for example, Townsend (1994), Townsend (1995), Ravallion and Chaudhuri (1997), Grimard (1997), and Fafchamps and Lund (2003). Morduch (1999), Townsend (1995), and Dercon (2004) provide reviews of the earlier studies. Attanasio and Pistaferri (2016) and Meyer and Sullivan (2023) provide recent discussions. The households are vulnerable to risk, and the idiosyncratic shocks impact the levels of their consumption because shocks are not fully insured in the markets. Therefore, consumption smoothing is an essential agenda for mitigating poverty, especially in developing economies. [^2] Smoothness of consumption is also an important factor in human capital formation. For example, Foster (1995) shows that children’s growth patterns are influenced by the inefficiency of the credit markets using the floods of 1988 in Bangladesh as a shock. Rose (1999) investigates the relationship between consumption smoothing and excess losses of girls, revealing that the favorable rainfall shocks in early life increased the survival probabilities of the girls in rural India. Gertler and Gruber (2002) also offers evidence that the smoothing of consumption against health shocks is fairly imperfect in Indonesia.
Historical economies offer experimental environments to investigate the means and degrees of consumption smoothing under less developed societies. Lessons from historical institutions in developed economies shall provide meaningful insights for today’s developing economies. Despite this advantage, little has been known about historical consumption smoothing behaviors, partially because of the lack of longitudinal household budget datasets. This study digitized a unique daily longitudinal budget survey on working-class households conducted in Tokyo around 1919 and employed systematic econometric methods to explore the consumption behaviors among factory worker households in a representative manufacturing area. The result shows that the full insurance hypothesis is strongly rejected for overall consumption expenditure. The estimated income elasticity is close to that for the rural households in southern India around 1980. This means that the factory worker households in historical metropolis suffered certain consumption losses when they faced idiosyncratic adverse income shocks. However, I also found evidence that while expenses on dispensable categories are likely to be cut in response to the idiosyncratic shocks, expenses on indispensable consumption categories are partially smoothed in the short-run. Mechanism analyses provide suggestive evidence that credit purchases in the local economy played an important role in mitigating adverse idiosyncratic shocks.
The first contribution of this study is to offer a case study of a metropolis in a developing economy. Urbanization is now a general phenomenon in the world, and the shares of urban population in many developing countries, especially in East Asia, are reaching a similar level to the developed countries (Henderson and Turner 2020). This fact raises the need for a deeper understanding of consumption behavior among urban working households in the process of economic development. Despite this, there is still little evidence for the urban worker households in the literature. While Townsend (1995) offers the results for Bangkok in Thailand between 1975 and 1990, a large part of the existing studies focused on the rural economies (Townsend 1994; Ravallion and Chaudhuri 1997; Gertler and Gruber 2002). [^3] Recent field experimental studies also focus on the village economies (Comola and Prina 2023; Macours et al. 2022). A few exceptional studies provide cases on the middle-sized contemporary economies (Skoufias and Quisumbing 2005; Aguila et al. 2017). In historical context, Ogasawara (2024) analyzes the factory workers’ households in Osaka City around 1920, which was roughly half the size of Tokyo City (Online Appendix Table A.1). This study considers a set of homogeneous households in a representative manufacturing area in prewar Tokyo, which provides a valuable empirical setting for testing the risk-sharing behaviors in the historical metropolis. [^4] Focusing on the sample sharing the similar occupational and regional characteristics systematically mitigates bias in estimating income elasticity under the standard risk-sharing regressions due to the heterogeneity in the risk preferences (Schulhofer-Wohl 2011; Mazzocco and Saini 2012). I found evidence of consumption smoothing using credit transactions with local retailers among working-class households.
Macroeconomics literature has investigated the theoretical mechanisms behind the false rejections of the null of the risk-sharing hypothesis. An essential factor disturbing the valid statistical inference is measurement error in aggregate data (Cochrane 1991; Nelson 1994; Dynarski et al. 1997; Gervais and Klein 2010). This study reveals a specific generating process of systematic measurement errors in the aggregated income and expenditure series in the survey dataset. In my empirical setting, the factory workers had two paydays in a month: one was set in the middle and the other at the end of the month. Hence, the panel dataset based on the calendar months leads to a substantial mismatch between the aggregate income and expenditure series. [^5] If payday is set soon after the initial date in a certain time cell, it does not cause a mismatch in income and expenditure. However, when payday is set just before the initial date in the cell, it leads to measurement errors because the consumption schedule in that cell shall based on the earnings in the previous cell. In other words, this way of matching assumes that the households in a certain cell make their consumption schedule based on future earnings, which is unrealistic. I found that the aggregation ignoring this mechanism causes false rejections in most consumption subcategories due to the attenuation bias. This evidence shows the importance of reflecting the timing of paydays in the aggregation and invokes careful reviews of the results from the aggregate budget datasets.
The second contribution is to add a new analytical perspective to economic history studies. How working-class households overcame their economic hardships during the late 19th to the early 20th century Britain has attracted wider attention. Weights of evidence from these studies have revealed the roles of historical financial institutions and labor supply adjustments in compensating income losses in working-class households. [^6] Horrell and Oxley (2000) found the important roles of the labor supply adjustments among family members in coping with the income losses of the head in working-class households in late 19th-century Britain. Scott and Walker (2012) also found that interwar British working-class households had access to various institutions such as clothing clubs, hire-purchase, contractural savings, and insurance. The households with less earnings were likely to use credit-based institutions like clubs and hire-purchase, whereas those with higher incomes could use the savings institutions. Kiesling (1996) and Boyer (1997) focused on the risk-coping behaviors among the cotton workers households in Lancashire during the recession due to Civil War. Another strand of literature in financial history provided supply-side (lender-side) evidence on the functions of the micro-financial institutions for the workers (O’Connell and Reid 2005). Despite this, there is little evidence of how those risk-coping strategies contributed to smooth household consumption. In addition, the previous studies have utilized the cross-sectional variations in income to estimate the elasticities for the temporary income sources. [^7] For example, Horrell and Oxley (2000) and James and Suto (2011) consider the deviations of actual earnings from the predicted earnings to sort out the shocks. While Scott and Walker (2012) considers the smoothness of expenditure rather than income based on the deviation from the mean expenditure, their analyses employed the cross-sectional variations in the pooled cross-sectional sample. An exceptional study is Saaritsa (2011), which shows the estimated elasticity to temporary income from a fixed effect one-way error component model using a longitudinal household budget survey in 1928 Helsinki. Hence, the estimated elasticity tells the mean tendency that the households with lower (higher) income were more (less) likely to use a specific risk-coping strategy. [^8] In other words, the elasticity is derived from the observations across different households with different income levels. Since risk-coping is a concept about the dynamic behavior of a specific household, however, comparing the different time cells would be more natural to evaluate the roles of the risk-coping strategies. By using the panel dataset, this study implements the stylized econometric method to systematically investigate the households’ smoothing mechanisms to offer comparable estimates.
Another historical contribution is adding micro-level evidence on consumer behaviors during industrialization. Francks (2009) provides a comprehensive view of the long-run consumption history in Japan from the Edo era to the twentieth century. [^9] There are several business history studies analyzing the relationship between consumption and production. A representative work is Gordon (2013), who investigated the expansion of the sewing machine from many aspects, including the production, distribution, and gender social norms in Japan. However, these studies have little interest in the households’ risk-coping behaviors. A few studies investigated on the relationships between household chores and consumption behaviors in interwar households (Tanimoto 2011, 2012). This study employs a different analytical view by focusing on short-run consumption behaviors. Looking into the extent and how the local society could ensure the households’ consumption would provide valuable insights into the consumption history literature.
The rest of this paper is organized as follows. Section 2 overviews the Japanese economy around the First World War (WWI) and reviews Tsukisima, the target area in my analysis. I also summarized the financial institutions that were accessible to working-class households. Section 3 introduces the documents used to construct my panel dataset and discusses the sample characteristics. Section 4 summarizes the empirical strategies and results. Section 5 discusses the roles of credit purchases in the local economy. Section 6 concludes.
2 Historical Background
2.1 Japanese Economy around the First World War
Japan had the foundations of an industrial nation by WWI. Between 1885 and 1915, real GNP grew annually by 2.6%, while population increased at 1.1%. [^10] Given that economic growth is faster than population growth, this implies that modern economic growth had started in this period as defined by Simon Kuznets (Miyamoto 2008, pp. 60; 64). In 1915, the percentage of the employed population in the primary, secondary, and tertiary industries were 63%, 20%, and 18%, respectively. Based on domestic genuine production, these industries accounted for just under 30%, over 30%, and approximately 40%, respectively.
Soon after the outbreak of WWI, exports from European countries declined. Japan increased exports to the Asian market on behalf of those European countries and exported munitions and foodstuffs to the European Allies, leading to a current account surplus. Given the declines in the export capacity of the European countries, import substitution had progressed. This expands the domestic market and leads to a rise in the machinery, metal, and chemical sectors. A worldwide shortage of ships also stimulated the development of shipbuilding, which proliferated as a leading industry. The development of the shipbuilding industry stimulated demand for the machinery and steel industries. Through this process of heavy and chemical industrialization, the total number of enterprises increased. [^11] Descriptions about the process of heavy chemical industrialization are based on Nakamura (1971, p. 128), Nakamura and Odaka (1989, p. 10), and Sawai and Tanimoto (2016, p. 250). Year-on-year growth in the number of incorporated firms was 17.8% in 1918, 22.3% in 1919, and 13.5% in 1920. These are the top three years of the 34 years from after the Russo-Japanese War to 1940 (Nakamura and Odaka 1989, p. 22). Figure 1 shows that the GNP growth from 1915-19 was 11.5% in nominal terms (4.4% in real terms), the highest growth throughout the 1910s and 1920s.
This rapid industrialization impacted the Japanese economy in many ways. Due to export-led economic growth, large inflows of gold increased domestic currency balances, causing prices to soar. Although wages began to rise around 1916 as a response to rising prices and labor demand due to the increased number of firms, wage increases failed to keep pace with soaring prices. By the end of the war, the number of labor disputes grew sharply, and the labor movement flourished. Furthermore, the rise of business accelerated urbanization as the rural population moved to cities in search of job opportunities. The share of large cities (with people of 50,000 or more) as a percentage of the total population was less than 10% in 1889 but increased to nearly 16% in 1920 (Online Appendix Figure A.1a). The growth rate was exceptionally high during the WWI period, indicating that structural changes through the war accelerated the rise of the urban population. Urbanization has increased public investment in large cities, especially in construction investment in social capital such as roads and water supply systems. [^12] Descriptions on the export-led growths, urbanization, and investments are from Nakamura and Odaka (1989, pp. 21–23) and Nakamura (1971, pp. 142; 147).
Exports in the 1920s were lower than before because of the appreciation of the yen against the real exchange rate due to the accumulation of domestic and foreign species during WWI. Since the current account deficit was settled through the disbursement of foreign species owned by the government and the Bank of Japan, there was no reduction in the domestic money supply despite the sizeable current account deficit. This resulted in the domestic prices, especially prices of nontradable goods, being kept high. The first half of the 1920s saw a temporary decline in industrial investment as the capital investment boom of the WWI period came to a halt under declining demand. [^13] On the other hand, construction investment by the private sector, including electric power and private railways, increased (Nakamura 1971, p. 145). Descriptions about the post-war economy in this paragraph are from Sawai and Tanimoto (2016, pp. 253–254).
Figure 1: Gross National Product per Capita and Personal Consumption Expenditure (in yen)
Note: This figure shows the gross national product per capita (solid line) and personal consumption expenditure (dashed line) between 1910 and 1926. The prices are deflated using the CPI in 1934-1935 price.
Sources: Created by the author using Ohkawa et. al. (1974, p. 237).
As described, there were short-term business cycles, such as the boom caused by WWI and the subsequent recession. However, taken as a whole, Japan achieved a high economic growth rate relative to other countries in the early 20th century and before WWII. As shown in Figure 1, the growth rates of personal consumption expenditure were particularly high during and immediately after WWI, with average growth rates reaching 112% in 1915-19 and 118% in 1920-24.141414Total consumption as a percentage of the GNP remained above 80%. Although this level is not apparent among other countries (Kuznets 1962), the growth rate of real personal consumption per capita was relatively high: from 1875 to 1939, the average growth rate of personal consumption was 1.36%, and the rates never turned negative except during the Russo-Japanese War period. This was a high level even by international standards. Given the rapid urbanization, therefore, it is plausible to focus on the consumption behavior of urban households when considering the Japanese economy in this period. However, progress has yet to be made. [^15] Descriptions on the personal consumpsion expenditure are from Nakamura (1971, pp. 2–5; 27; 30–31).
2.2 Tsukishima: A Representative Manufacturing Area
The population in Tokyo City has been growing since the end of the 19th century, and industrialization due to WWI spurred its growth. Tokyo City’s population share increased from 2.5% in 1889 to approximately 4% by 1920.161616In 1920, there were more than 10,000 municipalities in Japan, suggesting how huge Tokyo was. There was a difference of about 1 million between Tokyo and Osaka, the second-largest city by population. Online Appendix Figure A.1a summarizes the trends in the number of people in the top five cities between 1889 and 1920. Throughout the 1910s, the percentage of industrial and commercial workers increased in Tokyo City, building the city as a gigantic commercial and industrial place (Tanimoto 2002, p. 8). Column 1 of Panel A in Table 1 shows that nearly half of the male workers in Tokyo worked in the industrial sector as of 1920.
Figure 2: Administrative Wards in Tokyo City
Notes: The name in each lattice indicates the administrative ward in Tokyo. The gray ward indicates the Kyobashi Ward and an island in the Kyobashi Ward indicates Tsukishima. The border surrounded by Asakusa, Honjyo, Nihonbashi, and Fukagawa wards includes the Sumida River. The borders of each ward are based on the administrative district in 1920. Geographical coordinate system is based on JGD2000/(B, L).
Source: Created by the author using the official shapefile (Ministry of Land, Infrastructure, Transport and Tourism, database).
Figure 3: Factories in Tsukishima
Notes: Blocks in green and light green indicate the shipyards. Blocks in red and pink indicate the factories in the machinery sector. Blocks in blue and light blue indicate the factories in the other manufacturing sectors. Blocks in brown are the warehouses. Small circles colored in black and white show the workplaces in the machinery sector and the other manufacturing sectors, respectively.
Source: Department of Health, Ministry of the Interior 1923c, second map. The tone was adjusted by the author using Adobe Photoshop 24.7.0.
The subject of this paper is Tsukishima, an industrial area located southeast of Kyobashi Ward in Tokyo (Figure 2). [^17] The base of the island was reclaimed during the Edo and Meiji periods from a delta at the mouth of the Sumida River. Online Appendix A.3 summarizes the development history of Tsukishima in detail. Column (2) of Panel A in Table 1 shows that Kyobashi had an industrial composition similar to that of Tokyo City in 1920. Column (3) confirms that Tsukishima is a representative industrial area in the city; approximately 68% of male workers were employed in the manufacturing sector. In the same year, there were 214 factories on this island, and approximately 80% of them belonged to the machinery and related sectors (Department of Health, Ministry of the Interior 1923a, pp 385–389). [^18] According to the census of 1920, the number of workers in the metal and machinery equipment manufacturing industry ($27.1$%) was the highest compared to other industries in the entire city of Tokyo. The same tendency was observed in Kyobashi Ward ($35.5$%). Although similar statistics are not available for Tsukishima, this supports the evidence that Kyobashi was not an area with a unique industrial pattern (Statistics Bureau of the Cabinet 1929c, pp. 85–86; 108–109)). Figure 3 shows the spatial distribution of factories in Tsukishima as of June 30, 1918. Tsukishima had many machine factories and shipyards, and the workers’ residences were clustered there. [^19] For example, a representative shipbuilding plant was located on Ishikawajima, an island in the north of Tsukishima (Figure 3). Ships were strongly influenced by the import substitution that was underway at the beginning of the 20th century, and domestic production of ships was progressing (Section 2.1). The production of a hydraulic 80-ton hardened crane by Ishikawajima Shipyard in 1917 is an example of the domestic output of a production good that had previously relied on imports (Nakamura and Odaka 1989, pp. 10; 32). The focus of this study is the consumption behavior of these worker households in the industrial sector.
Table 1: Industrial and Family Structures in Tsukishima
Panel A: Industrial structure
Name of survey (1) 1920 Population census (2) 1920 Population census (3) 1920 Population census
Survey area Tokyo City Kyobashi Ward Tsukishima
Survey subject Complete survey Complete survey Complete survey
Survey month and year October 1920 October 1920 October 1920
Agriculture 1.0 0.6 0.4
Fisheries 0.1 0.3 0.8
Mining 0.4 0.4 0.8
Manufacturing 44.5 45.2 67.7
Commerce 32.4 34.0 16.9
Transport 7.6 10.2 6.1
Public service and professions 11.3 10.2 4.6
Housework 0.3 0.1 0.2
Other industry 2.4 2.0 2.5
Panel B: Demographic structure
Name of survey (1) 1920 Population census (2) 1920 Population census (3) 1920 Population census
Survey area Tokyo City Kyobashi Ward Tsukishima
Survey subject Complete survey Complete survey Complete survey
Survey month and year October 1920 October 1920 October 1920
Average household size (in person) 4.6 4.7 4.3
Share of married males (%) 38 36 40
Sex ratio (males/females) 1.2 1.2 1.3
0–13 years 1.1 1.1 1.1
14–19 years 1.5 1.7 2.2
20–29 years 1.3 1.4 1.6
30–39 years 1.2 1.2 1.3
40–49 years 1.2 1.2 1.3
50–59 years 1.1 1.1 1.2
60+ 0.8 0.8 0.9
Average age of males 25.3 25.3 25.1
Notes:
Panel A: This panel summarizes the industrial structures based on the occupations of male workers measured in the 1920 population census. Each share is calculated as the number of males worked as regular workers in each sector divided by the total number of male workers (in percentage). All the figures do not include the number of unemployed (mugyō sha) males.
Panel B: The average household size is the number of people divided by the number of households. The share of married males is the number of married males living with/without spouses divided by the total number of males (in percentage). The average age of males is calculated using the tables for population by age group reported in the census. Since the tables are open-ended, the class for over 60 years old is rounded as $64.5$ following the range of the former category (i.e., 50–59). All the figures in this panel do not include a small number of quasi-households (jyun setai), which include the person, such as the students in dormitories and the patients in hospitals.
The figures for Tsukishima listed in column 3 of both panels are based on the total sum of statistics of all the blocks (chōme) in Tsukishima (i.e., from Tsukudajima to Tsukishima dōri jyūni chōme).
Sources: Columns 1 and 2 of Panel A use the Statistics Bureau of the Cabinet (1929c, pp. 18–19). Columns 1 and 2 of Panel B use the Statistics Bureau of the Cabinet (1929c, pp. 38–43); Tokyo City Office (1922b, pp. 2–3; 42–46). Column 3 of Panel A uses the Tokyo City Office (1922c, pp. 42–47). Column 3 of Panel B uses the Tokyo City Office (1922a, pp. 26–31); Tokyo City Office (1922b, pp. 2–3; 42–46).
To look more closely at the characteristics of working-class households in Tsukishima, Panel B in Table 1 summarizes the population and occupational statistics from the census. First, the average household size in Tsukishima is $4.3$, similar but slightly smaller than the values in Tokyo City ($4.6$) and Kyobashi Ward ($4.7$). The percentage of the married male population is $40$%, which is in the similar range of the figures of Tokyo ($38$%) and Kyobashi ($36$%). The average age of men in Tsukishima is $25.1$, slightly lower than the figures for Tokyo and Kyobashi. While the sex ratio by age group in Tsukishima shows a similar trend to that of Tokyo and Kyobashi, the ratios for the 1414141419191919 and 2020202029292929 age bins are higher than those. Overall, these statistics reflect the fact that Tsukishima is a typical industrial area and many young male workers were employed there. The slightly smaller household size in Tsukishima can be attributed to the relatively large number of young factory worker households. The percentage of married men in Tsukishima suggests that a certain proportion of the households in their early 20s might have been couples before they started raising children.
Finally, I briefly overview the lives of the factory workers’ households in Tsukishima. Factories and workers’ residences were scattered throughout the island. [^20] Tsukishima is composed of Ishikawajima island at the northern end, the adjacent Tsukudajima and Shin-tsukudajima islands, and the rectangular Lots 1–3 in the south (Online Appendix Figure A.3). Specifically, factories and worker residences were scattered throughout Shin-tsukudajima and Lots 1 and 2, whereas Lot 3, a relatively new reclaimed site completed in 1913, was undeveloped. The factory workers had lived in similar tenements and used the commercial stores and restaurants in the central lot. [^21] Online Appendix A.4 summarizes the housing in Tsukishima in detail using a housing survey. Regarding the stores, there were 209 stores selling daily necessities, 109 selling clothing, 74 selling tools, and 120 selling food, drink, and entertainment. As for medical personnel, there were eight doctors, four dentists, eight obstetricians, and 14 drug stores. There were also ten bathhouses (Department of Health, Ministry of the Interior 1923a, pp. 41–47). There were two postal offices, 13 pawnshops, and five moneylenders, and those were scattered throughout the island (Online Appendix Figure A.2). The main routes to the mainland were the wooden Aioi Bridge (Aioi bashi) and three ferries (watashi bune). [^22] See Online Appendix Figure A.3 for these routes. The average daily total number of people using the ferries in 1918 was reported to be $53,318$, accounting for more than twice as many people as the Tsukishima population. [^23] See Department of Health, Ministry of the Interior (1923a, p. 30). The number of people and households of Tsukishima at the end of December 1918 is reported to be $24,399$ and $5,562$, respectively (National Police Agency 1920a, pp. 50–51). These descriptions illustrate Tsukishima as a busy machine industry area during the industrialization.
2.3 Risk-Coping Institutions
Working-class households had been exposed to several risks that led to losses of income at that time. Generally, social insurance, such as health, accident, and unemployment insurance, can help mitigate the idiosyncratic income shocks on the workers. [^24] Gertler and Gruber (2002) shows the impacts of illness on consumption smoothing behavior in early 1990s Indonesia and argues the importance of public support for the illness. Kantor and Fishback (1996) show that the installation of workers’ compensation reduced the precautionary savings in the interwar U.S.. Gruber (1997) provides evidence that unemployment insurance smooths individual consumption as the case of the U.S. from 1968–1987. However, comprehensive social insurance systems did not exist in the early 1920s in Japan. The former health insurance was enacted in 1927, and unemployment insurance had not been established throughout the prewar period. In addition, the Factory Act enacted in 1916 did also not specify compensation for the sickness of workers. Therefore, the workers had basically no public support for their illness, injuries, and joblessness. Instead, they might have used private financial institutions to deal with the idiosyncratic risks.
Several financial institutions were indeed available for working-class households. A representative savings institution for households was the postal savings (yūbin chokin). In Tokyo City, roughly $1.4$ million people had postal savings accounts, covering approximately $67$% of the entire citizens in 1920 (Online Appendix Table A.2). The average savings amount per capita in Kyobashi Ward was $44.5$ yen (Tokyo City Office 1921, pp. 892–893), which is less than one month’s earnings of the skilled-factory workers (Section 3.2). Savings bank (chochiku ginkō) is another famous savings institution among the working class in the 1910s (Ito and Saito 2019, pp. 79–81). The number of people with savings accounts was roughly $1.4$ million, accounting for approximately $58$% of total people in Tokyo (Online Appendix Table A.2). However, $84.8$% of all the depositors were the commerce and miscellaneous workers, and the manufacturing sector shared only $11.5$% in Kyobashi Ward (Online Appendix Table A.3). This reflects the fact that savings bank was an institution designed for lower-income workers than factory workers (Tokyo Institute for Municipal Research 1925b, p. 81). This is consistent with the fact that Tsukishima did not have any savings bank branches, while there were two postal offices (Section 2.2). Although another savings institution was the mutual loan association (mujin), it had been used among small business owners and merchants for running their businesses (Online Appendix A.2). [^25] Generally, the ordinary bank in prewar Japan was the institution for providing large loans to businesses (Teranishi 2011). There was no bank in Tsukishima, as workers did not borrow money from banks to make ends meet. The official report states that “a commercial bank is not necessary in a laboring area such as this Tsukishima, and even for large entrepreneurs such as factory owners, there is no particular need to establish a bank for their business transactions, and transactions with banks in other areas are sufficient” (Department of Health, Ministry of the Interior 1923a, p. 48).
There also be a few lending institutions. Credit purchases (kakegai) in retail shops were the most popular institution. Although the systematic statistics on credit purchases are unavailable, the THBS households often used credit purchases in various retail shops for rice, fish, vegetables, firewood, and charcoal. [^26] Cooperative societies were also available for worker households to buy daily commodities, albeit they were not lending institutions. There were 26 cooperative societies in Tokyo prefecture in 1924, of which 23 were organized by the citizens and workers (Central Federation of Industrial Associations 1925, p. 57–59). However, while the two largest societies had roughly $4,000$ members, the others were generally small societies with less than 100 members. This figure accounts for only a few percentage points in the workers in the city, given that the number of male workers was more than 768 thousand. Most of the THBS households had not used the cooperatives. I will describe this institution in detail in Section 5. The pawnshop (shichiya) was another lending institution. The accessibilities were higher than the other lending institutions: [^27] The number of pawnshops in Tokyo were $1,334$ in 1918 and $1,261$ in 1919, respectively. The total number of cases was $8,226,883$ and $7,573,406$ each year. These figures are more than three times greater than the entire population in Tokyo City. there were $13$ pawnshops in Tsukishima, more than twice the large number of moneylenders (Section 2.2). However, the core users of the pawnshops were the low-income working-class households rather than the factory workers (Tokyo City Social Affairs Bureau 1921, p. 9). According to the statistics on the pawnshop users in Tokyo prefecture in 1923, the most representative users were the day laborers and workers classified as “miscellaneous”, accounting for $63.4$% of all users. In contrast, those for agriculture, commerce, and manufacturing sectors were only $1.7$, $19.5$, and $15.6$%, respectively (Tokyo Institute for Municipal Research 1926, p. 25). The inexpensive article means that pawnshops were used to deal with very short-run necessities only. Among the pawnshops of Kyobashi, for example, the average amount per case was $7.8$ yen, which accounts for $12$% of the monthly disposable income of the THBS households (Panel A of Table 3). In this light, the pawnshop is a small lending institution for low-income working-class households who might have faced credit constraints. Lenders did not need to screen the borrower’s credit, and the borrowers did not worry about incurring heavy debt as the main articles were inexpensive clothes (Shibuya et al. 1982). In addition, the interest rates were regulated by the Pawnbroker Regulation Act of 1895, and the average redemption rate was substantially high under the lower interest rates. [^28] This means that the average redemption rate was substantially high: it was approximately $94$% in 1920 among $88$ pawnshops in Kyobashi ($401,222$ cases/$427,265$ cases) (Tokyo City Office 1922b, pp. 888–889).
Other lending institutions are money lenders (kinsen kashitsuke gyō) and credit unions (shinyō kumiai). Generally, moneylenders were used by business owners with very high-interest rates (Shibuya 2000, pp. 184; 248). Similarly, the coverage of the credit unions was substantially low, say a few percentage points of the male workers in the manufacturing sector (Online Appendix A.2). Informal gifts may be another type of risk-coping institution, although systematic statistics are not available. A household survey on the 185 factory-worker households in Tokyo suggests that the share of average monthly gifts in total income was approximately $3$% in November 1922 (Social Affairs Division 1925, pp. 58-59). While this means that the gifts comprised a few percentage points of the monthly income, it is unclear whether these informal income transfers had functioned as mutual aid.
Finally, life insurance was also not designed to compensate for temporary income losses. Generally, life insurance companies provided death insurance, meaning that temporary income reductions were not covered by insurance. Moreover, they did not target the working-class as customers (Usami 1984, pp. 87–88). Although postal life insurance (kani seimei hoken) operated by the government started in 1916, it only provided whole life insurance and endowment insurance. Thus, it also never covered the health shocks nor layoffs (Usami 1984, pp. 102–107).  [^29] There were also several employment agencies that introduced jobs to low-income job-seekers in Tokyo. However, a large part of them was unscrupulous and not commonly used at that time (Machida 2016).
To summarize, while there were several institutions, the available institutions for the factory workers were few in the practical sense. For savings institutions, the postal savings might have been the best available device for Tsukishima workers. Regarding lending institutions, credit purchase is the plausible device among factory workers. While pawnshops were accessible financial institutions, they might not provide sufficient money as the articles were usually inexpensive. The factory workers in Tsukishima might have combined these devices to deal with the idiosyncratic income shocks.
3 Data
This section is designed to understand the characteristics of the data used in the study. I give an overview of the Tsukishima Survey in Section 3.1. I then utilize the statistics from the complete surveys in Section 3.2 to assess the representativeness of the sample households. In Section 3.3, I explain my data aggregation method to overcome the potential measurement error issue. The variables used, the trend in income and expenditure, and variations in idiosyncratic shocks are described in Section 3.4.
3.1 Tsukishima Household Budget Survey
The Tsukishima Survey was the first urban social survey conducted in Japan by the Ministry of Home Affairs around 1919. The purpose was to investigate the status of the lives of urban worker households amid the rapid progression of industrialization throughout WWI. It was designed as a field survey consisting of multiple survey items, including the budget survey (hereafter THBS). [^30] Online Appendix B.1 summarizes detailed information about the Tsukishima Survey. The THBS was the first household survey in Japan to use the budget book method and is considered the prototype for subsequent household surveys in Japan (Sekiya 1970, p. 43). The Ministry of Interior published an official report on the Tsukishima Survey in 1921, which contains some aggregated information on the target households. However, household-level budget information is required to analyze consumption smoothing behaviors.
Fortunately, the household budget book (kinsendeiri hikaechō) gathered in the THBS is preserved in the archives of the Ohara Institute of Social Research (hereafter, OISR). [^31] The author is permitted to access the materials related to the Tsukishima Survey by the OISR: Archives of the Tsukishima Survey (THBS, unreleased). While most of the sources are unreleased, the official reports of the survey published by the Ministry of Home Affairs are included in the archives of Gonda Yasunosuke at the OISR and are available to the public (The OISR, Archives of Gonda Yasunosuke (7-2; 7-3; 7-4)). These budget books were transferred to the OISR by Iwasaburo Takano, who was responsible for the Tsukishima Survey as a member of the 7th Section of the Health and Hygiene Research Committee of the Ministry of Home Affairs and became the first director of the OISR. [^32] Iwasaburo Takano was a professor at Tokyo Imperial University and a specialist in social statistics and social policy. He emphasized gathering statistics about workers through social surveys and is known as a researcher who significantly contributed to the formation of social survey methodology in Japan. His contribution to the Tsukishima Survey is summarized in detail in the Online Appendix B.1.
I corrected and digitized all the 40 THBS budget books for the factory worker households that were included in the official report. [^33] The archives of the OISR also reserve several other household budget books of factory worker households. However, these budget books are generally in poor record keeping. Therefore, I do not use these incomplete books in this study. The characteristics of the remaining household budget books are summarized in detail in the Online Appendix B.5. During the study period, each household recorded their income and expenses daily in a budget book. Notably, a survey office was set up in Tsukishima, where a few surveyors lived and were closely connected with the survey households. Furthermore, weekly administrative meetings were held to discuss the progress of the survey (Miyoshi 1980, p. 38). This confirms that the information in these budget books is reliable and that the THBS budget dataset ensures a quality that allows quantitative analysis. [^34] Online Appendix B.1 provides finder details on how the Tsukishima Survey had been conducted carefully as a social survey project of Takano.
3.2 Sample Characteristics
Analytical Sample
The estimation strategy used for testing consumption smoothing and risk-coping behaviors requires the panel structures for aggregate income and expenditure (Section 4). Among the 40 THBS households, there are five cross-sectional units in our semi-monthly data. [^35] Section 3.3 provides details on the definitions of aggregate income and expenditure. I excluded these five cross-sectional units, leaving 35 THBS households. Next, a household with incomplete income information, which is necessary for my analysis, is removed. Finally, I exclude a household for which no information on payments to the credit purchases was recorded. Consequently, my THBS sample for the semi-monthly series includes 33 households with panel structures from 12th January to 11th November 1919.363636I use adjusted monthly panel dataset for subanalysis in Section 4. Following the same trimming steps, my adjusted monthly panel dataset includes 26 households for the same period. Note that several units have only three semi-month cells, say less than two adjusted months. These units have panel structure in the semi-monthly panel, but not in the adjusted monthly panel. Online Appendix B.3 summarizes the finer details of the original data structure and trimming.
Before assessing the representatives of the THBS sample, I test the potential selection bias due to attrition of the units. The panel units with shorter (longer) time-series observations might have different preferences on the consumption than those with longer (shorter) time-series observations. In other words, the households with shorter panels should have similar family size characteristics to those with more extended panels. To test this potential issue, I regressed an indicator variable for the shorter panel units on the family size variables. Online Appendix Table B.2 summarizes the results. All the estimated coefficients on the covariates are close to zero and statistically insignificant, and the Wald statistics support the null results. This result is robust to using different thresholds such as first quantile, median, and third quantile. This supports the evidence that the household’s preference is less likely to be correlated with unbalancedness.
Representativeness
Table 2: Assessing Representativeness: Tsukishima Household Budget Survey
Panel A: Manufacturing sector
Name of survey (1) 1919 Statistics of (2) 1920 Tsukishima Factory (3) The THBS
National Police Agency Survey
Survey area Tsukishima Tsukishima Tsukishima
Survey month and year December 1919 November 1920 1919
Survey subject All factories with All factories with # of 33 households
# of workers $\geq 15$ workers $\geq 15$ workers $<15$
Textile 0.2 2.9 0.0 0
Machinery 87.4 85.3 91.8 93.9
Chemical 1.1 5.9 5.5 0
Food 1.2 1.5 0.0 3.0
Miscellaneous 10.2 4.4 2.7 3.0
Observations $7,647$ workers 68 factories 146 factories 33 heads
Panel B: Family structure
Name of survey (1) 1920 Population (2) 1919 Statistics of (3) The THBS
Census National Police Agency
Survey area Tsukishima Tsukishima Tsukishima
Survey month and year October 1920 December 1919 1919
Survey subject Complete survey Complete survey 33 households
Average household size 4.3 4.3 4.3
Sex ratio 1.3 1.2 1.2
Household size (share)
       1 6.0 0
       2 16.7 (18.7) 15.2
       3–5 50.8 (57.1) 57.6
       6–8 21.5 (24.2) 27.3
       9+ 4.9 0
Pearson $\chi^{2}$ statistic $p$-value 0.948
Panel C: Monthly earnings
Name of survey (1) The THBS (2) Manufacturing Census
Survey area Tsukishima Tokyo City
Survey/equivalent year 1919 1919
Survey subject THBS heads Male factory workers
(Ave. age: 33.2) in machinery factories
(Age range: 30-40)
Average monthly earnings 59.1 (median = 56.2) 56.0
95%CI [53.7, 64.5]
Notes: Panel A: Column 1 shows the share of factory workers to the total factory workers in each sector (in percentage). The number of factory workers includes female and male factory workers in the 68 factories, with 15 workers and more. Column 2 indicates the share of factories to the total number of factories in each sector by factory size (in percentage). Column 3 summarizes the manufacturing sectors for the heads of the THBS. Columns 2 and 3 use the classification definition in the Statistics of the National Police Agency (column 1). Following this definition, several “woodworking” factories measured in the Tsukishima Factory Survey are included in the machinery sector in column 2, as it consists of shipbuilding and wooden pattern factories. The report also supports that the wooden patterns are used in the casting process and shall be regarded as one of the machinery sector (Department of Health, Ministry of the Interior 1923a, pp. 388–389).
Panel B: The first and second rows in columns 1 and 2 show the average household size and sex ratio measured in the 1920 Population Census and Statistics of the National Police Agency of 1919, respectively. The third to eighth rows in columns 1 and 3 show household size distributions in the census and the THBS sample, respectively. The figures in parentheses in Column (1) are the percentage share for family size bins between 2 and 8 people. The ninth row in column 3 shows the $p$-value from the Pearson $\chi^{2}$ test for the equality of the household size distributions between 2 and 8 people. The share in each household size bin measured in the census is used as the theoretical probability in calculating the test statistic (Online Appendix B.5).
Panel C: Column 1 lists the average monthly earnings of the THBS heads calculated using the adjusted monthly panel dataset. Of the 33 THBS households, information from 26 households for which adjusted monthly income can be traced back is used in this panel to follow the unit in the manufacturing census. Comparison based on the average semi-monthly income from all 33 households shows a similar result. Gaussian-based 95% confidence interval (CI) using bootstrap standard error is reported in the brackets. Column 2 shows the average monthly wage of the male factory workers in the skilled age range (30-40 years old) in machinery factories in Tokyo City. This figure is calculated using the average monthly wage and average monthly ancillary wage measured in the manufacturing censuses. All the wage statistics are deflated using the CPI for cities provided by Ohkawa et al. (1967, p. 255). Online Appendix B.6 summarizes the finer details of the process of this wage estimation.
Sources: Panel A: Column 1 uses the National Police Agency (1920b, p. 236). Column 2 uses the Department of Health, Ministry of the Interior (1923a, p. 385–386). Column 3 uses the Department of Health, Ministry of the Interior (1923a, p. 154). Panel B: Column 1 uses the Tokyo City Office (1922a, pp. 26–31; 262–283); Tokyo City Office (1922b, pp. 2–3; 42–46). Column 2 uses the National Police Agency (1920b, p. 53). The figures in column 3 are calculated by the author using the THBS dataset. Panel C: Column 1 uses the THBS dataset. Column 2 uses the Tokyo City Statistics Division (1926a, pp. 16; 244–245); Tokyo City Office (1921, pp. 726–757).
Next, I assess the representativeness of the THBS sample in Tsukishima using the manufacturing and population censuses,
Panel A of Table 2 summarizes the share of male workers in the manufacturing industry by sector. Statistics of the National Police Agency (hereafter SNPA) focusing on factories with more than or equal to 15 workers show that roughly 90% of male workers had been engaged in the machinery sector in 1919 (Column 1). Column 2 summarizes the statistics in the manufacturing census conducted in the Tsukishima Factory Survey. The proportion of machine factories is $85$% in larger factories and $92$% in smaller factories, indicating that most of the male workers worked in the machinery sector.373737 I have confirmed that the paternal occupational statistics for the primary school students in Tsukishima provide materially similar results (Online Appendix A.5). In Column 3, I confirm that the THBS heads’ occupations show a similar proportion: $94$% worked in the machinery sector. In fact, the report states that “It is not too much to say that most of the industry in Tsukishima is machinery. Therefore, to describe the labor situation in Tsukishima, it is sufficient to describe its machinery industry” (Department of Health, Ministry of the Interior 1923a, p. 389). Figure 3 illustrating a number of machinery factories supports this statement. As for the other sectors, the THBS includes a few heads who worked in food and miscellaneous sectors, and the proportions are similar to those listed in column 2.
Panel B of Table 2 summarizes the family structure. The average household sizes are identical in the Population Census (Column 1), SNPA (Column 2), and THBS (Column 3). The sex ratio is also similar across the surveys (1.21.21.21.21.31.31.31.3). The number of people in the THBS households ranges from 2 to 8, meaning the THBS sample does not include single or very large households. [^38] Those family households with 2–8 people cover most of the households in Tsukishima because the share of single and vary large (9+ people) households was roughly $10$% (Column 1 of Panel B). In other words, the THBS mainly focuses on the consumption behavior in couples and households with a few children. For this household size range, the Pearson $\chi^{2}$ test does not reject the null of equality to the household size distribution measured in the census ($p$-value = $0.948$), meaning that the THBS sample reasonably approximates the distribution of family households in Tsukishima. Online Appendix B.5 illustrates the household size distributions in finer detail. This confirms the validity of the statement in the report that the THBS households can be “regarded as representative of the family form in Tsukishima” (Department of Health, Ministry of the Interior 1923a, p. 145).
The official report of the Tsukishima Survey suggests that roughly nine out of ten factory workers in the skilled-age range were classified as skilled workers. Roughly six of those skilled workers were employed in the middle- to large-scale factories, and three worked in small-scale factories. [^39] Online Appendix A.5 summarizes this point in detail using the complete survey on the primary school students in 4th to 6th grades in Tsukishima. Given this, the THBS was designed to investigate the budgets in skilled workers’ households. [^40] Shortly, Iwasaburo Takano, a survey director, regarded that the skilled workers in the modern industrial sector will constitute the core of the labor force, labor movement, and citizens. The finer details in the background of this survey design are summarized in Online Appendix B.1. The average THBS head’s age is 33, which is indeed in the skilled workers’ representative age range in the machinery factories (Kitazawa 1924). [^41] Specific occupations are not available for most heads because they usually list “factory workers” in the occupation section. However, sub-categories available at a few heads indicate the fact that they were lathe operator and finisher, the typical skilled workers in machinery factories. Online Appendix A.6 summarizes the works in the machinery factories in Tsukishima. Column 1 of Panel C in Table 2 indicates that the average monthly earning of the THBS heads is $59$ yen (median = $56$ yen). In Column 2, I calculated the average monthly earning of male machinery factory workers in the skilled-age range using the manufacturing censuses, which shows similar monthly earnings ($56$ yen). [^42] Shortly, this figure is calculated using the average daily wage of male factory workers aged 30-40 and average annual working days measured in the manufacturing censuses. Online Appendix B.6 provides the finer details of the calculation steps. The average wages for the male factory workers had an inverted-U-shaped distribution with respect to age, taking the greatest figure in the 30s (Online Appendix Figure B.7). This is a similar life-cycle pattern to the factory workers observed in late 19th century England (Horrell and Oxley 2000, p. 42). This provides evidence that the monthly earnings of the THBS heads and skilled workers’ households in Tsukishima are materially similar. To summarize, this evidence indicates that the THBS sample is useful in inferring the mean tendency of those households.
3.3 Aggregation
Measurement Error
(a) Daily panel dataset:
income and expenditure
(b) Daily panel dataset:
net income
(c) Semi-monthly panel dataset:
income and expenditure
(d) Semi-monthly panel dataset:
net income
(e) Unadj. semi-monthly panel dataset: income and expenditure
(f) Unadj. semi-monthly panel dataset: net income
Figure 4: Relationships between Average Income and Expenditure
by Different Time-series Frequencies
Note: Figure 4a4c, and 4e illustrate the time-series plots of the average daily, semi-monthly, and unadjusted semi-monthly income and expenditure, respectively. Figure 4b4d, and 4f illustrate the time-series plots of the average daily, semi-monthly, and unadjusted semi-monthly net income (i.e., income minus expenditure). Figures 4a and 4b shows the daily data between 12th January and 11th November 1919. The (adjusted) semi-monthly series defines the 12-26th for the first and 27-11th for the second half. Figures 4c and 4d show the semi-monthly series calculated using the daily data between 12th January and 11th November 1919. The unadjusted semi-monthly series uses the 15th of each month as a threshold. Figures 4e and 4f show the unadjusted semi-monthly series calculated using the daily data between 15th January and 31st October 1919.
Source: Created by the author using the THBS sample.
It is necessary to summarize the daily budget information into a more extended time range because households need a set of time to smooth their consumption given a realized amount of income. To explore the reasonable way of aggregation, I first overview the daily time series on the total income and expenditure. Figure 4a shows that the head’s income had been paid twice monthly: the semi-monthly wages were paid on the middle (14–16th) and last day (30th or 31st). [^43] The head’s income was the dominant way in earnings because the breadwinner households had been typical at that time (Panel B in Table 3). Importantly, Figure 4a indicates that the households expensed much of their income just on the payday and spent the rest of the payment little by little until the next payday. In other words, the expenditures had never clearly increased just before the paydays. This means that the households had made their consumption decisions (schedules until the next payday) based on the income they obtained. Notice that calendar (semi-)monthly aggregation can not reflect this household behavior because calendar (semi-)month does not place the payday at the beginning of its cell. [^44] For example, a calendar semi-month cell in January, say the cell from 1st to 14th, starts on the 1st but its payday is set on the 14th. This assumes that the aggregate consumption in this semi-month cell is based on the earnings at the end of its cell. This is clearly an unrealistic assumption. Another source of measurement error is the shift of the payday: the payday for the latter half of a month was sometimes set on the first date of the following month. Although this shift did not occur frequently, the calendar month cell shall contain a substantial noise from the shift as it leads to the cells including more (three paydays) and fewer (one payday) payments than normal. Nelson (1994) suggests that this sort of mismatching in the timing of income and expenditure tends to induce the type II error in the test of consumption smoothing.
Adjustment
In order to deal with the mismatch, I aggregated the daily observations between the 12th and 26th for the first half and those between the 27th and 11th of the next month for the second half. [^45] This means that the number of days included in a semi-month ranges from 13 to 16. Note that this heterogeneity in the number of days over different semi-months is a cross-sectional constant. Thus, these differences are entirely captured using the semi-month fixed effects in the regression. This range is based on the fact that almost all of the paydays in the second half of a month measured in the THBS dataset, are 27th–31st. [^46] Note that I do not use the definition of semi-month starting from the 2nd (to the 16th) in each month because this way of stratification places the payday at the end of each semi-month cell. Online Appendix B.7 provides the evidence that while the definition using 2nd in each month as the threshold provides an improved series in income, the expenditure can not chase the income in the right way because the expenses measured under this definition are based on the income in the previous semi-month cell. Again, I confirmed that my preferred definition using the 12th in each month as the threshold offers a clean series of income and expenditure. Figure 4c shows the time-series plots of the income and expenditure from this adjusted semi-monthly panel dataset, whereas Figure 4e illustrates the unadjusted (calender) semi-monthly series using the 15th of each month as the threshold of the semi-month. Clearly, Figure 4e provides very rough measures of income and expenditure due to the measurement errors from the misadjustment of paydays. The most significant improvement in Figure 4c is that the expenditure can chase the income in each month, reflecting the households’ consumption behaviors based on their income measured in the nearly exact timing. Given the improvements in both income and expenditure measures, the net income series shows a much smoother trend in Figure 4d, which are not available under the unadjusted series (Figure 4f). In Section 4.1, I will demonstrate how these measurement errors in the unadjusted dataset attenuate the estimates.
3.4 Data Description
Variables
The THBS budget book has two primary sections: expenditure and income. I digitized records in each category following the categorization used in the official report. Panel A of Table 3 lists the $11$ subcategories for expenditures: food, housing, utilities, furniture, clothes, education, medical, entertainment, transportation, gifts, and miscellaneous. These are used to test the smoothness of the consumptions in each subcategory. To test the roles of risk-coping strategies, I consider five net income categories: savings, insurance, borrowing, credit purchase, and gifts. Net variables are defined as the income minus expenditure in each category; for example, net savings refers to the difference between the withdrawal amount and deposits to savings. I also consider the sales of miscellaneous goods and labor earnings from the family members except for the head to test the contribution of the asset sales and intrahousehold labor-supply adjustments. Panel B of Table 3 summarizes these variables. [^47] Online Appendix B.9 presents the results for the panel unit root tests. The null hypothesis of unit roots in all the panels is rejected at the conventional level for all these variables.
Alterations in family size may shift the household’s preferences (Jappelli and Pistaferri 2017). [^48] In village economies, landholding and asset stocks such as livestock and grain may also be correlated with the household’s risk-coping behaviors (e.g., Udry 1994, 1995). Since this study analyzes urban factory workers’ households, the roles of these assets can be discarded. Fortunately, the archives of the Tsukishima Survey include a set of sheets documenting the changes in the composition of the THBS households (Kazoku idō hyō). [^49] The Ohara Institute for Social Research, Archives of the Tsukishima Survey (THBS, unreleased). I found that the size of households had been stable over time: one household experienced changes in the number of family members due to the birth of a son during the sample periods. This supports the evidence that the preference shifts had rarely occurred among the THBS households. Although family size is included in all regressions to be conservative, it has little effect on the results. [^50] The family size variable shall be omitted in several regressions due to the multicollinearity in the fixed-effects models given its small within variations. Although the number of family members in different age bins is available from the THBS, those family composition variables are also time-constant and are absorbed by the household fixed effect. Panel C of Table 3 shows the summary statistics on the family size variable.
Table 3: Summary Statistics
Panel A: Variables for testing consumption smoothing (in yen)
Semi-monthly Panel Dataset Adj. Monthly Panel Dataset
Level Log Level Log
Mean Obs. Mean Obs. Mean Obs. Mean Obs.
Total expenditure 34.5 289 3.4 289 69.3 124 4.2 124
      Food 10.9 289 2.3 270 21.5 124 3.0 115
      Housing 2.7 289 1.6 134 5.8 124 1.8 105
      Utilities 1.2 289 -0.1 232 2.3 124 0.6 111
      Furniture 0.4 289 -1.6 177 0.8 124 -1.1 99
      Clothes 2.8 289 0.3 248 5.8 124 1.2 113
      Education 1.3 289 -0.0 245 2.6 124 0.6 111
      Medical 1.7 289 0.2 269 3.3 124 0.9 115
      Entertainment 1.3 289 -0.2 255 2.6 124 0.5 114
      Transportation 0.4 289 -1.3 195 0.9 124 -0.7 98
      Gift 3.1 289 0.5 243 6.3 124 1.3 114
      Miscellaneous 0.5 289 -1.3 222 1.0 124 -0.7 111
Disposable income 37.6 289 3.5 278 76.8 124 4.2 123
Panel B: Variables for testing risk-coping mechanisms (in yen)
Semi-monthly Panel Dataset Adj. Monthly Panel Dataset
Mean Std. Dev. Obs. Mean Std. Dev. Obs.
Panel B-1: Savings, insurance, borrowing, and gifts
      Net savings -0.01 3.97 289 0.08 5.78 124
      Net insurance -0.03 6.67 289 0.26 10.15 124
      Net borrowing 0.31 5.03 289 1.16 7.63 124
      Net credit purchase -2.47 11.33 289 -5.05 20.92 124
      Net gifts -2.04 6.11 289 -3.83 9.82 124
Panel B-2: Labor supply adjustments & sales of miscellaneous assets
      Other family member’s earning 3.74 6.11 289 8.21 16.44 124
      Sales of assets 0.02 0.12 289 0.06 0.17 124
Panel B-3: Income variable
Head’s earning 29.43 18.11 289 59.2 29.9 124
Panel C: Family size (people)
Semi-monthly panel dataset
Mean Std. dev. Min Max Obs.
Number of family members 4.3 1.6 2.0 8.0 289
Panel D: Health shocks
Semi-monthly panel dataset
Mean Std. dev. Min Max Obs.
Number of days with illness (head) 0.1 0.8 0 11 254
Number of days with illness 0.2 1.7 0 16 254
(the other family members)
Notes:
Panel A: This panel shows the summary statistics of the variables used to test the consumption smoothing for the 33-and 26-panel units in the semi-monthly and adjusted monthly panel datasets, respectively. Disposable income is the total income minus tax payments.
Panel B: This panel shows the summary statistics of the variables used to test the risk-coping strategies for the 33-and 26-panel units in the semi-monthly and adjusted monthly panel datasets, respectively. Each net income variable is defined as the difference between income and expenses. “Net savings” refers to the difference between the amount withdrawn and the total deposits. “Net insurance” refers to the amount received from insurance minus the expenses paid for insurance. “Net borrowing” refers to the amount of borrowing minus the debt payments. “Net credit purchase” refers to the amount of credit purchases minus credit redeemed. “Net gifts” refers to the total received amounts, including both pecuniary and non-pecuniary gifts, minus the payments for gifts. “Other family member’s earnings” refers to the total income earned by all family members except for the head of the household. “Sales of assets” includes the sale of daily miscellaneous goods such as newspapers and empty bottles.
Panel C: The summary statistics of the family size variable for the 33-panel units in the semi-monthly panel dataset are listed. The statistics for the adjusted monthly panel dataset are not reported because they are materially similar to those for the semi-monthly panels.
Panel D: The summary statistics of the number of days with illness for family members among the 33-panel units in the semi-monthly panel dataset are listed. Since both are lagged variables, the initial time period in all the units is omitted. Similarly, a household with two semi-month observations is excluded. Similarly, .
Sources: Created by the author from the THBS sample.
Trend
Next, I illustrate the time-series figures of the THBS dataset to overview the trend in income and expenditure. Figure 4c shows a weak increasing trend in the average income, which is consistent with the overall macroeconomic trend in the nominal income during this period. This trend in earnings was, however, offset by the steep inflation under the expansions after the war (Section 2.1). In Figure 4d, the deviation of expenditure from income in January must capture the expenses on the goods for the New Year events. The gaps in June to the first half of July may reflect the payments of bonuses. The humps in March may reflect the preparation for the new fiscal and academic year from April, suggesting a higher proportion of temporary income from withdrawals and loans. These support the evidence that the budget data from the THBS households can capture a large part of the macroeconomic trends.
Idiosyncratic Shocks
Online Appendix Figure B.11 illustrates the residuals from the regression in the first difference in semi-monthly income on the first difference in aggregate semi-monthly income. [^51] Let $n_{t}$ be the number of households in semi-month $t$ in the year 1919. For household $i$ in semi-month $t$ cell, the residual is then defined as $\Delta\hat{\upsilon}_{i,t}=\Delta\textit{Income}_{i,t}-\hat{\alpha}\Delta\overline{\textit{Income}}_{.,t}$, where $\overline{\textit{Income}}_{.,t}=\sum_{i=1}^{n_{t}}\frac{\textit{Income}_{i,t}}{n_{t}}$ and $\hat{\alpha}$ is the estimated coefficient. I used the semi-monthly panel dataset to overview the entire distribution of the idiosyncratic shocks because the daily panels include a set of consecutive cells with no income. While the residuals satisfy the zero-average property, they occasionally take large negative and positive values, suggesting that the workers had experienced idiosyncratic income shocks in some cells. Regardless of favorable or adverse shocks, these cells that deviated from mean tendency are helpful in investigating how those households coped with the idiosyncratic shocks. Despite this, it is helpful to focus on the illness as it is the least predictable shock (Gertler and Gruber 2002). Although joblessness might be another shock, the THBS heads rarely changed their occupation during the survey period as they are the skilled workers who were less likely to move to other jobs. [^52] Another possibility is that the THBS heads who changed their jobs might have been omitted from the survey. However, I confirmed that the households with longer and shorter time-series observations have similar family characteristics (Online Appendix B.3). Thus, this omission does not lead to selection bias in terms of the potential vulnerabilities to the idiosyncratic risk. Moreover, the number of dismissals at the Tsukishima factories seems to be small. In fact, the average monthly dismissal rate (i.e., the number of dismissals/workers) from June 1919 to May 1920 at the one factory for which statistics are available from the official report of the Tsukishima Survey was only 1.5% (Department of Health, Ministry of the Interior 1923a, p. 415). In contrast, fortunately, the illness among the THBS family is documented systematically. As explained in Section 2.3, neither the comprehensive health insurance system nor the factory’s compensation for sickness was available. Moreover, the sanitary conditions in the cities were low, leading to higher mortality rates. [^53] See Schneider and Ogasawara (2018) for the overviews on the sanitary conditions in interwar Japan. The official report mentioned that the workers in the machinery factories in Tsukishima had faced the risk of diseases under severe sanitary conditions (Department of Health, Ministry of the Interior 1923a, p. 14).
Considering this, I first provide the overall results from the specification using the changes in earnings as the idiosyncratic income shock in Sections 4.1 and 4.2. I then consider the structure behind the adverse income shocks using the heads’ temporary illness to investigate the households’ consumption and risk-coping behaviors in Section 4.3. Panel D of Table 3 shows the summary statistics on the health shock variables.
Risk Preferences
This study investigates consumption smoothing behaviors among the skilled factory workers’ households in the machinery sector in Tsukishima. This empirical setting is preferable because the different attitudes toward risks among households with different occupations and regions may lead to bias in the standard risk-sharing regressions (Schulhofer-Wohl 2011; Mazzocco and Saini 2012). To assess the homogeneity of the risk preference among THBS households, I tested whether the earnings of the heads who worked in a few specific factories or sectors showed different responses to the aggregate shocks in the spirit of Schulhofer-Wohl (2011). Specifically, I consider the case that the heads who worked in large-scale or governmental factories to have potentially different risk preferences because those workers might have lower risk tolerance than those who worked in smaller enterprises. I confirmed that the sensitivities to the aggregate shocks are materially similar in both groups in the statistical sense. [^54] Online Appendix B.10 summarizes the finer details of this analysis. In short, the heads’ earnings are regressed on the aggregate consumption, its interaction term with respect to the indicator variable for the large-scale and government factories, family size variable, and household fixed effects. The results are unchanged if I trim a few semi-months with a smaller number of cross-sectional observations. This result is unchanged if I assume that the heads who worked in the smithing factory or in the non-machinery sectors had different risk preferences. [^55] The skilled workers in the smithing sector might be a business owner (Online Appendix A.6). Similarly, the workers in the non-machinery sectors in Tsukishima, a representative machinery production area, might have had a different, presumably higher, risk-tolerance. This supports the evidence that conditional on the household fixed effects, it is plausible to assume that the THBS households shared similar risk preferences.
4 Empirical Analysis
4.1 Consumption Smoothing
Estimation Strategy
I consider a linear fixed-effects model designed to test the consumption smoothing. [^56] Another empirical specification used to test the consumption smoothing is the first-difference model, including the aggregate consumption to capture the macroeconomic shocks in the economy (Mace 1991). In other words, this alternative model requires the interpretation of the estimated coefficient on the aggregate measure of consumption. Although the consumption of the THBS households tends to capture the macroeconomic trend in the consumption (Section 3.3), those households do not cover entire households in the target economy, Tsukishima. This disturbs the interpretation of the coefficient on aggregate consumption. To be conservative, therefore, I use the two-way fixed-effects model to control for the macroeconomic shocks (Cochrane 1991; Ravallion and Chaudhuri 1997). Online Appendix C.1 provides the conceptual framework and the derivation of my empirical specification. For household $i$ in time $t$, the reduced form equation is characterized as follows:
$\log c_{i,t}=\gamma\log y_{i,t}+\delta x_{i,t}+\mu_{i}+\phi_{t}+u_{i,t},$ (1)
where $c_{i,t}$ is consumption, $y_{i,t}$ is disposable income, $x_{i,t}$ is the family size control, $\mu_{i}$ is the household fixed effect, $\phi_{t}$ is the semi-month fixed effect, and $u_{i,t}$ is a random error term. The household fixed effect captures the time-constant unobservable factors such as permanent income and consumption preference. The time-fixed effect effectively removes the unobservable macroeconomic shocks and trends described in Section 3.4. The estimate of $\gamma$ shows income elasticity, ranging from zero for perfect insurance to one for the absence of insurance. [^57] This is comparable to the estimate from the same conceptual framework: this comparability is important in the consumption smoothing literature. Note that relevant previous studies have not integrated the recent theoretical discussion on the fixed effect model with two-way error components under the difference-in-differences setting (TWFE-DID) yet. In fact, the bias-corrected estimator of the TWFE-DID with continuous treatment for multiple periods is underdeveloped yet (see de Chaisemartin and D’Haultfœuille (2023) for a comprehensive review). The cluster-robust variance-covariance matrix estimator is used to deal with heteroskedasticity and serial dependency (Arellano 1987).
Results
Table 4: Results of Estimating Income Elasticities
(1) Semi-monthly Panels (2) Adj. Monthly Panels (3) Unadj. Monthly Panels
Disposable income Obs. Disposable income Obs. Disposable income Obs.
Coef. Std. error Coef. Std. error Coef. Std. error
Total consumption 0.358 [0.033]*** 278 0.433 [0.059]*** 123 0.261 [0.054]*** 141
      Food 0.136 [0.030]*** 259 0.293 [0.066]*** 114 0.100 [0.072] 132
      Housing 0.099 [0.415] 134 0.013 [0.084] 105 -0.263 [0.167] 124
      Utilities -0.003 [0.138] 223 0.343 [0.236] 110 0.197 [0.334] 130
      Furniture 0.426 [0.223]* 169 -0.114 [0.706] 98 0.250 [0.339] 112
      Clothes 0.378 [0.190]* 238 0.347 [0.424] 112 0.806 [0.262]*** 132
      Education 0.101 [0.051]* 236 0.162 [0.128] 110 0.069 [0.146] 128
      Medical expenses 0.039 [0.066] 258 0.194 [0.120] 114 -0.128 [0.094] 132
      Entertainment expenses 0.307 [0.084]*** 250 0.716 [0.181]*** 113 0.194 [0.153] 132
      Transportation 0.278 [0.110]** 187 0.605 [0.264]** 97 0.362 [0.291] 115
      Gift 0.523 [0.106]*** 234 0.469 [0.197]** 113 0.622 [0.306]* 130
      Miscellaneous -0.076 [0.147] 214 0.522 [0.252]** 110 0.246 [0.267] 126
***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors in brackets are clustered at the household level.
Notes: This table shows the results of equation 1: the regressions of the 11 measures of log-transformed consumption on log-transformed disposable income as well as on the family size control, household-fixed effect, and time-fixed effect. Columns 1, 2, and 3 show the results for the semi-monthly, adjusted monthly, and unadjusted monthly panel datasets, respectively. The estimated coefficients on log-transformed disposable income are listed in the columns named “Coef.”.
Column (1) of Table 4 presents the results from the semi-monthly panel dataset. In the total consumption, the estimated coefficient is $0.358$. This means that a one percentage point decrease in disposable income decreases the total expenditure by $0.36$ percentage point, suggesting that the factory worker households could not thoroughly smooth their consumption. An important finding is that the estimate for food categories is smaller ($0.136$) but highly statistically significant. This implies that while the households tended to reduce their food expenses when they faced adverse income shocks, they might have had some means to deal with those shocks to smooth food consumption. The estimates for the entertainment and transportation expenses ($0.307$ and $0.278$) are also highly statistically significant. This may make sense given that those expenses include family travel and theatergoing. [^58] The workers can enter their factories by walking in Tsukishima. Thus, the costs of commuting to the factory are not usually documented in the budget books. Similarly, the estimate for gifts shows a large magnitude ($0.523$), reflecting that the expenses on the others are essentially optional. The results for the other subcategories are statistically insignificant or unclear. The estimates for furniture and clothes are weakly statistically significant. Although less than conclusive, this may reflect that both categories are luxury categories. The estimate for the education category is small but also weakly statistically significant ($0.101$). This may be because it includes the allowance for children, which could be easily reduced. [^59] Toyotaro Miyoshi served as a research assistant at the Tsukishima Survey while he was a student at the Tokyo Imperial University. He left important retrospectives on the survey in his works (Online Appendix B.1). According to his book, since there were no nurseries, children in these working-class households regularly received a small allowance to play outside while their parents were working (Miyoshi 1989, pp. 383–384; 386; 387).
I present the results for the adjusted monthly panel dataset in Column (2) to see whether the income elasticities changed in a longer span. Not surprisingly, the adjusted monthly series yields greater estimates than the semi-monthly series. The estimate for total consumption is now $0.433$, which is roughly $1.2$ times greater than that from the semi-monthly series. One potential explanation behind this difference is that the available sources of temporary income have been limited, and those sources are not useful for compensating for adverse income shocks lasting for at least one month. The most noticeable example is the food category, which comprises the largest portion of the expenditure (Panel A of Table 3). The estimate is $0.293$, more than two times greater than the estimate for the semi-monthly series ($0.136$). This implies that the households’ food expenditures are likely to be insured in a few weeks but not over one month. As I will delve into the mechanism behind this behavior in the following subsections, credit purchases may work to smooth food consumption in the short run. Similarly, the expenses for entertainment and transportation are much more sensitive ($0.716$ and $0.605$) to the shocks in the adjusted monthly series. The estimate for gifts is slightly smaller but still in a similar range. In short, while the household could compensate for their income losses for a few weeks, they could not do so for several weeks, presumably because of borrowing constraints and/or a limited amount of liquidity assets. The expenses on those unnecessary categories might particularly be easily cut in response to relatively persistent shocks.
In contrast, a few subcategories are not sensitive to the idiosyncratic shocks over different time-series frequencies. For example, the estimates for the housing, utilities, and medical categories are not statistically significant in both semi-monthly and adjusted monthly series. This may reflect that the rent and utility payments were relatively fixed. [^60] For example, rent payment is usually fixed and paid once per month. Online Appendix C.2 shows that the housing category should be analyzed based on the adjusted monthly series rather than the semi-monthly series. The medical expenses included the fee for bathing (yuya). People had taken both every few days at that time, which was less likely to be cut, albeit hairdressing for women (kamiyui) in the same category may be likely to be cut. The estimates for the furniture, clothes, and education categories are statistically insignificant in the adjusted monthly series. While this could partly be the efficiency issue under a smaller sample size, those expenses may be sensitive to unexpected shocks in consumption, like unanticipated losses of household goods, clothes, and stationaries including writing papers. [^61] An alternative explanation for the clothes category relates to seasonality. The workers had bought their work clothes for the new season (for example, the OISR, Archives of the Tsukishima Survey (THBS #40, 31st March 1919)). Since those clothes are rather necessary items, this sort of seasonal expense may attenuate the estimate under low frequencey series. This systematically attenuates the estimates for these specific categories in the adjusted monthly series. [^62] Generally, the panels with short-run time bins are more likely to be influenced by unexpected shocks in consumption, which tends to attenuate the income elasticity (Nelson 1994). If this mechanism worked, subcategories that are sensitive (insensitive) to unexpected shocks in consumption should be more (less) attenuated. In my result, however, the estimates for the adjusted monthly panels are greater than those for the semi-monthly panels across a large part of subcategories. Thus, on average, my result indicates that households might face more difficulties in smoothing their consumption in the longer run.
Finally, I assess the validity of the aggregation method used in this study. Adjusting the aggregate data to account for the timing of paydays reveals more transparent relationships between income and expenditure than the unadjusted data (Section 3.3). To demonstrate this mechanism, I listed the estimates from the unadjusted monthly panel dataset in Column (3). The estimate for total consumption is $0.26$, which is much lower than the estimate for the adjusted series ($0.43$). In addition, most subcategories show systematically smaller estimates than those from the adjusted data listed in Column (2). For example, the estimate for the food category is now $0.10$ and statistically insignificant. Despite this, several subcategories take values close to one or large negative values, which are no longer economically interpretable. [^63] For example, the estimate for housing is now negative. The households paid a fixed rent amount monthly. Given that the rent payday is fixed, for example, the shifts in wage paydays might create cells with less income with a fixed amount of rent. In the case of seasonal consumption categories, the shifts may coincide with the timing of seasonal spending. The THBS households tended to purchase work clothes in March to prepare for the start of the new fiscal year in April. They also spent more on clothing in June, likely to buy summer clothes. The shift of paydays to these particular months may have created a strong but false positive correlation. This means that measurement errors in the unadjusted aggregation lead to strong attenuation bias, which results in false negatives in most of the consumption categories and may cause critical issues in statistical inference for a few subcategories.
4.2 Risk-coping Mechanisms
Preceding results show that the skilled workers’ households in Tsukishima were less likely to smooth their consumption. The estimated elasticities, however, become smaller when I focus on the semi-monthly frequency. Specifically, food expenditure is likely to be insured within at least half a month. The workers might have dealt with their shorter-run shocks relatively well but could not cope with the longer-run shocks. In this section, I test the mechanisms behind their smoothing behavior using several variables on the risk-coping devices.
Estimation Strategy
To test the roles of risk-coping strategies, I employ the empirical specification suggested by Fafchamps and Lund (2003). For household $i$ at time $t$, the specification is characterized as follows:
$r_{i,t}=\kappa\tilde{y}_{i,t}+\eta x_{i,t}+\nu_{i}+\zeta_{t}+\epsilon_{i,t},$ (2)
where $\tilde{y}_{i,t}$ is head’s earning, $x_{i,t}$ is the family size control, $\nu_{i}$ is the household fixed effect, $\zeta_{t}$ is the time fixed effect, and $\epsilon_{i,t}$ is a random error term. Online Appendix C.1 summarizes the derivation of this specification. I prefer to use the head’s income instead of the disposable income in equation 2 to avoid the potential endogeneity. [^64] For instance, the household’s total income is the outcome of the intra-household responses to the shocks, such as the labor-supply adjustments. I confirmed that the results are largely unchanged if I use the disposable income as the idiosyncratic income shock variable instead of the head’s income (not reported). This is consistent with the fact that the labor-supply adjustments did not occur among the THBS households (Table 5). Despite this, I prefer to use the conservative setting to deal with potential biases and to set a comparable setting to the structural equation used in Section 4.3. See Asdrubali et al. (2020) for a recent study using the head’s income as the shock variable in testing the full risk-sharing. As for $r_{i,t}$, I consider several variables to overview various risk-coping strategies. These include the net income on savings, insurance, borrowing, credit purchases, and gifts. Each net variable is defined as the difference between income and expenses. The estimates of $\kappa$ are therefore expected to be negative since net income becomes negative in good times but positive in unfavorable times. I also consider the earnings of all the family members except for the head and the sales of miscellaneous goods because both intra-household labor supply adjustments and asset sales may play a role in mitigating the idiosyncratic shocks as alternative risk-coping strategies (Udry 1995; Blundell et al. 2016). [^65] Rosenzweig and Stark (1989) show that marriage cum migration in South Indian villages contributed to reducing the variability of household food consumption. Since the THBS focuses on the households living in Tsukishima throughout the sample period, I can discard addressing the ex-ante risk-coping strategy through migration. The estimated coefficients for the sales and labor supply adjustment variables may be negative, reflecting additional earnings when the head’s income is reduced.
Results
Table 5: Results of Testing the Risk-Coping Mechanisms
(1) Semi-monthly Panels (2) Adj. Monthly Panels
Head’s earning Obs. Head’s earning Obs.
Coef. Std. error Coef. Std. error
Panel A: Savings, insurance, borrowing, and gifts (yen)
Net savings -0.043 [0.018]** 289 -0.035 [0.028] 124
Net insurance 0.001 [0.013] 289 0.002 [0.033] 124
Net borrowing -0.021 [0.027] 289 -0.027 [0.017] 124
Net credit purchase -0.128 [0.039]*** 289 -0.011 [0.042] 124
Net gifts -0.052 [0.019]** 289 -0.006 [0.029] 124
Panel B: Labor supply adjustments & sales of miscellaneous assets (yen)
Other member’s earnings 0.001 [0.010] 289 0.036 [0.033] 124
Sales of assets -0.000 [0.000] 289 0.000 [0.001] 124
***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors in brackets are clustered at the household level.
Notes:
This table shows the results of equation 2: the family size control, household fixed effects, and time-fixed effects are included in all the regressions. Columns 1 and 2 show the estimates for the semi-monthly and adjusted monthly panel datasets, respectively. The estimated coefficients on the head’s earnings are listed in the “Coef.” columns.
Panel A: Each net income variable is defined as the difference between income and expenses. “Net savings” refers to the difference between the amount withdrawn and the total deposits. “Net insurance” refers to the amount received from insurance minus the expenses paid for insurance. “Net borrowing” refers to the amount of borrowing minus the debt payments. “Net credit purchase” refers to the amount of credit purchases minus credit redeemed. “Net gifts” refers to the total received amounts, including both pecuniary and non-pecuniary gifts, minus the payments for gifts.
Panel B: “Other member’s earnings” refers to the total income earned by all family members except for the head of the household. “Sales of assets” includes the sale of daily miscellaneous goods such as newspapers and empty bottles. All dependent variables are in yen.
Table 5 presents the results. Columns (1) and (2) summarize the results for the semi-monthly and adjusted monthly panel datasets, respectively. Panel A shows the results for the net income variables, whereas Panel B shows those for the labor supply adjustments and sales of assets.
The estimated coefficient for the net savings is negative and statistically significant in Column (1). The estimate suggests that a one yen decrease (increase) in the head’s earnings increases (decreases) the net income from savings by $0.043$ yen. The estimate from the monthly series is similar ($0.042$), albeit weakly statistically significant (Column (2)). This is basically consistent with the historical fact that many workers had savings accounts both in postal savings and savings banks at that time (Section 2.3). However, the estimated magnitude is small, suggesting that the withdrawals from savings did not compensate for the earnings losses.
The estimates for the insurance category are close to zero and statistically insignificant in both columns. The budget books indicate that the primary source in the category is the mutual loan associations (mujin). As explained in Section 2.2, these associations’ rules were unsuitable for compensating the idiosyncratic shocks. A suggestive case is that a THBS household won a lottery and got $100$ yen in a monthly cell, which is clearly independent of the worker’s earnings. [^66] The OISR, Archives of the Tsukishima Survey (THBS #5, 17th May 1919). The bidding rule may be used to support a specific member in the association if members have a common belief in the aid. However, most of the incomes from mujin measured in the budget books are less than one yen, which is not the case on average.
For the net borrowing category, the estimated coefficients are close to zero and statistically insignificant in both columns. Section 2.2 shows that pawnshop was the popular lending institution for low-income households. Although THBS households occasionally used pawnshops, they never used them frequently. [^67] The OISR, Archives of the Tsukishima Survey (THBS #2; 19; 22). This makes sense because skilled factory workers can not classified as the low-income working class.
The estimated coefficient for the net income from credit purchases is negative and statistically significant in Column (1). The estimate suggests that a one yen decrease (increase) in the head’s earnings increases (decreases) the net income from credit purchases by $0.13$ yen. This magnitude is larger among the other categories. According to the THBS budget books, the credit purchases were mostly used to buy rice, fish, vegetables, and fuels (firewood and charcoal). Since rice, a staple food, is usually an expensive item, credit purchases might have played an important role among these households. The credit purchases were occasionally used in the call doctor and pharmacy. This is consistent with my baseline result for medical expenses in the semi-monthly series (Column (1) in Table 4). While this provides suggestive evidence that credit purchase is the primary risk-coping strategy, Column (2) shows that the credit was less likely to work in the longer run. This seems to be reasonable because borrowers could make monthly or semi-monthly payments but did not borrow for more than a month (Section 5). Instead, the result for the net savings category suggests that precautionary savings have helped mitigate income declines lasting at least one month due to idiosyncratic shocks. Given the magnitudes, however, it is important to note that it could only partially compensate for the idiosyncratic shocks.
The estimates for the net income from gifts are negative and statistically significant in Column (1). The budget books tell that the gifts included seasonal gifts, celebration money, and return gifts. Among them, the return gifts can be regarded as mutual aid from personal networks, which may be reflected in this result. According to an official report of Tokyo, however, the share of average monthly gifts in total income was approximately $3$% (Social Affairs Division 1925, pp. 58-59). In my adjusted monthly panels, this share is calculated to be roughly $4$%. In this light, while the gifts comprised a few percentage points of the monthly income, they may not have provided sufficient aid for the households facing economic hardship. In fact, the estimate is no longer statistically significant in the adjusted monthly series (Column (2)).
Finally, Panel B tests the roles of the intra-household labor supply adjustment and sales of miscellaneous assets. The estimated coefficients in both columns are close to zero and statistically insignificant. Miyoshi (1989, pp. 383–386) documented that the wives regularly engaged in side jobs rather than worked temporarily to compensate for the income loss. In fact, although sewing was the most popular side job among THBS wives, their earnings were too small to replace the lost income from their breadwinning husbands. [^68] The OISR, Archives of the Tsukishima Survey (THBS#1, 13th January 1919). In the final row in Panel B, I find that selling miscellaneous assets such as newspapers, bottles, and clocks did not work as a risk-coping strategy.
To summarize, I find that credit purchases have played a role in mitigating the idiosyncratic shocks. The results also indicate that the savings institutions and personal networks might have been used as risk-coping institutions among the workers. However, these were basically available for short-term, say at most half a month. The estimated magnitudes indicate that the workers could only partially compensate for the temporary losses of income by using these institutions, which is indeed consistent with the findings on the partial risk-sharing in Section 4.1. Overall, the results are aligned with the descriptives from the historical documents on the financial institutions available in Tokyo City (Section 2.3). An interesting finding, which is not explicitly described in the historical sources, is the possibility that the credit purchases in the local society and the informal gifts from personal networks might have provided some aid to the factory workers.
4.3 Event-Study Analysis: Response to the Advers Health Shocks
The foregoing analyses used the within variations in the earnings for estimating the income elasticity, meaning that the identification depends both on the adverse and favorable idiosyncratic shocks. While it is useful to illustrate the overall responses of the households to the shocks, it does not offer a clear view of the pathways behind the risk-coping behaviors against the adverse idiosyncratic shocks. To clarify this point, I next focus on the temporary illness of the heads that cause temporary income losses.
Estimation Strategy
Figure 5: Number of Influenza Deaths in Tokyo between 1918 and 1920
Notes: This figure shows the time-series plots of the monthly death counts from influenza (ryūkōsei kanbō) in Tokyo prefecture between January 1918 and December 1920. The death statistics used in this figure were collected based on the comprehensive registration system.
Source: Created by the author using the 1918, 1919, and 1920 editions of the Statistics of Causes of Deaths of the Empire Japan (Statistics Bureau of the Cabinet 1921; 1922; 1923).
I begin my analysis by testing whether the illness of the head impacted the earnings. To do so, I decided the exact timing and duration of illnesses of the family members using the Kazoku idō hyō (Section 3.4). Figure 5 shows the number of monthly death counts from influenza (ryūkōsei kanbō) in Tokyo between 1918 and 1920. There appears to have been a moderate spike in early 1919, which is partially captured in the THBS period. In fact, a few THBS budget books suggest that the family members had caught the flu during this period. Five households experienced the illness during the sample period, and the illness affecting the heads was concentrated between February and May 1919, when flu exposure was most prevalent. [^69] Among four cells with health shocks on the heads, two are observed in February, during the peak of the epidemics (Figure 5), and the others are in March and May. In addition, the average duration for the heads is $5.8$ days (Std. Dev. $=3.6$), which is consistent with the fact that Type A influenza causes fever for less than $6$ days (Uyeki et al. 2022). Therefore, it is plausible to see that the influenza epidemics largely led to illnesses during this period, providing a quasi-experimental setting in analyzing to what extent the adverse idiosyncratic shock driven by illness influences the households’ consumption behaviors.
Earnings in a semi-month are based on the work done during the previous semi-month. This means that the head’s earnings shall mainly be influenced by the number of days suffering illness in the previous semi-month cell. Table 6 presents the results from the regression of the head’s earnings on the number of illness days in the previous semi-monthly cell. While the instantaneous health shock has little impact on the head’s earnings (Column 1), the lagged health shock has a statistically significant effect on his earnings (Column 2). This result is unchanged if I include both variables in the same regression in Column 3, supporting the evidence that the health shocks in the previous cell reduced the earnings of the head. The estimate in Column (2) suggests that a one-day increase in the illness decreases the head’s semi-monthly income by roughly $2.4$ yen. Notably, this is close to the average daily wage rate of skilled factory workers (Online Appendix B.6). Thus, this reflects the reductions in their salary due to the absence: Miyoshi (1989, pp. 383; 388) highlighted that income was accidentally reduced due to absence when the head was infected with a disease. Finally, Column (4) adds the health shocks of the other family members into the specification of Column (3). The estimated coefficients on these placebo variables are statistically insignificant, supporting the robustness of the result in Column (3). These results confirm that the head who suffered illness lost earnings. [^70] Given that the average number of days suffering the illness is roughly $2.0$, the temporary health shock loses more than $5$ yen in that month, accounting for roughly $10$% of the average monthly income. Although this is a rough estimate and the losses shall depend on the severity of the shocks, my result at least shows that the household income had reacted to the adverse health shock among the factory workers’ households.
Table 6: Results of Testing the Impacts of Adverse Health Shocks
on the Head’s Earning
Dependent Variable: Semi-monthly Earnings of the Head
(1) (2) (3) (4)
Health shocks on:
The head [level] $-0.061$ $-0.972$ 1.319-1.319*- 1.319 ∗
$[0.679]$ $[0.699]$ $[0.666]$
The head [lag] 2.371-2.371***- 2.371 ∗ ∗ ∗ 2.868-2.868***- 2.868 ∗ ∗ ∗ 2.994-2.994***- 2.994 ∗ ∗ ∗
$[0.644]$ $[0.734]$ $[0.765]$
The others [level] $0.889$
$[1.275]$
The others [lag] $3.675$
$[3.312]$
Observations 287 254 254 254
***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors in brackets are clustered at the household level.
Notes: This table shows the results from the regressions of the head’s earnings on the health shock variables using the semi-monthly panel dataset. A household with two semi-month observations is excluded from the sample in all the regressions because the lagged health shock variables are included. The family size variable, household fixed effects, and time-fixed effects are included in all the regressions. “Health shocks on the head” refers to the number of days that the head suffered illness. “Health shocks on the others” refers to the number of days that family members, excluding the head, were ill. “[level]” indicates the variable of the head’s earnings in the same semi-month cell. “[lag]” indicates the lagged variable (i.e., the shocks measured in the previous semi-month cell). Although correlation coefficient between the level and lag of the health shock variables for the other family members is substantially high (Pearson’s coefficient $=0.99$), including eather of the two variables does not change the finding on the neglisible impacts of the health shocks on the other family members.
Following the evidence, I characterize equation 1 as a structural consumption smoothing equation by using the first-stage reduced form equation as follows:
$\log y_{i,t}=\beta h_{i,t-1}+\psi x_{i,t}+\varpi_{i}+\tau_{t}+e_{i,t},$ (3)
where $h_{i,t-1}$ is the number of days of head’s illness in the previous semi-month cell, $\varpi_{i}$ is the household fixed effect, $\tau_{t}$ is the semi-month fixed effect, and $e_{i,t}$ is a random error term. For the risk-sharing mechanisms analysis under equation 2, the log-transformed disposable income in equation 3 ($y_{i,t}$) is replaced with the head’s earnings ($\tilde{y}_{i,t}$). The health shock variable is plausibly exogenous because the illnesses were from the influenza epidemics, as explained above. [^71] As noted in Section 2.3, the factories did not offer compensation for their sickness leave because the Factory Act did not specify benefits for workers’ illness (Study Group of the Factory Act 1916). Even in the largest factory in Tsukishima, the Ishikawajim Shipyard, any descriptions of the influenza epidemics are documented in the company’s history (Arai 1930). In fact, Miyoshi (1989) documented the issue that the head’s disease directly causes income loss due to the absence. In addition, there is a correlation between the health shock variable and the head’s earnings at a statistically valid level (Table 6). Both support the evidence that the exogeneity of the instrument and relevance condition basically hold under this just-identified system.
Results: To what extent do illnesses affect consumption?
Table 7: Results of Estimating Income Elasticities:
Smoothing Consumption against the Adverse Health Shocks
Semi-monthly Panel Dataset
(1) Reduced-form approach (2) Instrumental variable approach
Head’s illness Obs. Disposable income Obs. First-stage
Coef. Std. error Coef. Std. error $F$-statistics
Total consumption -0.118 [0.015]*** 254 0.543 [0.076]*** 245 17.6
      Food -0.075 [0.017]*** 236 0.355 [0.075]*** 227 18.4
      Housing -0.017 [0.070] 126 0.118 [0.477] 126 7.6
      Utilities -0.051 [0.066] 199 0.201 [0.331] 192 12.4
      Furniture -0.157 [0.106] 156 0.761 [0.801] 150 6.2
      Clothes -0.339 [0.050]*** 219 1.233 [0.222]*** 211 65.4
      Education -0.097 [0.024]*** 216 0.462 [0.101]*** 209 18.9
      Medical expenses -0.002 [0.028] 236 -0.001 [0.142] 227 18.4
      Entertainment expenses -0.094 [0.055]* 227 0.427 [0.501]** 218 20.2
      Transportation -0.100 [0.094] 170 0.436 [0.510] 162 20.5
      Gift -0.325 [0.093]*** 212 1.567 [0.797]* 204 17.8
      Miscellaneous -0.075 [0.027]*** 195 0.343 [0.101]*** 189 17.3
***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors in brackets are clustered at the household level.
Notes:
Column 1 shows the results of equation 1, in which the log-transformed disposable income is replaced with the head’s illness (i.e., the number of days with illness in the previous semi-month cell) as the idiosyncratic shock variable. The dependent variables are the 11 measures of log-transformed consumption, and all the regressions include the family size control, household fixed effects, and time fixed effects. The estimated coefficients on the health shock variable are listed in the sub-column named “Coef.”.
Column 2 shows the results of equation 1, in which the log-transformed disposable income is regarded as the endogenous variable. The head’s illness is an instrumental variable (i.e., excluded exogenous regressor) for the first-stage reduced-form equation (3). The estimated coefficients on the disposable income are listed in the sub-column named “Coef.”. The first-stage $F$-statistics on the health shock variable are listed in the final column.
I first consider a reduced-form specification using the head’s illness as the exposure variable in equation 1 in the spirit of Cochrane (1991). Column (1) in Table 7 shows the results. Overall, the coefficients on the exposure variable are negative, supporting evidence that adverse health shocks on the heads reduce consumption levels. Next, I explicitly model the structure behind the relationship between the health shocks and earnings using the reduced-form equation of 3. If the negative impacts of the health shocks on consumption were derived from the income losses, the result from the structural estimation approach is similar to that in Column (1). The estimates from the structural equation are listed in Column (2) of Table 7. As shown, the results are materially similar to the reduced-form results in Column (1). Importantly, the estimates are systematically larger than those reported in Table 4. The estimates for total and food consumption categories are $0.543$ and $0.355$, respectively, whereas the baseline estimates are $0.358$ and $0.136$. While this result partly reflects the bias reduction under the instrumental variable estimation, it is also important to remember that this system is designed to reduce the bias by specifying the channel from the adverse (instead of favorable) income shocks. For example, expenses on the allowance for children could be easily cut during unfavorable times but may not respond to favorable shocks. Similarly, the expenses for furniture are likely to be cut, but the households might not buy those frequently, meaning that there are no sufficient cases for reacting to the rare health shock events. [^72] In other words, the furniture may have been purchased only when the households had extra income. This may partly reflected in the smaller value of the first-stage $F$-statistics in the furniture subcategory. [^73] Note that the logarithmic transformation leaves the cells with expenses on the furniture. Thus, the reduced-form equation captures the relationship between the health shocks and earnings in the cells with any expenses on the furniture. While the first-stage $F$-statitics is $35.7$ in the linear-linear model, the second-stage result is robust against the functional form assumption. This may reflect the asymptotic theoretical advantage in the panel data model under nearly weak instruments case (Staiger and Stock 1997; Cai et al. 2012; Baltagi et al. 2012). The same argument is applicable to the housing category. A similar interpretation can be applicable to the transportation category because family travel is not a frequent event and thus, unlikely to coincide with the timing of illness. In contrast, the entertainment expenses are still likely to be cut in case of illness, which may reflect the curtailment of tobacco, theater-going, and the allowance for heads. This may be another explanation for the difference in the results from my baseline results in Table 4.
While the estimates for clothes and gifts categories now exceed one, the null of equality to the population mean of one is not statistically rejected for both cases. Although the large estimates may be partly due to the operation under the instrument based on the rare health shock events, the findings are materially similar to my baseline results: both categories are sensitive to the shocks as those are dispensable categories.
Results: How did the households compensate for the losses from the ill?
Table 8: Results of Testing the Risk-Coping Mechanisms:
Adverse Health Shock as the Idiosyncratic Shock
Semi-monthly Panels
(1) Reduced-form approach (2) Instrumental variable approach
Head’s illness Obs. Head’s earning Obs.
Coef. Std. error Coef. Std. error
Panel A: Savings, insurance, borrowing, and gifts (yen)
Net savings 0.397 [0.318] 254 -0.167 [0.117] 254
Net insurance -0.168 [0.204] 254 0.071 [0.100] 254
Net borrowing 0.116 [0.261] 254 -0.049 [0.103] 254
Net credit purchase 1.294 [0.674]* 254 -0.546 [0.203]*** 254
Net gifts 0.181 [0.177] 254 -0.077 [0.090] 254
Panel B: Labor supply adjustments & sales of miscellaneous assets (yen)
Other member’s earnings 0.001 [0.081] 254 -0.001 [0.034] 254
Sales of assets 0.002 [0.002] 254 -0.001 [0.001] 254
***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors in brackets are clustered at the household level.
Notes:
Column 1 shows the results of equation 2, in which the head’s earning is replaced with the head’s illness (i.e., the number of days with illness in the previous semi-month cell) as the idiosyncratic shock variable. The family size control, household fixed effects, and time-fixed effects are included in all the regressions. The estimated coefficients on the head’s illness are listed in the subcolumn named “Coef.”.
Column 2 shows the results of equation 2, in which the head’s earning is treated as the endogenous variable. The head’s illness is an instrumental variable (i.e., excluded exogenous regressor) in the first-stage reduced-form equation. The first-stage result is identical to the result listed in Column (2) of Table 6 with the first-stage $F$-statistics of $13.6$. The estimated coefficients on the head’s earning are listed in the subcolumn named “Coef.”.
Panel A: Each net income variable is defined as the difference between income and expenses. “Net savings” refers to the difference between the amount withdrawn and the total deposits. “Net insurance” refers to the amount received from insurance minus the expenses paid for insurance. “Net borrowing” refers to the amount of borrowing minus the debt payments. “Net credit purchase” refers to the amount of credit purchases minus credit redeemed. “Net gifts” refers to the total received amounts, including both pecuniary and non-pecuniary gifts, minus the payments for gifts.
Panel B: “Other member’s earnings” refers to the total income earned by all family members except for the head of the household. “Sales of assets” includes the sale of daily miscellaneous goods such as newspapers and empty bottles. All dependent variables are in yen.
I test the risk-coping strategies for mitigating adverse health shocks in Table 8. Column 1 lists the reduced-form results for the risk-coping strategies in the same manner in Table 5. The estimated coefficients on the net income variables are positive in most cases, indicating the households’ compensating responses against the shocks. The estimate for the credit purchases category is weakly statistically significant, with the largest magnitude. It suggests that the one-day increase in the head’s illness increased the net credit purchases by $1.29$ yen, which accounts for roughly half of their average daily wages ($1.29/2.56$). Column 2 presents the estimates from the structural equation using the head’s earnings as the endogenous variable. A similar but much clearer result is obtained: the estimate for the credit purchase category is roughly four times greater magnitude than that reported in Column (1) of Table 5. The estimate indicates that a one yen decrease in the head’s earnings increases the net income from credit purchases by $0.55$ yen. This magnitude is consistent with the direct effect of the health shocks on the net credit purchase estimated in Column (1) of Table 8.747474This means that $2.56\times 0.55=1.4$ yen, which is close to $1.3$ yen in Column (1). The estimates for the savings and gifts categories are still negative but now statistically insignificant, supporting the evidence that credit purchase is the primary risk-coping strategy for dealing with health shocks. Since it shall take more time (steps) from withdrawing and receiving gifts to use the money at the retailers, it is likely that the households prefer to use credit purchases.
Overall, the event-study analysis using adverse health shocks confirms baseline findings and provides a sharper illustration of consumption smoothing behaviors in the case of adverse idiosyncratic shock events. In short, my results show that the unique available response among the THBS households to the idiosyncratic income shock was to use credit purchases. The available compensation via the credit purchases accounts for roughly half of their temporary income losses, which partially contributes to smoothing their food consumption in the short-run.
5 Credit System in Local Economy
Foregoing results shed new light on the roles of credit institutions among skilled workers’ households during historical industrialization. Since the credit purchases are provided by local retailers, this also suggests the importance of the relationships between the workers and retail shops. To the best of my knowledge, however, there is little historical work on credit purchases in prewar Japan. Hence, the details of this credit institution are not yet clear: for example, how retailers get the credit information of their customers, whether the contracts are explicit or implicit, and whether the lenders enforce penalties for the debtors.
Table 9: Customers, Sales Methods, and Competitiveness in the Retailers in Tokyo City
Selling Item Number of Share of sales Social Class of Sales Method (%) Number of
stores for regular the Regular Customer (%) Cash Credit Monthly competitors
reported customer (%) Upper Middle Other purchase installment reported
White Rice 101 73.1 16 45.7 38.2 26.9 71.5 1.7 8.5
Fish 59 64.6 6.8 57.6 35.6 43.9 56.1 0 5.0
Greengrocer 66 45.4 9.4 40.4 50.2 65.8 34.2 0 6.2
Firewood and charcoal 100 51.1 15.7 50.9 33.4 20.4 78.4 1.3 8.6
Notes: This table summarizes the statistics on the sales in the $326$ retail shops selling white rice, fish, fruit and vegetables, and firewood and charcoal in Tokyo surveyed in 1935. The number of stores reported the share of regular customers (column 2) are $90$, $56$, $61$, and $93$ in the white rice, fish, greengrocer, and fuel (firewood and charcoal) categories, respectively. The “share of sales for regular customers” is defined as the percentage of sales for regular customers in the annual total sales in 1935. The definition of the social classes (“Upper”; “Middle”; “Other”) is not exactly defined in the report. The share of sales method is based on the percentage of sales using each method in the annual total sales in 1935. The number of competitors indicates the number of peers within a three-cho, approximately $327$ meters, radius. The department stores and retail markets are not included in this survey.
Source: Tokyo City Chamber of Commerce and Industry 1937, pp. legend; 20–21; questionnaire.
The information on the credit purchase is considerably personal, as in the case of state-contingent loans in the rural credit market (Udry 1994), and can hardly be documented. Therefore, systematic information on credit purchases in the prewar period is unavailable. Fortunately, however, a few official survey reports documented some information on this payment method. A survey report on retailers in 1930 (hereafter 1930 Survey Report) revealed that credit purchase was a popular payment method in transactions with consumers. [^75] This survey was conducted in February and March 1930 by Tokyo City and covered $3,892$ small and medium enterprises (Tokyo City Office 1932a, p. 1). Those include not only retailers but also wholesale and other small businesses, and the number of respondents varies depending on the question items. The number of respondents in the statistical table on the payment methods is $2,654$, of which $1,197$ are retailers (Tokyo City Office 1932b, p. 273). To the best of my knowledge, this is the most comprehensive survey of the retailers in Tokyo around 1930. For example, 99%(155/157) of grain retailers, 81%(35/43) of fish retailers, 78%(64/82) of greengrocers, and 94%(45/48) of fuel retailers used credit purchases for their selling method (Tokyo City Office 1932b, pp. 260; 263). A survey report on retail shops in 1935 published by the Tokyo City Chamber of Commerce and Industry provides more detailed information. Table 9 summarizes the information on the sales in those retail shops of rice, fish, fruit and vegetables, and fuels. [^76] This survey covers 939 shops and was conducted on 31st December 1935. I digitized the statistics for all the shops selling the representative items (rice, fish, fruit and vegetables, and fuels) in the credit purchases. The survey instrument includes detailed questions regarding sales and purchases, making it the most comprehensive of the available documents (Tokyo City Chamber of Commerce and Industry 1937, pp. legend, questionnaire). Online Appendix Table C.1 decomposes the statistics listed in Table 9 by the scale of businesses. I confirm that the main findings are materially similar. First, many annual sales are from their regular customers, most of whom are from the non-upper classes. For example, $73$% of annual sales in rice retailers are from their regular customers, and more than $80$% of them are from non-upper classes, including the factory workers. Second, credit purchase is a popular sales method, especially among rice and fuel retailers: more than $70$% of the sales were done with credit in both shops. [^77] These figures are systematically lower than those from the 1930 Survey Report. Given that both surveys show a similar tendency regarding the use of credit purchases, this may be due to the differences in the samples, the definition of the retailers in each category, and the definition of the credit purchase itself. Although cash was also used frequently among greengrocers, $34$% of the sales were settled using credit. This supports the evidence that credit purchase was a popular payment method, and regular customers settled a large part of it. [^78] The 1930 Survey reveals that the retailers had also commonly used credit purchases to buy their source goods from wholesalers (Tokyo City Office 1932a, pp. 154–157). This means that credit purchase was a familiar payment method for retailers as well as consumers. Although these statistics are from the mid-1930s, this confirms the possibility that the retailers were engaged in credit transactions in the context of a long-term relationship with their regular customers. [^79] A textbook for retailers published in 1924 explains an interesting mechanism behind this long-term relationship from an aspect of customer psychology: credit purchase is made not only for the convenience of the customer’s budget but also because of the “vanity” of being able to be a regular customer (Shimizu 1924, pp. 115–116). The final column of this table further provides suggestive evidence behind this. The survey asked retailers whether the sales area was competitive or not, and if yes, they were asked to report the number of competitors within a three-cho (roughly $330$ meters) radius. As shown, rice and fuel stores have 2-3 more competitors than the retailers selling fish, fruit, and vegetables. Given the relatively high utilization of credit in rice and fuel retailers, they may have responded to competition by allowing their customers to trade on credit.
(a) Daily panel dataset:
purchase and payment
(b) Daily panel dataset:
net purchase
(c) Semi-monthly panel dataset:
purchase and payment
(d) Semi-monthly panel dataset:
net purchase
Figure 6: Relationships between Average Credit Purchases and Payments
by Different Time-series Frequencies
Note: Figure 6a, and 6c illustrate the time-series plots of the average daily and semi-monthly credit purchase and payment, respectively. The blue and red dashed lines indicate the average credit purchases and payments, respectively. Figure 6b and 6d illustrate the time-series plots of the average daily and semi-monthly credit purchase minus payment. The red solid line highlights the zero value. Figures 6a and 6b show the daily data between 12th January and 11th November 1919. Figures 6c and 6d show the semi-monthly series calculated using the daily data between 12th January and 11th November 1919. All figures are based on the 29 THBS households that used any credit purchases.
Source: Created by the author using the THBS sample.
Next, I look into consumers’ credit purchase behavior using the THBS dataset. Most of the THBS households ($29/33$) used any credit purchases during their sample period, supporting the evidence that this was a common practice among factory workers in Tsukishima. Figure 6a illustrates the average daily credit purchases and payments. Generally, there is a peak per semi-month in payments, which reflects the households’ behavior of making the payments for the credit purchase on or soon after their paydays. The purchases series shows a similar tendency but with lower spikes and larger variance. This indicates that the households purchased goods by a certain amount around their paydays and then used it when they needed unconservable food such as fish, fruit, and vegetables. Figure 6b illustrating the net purchases confirms this behavior: the households used the credit on a daily basis and redeemed it near paydays. Figure 6c shows that the payment slightly exceeds purchase, suggesting that many households could more or less repay their loans. Although the difference between payment and purchase is relatively obvious after August 1919, Figure 6d indicates that this difference is not statistically significant in all the semi-months. [^80] Strictly, payment can include accrued liabilities that should have been paid prior to the previous period. It is, therefore, reasonable that the average payment is greater than the average purchase in most semi-month cells. Thus, this is not a systematic trend but may be due to the smaller available number of households in the later sample period (Online Appendix Figure B.4).
I then analyze whether the volatility of earnings restricts credit purchases. Since the credit of a worker is unobservable, I use the volatility of earnings as a proxy of credit risk in the following empirical specification:
$I(\overline{CP}_{i}=0)=\varkappa_{0}+\varkappa_{1}CV_{i}+\bar{\mathbf{x}}_{i}^{\prime}\mathbf{\chi}+v_{i},$ (4)
where $I(\cdot)$ is an indicator function, $\overline{CP}$ is the (unit-)average credit purchases, $CV$ is the coefficient of variation for head’s earnings, $\bar{\mathbf{x}}$ is a vector of the (unit-)average family composition variables, and $v$ is a random error term. All variables are calculated using the semi-monthly series. [^81] For example, the coefficient of variation for household $i$ is defined as $CV_{i}=\{\sum_{t}^{T_{i}}(\tilde{y}_{i,t}-\bar{\tilde{y}}_{i,.})/(T_{i}-1)\}^{\frac{1}{2}}/\bar{\tilde{y}}_{i,.}$, where $T_{i}$ indicates the number of time-series observations, and $\bar{\tilde{y}}_{i,.}$ indicates the time-series average of the head’s income. In this cross-sectional model, I include a set of family composition variables to control for unobservable preferences because the household-fixed effect is unable to be modeled as an error component. If retailers allowed households to use credit purchases regardless of the variability of earnings, the estimate of $\varkappa_{1}$ becomes close to zero. In contrast, if retailers did not allow households with high-income volatility to trade on credit, the estimate should be positive and reasonably large.
Table 10: Volatility of Earnings and Use of Credit Purchases
Dependent Variable: Indicator Variable for
Non-CP Users Non-CP + Low-CP Users
(1) (2) (3) (4) (5) (6)
CV of Head’s Earnings $0.003$ $-0.002$ $-0.027$ $0.019$ $0.008$ $-0.005$
$[0.012]$ $[0.014]$ $[0.050]$ $[0.013]$ $[0.017]$ $[0.049]$
Sample mean of the DV $0.12$ $0.12$ $0.12$ $0.24$ $0.24$ $0.24$
Model LPM LPM Probit LPM LPM Probit
Number of THBS households 33 33 33 33 33 33
Family Size Yes Yes Yes Yes Yes Yes
Children aged 6–12 (%) Yes Yes Yes Yes Yes Yes
Children aged 13–16 (%) Yes Yes Yes Yes Yes Yes
Adults aged 17+ (%) Yes Yes Yes Yes Yes Yes
Head’s age No Yes Yes No Yes Yes
Head’s earnings No Yes Yes No Yes Yes
Zero slope ($p$-value) 0.467 0.718 0.336 0.515 0.519 0.451
***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. Robust standard errors are reported in brackets.
Notes: The dependent variable in Columns (1)-(3) is the indicator variable for the households who did not use any credit purchases during the sample period (non-CP users). The dependent variable in Columns (4)-(6) is the indicator variable for the non-CP users and households less than the first quartile in the sample average of credit purchases. Columns (1)-(2) and (4)-(5) use the linear probability model (LPM), whereas Columns (3) and (6) use the Probit model. “CV of Head’s Earnings” indicates the coefficient of variations in the head’s semi-monthly earnings. The sample mean of the CV is $5.76$ (Std. Dev. = $5.72$). All the regressions include the family composition variables: average family size, share of children aged 6–12, share of children aged 13–16, and share of adults aged 17+ (share of children aged 0–5 is used as a reference group). The average head’s age and average head’s earnings are added to the control variables in Columns 3 and 6. $F$-statistics $p$-values for the null of the zero-slope hypothesis are reported in Columns (1)-(2) and (4)-(5). $\chi^{2}$-statistics $p$-values for the null of the zero-slope hypothesis are listed in Columns (3) and (6). The standard errors for the LPM are based on the HC2 estimator proposed by Horn et al. (1975).
Column (1) of Table 10 presents the result for equation 4. The estimated coefficient is close to zero and statistically insignificant. A potential concern herein may be the correlations between the volatility of earnings and the head’s age and earnings level. Column (2) shows the result from the extended model, including the head’s age and earnings in the controls, which shows the robustness of the baseline estimate in Column (1). Another potential issue is the functional form assumption in the linear probability model. Column (3) then lists an alternative non-linear model, including the full set of control variables used in Column (2). It confirms that the estimate is not statistically significantly different from zero. Next, I changed the threshold for the dependent indicator variable to consider a wider range of households using a relatively smaller amount of credit purchases, say less than the first quartile of the average credit purchases. Columns (4)-(5) summarize the results in the same column layouts as Columns (1)-(3), confirming that the estimates are still close to zero in all the specifications. Finally, the final row in this table shows that the null of the joint zero slope hypothesis is strongly rejected in all the regressions. Overall, there is no apparent trend in the characteristics of households to use credit purchases. This suggests that retailers may not have put weight on the volatility of earnings when trading with factory workers on credit. This is consistent with the retailers’ motivation to secure their local customers discussed earlier, and with the fact that factory workers commonly used credit purchases. [^82] Despite this, low-income households might have had more difficulty using credit purchases. A famous essay drawn by a primary school student in the 1930s depicts an example of a rice retailer who refused to buy rice under credit purchase (Toyoda 1969, pp. 346–355).
The results so far indicate that retailers have been generous in the earnings volatility risk. This, in turn, implies that retailers may factor risk in the form of interest into their loans or retail prices. Unfortunately, there are no systematic statistics on the interest in credit purchases. Thus, I calculate the risk premiums backward by differencing payments and purchases. This backward calculation technique requires information both on the price and type of retailers. However, budget books usually document price-by-item information but not price-by-retailer information. This is problematic because the same items were often sold in different types of retailers: for example, eggs were sold in greengrocers (yaoya) in many cases but sometimes sold in variety stores (aramonoya). Fortunately, I found that a specific month in a budget book has documented the price-by-retailer information. [^83] The OISR, Archives of the Tsukishima Survey (THBS #4, February 1919). According to this budget book, the reimbursements were made semi-monthly, as seen in Figure 6b. They bought various goods in greengrocers, fish stores, and variety stores between 1st and 14th February 1919 and paid exactly the same amounts on the 15th. [^84] While the retailers used credit purchases from wholesalers (footnote 78), the repayment period was often set around one month (Tokyo City Office 1932, pp. 157-159). Therefore, this makes sense for consumers to make repayments at semi-monthly or monthly intervals. In the latter half of February, the household paid the same amounts to greengrocers and variety stores on 28th, but less amount to the fish stores. Since the budget book for March does not have information on the type of retailers, it is difficult to know exactly how the reimbursement shortfall was handled. However, this case suggests that the risk premiums may have been included in the retail prices but not in the form of interest.
Finally, I assess the loss for the retailers in credit purchases. According to the statistics on the losses on credit sales documented in the 1930 Survey Report, almost all of the retailers had suffered losses in credit purchases. For example, 99% of rice and 97% of fuel retailers more or less lost their sales. This means that credit purchase was inherently a loss-making sales method for the retailers. Since systematic statistics are unavailable to clarify the percentage of total profits corresponding to these losses. Thus, it is impossible to discuss how severe these losses are for the retailers. However, an important fact is that these losses were usually capped by 5% of total sales. For example, the available statistics suggest that 87% of rice and 65% of fuel retailers suffered losses amounting to 5% of sales. This supports the evidence that while the retailers could not avoid losses in credit sales, they might have set prices that anticipated predicted losses to some extent. This seems to be a plausible sales strategy for retailers to deal with adverse selection issues under imperfect consumer information. Online Appendix C.4 summarizes the statistics on the losses in credit purchases available from the 1930 Survey Report in detail.
To summarize, the available documents show that retailers competed with other retailers in the same sales area. Under this condition, they provided regular consumers with the opportunity to trade on credit without interest for at least one month. The 1930 Survey Report indeed stated that “intense competition among competitors has led to a high frequency of credit purchases” (Tokyo City Office 1932b, p. 415). This accessibility to credit purchases could help workers manage consumption after unexpected individual setbacks in the short term. [^85] Another potential explanation for the popularity of credit purchases might be thefts. However, the number of thefts $100$ people in 1919 is $0.5$ in Tsukishima, which is close to or slightly lower than the city average of $1.3$ (National Police Agency 1920b, pp 53–54; 244–245). In fact, Tsukishima Police Department does not mention that Tsukshima was an area with a particularly high rate of thefts (Tsukishima Police Department 1976, pp. 244–245). This suggests that this crime channel is not relevant to explain the use of credit purchases.
6 Conclusion
This study investigates the consumption smoothing behaviors among working-class households in Tokyo. I digitized the daily longitudinal budget survey on the factory worker households conducted in Tsukishima and implemented a systematic econometric method to obtain comparable estimates for income elasticities. The estimated elasticity for total consumption indicates that the factory worker households suffered certain consumption losses when they faced idiosyncratic income shocks. Despite this, I found evidence that while expenses on dispensable categories are likely to be cut in response to the income losses, expenses on indispensable categories are partially smoothed.
The estimated elasticity for total expenditure is in a similar range to those in rural economies in developing countries, although only a few available references exist. According to Ravallion and Chaudhuri (1997), the elasticity for the rural households in Indian villages between 1975 and 1984 was roughly $0.5$. As for the urban households, the estimated elasticity is $0.4$ for Bangkok between 1975 and 1990, which had a similar GDP per capita to 1920 Japan (Townsend 1995). Since my estimate is from a specific occupation, skilled factory workers, one must be careful when making comparisons with these estimates based on urban samples, which include a variety of occupations. At least, however, the factory workers in the historical metropolis of Japan around WWI had been relatively vulnerable to the idiosyncratic income shocks.
A recent study found an income elasticity of $0.39$ for factory workers’ households in Osaka city circa 1920 (Ogasawara 2024). My estimate from the adjusted monthly panels is slightly larger than this estimate. Although the measurement errors in aggregate consumption and income in the monthly panel dataset may explain the discrepancy in these estimates, if taken literally, this implies that the Tsukishima factory workers were more vulnerable than those in Osaka. This indicates that higher levels of urbanization do not necessarily correspond to a greater availability of risk-mitigation institutions. Therefore, inter-city and cross-national comparisons of income elasticities and analysis of the institutional characteristics behind the consumptions there would be another subject of future work.
The finding on the attenuation bias due to measurement errors is potentially important because it may prompt a reinterpretation of existing results. Errors in the timing of payroll can result in inconsistencies in the aggregate data for income and expenditure, albeit the household-level studies often use aggregate data. In this case, modifying the aggregation period to align with wage payments can reduce attenuation bias. Even if the aggregation period is unadjustable, clarifying payroll timing and discussing potential bias is useful.
The results from mechanism analyses provide suggestive evidence that credit purchases played an important role in mitigating adverse idiosyncratic shocks on consumption. Although the institutional details of credit purchases should be investigated further, this finding offers a new historical view of the interdependence between consumers and local retailers. Understanding the local retail system would be important in comprehending regional consumption behavior and, in turn, in understanding the improvement in living standards in prewar Japan.
Appendices
Appendix Appendix A Background Appendix
A.1 Developments of the Cities
(a) Number of people in the large cities
(b) Share of people in the large cities (%)
Figure A.1: Developments of the Large Cities in Japan
Note: This figure illustrates the development of large cities in Japan between 1889 and 1920. Figure A.1a shows the number of people in Tokyo, Osaka, Kyoto, and Nagoya cities. Figure A.1b shows the share of people in Tokyo (diamond) and the cities with more than $50,000$ people (circle) in total population in Japan.
Sources: Umemura et al. (1983, pp. 303–305); Umemura et al. (1988, pp. 166; 168).
Table A.1: City Characteristics:
Tokyo, Osaka, Kyoto, and Nagoya
Name of cities
Tokyo Osaka Kyoto Nagoya
Panel A: Population and Workers
Population 2,173,201 1,252,983 591,323 429,997
Male workers 768,633 476,229 189,922 138,275
Panel B: Industrial Structure (share in male workers)
Agriculture 1.0 0.8 2.2 1.8
Fisheries 0.1 0.1 0.0 0.4
Mining 0.4 0.3 0.2 0.2
Manufacturing 44.5 45.6 49.0 45.8
Commerce 32.4 34.0 32.3 29.6
Transport 7.6 10.6 5.5 8.0
Public service and professions 11.3 6.7 9.0 11.3
Housework 0.3 1.8 0.1 0.0
Other industry 2.4 0.1 1.5 2.8
Panel C: Local Infrastructures
Number of electrical lights per people in 1921 0.9 0.8 0.9 0.9
Accessibility to modern water supply (%) 78.9 80.5 46.5 39.4
Panel D: Meteorological Features
January 1918 – December 1923
Average daily temperature 14.1 15.4 14.1 14.6
Average daily precipitation 4.7 3.9 4.6 4.8
February – April only
Average daily temperature 8.0 8.8 7.8 8.3
Average daily precipitation 4.2 3.4 4.1 4.3
Notes:
1. Panel A: The number of people and male workers measured in the Population Census conducted in October 1920 are listed.
2. Panel B: Occupations are classified based on the industrial classification of the 1920 Population Census. Unemployed people are not included in the calculation.
3. Panel C: The number of electrical lights measured at the end of 1921 is divided by the number of people in the 1920 Population Census. The number of houses accessible to water supply measured at the end of September 1920 is divided by the number of total households measured in the 1920 Population Census (%).
4. Panel D: The average daily temperature and precipitation between 1st January 1918 and 31st December 1923 and between 1st February 1920 and 30th April 1920 are listed.
Sources: The number of people in 1920 is from the 1920 Population Census (Statistics Bureau of the Cabinet 1928, pp. 186; 214; 222; 225). The number of male workers and their industrial structures are from the Statistics Bureau of the Cabinet (1929a, pp. 20–21; 26–29). The number of houses accessible to the water supply is from the Water Works Association (1922, pp. 69–70). The total number of households is from the 1920 Population Census (Statistics Bureau of the Cabinet 1928, pp. 186; 214; 222; 225). The number of electrical lights is from the Ministry of Posts and Telecommunications (1922, p. 558) The daily temperature and precipitation are from the Japan Meteorological Agency (database).
Figure A.1a illustrates the developments of the large cities in Japan between 1889 and 1920. Table A.1 summarizes the scales, industrial structures, infrastructures, and meteorological features of the large cities.
A.2 Financial Institutions in the Cities
Several financial institutions were available for working-class households in the cities around 1920.
A representative savings institution is the postal savings (yūbin chokin). Postal savings was initially designed as a small savings institution for low-income households, and the share of postal savings to total savings was roughly $10$% in the 1910s (Tanaka 2018, p. 29). However, the number of postal savings accounts per total population in the early 1910s exceeded $20$% (Tanaka 2018, p. 33). [^86] James and Suto (2011) show that the saving rate in prewar Japan was higher than that of the U.S. As shown in Panel A of Table A.2, roughly $1.4$ million people had accounts in Tokyo in 1920. This covers approximately $67$% of the entire population in Tokyo and the average savings amount per capita was $64.5$ yen. In Kyobashi Ward, the average figure for the postal savings accounts was $44.5$ yen at the end of 1919 (Tokyo City Office 1921, pp. 892–893). This amount is less than the average monthly earnings of the skilled-factory workers (Section 3.2). Unfortunately, the statistics on postal savings by depositors’ occupation are unavailable. [^87] There are some statistics for the entire Japan. According to the Postal Savings Bureau (1924, p. 72), the share of depositors in the manufacturing sector was $4.8$% (the share based on the deposit amount is $5.5$%). However, this implies that, on average, postal savings may not provide sufficient temporary income.
Savings bank (chochiku ginkō) is another famous savings institution for working-class households in the 1910s (Ito and Saito 2019, pp 79–81). Panel A of Table A.2 shows that the number of people having savings accounts was roughly $1.46$ million, accounting for approximately $67.3$% of total people in Tokyo at that time. Although this figure is similar to the postal savings, the workers in the manufacturing industry were less likely to rely on the savings banks than the workers in the other industrial sectors. Table A.3 summarizes the number of depositors and the amounts of deposits in savings banks by occupation in Tokyo City (Panel A) and Kyobashi Ward (Panel B). In Kyobashi, the commerce and miscellaneous sectors shared roughly $84.8$% of all the depositors, whereas the manufacturing sector had only $11.5$%. This is unchanged if I use the amounts of deposits to calculate the shares. Tokyo Institute for Municipal Research stated that this tendency captures the fact that savings banks are mainly designed as financial institutions for low-income households (Tokyo Institute for Municipal Research 1925b, p. 81). Thus, skilled factory worker households, like the THBS households, were less likely to use the savings banks. This is consistent with the fact that Tsukishima did not have any savings bank branches, while there were two postal offices. [^88] There were 123 savings banks in Tokyo City in 1922, and 10 of them were in Kyobashi Ward (Tokyo Institute for Municipal Research 1925, p. 76). However, Tsukishima, a representative manufacturing area, did not have savings banks at that time.
Mutual loan association (mujin) is another type of savings institution. The members of this association pay a fixed amount of money several times, and each member withdraws the total amount of payments each time under specific rules. The lotteries or bidding system decided which members could withdraw money each time. This means that the workers could not use this institution to deal with the acute negative income shocks. [^89] This is the most representative mutual loan system called tsūjyō (i.e., common) mujin. There was a different type called sueoki chokin (or tsumitate kai) that people could receive their stakes with interests after a fixed amount had been accumulated for a certain period (Tokyo Legislative Research Association 1915, pp. 3–4). Clearly, both are also not flexible for compensating the income shocks. Instead, it had been used to make a certain amount of money among small business owners and merchants. A report suggests that the workers in the manufacturing sector shared only $10$% of all the mutual loan associations in Tokyo (Tokyo City Office 1935, p. 25). Moreover, the Tokyo Chamber of Commerce stated that the mutual loan association was used for business and rarely for household consumption (Tokyo Chamber of Commerce 1918, p. 121).
Cooperative societies were also available for worker households. Cooperative societies in Tokyo used credit as their payment method: households could buy goods in advance and pay at the end of the month or twice (middle and the end of) a month. The commodity prices were similar to or slightly lower than those in the retail shops (Central Federation of Industrial Associations 1925, pp. 18–19). However, the cooperatives only covered a few citizens. There were 26 cooperative societies in Tokyo prefecture in 1924, and 23 of them were organized by the citizens and workers (Central Federation of Industrial Associations 1925, p. 57–59). Among these 23 cooperatives, two were large societies with roughly 4000 members, and the others were small societies with less than 100 members. The total number of accounts for only a few percentage points in the workers in the city, given that the number of male workers was more than 768 thousand (Table A.1). In fact, most of the THBS households had not used the cooperatives.
An alternative borrowing institution is the credit purchase in retail shops. The THBS households had frequently used credit purchases in various retail shops for rice, fish, vegetables, alcohol, firewood, charcoal, and medicines. The details of the credit purchase are little documented, and the systematic statistics on the credit purchase are also unavailable. This must be because the information on the transactions is personal. Despite this, an available document shows that credit purchases are used in various retail shops (Section 5). The THBS budget data show a consistent tendency for credit purchases to be widely used among factory worker households.
Pawnshops were the popular lending institution among low-income working-class households. Lenders did not need to screen the borrower’s credit, and the borrowers did not worry about incurring heavy debt as the primary articles were inexpensive clothes (Shibuya et al. 1982). In addition, the interest rates were regulated by the Pawnbroker Regulation Act of 1895, and the average redemption rate was substantially high under the lower interest rates. Among $88$ pawnshops in Kyobashi, the average redemption rate was approximately $94$% ($401,222$ cases/$427,265$ cases) in 1920 (Tokyo City Office 1922b, pp. 888–889). The accessibilities were higher than the other lending institutions. The number of pawnshops in Tokyo was $1,334$ in 1918 and $1,261$ in 1919, respectively. The total number of cases was $8,226,883$ and $7,573,406$ each year. These figures are more than three times greater than the entire population in Tokyo City. This means that the accessibility to a pawnshop was relatively high for the workers. As described in the main text, the number of pawnshops in Tsukishima was 13, whereas that of moneylenders was five.
However, the inexpensive article means that pawnshops were used to deal with very short-run necessities only. Among the pawnshops of Kyobashi, for example, the average amount per case was $7.8$ yen, which was $12$% of the monthly earnings of the THBS households. While the pawnshops were popular lending institutions, they could not offer a sufficient amount of temporary income for relatively large idiosyncratic shocks. According to the statistics on the pawnshop users in Tokyo prefecture in 1923, the share of users in the agriculture, commerce, and manufacturing sectors was 1.7%, 19.5%, and 15.6%, respectively. In turn, the most representative users were the day laborers and workers classified as “miscellaneous”, accounting for $63.4$% (Tokyo Institute for Municipal Research 1926, p. 25). In fact, the Tokyo City Office stated that the pawnshops were mainly used among low-income (poor) working-class households (Tokyo City Social Affairs Bureau 1921, p. 9). This indicates the evidence that the primary pawnshop users were not the skilled-worker households like the THBS households.
Another lending institution is the money lenders (kinsen kashitsuke gyō). Generally, business owners used the money lenders. The interest rates in the money lenders were very high compared with those in the pawnshops (Shibuya 2000, pp. 184; 248). Tokyo Institute for Municipal Research was concerned about the increasing number of unscrupulous companies arrested by the National Police Agency (Tokyo Institute for Municipal Research 1925b, pp. 106–108). Given this fact, while there were two money lenders in Tsukishima, they might have been less likely used by the factory worker households.
Credit unions (shinyō kumiai) are another institution available for working-class households. In Tokyo City, there were 215 unions in 1923. According to the survey of the 184 unions among them, however, the number of union members in the manufacturing sector was only $5,662$ (Tokyo Institute for Municipal Research 1925c, p. 20). This accounts for, at the most, only a few percentage points of the male workers in the manufacturing sector in Tokyo measured in the population census of 1920 (Online Appendix Table A.1).
Table A.2: Financial Institutions in Tokyo, Osaka, Kyoto, and Nagoya Cities in 1920
Name of cities
Tokyo Osaka Kyoto Nagoya
Panel A: Savings Institutions
Postal saving
Number of depositors 1,462,604 725,642 305,927 172,183
         Population percentage (%) 67.3 57.9 51.74 40.04
Total amounts (yen) 94,330,789 34,742,000 12,036,439 9,540,725
         Amount per people (yen) 64.50 47.88 39.34 55.41
Savings bank
Number of depositors 1,461,389 1,648,090 663,831 482,960
         Population percentage (%) 67.25 131.53 112.26 112.32
Total amounts (yen) 48,682,260 50,674,000 28,495,407 24,014,781
         Amount per people (yen) 33.31 30.75 42.93 49.72
Panel B: Pawnshop
Number of shops 1,253 867 441 412†
Number of units lent 6,042,501 2,027,690 1,251,283 319,609
Amount of loans (yen) 38,667,993 3,785,186 4,424,357 2,983,250
Amount per loan (yen) 6.40 1.87 3.54 9.33
Notes:
1. Panel A: The statistics on postal savings were measured at the end of March 1921. The statistics on savings banks were measured at the end of 1920.
[…]
Sources: Panel A: Tokyo City Office (1923, pp. 969; 977); Osaka City Office (1922, pp. 6(61); 6(67)); Kyoto Prefecture (1922, p. 257); Kyoto City Office (1926, pp. 188); Nagoya City Office (1922, pp. 286; 290). Panel B: Tokyo City Office (1923, pp. 888–889); Osaka City Office (1922, pp. 6(68)–(69)); Kyoto City Office (1922, p. 191); Tokyo City Statistics Division (1926b, p. 12).
Table A.3: Deposits in the Savings Banks in Tokyo City by Occupations
Occupations
Agriculture Commerce Manufacturing Miscellaneous Total
Pane A: Entire Tokyo City
      Number of depositors 76210 550784 166498 619963 1413455
      Amount of deposits 2002124 19112713 6663091 19809591 47587519
      Number of depositors (%) 5.4 39.0 11.8 43.9 100.0
      Amount of deposits (%) 4.2 40.2 14.0 41.6 100.0
Pane B: Kyobahi Ward
      Number of depositors 3809 34665 11664 51481 101619
      Amount of deposits 130781 1218081 450805 1696086 3495753
      Number of depositors (%) 3.7 34.1 11.5 50.7 100.0
      Amount of deposits (%) 3.7 34.8 12.9 48.5 100.0
Notes: This table summarizes the number of depositors and the amount of deposits in the savings banks by occupation. In the document, the miscellaneous category is not clearly defined but named as “Miscellaneous (zatsu gyō)”. These statistics were measured on 31st December 1919. Panel A and B list the statistics for Tokyo City and Kyobashi Word, respectively.
Source: Tokyo City Office (1922a, pp. 864–865).
Figure A.2: Pawnshops and Money Lenders
Notes: This map shows the location of the pawnshops and moneylenders in Tsukishima. The bullseye indicates the pawnshop (shichiya). The black circle shows the money lenders (kanekashi gyō). The white circle shows the rickshaw station (kuruma yado).
Source: Department of Health, Ministry of the Interior 1923c, tenth map. The tone was adjusted by the author using Adobe Photoshop 24.7.0.
A.3 Development of Tsukishima
Figure A.3: Tsukishima: Ishikawajima, Tsukudajima, Shin-tsukudajima, and Lots 1–3
Notes: Tsukishima contains Ishikawajima, Tsukudajima, Shin-tsukudajima, and Tsukishima Lots 1–3. This map does not illustrate the northernmost part of Ishikawa jima (see Figure 3 for this part). “Ferries” shows the Tsukishima to Kyobashi Ward sea routes: Tsukuda, Tsukishima, and Kachidoki routes (right to left). “Aioi bridge” links Shin tsukudajima and Fukagawa ward.
Source: Department of Health, Ministry of the Interior 1923c, 15th map. The author modified and adjusted the tone using Adobe Photoshop 24.7.0.
Tsukishima (transliterated as “moon island”) is a reclaimed land on a sandbar of the Sumida River. Ishikawajima and Tsukudajima (“jima” means island) at the northern end of Tsukishima are the starting point of its history (Figure A.3 for map). The following description of this section is based on Chūō Ward (1994, pp. 122–124).
Ishikawajima was built in 1626 by a boatman, Hachizaemon Ishikawa, who received a fief from the Shogunate. In 1790, the Ishikawajima Prison was built on Ishikawajima. This was a facility for the reintegration of minor offenders and those without lodgings and was called a Ninsoku Yoseba. After Matthew Perry arrived in Japan in 1853, the Shogunate ordered the Mito Clan to build a shipyard on Ishikawajima. A month later, the groundbreaking ceremony was held for a Western-style wooden sailing ship delivered to the Shogunate in 1856. In 1866, the Ishikawajima Lighthouse was built on the west bank of the Yoseba and served to signal the position of ships entering the Nihonbashi direction. In 1870, the Ninsoku Yoseba became a place of imprisonment (Toba) and was renamed the Ishikawajima Jail Station of the Metropolitan Police Department in 1877. The prison was finally moved to Sugamo in Kitatoshima County in Tokyo prefecture, and the penitentiary on Ishikawajima was closed in 1895.
Adjacent to Ishikawajima is called Tsukudajima, where fishermen from “Tsukuda” in Osaka prefecture settled in 1644 after receiving land from the Shogunate (i.e., the island was named after their hometown, Tsukuda). In 1645, the “Tsukuda ferry” began, and throughout the Edo period, Tsukudajima was known as a fishing island for whitefish for the Shogun’s household. In 1863, the Tsukudajima Battery was built on the south side of Tsukudajima, one of 13 batteries used to defend Edo Bay. This battery later became the starting point for the reclamation of Lot 1 of Tsukishima.
In 1884, the Port of Tokyo Waterway Dredging Project was initiated. This was to secure the passage of ships by sweeping away the sediment in the shipping channel, and the Tsukishima landfill was created using the deposit. The first Tsukishima site was completed in 1891, and the “Tsukishima ferry” began the same year. The second site was created in 1894; the Shin (new) tsukudajima was built in 1896. From the west shore of Shin tsukudajima, the island overlooks the Boso Peninsula in Chiba Prefecture, making it a favorite vacation spot for writers. Cultural figures such as Toson Shimazaki and Kaoru Osanai stayed here from 1907 to the early Taisho period. In 1903, the Aioi Bridge was completed, bringing water and electricity to Tsukishima and improving transportation convenience, which contributed to industry development. [^90] Until then, Suzukiiya had been selling water to supply the entire island using a ferryboat. The “Kachidoki Ferry” began to lead to Lot 2 in 1905. In 1913, Lot 3 was finally completed.
A.4 Housings in Tsukishima
Since the Tsukishima Survey included a housing survey (Table B.1), the housing types in the entire island can be determined. Table A.4 summarizes the results of this survey. The description in this section is based on the Department of Health, Ministry of the Interior (1923a, pp. 69–76).
$88$ ($46$%) of the detached homes are in Tsukudajima. Tsukudajima is reported to be “superior to other areas in terms of housing quality, and the percentage of inferior housing is meager”. Shin-tsukudajima had a low percentage of detached homes. The report states that it was indeed “inhabited by unskilled laborers” and “occupies the most inconvenient housing status of all the islands.” Lot 1 was a center of life in Tsukishima and is a cluster of numerous residences, mainly tenements. Lot 2 had a relatively high percentage of single-family and two-family houses and appears to have had a character as a “workers’ residential area” than Lot 1. These indicate that there might have been some geographic differences in income distribution regarding housing type. Overall, however, of the $989$ dwellings in Tsukishima, $799$ ($81\%$) are row houses, and only $190$ are single-family homes. This confirms that Tsukishima was an industrial area with many working families. The workers lived in similar tenements and with similar standards of living. Figure A.4 is a photo showing these tenement houses.
Table A.4: Housing Types in Tsukishima
Tenement (nagaya)
(A) Detached (B) 2 families 3+ families Total (A)+(B) %
2-story 1-story 2-story 1-story
Tsukuda jima 88 32 26 6 23 175 146 83.4
Shin-tsukuda jima 16 27 10 36 59 148 53 35.8
Lot 1 58 62 74 73 200 467 194 41.5
Lot 2 28 27 43 9 92 199 98 49.2
Total 190 148 153 124 374 989 491 49.6
Note: This table summarizes the housing types in Tsukishima surveyed on 30th June 1919.
Source: Department of Health, Ministry of the Interior 1923a, pp. 69–76.
(a) 2-story tenement
(b) 1-story tenement
Figure A.4: Photographs of the Tenements in Tsukishima
Note: Figures A.4a and A.4b show the two- and one-story tenements in Tsukishima, respectively.
Source: Department of Health, Ministry of the Interior 1923c, figures 62–63; 66.
A.5 Machinery Factories in Tsukishima
First, I use the results of the Primary School Children Survey included in the Tsukishima Survey to analyze the occupation (industry) of the household heads in Tsukishima. This is a survey of 4th-6th graders at Tsukishima 1st and 2nd Elementary Schools in July 1919. There were 1,173 children in these grades, and the survey included all 1,100 children who resided in Tsukishima. At that time, there were no other elementary schools besides those mentioned above. Thus, this survey is a complete survey for all children in grades 4-6 in Tsukishima. Although it does not cover households without children in grades 4-6, it is undoubtedly a beneficial survey that provides an overview of the households with children that are the main target of this study.
Panel A of Table A.5 shows the industrial structures. The composition ratios are similar to the census-based statistics reported in Panel A of Table 1. It does not specify whether brothers and sisters in the same household were counted twice. Despite this, this similarity confirms that the statistics obtained from this survey are useful in identifying overall trends in family households in Tsukishima.
Panel B of Table A.5 divides the “social classes” of the household heads working in the manufacturing sector into six categories. In the report, large business owners are classified into the capitalist class; small business owners, free enterprise owners, and directors are regarded as the middle class; workers are defined as the working class (Department of Health, Ministry of the Interior 1923a, p. 50). The table reveals at least three things. First, most heads of households are workers, with only 1% of the capitalist class and less than 20% of the middle class. While the definition of “executive” is unclear, “small business owners” must be small factory owners who could have similar earnings to the skilled worker households (Department of Health, Ministry of the Interior 1923a, p. 64).
Table A.5: Industrial Structure and Social Classes in Tsukishima:
Evidence from Tsukishima Primary School Survey
Panel A: Industrial Structure
Name of survey Tsukishima Survey
Survey area Tsukishima
Survey month and year July 1919
Survey subject Heads of all the students in grades 4–6
in Tsukishima 1st and 2nd Primary Schools
# of heads % share
Agriculture 1 0.1
Fisheries 13 1.2
Mining 2 0.2
Manufacturing 623 58.6
Commerce 259 24.4
Transport 72 6.8
Public service and professions 55 5.2
Other industry 38 3.6
Observations 1,100 100
Panel B: Social Class of Heads in Manufacturing Sector
Name of survey Tsukishima Survey
Survey area Tsukishima
Survey month and year July 1919
Survey subject All students in grades 4–6
in Tsukishima 1st and 2nd Primary Schools
Tsukudajima Shin-tsukudajima Lot 1 Lot 2 Total
Large business owner 0 0 5 3 8
Small business owner 1 10 49 13 73
Self-employed 0 0 0 0 0
Executive 3 13 28 7 51
Worker 23 94 265 108 490
Unemployed/Unknown 0 0 0 0 0
Total 27 117 347 131 622
% share of workers 85.2 80.3 76.4 82.4 78.8
Notes: This table summarizes the statistics from the complete survey on the 4th-6th graders in Tsukishima1st and 2nd Primary Schools in July 1919. Panel A summarizes the heads’ industries, and Panel B sorts the “social classes” of the heads in the manufacturing sector, respectively. The report does not specify the definition of each “social class”.
Source: Department of Health, Ministry of the Interior 1923a, pp. 386–387.
In November 1920, there were 214 factories in Tsukishima, 168 of them (roughly $80$%) were classified as machinery factories. The official report documented that the wooden pattern factories could be included in the machinery sector because they were incorporated into the casting processes (Department of Health, Ministry of the Interior 1923a, p. 388). The share of machinery factories becomes approximately $90$% if I include those 24 wooden pattern factories.
Panel A of Table A.6 uses the manufacturing survey in the Tsukishima Survey to classify the 168 factories by size and industrial sector. First, more than $90$% of the plants are in the three sectors of machinery and equipment manufacturing, can manufacturing, and blacksmithing, with only $15$ plants in the casting sector. 113 ($67$%) of the machinery factories had workers less than 15. Second, many smithing factories have fewer than 15 workers. Hence, almost all of Tsukishima’s smithing industry was considered to be small businesses. [^91] The report suggests that 43 of the 47 smithing factories were the “town factories” or “workplaces”. The definition of “town factory” is not clear in the report (p. 392). Since they were always counted more than the number of factories with 15 or fewer workers, they are likely to include smaller factories operated as small businesses.
Panel B of Table A.6 classifies the types of the 620 workers based on the results of the primary school survey used in Table A.5. There were 455 skilled and 165 unskilled workers. Notably, these 620 workers include non-industrial workers: 490 industrial workers, 13 commercial, 61 in transportation, 16 in public service free trade, 38 in other gainful employment, and 2 in agriculture and mining in total. Since the workers in the industrial sector should include factory workers, workers in the cottage industry, and handicraft workers, the day laborers are considered to be classified into non-industrial workers (in fact, $620$ workers $-120$ day laborers $=490$). Therefore, roughly 90% of factory workers ($455/490$) can be regarded as skilled workers, $65$% of whom worked in factories, $17$% in cottage industries, and $18$% as craftsmen. Since the workers in cottage and handicraft industries are considered to work in small factories, $65$% of these workers might have worked in the middle- to large-scale factories (with 15+ workers).
One must be careful that this survey only covers the heads with primary school children aged 9–11 (i.e., 4th to 6th grades), suggesting that most of these heads are in their 30s. Therefore, the share of skilled workers to the entire workers should be smaller than the $90$%. At least, however, it is plausible to see that most of the Tsukishima workers in their 30s were skilled workers at that time.
Table A.6: Scale of Machinery Factories and Type of Workers in Tsukishima
Panel A: Scale of Machinery Factories
Name of survey Tsukishima Survey
Survey area Tsukishima
Survey month and year November 1920
Survey subject All machinery factories
Machinery and Can Casting Smithing Total
equipment
# of workers $\geq 15$ 22 15 10 8 55
# of workers $<15$ 44 25 5 39 113
Total 66 40 15 47 168
% shares 39.3 23.8 8.9 28.0 100
Panel B: Type of Workers (All Industries)
Name of survey Tsukishima Survey
Survey area Tsukishima
Survey month and year July 1919
Survey subject All students in grades 4–6
in Tsukishima 1st and 2nd Primary Schools
Skilled workers Unskilled workers
Factory workers 295 Factory workers 42
Cottage industry 76 Cottage industry 3
Handicraft industry 84 Day laborer 120
Total 455 165
Notes: Panel A summarizes the number of machinery factories by the scale and type of the factory. Statistics are from the manufacturing census conducted in the Tsukishima Survey (1st November 1920). The threshold of the scale (i.e., 15 workers) is based on the application of the Factory Act enacted in 1916.
Panel B sorted the heads by their skill level and job type. Statistics are obtained from the Complete Survey on the 4th-6th graders in Tsukishima1st and 2nd Primary Schools in July 1919.
Source: Panel A: Department of Health, Ministry of the Interior 1923a, pp. 386–387; 410–414. Panel B: Department of Health, Ministry of the Interior 1923a, p. 149.
A.6 Works in the Machinery Factories
Tsukishima Survey includes the factory survey that counts the number of machinery factories in Tsukishima. Panel A of Table A.6 summarizes the number of machinery factories by type of plant. The machinery and equipment manufacturing factories shared 39% of all the machinery factories. Typical jobs for workers in machinery and equipment manufacturing plants are finishers (or fitters) and turners. A finisher is “a skilled worker responsible for finishing, assembling, and commissioning machine products” and is “one of the finest skilled workers because it requires skills in all aspects to correct defects in machine making” (Department of Health, Ministry of the Interior 1923a, p. 393). The turner is the “middle axis of machinists,” to which most semi-skilled and skilled workers belong. Becoming a full-fledged finisher and lathe operator takes at least five years. Generally, apprentices enter the workforce around 14–18, but they do not work in the same factory for many years within the apprenticeship system. After learning the skills for 1-2 years, they often move to another factory to improve their skills (Department of Health, Ministry of the Interior 1923a, p. 394). Wages are paid on a piece-work basis or on a time basis. In Tsukishima, a time contract was adopted, and payment was based on the “residual time” method. Under the “residual time” method, “if the work is completed within the contracted time, the remaining time is considered the bonus.” For example, if 12 hours are contracted, and the work is completed in 10 hours, 2 hours are paid as a bonus (pp. 396-397). Wages were paid by dividing one month into two periods (first and second semester). The average work hours are reported to be roughly 10 hours, with occasional all-night work (p. 402). Overtime hours were reportedly around 2 hours per day (Department of Health, Ministry of the Interior 1923a, p. 402).
Next, I summarize labor in the canning factory, which accounts for 24% of all Tsukishima factories. The canning factory is engaged in shipbuilding, bridges, steel frames for construction, and other works, including fabrication and repair of gas reservoirs (p. 406). Workers in a canning factory are called “boiler makers” and are typically involved in bending steel plates and joining them with rivets. Workers are divided into rivetter, holder-up, plater, angle-ironsmith, and caulking. The work is done in groups, with the skilled angle-ironsmith and caulking. The “bōsin” of the rivetter gives the instructions as the head of the group. Wages are paid on a piece-rate contract basis. Because of the contracted wage, the cannery workers’ income is unstable; thus, they are paid a fixed wage (daily wage amount) during the off-season.
Smithing is the second most common industry, accounting for 28% of the total. Panel A of Table A.6 reveals that most of the smithing industry is relatively small scale. They make boulders, nuts, rivets, lathe stops, jigs, and spanners for can manufacturing and construction. They were subcontractors who manufactured metal parts to supply the larger factories in Tsukishima (Department of Health, Ministry of the Interior 1923a, p. 412). There are three types of jobs: yokoza, sakite, and tatara-fuki (bellows-blower). The yokoza is self-employed skilled worker (oyakata), while the other two are young unskilled laborers. In small operations, wages were based on a percentage based on voluntary business discussions with the proprietor, and working hours were not regular (pp. 411–413).
Appendix Appendix B Data Appendix
B.1 Tsukishima Survey: The First Social Survey in Japan
Kawai (1980) summarizes the history of social research methodology. The methodology of social research in Japan was formed in the 1930s. Social research methodologies had been developed in the 1910s and 1920s (Kawai 1980, pp. 47–48). During this period, “numerous social surveys were conducted by various survey entities, including government agencies, research groups, local administrative bodies, private organizations, survey research institutes, and researchers” (p.48). This is because “around WWI, there were many problems in labor, lifestyle, urban, health and sanitation, public safety, and civic due to the heavy and chemical industrialization, the monopolization of capitalism, and urban expansion.” The Report on Field Survey in Tsukishima, Kyobashi Ward, Tokyo (called Tsukishima Survey) [^92] Takano himself, who was responsible for overseeing the survey, refers to this field survey as the “Tsukishima Survey” (Department of Health, Ministry of the Interior 1923a, p. 6). Considering this, I call the Tsukishima Survey throughout this paper. used in this study is the most pioneering of these surveys of urban communities. Sekiya (1970, p. 43) regards the Tsukishima Survey as a pioneering household survey in Japan.
First, I summarize the investigators, research questions, and subjects of the Tsukishima Survey. The person responsible for the Tsukishima Survey was Iwasaburō Takano (hereafter “Takano”), a professor of social statistics and social policy at the Law School of Tokyo Imperial University. Takano became a professor of economics at Tokyo Imperial University in 1919, and became the director of the OISR in 1920. Takano placed great importance on gathering objective statistics about workers through social surveys (Sekiya 1970, p. 33). He had the experience of conducting the first household survey in Japan in 1916.939393This is called the “Survey of Household Income and Expenditure of 20 Artisans and Workers in Tokyo” and was inspired by Charles Booth (Miyoshi 1980, p. 33). Although the original documents of this survey were preserved at Tokyo Imperial University, those were destroyed in the Great Kantō Earthquake (Takano 1933, pp. 727–728). More precisely, before this household survey, Takano conducted a field survey in a poor area of Tokyo when he was a student at Tokyo Imperial University. The findings were reported in April 1894 as an exercise entitled “East London in Tokyo” (Kawai 1981b, p. 3). Takano has been a member of the seventh subcommittee (Rural Sanitation Subcommittee) of the Ministry of Home Affairs’ Health and Sanitation Investigation Committee since 1916.949494The Health and Sanitation Investigation Committee of the Ministry of Home Affairs was established on June 27, 1916, under the second Shigenobu Okuma Cabinet by issuing Imperial Ordinance No. 172. Initially, 34 members were appointed, and the first meeting was held on July 8. The reason for the establishment of this committee was based on the need for a policy response regarding the establishment and institutionalization of various investigation organizations and the penetration of the capitalist economy during the Meiji period. Another reason is the requirement for a public health response to the rising mortality rate, exemplified by the tuberculosis epidemic (Kawai 1981a, pp. 10–13). In 1917, the Seventh Section evolved into the Section on Urban and Rural Sanitary Conditions (Kawai 1981a, pp. 12–13). On October 22, 1918, Takano submitted a proposal to the subcommittee for a field survey of urban sanitary conditions, which was approved. The survey aimed to determine the housing conditions, household budget status, and health status (child health, production, stillbirths, deaths, and diseases) in the areas where skilled-worker families congregated (Sekiya 1970, p. 6). Initially, the survey area was planned to be Yanagibashi Yokogawa-chō in Honjo Ward. However, at a committee meeting on November 23, the survey site was changed to Tsukishima in Kyobashi Ward, which had more skilled worker households than in Yokogawa-chō (Kawai 1980, pp. 55–59). [^95] Takeda (2015, pp. 75–76) provides more details on how the survey site was changed. Takano was looking for a commissioned researcher to conduct the survey and asked Sakuzō Yoshino, a colleague at Tokyo Imperial University, to recommend a suitable person. Yoshino approached his former student Kotora Tanahashi, who had become a member of the fraternity club, a mutual aid organization for workers, to accept the investigator position. The diary written by Kotora Tanahashi (“Tanahashi Kotora Diary”) is in the possession of the OISR, and Takeda used the diary as the basis for her analysis of the circumstances surrounding the change of survey location. Tanahashi eventually recommended Yoshitsuru Yamana, who was his classmate in high school and Tokyo Imperial University, as a researcher and considered Tsukishima to be the best research area because of the ease of selecting survey households using the fraternity’s network (Takeda 2015, pp. 76–77). On October 26, Tanahashi and Yamana, together with the Kyobashi Branch of the fraternity club, conducted an on-site inspection of Tsukishima. Based on this inspection, Tanahashi and Yamana approached Takano about locating the survey station in Tsukishima. Finally, Takano and Yamana conducted an on-site inspection of Tsukishima on October 30 and decided on Tsukishima as the survey site.
Although the statistical method of sampling was not yet common at that time, Takano selected the survey sites based on his visionary thinking. In fact, “There can be no dispute that we have chosen Tokyo as the target of our metropolitan survey. However, it is impossible to survey the entire Tokyo area. We have no choice but to conduct a partial survey. For this purpose, it would be the best idea to select a representative area in Tokyo and survey that area narrowly but deeply.” (Department of Health, Ministry of the Interior 1923a, pp. 2–3). Takano further explained the need to focus on the working-class, which makes up most of the citizens, rather than to see the political and commercial center of Tokyo: “Most of Tokyo’s citizens are wage earners who subsist on labor. As is usual in large cities, a large number of people are engaged in manual labor in Tokyo. However, the most genuine type of worker is the skilled worker. Therefore, the area where many families of skilled workers congregate can be regarded as a representative area of Tokyo. And it seemed to me that a study of the social and sanitary conditions in that area could explain a large part of the social and sanitary conditions in Tokyo.” (Department of Health, Ministry of the Interior 1923a, pp. 3).
It is necessary to delve into the point where Takano states that “the most genuine type of worker is the skilled workers” and the fact that the main target of the Tsukishima Survey was skilled workers. [^96] In other words, the Tsukishima Survey was not designed for poor households. Therefore, unlike the surveys conducted in the UK by C. Booth and B.S. Rowntree, no standard of living was set in the survey (Sekiya 1970, p. 42). The description by Kawai (1981a, pp. 13–18) is helpful in this regard. As summarized in Section 2, Japan’s heavy and chemical industrialization throughout WWI led to a large concentration of city workers. As a result, the ratio of small-scale conventional factories based on “master employees” as is often observed in the textile industry decreased. On the other hand, the number of factories with a direct “factory owner-employee” employment relationship increased, particularly in the machinery and equipment, metal, and chemical industries. These workers in the modern industrial sector accumulated skills, formed households, and settled in the cities. It can be said that a new urban worker lifestyle was being formed. In short, although the city had a class of workers engaged in traditional industries, a new type of workers in the modern industrial sector, and a lower class, Takano believed that the skilled workers in the modern industrial sector would play a central role in terms of the labor force, labor movement, and urban life.
In addition, Sekiya notes the rise of the labor movement. As noted in Section 2, while wages rose throughout WWI, the price surge caused relative wages to decline. In 1916, Japan’s first labor protection legislation, the Factory Law, was enacted. However, the number of working days and hours did not decrease, and the number of labor disputes steadily increased (Sekiya 1970, pp. 24–26). Given this, Takano believed that workers should organize labor unions to help themselves with labor problems that could not be resolved through labor protection legislation. Sekiya (1970, pp. 26–40) speculates that Takano’s intention in conducting the Tsukishima Survey was to obtain basic data for solving labor problems, specifically to identify the scope of matters to be solved by workers on their own initiative (or covered by labor protection legislation). His insights suggest that this historical background should be addressed as a reason for Takano’s focus on skilled workers.
Kawai (1981b, pp. 15; 17) assessed the 40 households covered by the household survey included in the Tsukishima Survey as “the representativeness of the sample has remained ambiguous.” There is some mention but no clear description of how the sample households were characterized among the worker households in Tsukishima. It is reported that “to find suitable persons to fill in the form, we must ask for the assistance of school principals, police officers, physicians, and laborers, and we have requested a meeting of these persons at the survey office to explain the purpose of the survey, and to solicit applicants for the distribution of the household account book” (Department of Health, Ministry of the Interior 1923a, pp. 8–9). As explained, the target households were not randomly selected because the statistical sampling method was not prevalent. However, it is plausible that Takano designed this household budget survey in the sampling framework to capture the representative households in Tsukishima. As described in Section 3, the THBS households can be regareded as the standard skilled factory worker households in Tsukishima.
Table B.1: Survey Categories in the Field Study of Tsukishima Survey
Periods of the Chapter in the
field studies published report
1. Social mapping Nov. 1918–May 1919 A2
2. Physical examination of children Jan. 1919–July 1919 3
3. Physical examination of workers June 1919 3
4. Nutrition survey of the worker households May 1919 3
5. Survey of tenement houses Autumn, 1919 3
6. Survey of sanitation occupation June 1919 3
7. Sanitary survey in elementary schools Unknown 3
8. Factory labor survey Unknown 4
9. Labor household survey Nov. 1918–Jan.1920 2
10. Elementary school survey July 1919 2
11. Restaurant survey Dec. 1919 & April 1920 2
12. Field survey of vaudeville Dec. 1919 2
13. Survey of stalls and passersby Dec. 1919 2
14. Photographing Oct. 1920–Nov.1920 A2
Notes: The years/months listed in the first column (“Periods of the field studies”) are based on the description in the Department of Health, Ministry of the Interior (1923a, pp. 10–17). Figures listed in the second column (“Chapter in the published report” ) are based on the table of contents in the Department of Health, Ministry of the Interior (1923a, pp. 1–12; 1923b, pp.1–2). “A2” indicates the Appendix 2.
Sources: Department of Health, Ministry of the Interior (1923a, 1923b, 1923c).
Next, I overview the survey items and methodology. The investigation was primarily divided into “documentary” and “field” surveys. The former was an analysis of the characteristics of Tsukishima, using published statistics such as the Police Department Statistics Book (Department of Health, Ministry of the Interior 1923a, p. 10). This was a pioneering attempt at the time and had influenced subsequent urban studies (Kawai 1980, pp. 61-62). The latter was a field survey including 14 survey items (Table B.1). It is not possible to describe all the survey items in detail, but following the report, they can be divided into four large categories: 1. survey on workers’ livelihood (chapter 2); 2. survey on sanitary conditions (chapter 3); 3. survey on labor conditions (chapter 4); and 4. social mapping and photography (appendix). The survey was conducted over two years, generally from November 1918 to November 1920, although the timing varied depending on the survey items (Table B.1).
A unique feature of the survey methodology is using young researchers as commissioned researchers. The household budget survey and leisure life survey were conducted by Yasunosuke Gonda, a faculty member of the Doitsugaku kyōkai gakkō (Association for German Sciences), who had researched activity photography and entertainment. Tetsuo Hoshino, a medical doctor, was responsible for the sanitary survey. The labor survey was handled by Yoshitsuru Yamana, a graduate of Tokyo Imperial University who contributed to the organization of labor unions. In addition, Toyotaro Miyoshi, a sociology student at the Tokyo Imperial University, and others became temporary employees. In other words, although the Tsukishima Survey was initially planned as an insurance and sanitation survey by the Ministry of Home Affairs, in reality, it was a long-term survey conducted as a joint study by young researchers in different fields, led by Takano (Kawai 1980, p. 63). At the time, the Health Bureau of the Ministry of the Interior was conducting several sanitary surveys of rural areas. Takano may have designed the urban sanitary survey from this background (Kawai 1981a, pp. 8–9). However, based on the idea that “health and sanitation surveys can at least encompass economic and social surveys,” Takano included a wide range of survey items, not limited to sanitation surveys (Department of Health, Ministry of the Interior 1923a, p. 56). Kawai (1980, p. 55) credits the Tsukishima Survey as a noteworthy pioneering concept: “going beyond a mere insurance and sanitation survey, it attempted to accurately depict the state of urban workers’ lives as they became established in the city during modernization, with skilled workers as its core.”
(a) Outside of the office
(b) Inside of the office
Figure B.1: Photographs of the Survey Office in Tsukishima
Note: Figures B.1a and B.1b show the outside and inside of the survey office in Tsukishima, respectively. Source: Department of Health, Ministry of the Interior 1923a, figures 1 and 2.
Takano stated, “To be as complete as possible, it is best to continue the survey over as long a period as possible and to obtain accurate facts by recording the workers’ household income and expenditures over a long period.” For this reason, he believed that “the best method of investigation would be to establish a survey station directly in the survey area and to have a full-time investigator reside there as much as possible, familiarize himself with the area and the people, and conduct the field survey” (Department of Health, Ministry of the Interior 1923a, p. 4). In fact, through the mediation of the Tsukishima police station, a survey office was rented at 9-3 Higashinakadori, Tsukishima, where Yoshitsuru Yamana moved in from November 1918 and contacted worker households (Department of Health, Ministry of the Interior 1923a, p. 7; Takeda 2015, p. 77). The survey office was open from November 1918 to December 1920, including the entire survey period. Figure B.1 shows the photographs of the survey office.
It is also important to note that Takano and his team held weekly meetings at the research center during the survey period. Toyotarō Miyoshi, who was an assistant in the survey, recalls the situation at that time as follows:
After the start of the survey office, all the members gathered once a week with Iwasaburo Takano present to discuss the implementation and progress of the survey and administrative matters. Iwasaburo Takano would always attend each meeting, happily listening to the results of the week’s research, inspecting the materials, providing detailed and thoughtful guidance, and usually ending the study session in a friendly atmosphere. Even now, I can recall the warm face of sensei (i.e., Professor Takano) of those days. It was always a friendly, laboratory-like atmosphere in which the C. Booth and S. Rowntree investigations were occasionally discussed.
Miyoshi (1980, p. 38)
He also states the following about Takano:
When the Tsukishima Survey was launched, the Faculty of Economics (in the Tokyo Imperial University) was becoming independent, and he was swamped both inside and outside the university. However, his enthusiasm for the Tsukishima Survey was extreme, and he was always present on his business trips to Tsukishima. Moreover, he made every effort to carry out his duties with precision.
Miyoshi (1980, p. 40)
Kawai (1981b, p. 28) assessed that there were no interconnections among the survey items mentioned above and that the survey “ended up being several fragmentary surveys rather than a comprehensive and systematic survey of workers’ lives.” This difficulty may be why previous studies have not frequently used the Tsukishima Survey. However, as already explained, Takano did not intend to conduct a large-scale social survey in Tsukishima in the first place. Instead, as he noted, the best method is to investigate a representative area “narrowly but deeply,” the goal was to improve the quality of the investigation rather than the scale of it. In this light, the fact that the researcher lived in the survey area and was constantly organizing and revising the survey results should be appreciated. Even though the Tsukishima Survey is the early social survey in Japan, it is of high quality as a prewar household survey.
Figure B.2: Official Published Report of the Tsukishima Survey
Notes: This photograph shows an example of the official report of the Tsukishima Survey published by the Department of Health, Ministry of the Interior, in 1923.
Source: Department of Health, Ministry of the Interior 1923b, title page. The tone was adjusted by the author using Adobe Photoshop 24.7.0.
Finally, I summarize the format of the published report and the unpublished materials. The results of the Tsukishima Survey were officially summarized in one report published in December 1921 by the Sanitary Bureau of the Ministry of the Interior and in two appendices (Figure B.2 is an example of the front cover). The report was published under the name “Report on Field Survey in Tsukishima, Kyobashi, Tokyo,” and consists of four sections. The first section is a comprehensive review by Takano and outlines the survey. The second section summarizes the living conditions of the workers and was written by Gonda (partly by Yamana). Hoshino wrote the third section and summarized the sanitary conditions of workers. In the fourth section, Yamana summarizes the labor situation in Tsukishima. As for the two appendices, one contains statistical tables related to the above report, and the other organizes social maps and photographs. The correspondence between the survey items and the report is summarized in Table B.1. The Ministry of Home Affairs has already published these materials, which are available at the Gonda Yasunosuke Library of the OISR and the University of Tokyo Library.979797 In the OISR, these are included in the Archives of Gonda Yasunosuke (7-2; 7-3; 7-4). In addition, a second manuscript of the Tsukishima Survey report exists. This is the manuscript of the second series, which was preserved in the home of Gonda Hayao, the second son of Gonda Yasunosuke, and was discovered by Terade Koji. The second series includes a summary of the results of a household survey of elementary school teachers’ households in Tokyo, conducted as a supplementary survey to the Tsukishima Survey (see Terade (1982) for details). The manuscript of this second series was later donated to the OISR and is now available as part of the Gonda Yasunosuke Collection at the Institute (The OISR, Archives of Gonda Yasunosuke (8-12; 8-13; 8-14; 8-15)).
B.2 Original Micro-spreadsheets from the Household Budget Survey in Tsukishima Survey
To analyze the consumption smoothing behavior of THBS households, it is necessary to use the original micro-spreadsheets. The official report does not include household-level information, and a large part of the original forms used to prepare the report have little survived. Fortunately, however, some of the original micro-spreadsheets are kept in the archives of the OISR because Takano was its first director. This study uses the original micro-spreadsheets from the household budget survey (No. 9 in Table B.1), which was used in writing the second section of the official report by Gonda. While they have yet to be publicly available, I received permission from the OISR to use the original forms stored in the archives: the OISR, Archives of the Tsukishima Survey (THBS, unreleased). [^98] I would like to express my sincere thanks to Professor Kazue Enoki, Deputy Director of the OISR, for allowing me access to the documents. I wish to thank Ms. Mika Nakamura, an archivist at the institute, for her careful assistance during the research.
Figure B.3: An Example Page of the Budget Book in the Tsukishima Household Budget Survey
Notes: This photograph depicts an example page from the household expenses account of the THBS. The right page documents the details of costs. The left page includes the details of revenue and receiving and making payments on the goods (actual things, including gifts).
Source: The THBS, account #6, 15th February 1919. The tone was adjusted by the author using Adobe Photoshop 24.7.0.
The budget book used for the household budget survey is named the Kinsendeiri hikaechō (money receipts and disbursements account book). This budget book is designed based on the form used in Takano’s first household survey (Online Appendix B.1). [^99] Department of Health, Ministry of the Interior (1923a, p. 152). The budget book is to be filled out daily and includes four sections: income, expenses, receipts from credit purchases, and receipts and disbursements of actual goods. Most of the entries are in the income and expenses sections, indicating that living was basically conducted through wage receipt and consumption in cash. Figure B.3 shows an example of the budget book.
As described, the method used to select households for the survey was not recorded in detail (Section B.1). According to the report, they initially recruited just over 90 households through the factory, police station, and elementary school in Tsukishima. Of these, 50 households completed the book for at least one month, but 10 were deemed incomplete and not included in the report. In addition to the 40 budget books used in preparing the report, therefore, the OISR has a set of other budget books marked as “discarded”. According to my survey, about six incomplete books of the worker households have been determined to be in this “discarded” category. Therefore, based on the remaining conditions of the books, it is plausible to believe that roughly 50 households participated in the survey. Of the 40 households included in the report tally, two households had recorded for more than one year, 11 households had recorded for 6–12 months, and the remaining 27 households had registered for less than six months (Department of Health, Ministry of the Interior 1923a, pp. 152– 153). Initially, the survey period was targeted to be six months, but the average bookkeeping period for the 40 households was four months.
B.3 Trimming
(a) Daily panel dataset
(b) Semi-monthly panel dataset
(c) Adj. monthly panel dataset
Figure B.4: Structures of the THBS Datasets by Different Time Frequencies
Note: Figure B.4aB.4b, and B.4c show the observations in the daily, semi-monthly, and adjusted monthly panel datasets, respectively. For example, a marker in Figure B.4a indicates an observation in a specific year-month-date cell. Strictly, semi-monthly, and monthly datasets include both panel and cross-sectional units because aggregation shall trim fractions in the edges of the survey period. Figure B.4b shows 35-panel units and 5 cross-sectional units (#29; #31-33; #35). Figure B.4c shows 28-panel units and 6 cross-sectional units (#20; #25; #26; #28; #34; #40). 5 cross-sectional units in the semi-monthly panel do not appear in this figure (#29; #31-33; #35). Note that unit #30 having two semi-monthly cells is not included because those cells are separated into different adjusted months (February and March, respectively).
Source: Created by the author using the THBS sample.
To check whether the statistics from THBS budget books contained any missing information, I digitized all 40 household budget books. Throughout this digitization process, I found that there are a few misreporting issues in the budget books.
First, there is a household that does not include income information (household #37). Although the official report by Gonda reported the average monthly income for this household (Department of Health, Ministry of the Interior 1923a, p. 183), no documentation would allow me to know how this average income was calculated. Hence, I excluded this household’s budget book. Similarly, a household does not include information on the payments to the credit purchases (household #39). Although this household’s budget book documented the amount purchased in credit purchases, I excluded this household’s budget book because the expenditure series becomes systematically smaller due to the omission of the payment information.
Second, the final pages of the budget books are sometimes omitted. This may be because the vendor who bound the budget books accidentally truncated the last page. All these incomplete sheets can not be included in the analysis.
Third, several households did not complete their budget books in the first and/or last month of the survey period. For example, the budget book of household #1 has statistics from 1st January 1919 to 12th January 1920. Similarly, the book of household #29 contains statistics from 26th January to 23rd February 1919. In semi-monthly and monthly datasets, these censored budget books may not be included in the aggregation. In the latter case, for instance, the budget sheets of 26th January and after 11th February 1919 are not included in a semi-month cell because the semi-month cell is defined from the 27th to 11th of the next month in this case. Thus, household #29 only has a cross-sectional observation and can not be included in my analytical sample.
Figure B.4 summarizes the structure of the 40 THBS households. Figures B.4a and B.4b show the observations in the daily and semi-monthly datasets, respectively. As explained, there are five cross-sectional units (#29; #31-33; #35) in the semi-monthly dataset. Therefore, there are 35 units with a panel structure. After excluding two units without complete information on income and credit purchases (#37; #39), I get the 33 THBS households as the analytical sample for my panel data analysis.
Regarding the time dimension, I trimmed all the observations before 12th January 1919 and after 11th November 1919 because the number of surveyed households is substantially small. For example, Figure B.4b shows that the number of cross-sectional units is three or less in both periods, causing substantial fluctuations in the net income. Consequently, my analytical sample includes 33 households measured between 12th January and 11th November 1919. The average number of semi-months per unit is $8.8$ (Std. Dev. $=5.0$). The characteristics of the 33 THBS households are summarized in detail in Section 3.
Regarding the adjusted monthly series, we further exclude several units that have insufficient observations for constructing the monthly panel: 26 THBS households are included in my adjusted monthly panel dataset. Figure B.4c shows the observations in the adjusted monthly dataset.
B.4 Testing the Potential Influence of the Unbalancedness
Table B.2: Results for the Balancing Tests
DV: Indicator variable for the households
with shorter time-series observations
Threshold:
(1) Median (2) 75 percentile (3) 25percentile
($150$ days) ($184$ days) ($79$ days)
Size -0.243 0.238 -0.135
[0.244] [0.254] [0.294]
Children aged 6–12 (%) 0.017 -0.011 0.026
[0.020] [0.019] [0.022]
Children aged 13–16 (%) 0.029 -0.050 0.057
[0.038] [0.045] [0.041]
Men aged 17+ (%) -0.008 0.008 -0.010
[0.019] [0.018] [0.026]
Intercept 1.149 -0.576 -0.164
[2.107] [2.048] [2.805]
Zero slope ($p$-value) 0.710 0.753 0.160
Maximized Log-likelihood -21.8 -18.5 -15.2
Pseudo $R$-squared 0.046 0.046 0.171
Number of households 33 33 33
The results from Probit models are reported. Robust standard errors are in brackets.
Notes: The dependent variable is an indicator variable that takes one if the household has shorter time-series observations. Column 1 uses the median of the number of days per unit as the threshold. The result from the regression using the average number of days per unit ($144$ days) as the threshold is identical to that listed in column 1. Column 2 uses the 75 percentile of the number of days per unit as the threshold. Column 3 uses the 25 percentile of the number of days per unit as the threshold. All family size variables are the time-series average in each unit over the sample period. The proportion of children aged 0–5 years (%) is used as the reference group. Wald $\chi^{2}$ statistics $p$-values for the null of the zero-slope hypothesis are reported in the sixth row. The results from the expanded regressions including the head’s age and earnings are materially similar. The null of the zero-slope hypothesis is not rejected in all the specifications at the conventional level (not reported).
In this subsection, I test whether attrition leads to selection bias in the THBS sample. If the difference in the preference for consumptions predicted the large attritions, there must be statistically significant correlations between the family size variables and attritions. In other words, the units with shorter (longer) time-series observations have different preferences from those with longer (shorter) time-series observations. Table B.2 summarizes the results of the balancing tests (Section 3.2). The dependent variable is an indicator variable that takes one if the household has time-series observations less than the median (column 1), 75 percentile (column 2), and 25 percentile (column 3). In Column 1, all the estimated coefficients on the family size variables are close to zero and statistically insignificant, and the Wald statistics $p$-value suggests the null results. This is unchanged if I use the third and first quantiles in columns (2) and (3), respectively. Therefore, the family characteristics are similar between the households with short-term observations and those with long-term observations. This result supports the evidence that the unbalancedness in the THBS dataset is unlikely to lead to selection bias.
B.5 Comparing Distributions of the Household Size
Figure B.5 illustrates the household size distributions measured in the census and THBS sample. Pearson chi-squared test does not reject the null hypothesis of the equality of distributions with $p$-value of $0.948$. This supports the evidence that the THBS sample has a similar household size distribution to the population of Tsukishima (Section 3.1). As described in Section 3.2, while the THBS sample does not cover households with one person and with a very large number of people, it covers most of the classes in the household sizes with reasonable approximation to the population distribution.100100100The number of households with one people is $362$, accounting for only $6$% of the total number of households. Similarly, the number of households with nine or more people is $294$, accounting for only $5$% (Section 3.2).
Figure B.5: Distribution of the Household Size:
Population Census v. the THBS sample
Notes: This figure shows the share of the household in each household size category to the total number of households in Tsukishima measured in the 1920 Population Census and the THBS samples. Pearson chi-squared test does not reject the null hypothesis of the equality of distributions with $p$-value of $0.948$. The share in each household size measured in the census is used as the theoretical probability in calculating $\chi^{2}$ test statistic (Pearson 1900).
Sources: Created by the author using the Tokyo City Office (1922a, pp. 262–283) and the THBS dataset.
B.6 Estimating Average Head’s Monthly Income
Comprehensive estimation effort on the time series figures of average wages in prewar Japan was done in the Long-Term Economic Statistics (hereafter LTES). Specifically, the LTES estimates the average national salary series of “general occupations,” “manufacturing sector,” and “manufacturing occupations (by medium classification).” For the manufacturing occupations category, the LTES reports the average wage per day of factory workers in Tokyo City and Osaka cities from 1917 to 1922 (Ohkawa et al. 1967, p. 255). These values were estimated using the “Annual Statistics of the City of Tokyo” (hereafter, the ASCT) and “Osaka City Statistical Table.” However, the household survey includes the household head’s monthly income but no information on the daily wage. Hence, the “average daily wage per factory worker” estimated in the LTES is not useful for comparing the heads’ average monthly income from the THBS. In this section, I use available wage and labor statistics to estimate the average monthly income of male factory worker households in Tokyo and compare it to the THBS heads’ average monthly earnings.
To clarify the problem, I first estimate average monthly earnings most simply. Specifically, I use the ASCT (volume 17), which provides wage statistics as of the end of 1918, as a case study. For each industrial sector, the ASCT (volume 17) provides the “average daily wage of male factory workers aged 15 and over,” “the number of male factory workers aged 15 and over,” and “average number of working days per year” for four factory categories: factories with ten or fewer workers, and factories using or not using motive power. Using the “number of workers over 15 years old” in each factory category cell ($c$) as weights, I can get the average wage per month for a worker, for example in the machinery sector, as:
$\textit{Wage}^{15+}_{Machine}=\frac{\sum_{c}\{(\textit{Daily Wage}^{15+}_{Machine,c}\times\textit{Working Days}_{Machine,c})\times\textit{Workers}^{15+}_{Machine,c}\}}{12\sum_{c}\textit{Workers}^{15+}_{Machine,c}}.$ (5)
This value is estimated to be $27.8$ yen. This is the most straightforward estimate using daily wages, the number of workers, and the number of working days. Thus, it has a few critical issues that cause a downward bias in the estimate.
Figure B.6: Distribution of the THBS Heads’ Age
Notes: This figure shows the density of the THBS heads’ age calculated from the THBS daily panel dataset. The dashed line (red) indicates the average age of all the heads (33.2 years old). The dashed line (blue) shows the average age of the heads, excluding an elder head aged 67.6 (32.1 years old). The difference between the two average figures is statistically insignificant with $p=0.4087$.
Source: Created by the author using the THBS dataset.
The first issue is that the wages of factory workers are aggregated in a “15 years old or older” category in the ASCT (volume 17). Figure B.6 shows the THBS heads’ age distribution. The sample mean is $33.2$ years old, which is unchanged if I exclude a head aged more than $60$ (i.e., $32.1$ years old). This means that I have to refer to the average daily wage for the 30s to 40s for comparison. However, since many workers are in their 20s in the distribution of entire workers, the average wage reported in the ASCT shall be considerably lower than the average wage for factory workers around 30 years old. Thus, I need the wage profiles by age to correct for this downward bias.101101101 The ASCT (volume 18), which contains statistics for 1919, does not list such wages by age group but only average wages. As is clear, the same problem arises in this case.
Second, there is no consideration of ancillary pay such as bonuses and various allowances. As explained later, these wages shared a non-negligible proportion of the head’s earnings. Thus, estimation without considering ancillary wages would have a strong downward bias. In the following, I summarize the procedure for correcting these downward biases.
B.6.1 Daily Wage Distribution by Ages
Since the early wage surveys focused on average wages, official statistics in the 1910s do not include average wage statistics by age (bin). By the “Statistical Data Field Survey Law” (Law No. 52) in April 1922, average wages by age group were surveyed for the first time in the “Labor Statistics Field Survey” in October 1924 conducted by the Social Affairs Bureau of the Ministry of Home Affairs (Committee for the History of the Labor Movement 1964, p. 14). The result of this survey in Tokyo was published as the “Labor Statistics Table for Tokyo City and Suburbs” (hereafter, LST) by the Tokyo City Statistics Division in 1926.
(a) 1918
(b) 1919
Figure B.7: Average Daily Wage by Age Bins
Notes: This figure shows the average daily wage distribution by age bins based on the manufacturing census in 1924. Figures B.7a and B.7b show the distributions under 1918 and 1919 prices, respectively. The weighted average daily wage for a 30-39-year-old bin is shown in red. The weighted averages for all and 14+ bins are shown in gray and blue lines, respectively. The fitted fractional polynomial models are shown in the dashed lines. For the 1918 price data, the fitted model is $\hat{y}^{\textit{Daily Wage}}=-10.01+4.15a^{1/2}-0.33a^{2}$ with $R^{2}=0.99$. For the 1919 price data, the fitted model is $\hat{y}^{\textit{Daily Wage}}=-13.18+5.47a^{1/2}-0.43a^{2}$ with $R^{2}=0.99$.
Sources: Created by the author using the Tokyo City Statistics Division (1926a, pp. 16; 244-245). Data on the consumer price index for the cities are obtained from Ohkawa et al. (1967, p. 135).
Based on the LST, Figure B.7 shows the average daily wages of male factory workers in Tokyo City’s machinery and equipment manufacturing industry by age group. In the LST, there are 13 age bins, and the number of workers and average wages are listed for each bin. Figure B.7a shows the distribution of wages adjusted to 1918 prices using the consumer price index (CPI) for urban areas (Ohkawa et al. 1967, p. 135). The weighted average for ages $14$ and older is $2.58$ yen, which is about the same as the weighted average for daily wages of $2.56$ yen. The $2.58$ yen is between the average values for the 20-24 age bin and the 25-29 age bin. Thus, it likely to represent the daily wage of workers around 25. This is consistent with the fact that workers are concentrated in the 20s. However, the weighted average for the 30-39 age bin is $3.02$ yen, which deviates significantly from $2.58$ yen. The distribution of wages at 1919 prices shows a similar result (Figure B.7b). Using weighted averages for earnings comparisons leads to the erroneous conclusion that the sample households are upwardly biased.
To address this issue, I use the following steps to estimate the distribution of wages of the factory workers in Tokyo City in 1918-1919, the observation period of the THBS sample. Since the same estimation method is applied to both years, I use 1918 as an example below.
First, the distribution of daily wages by age group was projected from statistics obtained from the LST. Following Royston and Altman (1994), I consider a fractional polynomial of degree $m$ as follows:
$\phi_{m}(a;\bm{\varphi},\mathbf{p})=\varphi_{0}+\sum_{j=1}^{m}\varphi_{j}a^{(p_{j})},$ (6)
where $a$ ($>0$) denotes class value of age bin, $\bm{\varphi}$ is a vector of coefficients and $\mathbf{p}$ is a vector of powers, satisfying $p_{1}<p_{2}<...<p_{m}$. $a^{(p_{j})}$ indicates the Box-Tidwell transformation, taking $\text{ln}(a)$ if $p_{j}=0$ and $a^{p_{j}}$ if $p_{j}\neq 0$. As suggested in Royston and Altman (1994, p. 433), the models requiring a degree higher than two ($m>2$) are rare in practice. The wage by age bin distribution illustrated in Figure B.7 clearly shows a simple quadratic curve. Thus, I consider a fractional polynomial function with degree two as follows:
$\phi_{2}(a;\varphi_{0},\varphi_{1},\varphi_{2},p_{1},p_{2})=\varphi_{0}+\varphi_{1}a^{(p_{1})}+\varphi_{2}a^{(p_{2})}.$ (7)
To model the function, I consider a set of powers $\mathscr{P}=\{-2,-1,-0.5,0,0.5,1,2,3\}$ in the spirit of Royston and Altman (1994, p. 434).
The maximum likelihood estimation is used to search for the most suitable candidate of the coefficients for a given $m$. Since $m=2$, there are ${}_{8}\mathrm{H}_{2}$ ($=36$) repeated combinations for the candidates of the parameters. By adding the $8$ simple candidates under the nested models (i.e., the models with degree-1 fractional polynomial) to be conservative, I run $44$ regressions in total to find out the model with the highest maximized log-likelihood. The best power vector suggested is $(p_{1},p_{2})=(0.5,1)$.102102102The deviance, defined as the twice negative (maximized) log-likelihood, is calculated to be $-31.5$. Royston and Altman (1994) proposed the deviance difference, $D(m,\mathbf{p})-D(m,\mathbf{\tilde{p}})$, where $\mathbf{\tilde{p}}$ is the best power vector to compare the different models. This statistic has an asymptotic chi-squared distribution with $m$ degree of freedom. The estimated deviance difference with the simplest linear model (i.e., $m=1$ and $p_{1}=1$) is $54.6$ with $p$-value = $0.000$. Similarly, the difference with an another non-linear model with $m=1$ and $p_{1}=-2$ is $34.2$ with $p$-value = $0.000$. The projection model fitted using the best power vector is as follows:
$\hat{y}^{\textit{Daily Wage}}=-10.01+4.15a^{\frac{1}{2}}-0.33a^{1},\quad R^{2}=0.99.$ (8)
The dashed curve in Figure B.7 is the predicted distribution in equation 8. The fitting suggests that the predicted distribution works properly, especially below the 50-54 bin.
Next, I shift the predicted distribution based on the difference in the average daily wages in the ASCT and LST. The weighted average of the daily wage is $2.56$ yen for the LST, whereas that for the ASTC is $1.05$ yen, meaning that there is roughly $1.5$ yen difference.103103103For the LST, I use the number of factory workers in each age bin as weight. The ASCT reports the average daily wage for four different categories of factories in each manufacturing sector: the factory scale (10+ or below ten workers) and engines (use or not). I calculated the weighted average of the daily wage using the number of factory workers in each age bin (12–14 and 15+) and in each category as weights. Thus, I shift the functional polynomial function by this amount of difference to yield the proxy distribution for 1918.104104104This operation, therefore, assumes that the wage-by-age distributions in 1918 and 1924 are reasonably similar. In the LST (i.e., the statistic on which the projected distribution is based), the weighted average for those aged 14 and older was 3.25 yen. In comparison, the overall weighted average was 3.23 yen, a difference of only 0.02 yen. A similar trend can be observed in the statistics obtained from the ASCT: the weighted average for those aged 15 and older is 1.14, while the overall average is 1.13, a difference of only 0.01 yen. At least, in the light of the first moment, this confirms that the wage distribution in the ASCT is similar to that in the LST. Figure B.8a illustrates this operation: dotted curve is a fractional polynomial shifted from the original fractional polynomial shown as a short dashed curve.
The average daily wage for 12–14-year-old bin is reported to be $0.42$ yen in the ASCT, whereas that for 12–13-year-old bin in the LST is $0.60$ yen. This difference (i.e., $0.18$ yen) is smaller than the difference in the weighted average for all age bins (i.e., $1.5$ yen). This implies that the daily wage by age distribution in the ASCT could be smoother (i.e., flatter) than that of the LST. Thus, the shifted fractional polynomial would provide stronger shrinkage for those bins of younger age. In other words, the shifted curve shall get closer to the actual (unobservable) wage distribution as age increases.
The long dashed curve in Figure B.8a examines this relationship by fitting another fractional polynomial model using the 12-14-year old bin figure of the ASCT and the predicted wages for 25–65 from the shifted fractional polynomial.105105105 The fitted model is as follows: $\hat{y}^{\textit{Daily Wage}}=-1.19+0.14a^{1}-0.002a^{2},\quad R^{2}=0.98.$ The deviance is reported to be $-149$. I used the 25–65 age range for the prediction based on the age bins used in the LST. Let this curve predict wage distribution for the true distribution. The average daily wage for the 30–40-year-old bin in the shifted fractional polynomial is $1.55$ yen, which is similar to that for the same-age bin in the predicted wage distribution (i.e., $1.52$ yen). However, the wage gaps for the younger ages show more significant differences. As long as our interest is in the late 20s to the early 40s, therefore, it is plausible to employ the predicted wages for the wage comparison in Section 3.
Figure B.8b illustrates the result of the same analysis from the wage data of 1919 price, reflecting the increasing trend in price between 1918 and 1919. The average daily wage for a 30–40-year-old bin in the shifted fractional polynomial is $2.46$ yen. Again, this is similar to that for the predicted wage distribution (i.e., $2.42$ yen).
(a) 1918
(b) 1919
Figure B.8: Daily Wage by Ages Distribution:
Shifted Fractional Polynomial
Notes: The short dashed curve illustrates the fractional polynomial (FP) in Figure B.7. The dotted curve indicates the FP shifted by the difference in the average daily wage between the ASCT and LST. The ASCT documents the wage statistics for four different categories for each manufacturing sector: the scale of the factory (10+ or below ten workers) and engines (use or not). I calculated the weighted average of the daily wage using the number of factory workers in each age bin and the category as weights. The long-dashed curve shows the FP refitted using the average daily wage for 12-14-year-old bin ($0.42$ yen) measured in the ASCT and the daily wages for 25-65-year-old suggested in the shifted FP. The average daily wage for a 12–14-year-old bin is only available for the ASCT of 1918. Therefore, this figure for the 1919 price is calculated using the ratio between the average daily wage ($1.05$ yen) and that for 12–14 years old bin ($0.42$) in 1918 as $1.80\times(0.42/1.05)$, where $1.80$ is the average daily wage measured in the ASCT of 1919. The red dashed line is the average daily wage for 30–40-year-old in the shifted FP. The blue dashed line is the average daily wage for 30–40-year-old in the refitted FP.
Sources: Created by the author using the Tokyo City Office (1921, pp. 726–757); Tokyo City Statistics Division (1926a, pp. 16; 244-245).
B.6.2 Ancillary Wage
Figure B.9: Log-Relationship between the Number of Workers and Share of Ancillary Wage: Original Data from the Factory Survey by Kitazawa (1924)
Notes: This scatter plot shows the relationship between the log-transformed number of factory workers and the log-transformed share of ancillary wage. The raw data are listed in Table B.3. The dashed line indicates the linear fit using the log-transformed variables: the fitted regression is $\hat{y}=0.747+0.428x$ with $R^{2}=0.82$. The predictions are materially similar if I use the fractional polynomial using the raw data on the number of factory workers and share of ancillary wage in Table B.3. In the case of fitting based on the raw data, the fitted fractional polynomial model is $\hat{y}^{\textit{Daily Wage}}=53.6-399.3a^{-1/2}+(2.84\times 10^{-10})a^{3}$ with $R^{2}=0.87$. I prefer to use linear regression herein because it can provide more precise fitted values for factories with a small number of workers.
Source: Kitazawa (1924, pp. 20; 58–62).
The daily wage estimated in Section B.6.1 is a fixed salary and thus, does not include additional salaries such as bonuses and allowances (hereafter, wages other than regular salaries are collectively referred to as ancillary wages). However, there are no systematic statistics on these ancillary wages throughout the prewar period (Committee for the History of the Labor Movement 1964, pp. 14–15). Moreover, since the Tsukishima Survey targets skilled worker households, information on the ancillary wages paid to skilled workers is necessary. Fortunately, I found a survey report named “Wage Survey Report” edited by Kitazawa (1924). This report includes the result of a survey on the wages for machine finishers (fitters) and turners at six machine factories in Tokyo from May–October 1921. Thus, this is particularly useful material to know the wage structure of skilled factory workers.
According to Kitazawa (1924), there are two methods of paying wages to workers in the machine industry: one is to use hourly wages to calculate fixed salaries, and the other is to combine hourly wages with piece-rate wages, with the former being the most common (p. 33). Despite this, wages paid in both methods are usually divided into fixed salaries and other ancillary wages, and there is not much difference in essence (p. 39). Fixed wages are salaries added from a fixed hourly wage rate, while ancillary wages include bonuses and overtime pay, night shift pay, and labor premiums (profit sharing). In the case of a combination of hourly and piece-rate pay systems, the accessory wage includes a share of the piece-rate profit.
Panel A in Table B.3 summarizes the survey results. The surveyed factories included 86, 157, 376, 623, 928, and 2,837 workers, ranging from small to large plants. The percentage of skilled workers (finishers and turners) in the surveyed factories ranged from $27$% to $77$%, and the portion of monthly ancillary wages was 12–53%. Figure B.9 is a linear regression fit using log-transformed data on the percentage of ancillary wages and the number of workers. There is a positive correlation between the factory size and ancillary wages. This is consistent with the historical fact that larger factories have more extensive benefits than smaller ones (Odaka 1999).
Next, I estimate the percentage of ancillary wages at the Tsukishima factories. One of the most frequently used statistics on factories in Tokyo is the Handbook of Factories (kōjyō tsūran), which covers all the factories with ten or more workers as of January 1, 1920. This handbook is the most comprehensive source for ascertaining the number of factories and their sizes around 1919. This document does not cover the number of tiny working places with less than nine workers. Importantly, however, the purpose of Panel C in Table 2 is to compare the average monthly wage of the skilled factory workers in Tsukishima and the THBS heads’ average monthly earnings. Since the THBS heads did not work in the tiny workplaces, the censoring at the left tail of the factory size distribution is unlikely to influence this comparison.
Panel B of Table B.3 summarizes the distributions of the number of factories, the total number of workers, and the estimated average share of ancillary wage in Tokyo City (Column 1) and Tsukishima (Column 2). The average ancillary wage at the Tsukishima factories is 12%, which is slightly larger than the value for the entire city of Tokyo (9%). The percentage shares of the number of factory workers by each bin are listed in Columns (1) and (2) of Panel C in Table B.3. As shown, Tsukishima factories tended to have larger factories than those of the entire city, which is reflected in the slightly larger estimate in the average ancillary wage for the Tsukishima factories. In Column (3) of Table B.3, the number of THBS heads whose occupation, factory name, and factory size are available are listed. While the percentage shares of the heads in each bin show a similar distribution to that of the Tsukishima factories, the share of middle (large) sized factories is larger (smaller) than the population. If interpreted literally, this tendency implies that the (unobservable) true average ancillary wage for the THBS heads may be slightly lower than the population mean. Note again that, however, the purpose of Panel C in Table 2 is to compare the average monthly wage of the skilled factory workers employed in the Tsukishima factories to the THBS heads’ average monthly earnings. Therefore, the purpose herein is to estimate the average ancillary wage for the target population, say the Tsukishima factories but not to calculate the average ancillary wage for the THBS sample itself.
Table B.3: Estimating the Average Share of Monthly Ancillary Wage
Panel A: Original Data on the Share of Ancillary Wage
Document: Report of Factory Survey by Kitazawa (1924)
Survey subject: Six Machinery Factories in Tokyo City
Survey month and year: May to October 1921
Number of Share of Share of Share of
factory skilled fixed ancillary
Factory workers workers (%) wage (%) wage (%)
A 157 44 83 17
B 86 53 88 12
C 376 27 57 43
D 623 77 68 32
E 2837 NA 47 53
F 928 41 60 40
Panel B: Estimating the Average Share of Ancillary Wage
Document: Handbook of Factories (1921 edition)
Survey subject: Machinery Factories with 10+ workers in Tokyo City
Survey month and year: January 1920
Machinery Factories
(1) Tokyo City (2) Tsukishima
Number Number Average share Number Number Average share
of factories of factory of ancillary of factories of factory of ancillary
Factory size workers wage (%) workers wage (%)
10–99 workers 760 18,760 8 39 1,252 9
100–199 workers 52 7,155 17 5 713 18
200+ workers 34 21,764 31 3 3,851 38
Total 846 47,679 9 47 5,816 12
Panel C: Comparisons of the Distributions of Factory Size
Machinery Factories
(1) Tokyo City (2) Tsukishima (3) THBS
Number % share Number % share Number % share
of factory of factory of THBS
Factory size workers workers heads
10–99 workers 18,760 39 1,252 22 6 21
100–199 workers 7,155 15 713 12 7 25
200+ workers 21,764 46 3,851 66 15 54
Total 47,679 100 5,816 100 28 100
Notes:
Panel A: This panel shows the size of factories and the share of fixed and ancillary monthly wages in six factories surveyed in Kitazawa (1924). The number of factory workers surveyed in factories A–F are 14, 23, 9, 11, 11, and 16, respectively. The share of these surveyed skilled workers to all skilled workers is 20.3 in Factory A, 50.0 in Factory B, 5.3 in Factory C, 2.3 in Factory D, and 4.2 in Factory F, respectively. This figure for factory E is unavailable because the number of skilled workers is unknown (Kitazawa 1924, p. 21).
Panel B: Subcolumns 1–2 and 4–5 summarize the number of factories and total number of workers in Tokyo City and Tsukishima, respectively (Handbook of Factories 1921). Subcolumns 3 and 6 show the average share of ancillary monthly wage predicted in Figure B.9 for Tokyo City and Tsukishima, respectively.
Panel C: Columns 1–3 summarize the number of factory workers and its share in percentage points in Tokyo City, Tsukishima, THBS sample, respectively. The statistics in Columns 1 and 2 (same figures used in Panel B) are from the Handbook of Factories (1921). In Column 3, the THBS heads who worked in the machinery factories are considered. There are 28 heads whose factory names and size are available. For a blacksmith included in these heads, I assume the number of workers to be 10, based on the report (Department of Health, Ministry of the Interior 1923a, p. 412).
Sources: Panel A: Kitazawa (1924, pp. 20; 58–62). Panel B: The number of factory workers is taken from the Ministry of Agriculture and Commerce (1921, pp. 613-629; 674). The average shares of ancillary wage are calculated using Kitazawa (1924, pp. 20; 58–62). See also Figure B.9. Panel C: Data on the number of factory workers and factories are from the Ministry of Agriculture and Commerce (1921, pp. 613-629; 674). Data on the number of heads are from the THBS sample.
B.6.3 Estimating Monthly Income of Adult Male Factory Workers
Table B.4: Estimating the Monthly Income of Adult Male
Factory Workers in Machinery Sector
Weight for the Ancillary Wage
(1) Tsukishima (2) Tokyo City
$\vartheta=0.12$ $\vartheta=0.09$
Panel A: Components in equation 14: $w^{Daily}_{j}$, $l_{j}$, and $n_{m}$
Average daily wage ($w^{Daily}_{j}$)
$w^{Daily}_{1918}$ 1.52 1.52
$w^{Daily}_{1919}$ 2.42 2.42
Average annual working days ($l_{j}$)
$l_{1918}$ 313 313
$l_{1919}$ 322 322
Frequency weights ($n_{m}$)
$n_{1}$ 5 5
$n_{2}$ 11 11
$n_{3}$ 19 19
$n_{4}$ 22 22
$n_{5}$ 21 21
$n_{6}$ 18 18
$n_{7}$ 11 11
$n_{8}$ 8 8
$n_{9}$ 5 5
$n_{10}$ 4 4
Panel B: Estimated monthly earning: $w^{\textit{Total}}$
$w^{\textit{Total}}$ 56.0 54.5
Notes:
1. Average daily wage ($w^{Daily}_{j}$): The average daily wage of male factory workers in the skilled age range (aged 30-40) in the machinery sector is calculated using two official reports of the manufacturing censuses documented in the ASCT and LST. Details in the calculation steps are summarized in Online Appendix B.6.1.
2. Average annual working days ($l_{j}$): The ASCT documents the number of factories and average annual working days for each machinery sub-sector in 2 factory-size bins. I used the number of factories in each machinery sub-sector and factory size bin as the weights to calculate the weighted average of the annual working days in the entire machinery sector.
3. Weight for the ancillary wage ($\vartheta$): The average share of monthly ancillary wage in the machinery sector is calculated based on the factory survey report in Kitazawa (1924). The calculation steps are described in Online Appendix B.6.2. Column 1 uses the share based on the factory size distribution in Tsukishima, whereas column 2 uses that in the entire Tokyo City.
4. Frequency weights ($n_{m}$) are based on the number of cross-sectional observations between January and October 1919 in the adjusted monthly panel dataset.
5. $w^{\textit{Total}}$ is calculated using equation 14.
Sources: Data used to calculate the average daily wage are from the Tokyo City Office (1921, pp. 726-757); Tokyo City Office (1922, pp. 732-733); Tokyo City Statistics Division (1926a, pp. 16; 244-245). Data used to calculate the average annual working days are from the Tokyo City Office (1921, pp. 726-741) and Tokyo City Office (1922, pp. 716-725). Data used to calculate the average share of monthly ancillary wage are from Kitazawa (1924, pp. 20; 58-62) and the Ministry of Agriculture and Commerce (1921, pp. 613-629; 674).
Following the institution described in Online Appendix B.6.2, the monthly wage of the male workers in the skilled age range ($w^{\textit{Total}}$) can be defined as follows:
$w^{\textit{Total}}=w^{\textit{Regular}}+w^{\textit{Ancillary}},$ (9)
where $w^{\textit{Regular}}>0$ is the monthly regular wage and $w^{\textit{Ancillary}}\geq 0$ indicates the monthly ancillary wage. Online Appendix B.6.2 shows that $w^{\textit{Ancillary}}$ can be calculated as the product of the monthly regular wage and weight for the ancillary wage, $\vartheta\in(0,1)$.106106106Strictly, this weight is not required to be in this range. However, all the cases reported in Kitazawa (1924) take the weights within this range. The monthly wage equation is then specified as follows:
$w^{\textit{Total}}=(1+\vartheta)w^{\textit{Regular}}.$ (10)
To calculate equation 10, I have to conduct a few final tunings in my wage data. First, I fixed the deviations in the timing of the manufacturing censuses and the THBS observations. Let $w^{\textit{Regular}}_{j}$ be the monthly regular wage calculated using wage statistics measured in year $j\in\{1918,1919\}$ as follows:
$w^{\textit{Regular}}_{j}=w^{\textit{Daily}}_{j}l_{j}/12,$ (11)
where $w^{\textit{Daily}}_{j}$ is the average daily wage of male factory workers in the skilled age range and $l_{j}$ is the average annual working days (Online Appendix B.6.1). The THBS were conducted from November 1918 to January 1920, and my THBS sample was observed from December 1918 to December 1919 (Section 3). However, $w^{\textit{Regular}}_{j}$ is calculated using the wage statistics measured in the manufacturing census conducted in December of 1918 and 1919. This means that if I use $w^{\textit{Regular}}_{1919}$ alone for equation 10, it shall provide upward-biased wage due to the deviation between the timing of the THBS observations and of the manufacturing census. To cope with this issue, I use the weighted average between $w^{\textit{Regular}}_{1918}$ and $w^{\textit{Regular}}_{1919}$ to yield monthly wage series as follows:
$w^{\textit{Regular}}_{m}=\{(12-m)w^{\textit{Regular}}_{1918}+mw^{\textit{Regular}}_{1919}\}/12,$ (12)
where $m\in\{0,1,2,...,12\}$ corresponds to the year-month cells from December 1918 to December 1919. This way of weighting assumes that the average daily wage of machinery workers had increased during this period. This assumption is plausible given the historical fact that the average wage in the manufacturing industry indeed showed a secular increasing trend around 1920 in Tokyo City (Ohkawa et al. 1967, p. 255).
Secondly, I use the number of observations in each year-month cell as the frequency weight. Let $n_{m}$ be the number of households measured in year-month cell $m$. Then, the average monthly wage can be defined as:
$w^{\textit{Regular}}=\frac{\sum_{m}n_{m}w^{\textit{Regular}}_{m}}{\sum_{m}n_{m}},$ (13)
where $m\in\{\text{December}~{}1918,\text{January}~{}1919,...,\text{December}~{}1919\}$ is the year-month cell and $n_{m}$ is the frequency weight for year-month cell $m$.
Substitute equation 1112, and 13 into equation 10 to yield the final specification:
$w^{\textit{Total}}=(1+\vartheta)\frac{\sum_{m}n_{m}\{(12-m)w^{\textit{Daily}}_{1918}l_{1918}+mw^{\textit{Daily}}_{1919}l_{1919}\}}{12^{2}\sum_{m}n_{m}}.$ (14)
Panel A in Table B.4 summarizes all the components in equation 14. In column 1 of Panel B in Table B.4, I calculate the average monthly wage under the share of the ancillary wage for the factories in Tsukishima ($\vartheta$). The estimate is $56$ yen, which is similar to the average and median of monthly earnings in the THBS heads ($59$ and $56$ yen, respectively; see Panel C of Table 2). Therefore, I can conclude that the deviation from the population mean ($56$ yen) is not too large to infer the mean household behaviors of the skilled workers in the machinery factories in Tsukishima. Given that the average daily wage of the male factory workers in the skilled age range is $2.5$ yen (Figure B.8), $3$ yen difference from the average monthly income is roughly one day’s earnings. This is consistent with the findings in my descriptive analyses that show the similarities in the distribution of family structure, average household size, and share of ancillary wage between the THBS households and Tsukishima households (Section 3; Online Appendix B.6.2).
Column 2 uses the share of the ancillary wage for the entire Tokyo City ($\vartheta=0.09$) to calculate the average monthly wage. The estimate is $54.5$ yen, which is slightly smaller than that in Column 1, reflecting the difference in the weights for the ancillary wages. If I see this figure as an unbiased estimate for the entire city, the difference in the average factory size between the entire Tokyo City and Tsukishima may be reflected in this deviation (Panel B in Table B.3).
B.7 Measurement Errors in the Unadjusted Panel Dataset
(a) Unadjusted Semi-monthly Panel Dataset: Income and expenditure
(b) Unadjusted Semi-monthly Panel Dataset: Net income
(c) Shifted Semi-monthly Panel Dataset: Income and expenditure
(d) Shifted Semi-monthly Panel Dataset: Net income
(e) Adjusted Semi-monthly Panel Dataset: Income and expenditure
(f) Adjusted Semi-monthly Panel Dataset: Net income
Figure B.10: Comparisons of Average Semi-monthly Income and Expenditure
from Different Aggregations
Note: Figures B.10aB.10cB.10e show the average semi-monthly income and expenditure calculated from unadjusted, shifted, and adjusted panel datasets, respectively. Unadjusted series defines the days from the 1st to the 15th of each month as the first half, whereas the days from the 16th to the end of each month as the second half. The shifted series defines the days from the 2nd to the 16th of each month as the first half, whereas the days from the 17th to the 1st of the following month are the second half. Adjusted series is the baseline definition in the main text, which uses the days from 12th to 26th in each month as the “first-half” and those from 27th to 11th in the next month as the “second-half”.
As explained in Section 3, the raw series contains considerable measurement errors due to the timing of payday. The shifted series can deal with the measurement errors from the timing of payday but can not adjust the difference in the payday timing and expenses. In other words, the shifted series ignores the households’ decision on the consumption schedule in each semi-month. The adjusted series can deal with measurement errors and take the timing of the household expenses into account.
Figures B.10bB.10dB.10f illustrate the average semi-monthly income minus expenditure calculated from unadjusted, shifted, and adjusted panel datasets, respectively.
Source: Created by the author using the THBS sample.
Figure B.10 explains how different ways of aggregation influence the measured time-series in income and expenditure. Figures B.10aB.10cB.10e show the average semi-monthly income and expenditure calculated from raw, shifted, and adjusted panel datasets, respectively.
The unadjusted (calender-month) series defines the days from the 1st to the 15th of each month as the first half and the days from the 16th to the end of each month as the second half. This series shall contain substantial measurement errors due to the mismatches in the income and expenditure data explained in Section 3.3. The time-series plots of average income in Figure B.10a have substantially large variances and fluctuate over time. This is because some semi-monthly income is assigned to the wrong semi-month cells. In these wrong cells, the measured incomes shall take unnaturally greater values than usual. Moreover, the time-series plots of average expenditure do not chase those of the income. For example, the average expenditure in the second half of June greatly differs from the average income in the same cell. Consequently, the average net income (i.e., income minus expenditure) in each semi-month is relatively flat and takes a few odd cells with large positive values (Figure B.10b).
The shifted series defines the days from the 2nd to the 16th of each month as the first half and the days from the 17th to the 1st of the following month as the second half. Although the shifted series can partly deal with the measurement errors in income, it does not adjust the difference in the payday timing and expenses. Figure B.10c indeed gets a much smaller variance in the average incomes over time because all the paydays are now fixed to assign the suitable semi-month cells. However, the expenditure series still does not chase the income series well because the payday is placed on the latter edge in each semi-month cell. In other words, the average expenditure in each semi-month reflects a large part of the consumption based on the income of the previous semi-month cell. Since the average incomes take a better series than Figure B.10a, the average net income offers better time-series plots (Figure B.10d). However, it still has a few cells with large negative and large values due to the deviations of expenditure from income.
The adjusted series uses the days from the 12th to 26th of each month as the “first half” and those from the 27th to 11th of the following month as the “second half”. By shifting the entire semi-month cells, this adjusted series can deal with the measurement errors from the payday shifts and fix the miss-matching in the timing of income and expenditure. In other words, each payday is assigned to a suitable semi-month cell, and each semi-month cell reflects the household’s decision on consumption given the latest income. Figure B.10e shows similar time-series plots of average income in Figure B.10c, in which has a corrected semi-monthly average income series only. Notably, the expenditure series are now chasing the income series in Figure B.10e because each semi-month cell’s expenditure is based on the income in the same cell. Accordingly, Figure B.10f now shows a smoother series in the net income. It shows a clear increasing trend in the net income and has a smaller variance over time.
These illustrations support the evidence that the adjusted semi-monthly datasets can offer cleaner within variations in income and expenditure because they can fix the measurement errors and can reflect the households’ consumption behaviors based on their latest incomes measured in the same semi-month cell.
B.8 Distribution of the Idiosyncratic Shocks
Figure B.11: Distribution of the Idiosyncratic Shocks
Notes: This figure shows the distribution of idiosyncratic shocks in the semi-monthly THBS panel dataset. The idiosyncratic shocks are estimated as the residuals from the regression of the first difference in income on the first difference in aggregate income. The red line indicates the mean of the idiosyncratic shocks ($-0.235$).
Source: Created by the author using the THBS sample.
Figure B.11 illustrates the distribution of the idiosyncratic shocks predicted using the semi-monthly panel dataset.107107107I used the semi-monthly panel dataset to overview the entire distribution of the idiosyncratic shocks rather than the daily panel datasets. This is because the denser panels include a set of consecutive household-date cells with no income. These cells can not be used to predict the idiosyncratic shocks as the residuals economically meaningfully, and thus it is practically impossible to illustrate the idiosyncratic shocks based on the daily dataset. The idiosyncratic shocks are calculated as the residuals from the regression in the first difference in semi-monthly income on the first difference in aggregate semi-monthly income. Let $n_{t}$ be the number of households in semi-month $t$ in the year 1919. For household $i$ in semi-month $t$ cell, the residual is then defined as follows:
$\Delta\hat{\upsilon}_{i,t}=\Delta\textit{Income}_{i,t}-\hat{\alpha}\Delta\overline{\textit{Income}}_{.,t},$ (15)
where $\overline{\textit{Income}}_{.,t}=\sum_{i=1}^{n_{t}}\frac{\textit{Income}_{i,t}}{n_{t}}$ and $\hat{\alpha}$ is the estimated coefficient. Figure B.11 shows that the residuals satisfy the zero-average property. However, this figure also indicates that the residuals have rich variations, supporting that the THBS households had experienced both favorable and adverse idiosyncratic shocks during the survey periods.
B.9 Unit Root and Autocorrelations
Table B.5 presents the results of the Fisher-type panel unit root tests for the total expenditure in the semi-monthly (Column 1) and adjusted monthly (Column 2) panel datasets. Both columns show that the null hypothesis of unit roots in all the panels is rejected at the conventional level. The results of the unit root tests for consumption subcategories and variables used to test the risk-coping mechanisms are not reported in this table because those are materially similar results. In Section 4, I used the cluster-robust variance-covariance estimator in the regressions to allow arbitrary autocorrelations (Arellano 1987). Despite this, I confirmed that the null hypothesis of no serial correlation is not rejected at the conventional level in Wooldridge’s autocorrelation test for almost of the consumption and net income variables (Wooldridge 2002).
Table B.5: Results of the Unit Root Tests
Test statistics (1) Semi-Monthly (2) Adjusted Monthly
$P$-statistic $p$-value 0.0000 0.0001
$Z$-statistic $p$-value 0.0000 0.0000
$L^{*}$-statistic $p$-value 0.0000 0.0000
$P_{m}$-statistic $p$-value 0.0000 0.0000
The results of the Fisher-type panel unit root tests based on augmented Dickey-Fuller (ADF) tests are reported in this table. The null hypothesis is that all the panels contain unit roots, whereas the alternative hypothesis is that at least one panel is stationary. In all the specifications, the process under the null hypothesis is assumed to be a random walk with drift. The demeaned data are used to address the effect of cross-sectional dependence. The number of lagged differences in the ADF regression equation reported is set as one because including two or more two lags reduces a large part of the panel units available for computing the test statistics due to the unbalancedness of the panels. I confirmed that including two lags in the semi-monthly panels does not change the results. Including two lags in the adjusted monthly panels and three lags in the semi-monthly panels is impossible. See Choi (2001) for the details of the tests. These tests need both the first-differenced dependent variable and the lagged dependent variable. Therefore, the units with less than four time-series observations are excluded in the regressions.
Source: Calculated by the author using the THBS sample.
B.10 Risk Preferences
The potential influences of heterogeneous risk preference on standard risk-sharing regressions are discussed in a few works (Schulhofer-Wohl 2011; Mozzoco and Saini 2012). In my empirical setting, the heterogeneity in the risk preference is assumed to be removed by using the household fixed-effect. When the risk preferences are difference over households, say $\sigma=\sigma_{i}$, equation 26 becomes:
$\log c_{i,t}=\log c_{t}^{a}+\frac{1}{\sigma_{i}}(\log\omega_{i}-\omega^{a})+(\theta_{i,t}-\theta_{t}^{a}).$ (16)
The first differencing removes the second term with the coefficient of relative risk aversion in the right-hand side of this equation, even though it varies over the households.
Schulhofer-Wohl (2011) shows that the heterogeneous risk preference may lead to bias in the estimation of the income elasticity under the standard risk-sharing regressions and proposes a parametric test for that heterogeneity. This test is valid under a particular utility function. Specifically, the bias could be raised under the case that the coefficient of relative risk aversion interacts with the aggregate shock in the first-order condition for consumption because the interaction means that the aggregate shock can not be removed using the time-fixed effect in the risk-sharing regressions.108108108Schulhofer-Wohl (2011, pp. 928–929) considers the utility maximization problem under the following weighted sum of discounted expected utilities: $\sum_{i}\dot{\alpha}_{i}E\left[\sum_{t}\dot{\beta}^{t}\frac{[c_{i,t}(s_{t})]^{1-\dot{\gamma}_{i}}}{1-\dot{\gamma}_{i}}\right],$ (17) where $\dot{\alpha}_{i}$ is Parate weight and $\dot{\gamma}_{i}$ is the coefficient of relative risk aversion for household $i$. The log-transformed first-order condition becomes: $\text{log}c_{i,t}=\frac{\text{log}\dot{\alpha}_{i}}{\dot{\gamma}_{i}}-\frac{\text{log}\dot{\lambda}_{t}}{\dot{\gamma}_{i}},$ (18) where $\dot{\lambda}_{t}$ indicates the Lagrangian multiplier. This result implies that the effects of the aggregate shocks ($\dot{\lambda}_{t}$) depend on the household’s attitudes toward risk and that the first-differencing does not remove the second term in equation 18. As explained, this setting differs from the one used in this current study. Mozzoco and Saini (2012) consider a more general case for testing the homogeneity of risk preferences. Their test requires a panel dataset with a very long time dimension that can not be applicable to my THBS dataset.109109109The data used in their study covers the household-level monthly panel dataset surveyed in rural India between 1975 and 1985. The households with fewer than 80 data points are not included in the homogeneity tests (Mozzoco and Saini 2012, p. 453). See also Shrinivas and Fafchamps (2018) and Mozzoco and Saini (2018) for the subsequent discussions.
An alternative way to test potential heterogeneity in the risk preference is to use direct measures on the risk preferences of workers. Schulhofer-Wohl (2011) uses the results of experimental risk-tolerance questions on the workers’ risk preferences in the Health and Retirement Study to test whether the correlations between earnings and aggregate shocks differ across high- and low-risk-tolerance workers.110110110This question is asking if the respondents would be willing to take a risk to get a new job with a higher income, even if there’s a 50-50 chance of a reduction in income by a third (Schulhofer-Wohl 2011, P. 931). They found that the high-risk-tolerance workers have more procyclical incomes than the low-risk-tolerance workers.
Since the THBS has no experimental questions, it is impossible to apply the same test. Given this limitation that the direct test using experimental risk-tolerance question is unavailable, I compare the correlations between the heads’ income and aggregate consumption over sub-groups to get insights into the potential difference in the heads’ attitudes toward risks. For household $i$ and semi-month $t$, the specification is characterized as follows:
$\log\tilde{y}_{i,t}=\xi_{1}\log c_{t}^{a}+\xi_{2}d_{i}\log c_{t}^{a}+ox_{i,t}+\iota_{i}+\varepsilon_{i,t},$ (19)
where $\tilde{y}_{i,t}$ is head’s earnings, $c_{t}^{a}$ is aggregate consumption, $x_{i,t}$ is the household size, $\iota_{i}$ is the household fixed-effect, and $\varepsilon_{i,t}$ is a random error term. The aggregate consumption is calculated as the cross-sectional average of the total expenditure among the THBS sample, given that the population mean of consumption expenditures among all the skilled factory workers’ households in Tsukishima is unavailable.
$d_{i}$ is an indicator variable that takes one if household $i$ belongs to a specific group that might have had different preferences on their jobs than the other heads in the sample. Firstly, I consider the case that the heads who worked in the large-scale factory had different risk preferences. Although there are no specific anecdotes behind this, the workers in the large-scale factory might have had a lower risk tolerance than those in the small- and medium-scale enterprises. 13 heads who worked in the Ishikawajima shipyard and governmental factory are classified into this group. Secondly, I assume that the heads who worked in the smithing and the non-machinery factories had different risk preferences. As described in Online Appendix A.5, skilled workers in the smithing sector could be small business owners.111111111Workers in the canning factories might have had a different preference on the risk preferences (Online Appendix A.6). However, my THBS sample does not include the workers in canning factories. The workers in non-machinery factories might also have different job preferences than the machinery factory workers. This group has only three heads: one worked in a smithing factory, and two worked in food and paper-making factories.
Table B.6 presents the results. In Column (1), the estimated coefficient on the interaction term is close to zero and statistically insignificant. Thus, the sensitivities to aggregate consumption are similar across different scales of the factories. Column (2) further shows that the estimated coefficient on the interaction term is not statistically significant under an alternative proxy of risk tolerance. It suggests that the heads worked in the smithing and non-machinery factories show similar responses to the aggregate shock. Since the aggregate consumption is the cross-sectional sample average, the aggregate variable may contain noise in the period with the small number of cross-sectional observations. To deal with this potential issue, I exclude October and November 1919 in Columns (3) and (4). As shown, the estimated coefficients on the interaction terms are largely unchanged in both columns.
This finding is consistent with the fact that these heads have similar characteristics in nature because the THBS focused on skilled factory workers in a representative manufacturing area, Tsukishima. Therefore, it is plausible to see that the THBS households had more or less similar attitudes toward the risks.
Table B.6: Results for Testing Heterogeneous Risk Preferences
DV: ln(head’s earnings)
Analytical Period
(1) Full (2) Full Jan. to Sept. Jan. to Sept.
Aggregate Consumption 0.277* 0.273 0.361** 0.362*
[0.158] [0.198] [0.164] [0.198]
Aggregate Consumption -0.041 -0.061
$\times$ Large Scale Factory [0.453] [0.415]
Aggregate Consumption -0.089 -0.178
$\times$ Smithing/Non-machinery [0.215] [0.215]
Observations 261 261 239 239
***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors in brackets are clustered at the household level.
Notes: This table shows the results of equation 19 for the semi-monthly panel dataset. ‘Aggregate Consumption’ indicates the cross-sectional sample average of the total expenditures. ‘Large Scale Factory’ indicates an indicator variable that takes one if the head worked in the Ishikawajima Shipyard or governmental factory. ‘Smithing/Non-machinery’ indicates an indicator variable that takes one if the head worked in the smithing, food, or paper-making factories. All the regressions include the family size variable and household fixed-effect.
Columns 1–2 show the results for the entire sample period. Columns 2–3 show the results for the period between January and September 1919 to remove year-month cells with the smaller (cross-sectional) number of households (Online Appendix Figure B.4).
Appendix Appendix C Empirical Analysis Appendix
C.1 Conceptual Framework
Consider the economy where household $i=1,...,N$ receives an uncertain income at time $t$ depending on the state of the world $s_{t}\in S_{t}$ as $y_{i,t}(s_{t})$. For infinite time horizon, the state $s_{t}$ realizes with probability $\pi(s_{t})\in[0,1]$. The weighted sum of the expected lifetime utility of $N$ households in the economy is expressed as:
$\sum_{i=1}^{N}\omega_{i}\sum_{t=0}^{\infty}\rho^{t}\sum_{s_{t}\in S_{t}}\pi(s_{t})u(c_{i,t}(s_{t}),\theta_{i,t}(s_{t})),$ (20)
where $\omega_{i}$ is the social planner’s weight representing the reciprocal of the marginal utility of each household ($10<\omega_{i}<$ 1 and $\sum_{i=1}^{N}\omega_{i}=1$), $10<\rho^{t}<$ 1 is the discount factor, $c_{i,t}(s_{t})$ is consumption, and $\theta_{i,t}(s_{t})$ is a preference shock. For all states $s_{t}$, the aggregate resource constraint is given by:
$\sum_{i=1}^{N}c_{i,t}(s_{t})=\sum_{i=1}^{N}e_{i,t}(s_{t}),$ (21)
where $e_{i,t}(s_{t})$ is household $i$’s exogenous endowment at time $t$.
Following Mace (1991), I assume that the aggregate idiosyncratic shocks approach zero as $N$ gets large. Consider the situation that the household’s endowment can be decomposed into the permanent component ($e_{i,t}^{P}$), shock linked to macroeconomic trends ($\varrho_{i,t}^{a}(s_{t})$), and idiosyncratic shock ($\varsigma_{i,t}(s_{t})$) as:
$e_{i,t}(s_{t})=e_{i,t}^{P}+\varrho_{i,t}^{a}(s_{t})+\varsigma_{i,t}(s_{t}).$ (22)
[…]
Subject to equation(21), the social planner maximizes equation(20) by choosing an allocation of consumption across households. The first-order condition at time $t$ is derived as follows:
$\omega_{i}\rho^{t}\pi(s_{t})u_{c}(c_{i,t}(s_{t}),\theta_{i,t}(s_{t}))=\lambda(s_{t}),$ (23)
where $\lambda$ is the Lagrange multiplier for the single lifetime resource constraint. To derive my empirical specification, I assume a constant relative risk aversion preference as:
$u(c_{i,t}(s_{t}),\theta_{i,t}(s_{t}))=\frac{c_{i,t}^{1-\sigma}(s_{t})}{1-\sigma}\exp(\sigma\theta_{i,t}(s_{t})),$ (24)
where $\sigma>0$ is the coefficient of relative risk aversion. Below, I consider a particular history of the state of the world so that the subscript $t$ can represent the dependence on the state (Hayashi et al. 1996). This simplifies the notation in equation 23 as:
$\omega_{i}\rho^{t}\pi_{t}c_{i,t}^{-\sigma}\text{exp}(\sigma\theta_{i,t})=\lambda_{t}.$ (25)
Taking the log of equation (25) and aggregating over households, consumption for household $i$ is expressed as:
$\log c_{i,t}=\log c_{t}^{a}+\frac{1}{\sigma}(\log\omega_{i}-\omega^{a})+(\theta_{i,t}-\theta_{t}^{a}),$ (26)
where
$c_{t}^{a}=\exp\left(\frac{1}{N}\sum_{i=1}^{N}\log c_{i,t}\right),~{}~{}~{}~{}~{}\omega^{a}=\frac{1}{N}\sum_{i=1}^{N}\log\omega_{i},~{}~{}~{}~{}~{}\theta_{t}^{a}=\frac{1}{N}\sum_{i=1}^{N}\theta_{i,t}.$ (27)
The first-difference in equation (26) eliminates the household time-constant effect to yield:
$\log c_{i,t}-\log c_{i,t-1}=\log c_{t}^{a}-\log c_{t-1}^{a}+(\theta_{i,t}-\theta_{i,t-1})+(\theta_{t}^{a}-\theta_{t-1}^{a}).$ (28)
Individual income is a conventional proxy variable for the preference shock (Mace 1991). Using the growth rate of household income $\log y_{i,t}-\log y_{i,t-1}$ as a proxy for idiosyncratic shock and adding a measurement error in consumption $\epsilon_{i,t}$, the empirical specification can be characterized as follows:
$\log c_{i,t}-\log c_{i,t-1}=\alpha_{1}(\log c_{t}^{a}-\log c_{t-1}^{a})+\alpha_{2}(\log y_{i,t}-\log y_{i,t-1})+\epsilon_{i,t}.$ (29)
The aggregate consumption growth in equation(29) is assumed to reflect macroeconomic shocks in Tsukishima in the case of my empirical setting. As discussed in Section 3.4, the consumption of the THBS households indeed captures the overall macroeconomic trend. Despite this, interpreting $\alpha_{1}$ is not easy because the THBS households include the factory workers’ households in Tsukishima alone. To deal with this potential issue, I use a two-way fixed-effects model instead of the first-difference model in the spirit of Cochrane (1991). Another potential issue is that $\epsilon_{i,t}$ may include the household-specific time-varying preference shock, which might be correlated with the idiosyncratic income shock. Family size influences the household’s preference because the marginal utility of consumption increases with family size (Jappelli and Pistaferri 2017). As explained, the family size had been fairly stable during the sample period in the THBS households, meaning that the preference shifts had rarely occurred (Section 3.4). To be conservative, however, I include the number of family members as a control variable for representing potential preference shock.
Considering these modifications, the empirical specification can finally be characterized as follows:
$\log c_{i,t}=\gamma\log y_{i,t}+\delta x_{i,t}+\mu_{i}+\phi_{t}+u_{i,t},$ (30)
where $c_{i,t}$ is consumption, $y_{i,t}$ is disposable income, $x_{i,t}$ is family size, $\mu_{i}$ is the household fixed effect, $\phi_{t}$ is the time fixed effect, and $u_{i,t}$ is a random error term. The estimate of $\gamma$ shall range from zero (i.e., idiosyncratic shocks are perfectly insured) to one (for the absence of insurance).
In order to test the risk-coping mechanisms, I consider the following representation of individual consumption in the spirit of Fufchamps and Lund (2003).
$c_{i,t}=y_{i}^{P}+y_{i,t}^{T}+r_{i,t},$ (31)
where $y_{i}^{P}$ is permanent income and $y_{i,t}^{T}$ is transitory income. $r_{i,t}=\bm{z}^{\prime}_{i,t}\bm{1}$ is a linear combination of the net temporary incomes expressed using a $k\times 1$ net income vector ($\bm{z}$) and a vector of ones. Substituting equation 31 into equation 26 and taking the first-difference removes the time-constant components, including the permanent income. It follows that my empirical specification can be characterized as follows:
$r_{i,t}=\kappa\tilde{y}_{i,t}+\eta x_{i,t}+\nu_{i}+\zeta_{t}+\epsilon_{i,t},$ (32)
where $\tilde{y}_{i,t}$ is head’s earning, $x_{i,t}$ is the family size control, $\nu_{i}$ indicates the household-fixed effect, $\zeta_{t}$ indicates the time-fixed effect, and $\epsilon_{i,t}$ is a random error term. I herein use the head’s earnings ($\tilde{y}_{i,t}$) as the idiosyncratic shock variable instead of the disposable income.112112112In the analysis, I consider the labor-supply adjustment variable and sales of miscellaneous assets as the additional dependent variables in addition to the net income variables $r_{i,t}$ (Section 4.2). The former variable is the total earnings of the family members except for the head, meaning that the simultaneous issue rises if I use the disposable income in this equation. The regression is separately run for each component in the net income vector ($\bm{z}$).
C.2 Raw Relationship between Income and Expenditures
Figure C.1 summarizes the densities of the log-differences (LD) of the semi-monthly expenditures. Figure C.1a illustrates the LD distribution of the total expenditure, whereas Figures C.1bC.1l summarize the LD distributions of the 11 subcategories of the expenditure. These figures confirm no specific outliers in all these expenditure categories.
Figure C.2 summarizes the scatter plots between the LD of the semi-monthly disposable income and expenditures. Figure C.2a illustrates the relationship between the LD of the semi-monthly disposable income and total expenditure. Figures C.2bC.2l illustrate the correlations between the LD of the semi-monthly disposable income and 11 expenditure subcategories.
Generally, rent was paid once per month. This means that the semi-monthly dataset fails to capture the variations in rent payment because it includes some semi-monthly cells with zero values. In addition, the water bills are included in the rent subcategory but not in the utility subcategory. This means that a semi-month cell takes a minimal value when the payment of rent is not included in that cell. Figure C.1c indeed shows that the LD of rent generates some larger values in the semi-monthly panel dataset due to those water bill payments. These cells taking large LD values substantially increased the slope, as shown in Figure C.2c.
Given this technical issue, the adjusted monthly panel dataset is suitable for estimating the income elasticity for the rent category. In fact, Figure C.4c confirms that the scatter plot from the adjusted monthly variation shows a flatter relationship.
(a) Total expenditure
(b) Food
(c) Housing
(d) Utilities
(e) Furniture
(f) Clothing
(g) Education
(h) Medical
(i) Entertainment
(j) Transportation
(k) Gift
(l) Miscellaneous
Figure C.1: Distributions of the Log-differences of the Semi-monthly Expenditures
Notes: The distribution of the log-difference in semi-monthly total expenditure is shown in Figure C.1a. Figures C.1bC.1l show the distributions of the log-differences in semi-monthly expenditure for the 11 subcategories listed in panel A of Table 3.
Sources: Created by the author using the THBS sample.
(a) Total expenditure
(b) Food
(c) Housing
(d) Utilities
(e) Furniture
(f) Clothing
(g) Education
(h) Medical
(i) Entertainment
(j) Transportation
(k) Gift
(l) Miscellaneous
Figure C.2: Raw Relationship between the Log-differences
in Semi-monthly Income and Expenditure
Notes: Figure C.2a illustrates the relationship between the log-differences in semi-monthly income and expenditure. Figures C.2bC.2l illustrate the relationships between the log-differences in semi-monthly disposable income and the 11 expenditure subcategories listed in Panel A in Table 3. The range of the y-axis is fixed across all the figures for comparability.
Sources: Created by the author using the THBS sample.
(a) Total expenditure
(b) Food
(c) Housing
(d) Utilities
(e) Furniture
(f) Clothing
(g) Education
(h) Medical
(i) Entertainment
(j) Transportation
(k) Gift
(l) Miscellaneous
Figure C.3: Distributions of the Log-differences of the Monthly Expenditures
Notes: The distribution of the log-difference in adjusted monthly total expenditure is shown in Figure C.3a. Figures C.3bC.3l show the distributions of the log-differences in adjusted monthly expenditure for the 11 subcategories listed in panel A of Table 3.
Sources: Created by the author using the THBS sample.
(a) Total expenditure
(b) Food
(c) Housing
(d) Utilities
(e) Furniture
(f) Clothing
(g) Education
(h) Medical
(i) Entertainment
(j) Transportation
(k) Gift
(l) Miscellaneous
Figure C.4: Raw Relationship between the Log-differences
in Monthly Income and Expenditure
Notes: Figure C.4a illustrates the relationship between the log-differences in semi-monthly income and expenditure. Figures C.4bC.4l illustrate the relationships between the log-differences in semi-monthly disposable income and the 11 expenditure subcategories listed in Panel A in Table 3. The range of the y-axis is fixed across all the figures for comparability.
Sources: Created by the author using the THBS sample.
C.3 Retail Sales in Tokyo City: A Retail Store Survey of 1935
Table C.1 summarizes the contribution of regular customers to their total sales, social classes of the customers, and sales methods in the retailers selling white rice, fish, greengrocers, and fuels (firewood and charcoal) by scale of business. In the original document, the figures for “Total” in the firewood and charcoal category are missing in “Sales Method (%)”. For these three cells, I filled the averages over the three business-size categories in each sales method: 20.4% for “Cash”, 78.4% for “Credit purchase”, and 1.3% for “Monthly installment”.
Table C.1: Customers, Sales Methods, and Competitiveness in the Retailers in Tokyo City
Selling Item Scale of Number of Share of sales Social Class of Sales Method (%) Number of
Business stores for regular the Regular Customer (%) Cash Credit Monthly competitors
reported customer (%) Upper Middle Other purchase installment reported
White Rice Small 19 69.7 12.1 49.5 38.4 29.5 70.2 0.3 7.3
Medium 55 69.5 17.5 41.6 40.9 26.9 72.6 0.5 8.7
Large 27 81.7 15.9 51.5 32.6 26.7 71.0 2.3 8.8
Total 101 73.1 16 45.7 38.2 26.9 71.5 1.7 8.5
Fish Small 34 64.5 5.9 53.2 40.9 42.7 57.3 0 5.1
Medium 20 62.6 10 58.0 32.0 43.8 56.2 0 5.3
Large 5 72.0 0 86.0 14.0 44.8 55.2 0 3.4
Total 59 64.6 6.8 57.6 35.6 43.9 56.1 0 5.0
Greengrocer Small 46 42.2 8 39.2 52.7 59.7 40.3 0 6.5
Medium 17 57.1 8.8 43.5 47.6 65.0 35.0 0 6.1
Large 3 36.7 33.3 40.0 26.7 79.7 20.3 0 3.7
Total 66 45.4 9.4 40.4 50.2 65.8 34.2 0 6.2
Firewood and charcoal Small 78 68.5 13.5 50.7 35.8 25.0 74.2 0.8 8.4
Medium 17 77.7 23.5 49.1 27.4 19.3 79.1 1.6 8.9
Large 5 33.3 23.0 60.0 17.0 16.8 81.8 1.4 10.6
Total 100 51.1 15.7 50.9 33.4 20.4 78.4 1.3 8.6
Notes:
This table summarizes the statistics on the sales in the retail shops selling white rice, fish, fruit and vegetables, and firewood and charcoal in Tokyo surveyed in 1935.
1. Small, middle, and large-scale businesses include retailers with annual sales of less than 10,000, 10,000–30,000, and 30,000+ yen, respectively.
2. The number of stores reported the share of regular customers (%) are 90, 56, 61, and 93 in the white rice, fish, greengrocers, and firewood and charcoal categories, respectively.
3. The share of regular customers is defined as the percentage of sales for regular customers in the annual total sales in 1935.
4. The definition of the social classes (“Upper”; “Middle”; “Other”) is not clearly defined in the report. Therefore, this must be based on the subjective judgment of the retailers.
5. The share of sales method is based on the percentage of sales using each method in the annual total sales in 1935.
6. The number of competitors indicates the number of peers within a three-cho (approximately 327 meters) radius.
Source: Tokyo City Chamber of Commerce and Industry 1937, pp. legend; 20–21; questionnaire.
C.4 Retail Sales in Tokyo City: A Small and Medium Scale Enterprises Survey of 1930
Table C.2 summarizes the use of credit purchase and the losses on the credit sales documented in the 1930 Survey Report. Following the report’s style, the item categories are divided into 6 categories: white rice, other grains, vegetables, fruit, fish, and fuel. The total number of retailers in each category is 157, 7, 67, 16, 46, and 64, respectively. As explained in Section 5, most of them used credit purchases as their selling method. Importantly, among the retailers using credit purchases, almost all of them suffered losses in sales. For example, 99% of white rice retailers experienced any losses. The minimum value of this share is still 97% for the fuel retailers. This means that credit purchase was inherently a loss-making sales method for the retailers.
The statistics in Table C.2 also indicate that many cases have unknown loss amounts. Overall, however, most of the cases seem to be grouped into bins of less than 5% of losses. For example, 87% (97 retailers) of the 111 available rice retailers suffered losses amounting to 5% of sales. Similarly, 65% (32 retailers) of the 49 available fuel retailers are grouped into the same loss category. Systematic statistics are unavailable to indicate the percentage of total profits corresponding to these losses. However, it is worth mentioning that a large part of losses was capped by 5% of total sales in all these different retailers. In other words, this implies that while the retailers could not avoid losses, they might have set prices that anticipated predicted losses.
Table C.2: Sales Methods and Losses on Sales under Credit Purchases in Tokyo City
Selling Item
White rice Other grains Vegetables Fruit Fish Fuel
Total number of retailers 157 7 67 16 46 64
Cash payment only 2 0 10 8 8 3
(never used the credit purchase)
Retailers using credit purchases 155 7 57 8 58 61
      Losses among sales amount (%)
              $\leq$ 1% 59 1 14 4 10 18
              2–5% 38 1 19 2 8 14
              6–10% 8 1 4 0 3 12
              11–20% 3 0 0 0 1 3
              21–30% 2 0 1 0 0 0
              30+0 0 0 2 0 0
              Unknown 44 4 19 0 16 12
      No loss 1 0 0 0 0 2
Retailers using credit purchases (%) 99 100 85 50 79 95
Share of retailers experienced losses (%) 99 100 100 100 100 97
Notes: This table summarizes the statistics on the sales methods and losses on sales under credit purchases among the retailers measured in Tokyo City surveyed in 1930. Source: Tokyo City Office 1932, pp. 454–469.
References used in the Appendices
  • (1)
  • (2) Arellano, Manuel. “Computing standard errors for robust within-groups estimators.” Oxford Bulletin of Economics and Statistics 49, no. 4 (1987): 431–434.
  • (3)
  • (4) Cochrane, John H. “A Simple Test of Consumption Insurance.” Journal of Political Economy 99, no. 5 (1991): 957–976.
  • (5)
  • (6) Choi, In “Unit root tests for panel data”. Journal of International Money and Finance 20, no. 2 (2001): 249–272.
  • (7)
  • (8) James, John A., and Isao Suto. “Early Twentieth-century Japanese Worker Saving: Precautionary Behaviour before a Social Safety Net.” Cliometrica 5, no. 1 (2011): 1–25.
  • (9)
  • (10) Jappelli, Tullio, and Luigi Pistaferri. The economics of consumption: Theory and evidence. New York, NY: Oxford University Press, 2017.
  • (11)
  • (12) Kawai, Takao. “Shakai chōsa houhoushi ni tsuite (On history of social research methods)” [in Japanese]. Hōgaku Kenkyū 53, no. 9 (1980): 41–67.
  • (13)
  • (14) Kawai, Takao. “Tsukishima chōsa saikōsatsu, ichi (Tsukishima Survey reconsidered (1))” [in Japanese]. Hōgaku Kenkyū 54, no. 8 (1981a): 1–26.
  • (15)
  • (16) Kawai, Takao. “Tsukishima chōsa saikōsatsu, ni kan (Tsukishima Survey reconsidered (2, end))” [in Japanese]. Hōgaku Kenkyū 54, no. 9 (1981b): 1–33.
  • (17)
  • (18) Mace, Barbara J. “Full Insurance in the Presence of Aggregate Uncertainty.” Journal of Political Economy 99, no. 5 (1991): 928–956.
  • (19)
  • (20) Mazzocco, Maurizio and Shiv Saini. “Testing Efficient Risk Sharing with Heterogeneous Risk Preferences.” The American Economic Review 102, no. 1 (2012): 428–468.
  • (21)
  • (22) Mazzocco, Maurizio and Shiv Saini. “Testing Efficient Risk Sharing with Heterogeneous Risk Preferences: Reply.” The American Economic Review 108, no. 10 (2018): 3114–3115.
  • (23)
  • (24) Miyoshi, Toyotaro. “Tsukishima chōsa no seiritsu to sonokeika ni tsuite (The establishment and progress of the Tsukishima Survey)” [in Japanese]. Meiseidaigaku Kenkyū Kiyō 16, (1980): 31–48.
  • (25)
  • (26) Takano, Iwasaburo. Honpō shakaitōkei ron (Social Statistics in Japan) [in Japanese]. Tokyo: Kaizō sha.
  • (27)
  • (28) Pearson, Karl. On the criterion that a given system of deviations from the probable in the case of a correlated system of variables is such that it can be reasonably supposed to have arisen from random sampling. Philosophical Magazine 50, no. 302, (1900): 157–175.
  • (29)
  • (30) Schulhofer-Wohl, Sam. “Heterogeneity and Tests of Risk Preference”. Journal of Political Economy 119, no.5, (2011): 925–958.
  • (31)
  • (32) Sekiya, Koichi. Seikatsu koten sōsho, volume 6: Tsukishima chōsa (A classic series of books about life, volume 6: Tsukishima Survey) [in Japanese]. Tokyo: Kōseikan.
  • (33)
  • (34) Shrinivas, Aditya and Marcel Fafchamps. “Testing Efficient Risk Sharing with Heterogeneous Risk Preferences: Comment.” The American Economic Review 108, no. 10 (2018): 3104–3113.
  • (35)
  • (36) Shibuya, Ryuichi, Kameji Suzuki, and Shojiro Ishiyama. Nihon no Shichiya (Pawn Shops in Japan) [in Japanese]. Tokyo: Waseda University Press, 1982.
  • (37)
  • (38) Shibuya, Ryuichi. Kourigashikinyū no tenkai kōzō (Development and Structure of Usuries) [in Japanese]. Tokyo: Nihon Tosho Center, 2000.
  • (39)
  • (40) Takeda, Naoko. “Shitsuteki chōsa data no nijibunseki (Qualitative Secondary Analysis and Reuse of the Original Research Data Set of the Japanese Sociological Classic)” [in Japanese]. Nihonrōdōkenkyūzasshi 57, no. 12 (2015): 70–80.
  • (41)
  • (42) Tanaka, Hikaru. Mouhitotsu no kinyū system (Another Financial System) [in Japanese]. Nagoya: Nagoya University Press, 2018.
  • (43)
  • (44) Terade, Kouji. “Tsukishima hōkokusho dai nishū (A study of “The volume II of the report on the investigation at Tsukishima”)” [in Japanese]. Mitagakkaizasshi 75, no. 6 (1982): 94–108.
  • (45)
  • (46) Odaka, Konosuke. “Japanese-style Labour Relations.” In The Japanese Economic System and Its Historical Origins, edited by Tetsuji Okazaki and Masahiro Okuno-Fujiwara. Oxford: Oxford University Press, 1999.
  • (47)
  • (48) Patrick, Royston and Douglas G. Altman. “Regression using Fractional Polynomials of Continuous Covariates: Parsimonious Parametric Modeling.” Journal of the Royal Statistical Society. Series C (Applied Statistics) 43 no. 3 (1994): 429–467.
  • (49)
  • (50) Wooldridge, Jeffrey M. Econometric Analysis of Cross Section and Panel Data. Cambridge: MIT Press, 2002.
  • (51)
Documents and Database used in the Appendices
  • (1)
  • (2) Central Federation of Industrial Associations. Sangyō kumiai chōsa siryō, volume 9 (Survey report of the industrial association, volume.9). [in Japanese] Tokyo: Central Federation of Industrial Associations, 1925.
  • (3)
  • (4) Committee for the History of the Labor Movement. Nihon rōdōundō shiryō (Historical Documents of the Japanese Labor Movement, volume 10). [in Japanese] Tokyo: Committee for the History of the Labor Movement, 1964.
  • (5)
  • (6) Chūō Ward. Chūō ku enkaku zushū, Tsukishima hen (Chūō Ward History Chart, Tsukishima edition). [in Japanese] Tokyo: Jinbun sha, 1994.
  • (7)
  • (8) Department of Health, Ministry of the Interior. Tokyo-shi Kyobashi-ku Tsukishima ni okeru jicchichōsahōkoku (dai isshū) (Report of the Field Study in Tsukishima in Kyobashi Ward of Tokyo City, Volume 1) [in Japanese]. Tokyo: Department of Health, Ministry of the Interior, 1923a.
  • (9)
  • (10) Department of Health, Ministry of the Interior. Tokyo-shi Kyobashi-ku Tsukishima ni okeru jicchichōsahōkoku (dai isshū, furoku no ichi) (Report of the Field Study in Tsukishima in Kyobashi Ward of Tokyo City, Appendix 1 of Volume 1) [in Japanese]. Tokyo: Department of Health, Ministry of the Interior, 1923b.
  • (11)
  • (12) Department of Health, Ministry of the Interior. Tokyo-shi Kyobashi-ku Tsukishima ni okeru jicchichōsahōkoku (dai isshū, furoku no ni) (Report of the Field Study in Tsukishima in Kyobashi Ward of Tokyo City, Appendix 2 of Volume 1) [in Japanese]. Tokyo: Department of Health, Ministry of the Interior, 1923c.
  • (13)
  • (14) Ito, Osamau, and Saito Nao. Kinyū gyō(Financial Industry) [in Japanese]. Tokyo: Nihon Keieishi Kenkyūjyo, 2019.
  • (15)
  • (16) Japan Meteorological Agency (database) [in Japanese]. https://www.data.jma.go.jp/gmd/risk/obsdl/index.php.
  • (17)
  • (18) Kitazawa, Shinjiro. Tokyo niokeru kikaikogyō no jyukurenshokkō toshiteno shiagekō narabini senbankō no chinginhōkoku (The report on a survey of wages of fitters and turners as skilled-workers in the machine industry in Tokyo). [in Japanese] Tokyo: Ohara Institute of Social Research, 1924.
  • (19)
  • (20) Kyoto City Office.Kyōto shi tōkeisho, dai jyūsan kai (Statistical Table of Kyoto City, volume. 13). [in Japanese] Kyoto: Kyoto City Office, 1922.
  • (21)
  • (22) Kyoto Prefecture.Kyōto fu tōkeisho, taishō kyū nen, dai san hen (Statistical Table of Kyoto Prefecture, 1920 edition, vol. 3). [in Japanese] Kyoto: Kyoto Prefecture, 1922.
  • (23)
  • (24) Ministry of Agriculture and Commerce. Kōjyō Tsūran (Taishō jyūnen jyūgatsu) (Handbook of Factories, November 1921 edition) [in Japanese]. Tokyo: Ministry of Agriculture and Commerce, 1921.
  • (25)
  • (26) Ministry of Posts and Telecommunications. Denki jigyō yōran, dai jyūyon kai (Handbook of Electricity Industry, vol. 14). [in Japanese] Tokyo: Electric Association, 1922.
  • (27)
  • (28) Nagoya City Office.Nagoya shi tōkeisho, dai nijyūni kai (Statistical Table of Kyoto City, volume. 22). [in Japanese] Nagoya: Nagoya City Office, 1922.
  • (29)
  • (30) Ohkawa, Kazushi, Miyohei Shinohara, and Mataji Umemura. Estimates of Long-term Economic Statistics of Japan since 1868, volume.8. [in Japanese] Tokyo: Toyo Keizai Shinposha, 1967.
  • (31)
  • (32) Osaka City Office. Osakashi tōkeisho, Taishō kyūnen (Statistical Table of Osaka City, 1920 Edition). [in Japanese] Osaka: Osaka City Office, 1922.
  • (33)
  • (34) Postal Savings Bureau. Chochikukyoku nenpō, volume 30 (Annual Statistical Report of the Postal Savings Bureau, volume 30). [in Japanese] Postal Savings Bureau, 1924.
  • (35)
  • (36) Statistics Bureau of the Cabinet. Taishō kyūnen kokusei chōsa hōkoku, zenkoku no bu, dai ikkan (Population Census of Japan, Vol. 1, 1920 Edition). [in Japanese] Tokyo: Statistics Bureau of the Cabinet, 1928.
  • (37)
  • (38) Statistics Bureau of the Cabinet. Taoshō kyūnen kokuseichōsahōkoku, zenkoku no bu, dai nikan (Population Census of Japan, Vol. 2, 1920 Edition). [in Japanese] Tokyo: Statistics Bureau of the Cabinet, 1929a.
  • (39)
  • (40) Tokyo Chamber of Commerce.Tokyoshinai shōshōkōgyō kinyūchōsa siryō (Survey report on the finance for small commercial and industrial enterprises in Tokyo City) [in Japanese]. Tokyo Chamber of Commerce, 1918.
  • (41)
  • (42) Tokyo Chamber of Commerce and Industry.Tokyoshinai niokeu kourigyōkeiei narabini kinyūchōsa (Survey on retail business management and finance in Toky City) [in Japanese]. Tokyo Chamber of Commerce and Industry, 1937.
  • (43)
  • (44) Tokyo City Social Affairs Bureau.Tokyo shinai saimin no nyūshichi nikansuru chōsa (Survey on the use of pawnshops among the poor in Tokyo City) [in Japanese]. Tokyo City Social Affairs Bureau, 1921.
  • (45)
  • (46) Tokyo City Office.Tokyo shi tōkei nenpyō, dai jyūnanakai (Annual Statistics of the City of Tokyo, vol.17) [in Japanese]. Tokyo: Tokyo City Office, 1921.
  • (47)
  • (48) Tokyo City Office.Tokyo shi tōkeisho, dai jyūhachi kai (Annual Statistics of the City of Tokio, volume. 18). [in Japanese] Tokyo: Tokyo City Office, 1922b.
  • (49)
  • (50) Tokyo City Office.Tokyo shi tōkeisho, dai jyūkyū kai (Annual Statistics of the City of Tokio, volume. 19). [in Japanese] Tokyo: Tokyo City Office, 1923.
  • (51)
  • (52) Tokyo City Office. Tokyo shi shisei tōkei genpyō, dai ikkan (Tokyo City Municipal Statistics Original Table, vol.1) [in Japanese]. Tokyo: Tokyo City Office, 1922a.
  • (53)
  • (54) Tokyo City Office. Chūshō shōkōgyōsha no kinyūmondai (Chūshō shōkōgyōsha shinkō chōsakai siryō, daini) (Financial issues for small and medium commercial and industrial enterprises (Material of the Committee for the Promotion of Small and Medium Commercial and Industrial Development, volume 2)) [in Japanese]. Tokyo: Tokyo City Office, 1935.
  • (55)
  • (56) Tokyo City Statistics Division. Tokyo shi oyobi kinkō chōson rōdōtōkeigenpyō (dai ikkai, dai nikan) (Labor Statistics Table for Tokyo City and Suburbs, issue 2 of volume 1) [in Japanese]. Tokyo: Tokyo City Office, 1926a.
  • (57)
  • (58) Tokyo City Statistics Division. Shichiyagyō no tōkei chōsa (Statistical Survey of the Pawnshop Business) [in Japanese]. Tokyo: Tokyo City Office, 1926b.
  • (59)
  • (60) Tokyo Institute for Municipal Research. Toshishominkinyū nikansuru chōsa, daiissatsu toshikinyū no gaikyō (Survey of Financial Institutions among Common People, volume. 1, Overview of the Financial Institutions in the Cities) [in Japanese]. Tokyo Institute for Municipal Research, 1925a.
  • (61)
  • (62) Tokyo Institute for Municipal Research. Toshishominkinyū nikansuru chōsa, dainisatsu toshichochiku ginkō (Survey of Financial Institution among Common People, volume. 2, Municipal Savings Banks) [in Japanese]. Tokyo Institute for Municipal Research, 1925b.
  • (63)
  • (64) Tokyo Institute for Municipal Research. Toshishominkinyū nikansuru chōsa, daisansatsu shigaichi shinyōkumiai (Survey of Financial Institutions among Common People, volume. 2, Credit Unions in the cities) [in Japanese]. Tokyo Institute for Municipal Research, 1925c.
  • (65)
  • (66) Tokyo Institute for Municipal Research. Toshishominkinyū nikansuru chōsa, daiyonsatsu kousetsu shichiho (Survey of Financial Institutions among Common People, volume. 2, Pubic pawnshops) [in Japanese]. Tokyo Institute for Municipal Research, 1926.
  • (67)
  • (68) Tokyo Legislative Research Association. Chokin to Mujin (Saving and Mutual Loan Association) [in Japanese]. Tokyo: Tokyo Legislative Research Association, 1915.
  • (69)
  • (70) Umemura, Mataji, Takamatsu, Nobukiyo, and Itoh, Shigeru. Estimates of Long-term Economic Statistics of Japan since 1868, volume.13. [in Japanese] Tokyo: Toyo Keizai Shinposha, 1983.
  • (71)
  • (72) Umemura, Mataji, Akasaka, Keiko, Minami, Ryoshin, Takamatsu, Nobukiyo, Arai, Kurotake, and Itoh, Shigeru. Estimates of Long-term Economic Statistics of Japan since 1868, volume.2. [in Japanese] Tokyo: Toyo Keizai Shinposha, 1988.
  • (73)
  • (74) Water Works Association. Jyōsuidō tōkei oyobi hōkoku (Statistics and Reports of the Water Supply, Vol. 1). [in Japanese] Tokyo: Water Works Association, 1922.
  • (75)
Archival Sources
  • (1)
  • (2) The Ohara Institute for Social Research, Archives of the Tsukishima Survey (THBS, unreleased).
  • (3)
  • (4) The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (7-2).
  • (5) The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (7-3).
  • (6) The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (7-4).
  • (7)
  • (8) The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (8-12).
  • (9) The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (8-13).
  • (10) The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (8-14).
  • (11) The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (8-15).
  • (12)

### v10

@@
-Consumption Smoothing in Metropolis: Evidence from the Working-class Households in Prewar Tokyo
+Consumption Smoothing in Metropolis: Evidence from Working-class Households in Prewar Tokyo
 Abstract
-I analyze the risk-coping behaviors among factory worker households in early 20th-century Tokyo.
-I have digitized a unique daily longitudinal survey dataset on household budgets to analyze how consumption is impacted by idiosyncratic shocks.
-I find that while the households were so vulnerable that the shocks impacted their consumption levels, the income elasticity for food consumption is relatively low in the short-run perspective.
-The result from mechanism analysis suggests that credit purchases with local retailers smoothed short-run food consumption.
-The event-study analysis using the adverse health shock as the idiosyncratic income shock confirms the robustness of the results.
-I also find evidence that the misassignment of payday in data aggregation results in a systematic attenuation bias due to measurement error in the standard consumption smoothing regressions.
-Keywords: consumption smoothing; credit purchases; daily budget data; local retailers; measurement error; risk-coping strategy; risk-sharing
+I analyzed the risk-coping strategies among factory worker households in early 20th-century Tokyo.
+I digitized and analyzed a unique daily longitudinal household budget survey to determine how consumption was impacted by idiosyncratic shocks.
+I found that while the households were so vulnerable that the shocks impacted their consumption levels, the estimated income elasticity for food consumption was relatively low in the short run.
+The event-study analysis using adverse health shock confirms the robustness of the results.
+The result of mechanism analysis suggests that credit purchases with local retailers smoothed short-run food consumption.
+Despite the potential loss of profit, retailers in a competitive situation allowed consumers to trade on credit.
+This shows the roles of informal credit institutions in mitigating vulnerability among urban worker households.
+Keywords: Consumption smoothing; credit institution; urban local community; measurement error; risk-coping strategy; risk-sharing
 JEL Codes: E21; N35
@@
 Economic theory illustrates the mechanism behind the smoothing of consumption in complete markets (Arrow and Debreu 1954).
-A series of empirical studies have tested the smoothness of consumption among households in various economies under systematic empirical specification.
-Evidence shows that markets in contemporary developing countries are often imperfect.111 The early representative literature includes, for example, Townsend (1994), Townsend (1995), Ravallion and Chaudhuri (1997), Grimard (1997), and Fafchamps and Lund (2003).
+Additionally, a series of empirical studies have tested the smoothness of consumption among households in various economies under systematic empirical specification.
+Evidence shows that markets in contemporary developing countries are often imperfect.111 The early representative literature includes, for example, Townsend (1994), Townsend (1995), and Ravallion and Chaudhuri (1997).
 Morduch (1999), Townsend (1995), and Dercon (2004) provide reviews of the earlier studies.
 Attanasio and Pistaferri (2016) and Meyer and Sullivan (2023) provide recent discussions.
-The households are vulnerable to risk, and the idiosyncratic shocks impact the levels of their consumption because shocks are not fully insured in the markets.
-Therefore, consumption smoothing is an essential agenda for mitigating poverty, especially in developing economies.
+The households are vulnerable to risk, and idiosyncratic shocks impact the levels of their consumption because shocks are not fully insured in the markets.
+Therefore, consumption smoothing is an essential mechanism for mitigating poverty, especially in developing economies.
 [^2] Smoothness of consumption is also an important factor in human capital formation.
-For example, Foster (1995) shows that children’s growth patterns are influenced by the inefficiency of the credit markets using the floods of 1988 in Bangladesh as a shock.
-Rose (1999) investigates the relationship between consumption smoothing and excess losses of girls, revealing that the favorable rainfall shocks in early life increased the survival probabilities of the girls in rural India.
-Gertler and Gruber (2002) also offers evidence that the smoothing of consumption against health shocks is fairly imperfect in Indonesia.
-Historical economies offer experimental environments to investigate the means and degrees of consumption smoothing under less developed societies.
-Lessons from historical institutions in developed economies shall provide meaningful insights for today’s developing economies.
-Despite this advantage, little has been known about historical consumption smoothing behaviors, partially because of the lack of longitudinal household budget datasets.
-This study digitized a unique daily longitudinal budget survey on working-class households conducted in Tokyo around 1919 and employed systematic econometric methods to explore the consumption behaviors among factory worker households in a representative manufacturing area.
-The result shows that the full insurance hypothesis is strongly rejected for overall consumption expenditure.
-The estimated income elasticity is close to that for the rural households in southern India around 1980.
-This means that the factory worker households in historical metropolis suffered certain consumption losses when they faced idiosyncratic adverse income shocks.
-However, I also found evidence that while expenses on dispensable categories are likely to be cut in response to the idiosyncratic shocks, expenses on indispensable consumption categories are partially smoothed in the short-run.
-Mechanism analyses provide suggestive evidence that credit purchases in the local economy played an important role in mitigating adverse idiosyncratic shocks.
+For example, Foster (1995) investigates the shock of the 1988 floods in Bangladesh and shows that children’s growth patterns are influenced by the inefficiency of the credit markets.
+Rose (1999) investigates the relationship between consumption smoothing and excess losses of girls, revealing that the favorable rainfall events in early life increased the survival probabilities of girls in rural India.
+Gertler and Gruber (2002) also shows that consumption smoothing against health shocks is imperfect in Indonesia.
+Credit institutions in local society can mitigate idiosyncratic income shocks, especially when the availability of formal financial institutions is limited.
+Udry (1994) showed that state-contingent loans among relatives and neighbors could partially insure against idiosyncratic shocks on rural households in northern Nigeria.
+A detailed field survey revealed that asymmetric information was no longer important in those private credit contracts (Udry 1990).
+Fafchamps and Lund (2003) also presented evidence that the households in the rural Philippines used informal networks with relatives and friends as a risk-mitigation strategy for idiosyncratic shocks.
+Related studies consider risk-sharing among households with similar cultural and institutional backgrounds.
+Grimard (1997) showed that insurance among households in the same ethnic group could partially mitigate idiosyncratic shocks in the multiethnic economy of Côte d’Ivoire.
+Munshi and Rosenzweig (2016) presented the importance of rural insurance networks among workers in the same caste in the context of spatial wage inequality in India.
+The evidence suggests that the availability of informal credit institutions relates to the stability of consumption in rural economies.
+This study assesses risk-coping behavior in informal credit transactions in the urban community in early twentieth-century Tokyo.
+I digitized and analyzed a unique daily longitudinal budget survey on factory worker households conducted in a representative manufacturing area.
+I found that the full insurance hypothesis is strongly rejected for overall consumption expenditure, meaning that the factory worker households in the historical metropolis suffered certain consumption losses when they faced idiosyncratic adverse income shocks.
+However, expenses on indispensable consumption categories were partially smoothed in the short run.
+Credit purchases in local retailers played an important role in mitigating adverse idiosyncratic shocks, whereas the other risk-mitigation strategies were impractical.
+Credit purchases did not require an explicit contract and thus often led to a loss in sales for the retailers.
+However, to remain competitive, retailers allowed credit transactions for their customers to remain regular customers.
+This shows that informal credit institutions in urban local communities functioned as informal insurance networks in rural economies.
 The first contribution of this study is to offer a case study of a metropolis in a developing economy.
-Urbanization is now a general phenomenon in the world, and the shares of urban population in many developing countries, especially in East Asia, are reaching a similar level to the developed countries (Henderson and Turner 2020).
-This fact raises the need for a deeper understanding of consumption behavior among urban working households in the process of economic development.
-Despite this, there is still little evidence for the urban worker households in the literature.
-While Townsend (1995) offers the results for Bangkok in Thailand between 1975 and 1990, a large part of the existing studies focused on the rural economies (Townsend 1994; Ravallion and Chaudhuri 1997; Gertler and Gruber 2002).
-[^3] Recent field experimental studies also focus on the village economies (Comola and Prina 2023; Macours et al.
+Urbanization is now a worldwide phenomenon.
+Moreover, the proportions of urban population in many developing countries, especially in East Asia, are reaching a similar level to the developed countries (Henderson and Turner 2020).
+This requires a deeper understanding of consumption behavior among urban working households in the economic development process.
+However, risk-mitigation strategies among urban worker households have been understudied.
+[^3] Extant studies have focused particularly on the rural economies (e.g., Townsend 1994; Ravallion and Chaudhuri 1997; Gertler and Gruber 2002).
+Recent field experimental studies also focus on village economies (Comola and Prina 2023; Macours et al.
 2022).
-A few exceptional studies provide cases on the middle-sized contemporary economies (Skoufias and Quisumbing 2005; Aguila et al.
-2017).
-In historical context, Ogasawara (2024) analyzes the factory workers’ households in Osaka City around 1920, which was roughly half the size of Tokyo City (Online Appendix Table A.1).
-This study considers a set of homogeneous households in a representative manufacturing area in prewar Tokyo, which provides a valuable empirical setting for testing the risk-sharing behaviors in the historical metropolis.
-[^4] Focusing on the sample sharing the similar occupational and regional characteristics systematically mitigates bias in estimating income elasticity under the standard risk-sharing regressions due to the heterogeneity in the risk preferences (Schulhofer-Wohl 2011; Mazzocco and Saini 2012).
-I found evidence of consumption smoothing using credit transactions with local retailers among working-class households.
-Macroeconomics literature has investigated the theoretical mechanisms behind the false rejections of the null of the risk-sharing hypothesis.
-An essential factor disturbing the valid statistical inference is measurement error in aggregate data (Cochrane 1991; Nelson 1994; Dynarski et al.
-1997; Gervais and Klein 2010).
-This study reveals a specific generating process of systematic measurement errors in the aggregated income and expenditure series in the survey dataset.
-In my empirical setting, the factory workers had two paydays in a month: one was set in the middle and the other at the end of the month.
-Hence, the panel dataset based on the calendar months leads to a substantial mismatch between the aggregate income and expenditure series.
-[^5] If payday is set soon after the initial date in a certain time cell, it does not cause a mismatch in income and expenditure.
-However, when payday is set just before the initial date in the cell, it leads to measurement errors because the consumption schedule in that cell shall based on the earnings in the previous cell.
-In other words, this way of matching assumes that the households in a certain cell make their consumption schedule based on future earnings, which is unrealistic.
-I found that the aggregation ignoring this mechanism causes false rejections in most consumption subcategories due to the attenuation bias.
-This evidence shows the importance of reflecting the timing of paydays in the aggregation and invokes careful reviews of the results from the aggregate budget datasets.
-The second contribution is to add a new analytical perspective to economic history studies.
-How working-class households overcame their economic hardships during the late 19th to the early 20th century Britain has attracted wider attention.
-Weights of evidence from these studies have revealed the roles of historical financial institutions and labor supply adjustments in compensating income losses in working-class households.
-[^6] Horrell and Oxley (2000) found the important roles of the labor supply adjustments among family members in coping with the income losses of the head in working-class households in late 19th-century Britain.
-Scott and Walker (2012) also found that interwar British working-class households had access to various institutions such as clothing clubs, hire-purchase, contractural savings, and insurance.
-The households with less earnings were likely to use credit-based institutions like clubs and hire-purchase, whereas those with higher incomes could use the savings institutions.
-Kiesling (1996) and Boyer (1997) focused on the risk-coping behaviors among the cotton workers households in Lancashire during the recession due to Civil War.
-Another strand of literature in financial history provided supply-side (lender-side) evidence on the functions of the micro-financial institutions for the workers (O’Connell and Reid 2005).
-Despite this, there is little evidence of how those risk-coping strategies contributed to smooth household consumption.
-In addition, the previous studies have utilized the cross-sectional variations in income to estimate the elasticities for the temporary income sources.
-[^7] For example, Horrell and Oxley (2000) and James and Suto (2011) consider the deviations of actual earnings from the predicted earnings to sort out the shocks.
-While Scott and Walker (2012) considers the smoothness of expenditure rather than income based on the deviation from the mean expenditure, their analyses employed the cross-sectional variations in the pooled cross-sectional sample.
-An exceptional study is Saaritsa (2011), which shows the estimated elasticity to temporary income from a fixed effect one-way error component model using a longitudinal household budget survey in 1928 Helsinki.
-Hence, the estimated elasticity tells the mean tendency that the households with lower (higher) income were more (less) likely to use a specific risk-coping strategy.
-[^8] In other words, the elasticity is derived from the observations across different households with different income levels.
-Since risk-coping is a concept about the dynamic behavior of a specific household, however, comparing the different time cells would be more natural to evaluate the roles of the risk-coping strategies.
-By using the panel dataset, this study implements the stylized econometric method to systematically investigate the households’ smoothing mechanisms to offer comparable estimates.
-Another historical contribution is adding micro-level evidence on consumer behaviors during industrialization.
+Although several studies considered consumption smoothing among urban households in postwar Bangkok and prewar Osaka, the risk-mitigation mechanisms, particularly via credit institutions, are not fully evaluated.
+[^4] See Townsend (1995) and Ogasawara (2024), respectively.
+Osaka was middle scale city of approximately half the size of Tokyo City circa 1920 (Online Appendix Table A.1).
+(Skoufias and Quisumbing 2005; Aguila et al.
+2017) also provided case studies on middle-sized contemporary economies.
+However, they did not consider the roles of local credit institutions.
+This study considers a set of homogeneous households in a representative manufacturing area in prewar Tokyo, which provides a valuable empirical setting for testing risk-sharing strategies in the historical metropolis.
+[^5] Focusing on a sample that shares similar occupational and regional characteristics systematically mitigates bias in estimating income elasticity under the standard risk-sharing regressions due to the heterogeneity in the risk preferences (Schulhofer-Wohl 2011; Mazzocco and Saini 2012).
+My findings show that informal credit institutions in the local urban community could mitigate vulnerability among households similar to rural developing economies (Udry 1994; Fafchamps and Lund 2003).
+The second contribution is to add analytical perspectives to economic history studies.
+First, studies of how working-class households overcame their economic hardships during the late 19th to the early 20th century have focused primarily on Britain.
+The weight of evidence supports the roles of historical financial institutions and labor supply adjustments in compensating for income loss in working-class households.
+[^6] Horrell and Oxley (2000) highlight the important role of labor supply adjustments among family members to manage the income losses of the head in working-class households in late 19th-century Britain.
+Scott and Walker (2012) also found that interwar British working-class households had access to various institutions such as clothing clubs, hire-purchase, contractual savings, and insurance.
+The households with less earnings were likely to use credit-based institutions like clubs and hire-purchase, whereas those with higher incomes could use savings institutions.
+Kiesling (1996) and Boyer (1997) focused on the risk-mitigation strategies among the cotton workers households in Lancashire during the recession caused by the Civil War.
+However, there is little evidence of how those risk-coping strategies smoothed their consumption.
+Similarly, previous studies have used cross-sectional variations in income to estimate the elasticities of the temporary income sources.
+[^7] For example, Horrell and Oxley (2000) and James and Suto (2011) consider the deviations of actual earnings from the predicted earnings as the shocks.
+Scott and Walker (2012) employed the cross-sectional variations in the pooled cross-sectional sample to analyze the smoothness of expenditure (deviation from mean expenditure).
+An exceptional study is Saaritsa (2011), which shows the estimated elasticity of temporary income from a fixed effect one-way error component model using a panel household budget survey in 1928 Helsinki.
+The estimated elasticity is, therefore, potentially disturbed by household-specific unobservable confounding factors such as preference and learning ability.
+As risk-mitigation is a concept in the dynamic behavior of a specific household, the estimator derived from within variations of the units is more suitable for evaluating the roles of risk-mitigation strategies.
+This study implements a stylized econometric method for the panel dataset to offer comparable estimates for future studies.
+Second, this study adds micro-level evidence on consumer strategies during industrialization.
 Francks (2009) provides a comprehensive view of the long-run consumption history in Japan from the Edo era to the twentieth century.
-[^9] There are several business history studies analyzing the relationship between consumption and production.
-A representative work is Gordon (2013), who investigated the expansion of the sewing machine from many aspects, including the production, distribution, and gender social norms in Japan.
-However, these studies have little interest in the households’ risk-coping behaviors.
-A few studies investigated on the relationships between household chores and consumption behaviors in interwar households (Tanimoto 2011, 2012).
-This study employs a different analytical view by focusing on short-run consumption behaviors.
-Looking into the extent and how the local society could ensure the households’ consumption would provide valuable insights into the consumption history literature.
-The rest of this paper is organized as follows.
-Section 2 overviews the Japanese economy around the First World War (WWI) and reviews Tsukisima, the target area in my analysis.
-I also summarized the financial institutions that were accessible to working-class households.
-Section 3 introduces the documents used to construct my panel dataset and discusses the sample characteristics.
-Section 4 summarizes the empirical strategies and results.
-Section 5 discusses the roles of credit purchases in the local economy.
-Section 6 concludes.
+[^8] Several business history studies analyze the relationship between consumption and production.
+A representative work is Gordon (2013), which investigated the expansion of the sewing machine from many perspectives, including production, distribution, and gender social norms in Japan.
+However, these studies have not focused on the households’ risk-mitigation strategies.
+Some studies investigated the relationships between household chores and consumption strategies in interwar households (Tanimoto 2011, 2012).
+This study employs an alternative analytical view by focusing on short-run consumption strategies.
+An examination of how, and the extent to which, the local society could ensure household consumption provides valuable insights into consumption history literature.
+Third, several financial historical studies provide supply-side (lender-side) evidence on the functions of the micro-financial institutions for the workers (O’Connell and Reid 2005).
+This study is the first to document the function of informal credit transactions by local retailers in the context of consumption smoothing.
+Munshi and Rosenzweig (2016) linked spatial wage inequality and men’s mobility in India with the availability of risk-mitigation networks among workers in the same caste.
+From this perspective, lower barriers of local credit institutions may motivate rural workers to urbanize, which may partially explain why Tokyo grew so fast circa WWI.
+This hypothesis requires testing in future research.
 2 Historical Background
-2.1 Japanese Economy around the First World War
-Japan had the foundations of an industrial nation by WWI.
-Between 1885 and 1915, real GNP grew annually by 2.6%, while population increased at 1.1%.
-[^10] Given that economic growth is faster than population growth, this implies that modern economic growth had started in this period as defined by Simon Kuznets (Miyamoto 2008, pp.
+2.1 Japanese Economy circa the First World War
+Between 1885 and 1915, the Japanese real GNP grew annually by 2.6%, whereas the population increased by 1.1%.
+[^9] Faster economic than population growth implies that modern economic growth had started in this period as defined by Simon Kuznets (Miyamoto 2008, pp.
 60; 64).
@@
 Japan increased exports to the Asian market on behalf of those European countries and exported munitions and foodstuffs to the European Allies, leading to a current account surplus.
-Given the declines in the export capacity of the European countries, import substitution had progressed.
-This expands the domestic market and leads to a rise in the machinery, metal, and chemical sectors.
-A worldwide shortage of ships also stimulated the development of shipbuilding, which proliferated as a leading industry.
-The development of the shipbuilding industry stimulated demand for the machinery and steel industries.
+The import substitution expanded the domestic market and led to a rise in the machinery, metal, and chemical sectors.
+A worldwide shortage of ships also stimulated the development of shipbuilding, which proliferated.
+The development of the shipbuilding industry further stimulated demand for the machinery and steel industries.
 Through this process of heavy and chemical industrialization, the total number of enterprises increased.
@@
 250).
-Year-on-year growth in the number of incorporated firms was 17.8% in 1918, 22.3% in 1919, and 13.5% in 1920.
-These are the top three years of the 34 years from after the Russo-Japanese War to 1940 (Nakamura and Odaka 1989, p.
+Year-on-year growth in the number of incorporated firms was 17.8%, 22.3%, and 13.5% in 1918, 1919, and 1920, respectively.
+These are the top three of the 34 years from after the Russo-Japanese War to 1940 (Nakamura and Odaka 1989, p.
 22).
-Figure 1 shows that the GNP growth from 1915-19 was 11.5% in nominal terms (4.4% in real terms), the highest growth throughout the 1910s and 1920s.
+Figure 1 shows that the GNP growth from 1915–19 was 11.5% in nominal terms (4.4% in real terms), the highest growth throughout the 1910s and 1920s.
 This rapid industrialization impacted the Japanese economy in many ways.
 Due to export-led economic growth, large inflows of gold increased domestic currency balances, causing prices to soar.
-Although wages began to rise around 1916 as a response to rising prices and labor demand due to the increased number of firms, wage increases failed to keep pace with soaring prices.
+Although wages began to rise circa 1916 as a response to rising prices and labor demand due to the increased number of firms, wage increases failed to keep pace with soaring prices.
 By the end of the war, the number of labor disputes grew sharply, and the labor movement flourished.
 Furthermore, the rise of business accelerated urbanization as the rural population moved to cities in search of job opportunities.
-The share of large cities (with people of 50,000 or more) as a percentage of the total population was less than 10% in 1889 but increased to nearly 16% in 1920 (Online Appendix Figure A.1a).
-The growth rate was exceptionally high during the WWI period, indicating that structural changes through the war accelerated the rise of the urban population.
+The share of large cities (with a population of $50,000$ or more) as a percentage of the total population was less than $10$% in 1889 but increased to nearly $16$% in 1920 (Online Appendix Figure A.1a).
+The growth rate was exceptionally high during WWI, indicating that structural changes through the war accelerated the rise of the urban population.
 Urbanization has increased public investment in large cities, especially in construction investment in social capital such as roads and water supply systems.
-[^12] Descriptions on the export-led growths, urbanization, and investments are from Nakamura and Odaka (1989, pp.
+[^11] Descriptions of the export-led growths, urbanization, and investments are from Nakamura and Odaka (1989, pp.
 21–23) and Nakamura (1971, pp.
 142; 147).
-Exports in the 1920s were lower than before because of the appreciation of the yen against the real exchange rate due to the accumulation of domestic and foreign species during WWI.
-Since the current account deficit was settled through the disbursement of foreign species owned by the government and the Bank of Japan, there was no reduction in the domestic money supply despite the sizeable current account deficit.
-This resulted in the domestic prices, especially prices of nontradable goods, being kept high.
+Exports in the 1920s were lower because of the appreciation of the yen against the real exchange rate due to the accumulation of domestic and foreign currencies during WWI.
+As the current account deficit was settled through the disbursement of foreign currencies owned by the government and the Bank of Japan, there was no reduction in the domestic money supply despite the sizeable current account deficit.
+This resulted in the domestic prices (especially prices of non-tradable goods) remaining high.
 The first half of the 1920s saw a temporary decline in industrial investment as the capital investment boom of the WWI period came to a halt under declining demand.
-[^13] On the other hand, construction investment by the private sector, including electric power and private railways, increased (Nakamura 1971, p.
+[^12] Conversely, construction investment by the private sector, including electric power and private railways, increased (Nakamura 1971, p.
 145).
@@
 Note: This figure shows the gross national product per capita (solid line) and personal consumption expenditure (dashed line) between 1910 and 1926.
-The prices are deflated using the CPI in 1934-1935 price.
-Sources: Created by the author using Ohkawa et. al. (1974, p. 237).
-As described, there were short-term business cycles, such as the boom caused by WWI and the subsequent recession.
-However, taken as a whole, Japan achieved a high economic growth rate relative to other countries in the early 20th century and before WWII.
-As shown in Figure 1, the growth rates of personal consumption expenditure were particularly high during and immediately after WWI, with average growth rates reaching 112% in 1915-19 and 118% in 1920-24.141414Total consumption as a percentage of the GNP remained above 80%.
-Although this level is not apparent among other countries (Kuznets 1962), the growth rate of real personal consumption per capita was relatively high: from 1875 to 1939, the average growth rate of personal consumption was 1.36%, and the rates never turned negative except during the Russo-Japanese War period.
-This was a high level even by international standards.
-Given the rapid urbanization, therefore, it is plausible to focus on the consumption behavior of urban households when considering the Japanese economy in this period.
-However, progress has yet to be made.
-[^15] Descriptions on the personal consumpsion expenditure are from Nakamura (1971, pp.
+The prices are deflated using the CPI in 1934–1935 price.
+Source: Created by the author using Ohkawa et. al. (1974, p. 237).
+As described, there was a short-term business cycle characterized by the boom caused by WWI and the subsequent recession.
+However, as a whole, Japan achieved a high economic growth rate relative to other countries in the early 20th century.
+As shown in Figure 1, the growth rates of personal consumption expenditure were particularly high during and immediately after WWI, with average growth rates reaching 112% in 1915–19 and 118% in 1920–24.
+This was a high level by international standards.
+[^13] Total consumption as a percentage of the GNP remained above 80%.
+Although this level is not apparent among other countries (Kuznets 1962), the growth rate of real personal consumption per capita was relatively high: from 1875 to 1939, the average growth rate of personal consumption was 1.36%, and the rates never became negative except during the Russo-Japanese War period.
+Descriptions on the personal consumption expenditure are from Nakamura (1971, pp.
 2–5; 27; 30–31).
+This highlights the importance of revealing the consumption behavior of urban households to understand the mechanisms underpinning the rapid Japanese economic growth in this period.
 2.2 Tsukishima: A Representative Manufacturing Area
 The population in Tokyo City has been growing since the end of the 19th century, and industrialization due to WWI spurred its growth.
-Tokyo City’s population share increased from 2.5% in 1889 to approximately 4% by 1920.161616In 1920, there were more than 10,000 municipalities in Japan, suggesting how huge Tokyo was.
+Tokyo City’s population share increased from $2.5$% in 1889 to approximately $4$% by 1920.141414In 1920, there were more than $10,000$ municipalities in Japan, suggesting how large Tokyo was.
 There was a difference of about 1 million between Tokyo and Osaka, the second-largest city by population.
@@
 The border surrounded by Asakusa, Honjyo, Nihonbashi, and Fukagawa wards includes the Sumida River.
-The borders of each ward are based on the administrative district in 1920.
+The borders of each ward are based on the administrative districts in 1920.
 Geographical coordinate system is based on JGD2000/(B, L).
@@
 Blocks in brown are the warehouses.
-Small circles colored in black and white show the workplaces in the machinery sector and the other manufacturing sectors, respectively.
+Small circles colored in black and white show the workplaces in the machinery and other manufacturing sectors, respectively.
 Source: Department of Health, Ministry of the Interior 1923c, second map. The tone was adjusted by the author using Adobe Photoshop 24.7.0.
@@
 Online Appendix A.3 summarizes the development history of Tsukishima in detail.
-Column (2) of Panel A in Table 1 shows that Kyobashi had an industrial composition similar to that of Tokyo City in 1920.
+Column (2) of Panel A in Table 1 shows that Kyobashi had an industrial composition like that of Tokyo City in 1920.
 Column (3) confirms that Tsukishima is a representative industrial area in the city; approximately $68$% of male workers were employed in the manufacturing sector.
@@
 [^17] For example, a representative shipbuilding plant was located on Ishikawajima, an island in the north of Tsukishima (Figure 3).
-Ships were strongly influenced by the import substitution that was underway at the beginning of the 20th century, and domestic production of ships was progressing (Section 2.1).
-The production of a hydraulic 80-ton hardened crane by Ishikawajima Shipyard in 1917 is an example of the domestic output of a production good that had previously relied on imports (Nakamura and Odaka 1989, pp.
+Ships were strongly influenced by the import substitution that was underway at the beginning of the 20th century, and domestic production of ships progressed (Section 2.1).
+The production of a hydraulic 80-ton hardened crane by Ishikawajima Shipyard in 1917 is an example of the domestic output of a product that had previously relied on imports (Nakamura and Odaka 1989, pp.
 10; 32).
@@
 Panel A: This panel summarizes the industrial structures based on the occupations of male workers measured in the 1920 population census.
-Each share is calculated as the number of males worked as regular workers in each sector divided by the total number of male workers (in percentage).
+Each share is calculated as the number of males who worked as regular workers in each sector divided by the total number of male workers (in percentage).
 All the figures do not include the number of unemployed (mugyō sha) males.
@@
 The average age of males is calculated using the tables for population by age group reported in the census.
-Since the tables are open-ended, the class for over 60 years old is rounded as $64.5$ following the range of the former category (i.e., 50–59).
+As the tables are open-ended, the class for over 60 years old is rounded as $64.5$ following the range of the former category (i.e., 50–59).
 All the figures in this panel do not include a small number of quasi-households (jyun setai), which include the person, such as the students in dormitories and the patients in hospitals.
@@
 2–3; 42–46).
-To look more closely at the characteristics of working-class households in Tsukishima, Panel B in Table 1 summarizes the population and occupational statistics from the census.
-First, the average household size in Tsukishima is $4.3$, similar but slightly smaller than the values in Tokyo City ($4.6$) and Kyobashi Ward ($4.7$).
-The percentage of the married male population is $40$%, which is in the similar range of the figures of Tokyo ($38$%) and Kyobashi ($36$%).
+Panel B in Table 1 summarizes the population and occupational statistics from the census to examine the characteristics of working-class households in Tsukishima.
+First, the average household size in Tsukishima is $4.3$—similar to but slightly smaller than the values in Tokyo City ($4.6$) and Kyobashi Ward ($4.7$).
+The percentage of the married male population is $40$%, which is similar to the range of the figures for Tokyo ($38$%) and Kyobashi ($36$%).
 The average age of men in Tsukishima is $25.1$, slightly lower than the figures for Tokyo and Kyobashi.
 While the sex ratio by age group in Tsukishima shows a similar trend to that of Tokyo and Kyobashi, the ratios for the 1414141419191919 and 2020202029292929 age bins are higher than those.
-Overall, these statistics reflect the fact that Tsukishima is a typical industrial area and many young male workers were employed there.
+Overall, these statistics reflect the fact that Tsukishima was a typical industrial area and many young male workers were employed there.
 The slightly smaller household size in Tsukishima can be attributed to the relatively large number of young factory worker households.
 The percentage of married men in Tsukishima suggests that a certain proportion of the households in their early 20s might have been couples before they started raising children.
-Finally, I briefly overview the lives of the factory workers’ households in Tsukishima.
+Finally, I provide a brief overview of the lives of the factory workers’ households in Tsukishima.
 Factories and workers’ residences were scattered throughout the island.
+The factory workers lived in similar tenements and used the commercial stores and restaurants in the central lot.
 [^18] Tsukishima is composed of Ishikawajima island at the northern end, the adjacent Tsukudajima and Shin-tsukudajima islands, and the rectangular Lots 1–3 in the south (Online Appendix Figure A.3).
 Specifically, factories and worker residences were scattered throughout Shin-tsukudajima and Lots 1 and 2, whereas Lot 3, a relatively new reclaimed site completed in 1913, was undeveloped.
-The factory workers had lived in similar tenements and used the commercial stores and restaurants in the central lot.
-[^21] Online Appendix A.4 summarizes the housing in Tsukishima in detail using a housing survey.
-Regarding the stores, there were 209 stores selling daily necessities, 109 selling clothing, 74 selling tools, and 120 selling food, drink, and entertainment.
-As for medical personnel, there were eight doctors, four dentists, eight obstetricians, and 14 drug stores.
+Online Appendix A.4 summarizes the housing in Tsukishima in detail using a housing survey.
+There were $209$, $109$, $74$, and $120$ stores selling daily necessities, clothing, tools, and food, drink, and entertainment, respectively.
+Regarding medical personnel, there were eight doctors, four dentists, eight obstetricians, and $14$ drug stores.
 There were also ten bathhouses (Department of Health, Ministry of the Interior 1923a, pp.
 41–47).
-There were two postal offices, 13 pawnshops, and five moneylenders, and those were scattered throughout the island (Online Appendix Figure A.2).
+There were two postal offices, $13$ pawnshops, and five moneylenders, which were scattered throughout the island (Online Appendix Figure A.2).
 The main routes to the mainland were the wooden Aioi Bridge (Aioi bashi) and three ferries (watashi bune).
-[^22] See Online Appendix Figure A.3 for these routes.
 The average daily total number of people using the ferries in 1918 was reported to be $53,318$, accounting for more than twice as many people as the Tsukishima population.
@@
 30).
-The number of people and households of Tsukishima at the end of December 1918 is reported to be $24,399$ and $5,562$, respectively (National Police Agency 1920a, pp.
+The number of people and households in Tsukishima at the end of December 1918 was reported to be $24,399$ and $5,562$, respectively (National Police Agency 1920a, pp.
 50–51).
-These descriptions illustrate Tsukishima as a busy machine industry area during the industrialization.
+Online Appendix Figure A.3 shows the routes of the ferries.
+These descriptions illustrate Tsukishima as a busy machine manufacturing area during industrialization.
 2.3 Risk-coping Institutions
-Working-class households had been exposed to several risks that led to losses of income at that time.
-Generally, social insurance, such as health, accident, and unemployment insurance, can help mitigate the idiosyncratic income shocks on the workers.
-[^24] Gertler and Gruber (2002) shows the impacts of illness on consumption smoothing behavior in early 1990s Indonesia and argues the importance of public support for the illness.
-Kantor and Fishback (1996) show that the installation of workers’ compensation reduced the precautionary savings in the interwar U.S..
-Gruber (1997) provides evidence that unemployment insurance smooths individual consumption as the case of the U.S.
-from 1968–1987.
+Working-class households were exposed to several risks that led to losses of income at that time.
+Generally, social insurance, such as health, accident, and unemployment insurance, can mitigate idiosyncratic income shocks on workers.
+[^20] Gertler and Gruber (2002) shows the impact of illness on consumption-smoothing behavior in the early 1990s in Indonesia and argues for the importance of public compensation for illness.
+Kantor and Fishback (1996) show that the installation of workers’ compensation reduced the precautionary savings in the interwar US.
+Gruber (1997) provides evidence that unemployment insurance smooths individual consumption as was the case in the US from 1968–1987.
 However, comprehensive social insurance systems did not exist in the early 1920s in Japan.
-The former health insurance was enacted in 1927, and unemployment insurance had not been established throughout the prewar period.
-In addition, the Factory Act enacted in 1916 did also not specify compensation for the sickness of workers.
-Therefore, the workers had basically no public support for their illness, injuries, and joblessness.
-Instead, they might have used private financial institutions to deal with the idiosyncratic risks.
-Several financial institutions were indeed available for working-class households.
-A representative savings institution for households was the postal savings (yūbin chokin).
-In Tokyo City, roughly $1.4$ million people had postal savings accounts, covering approximately $67$% of the entire citizens in 1920 (Online Appendix Table A.2).
+The previous health insurance was enacted in 1927, and unemployment insurance was not established throughout the prewar period.
+Additionally, the Factory Act enacted in 1916 did not specify compensation for the sickness of workers.
+Therefore, they had no public support for their illness, injuries, and unemployment.
+Instead, they might have used private financial institutions to manage idiosyncratic risks.
+Several financial institutions were available for working-class households.
+A representative savings institution was the postal savings (yūbin chokin).
+In Tokyo City, approximately $1.4$ million people had postal savings accounts, covering approximately $67$% of the citizens in 1920 (Online Appendix Table A.2).
 The average savings amount per capita in Kyobashi Ward was $44.5$ yen (Tokyo City Office 1921, pp.
-892–893), which is less than one month’s earnings of the skilled-factory workers (Section 3.2).
-Savings bank (chochiku ginkō) is another famous savings institution among the working class in the 1910s (Ito and Saito 2019, pp.
+892–893), which was less than one month’s earnings for the skilled-factory workers (Section 3.2).
+Savings bank (chochiku ginkō) was another famous savings institution among the working class in the 1910s (Ito and Saito 2019, pp.
 79–81).
-The number of people with savings accounts was roughly $1.4$ million, accounting for approximately $58$% of total people in Tokyo (Online Appendix Table A.2).
-However, $84.8$% of all the depositors were the commerce and miscellaneous workers, and the manufacturing sector shared only $11.5$% in Kyobashi Ward (Online Appendix Table A.3).
-This reflects the fact that savings bank was an institution designed for lower-income workers than factory workers (Tokyo Institute for Municipal Research 1925b, p.
+The number of people with savings accounts was approximately $1.4$ million, accounting for approximately $58$% of total people in Tokyo (Online Appendix Table A.2).
+However, $84.8$% of all the depositors were commercial and miscellaneous workers, and the manufacturing sector shared only $11.5$% in Kyobashi Ward (Online Appendix Table A.3).
+This reflects the fact that savings banks were institutions designed for lower-income workers rather than factory workers (Tokyo Institute for Municipal Research 1925b, p.
 81).
 This is consistent with the fact that Tsukishima did not have any savings bank branches, while there were two postal offices (Section 2.2).
-Although another savings institution was the mutual loan association (mujin), it had been used among small business owners and merchants for running their businesses (Online Appendix A.2).
-[^25] Generally, the ordinary bank in prewar Japan was the institution for providing large loans to businesses (Teranishi 2011).
+Although another savings institution was the mutual loan association (mujin), it was used among small business owners and merchants for running their businesses (Online Appendix A.2).
+[^21] Ordinary banks in prewar Japan provided large loans to businesses (Teranishi 2011).
 There was no bank in Tsukishima, as workers did not borrow money from banks to make ends meet.
-The official report states that “a commercial bank is not necessary in a laboring area such as this Tsukishima, and even for large entrepreneurs such as factory owners, there is no particular need to establish a bank for their business transactions, and transactions with banks in other areas are sufficient” (Department of Health, Ministry of the Interior 1923a, p.
+The official report states that “a commercial bank is not necessary in a laboring area such as Tsukishima, and even for large entrepreneurs such as factory owners, there is no particular need to establish a bank for their business transactions; transactions with banks in other areas are sufficient” (Department of Health, Ministry of the Interior 1923a, p.
 48).
-There also be a few lending institutions.
-Credit purchases (kakegai) in retail shops were the most popular institution.
-Although the systematic statistics on credit purchases are unavailable, the THBS households often used credit purchases in various retail shops for rice, fish, vegetables, firewood, and charcoal.
-[^26] Cooperative societies were also available for worker households to buy daily commodities, albeit they were not lending institutions.
+There were also few lending institutions.
+Credit purchases (kakegai) with retailers were the most popular institution.
+Although systematic statistics of credit purchases are unavailable, the THBS households often used credit purchases in various retailers for rice, fish, vegetables, firewood, and charcoal.
+[^22] Cooperative societies were also available for worker households to buy daily commodities, although they were not lending institutions.
 There were $26$ cooperative societies in Tokyo prefecture in 1924, of which 23 were organized by the citizens and workers (Central Federation of Industrial Associations 1925, p.
 57–59).
-However, while the two largest societies had roughly $4,000$ members, the others were generally small societies with less than 100 members.
-This figure accounts for only a few percentage points in the workers in the city, given that the number of male workers was more than 768 thousand.
-Most of the THBS households had not used the cooperatives.
+However, while the two largest societies had approximately $4,000$ members, the others were generally small societies with less than 100 members.
+This figure accounts for only a few percentage of the workers in the city, given that the number of male workers was more than $768$ thousand.
+In fact, the THBS households rarely used the cooperatives.
 I will describe this institution in detail in Section 5.
 The pawnshop (shichiya) was another lending institution.
-The accessibilities were higher than the other lending institutions: [^27] The number of pawnshops in Tokyo were $1,334$ in 1918 and $1,261$ in 1919, respectively.
+The accessibility was higher than for the other lending institutions.
+[^23] The number of pawnshops in Tokyo was $1,334$ and $1,261$ in 1918 and 1919, respectively.
 The total number of cases was $8,226,883$ and $7,573,406$ each year.
 These figures are more than three times greater than the entire population in Tokyo City.
-there were $13$ pawnshops in Tsukishima, more than twice the large number of moneylenders (Section 2.2).
-However, the core users of the pawnshops were the low-income working-class households rather than the factory workers (Tokyo City Social Affairs Bureau 1921, p.
+There were $13$ pawnshops in Tsukishima, more than twice the number of moneylenders (Section 2.2).
+However, the primary users of the pawnshops were the low-income working-class households rather than the factory workers (Tokyo City Social Affairs Bureau 1921, p.
 9).
-According to the statistics on the pawnshop users in Tokyo prefecture in 1923, the most representative users were the day laborers and workers classified as “miscellaneous”, accounting for $63.4$% of all users.
-In contrast, those for agriculture, commerce, and manufacturing sectors were only $1.7$, $19.5$, and $15.6$%, respectively (Tokyo Institute for Municipal Research 1926, p.
+According to the statistics on the pawnshop users in Tokyo prefecture in 1923, the most representative users were the day laborers and workers classified as “miscellaneous,” accounting for $63.4$% of all users.
+By contrast, those for agriculture, commerce, and manufacturing sectors were only $1.7$, $19.5$, and $15.6$%, respectively (Tokyo Institute for Municipal Research 1926, p.
 25).
-The inexpensive article means that pawnshops were used to deal with very short-run necessities only.
+The inexpensive articles pawned suggest that the pawnshops were only used to dealing with very short-run loans for necessities.
 Among the pawnshops of Kyobashi, for example, the average amount per case was $7.8$ yen, which accounts for $12$% of the monthly disposable income of the THBS households (Panel A of Table 3).
-In this light, the pawnshop is a small lending institution for low-income working-class households who might have faced credit constraints.
-Lenders did not need to screen the borrower’s credit, and the borrowers did not worry about incurring heavy debt as the main articles were inexpensive clothes (Shibuya et al.
+Thus, the pawnshop was a small lending institution for low-income working-class households who needed credit.
+Indeed, lenders did not need to screen the borrower’s credit, and the borrowers did not worry about incurring heavy debt as the main articles pawned were inexpensive clothes (Shibuya et al.
 1982).
-In addition, the interest rates were regulated by the Pawnbroker Regulation Act of 1895, and the average redemption rate was substantially high under the lower interest rates.
-[^28] This means that the average redemption rate was substantially high: it was approximately $94$% in 1920 among $88$ pawnshops in Kyobashi ($401,222$ cases/$427,265$ cases) (Tokyo City Office 1922b, pp.
+Additionally, the interest rates were regulated by the Pawnbroker Regulation Act of 1895, and the average redemption rate was substantially high under the lower interest rates.
+[^24] The average redemption rate was approximately $94$% in 1920 among $88$ pawnshops in Kyobashi ($401,222$ cases/$427,265$ cases) (Tokyo City Office 1922b, pp.
 888–889).
-Other lending institutions are money lenders (kinsen kashitsuke gyō) and credit unions (shinyō kumiai).
-Generally, moneylenders were used by business owners with very high-interest rates (Shibuya 2000, pp.
+Other lending institutions were money lenders (kinsen kashitsuke gyō) and credit unions (shinyō kumiai).
+Generally, moneylenders were used by business owners with very high interest rates (Shibuya 2000, pp.
 184; 248).
-Similarly, the coverage of the credit unions was substantially low, say a few percentage points of the male workers in the manufacturing sector (Online Appendix A.2).
-Informal gifts may be another type of risk-coping institution, although systematic statistics are not available.
+Similarly, the coverage of the credit unions was substantially low, possibly a few percentage points of the male workers in the manufacturing sector (Online Appendix A.2).
+Informal gifts might have been another type of risk-mitigation institution, although systematic statistics are not available.
 A household survey on the $185$ factory-worker households in Tokyo suggests that the share of average monthly gifts in total income was approximately $3$% in November 1922 (Social Affairs Division 1925, pp.
-58-59).
-While this means that the gifts comprised a few percentage points of the monthly income, it is unclear whether these informal income transfers had functioned as mutual aid.
-Finally, life insurance was also not designed to compensate for temporary income losses.
-Generally, life insurance companies provided death insurance, meaning that temporary income reductions were not covered by insurance.
-Moreover, they did not target the working-class as customers (Usami 1984, pp.
+58–59).
+While this means that the gifts comprised a few percentage points of the monthly income, it is unclear whether these informal income transfers functioned as mutual aid.
+Finally, life insurance was not designed to compensate for temporary income losses.
+Generally, life insurance companies provided death insurance, meaning that temporary income reductions were not covered.
+Moreover, they did not target the working class as customers (Usami 1984, pp.
 87–88).
-Although postal life insurance (kani seimei hoken) operated by the government started in 1916, it only provided whole life insurance and endowment insurance.
-Thus, it also never covered the health shocks nor layoffs (Usami 1984, pp.
+Although postal life insurance (kani seimei hoken) operated by the government started in 1916, it only provided life and endowment insurance.
+Thus, it never covered illness or unemployment (Usami 1984, pp.
 102–107).
 [^25] There were also several employment agencies that introduced jobs to low-income job-seekers in Tokyo.
-However, a large part of them was unscrupulous and not commonly used at that time (Machida 2016).
-To summarize, while there were several institutions, the available institutions for the factory workers were few in the practical sense.
-For savings institutions, the postal savings might have been the best available device for Tsukishima workers.
-Regarding lending institutions, credit purchase is the plausible device among factory workers.
-While pawnshops were accessible financial institutions, they might not provide sufficient money as the articles were usually inexpensive.
-The factory workers in Tsukishima might have combined these devices to deal with the idiosyncratic income shocks.
+However, many of them were unscrupulous, and not commonly used at that time (Machida 2016).
+To summarize, while there were several lending institutions, few were available to the factory workers.
+Postal savings might have been the best available device for them.
+Credit purchase was the most plausible device regarding lending institutions.
+While pawnshops were accessible, they might not have provided sufficient money as the articles pawned were usually inexpensive.
+The factory workers in Tsukishima might have combined these devices to manage idiosyncratic income shocks.
 3 Data
-This section is designed to understand the characteristics of the data used in the study.
-I give an overview of the Tsukishima Survey in Section 3.1.
-I then utilize the statistics from the complete surveys in Section 3.2 to assess the representativeness of the sample households.
-In Section 3.3, I explain my data aggregation method to overcome the potential measurement error issue.
-The variables used, the trend in income and expenditure, and variations in idiosyncratic shocks are described in Section 3.4.
 3.1 Tsukishima Household Budget Survey
-The Tsukishima Survey was the first urban social survey conducted in Japan by the Ministry of Home Affairs around 1919.
-The purpose was to investigate the status of the lives of urban worker households amid the rapid progression of industrialization throughout WWI.
-It was designed as a field survey consisting of multiple survey items, including the budget survey (hereafter THBS).
-[^30] Online Appendix B.1 summarizes detailed information about the Tsukishima Survey.
-The THBS was the first household survey in Japan to use the budget book method and is considered the prototype for subsequent household surveys in Japan (Sekiya 1970, p.
+The Tsukishima Survey was the first urban social survey conducted in Japan by the Ministry of Home Affairs circa 1919.
+The purpose was to investigate the status of the lives of urban worker households amid the rapid industrialization throughout WWI.
+It was designed as a field survey consisting of multiple survey items, including the budget survey (hereinafter THBS).
+[^26] The THBS was the first household survey in Japan to use the budget book method and is considered the prototype for subsequent household surveys in Japan (Sekiya 1970, p.
 43).
-The Ministry of Interior published an official report on the Tsukishima Survey in 1921, which contains some aggregated information on the target households.
-However, household-level budget information is required to analyze consumption smoothing behaviors.
+Online Appendix B.1 summarizes detailed information about the Tsukishima Survey.
+The Ministry of Home Affairs published a few official reports on the Tsukishima Survey in 1921, which contained some aggregated information on the target households.
+However, household-level budget information is required to analyze consumption-smoothing strategies.
 Fortunately, the household budget book (kinsendeiri hikaechō) gathered in the THBS is preserved in the archives of the Ohara Institute of Social Research (hereafter, OISR).
-[^31] The author is permitted to access the materials related to the Tsukishima Survey by the OISR: Archives of the Tsukishima Survey (THBS, unreleased).
-While most of the sources are unreleased, the official reports of the survey published by the Ministry of Home Affairs are included in the archives of Gonda Yasunosuke at the OISR and are available to the public (The OISR, Archives of Gonda Yasunosuke (7-2; 7-3; 7-4)).
+[^27] The author was permitted to access those unreleased materials related to the Tsukishima Survey by the OISR: Archives of the Tsukishima Survey (THBS, unreleased).
+The official reports of the survey published in 1921 by the Ministry of Home Affairs are included in the archives of Gonda Yasunosuke at the OISR and are available to the public (The OISR, Archives of Gonda Yasunosuke (7-2; 7-3; 7-4)).
 These budget books were transferred to the OISR by Iwasaburo Takano, who was responsible for the Tsukishima Survey as a member of the 7th Section of the Health and Hygiene Research Committee of the Ministry of Home Affairs and became the first director of the OISR.
@@
 His contribution to the Tsukishima Survey is summarized in detail in the Online Appendix B.1.
-I corrected and digitized all the 40 THBS budget books for the factory worker households that were included in the official report.
-[^33] The archives of the OISR also reserve several other household budget books of factory worker households.
-However, these budget books are generally in poor record keeping.
+I corrected and digitized all forty THBS budget books for the factory worker households that were included in the official report.
+[^29] The archives of the OISR also preserve several other household budget books of factory worker households.
+However, these budget books contain poor record-keeping.
 Therefore, I do not use these incomplete books in this study.
@@
 38).
-This confirms that the information in these budget books is reliable and that the THBS budget dataset ensures a quality that allows quantitative analysis.
-[^34] Online Appendix B.1 provides finder details on how the Tsukishima Survey had been conducted carefully as a social survey project of Takano.
+This confirms that the information in these budget books is reliable and that the THBS budget dataset ensures a quality that enables quantitative analysis.
+[^30] Online Appendix B.1 provides finder details on how the Tsukishima Survey was conducted carefully as a social survey project of Takano.
 3.2 Sample Characteristics
 Analytical Sample
-The estimation strategy used for testing consumption smoothing and risk-coping behaviors requires the panel structures for aggregate income and expenditure (Section 4).
+The estimation strategy used for testing consumption-smoothing and risk-coping strategies requires panel structures for aggregate income and expenditure (Section 4).
+I use the semi-monthly series for baseline analysis, given that the head’s income was paid twice monthly.
+[^31] Section 3.3 provides details on the definitions of aggregate income and expenditure.
 Among the $40$ THBS households, there are five cross-sectional units in our semi-monthly data.
-[^35] Section 3.3 provides details on the definitions of aggregate income and expenditure.
-I excluded these five cross-sectional units, leaving 35 THBS households.
-Next, a household with incomplete income information, which is necessary for my analysis, is removed.
-Finally, I exclude a household for which no information on payments to the credit purchases was recorded.
-Consequently, my THBS sample for the semi-monthly series includes 33 households with panel structures from 12th January to 11th November 1919.363636I use adjusted monthly panel dataset for subanalysis in Section 4.
-Following the same trimming steps, my adjusted monthly panel dataset includes 26 households for the same period.
-Note that several units have only three semi-month cells, say less than two adjusted months.
-These units have panel structure in the semi-monthly panel, but not in the adjusted monthly panel.
+I excluded these five cross-sectional units, leaving $35$ households.
+Next, I removed a household with incomplete income information.
+Finally, I excluded a household for which no information on payments to the credit purchases was recorded.
+Consequently, my THBS sample for the semi-monthly series included $33$ households with panel structures from January 12 to November 11, 1919.323232I also used the adjusted monthly panel dataset for sub-analysis in Section 4.
+Following the same trimming steps, my adjusted monthly panel dataset included $26$ households for the same period.
+Note that several units have only three semi-month cells, for example, less than two adjusted months.
+These units have a panel structure in the semi-monthly panel dataset, but not in the adjusted monthly panel.
 Online Appendix B.3 summarizes the finer details of the original data structure and trimming.
-Before assessing the representatives of the THBS sample, I test the potential selection bias due to attrition of the units.
+Before assessing the representatives of the THBS sample, I tested the potential selection bias due to attrition of the units.
 The panel units with shorter (longer) time-series observations might have different preferences on the consumption than those with longer (shorter) time-series observations.
@@
 Online Appendix Table B.2 summarizes the results.
-All the estimated coefficients on the covariates are close to zero and statistically insignificant, and the Wald statistics support the null results.
-This result is robust to using different thresholds such as first quantile, median, and third quantile.
-This supports the evidence that the household’s preference is less likely to be correlated with unbalancedness.
+All the estimated coefficients on the covariates were close to zero and statistically insignificant, and the Wald statistics support the null results.
+This is robust to using different thresholds such as first quantile, median, and third quantile.
+This supports the evidence that the household’s preference was less likely to be correlated with a lack of balance.
 Representativeness
@@
 Following this definition, several “woodworking” factories measured in the Tsukishima Factory Survey are included in the machinery sector in column 2, as it consists of shipbuilding and wooden pattern factories.
-The report also supports that the wooden patterns are used in the casting process and shall be regarded as one of the machinery sector (Department of Health, Ministry of the Interior 1923a, pp.
+The report also supports that the wooden patterns are used in the casting process and shall be included in the machinery sector (Department of Health, Ministry of the Interior 1923a, pp.
 388–389).
 Panel B: The first and second rows in columns 1 and 2 show the average household size and sex ratio measured in the 1920 Population Census and Statistics of the National Police Agency of 1919, respectively.
-The third to eighth rows in columns 1 and 3 show household size distributions in the census and the THBS sample, respectively.
+The fourth to eighth rows in columns 1 and 3 show household size distributions in the census and the THBS sample, respectively.
 The figures in parentheses in Column (1) are the percentage share for family size bins between 2 and 8 people.
@@
 Gaussian-based 95% confidence interval (CI) using bootstrap standard error is reported in the brackets.
-Column 2 shows the average monthly wage of the male factory workers in the skilled age range (30-40 years old) in machinery factories in Tokyo City.
+Column 2 shows the average monthly wage of the male factory workers in the skilled age range (30–40 years old) in machinery factories in Tokyo City.
 This figure is calculated using the average monthly wage and average monthly ancillary wage measured in the manufacturing censuses.
@@
 726–757).
-Next, I assess the representativeness of the THBS sample in Tsukishima using the manufacturing and population censuses,
+Next, I assess the representativeness of the THBS sample in Tsukishima using the manufacturing and population censuses.
 Panel A of Table 2 summarizes the share of male workers in the manufacturing industry by sector.
-Statistics of the National Police Agency (hereafter SNPA) focusing on factories with more than or equal to 15 workers show that roughly 90% of male workers had been engaged in the machinery sector in 1919 (Column 1).
+Statistics of the National Police Agency (hereafter SNPA) covering factories with more than or equal to 15 workers show that approximately 90% of male workers were engaged in the machinery sector in 1919 (Column 1).
 Column 2 summarizes the statistics in the manufacturing census conducted in the Tsukishima Factory Survey.
-The proportion of machine factories is $85$% in larger factories and $92$% in smaller factories, indicating that most of the male workers worked in the machinery sector.373737 I have confirmed that the paternal occupational statistics for the primary school students in Tsukishima provide materially similar results (Online Appendix A.5).
+The proportion of machine factories is $85$% in larger factories and $92$% in smaller factories, indicating that most of the male workers worked in the machinery sector.
+[^33] I have confirmed that the paternal occupational statistics for the primary school students in Tsukishima provide materially similar results (Online Appendix A.5).
 In Column 3, I confirm that the THBS heads’ occupations show a similar proportion: $94$% worked in the machinery sector.
-In fact, the report states that “It is not too much to say that most of the industry in Tsukishima is machinery.
+In fact, the report states that “It is not too much to say that most of the industry in Tsukishima is machinery manufacturing.
 Therefore, to describe the labor situation in Tsukishima, it is sufficient to describe its machinery industry” (Department of Health, Ministry of the Interior 1923a, p.
@@
 The number of people in the THBS households ranges from 2 to 8, meaning the THBS sample does not include single or very large households.
-[^38] Those family households with 2–8 people cover most of the households in Tsukishima because the share of single and vary large (9+ people) households was roughly $10$% (Column 1 of Panel B).
+[^34] Those family households with 2–8 people cover most of the households in Tsukishima because the share of single and very large (9+ people) households was approximately $10$% (Column 1 of Panel B).
 In other words, the THBS mainly focuses on the consumption behavior in couples and households with a few children.
@@
 145).
-The official report of the Tsukishima Survey suggests that roughly nine out of ten factory workers in the skilled-age range were classified as skilled workers.
-Roughly six of those skilled workers were employed in the middle- to large-scale factories, and three worked in small-scale factories.
-[^39] Online Appendix A.5 summarizes this point in detail using the complete survey on the primary school students in 4th to 6th grades in Tsukishima.
+The official report of the Tsukishima Survey suggests that approximately nine out of ten factory workers in the skilled-age range were classified as skilled workers.
+Approximately six of those skilled workers were employed in the middle- to large-scale factories, and three worked in small-scale factories.
+[^35] Online Appendix A.5 summarizes this point in detail using the complete survey of the primary school students in the 4th to 6th grades in Tsukishima.
 Given this, the THBS was designed to investigate the budgets in skilled workers’ households.
-[^40] Shortly, Iwasaburo Takano, a survey director, regarded that the skilled workers in the modern industrial sector will constitute the core of the labor force, labor movement, and citizens.
+[^36] Shortly, Iwasaburo Takano, a survey director, regarded that the skilled workers in the modern industrial sector constituted the core of the labor force, labor movement, and citizens.
 The finer details in the background of this survey design are summarized in Online Appendix B.1.
 The average THBS head’s age is $33$, which is indeed in the skilled workers’ representative age range in the machinery factories (Kitazawa 1924).
-[^41] Specific occupations are not available for most heads because they usually list “factory workers” in the occupation section.
-However, sub-categories available at a few heads indicate the fact that they were lathe operator and finisher, the typical skilled workers in machinery factories.
-Online Appendix A.6 summarizes the works in the machinery factories in Tsukishima.
+[^37] Specific occupations are not available for most heads because they are usually listed as “factory workers” in the occupation section.
+However, sub-categories available for some heads indicate that they were the typical skilled workers in machinery factories such as lathe operators and finishers.
+Online Appendix A.6 summarizes the work in the machinery factories in Tsukishima.
 Column 1 of Panel C in Table 2 indicates that the average monthly earning of the THBS heads is $59$ yen (median = $56$ yen).
 In Column 2, I calculated the average monthly earning of male machinery factory workers in the skilled-age range using the manufacturing censuses, which shows similar monthly earnings ($56$ yen).
-[^42] Shortly, this figure is calculated using the average daily wage of male factory workers aged 30-40 and average annual working days measured in the manufacturing censuses.
+[^38] Shortly, this figure is calculated using the average daily wage of male factory workers aged 30–40 and average annual working days measured in the manufacturing censuses.
 Online Appendix B.6 provides the finer details of the calculation steps.
@@
 42).
-This provides evidence that the monthly earnings of the THBS heads and skilled workers’ households in Tsukishima are materially similar.
-To summarize, this evidence indicates that the THBS sample is useful in inferring the mean tendency of those households.
+This provides evidence that the monthly earnings of the THBS heads and skilled workers’ households in Tsukishima were materially similar.
+To summarize, this evidence indicates that the THBS sample is useful for inferring the mean tendency of those households.
 3.3 Aggregation
@@
 by Different Time-series Frequencies
-Note: Figure 4a4c, and 4e illustrate the time-series plots of the average daily, semi-monthly, and unadjusted semi-monthly income and expenditure, respectively.
+Notes: Figure 4a4c, and 4e illustrate the time-series plots of the average daily, semi-monthly, and unadjusted semi-monthly income and expenditure, respectively.
 Figure 4b4d, and 4f illustrate the time-series plots of the average daily, semi-monthly, and unadjusted semi-monthly net income (i.e., income minus expenditure).
-Figures 4a and 4b shows the daily data between 12th January and 11th November 1919.
-The (adjusted) semi-monthly series defines the 12-26th for the first and 27-11th for the second half.
-Figures 4c and 4d show the semi-monthly series calculated using the daily data between 12th January and 11th November 1919.
+Figures 4a and 4b show the daily data between January 12 and November 11, 1919.
+Figures 4c and 4d show the semi-monthly series calculated using the daily data between January 12 and November 11, 1919.
+The (adjusted) semi-monthly series defines the 12–26th for the first and 27–11th for the second half.
+Figures 4e and 4f show the unadjusted semi-monthly series calculated using the daily data between January 15 and October 31, 1919.
 The unadjusted semi-monthly series uses the 15th of each month as a threshold.
-Figures 4e and 4f show the unadjusted semi-monthly series calculated using the daily data between 15th January and 31st October 1919.
 Source: Created by the author using the THBS sample.
 It is necessary to summarize the daily budget information into a more extended time range because households need a set of time to smooth their consumption given a realized amount of income.
-To explore the reasonable way of aggregation, I first overview the daily time series on the total income and expenditure.
-Figure 4a shows that the head’s income had been paid twice monthly: the semi-monthly wages were paid on the middle (14–16th) and last day (30th or 31st).
-[^43] The head’s income was the dominant way in earnings because the breadwinner households had been typical at that time (Panel B in Table 3).
-Importantly, Figure 4a indicates that the households expensed much of their income just on the payday and spent the rest of the payment little by little until the next payday.
-In other words, the expenditures had never clearly increased just before the paydays.
-This means that the households had made their consumption decisions (schedules until the next payday) based on the income they obtained.
-Notice that calendar (semi-)monthly aggregation can not reflect this household behavior because calendar (semi-)month does not place the payday at the beginning of its cell.
-[^44] For example, a calendar semi-month cell in January, say the cell from 1st to 14th, starts on the 1st but its payday is set on the 14th.
-This assumes that the aggregate consumption in this semi-month cell is based on the earnings at the end of its cell.
+To explore a reasonable method of aggregation, I first overviewed the daily time series on the total income and expenditure.
+Figure 4a shows that the head’s income was paid twice monthly: the semi-monthly wages were paid on the middle (14–16th) and last day (30th or 31st).
+[^39] The head’s income was the dominant method for earnings because the breadwinner households were typical at that time (Panel B in Table 3).
+Importantly, Figure 4a indicates that the households expensed most of their income on the payday and spent the rest of the payment little by little until the next payday.
+In other words, the expenditure never clearly increased just before the paydays.
+This means that the households made their consumption decisions (schedules until the next payday) based on the income they obtained.
+Note that calendar (semi-)monthly aggregation cannot reflect this household behavior because calendar (semi-)month does not place the payday at the beginning of its cell.
+[^40] For example, a calendar semi-month cell in January, for example, the cell from 1st to 14th, starts on the 1st but its payday is set for the 14th.
+Accordingly, this assumes that the aggregate consumption in this semi-month cell is based on the earnings at the end of its cell.
 This is clearly an unrealistic assumption.
 Another source of measurement error is the shift of the payday: the payday for the latter half of a month was sometimes set on the first date of the following month.
-Although this shift did not occur frequently, the calendar month cell shall contain a substantial noise from the shift as it leads to the cells including more (three paydays) and fewer (one payday) payments than normal.
-Nelson (1994) suggests that this sort of mismatching in the timing of income and expenditure tends to induce the type II error in the test of consumption smoothing.
+Although this shift did not occur frequently, the calendar month cell might contain substantial interference from the shift as it leads to the cells including more (three paydays) and fewer (one payday) payments than normal.
+Nelson (1994) suggests that this sort of mismatching in the timing of income and expenditure induces type II errors when testing consumption smoothing.
 Adjustment
-In order to deal with the mismatch, I aggregated the daily observations between the 12th and 26th for the first half and those between the 27th and 11th of the next month for the second half.
-[^45] This means that the number of days included in a semi-month ranges from 13 to 16.
+To manage the mismatch, I aggregated the daily observations between the 12th and 26th for the first half and those between the 27th and 11th of the next month for the second half.
+[^41] This means that the number of days included in a semi-month ranges from 13 (i.e., in February) to 16 (e.g., January).
 Note that this heterogeneity in the number of days over different semi-months is a cross-sectional constant.
 Thus, these differences are entirely captured using the semi-month fixed effects in the regression.
-This range is based on the fact that almost all of the paydays in the second half of a month measured in the THBS dataset, are 27th–31st.
-[^46] Note that I do not use the definition of semi-month starting from the 2nd (to the 16th) in each month because this way of stratification places the payday at the end of each semi-month cell.
-Online Appendix B.7 provides the evidence that while the definition using 2nd in each month as the threshold provides an improved series in income, the expenditure can not chase the income in the right way because the expenses measured under this definition are based on the income in the previous semi-month cell.
-Again, I confirmed that my preferred definition using the 12th in each month as the threshold offers a clean series of income and expenditure.
+This range is based on the fact that paydays in the second half of a month measured in the THBS dataset ranged from 27th to 31st.
+[^42] Note that I do not use the definition of semi-month starting from the 2nd (to the 16th) in each month because this method of stratification places the payday at the end of each semi-month cell.
+Online Appendix B.7 provides the evidence that while the definition using 2nd in each month as the threshold provides an improved series in income, the expenditure cannot chase the income correctly because the expenses measured under this definition are based on the income in the previous semi-month cell.
+My preferred definition was using the 12th of each month as this threshold offers a clean series of income and expenditure.
 Figure 4c shows the time-series plots of the income and expenditure from this adjusted semi-monthly panel dataset, whereas Figure 4e illustrates the unadjusted (calender) semi-monthly series using the 15th of each month as the threshold of the semi-month.
-Clearly, Figure 4e provides very rough measures of income and expenditure due to the measurement errors from the misadjustment of paydays.
-The most significant improvement in Figure 4c is that the expenditure can chase the income in each month, reflecting the households’ consumption behaviors based on their income measured in the nearly exact timing.
-Given the improvements in both income and expenditure measures, the net income series shows a much smoother trend in Figure 4d, which are not available under the unadjusted series (Figure 4f).
+Clearly, Figure 4e provides very rough measures of income and expenditure due to the measurement errors from the maladjustment of paydays.
+The most significant improvement in Figure 4c is that the expenditure can chase the income in each month, reflecting the households’ consumption strategies based on their income measured in the approximate exact timing.
+Given the improvements in both income and expenditure measures, the net income series shows a much smoother trend in Figure 4d, which is unavailable under the unadjusted series (Figure 4f).
 In Section 4.1, I will demonstrate how these measurement errors in the unadjusted dataset attenuate the estimates.
@@
 These are used to test the smoothness of the consumptions in each subcategory.
-To test the roles of risk-coping strategies, I consider five net income categories: savings, insurance, borrowing, credit purchase, and gifts.
+To test the roles of risk-coping strategies, I considered five net income categories—savings, insurance, borrowing, credit purchase, and gifts.
 Net variables are defined as the income minus expenditure in each category; for example, net savings refers to the difference between the withdrawal amount and deposits to savings.
-I also consider the sales of miscellaneous goods and labor earnings from the family members except for the head to test the contribution of the asset sales and intrahousehold labor-supply adjustments.
+I also considered the sales of miscellaneous goods and labor earnings from the family members (except for the head) to test the contribution of the sales of small assets and intrahousehold labor-supply adjustments.
 Panel B of Table 3 summarizes these variables.
@@
 Alterations in family size may shift the household’s preferences (Jappelli and Pistaferri 2017).
-[^48] In village economies, landholding and asset stocks such as livestock and grain may also be correlated with the household’s risk-coping behaviors (e.g., Udry 1994, 1995).
-Since this study analyzes urban factory workers’ households, the roles of these assets can be discarded.
+[^44] In village economies, landholding and asset stocks such as livestock and grain may also be correlated with the household’s risk-coping strategies (e.g., Udry 1994, 1995).
+As this study analyzes urban factory workers’ households, the roles of these assets can be discarded.
 Fortunately, the archives of the Tsukishima Survey include a set of sheets documenting the changes in the composition of the THBS households (Kazoku idō hyō).
 [^45] The Ohara Institute for Social Research, Archives of the Tsukishima Survey (THBS, unreleased).
-I found that the size of households had been stable over time: one household experienced changes in the number of family members due to the birth of a son during the sample periods.
-This supports the evidence that the preference shifts had rarely occurred among the THBS households.
-Although family size is included in all regressions to be conservative, it has little effect on the results.
-[^50] The family size variable shall be omitted in several regressions due to the multicollinearity in the fixed-effects models given its small within variations.
-Although the number of family members in different age bins is available from the THBS, those family composition variables are also time-constant and are absorbed by the household fixed effect.
+I found that the size of households was stable over time: one household experienced changes in the number of family members due to the birth of a son during the sample periods.
+This supports the evidence that the preference shifts rarely occurred among the THBS households.
+Although family size was included in all regressions, to be conservative, it had little effect on the results.
+[^46] The family size variable was omitted in several regressions due to the multicollinearity in the fixed-effects models given its small within variations.
+Although the number of family members in different age bins is available from the THBS, those family composition variables are also time-constant and were absorbed by the household fixed effect.
 Panel C of Table 3 shows the summary statistics on the family size variable.
@@
       Transportation 0.4 289 -1.3 195 0.9 124 -0.7 98
-      Gift 3.1 289 0.5 243 6.3 124 1.3 114
+      Gifts 3.1 289 0.5 243 6.3 124 1.3 114
       Miscellaneous 0.5 289 -1.3 222 1.0 124 -0.7 111
@@
       Other family member’s earning 3.74 6.11 289 8.21 16.44 124
-      Sales of assets 0.02 0.12 289 0.06 0.17 124
+      Sales of miscellaneous assets 0.02 0.12 289 0.06 0.17 124
 Panel B-3: Income variable
-Head’s earning 29.43 18.11 289 59.2 29.9 124
+Head’s earnings 29.43 18.11 289 59.2 29.9 124
 Panel C: Family size (people)
@@
 Number of days with illness 0.2 1.7 0 16 254
-(the other family members)
+(other family members)
 Notes:
@@
 “Other family member’s earnings” refers to the total income earned by all family members except for the head of the household.
-“Sales of assets” includes the sale of daily miscellaneous goods such as newspapers and empty bottles.
+“Sales of miscellaneous assets” includes the sale of miscellaneous goods such as old newspapers and empty bottles.
 Panel C: The summary statistics of the family size variable for the 33-panel units in the semi-monthly panel dataset are listed.
-The statistics for the adjusted monthly panel dataset are not reported because they are materially similar to those for the semi-monthly panels.
+The statistics for the adjusted monthly panel dataset are not reported because they are materially similar.
 Panel D: The summary statistics of the number of days with illness for family members among the 33-panel units in the semi-monthly panel dataset are listed.
-Since both are lagged variables, the initial time period in all the units is omitted.
-Similarly, a household with two semi-month observations is excluded.
-Similarly, .
-Sources: Created by the author from the THBS sample.
+As those are lagged variables, the initial period is omitted in all the units.
+Accordingly, a household with two semi-month observations is excluded.
+Source: Created by the author from the THBS sample.
 Trend
 Next, I illustrate the time-series figures of the THBS dataset to overview the trend in income and expenditure.
-Figure 4c shows a weak increasing trend in the average income, which is consistent with the overall macroeconomic trend in the nominal income during this period.
-This trend in earnings was, however, offset by the steep inflation under the expansions after the war (Section 2.1).
-In Figure 4d, the deviation of expenditure from income in January must capture the expenses on the goods for the New Year events.
+Figure 4c shows a weak increasing trend in the average income, which is consistent with the overall macroeconomic trend in nominal income during this period.
+This trend in earnings was, however, offset by steep inflation caused by expansions after the war (Section 2.1).
+In Figure 4d, the deviation of expenditure from income in January captures the expenses on the goods for the New Year events.
 The gaps in June to the first half of July may reflect the payments of bonuses.
-The humps in March may reflect the preparation for the new fiscal and academic year from April, suggesting a higher proportion of temporary income from withdrawals and loans.
-These support the evidence that the budget data from the THBS households can capture a large part of the macroeconomic trends.
+The peaks in March may reflect the preparation for the new fiscal and academic year from April, suggesting a higher proportion of temporary income from withdrawals and loans.
+These show that the budget data from the THBS households can capture a substantial aspect of the macroeconomic trends.
 Idiosyncratic Shocks
@@
 For household $i$ in semi-month $t$ cell, the residual is then defined as $\Delta\hat{\upsilon}_{i,t}=\Delta\textit{Income}_{i,t}-\hat{\alpha}\Delta\overline{\textit{Income}}_{.,t}$, where $\overline{\textit{Income}}_{.,t}=\sum_{i=1}^{n_{t}}\frac{\textit{Income}_{i,t}}{n_{t}}$ and $\hat{\alpha}$ is the estimated coefficient.
-I used the semi-monthly panel dataset to overview the entire distribution of the idiosyncratic shocks because the daily panels include a set of consecutive cells with no income.
-While the residuals satisfy the zero-average property, they occasionally take large negative and positive values, suggesting that the workers had experienced idiosyncratic income shocks in some cells.
-Regardless of favorable or adverse shocks, these cells that deviated from mean tendency are helpful in investigating how those households coped with the idiosyncratic shocks.
-Despite this, it is helpful to focus on the illness as it is the least predictable shock (Gertler and Gruber 2002).
-Although joblessness might be another shock, the THBS heads rarely changed their occupation during the survey period as they are the skilled workers who were less likely to move to other jobs.
-[^52] Another possibility is that the THBS heads who changed their jobs might have been omitted from the survey.
-However, I confirmed that the households with longer and shorter time-series observations have similar family characteristics (Online Appendix B.3).
-Thus, this omission does not lead to selection bias in terms of the potential vulnerabilities to the idiosyncratic risk.
-Moreover, the number of dismissals at the Tsukishima factories seems to be small.
-In fact, the average monthly dismissal rate (i.e., the number of dismissals/workers) from June 1919 to May 1920 at the one factory for which statistics are available from the official report of the Tsukishima Survey was only 1.5% (Department of Health, Ministry of the Interior 1923a, p.
-415).
-In contrast, fortunately, the illness among the THBS family is documented systematically.
-As explained in Section 2.3, neither the comprehensive health insurance system nor the factory’s compensation for sickness was available.
+While the residuals satisfy the zero-average property, they occasionally take large negative and positive values, suggesting that the workers experienced idiosyncratic income shocks in some cells.
+Regardless of favorable or adverse shocks, these cells that deviated from the mean tendency help investigate how those households coped with the idiosyncratic shocks.
+Despite this, it is useful to consider illness as it is the least predictable shock (Gertler and Gruber 2002).
+Fortunately, the illness among the THBS families is documented systematically and thus, available for analysis.
+As explained in Section 2.3, neither a comprehensive health insurance system nor compensation for illness from the factories was available.
 Moreover, the sanitary conditions in the cities were low, leading to higher mortality rates.
 [^48] See Schneider and Ogasawara (2018) for the overviews on the sanitary conditions in interwar Japan.
-The official report mentioned that the workers in the machinery factories in Tsukishima had faced the risk of diseases under severe sanitary conditions (Department of Health, Ministry of the Interior 1923a, p.
+The official report mentioned that the workers in the machinery factories in Tsukishima faced the risk of diseases under severe sanitary conditions (Department of Health, Ministry of the Interior 1923a, p.
 14).
-Considering this, I first provide the overall results from the specification using the changes in earnings as the idiosyncratic income shock in Sections 4.1 and 4.2.
-I then consider the structure behind the adverse income shocks using the heads’ temporary illness to investigate the households’ consumption and risk-coping behaviors in Section 4.3.
+Although unemployment might be another shock, the THBS heads rarely changed their occupation during the survey period as they were skilled workers who were less likely to move to other jobs.
+[^49] Another possibility is that the THBS heads who changed their jobs might have been omitted from the survey.
+However, I confirmed that the households with longer and shorter time-series observations had similar family characteristics (Online Appendix B.3).
+Thus, this omission does not lead to selection bias in terms of the potential vulnerabilities to the idiosyncratic risk.
+Moreover, the number of dismissals at the Tsukishima factories would be small.
+In fact, the average monthly dismissal rate (i.e., the number of dismissals/workers) from June 1919 to May 1920 at the one factory available from the official report of the Tsukishima Survey was only $1.5$% (Department of Health, Ministry of the Interior 1923a, p.
+415).
 Panel D of Table 3 shows the summary statistics on the health shock variables.
 Risk Preferences
-This study investigates consumption smoothing behaviors among the skilled factory workers’ households in the machinery sector in Tsukishima.
-This empirical setting is preferable because the different attitudes toward risks among households with different occupations and regions may lead to bias in the standard risk-sharing regressions (Schulhofer-Wohl 2011; Mazzocco and Saini 2012).
-To assess the homogeneity of the risk preference among THBS households, I tested whether the earnings of the heads who worked in a few specific factories or sectors showed different responses to the aggregate shocks in the spirit of Schulhofer-Wohl (2011).
-Specifically, I consider the case that the heads who worked in large-scale or governmental factories to have potentially different risk preferences because those workers might have lower risk tolerance than those who worked in smaller enterprises.
-I confirmed that the sensitivities to the aggregate shocks are materially similar in both groups in the statistical sense.
-[^54] Online Appendix B.10 summarizes the finer details of this analysis.
-In short, the heads’ earnings are regressed on the aggregate consumption, its interaction term with respect to the indicator variable for the large-scale and government factories, family size variable, and household fixed effects.
+The different attitudes toward risks among households with different occupations and regions may lead to bias in the standard risk-sharing regressions (Schulhofer-Wohl 2011; Mazzocco and Saini 2012).
+In contrast, this study investigates consumption-smoothing strategies among homogeneous households — skilled factory workers’ households in the machinery manufacturing sector in Tsukishima.
+To assess the homogeneity of the risk preference among THBS households, I tested whether the earnings of the heads who worked in a few specific factories or sectors showed different responses to the aggregate shocks, following Schulhofer-Wohl (2011).
+Specifically, I consider the case that the heads who worked in large-scale or governmental factories might have had different risk preferences because those workers might have had lower risk tolerance than those who worked in smaller enterprises.
+I confirmed that the sensitivities to the aggregate shocks are statistically similar.
+[^50] Online Appendix B.10 summarizes the finer details of the analysis.
+In short, the heads’ earnings are regressed on the aggregate consumption, its interaction term with respect to the indicator variable for the large-scale and governmental factories, family size variable, and household fixed effects.
 The results are unchanged if I trim a few semi-months with a smaller number of cross-sectional observations.
-This result is unchanged if I assume that the heads who worked in the smithing factory or in the non-machinery sectors had different risk preferences.
-[^55] The skilled workers in the smithing sector might be a business owner (Online Appendix A.6).
-Similarly, the workers in the non-machinery sectors in Tsukishima, a representative machinery production area, might have had a different, presumably higher, risk-tolerance.
-This supports the evidence that conditional on the household fixed effects, it is plausible to assume that the THBS households shared similar risk preferences.
+This result is unchanged if I assume that the heads who worked in the smithing factory or in the non-machinery manufacturing sectors had different risk preferences.
+[^51] The skilled workers in the smithing sector could be business owners (Online Appendix A.6).
+Similarly, the workers in the non-machinery manufacturing sectors in Tsukishima—a representative machinery production area—could have a different, presumably higher, risk tolerance.
+This shows that it is plausible to assume that the THBS households shared similar risk preferences, conditional on the household fixed effects.
 4 Empirical Analysis
@@
 Estimation Strategy
-I consider a linear fixed-effects model designed to test the consumption smoothing.
-[^56] Another empirical specification used to test the consumption smoothing is the first-difference model, including the aggregate consumption to capture the macroeconomic shocks in the economy (Mace 1991).
-In other words, this alternative model requires the interpretation of the estimated coefficient on the aggregate measure of consumption.
-Although the consumption of the THBS households tends to capture the macroeconomic trend in the consumption (Section 3.3), those households do not cover entire households in the target economy, Tsukishima.
-This disturbs the interpretation of the coefficient on aggregate consumption.
-To be conservative, therefore, I use the two-way fixed-effects model to control for the macroeconomic shocks (Cochrane 1991; Ravallion and Chaudhuri 1997).
+I consider a linear fixed-effects model to test consumption smoothing.
+[^52] Another empirical specification used to test consumption smoothing is the first-difference model, including the aggregate consumption to capture macroeconomic shocks in the economy (Mace 1991).
+This alternative model requires the interpretation of the estimated coefficient on the aggregate measure of consumption.
+However, the THBS households do not include entire households in the target economy—Tsukishima.
+This may disturb the interpretation of the coefficient on aggregate consumption.
+To be conservative, therefore, I use the two-way fixed-effects model to control for macroeconomic shocks (Cochrane 1991; Ravallion and Chaudhuri 1997).
 Online Appendix C.1 provides the conceptual framework and the derivation of my empirical specification.
@@
 The estimate of $\gamma$ shows income elasticity, ranging from zero for perfect insurance to one for the absence of insurance.
-[^57] This is comparable to the estimate from the same conceptual framework: this comparability is important in the consumption smoothing literature.
-Note that relevant previous studies have not integrated the recent theoretical discussion on the fixed effect model with two-way error components under the difference-in-differences setting (TWFE-DID) yet.
-In fact, the bias-corrected estimator of the TWFE-DID with continuous treatment for multiple periods is underdeveloped yet (see de Chaisemartin and D’Haultfœuille (2023) for a comprehensive review).
-The cluster-robust variance-covariance matrix estimator is used to deal with heteroskedasticity and serial dependency (Arellano 1987).
+[^53] This is comparable to the estimate from the same conceptual framework: this comparability is important in the consumption-smoothing literature.
+Note that relevant previous studies have not integrated the recent theoretical discussion on the fixed effect model with two-way error components under the difference-in-differences (TWFE-DID) setting yet.
+In fact, the bias-corrected estimator for the TWFE-DID model with continuous treatment for multiple periods is underdeveloped (see de Chaisemartin and D’Haultfœuille (2023) for a comprehensive review).
+The cluster-robust variance-covariance matrix estimator is used to manage heteroskedasticity and serial dependency (Arellano 1987).
 Results
@@
       Transportation 0.278 [0.110]** 187 0.605 [0.264]** 97 0.362 [0.291] 115
-      Gift 0.523 [0.106]*** 234 0.469 [0.197]** 113 0.622 [0.306]* 130
+      Gifts 0.523 [0.106]*** 234 0.469 [0.197]** 113 0.622 [0.306]* 130
       Miscellaneous -0.076 [0.147] 214 0.522 [0.252]** 110 0.246 [0.267] 126
 ***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors in brackets are clustered at the household level.
-Notes: This table shows the results of equation 1: the regressions of the 11 measures of log-transformed consumption on log-transformed disposable income as well as on the family size control, household-fixed effect, and time-fixed effect.
+Notes: This table shows the results of Equation 1.
+The regressions of the 11 measures of log-transformed consumption on log-transformed disposable income, the family size control, household-fixed effect, and time-fixed effect are listed.
 Columns 1, 2, and 3 show the results for the semi-monthly, adjusted monthly, and unadjusted monthly panel datasets, respectively.
@@
 An important finding is that the estimate for food categories is smaller ($0.136$) but highly statistically significant.
-This implies that while the households tended to reduce their food expenses when they faced adverse income shocks, they might have had some means to deal with those shocks to smooth food consumption.
+This implies that while the households tended to reduce their food expenses when they faced adverse income shocks, they might have had some means to manage those shocks to smooth food consumption.
 The estimates for the entertainment and transportation expenses ($0.307$ and $0.278$) are also highly statistically significant.
-This may make sense given that those expenses include family travel and theatergoing.
-[^58] The workers can enter their factories by walking in Tsukishima.
-Thus, the costs of commuting to the factory are not usually documented in the budget books.
-Similarly, the estimate for gifts shows a large magnitude ($0.523$), reflecting that the expenses on the others are essentially optional.
+This may be understandable because those expenses include family travel and theatergoing.
+[^54] The workers could walk to the factories in Tsukishima, so the costs of commuting were not usually documented in the budget books.
+Similarly, the estimate for gifts shows a large amount ($0.523$), reflecting that the expenses on the others were optional.
 The results for the other subcategories are statistically insignificant or unclear.
 The estimates for furniture and clothes are weakly statistically significant.
-Although less than conclusive, this may reflect that both categories are luxury categories.
+Although not conclusive, this may reflect that both categories were luxury categories.
 The estimate for the education category is small but also weakly statistically significant ($0.101$).
-This may be because it includes the allowance for children, which could be easily reduced.
+This may be because it includes an allowance for children, which could be easily decreased.
 [^55] Toyotaro Miyoshi served as a research assistant at the Tsukishima Survey while he was a student at the Tokyo Imperial University.
@@
 383–384; 386; 387).
-I present the results for the adjusted monthly panel dataset in Column (2) to see whether the income elasticities changed in a longer span.
+I present the results for the adjusted monthly panel dataset in Column (2) to see whether the income elasticities changed over a longer span.
 Not surprisingly, the adjusted monthly series yields greater estimates than the semi-monthly series.
-The estimate for total consumption is now $0.433$, which is roughly $1.2$ times greater than that from the semi-monthly series.
-One potential explanation behind this difference is that the available sources of temporary income have been limited, and those sources are not useful for compensating for adverse income shocks lasting for at least one month.
-The most noticeable example is the food category, which comprises the largest portion of the expenditure (Panel A of Table 3).
-The estimate is $0.293$, more than two times greater than the estimate for the semi-monthly series ($0.136$).
-This implies that the households’ food expenditures are likely to be insured in a few weeks but not over one month.
-As I will delve into the mechanism behind this behavior in the following subsections, credit purchases may work to smooth food consumption in the short run.
-Similarly, the expenses for entertainment and transportation are much more sensitive ($0.716$ and $0.605$) to the shocks in the adjusted monthly series.
-The estimate for gifts is slightly smaller but still in a similar range.
+The estimate for total consumption was $0.433$, which is approximately $1.2$ times greater than that from the semi-monthly series.
+One potential explanation for this difference is that the available sources of temporary income might have been limited, and those sources were not useful for compensating for adverse income shocks that lasted for at least one month.
+The most noticeable example is the food category, which comprised the largest portion of the expenditure (Panel A of Table 3).
+The estimate is $0.293$, more than twice the estimate for the semi-monthly series ($0.136$).
+This implies that the households’ food expenditures were likely to be insured for a few weeks but not over one month.
+I will investigate the mechanism underlying this behavior in the following subsections, and credit purchases may be used to smooth food consumption in the short run.
+Similarly, the expenses for entertainment and transportation were much more sensitive ($0.716$ and $0.605$) to the shocks in the adjusted monthly series.
+The estimate for gifts is slightly smaller but in a similar range.
 In short, while the household could compensate for their income losses for a few weeks, they could not do so for several weeks, presumably because of borrowing constraints and/or a limited amount of liquidity assets.
-The expenses on those unnecessary categories might particularly be easily cut in response to relatively persistent shocks.
-In contrast, a few subcategories are not sensitive to the idiosyncratic shocks over different time-series frequencies.
+The expenses on those unnecessary categories might particularly be cut in response to relatively persistent shocks.
+In contrast, a few subcategories were not sensitive to the idiosyncratic shocks over different time-series frequencies.
 For example, the estimates for the housing, utilities, and medical categories are not statistically significant in both semi-monthly and adjusted monthly series.
-This may reflect that the rent and utility payments were relatively fixed.
-[^60] For example, rent payment is usually fixed and paid once per month.
+This may reflect the fact that the rent and utility payments were relatively fixed.
+[^56] For example, rent payment was usually fixed and paid once per month.
 Online Appendix C.2 shows that the housing category should be analyzed based on the adjusted monthly series rather than the semi-monthly series.
 The medical expenses included the fee for bathing (yuya).
-People had taken both every few days at that time, which was less likely to be cut, albeit hairdressing for women (kamiyui) in the same category may be likely to be cut.
+People took baths every few days at that time, which was less likely to be cut, although hairdressing for women (kamiyui) in the same category was likely to be cut.
 The estimates for the furniture, clothes, and education categories are statistically insignificant in the adjusted monthly series.
-While this could partly be the efficiency issue under a smaller sample size, those expenses may be sensitive to unexpected shocks in consumption, like unanticipated losses of household goods, clothes, and stationaries including writing papers.
+While this could partly be the efficiency issue under a smaller sample size, those expenses may be sensitive to unexpected shocks in consumption, such as unanticipated losses of household goods, clothes, and stationery, including writing paper.
 [^57] An alternative explanation for the clothes category relates to seasonality.
-The workers had bought their work clothes for the new season (for example, the OISR, Archives of the Tsukishima Survey (THBS #40, 31st March 1919)).
-Since those clothes are rather necessary items, this sort of seasonal expense may attenuate the estimate under low frequencey series.
-This systematically attenuates the estimates for these specific categories in the adjusted monthly series.
-[^62] Generally, the panels with short-run time bins are more likely to be influenced by unexpected shocks in consumption, which tends to attenuate the income elasticity (Nelson 1994).
-If this mechanism worked, subcategories that are sensitive (insensitive) to unexpected shocks in consumption should be more (less) attenuated.
-In my result, however, the estimates for the adjusted monthly panels are greater than those for the semi-monthly panels across a large part of subcategories.
-Thus, on average, my result indicates that households might face more difficulties in smoothing their consumption in the longer run.
+The workers had bought their work clothes for the new season (e.g., the OISR, Archives of the Tsukishima Survey (THBS #40, March 31, 1919)).
+As those clothes were necessary items, those seasonal expenses may attenuate the estimate.
+This may systematically attenuate the estimates for the specific categories in the adjusted monthly series.
+[^58] Generally, the panels with short-run time bins were more likely to be influenced by unexpected consumption shocks, which attenuated income elasticity (Nelson 1994).
+If this mechanism worked, subcategories that were sensitive (insensitive) to unexpected consumption shocks should be more (less) attenuated.
+In my result, however, the estimates for the adjusted monthly panels are greater than those for the semi-monthly panels in a large part of subcategories, suggesting that households faced more difficulties with smoothing their consumption in the long run.
 Finally, I assess the validity of the aggregation method used in this study.
@@
 For example, the estimate for the food category is now $0.10$ and statistically insignificant.
-Despite this, several subcategories take values close to one or large negative values, which are no longer economically interpretable.
-[^63] For example, the estimate for housing is now negative.
+This means that measurement errors in the unadjusted aggregation lead to strong attenuation bias, which results in false negatives (i.e., false rejections of the null of full risk-sharing hypothesis) in most of the consumption categories.
+I, therefore, specified a generating process of systematic measurement errors in the aggregate survey data, which has attracted attention in the macroeconomics literature (Cochrane 1991; Nelson 1994; Dynarski et al.
+1997; Gervais and Klein 2010).
+Moreover, several subcategories select values close to one or large negative values, which are no longer economically interpretable.
+[^59] For example, the estimate for housing was negative.
 The households paid a fixed rent amount monthly.
-Given that the rent payday is fixed, for example, the shifts in wage paydays might create cells with less income with a fixed amount of rent.
+Given that the rent payday was fixed, the shifts in wage paydays might create cells with less income with a fixed amount of rent.
 In the case of seasonal consumption categories, the shifts may coincide with the timing of seasonal spending.
@@
 They also spent more on clothing in June, likely to buy summer clothes.
-The shift of paydays to these particular months may have created a strong but false positive correlation.
-This means that measurement errors in the unadjusted aggregation lead to strong attenuation bias, which results in false negatives in most of the consumption categories and may cause critical issues in statistical inference for a few subcategories.
+The shift of paydays to these months may have created a strong but false positive correlation.
+This means that the miss-assignment of paydays in the unadjusted data may cause critical statistical inference issues in a few subcategories.
 4.2 Risk-coping Mechanisms
-Preceding results show that the skilled workers’ households in Tsukishima were less likely to smooth their consumption.
-The estimated elasticities, however, become smaller when I focus on the semi-monthly frequency.
-Specifically, food expenditure is likely to be insured within at least half a month.
-The workers might have dealt with their shorter-run shocks relatively well but could not cope with the longer-run shocks.
-In this section, I test the mechanisms behind their smoothing behavior using several variables on the risk-coping devices.
+The preceding results show that the skilled workers’ households in Tsukishima were less likely to smooth their consumption.
+The estimated elasticities, however, became smaller when I focused on the semi-monthly frequency.
+Specifically, food expenditure was likely to be insured within at least half a month.
+The workers might have coped with their shorter-run shocks relatively well but could not manage longer-run shocks.
+In this section, I test the mechanisms behind their smoothing behavior using several variables in the risk-coping strategies.
 Estimation Strategy
-To test the roles of risk-coping strategies, I employ the empirical specification suggested by Fafchamps and Lund (2003). For household $i$ at time $t$, the specification is characterized as follows:
+To test the roles of risk-coping strategies, I employed the empirical specification suggested by Fafchamps and Lund (2003).
+For household $i$ at time $t$, the specification is characterized as follows:
 $r_{i,t}=\kappa\tilde{y}_{i,t}+\eta x_{i,t}+\nu_{i}+\zeta_{t}+\epsilon_{i,t},$ (2)
@@
 Online Appendix C.1 summarizes the derivation of this specification.
-I prefer to use the head’s income instead of the disposable income in equation 2 to avoid the potential endogeneity.
+I used the head’s income instead of the disposable income in Equation 2 to avoid the potential endogeneity.
 [^60] For instance, the household’s total income is the outcome of the intra-household responses to the shocks, such as the labor-supply adjustments.
-I confirmed that the results are largely unchanged if I use the disposable income as the idiosyncratic income shock variable instead of the head’s income (not reported).
+I confirmed that the results were largely unchanged if I used the disposable income as the idiosyncratic income shock variable instead of the head’s income (not reported).
 This is consistent with the fact that the labor-supply adjustments did not occur among the THBS households (Table 5).
-Despite this, I prefer to use the conservative setting to deal with potential biases and to set a comparable setting to the structural equation used in Section 4.3.
+Despite this, I preferred to use the conservative setting to manage potential biases and to set a comparable setting to the structural equation used in Section 4.3.
 See Asdrubali et al.
 (2020) for a recent study using the head’s income as the shock variable in testing the full risk-sharing.
-As for $r_{i,t}$, I consider several variables to overview various risk-coping strategies.
-These include the net income on savings, insurance, borrowing, credit purchases, and gifts.
-Each net variable is defined as the difference between income and expenses.
-The estimates of $\kappa$ are therefore expected to be negative since net income becomes negative in good times but positive in unfavorable times.
-I also consider the earnings of all the family members except for the head and the sales of miscellaneous goods because both intra-household labor supply adjustments and asset sales may play a role in mitigating the idiosyncratic shocks as alternative risk-coping strategies (Udry 1995; Blundell et al.
+Regarding $r_{i,t}$, I considered several variables to overview various risk-coping strategies.
+These included the net income on savings, insurance, borrowing, credit purchases, and gifts.
+Each net variable was defined as the difference between income and expenses.
+The estimates of $\kappa$ were therefore expected to be negative since net income becomes negative in good times but positive in unfavorable times.
+I also considered the earnings of the family members (except for the head) and the sales of miscellaneous goods because both intra-household labor supply adjustments and asset sales might alleviate idiosyncratic shocks as alternative risk-mitigation strategies (Udry 1995; Blundell et al.
 2016).
 [^61] Rosenzweig and Stark (1989) show that marriage cum migration in South Indian villages contributed to reducing the variability of household food consumption.
-Since the THBS focuses on the households living in Tsukishima throughout the sample period, I can discard addressing the ex-ante risk-coping strategy through migration.
+Since the THBS focused on the households living in Tsukishima throughout the sample period, I did not need to address the ex-ante risk-mitigation strategy through migration.
 The estimated coefficients for the sales and labor supply adjustment variables may be negative, reflecting additional earnings when the head’s income is reduced.
@@
 Other member’s earnings 0.001 [0.010] 289 0.036 [0.033] 124
-Sales of assets -0.000 [0.000] 289 0.000 [0.001] 124
+Sales of miscellaneous assets -0.000 [0.000] 289 0.000 [0.001] 124
 ***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors in brackets are clustered at the household level.
 Notes:
-This table shows the results of equation 2: the family size control, household fixed effects, and time-fixed effects are included in all the regressions.
+This table shows the results of equation 2.
+The family size control, household fixed effects, and time-fixed effects are included in all the regressions.
 Columns 1 and 2 show the estimates for the semi-monthly and adjusted monthly panel datasets, respectively.
@@
 Panel B: “Other member’s earnings” refers to the total income earned by all family members except for the head of the household.
-“Sales of assets” includes the sale of daily miscellaneous goods such as newspapers and empty bottles.
+“Sales of miscellaneous assets” includes the sale of daily miscellaneous goods such as newspapers and empty bottles.
 All dependent variables are in yen.
@@
 Columns (1) and (2) summarize the results for the semi-monthly and adjusted monthly panel datasets, respectively.
-Panel A shows the results for the net income variables, whereas Panel B shows those for the labor supply adjustments and sales of assets.
+Panel A shows the results for the net income variables, whereas Panel B shows those for the labor supply adjustments and sales of small assets.
 The estimated coefficient for the net savings is negative and statistically significant in Column (1).
-The estimate suggests that a one yen decrease (increase) in the head’s earnings increases (decreases) the net income from savings by $0.043$ yen.
-The estimate from the monthly series is similar ($0.042$), albeit weakly statistically significant (Column (2)).
-This is basically consistent with the historical fact that many workers had savings accounts both in postal savings and savings banks at that time (Section 2.3).
-However, the estimated magnitude is small, suggesting that the withdrawals from savings did not compensate for the earnings losses.
+The estimate suggests that a one-yen decrease (increase) in the head’s earnings increases (decreases) the net income from savings by $0.043$ yen.
+The estimate from the adjusted monthly series is similar ($0.035$), albeit statistically insignificant (Column (2)).
+This is consistent with the historical fact that many workers had savings accounts in postal savings and savings banks at that time (Section 2.3).
+However, the estimated magnitude was small, suggesting that the withdrawals from savings did not compensate for the earnings losses.
 The estimates for the insurance category are close to zero and statistically insignificant in both columns.
-The budget books indicate that the primary source in the category is the mutual loan associations (mujin).
-As explained in Section 2.2, these associations’ rules were unsuitable for compensating the idiosyncratic shocks.
-A suggestive case is that a THBS household won a lottery and got $100$ yen in a monthly cell, which is clearly independent of the worker’s earnings.
-[^66] The OISR, Archives of the Tsukishima Survey (THBS #5, 17th May 1919).
+The budget books indicate that the primary source in this category was the mutual loan associations (mujin).
+As explained in Section 2.2, these associations’ rules were unsuitable for compensating for the idiosyncratic shocks.
+For example, a THBS household won a lottery and received $100$ yen in a monthly cell, which is independent of the worker’s earnings.
+[^62] The OISR, Archives of the Tsukishima Survey (THBS #5, May 17, 1919).
 The bidding rule may be used to support a specific member in the association if members have a common belief in the aid.
-However, most of the incomes from mujin measured in the budget books are less than one yen, which is not the case on average.
+However, most of the income from mujin measured in the budget books was less than one yen, which is not the case on average.
 For the net borrowing category, the estimated coefficients are close to zero and statistically insignificant in both columns.
-Section 2.2 shows that pawnshop was the popular lending institution for low-income households.
-Although THBS households occasionally used pawnshops, they never used them frequently.
+Section 2.2 shows that pawnshops were a popular lending institution for low-income households.
+Although THBS households occasionally used pawnshops, this was not frequent.
 [^63] The OISR, Archives of the Tsukishima Survey (THBS #2; 19; 22).
-This makes sense because skilled factory workers can not classified as the low-income working class.
-The estimated coefficient for the net income from credit purchases is negative and statistically significant in Column (1).
-The estimate suggests that a one yen decrease (increase) in the head’s earnings increases (decreases) the net income from credit purchases by $0.13$ yen.
-This magnitude is larger among the other categories.
-According to the THBS budget books, the credit purchases were mostly used to buy rice, fish, vegetables, and fuels (firewood and charcoal).
-Since rice, a staple food, is usually an expensive item, credit purchases might have played an important role among these households.
-The credit purchases were occasionally used in the call doctor and pharmacy.
+This is understandable because skilled factory workers were not classified as low-income working class at that time.
+The estimated coefficient for net income from credit purchases is negative and strongly statistically significant in Column (1).
+The estimate suggests that a one-yen decrease (increase) in the head’s earnings increases (decreases) the net income from credit purchases by $0.13$ yen.
+This magnitude is the largest in all the categories.
+According to the THBS budget books, the credit purchases were mainly used to buy rice, fish, vegetables, and fuels (firewood and charcoal).
+As rice—a staple food—is usually an expensive item, credit purchases might have played an important role among these households.
+In addition, credit purchase was occasionally used for the call doctor and pharmacy.
 This is consistent with my baseline result for medical expenses in the semi-monthly series (Column (1) in Table 4).
-While this provides suggestive evidence that credit purchase is the primary risk-coping strategy, Column (2) shows that the credit was less likely to work in the longer run.
-This seems to be reasonable because borrowers could make monthly or semi-monthly payments but did not borrow for more than a month (Section 5).
-Instead, the result for the net savings category suggests that precautionary savings have helped mitigate income declines lasting at least one month due to idiosyncratic shocks.
-Given the magnitudes, however, it is important to note that it could only partially compensate for the idiosyncratic shocks.
-The estimates for the net income from gifts are negative and statistically significant in Column (1).
-The budget books tell that the gifts included seasonal gifts, celebration money, and return gifts.
+While these results suggest that credit purchase was the primary risk-mitigation strategy, Column (2) shows that credit was no longer effective in the long term.
+This is reasonable because borrowers could make monthly or semi-monthly payments but did not borrow for more than a month (Section 5).
+Given the estimated magnitude ($0.13$), however, it is noteworthy that the credit could only partially compensate for the idiosyncratic shocks.
+The estimates for the gifts category are negative and statistically significant in Column (1).
+The budget books show that the gifts included seasonal gifts, celebration money, and return gifts.
 Among them, the return gifts can be regarded as mutual aid from personal networks, which may be reflected in this result.
@@
 58-59).
-In my adjusted monthly panels, this share is calculated to be roughly $4$%.
+In my adjusted monthly panels, this share is calculated to be approximately $4$%.
 In this light, while the gifts comprised a few percentage points of the monthly income, they may not have provided sufficient aid for the households facing economic hardship.
@@
 Finally, Panel B tests the roles of the intra-household labor supply adjustment and sales of miscellaneous assets.
-The estimated coefficients in both columns are close to zero and statistically insignificant.
+The estimated coefficients in both columns of the first row are close to zero and statistically insignificant.
 Miyoshi (1989, pp.
-383–386) documented that the wives regularly engaged in side jobs rather than worked temporarily to compensate for the income loss.
+383–386) documented that the wives regularly engaged in side jobs rather than working temporarily to compensate for the income loss.
 In fact, although sewing was the most popular side job among THBS wives, their earnings were too small to replace the lost income from their breadwinning husbands.
-[^68] The OISR, Archives of the Tsukishima Survey (THBS#1, 13th January 1919).
-In the final row in Panel B, I find that selling miscellaneous assets such as newspapers, bottles, and clocks did not work as a risk-coping strategy.
-To summarize, I find that credit purchases have played a role in mitigating the idiosyncratic shocks.
+[^64] The OISR, Archives of the Tsukishima Survey (THBS#1, January 13, 1919).
+In the second row, I show that selling miscellaneous assets such as newspapers, bottles, and clocks was not an effective risk-mitigation strategy.
+To summarize, the evidence suggests that credit purchases mitigated idiosyncratic shocks.
 The results also indicate that the savings institutions and personal networks might have been used as risk-coping institutions among the workers.
-However, these were basically available for short-term, say at most half a month.
-The estimated magnitudes indicate that the workers could only partially compensate for the temporary losses of income by using these institutions, which is indeed consistent with the findings on the partial risk-sharing in Section 4.1.
+However, these were available for the short term —at most half a month.
+The estimated magnitudes indicate that the workers could only partially compensate for the temporary losses of income by using these institutions.
+This supports the findings on partial risk-sharing in Section 4.1.
 Overall, the results are aligned with the descriptives from the historical documents on the financial institutions available in Tokyo City (Section 2.3).
-An interesting finding, which is not explicitly described in the historical sources, is the possibility that the credit purchases in the local society and the informal gifts from personal networks might have provided some aid to the factory workers.
-4.3 Event-Study Analysis: Response to the Advers Health Shocks
-The foregoing analyses used the within variations in the earnings for estimating the income elasticity, meaning that the identification depends both on the adverse and favorable idiosyncratic shocks.
-While it is useful to illustrate the overall responses of the households to the shocks, it does not offer a clear view of the pathways behind the risk-coping behaviors against the adverse idiosyncratic shocks.
-To clarify this point, I next focus on the temporary illness of the heads that cause temporary income losses.
+An interesting finding—which is not explicitly described in the historical sources—is that the credit purchases in the local society and the informal gifts from personal networks might have provided some aid to the factory workers.
+4.3 Event-Study Analysis: Response to Adverse Health Shocks
+The foregoing analyses used the within variations in the earnings to estimate income elasticity, meaning that the identification depends both on the adverse and favorable idiosyncratic shocks.
+While it is useful to illustrate the overall responses of the households to the shocks, it does not reveal the pathways behind the risk-coping strategies.
+Next, I focus on temporary illness of household heads that could cause temporary income losses.
 Estimation Strategy
@@
 Source: Created by the author using the 1918, 1919, and 1920 editions of the Statistics of Causes of Deaths of the Empire Japan (Statistics Bureau of the Cabinet 1921; 1922; 1923).
-I begin my analysis by testing whether the illness of the head impacted the earnings.
+I began my analysis by testing whether the illness of the head impacted the earnings.
 To do so, I decided the exact timing and duration of illnesses of the family members using the Kazoku idō hyō (Section 3.4).
-Figure 5 shows the number of monthly death counts from influenza (ryūkōsei kanbō) in Tokyo between 1918 and 1920.
+Figure 5 shows the number of monthly deaths from influenza (ryūkōsei kanbō) in Tokyo between 1918 and 1920.
 There appears to have been a moderate spike in early 1919, which is partially captured in the THBS period.
-In fact, a few THBS budget books suggest that the family members had caught the flu during this period.
-Five households experienced the illness during the sample period, and the illness affecting the heads was concentrated between February and May 1919, when flu exposure was most prevalent.
+In fact, a few THBS budget books indicate that the family members contracted influenza during this period.
+Five households experienced the illness during the sample period, and illness affecting the heads was concentrated between February and May 1919, when influenza exposure was most prevalent.
 [^65] Among four cells with health shocks on the heads, two are observed in February, during the peak of the epidemics (Figure 5), and the others are in March and May.
@@
 Dev.
-$=3.6$), which is consistent with the fact that Type A influenza causes fever for less than $6$ days (Uyeki et al.
+$=3.6$), which is consistent with the fact that Type A influenza caused fever for less than six days (Uyeki et al.
 2022).
-Therefore, it is plausible to see that the influenza epidemics largely led to illnesses during this period, providing a quasi-experimental setting in analyzing to what extent the adverse idiosyncratic shock driven by illness influences the households’ consumption behaviors.
+Therefore, it is plausible that influenza was responsible for illnesses during this period, providing a quasi-experimental setting for analyzing to what extent the adverse idiosyncratic shock of illness influences the households’ consumption strategies.
 Earnings in a semi-month are based on the work done during the previous semi-month.
-This means that the head’s earnings shall mainly be influenced by the number of days suffering illness in the previous semi-month cell.
+This means that the head’s earnings were influenced by the number of illness days in the previous semi-month cell.
 Table 6 presents the results from the regression of the head’s earnings on the number of illness days in the previous semi-monthly cell.
-While the instantaneous health shock has little impact on the head’s earnings (Column 1), the lagged health shock has a statistically significant effect on his earnings (Column 2).
+While the instantaneous health shock has little impact on the head’s earnings (Column 1), the lagged health shock has a statistically significant effect on their earnings (Column 2).
 This result is unchanged if I include both variables in the same regression in Column 3, supporting the evidence that the health shocks in the previous cell reduced the earnings of the head.
-The estimate in Column (2) suggests that a one-day increase in the illness decreases the head’s semi-monthly income by roughly $2.4$ yen.
+The estimate in Column (2) suggests that a one-day increase in illness decreases the head’s semi-monthly income by approximately $2.4$ yen.
 Notably, this is close to the average daily wage rate of skilled factory workers (Online Appendix B.6).
-Thus, this reflects the reductions in their salary due to the absence: Miyoshi (1989, pp.
-383; 388) highlighted that income was accidentally reduced due to absence when the head was infected with a disease.
+This reflects the reductions in their salary due to absence: Miyoshi (1989, pp.
+383; 388) highlighted that income was reduced due to absence when the head was infected with a disease.
 Finally, Column (4) adds the health shocks of the other family members into the specification of Column (3).
@@
 These results confirm that the head who suffered illness lost earnings.
-[^70] Given that the average number of days suffering the illness is roughly $2.0$, the temporary health shock loses more than $5$ yen in that month, accounting for roughly $10$% of the average monthly income.
-Although this is a rough estimate and the losses shall depend on the severity of the shocks, my result at least shows that the household income had reacted to the adverse health shock among the factory workers’ households.
+[^66] Given that the average number of illness days was approximately $2.0$, the temporary health shock losses were more than five yen in that month, accounting for approximately $10$% of the average monthly income.
+Although this is a rough estimate and the losses depend on the severity of the shocks, my result shows that the household income responded to the adverse health shock of illness among the factory workers’ households.
 Table 6: Results of Testing the Impacts of Adverse Health Shocks
@@
 “[lag]” indicates the lagged variable (i.e., the shocks measured in the previous semi-month cell).
-Although correlation coefficient between the level and lag of the health shock variables for the other family members is substantially high (Pearson’s coefficient $=0.99$), including eather of the two variables does not change the finding on the neglisible impacts of the health shocks on the other family members.
-Following the evidence, I characterize equation 1 as a structural consumption smoothing equation by using the first-stage reduced form equation as follows:
+Although the correlation coefficient between the level and lag of the health shock variables for the other family members is substantially high (Pearson’s coefficient $=0.99$), including either of the two variables does not change the finding on the negligible impacts of the health shocks on the other family members.
+Following the evidence, I characterize Equation 1 as a structural consumption-smoothing equation by using the first-stage reduced form equation as follows:
 $\log y_{i,t}=\beta h_{i,t-1}+\psi x_{i,t}+\varpi_{i}+\tau_{t}+e_{i,t},$ (3)
-where $h_{i,t-1}$ is the number of days of head’s illness in the previous semi-month cell, $\varpi_{i}$ is the household fixed effect, $\tau_{t}$ is the semi-month fixed effect, and $e_{i,t}$ is a random error term.
+where $h_{i,t-1}$ is the number of days of the head’s illness in the previous semi-month cell, $\varpi_{i}$ is the household fixed effect, $\tau_{t}$ is the semi-month fixed effect, and $e_{i,t}$ is a random error term.
 For the risk-sharing mechanisms analysis under equation 2, the log-transformed disposable income in equation 3 ($y_{i,t}$) is replaced with the head’s earnings ($\tilde{y}_{i,t}$).
 The health shock variable is plausibly exogenous because the illnesses were from the influenza epidemics, as explained above.
-[^71] As noted in Section 2.3, the factories did not offer compensation for their sickness leave because the Factory Act did not specify benefits for workers’ illness (Study Group of the Factory Act 1916).
-Even in the largest factory in Tsukishima, the Ishikawajim Shipyard, any descriptions of the influenza epidemics are documented in the company’s history (Arai 1930).
-In fact, Miyoshi (1989) documented the issue that the head’s disease directly causes income loss due to the absence.
+[^67] As noted in Section 2.3, the factories did not offer compensation for sick leave because the Factory Act did not specify benefits for workers’ illness (Study Group of the Factory Act 1916).
+Even in the largest factory in Tsukishima, the Ishikawajima Shipyard, any descriptions of the influenza epidemics are documented in the company’s history (Arai 1930).
+In fact, Miyoshi (1989) documented the issue that the head’s disease directly caused income loss due to the absence at that time.
 In addition, there is a correlation between the health shock variable and the head’s earnings at a statistically valid level (Table 6).
-Both support the evidence that the exogeneity of the instrument and relevance condition basically hold under this just-identified system.
+Both support the evidence that the exogeneity of the instrument and relevance condition hold under this just-identified system.
 Results: To What Extent do Illnesses Affect Consumption?
+I first consider a reduced-form specification using the head’s illness as the exposure variable in equation 1 following Cochrane (1991).
+Column (1) in Table 7 shows the results.
+Overall, the coefficients on the exposure variable are negative, supporting evidence that adverse health shocks on the heads reduce consumption levels.
+Next, I explicitly model the structure underlying the relationship between the health shocks and earnings using the reduced-form equation of (3).
+If the negative impacts of the health shocks on consumption were derived from the income losses, the results from the structural estimation approach correspond to those in Column (1).
+The estimates from the structural equation are listed in Column (2) of Table 7.
+As shown, the results are congruent with the reduced-form results in Column (1).
+Importantly, the estimates are systematically larger than those reported in Table 4.
+The estimates for total and food consumption categories are $0.543$ and $0.355$, respectively, whereas the baseline estimates are $0.358$ and $0.136$.
+While this result partly reflects the bias reduction under the instrumental variable estimation, it is also noteworthy that this system is designed to reduce the bias by specifying the channel from adverse (instead of favorable) income shocks.
+For example, expenses on the allowance for children could be easily cut during unfavorable times but may not respond to favorable shocks.
+Similarly, the expenses for furniture are likely to be cut, but the households might not buy those frequently, meaning that there are no sufficient cases for reacting to rare health shock events.
+[^68] In other words, furniture may have only been purchased when the households had extra income.
+This may partly be reflected in the smaller value of the first-stage $F$-statistics in the furniture subcategory.
+[^69] Note that the logarithmic transformation leaves the cells with expenses on the furniture.
+Thus, the reduced-form equation captures the relationship between the health shocks and earnings in the cells with any expenses on furniture.
+While the first-stage $F$-statitics is $35.7$ in the linear-linear model, the second-stage result is robust against the functional form assumption.
+This may reflect the asymptotic theoretical advantage in the panel data model under nearly weak instruments case (Staiger and Stock 1997; Cai et al.
+2012; Baltagi et al.
+2012).
+The same argument is applicable to the housing category.
+A similar interpretation can be applied to the transportation category because family travel was not a frequent event and thus, unlikely to coincide with the timing of illness.
+By contrast, the entertainment expenses were still likely to be cut in case of illness, which might reflect the curtailment of tobacco, theater-going, and the allowance for heads.
+This may be another explanation for the difference in the results from my baseline results in Table 4.
+While the estimates for the clothes and gifts categories now exceed one, the null of equality to the population mean of one is not statistically rejected for both cases.
+Although the large estimates may be partly due to the operation under the instrument based on rare health shock events, the findings are materially like my baseline results—both categories are sensitive to the shocks as are those of the dispensable categories.
 Table 7: Results of Estimating Income Elasticities:
-Smoothing Consumption against the Adverse Health Shocks
+Smoothing Consumption against Adverse Health Shocks
 Semi-monthly Panel Dataset
@@
       Transportation -0.100 [0.094] 170 0.436 [0.510] 162 20.5
-      Gift -0.325 [0.093]*** 212 1.567 [0.797]* 204 17.8
+      Gifts -0.325 [0.093]*** 212 1.567 [0.797]* 204 17.8
       Miscellaneous -0.075 [0.027]*** 195 0.343 [0.101]*** 189 17.3
@@
 Notes:
-Column 1 shows the results of equation 1, in which the log-transformed disposable income is replaced with the head’s illness (i.e., the number of days with illness in the previous semi-month cell) as the idiosyncratic shock variable.
-The dependent variables are the 11 measures of log-transformed consumption, and all the regressions include the family size control, household fixed effects, and time fixed effects.
-The estimated coefficients on the health shock variable are listed in the sub-column named “Coef.”.
+Column 1 shows the results of Equation 1, in which the log-transformed disposable income is replaced with the head’s illness (i.e., the number of days with illness in the previous semi-month cell).
+The dependent variables are the 11 measures of log-transformed consumption, and all the regressions include the family size control, household-fixed effects, and time-fixed effects.
+The estimated coefficient on the health shock variable is listed in the sub-column named “Coef.”.
 Column 2 shows the results of equation 1, in which the log-transformed disposable income is regarded as the endogenous variable.
 The head’s illness is an instrumental variable (i.e., excluded exogenous regressor) for the first-stage reduced-form equation (3).
-The estimated coefficients on the disposable income are listed in the sub-column named “Coef.”.
+The estimated coefficient of the disposable income is listed in the sub-column named “Coef.”.
 The first-stage $F$-statistics on the health shock variable are listed in the final column.
-I first consider a reduced-form specification using the head’s illness as the exposure variable in equation 1 in the spirit of Cochrane (1991).
-Column (1) in Table 7 shows the results.
-Overall, the coefficients on the exposure variable are negative, supporting evidence that adverse health shocks on the heads reduce consumption levels.
-Next, I explicitly model the structure behind the relationship between the health shocks and earnings using the reduced-form equation of 3.
-If the negative impacts of the health shocks on consumption were derived from the income losses, the result from the structural estimation approach is similar to that in Column (1).
-The estimates from the structural equation are listed in Column (2) of Table 7.
-As shown, the results are materially similar to the reduced-form results in Column (1).
-Importantly, the estimates are systematically larger than those reported in Table 4.
-The estimates for total and food consumption categories are $0.543$ and $0.355$, respectively, whereas the baseline estimates are $0.358$ and $0.136$.
-While this result partly reflects the bias reduction under the instrumental variable estimation, it is also important to remember that this system is designed to reduce the bias by specifying the channel from the adverse (instead of favorable) income shocks.
-For example, expenses on the allowance for children could be easily cut during unfavorable times but may not respond to favorable shocks.
-Similarly, the expenses for furniture are likely to be cut, but the households might not buy those frequently, meaning that there are no sufficient cases for reacting to the rare health shock events.
-[^72] In other words, the furniture may have been purchased only when the households had extra income.
-This may partly reflected in the smaller value of the first-stage $F$-statistics in the furniture subcategory.
-[^73] Note that the logarithmic transformation leaves the cells with expenses on the furniture.
-Thus, the reduced-form equation captures the relationship between the health shocks and earnings in the cells with any expenses on the furniture.
-While the first-stage $F$-statitics is $35.7$ in the linear-linear model, the second-stage result is robust against the functional form assumption.
-This may reflect the asymptotic theoretical advantage in the panel data model under nearly weak instruments case (Staiger and Stock 1997; Cai et al.
-2012; Baltagi et al.
-2012).
-The same argument is applicable to the housing category.
-A similar interpretation can be applicable to the transportation category because family travel is not a frequent event and thus, unlikely to coincide with the timing of illness.
-In contrast, the entertainment expenses are still likely to be cut in case of illness, which may reflect the curtailment of tobacco, theater-going, and the allowance for heads.
-This may be another explanation for the difference in the results from my baseline results in Table 4.
-While the estimates for clothes and gifts categories now exceed one, the null of equality to the population mean of one is not statistically rejected for both cases.
-Although the large estimates may be partly due to the operation under the instrument based on the rare health shock events, the findings are materially similar to my baseline results: both categories are sensitive to the shocks as those are dispensable categories.
-Results: How did the households compensate for the losses from the ill?
+Results: How did the Households Compensate for Losses from illness?
+I tested the risk-coping strategies for mitigating adverse health shocks in Table 8.
+Column 1 lists the reduced-form results in the same manner as Table 5.
+The estimated coefficients on the net income variables are positive in most cases, indicating the households’ compensating strategies against the shocks.
+The estimate for the credit purchase category is weakly statistically significant, with the largest magnitude.
+It suggests that the one-day increase in the head’s illness increased the net credit purchases by $1.29$ yen, which accounts for approximately half of their average daily wages ($1.29/2.56$).
+Column 2 presents the estimates from the structural equation using the head’s earnings as the endogenous variable.
+A similar but clearer result is obtained—the estimate for the credit purchase category is approximately four times greater than that reported in Column (1) of Table 5.
+It indicates that a one yen decrease in the head’s earnings increases the net income from credit purchases by $0.55$ yen.
+This magnitude is consistent with the direct effect of the health shocks on the net credit purchase estimated in Column (1) of Table 8.707070This means that $2.56\times 0.55=1.4$ yen, which is close to $1.3$ yen in Column (1).
+The estimates for the savings and gifts categories are still negative but now statistically insignificant, supporting the evidence that credit purchase is the primary risk-mitigation strategy for dealing with health shocks.
+Since withdrawing and receiving gifts to use the money at the retailers takes more time (steps), the households likely prefer to use credit purchases.
+Overall, the event-study analysis using adverse health shocks confirms the baseline findings and provides a sharper illustration of consumption-smoothing strategies in the case of adverse idiosyncratic shock events.
+My results show that the unique available response to idiosyncratic income shocks among the THBS households was to use credit purchases.
+The available compensation via credit purchases accounts for approximately half of their temporary income losses, which partially contributes to smoothing their food consumption in the short term.
 Table 8: Results of Testing the Risk-coping Mechanisms:
@@
 Other member’s earnings 0.001 [0.081] 254 -0.001 [0.034] 254
-Sales of assets 0.002 [0.002] 254 -0.001 [0.001] 254
+Sales of miscellaneous assets 0.002 [0.002] 254 -0.001 [0.001] 254
 ***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors in brackets are clustered at the household level.
 Notes:
-Column 1 shows the results of equation 2, in which the head’s earning is replaced with the head’s illness (i.e., the number of days with illness in the previous semi-month cell) as the idiosyncratic shock variable.
+Column 1 shows the results of equation 2, in which the head’s earning is replaced with the head’s illness (i.e., the number of days with illness in the previous semi-month cell).
 The family size control, household fixed effects, and time-fixed effects are included in all the regressions.
-The estimated coefficients on the head’s illness are listed in the subcolumn named “Coef.”.
+The estimated coefficients on the head’s illness are listed in the sub-column named “Coef.”.
 Column 2 shows the results of equation 2, in which the head’s earning is treated as the endogenous variable.
@@
 The first-stage result is identical to the result listed in Column (2) of Table 6 with the first-stage $F$-statistics of $13.6$.
-The estimated coefficients on the head’s earning are listed in the subcolumn named “Coef.”.
+The estimated coefficients on the head’s earnings are listed in the sub-column named “Coef.”.
 Panel A: Each net income variable is defined as the difference between income and expenses.
@@
 Panel B: “Other member’s earnings” refers to the total income earned by all family members except for the head of the household.
-“Sales of assets” includes the sale of daily miscellaneous goods such as newspapers and empty bottles.
+“Sales of miscellaneous assets” includes the sale of daily miscellaneous goods such as newspapers and empty bottles.
 All dependent variables are in yen.
-I test the risk-coping strategies for mitigating adverse health shocks in Table 8.
-Column 1 lists the reduced-form results for the risk-coping strategies in the same manner in Table 5.
-The estimated coefficients on the net income variables are positive in most cases, indicating the households’ compensating responses against the shocks.
-The estimate for the credit purchases category is weakly statistically significant, with the largest magnitude.
-It suggests that the one-day increase in the head’s illness increased the net credit purchases by $1.29$ yen, which accounts for roughly half of their average daily wages ($1.29/2.56$).
-Column 2 presents the estimates from the structural equation using the head’s earnings as the endogenous variable.
-A similar but much clearer result is obtained: the estimate for the credit purchase category is roughly four times greater magnitude than that reported in Column (1) of Table 5.
-The estimate indicates that a one yen decrease in the head’s earnings increases the net income from credit purchases by $0.55$ yen.
-This magnitude is consistent with the direct effect of the health shocks on the net credit purchase estimated in Column (1) of Table 8.747474This means that $2.56\times 0.55=1.4$ yen, which is close to $1.3$ yen in Column (1).
-The estimates for the savings and gifts categories are still negative but now statistically insignificant, supporting the evidence that credit purchase is the primary risk-coping strategy for dealing with health shocks.
-Since it shall take more time (steps) from withdrawing and receiving gifts to use the money at the retailers, it is likely that the households prefer to use credit purchases.
-Overall, the event-study analysis using adverse health shocks confirms baseline findings and provides a sharper illustration of consumption smoothing behaviors in the case of adverse idiosyncratic shock events.
-In short, my results show that the unique available response among the THBS households to the idiosyncratic income shock was to use credit purchases.
-The available compensation via the credit purchases accounts for roughly half of their temporary income losses, which partially contributes to smoothing their food consumption in the short-run.
 5 Credit System in Local Economy
-Foregoing results shed new light on the roles of credit institutions among skilled workers’ households during historical industrialization.
-Since the credit purchases are provided by local retailers, this also suggests the importance of the relationships between the workers and retail shops.
-To the best of my knowledge, however, there is little historical work on credit purchases in prewar Japan.
-Hence, the details of this credit institution are not yet clear: for example, how retailers get the credit information of their customers, whether the contracts are explicit or implicit, and whether the lenders enforce penalties for the debtors.
+The foregoing results enlighten the roles of credit institutions among skilled workers’ households during industrialization.
+As the credit for purchases was provided by local retailers, this suggests the importance of the relationships between the workers and retailers.
+To the best of my knowledge, however, there is little historical research on credit purchases in prewar Japan.
+Hence, the details of this credit institution are not yet clear.
 Table 9: Customers, Sales Methods, and Competitiveness in the Retailers in Tokyo City
@@
 Firewood and charcoal 100 51.1 15.7 50.9 33.4 20.4 78.4 1.3 8.6
-Notes: This table summarizes the statistics on the sales in the $326$ retail shops selling white rice, fish, fruit and vegetables, and firewood and charcoal in Tokyo surveyed in 1935.
-The number of stores reported the share of regular customers (column 2) are $90$, $56$, $61$, and $93$ in the white rice, fish, greengrocer, and fuel (firewood and charcoal) categories, respectively.
-The “share of sales for regular customers” is defined as the percentage of sales for regular customers in the annual total sales in 1935.
-The definition of the social classes (“Upper”; “Middle”; “Other”) is not exactly defined in the report.
-The share of sales method is based on the percentage of sales using each method in the annual total sales in 1935.
+Notes: This table summarizes the statistics on the sales in the $326$ retailers selling white rice, fish, fruit and vegetables, and firewood and charcoal in Tokyo surveyed in 1935.
+The number of stores reported the share of regular customers (column 2) is $90$, $56$, $61$, and $93$ in the white rice, fish, greengrocer, and fuel categories, respectively.
+The “share of sales for regular customers” is defined as the percentage of sales for regular customers in the annual sales.
+The definition of the social classes (“Upper”; “Middle”; “Other”) is not defined in the report.
+The share of sales method is based on the percentage of sales using each method in the annual sales.
 The number of competitors indicates the number of peers within a three-cho, approximately $327$ meters, radius.
@@
 Source: Tokyo City Chamber of Commerce and Industry 1937, pp. legend; 20–21; questionnaire.
-The information on the credit purchase is considerably personal, as in the case of state-contingent loans in the rural credit market (Udry 1994), and can hardly be documented.
-Therefore, systematic information on credit purchases in the prewar period is unavailable.
-Fortunately, however, a few official survey reports documented some information on this payment method.
+The information on the credit purchases is personal, as in the case of state-contingent loans in the rural credit market (Udry 1994) and thus, not well documented.
+Therefore, systematic information on credit purchases is unavailable.
+Fortunately, however, I found that some official survey reports documented information on this credit institution.
 A survey report on retailers in 1930 (hereafter 1930 Survey Report) revealed that credit purchase was a popular payment method in transactions with consumers.
@@
 1).
-Those include not only retailers but also wholesale and other small businesses, and the number of respondents varies depending on the question items.
-The number of respondents in the statistical table on the payment methods is $2,654$, of which $1,197$ are retailers (Tokyo City Office 1932b, p.
+Those included retailers as well as wholesale and other small businesses.
+The number of respondents varied depending on the question items.
+The number of respondents in the statistical table on the payment methods was $2,654$, of which $1,197$ were retailers (Tokyo City Office 1932b, p.
 273).
-To the best of my knowledge, this is the most comprehensive survey of the retailers in Tokyo around 1930.
-For example, 99%(155/157) of grain retailers, 81%(35/43) of fish retailers, 78%(64/82) of greengrocers, and 94%(45/48) of fuel retailers used credit purchases for their selling method (Tokyo City Office 1932b, pp.
+To the best of my knowledge, this is the most comprehensive survey of the retailers in Tokyo circa 1930.
+In fact, $99$%($155/157$) of grain retailers, $81$%($35/43$) of fish retailers, $78$%($64/82$) of greengrocers, and $94$%($45/48$) of fuel retailers used credit purchases for their selling method (Tokyo City Office 1932b, pp.
 260; 263).
-A survey report on retail shops in 1935 published by the Tokyo City Chamber of Commerce and Industry provides more detailed information.
-Table 9 summarizes the information on the sales in those retail shops of rice, fish, fruit and vegetables, and fuels.
-[^76] This survey covers 939 shops and was conducted on 31st December 1935.
-I digitized the statistics for all the shops selling the representative items (rice, fish, fruit and vegetables, and fuels) in the credit purchases.
-The survey instrument includes detailed questions regarding sales and purchases, making it the most comprehensive of the available documents (Tokyo City Chamber of Commerce and Industry 1937, pp.
+A survey of retailers in 1935 published by the Tokyo City Chamber of Commerce and Industry provides more detailed information.
+Table 9 summarizes the information on the retailers’ sales measured in this survey.
+[^72] This survey covered $939$ retailers, and its report was edited on December 31, 1935.
+I digitized the statistics for all the retailers selling the representative items (rice, fish, fruit and vegetables, and fuels).
+The survey instrument included detailed questions regarding sales and purchases (Tokyo City Chamber of Commerce and Industry 1937, pp.
 legend, questionnaire).
-Online Appendix Table C.1 decomposes the statistics listed in Table 9 by the scale of businesses.
+Online Appendix Table C.1 decomposes the statistics listed in Table 9 by scale of businesses.
 I confirm that the main findings are materially similar.
-First, many annual sales are from their regular customers, most of whom are from the non-upper classes.
-For example, $73$% of annual sales in rice retailers are from their regular customers, and more than $80$% of them are from non-upper classes, including the factory workers.
-Second, credit purchase is a popular sales method, especially among rice and fuel retailers: more than $70$% of the sales were done with credit in both shops.
+First, many annual sales were from their regular customers, most of whom were from the non-upper classes.
+For example, $73$% of annual sales of rice retailers were from their regular customers, and more than $80$% of them were from non-upper classes, including factory workers.
+Second, credit purchase was a popular sales method, especially among rice and fuel retailers—more than $70$% of the sales were conducted with credit in both shops.
 [^73] These figures are systematically lower than those from the 1930 Survey Report.
-Given that both surveys show a similar tendency regarding the use of credit purchases, this may be due to the differences in the samples, the definition of the retailers in each category, and the definition of the credit purchase itself.
-Although cash was also used frequently among greengrocers, $34$% of the sales were settled using credit.
-This supports the evidence that credit purchase was a popular payment method, and regular customers settled a large part of it.
-[^78] The 1930 Survey reveals that the retailers had also commonly used credit purchases to buy their source goods from wholesalers (Tokyo City Office 1932a, pp.
+Given that both surveys show a systematic tendency toward credit purchases, this may be due to the differences in the samples, the definition of the retailers in each category, and the definition of the credit purchase itself.
+Although cash was also used among greengrocers, $34$% of the sales were settled using credit.
+This suggests that credit purchase was a popular payment method and that regular customers settled a large part of it.
+[^74] The 1930 Survey reveals that the retailers also commonly used credit purchases to buy their source goods from wholesalers (Tokyo City Office 1932a, pp.
 154–157).
 This means that credit purchase was a familiar payment method for retailers as well as consumers.
-Although these statistics are from the mid-1930s, this confirms the possibility that the retailers were engaged in credit transactions in the context of a long-term relationship with their regular customers.
-[^79] A textbook for retailers published in 1924 explains an interesting mechanism behind this long-term relationship from an aspect of customer psychology: credit purchase is made not only for the convenience of the customer’s budget but also because of the “vanity” of being able to be a regular customer (Shimizu 1924, pp.
+Although these statistics are from the mid-1930s, this confirms the possibility that the retailers were engaged in credit transactions with long-term relationships with their regular customers.
+[^75] A textbook for retailers published in 1924 explains the psychology behind this long-term relationship—credit purchase was provided for the convenience of the customer’s budget but also for the “vanity” (status) of being regular customers (Shimizu 1924, pp.
 115–116).
-The final column of this table further provides suggestive evidence behind this.
-The survey asked retailers whether the sales area was competitive or not, and if yes, they were asked to report the number of competitors within a three-cho (roughly $330$ meters) radius.
-As shown, rice and fuel stores have 2-3 more competitors than the retailers selling fish, fruit, and vegetables.
+The final column of this table further provides suggestive evidence for this.
+The survey asked retailers whether the sales area was competitive or not, and if yes, they were asked to report the number of competitors within a three-cho (approximately $330$ meters) radius.
+As shown, rice and fuel stores have two to three more competitors than the retailers selling fish, fruit, and vegetables.
 Given the relatively high utilization of credit in rice and fuel retailers, they may have responded to competition by allowing their customers to trade on credit.
@@
 by Different Time-series Frequencies
-Note: Figure 6a, and 6c illustrate the time-series plots of the average daily and semi-monthly credit purchase and payment, respectively.
+Notes: Figure 6a, and 6c illustrate the time-series plots of the average daily and semi-monthly credit purchase and payment, respectively.
 The blue and red dashed lines indicate the average credit purchases and payments, respectively.
@@
 The red solid line highlights the zero value.
-Figures 6a and 6b show the daily data between 12th January and 11th November 1919.
-Figures 6c and 6d show the semi-monthly series calculated using the daily data between 12th January and 11th November 1919.
+Figures 6a and 6b show the daily data between January 12 and November 11, 1919.
+Figures 6c and 6d show the semi-monthly series calculated using the daily data between January and November 11, 1919.
 All figures are based on the 29 THBS households that used any credit purchases.
 Source: Created by the author using the THBS sample.
-Next, I look into consumers’ credit purchase behavior using the THBS dataset.
-Most of the THBS households ($29/33$) used any credit purchases during their sample period, supporting the evidence that this was a common practice among factory workers in Tsukishima.
+Next, I examined the credit purchases included in the THBS dataset.
+Most THBS households ($29/33$) used any credit purchases during their sample period, confirming that credit purchase was a common practice among factory workers.
 Figure 6a illustrates the average daily credit purchases and payments.
-Generally, there is a peak per semi-month in payments, which reflects the households’ behavior of making the payments for the credit purchase on or soon after their paydays.
-The purchases series shows a similar tendency but with lower spikes and larger variance.
-This indicates that the households purchased goods by a certain amount around their paydays and then used it when they needed unconservable food such as fish, fruit, and vegetables.
-Figure 6b illustrating the net purchases confirms this behavior: the households used the credit on a daily basis and redeemed it near paydays.
-Figure 6c shows that the payment slightly exceeds purchase, suggesting that many households could more or less repay their loans.
+Generally, there is a peak per semi-month in payments, which reflects the practice of making the payments for the credit purchase on or soon after their paydays.
+The purchase series shows a similar tendency but with lower spikes and larger variance.
+This indicates that the households first purchased goods for a certain amount on or soon after their paydays, and then used purchased perishable foods such as fish, fruit, and vegetables when they needed them.
+Figure 6b illustrating the net purchase confirms this behavior: the households used the credit on a daily basis and redeemed it near paydays.
+Figure 6c shows that the payment slightly exceeded purchase, suggesting that many households could approximately repay their loans.
 Although the difference between payment and purchase is relatively obvious after August 1919, Figure 6d indicates that this difference is not statistically significant in all the semi-months.
@@
 Thus, this is not a systematic trend but may be due to the smaller available number of households in the later sample period (Online Appendix Figure B.4).
-I then analyze whether the volatility of earnings restricts credit purchases.
-Since the credit of a worker is unobservable, I use the volatility of earnings as a proxy of credit risk in the following empirical specification:
+I then analyzed whether the volatility of earnings restricted credit purchases.
+Since credit of worker was unobservable, I used the volatility of earnings as a proxy of credit risk in the following empirical specification:
 $I(\overline{CP}_{i}=0)=\varkappa_{0}+\varkappa_{1}CV_{i}+\bar{\mathbf{x}}_{i}^{\prime}\mathbf{\chi}+v_{i},$ (4)
@@
 [^77] For example, the coefficient of variation for household $i$ is defined as $CV_{i}=\{\sum_{t}^{T_{i}}(\tilde{y}_{i,t}-\bar{\tilde{y}}_{i,.})/(T_{i}-1)\}^{\frac{1}{2}}/\bar{\tilde{y}}_{i,.}$, where $T_{i}$ indicates the number of time-series observations, and $\bar{\tilde{y}}_{i,.}$ indicates the time-series average of the head’s income.
-In this cross-sectional model, I include a set of family composition variables to control for unobservable preferences because the household-fixed effect is unable to be modeled as an error component.
+In this cross-sectional model, I include a set of family composition variables to control for unobservable preferences because the household-fixed effect cannot be modeled as an error component.
 If retailers allowed households to use credit purchases regardless of the variability of earnings, the estimate of $\varkappa_{1}$ becomes close to zero.
-In contrast, if retailers did not allow households with high-income volatility to trade on credit, the estimate should be positive and reasonably large.
+By contrast, if retailers did not allow households with high-income volatility to trade on credit, the estimate should be positive and reasonably large.
 Table 10: Volatility of Earnings and Use of Credit Purchases
@@
 ***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. Robust standard errors are reported in brackets.
-Notes: The dependent variable in Columns (1)-(3) is the indicator variable for the households who did not use any credit purchases during the sample period (non-CP users).
-The dependent variable in Columns (4)-(6) is the indicator variable for the non-CP users and households less than the first quartile in the sample average of credit purchases.
-Columns (1)-(2) and (4)-(5) use the linear probability model (LPM), whereas Columns (3) and (6) use the Probit model.
+Notes: The dependent variable in Columns (1)–(3) is the indicator variable for the households who did not use any credit purchases during the sample period (non-CP users).
+The dependent variable in Columns (4)–(6) is the indicator variable for the non-CP users and households less than the first quartile in the sample average of credit purchases.
+Columns (1)–(2) and (4)–(5) use the linear probability model (LPM), whereas Columns (3) and (6) use the Probit model.
 “CV of Head’s Earnings” indicates the coefficient of variations in the head’s semi-monthly earnings.
@@
 The average head’s age and average head’s earnings are added to the control variables in Columns 3 and 6.
-$F$-statistics $p$-values for the null of the zero-slope hypothesis are reported in Columns (1)-(2) and (4)-(5).
+$F$-statistics $p$-values for the null of the zero-slope hypothesis are reported in Columns (1)–(2) and (4)–(5).
 $\chi^{2}$-statistics $p$-values for the null of the zero-slope hypothesis are listed in Columns (3) and (6).
@@
 The estimated coefficient is close to zero and statistically insignificant.
-A potential concern herein may be the correlations between the volatility of earnings and the head’s age and earnings level.
+A potential concern may be the correlations between the volatility of earnings and the head’s age and earnings level.
 Column (2) shows the result from the extended model, including the head’s age and earnings in the controls, which shows the robustness of the baseline estimate in Column (1).
@@
 It confirms that the estimate is not statistically significantly different from zero.
-Next, I changed the threshold for the dependent indicator variable to consider a wider range of households using a relatively smaller amount of credit purchases, say less than the first quartile of the average credit purchases.
-Columns (4)-(5) summarize the results in the same column layouts as Columns (1)-(3), confirming that the estimates are still close to zero in all the specifications.
-Finally, the final row in this table shows that the null of the joint zero slope hypothesis is strongly rejected in all the regressions.
-Overall, there is no apparent trend in the characteristics of households to use credit purchases.
-This suggests that retailers may not have put weight on the volatility of earnings when trading with factory workers on credit.
-This is consistent with the retailers’ motivation to secure their local customers discussed earlier, and with the fact that factory workers commonly used credit purchases.
-[^82] Despite this, low-income households might have had more difficulty using credit purchases.
+Next, I changed the threshold for the dependent indicator variable to consider a wider range of households including the households with less than the first quartile of the average credit purchases.
+Columns (4)–(6) summarize the results in the same column layouts as Columns (1)–(3), confirming that the estimates are still close to zero in all the specifications.
+Finally, the last row in this table shows that the null of the joint zero slope hypothesis is strongly rejected in all the regressions.
+Overall, there is no apparent trend in the characteristics of households that use credit for their purchases.
+This suggests that retailers may not have considered the volatility of earnings when providing credit to factory workers.
+This is consistent with the retailers’ motivation to secure their local customers discussed earlier, and with the fact that factory workers commonly used credit for purchases.
+[^78] Despite this, low-income households might have had more difficulty using credit for purchases.
 A famous essay drawn by a primary school student in the 1930s depicts an example of a rice retailer who refused to buy rice under credit purchase (Toyoda 1969, pp.
 346–355).
-The results so far indicate that retailers have been generous in the earnings volatility risk.
-This, in turn, implies that retailers may factor risk in the form of interest into their loans or retail prices.
-Unfortunately, there are no systematic statistics on the interest in credit purchases.
-Thus, I calculate the risk premiums backward by differencing payments and purchases.
-This backward calculation technique requires information both on the price and type of retailers.
+The results so far indicate that retailers were generous in considering the earnings volatility risk.
+This, in turn, implies that retailers may factor risk into retail prices or by charging interest on their loans.
+Unfortunately, there are no systematic statistics on the interest charged on credit purchases.
+It may be possible to reverse calculate the risk premiums by differencing payments and purchases.
+This reverse calculation technique requires information on both the price and type of retailers.
 However, budget books usually document price-by-item information but not price-by-retailer information.
-This is problematic because the same items were often sold in different types of retailers: for example, eggs were sold in greengrocers (yaoya) in many cases but sometimes sold in variety stores (aramonoya).
-Fortunately, I found that a specific month in a budget book has documented the price-by-retailer information.
+This is problematic because the same items were often sold in different types of retailers: for example, eggs were frequently sold in greengrocers (yaoya) but sometimes also in general stores (aramonoya).
+Fortunately, I found a specific month in a budget book that has documented the price-by-retailer information.
 [^79] The OISR, Archives of the Tsukishima Survey (THBS #4, February 1919).
 According to this budget book, the reimbursements were made semi-monthly, as seen in Figure 6b.
-They bought various goods in greengrocers, fish stores, and variety stores between 1st and 14th February 1919 and paid exactly the same amounts on the 15th.
-[^84] While the retailers used credit purchases from wholesalers (footnote 78), the repayment period was often set around one month (Tokyo City Office 1932, pp.
+They bought various goods in greengrocers, fish stores, and general stores between February 1 and February 14, 1919, and paid exactly the same amounts on the 15th.
+[^80] The repayment period in credit purchases among retailers and wholesalers (footnote 74) was often set at approximately one month (Tokyo City Office 1932, pp.
 157-159).
-Therefore, this makes sense for consumers to make repayments at semi-monthly or monthly intervals.
-In the latter half of February, the household paid the same amounts to greengrocers and variety stores on 28th, but less amount to the fish stores.
-Since the budget book for March does not have information on the type of retailers, it is difficult to know exactly how the reimbursement shortfall was handled.
+Therefore, customers likely made repayments at semi-monthly or monthly intervals.
+In the latter half of February, the household paid the same amounts to greengrocers and general stores on the 28th, but paid less to the fish stores.
+As the budget book for March does not have information on the type of retailers, it is difficult to determine exactly how the reimbursement shortfall was managed.
 However, this case suggests that the risk premiums may have been included in the retail prices but not in the form of interest.
-Finally, I assess the loss for the retailers in credit purchases.
-According to the statistics on the losses on credit sales documented in the 1930 Survey Report, almost all of the retailers had suffered losses in credit purchases.
-For example, 99% of rice and 97% of fuel retailers more or less lost their sales.
+Finally, I assessed the losses suffered by retailers for credit purchases.
+According to the 1930 Survey Report, almost all the retailers suffered losses for credit purchases.
+For example, $99$% of rice and $97$% of fuel retailers lost their sales.
 This means that credit purchase was inherently a loss-making sales method for the retailers.
-Since systematic statistics are unavailable to clarify the percentage of total profits corresponding to these losses.
-Thus, it is impossible to discuss how severe these losses are for the retailers.
-However, an important fact is that these losses were usually capped by 5% of total sales.
+As systematic statistics are unavailable to clarify the percentage of total profits corresponding to these losses, it is impossible to assess how severe these losses were for the retailers.
+However, an important fact is that these losses were usually capped at 5% of total sales.
 For example, the available statistics suggest that $87$% of rice and $65$% of fuel retailers suffered losses amounting to $5$% of sales.
 This supports the evidence that while the retailers could not avoid losses in credit sales, they might have set prices that anticipated predicted losses to some extent.
-This seems to be a plausible sales strategy for retailers to deal with adverse selection issues under imperfect consumer information.
+This seems to be a plausible sales strategy for retailers to manage adverse selection issues under imperfect consumer information.
 Online Appendix C.4 summarizes the statistics on the losses in credit purchases available from the 1930 Survey Report in detail.
 To summarize, the available documents show that retailers competed with other retailers in the same sales area.
-Under this condition, they provided regular consumers with the opportunity to trade on credit without interest for at least one month.
+Under this circumstance, they provided regular consumers with the opportunity to trade on credit without interest for at least one month.
 The 1930 Survey Report indeed stated that “intense competition among competitors has led to a high frequency of credit purchases” (Tokyo City Office 1932b, p.
@@
 This accessibility to credit purchases could help workers manage consumption after unexpected individual setbacks in the short term.
-[^85] Another potential explanation for the popularity of credit purchases might be thefts.
-However, the number of thefts $100$ people in 1919 is $0.5$ in Tsukishima, which is close to or slightly lower than the city average of $1.3$ (National Police Agency 1920b, pp 53–54; 244–245).
-In fact, Tsukishima Police Department does not mention that Tsukshima was an area with a particularly high rate of thefts (Tsukishima Police Department 1976, pp.
+[^81] Another potential explanation for the popularity of credit purchases might be theft.
+However, the number of thefts in $100$ people in 1919 was $0.5$ in Tsukishima, which is close to or even slightly lower than the city average of $1.3$ (National Police Agency 1920b, pp 53–54; 244–245).
+In fact, Tsukishima Police Department does not mention that Tsukshima was an area with a particularly high rate of theft (Tsukishima Police Department 1976, pp.
 244–245).
-This suggests that this crime channel is not relevant to explain the use of credit purchases.
+Therefore, the crime channel is not relevant to explain the use of credit purchases.
 6 Conclusion
-This study investigates the consumption smoothing behaviors among working-class households in Tokyo.
-I digitized the daily longitudinal budget survey on the factory worker households conducted in Tsukishima and implemented a systematic econometric method to obtain comparable estimates for income elasticities.
+This study investigates the consumption-smoothing strategies used by working-class households in prewar Tokyo.
+I digitized the daily longitudinal budget survey on factory worker households conducted in Tsukishima in 1919.
+I carefully assessed the potential measurement error due to the miss-assignments of paydays and proposed a way to avoid the assignment errors.
 The estimated elasticity for total consumption indicates that the factory worker households suffered certain consumption losses when they faced idiosyncratic income shocks.
-Despite this, I found evidence that while expenses on dispensable categories are likely to be cut in response to the income losses, expenses on indispensable categories are partially smoothed.
-The estimated elasticity for total expenditure is in a similar range to those in rural economies in developing countries, although only a few available references exist.
-According to Ravallion and Chaudhuri (1997), the elasticity for the rural households in Indian villages between 1975 and 1984 was roughly $0.5$.
-As for the urban households, the estimated elasticity is $0.4$ for Bangkok between 1975 and 1990, which had a similar GDP per capita to 1920 Japan (Townsend 1995).
-Since my estimate is from a specific occupation, skilled factory workers, one must be careful when making comparisons with these estimates based on urban samples, which include a variety of occupations.
-At least, however, the factory workers in the historical metropolis of Japan around WWI had been relatively vulnerable to the idiosyncratic income shocks.
+The estimated elasticity for total expenditure ($0.43$) is in a similar range to those in rural economies in developing countries.
+According to Ravallion and Chaudhuri (1997), the elasticity for the rural households in Indian villages between 1975 and 1984 was approximately $0.5$.
+As for the urban households, the estimated elasticity was $0.4$ for Bangkok between 1975 and 1990, which had a similar GDP per capita to 1920 Japan (Townsend 1995).
+Since my estimate is from a specific occupation, skilled factory workers, we need to be careful when making comparisons with these estimates based on urban samples which include a variety of occupations.
+Nevertheless, the factory workers in the historical metropolis of Japan circa WWI were vulnerable to idiosyncratic income shocks.
 A recent study found an income elasticity of $0.39$ for factory workers’ households in Osaka city circa 1920 (Ogasawara 2024).
 My estimate from the adjusted monthly panels is slightly larger than this estimate.
-Although the measurement errors in aggregate consumption and income in the monthly panel dataset may explain the discrepancy in these estimates, if taken literally, this implies that the Tsukishima factory workers were more vulnerable than those in Osaka.
+If taken literally, this implies that the Tsukishima factory workers were slightly more vulnerable than those in Osaka.
 This indicates that higher levels of urbanization do not necessarily correspond to a greater availability of risk-mitigation institutions.
-Therefore, inter-city and cross-national comparisons of income elasticities and analysis of the institutional characteristics behind the consumptions there would be another subject of future work.
-The finding on the attenuation bias due to measurement errors is potentially important because it may prompt a reinterpretation of existing results.
-Errors in the timing of payroll can result in inconsistencies in the aggregate data for income and expenditure, albeit the household-level studies often use aggregate data.
-In this case, modifying the aggregation period to align with wage payments can reduce attenuation bias.
-Even if the aggregation period is unadjustable, clarifying payroll timing and discussing potential bias is useful.
-The results from mechanism analyses provide suggestive evidence that credit purchases played an important role in mitigating adverse idiosyncratic shocks on consumption.
-Although the institutional details of credit purchases should be investigated further, this finding offers a new historical view of the interdependence between consumers and local retailers.
-Understanding the local retail system would be important in comprehending regional consumption behavior and, in turn, in understanding the improvement in living standards in prewar Japan.
+Therefore, inter-city and cross-national comparisons of income elasticities and analysis of the institutional characteristics behind consumption are subjects for future work.
+Despite the vulnerability, the results from mechanism analyses show that credit purchases were commonly used among workers and played a role in mitigating adverse idiosyncratic shocks in several consumption subcategories, particularly food.
+My findings offer a new historical view of the interdependence between consumers and local retailers.
+As researchers build an understanding of local credit institutions we can comprehend regional consumption behavior and may understand how living standards improved in prewar Japan.
 Appendices
 Appendix Appendix A Background Appendix
-A.1 Developments of the Cities
+A.1 Development of the Cities
 (a) Number of people in the large cities
 (b) Share of people in the large cities (%)
-Figure A.1: Developments of the Large Cities in Japan
-Note: This figure illustrates the development of large cities in Japan between 1889 and 1920.
+Figure A.1: Development of the Large Cities in Japan
+Notes: This figure illustrates the development of large cities in Japan between 1889 and 1920.
 Figure A.1a shows the number of people in Tokyo, Osaka, Kyoto, and Nagoya cities.
-Figure A.1b shows the share of people in Tokyo (diamond) and the cities with more than $50,000$ people (circle) in total population in Japan.
+Figure A.1b shows the share of people in Tokyo City (diamond) and the cities with more than $50,000$ people (circle) in total population in Japan.
 Sources: Umemura et al. (1983, pp. 303–305); Umemura et al. (1988, pp. 166; 168).
@@
 Notes:
-1. Panel A: The number of people and male workers measured in the Population Census conducted in October 1920 are listed.
-2. Panel B: Occupations are classified based on the industrial classification of the 1920 Population Census. Unemployed people are not included in the calculation.
-3.
+Panel A: The number of people and male workers measured in the Population Census conducted in October 1920 are listed.
+Panel B: Occupations are classified based on the industrial classification of the 1920 Population Census. Unemployed people are not included in the calculation.
 Panel C: The number of electrical lights measured at the end of 1921 is divided by the number of people in the 1920 Population Census.
 The number of houses accessible to water supply measured at the end of September 1920 is divided by the number of total households measured in the 1920 Population Census (%).
-4. Panel D: The average daily temperature and precipitation between 1st January 1918 and 31st December 1923 and between 1st February 1920 and 30th April 1920 are listed.
+Panel D: The average daily temperature and precipitation between January 1, 1918, and December 31, 1923, and between February 1, 1920, and April 30, 1920 are listed.
 Sources: The number of people in 1920 is from the 1920 Population Census (Statistics Bureau of the Cabinet 1928, pp.
@@
 A.2 Financial Institutions in the Cities
-Several financial institutions were available for working-class households in the cities around 1920.
-A representative savings institution is the postal savings (yūbin chokin).
-Postal savings was initially designed as a small savings institution for low-income households, and the share of postal savings to total savings was roughly $10$% in the 1910s (Tanaka 2018, p.
+Several financial institutions were available for working-class households in the cities circa 1920.
+A representative savings institution was the postal savings (yūbin chokin).
+Postal savings was initially designed as a small savings institution for low-income households, and the share of postal savings to total savings was approximately $10$% in the 1910s (Tanaka 2018, p.
 29).
@@
 33).
-[^86] James and Suto (2011) show that the saving rate in prewar Japan was higher than that of the U.S.
-As shown in Panel A of Table A.2, roughly $1.4$ million people had accounts in Tokyo in 1920.
-This covers approximately $67$% of the entire population in Tokyo and the average savings amount per capita was $64.5$ yen.
+James and Suto (2011) indeed show that the saving rate in prewar Japan was higher than that of the US.
+As shown in Panel A of Table A.2, approximately $1.4$ million people had accounts in Tokyo in 1920.
+This covers approximately $67$% of the entire population in Tokyo City and the average savings amount per capita was $64.5$ yen.
 In Kyobashi Ward, the average figure for the postal savings accounts was $44.5$ yen at the end of 1919 (Tokyo City Office 1921, pp.
 892–893).
-This amount is less than the average monthly earnings of the skilled-factory workers (Section 3.2).
+This amount was less than the average monthly earnings of the skilled-factory workers (Section 3.2).
 Unfortunately, the statistics on postal savings by depositors’ occupation are unavailable.
-[^87] There are some statistics for the entire Japan.
+[^82] There are some statistics for the whole of Japan.
 According to the Postal Savings Bureau (1924, p.
-72), the share of depositors in the manufacturing sector was $4.8$% (the share based on the deposit amount is $5.5$%).
-However, this implies that, on average, postal savings may not provide sufficient temporary income.
-Savings bank (chochiku ginkō) is another famous savings institution for working-class households in the 1910s (Ito and Saito 2019, pp 79–81).
-Panel A of Table A.2 shows that the number of people having savings accounts was roughly $1.46$ million, accounting for approximately $67.3$% of total people in Tokyo at that time.
+72), the share of depositors in the manufacturing sector was $4.8$% (the share based on the deposit amount was $5.5$%).
+However, this implies that, while a large proportion of the workers might have had accounts, postal savings may not have provided sufficient temporary income.
+Savings bank (chochiku ginkō) was another popular savings institution for working-class households in the 1910s (Ito and Saito 2019, pp 79–81).
+Panel A of Table A.2 shows that the number of people who had savings accounts was approximately $1.46$ million, accounting for approximately $67.3$% of the total people in Tokyo City at that time.
 Although this figure is similar to the postal savings, the workers in the manufacturing industry were less likely to rely on the savings banks than the workers in the other industrial sectors.
 Table A.3 summarizes the number of depositors and the amounts of deposits in savings banks by occupation in Tokyo City (Panel A) and Kyobashi Ward (Panel B).
-In Kyobashi, the commerce and miscellaneous sectors shared roughly $84.8$% of all the depositors, whereas the manufacturing sector had only $11.5$%.
+In Kyobashi, the commerce and miscellaneous sectors shared approximately $84.8$% of all the depositors, whereas the manufacturing sector had only $11.5$%.
 This is unchanged if I use the amounts of deposits to calculate the shares.
-Tokyo Institute for Municipal Research stated that this tendency captures the fact that savings banks are mainly designed as financial institutions for low-income households (Tokyo Institute for Municipal Research 1925b, p.
+The Tokyo Institute for Municipal Research confirmed that savings banks were designed as financial institutions for low-income households (Tokyo Institute for Municipal Research 1925b, p.
 81).
-Thus, skilled factory worker households, like the THBS households, were less likely to use the savings banks.
+Thus, skilled factory worker households—like the THBS households—were less likely to use the savings banks.
 This is consistent with the fact that Tsukishima did not have any savings bank branches, while there were two postal offices.
-[^88] There were 123 savings banks in Tokyo City in 1922, and 10 of them were in Kyobashi Ward (Tokyo Institute for Municipal Research 1925, p.
+[^83] There were 123 savings banks in Tokyo City in 1922, and ten of them were in Kyobashi Ward (Tokyo Institute for Municipal Research 1925, p.
 76).
-However, Tsukishima, a representative manufacturing area, did not have savings banks at that time.
-Mutual loan association (mujin) is another type of savings institution.
-The members of this association pay a fixed amount of money several times, and each member withdraws the total amount of payments each time under specific rules.
+However, Tsukishima—a representative manufacturing area—did not have savings banks at that time.
+Mutual loan association (mujin) was another type of savings institution.
+The members of this association paid a fixed amount of money several times, and each member withdrew the total amount of payments each time under specific rules.
 The lotteries or bidding system decided which members could withdraw money each time.
-This means that the workers could not use this institution to deal with the acute negative income shocks.
+This means that the workers could not use this institution to manage acute adverse income shocks.
 [^84] This is the most representative mutual loan system called tsūjyō (i.e., common) mujin.
-There was a different type called sueoki chokin (or tsumitate kai) that people could receive their stakes with interests after a fixed amount had been accumulated for a certain period (Tokyo Legislative Research Association 1915, pp.
+There was a different type called sueoki chokin (or tsumitate kai) from which people could receive their stakes with interests after a fixed amount had been accumulated for a certain period (Tokyo Legislative Research Association 1915, pp.
 3–4).
-Clearly, both are also not flexible for compensating the income shocks.
-Instead, it had been used to make a certain amount of money among small business owners and merchants.
-A report suggests that the workers in the manufacturing sector shared only $10$% of all the mutual loan associations in Tokyo (Tokyo City Office 1935, p.
+However, these were also not flexible for compensating for income shocks.
+Instead, these were used by small business owners and merchants to generate money.
+A report suggests that the workers in the manufacturing sector shared only $10$% of all the mutual loan associations in Tokyo City (Tokyo City Office 1935, p.
 25).
@@
 Cooperative societies were also available for worker households.
-Cooperative societies in Tokyo used credit as their payment method: households could buy goods in advance and pay at the end of the month or twice (middle and the end of) a month.
-The commodity prices were similar to or slightly lower than those in the retail shops (Central Federation of Industrial Associations 1925, pp.
+Cooperative societies in Tokyo used credit as their payment method—households could buy goods in advance and pay at the end of the month or twice (in the middle and at the end of) a month.
+The commodity prices were similar to or slightly lower than those in the retailers (Central Federation of Industrial Associations 1925, pp.
 18–19).
-However, the cooperatives only covered a few citizens.
+However, the cooperatives covered a small proportion of citizens.
 There were $26$ cooperative societies in Tokyo prefecture in 1924, and 23 of them were organized by the citizens and workers (Central Federation of Industrial Associations 1925, p.
 57–59).
-Among these 23 cooperatives, two were large societies with roughly 4000 members, and the others were small societies with less than 100 members.
-The total number of accounts for only a few percentage points in the workers in the city, given that the number of male workers was more than 768 thousand (Table A.1).
-In fact, most of the THBS households had not used the cooperatives.
-An alternative borrowing institution is the credit purchase in retail shops.
-The THBS households had frequently used credit purchases in various retail shops for rice, fish, vegetables, alcohol, firewood, charcoal, and medicines.
-The details of the credit purchase are little documented, and the systematic statistics on the credit purchase are also unavailable.
-This must be because the information on the transactions is personal.
-Despite this, an available document shows that credit purchases are used in various retail shops (Section 5).
-The THBS budget data show a consistent tendency for credit purchases to be widely used among factory worker households.
-Pawnshops were the popular lending institution among low-income working-class households.
-Lenders did not need to screen the borrower’s credit, and the borrowers did not worry about incurring heavy debt as the primary articles were inexpensive clothes (Shibuya et al.
+Among these $23$ cooperatives, two were large societies with approximately $4,000$ members, and the others were small societies with less than $100$ members.
+Given that the number of male workers was more than $768$ thousand (Table A.1), these account for only a few percentage points of the total workers in the city.
+In fact, the THBS households rarely used the cooperatives.
+An alternative borrowing institution was credit purchase in retailers.
+The details of the credit purchase are little documented, and the systematic statistics on credit purchases are also unavailable.
+I surmise that this was because the information on the transactions was personal.
+Despite this, a few available documents show that credit purchases were used in various retailers at that time (Section 5).
+The THBS households frequently used credit for purchases in various retailers for rice, fish, vegetables, alcohol, firewood, charcoal, and medicines.
+Pawnshops were the most popular lending institution among low-income working-class households.
+Lenders did not need to screen the borrower’s credit, and the borrowers did not worry about incurring heavy debt as the primary articles pawned were inexpensive clothes (Shibuya et al.
 1982).
-In addition, the interest rates were regulated by the Pawnbroker Regulation Act of 1895, and the average redemption rate was substantially high under the lower interest rates.
+In addition, the interest rates were regulated by the Pawnbroker Regulation Act of 1895.
+Thus, the average redemption rate was substantially high under the lower interest rates.
 Among $88$ pawnshops in Kyobashi, the average redemption rate was approximately $94$% ($401,222$ cases/$427,265$ cases) in 1920 (Tokyo City Office 1922b, pp.
 888–889).
-The accessibilities were higher than the other lending institutions.
-The number of pawnshops in Tokyo was $1,334$ in 1918 and $1,261$ in 1919, respectively.
+The number of pawnshops in Tokyo was $1,334$ and $1,261$ in 1918 and 1919, respectively.
 The total number of cases was $8,226,883$ and $7,573,406$ each year.
-These figures are more than three times greater than the entire population in Tokyo City.
-This means that the accessibility to a pawnshop was relatively high for the workers.
-As described in the main text, the number of pawnshops in Tsukishima was 13, whereas that of moneylenders was five.
-However, the inexpensive article means that pawnshops were used to deal with very short-run necessities only.
-Among the pawnshops of Kyobashi, for example, the average amount per case was $7.8$ yen, which was $12$% of the monthly earnings of the THBS households.
-While the pawnshops were popular lending institutions, they could not offer a sufficient amount of temporary income for relatively large idiosyncratic shocks.
+These figures were more than three times greater than the entire population in Tokyo City.
+This means that pawnshops were more accessible to workers than other institutions.
+As described in the main text, there were $13$ pawnshops in Tsukishima, whereas there were five moneylenders.
+However, the inexpensive articles pawned meant that pawnshops were available only for short-run necessities.
+Among the pawnshops of Kyobashi, for example, the average amount per case was $7.8$ yen, which was approximately $10$% of the monthly earnings of the THBS households.
+While the pawnshop was a popular lending institution, it could not offer enough temporary income for relatively large idiosyncratic shocks.
 According to the statistics on the pawnshop users in Tokyo prefecture in 1923, the share of users in the agriculture, commerce, and manufacturing sectors was $1.7$%, $19.5$%, and $15.6$%, respectively.
-In turn, the most representative users were the day laborers and workers classified as “miscellaneous”, accounting for $63.4$% (Tokyo Institute for Municipal Research 1926, p.
+In turn, the most representative users were the day laborers and workers classified as “miscellaneous,” accounting for $63.4$% (Tokyo Institute for Municipal Research 1926, p.
 25).
@@
 This indicates the evidence that the primary pawnshop users were not the skilled-worker households like the THBS households.
-Another lending institution is the money lenders (kinsen kashitsuke gyō).
-Generally, business owners used the money lenders.
-The interest rates in the money lenders were very high compared with those in the pawnshops (Shibuya 2000, pp.
+Another lending institution was the money lenders (kinsen kashitsuke gyō).
+Generally, business owners used money lenders.
+The interest rates of the money lenders were very high compared with those in the pawnshops (Shibuya 2000, pp.
 184; 248).
-Tokyo Institute for Municipal Research was concerned about the increasing number of unscrupulous companies arrested by the National Police Agency (Tokyo Institute for Municipal Research 1925b, pp.
+Tokyo Institute for Municipal Research was indeed concerned about the increasing number of unscrupulous companies arrested by the National Police Agency (Tokyo Institute for Municipal Research 1925b, pp.
 106–108).
-Given this fact, while there were two money lenders in Tsukishima, they might have been less likely used by the factory worker households.
-Credit unions (shinyō kumiai) are another institution available for working-class households.
+Given this fact, while there were two money lenders in Tsukishima, they were less likely to be used by the factory worker households.
+Credit union (shinyō kumiai) was another institution available for working-class households.
 In Tokyo City, there were $215$ unions in 1923.
@@
 20).
-This accounts for, at the most, only a few percentage points of the male workers in the manufacturing sector in Tokyo measured in the population census of 1920 (Online Appendix Table A.1).
+This accounts for only a few percentage points of the male workers in the manufacturing sector in Tokyo measured in the population census of 1920 (Online Appendix Table A.1).
 Table A.2: Financial Institutions in Tokyo, Osaka, Kyoto, and Nagoya Cities in 1920
@@
 Panel B: Pawnshop
-Number of shops 1,253 867 441 412†
+Number of retailers 1,253 867 441 412†
 Number of units lent 6,042,501 2,027,690 1,251,283 319,609
@@
 Pane A: Entire Tokyo City
-      Number of depositors 76210 550784 166498 619963 1413455
-      Amount of deposits 2002124 19112713 6663091 19809591 47587519
+      Number of depositors 76,210 550,784 166,498 619,963 1,413,455
+      Amount of deposits 2,002,124 19,112,713 6,663,091 19,809,591 47,587,519
       Number of depositors (%) 5.4 39.0 11.8 43.9 100.0
@@
 Pane B: Kyobahi Ward
-      Number of depositors 3809 34665 11664 51481 101619
-      Amount of deposits 130781 1218081 450805 1696086 3495753
+      Number of depositors 3,809 34,665 11,664 51,481 101,619
+      Amount of deposits 130,781 1218,081 450,805 1,696,086 3,495,753
       Number of depositors (%) 3.7 34.1 11.5 50.7 100.0
       Amount of deposits (%) 3.7 34.8 12.9 48.5 100.0
-Notes: This table summarizes the number of depositors and the amount of deposits in the savings banks by occupation.
-In the document, the miscellaneous category is not clearly defined but named as “Miscellaneous (zatsu gyō)”.
-These statistics were measured on 31st December 1919.
+Notes: This table summarizes the number of depositors and the total deposits in the savings banks by occupation.
+In the document, the miscellaneous category is not clearly defined but named as “Miscellaneous (zatsu gyō).” These statistics were measured on December 31, 1919.
 Panel A and B list the statistics for Tokyo City and Kyobashi Word, respectively.
@@
 This map does not illustrate the northernmost part of Ishikawa jima (see Figure 3 for this part).
-“Ferries” shows the Tsukishima to Kyobashi Ward sea routes: Tsukuda, Tsukishima, and Kachidoki routes (right to left).
+“Ferries” shows the Tsukishima to the mainland of Kyobashi Ward sea routes called Tsukuda, Tsukishima, and Kachidoki routes (right to left).
 “Aioi bridge” links Shin tsukudajima and Fukagawa ward.
@@
 Tsukishima (transliterated as “moon island”) is a reclaimed land on a sandbar of the Sumida River.
-Ishikawajima and Tsukudajima (“jima” means island) at the northern end of Tsukishima are the starting point of its history (Figure A.3 for map).
-The following description of this section is based on Chūō Ward (1994, pp.
+Ishikawajima and Tsukudajima islands at the northern end of Tsukishima (Figure A.3) are the starting point of its history.
+The following description in this section is based on Chūō Ward (1994, pp.
 122–124).
 Ishikawajima was built in 1626 by a boatman, Hachizaemon Ishikawa, who received a fief from the Shogunate.
-In 1790, the Ishikawajima Prison was built on Ishikawajima.
-This was a facility for the reintegration of minor offenders and those without lodgings and was called a Ninsoku Yoseba.
+In 1790, the Ishikawajima Prison (ninsoku yoseba) was built on Ishikawajima.
+This was a facility for the reintegration of minor offenders and those without lodgings.
 After Matthew Perry arrived in Japan in 1853, the Shogunate ordered the Mito Clan to build a shipyard on Ishikawajima.
-A month later, the groundbreaking ceremony was held for a Western-style wooden sailing ship delivered to the Shogunate in 1856.
-In 1866, the Ishikawajima Lighthouse was built on the west bank of the Yoseba and served to signal the position of ships entering the Nihonbashi direction.
-In 1870, the Ninsoku Yoseba became a place of imprisonment (Toba) and was renamed the Ishikawajima Jail Station of the Metropolitan Police Department in 1877.
+A month later, a groundbreaking ceremony was held for a Western-style wooden sailing ship delivered to the Shogunate in 1856.
+In 1866, the Ishikawajima Beacon was built on the west bank of the Ishikawajima Prison and served to signal the position of ships entering the Nihonbashi direction.
+In 1870, the Ishikawajima Prison became a place of imprisonment (toba) and was renamed the Ishikawajima Jail Station of the Metropolitan Police Department in 1877.
 The prison was finally moved to Sugamo in Kitatoshima County in Tokyo prefecture, and the penitentiary on Ishikawajima was closed in 1895.
-Adjacent to Ishikawajima is called Tsukudajima, where fishermen from “Tsukuda” in Osaka prefecture settled in 1644 after receiving land from the Shogunate (i.e., the island was named after their hometown, Tsukuda).
+The area adjacent to Ishikawajima is called Tsukudajima, where fishermen from “Tsukuda” in Osaka prefecture settled in 1644 after receiving land from the Shogunate.
 In 1645, the “Tsukuda ferry” began, and throughout the Edo period, Tsukudajima was known as a fishing island for whitefish for the Shogun’s household.
-In 1863, the Tsukudajima Battery was built on the south side of Tsukudajima, one of 13 batteries used to defend Edo Bay.
+In 1863, the Tsukudajima Battery was built on the south side of Tsukudajima—one of $13$ batteries used to defend Edo Bay.
 This battery later became the starting point for the reclamation of Lot 1 of Tsukishima.
@@
 The first Tsukishima site was completed in 1891, and the “Tsukishima ferry” began the same year.
-The second site was created in 1894; the Shin (new) tsukudajima was built in 1896.
+The second site was created in 1894; the Shin Tsukudajima (namely, new Tsukudajima) was built in 1896.
 From the west shore of Shin Tsukudajima, the island overlooks the Boso Peninsula in Chiba Prefecture, making it a favorite vacation spot for writers.
 Cultural figures such as Toson Shimazaki and Kaoru Osanai stayed here from 1907 to the early Taisho period.
-In 1903, the Aioi Bridge was completed, bringing water and electricity to Tsukishima and improving transportation convenience, which contributed to industry development.
-[^90] Until then, Suzukiiya had been selling water to supply the entire island using a ferryboat.
-The “Kachidoki Ferry” began to lead to Lot 2 in 1905.
-In 1913, Lot 3 was finally completed.
-A.4 Housings in Tsukishima
-Since the Tsukishima Survey included a housing survey (Table B.1), the housing types in the entire island can be determined.
-Table A.4 summarizes the results of this survey.
-The description in this section is based on the Department of Health, Ministry of the Interior (1923a, pp.
+In 1903, the Aioi Bridge was completed, bringing water and electricity to Tsukishima and improving transportation convenience, which contributed to development of industry.
+[^85] Until then, a retailer named Suzukiiya had been selling water to supply the entire island using a ferryboat.
+The “Kachidoki Ferry” began to deliver to Lot 2 in 1905.
+Lot 3 was finally completed in 1913.
+A.4 Housing in Tsukishima
+The Tsukishima Survey included a housing survey (Table B.1).
+Table A.4 summarizes the housing types measured in this survey.
+As shown, $88$ ($46$%) of the detached homes are in Tsukudajima.
+Tsukudajima is reported to be “superior to other areas in terms of housing quality, and the percentage of inferior housing is meager.” [^86] The description in this section is based on the Department of Health, Ministry of the Interior (1923a, pp.
 69–76).
-$88$ ($46$%) of the detached homes are in Tsukudajima.
-Tsukudajima is reported to be “superior to other areas in terms of housing quality, and the percentage of inferior housing is meager”.
 Shin-tsukudajima had a low percentage of detached homes.
-The report states that it was indeed “inhabited by unskilled laborers” and “occupies the most inconvenient housing status of all the islands.” Lot 1 was a center of life in Tsukishima and is a cluster of numerous residences, mainly tenements.
-Lot 2 had a relatively high percentage of single-family and two-family houses and appears to have had a character as a “workers’ residential area” than Lot 1.
+The report states that it was indeed “inhabited by unskilled laborers” and “occupied the most inconvenient housing status of all the islands.” Lot 1 was the center of life in Tsukishima and was a cluster of numerous residences, mainly tenements.
+Lot 2 had a relatively high percentage of single-family and two-family houses and appears to be characterized as a “workers’ residential area” than Lot 1.
 These indicate that there might have been some geographic differences in income distribution regarding housing type.
-Overall, however, of the $989$ dwellings in Tsukishima, $799$ ($81\%$) are row houses, and only $190$ are single-family homes.
+Overall, however, of the $989$ dwellings in Tsukishima, $799$ ($81\%$) were row houses, and only $190$ were single-family homes.
 This confirms that Tsukishima was an industrial area with many working families.
@@
 Total 190 148 153 124 374 989 491 49.6
-Note: This table summarizes the housing types in Tsukishima surveyed on 30th June 1919.
+Note: This table summarizes the housing types in Tsukishima surveyed on June 30, 1919.
 Source: Department of Health, Ministry of the Interior 1923a, pp. 69–76.
@@
 A.5 Machinery Factories in Tsukishima
-First, I use the results of the Primary School Children Survey included in the Tsukishima Survey to analyze the occupation (industry) of the household heads in Tsukishima.
+First, I used the results of the Primary School Children Survey included in the Tsukishima Survey to analyze the occupation (industry) of the household heads in Tsukishima.
 This is a survey of 4th-6th graders at Tsukishima 1st and 2nd Elementary Schools in July 1919.
 There were $1,173$ children in these grades, and the survey included all $1,100$ children who resided in Tsukishima.
-At that time, there were no other elementary schools besides those mentioned above.
-Thus, this survey is a complete survey for all children in grades 4-6 in Tsukishima.
-Although it does not cover households without children in grades 4-6, it is undoubtedly a beneficial survey that provides an overview of the households with children that are the main target of this study.
+At that time, there were no other elementary schools.
+Thus, this survey is a complete survey for all children in grades 4–6 in Tsukishima.
+Although it does not include households without children in grades 4–6, it is undoubtedly a beneficial survey that provides an overview of the households with children—the main target of this study.
 Panel A of Table A.5 shows the industrial structures.
@@
 It does not specify whether brothers and sisters in the same household were counted twice.
-Despite this, this similarity confirms that the statistics obtained from this survey are useful in identifying overall trends in family households in Tsukishima.
+Despite this, the similarity confirms that the statistics obtained from this survey are useful to identify overall trends in family households in Tsukishima.
 Panel B of Table A.5 divides the “social classes” of the household heads working in the manufacturing sector into six categories.
-In the report, large business owners are classified into the capitalist class; small business owners, free enterprise owners, and directors are regarded as the middle class; workers are defined as the working class (Department of Health, Ministry of the Interior 1923a, p.
+In the report, large business owners were classified into the capitalist class; small business owners, free enterprise owners, and directors were regarded as the middle class; workers were defined as the working class (Department of Health, Ministry of the Interior 1923a, p.
 50).
-The table reveals at least three things.
-First, most heads of households are workers, with only 1% of the capitalist class and less than 20% of the middle class.
-While the definition of “executive” is unclear, “small business owners” must be small factory owners who could have similar earnings to the skilled worker households (Department of Health, Ministry of the Interior 1923a, p.
+Most heads of households were workers, with only 1% of the capitalist class and less than 20% of the middle class.
+While the definition of “executive” is unclear, “small business owners” must be small factory owners who could have had similar earnings to the skilled worker households (Department of Health, Ministry of the Interior 1923a, p.
 64).
@@
 Observations 1,100 100
-Panel B: Social Class of Heads in Manufacturing Sector
+Panel B: Worker Types of Heads in Manufacturing Sector
 Name of survey Tsukishima Survey
@@
 % share of workers 85.2 80.3 76.4 82.4 78.8
-Notes: This table summarizes the statistics from the complete survey on the 4th-6th graders in Tsukishima1st and 2nd Primary Schools in July 1919.
-Panel A summarizes the heads’ industries, and Panel B sorts the “social classes” of the heads in the manufacturing sector, respectively.
-The report does not specify the definition of each “social class”.
+Notes: This table summarizes the statistics from the complete survey on the 4th–6th graders in Tsukishima1st and 2nd Primary Schools in July 1919.
+Panel A summarizes the heads’ industries, and Panel B sorts the “worker type” of the heads in the manufacturing sector, respectively.
 Source: Department of Health, Ministry of the Interior 1923a, pp. 386–387.
-In November 1920, there were 214 factories in Tsukishima, 168 of them (roughly $80$%) were classified as machinery factories.
+In November 1920, there were $214$ factories in Tsukishima, $168$ of them (approximately $80$%) were classified as machinery factories.
 The official report documented that the wooden pattern factories could be included in the machinery sector because they were incorporated into the casting processes (Department of Health, Ministry of the Interior 1923a, p.
@@
 The share of machinery factories becomes approximately $90$% if I include those $24$ wooden pattern factories.
-Panel A of Table A.6 uses the manufacturing survey in the Tsukishima Survey to classify the 168 factories by size and industrial sector.
+Panel A of Table A.6 uses the manufacturing survey in the Tsukishima Survey to classify the $168$ machinery factories by size and industrial sector.
 First, more than $90$% of the plants are in the three sectors of machinery and equipment manufacturing, can manufacturing, and blacksmithing, with only $15$ plants in the casting sector.
-113 ($67$%) of the machinery factories had workers less than 15.
+$113$ ($67$%) of the machinery factories have less than $15$ workers.
 Second, many smithing factories have fewer than $15$ workers.
-Hence, almost all of Tsukishima’s smithing industry was considered to be small businesses.
-[^91] The report suggests that 43 of the 47 smithing factories were the “town factories” or “workplaces”.
-The definition of “town factory” is not clear in the report (p.
+Hence, almost all of Tsukishima’s smithing industry was classified as small businesses.
+[^87] The report suggests that $43$ of the $47$ smithing factories were “town factories” or “workplaces.” The definition of “town factory” is not clear in the report (p.
 392).
@@
 Panel B of Table A.6 classifies the types of the $620$ workers based on the results of the primary school survey used in Table A.5.
-There were 455 skilled and 165 unskilled workers.
-Notably, these 620 workers include non-industrial workers: 490 industrial workers, 13 commercial, 61 in transportation, 16 in public service free trade, 38 in other gainful employment, and 2 in agriculture and mining in total.
-Since the workers in the industrial sector should include factory workers, workers in the cottage industry, and handicraft workers, the day laborers are considered to be classified into non-industrial workers (in fact, $620$ workers $-120$ day laborers $=490$).
-Therefore, roughly 90% of factory workers ($455/490$) can be regarded as skilled workers, $65$% of whom worked in factories, $17$% in cottage industries, and $18$% as craftsmen.
-Since the workers in cottage and handicraft industries are considered to work in small factories, $65$% of these workers might have worked in the middle- to large-scale factories (with 15+ workers).
-One must be careful that this survey only covers the heads with primary school children aged 9–11 (i.e., 4th to 6th grades), suggesting that most of these heads are in their 30s.
-Therefore, the share of skilled workers to the entire workers should be smaller than the $90$%.
-At least, however, it is plausible to see that most of the Tsukishima workers in their 30s were skilled workers at that time.
+There are $455$ skilled and $165$ unskilled workers.
+The same report shows that $490$ out of $620$ workers are industrial workers.
+The day laborers are classified as non-industrial workers.
+In fact, $620$ workers minus $120$ day laborers become $490$, which is identical to the number of industrial workers documented.
+Therefore, approximately $90$% of factory workers ($455/490$) can be regarded as skilled workers in the industrial sector: $65$% work in factories, $17$% in cottage industries, and $18$% are craftsmen.
+Since workers in cottage and handicraft industries are classified as working in small factories, $65$% of these workers might have worked in the middle- to large-scale factories (with $15+$ workers).
+Notably, this survey only includes household heads with primary school children aged 9–11 (i.e., 4th to 6th grades), suggesting that most of these heads are in their thirties, in the skilled workers age range.
+Therefore, the proportion of skilled workers to the entire worker group could be smaller than the $90$%.
+Nevertheless, this confirms that most of the Tsukishima workers in their 30s were skilled workers at that time.
 Table A.6: Scale of Machinery Factories and Type of Workers in Tsukishima
@@
 Notes: Panel A summarizes the number of machinery factories by the scale and type of the factory.
-Statistics are from the manufacturing census conducted in the Tsukishima Survey (1st November 1920).
+Statistics are from the manufacturing census conducted in the Tsukishima Survey (November 1, 1920).
 The threshold of the scale (i.e., 15 workers) is based on the application of the Factory Act enacted in 1916.
-Panel B sorted the heads by their skill level and job type. Statistics are obtained from the Complete Survey on the 4th-6th graders in Tsukishima1st and 2nd Primary Schools in July 1919.
-Source: Panel A: Department of Health, Ministry of the Interior 1923a, pp. 386–387; 410–414. Panel B: Department of Health, Ministry of the Interior 1923a, p. 149.
+Panel B sorted the heads by their skill level and job type. Statistics are obtained from the Complete Survey on the 4th–6th graders in Tsukishima1st and 2nd Primary Schools in July 1919.
+Sources: Panel A: Department of Health, Ministry of the Interior 1923a, pp. 386–387; 410–414. Panel B: Department of Health, Ministry of the Interior 1923a, p. 149.
 A.6 Works in the Machinery Factories
@@
 Typical jobs for workers in machinery and equipment manufacturing plants are finishers (or fitters) and turners.
-A finisher is “a skilled worker responsible for finishing, assembling, and commissioning machine products” and is “one of the finest skilled workers because it requires skills in all aspects to correct defects in machine making” (Department of Health, Ministry of the Interior 1923a, p.
+A finisher is “a skilled worker responsible for finishing, assembling, and commissioning machine components” and is “one of the finest skilled workers because it requires skills in all aspects to correct defects in machine making” (Department of Health, Ministry of the Interior 1923a, p.
 393).
@@
 Becoming a full-fledged finisher and lathe operator takes at least five years.
-Generally, apprentices enter the workforce around 14–18, but they do not work in the same factory for many years within the apprenticeship system.
-After learning the skills for 1-2 years, they often move to another factory to improve their skills (Department of Health, Ministry of the Interior 1923a, p.
+Generally, apprentices enter the workforce at approximately 14–18, but they do not work in the same factory for many years within the apprenticeship system.
+After learning the skills for 1–2 years, they often move to another factory to improve their skills (Department of Health, Ministry of the Interior 1923a, p.
 394).
-Wages are paid on a piece-work basis or on a time basis.
-In Tsukishima, a time contract was adopted, and payment was based on the “residual time” method.
-Under the “residual time” method, “if the work is completed within the contracted time, the remaining time is considered the bonus.” For example, if 12 hours are contracted, and the work is completed in 10 hours, 2 hours are paid as a bonus (pp.
+Wages are usually paid on a piece-work basis or on a time basis.
+In Tsukishima, a time contract is broadly adopted, and payment is based on the “residual time” method.
+The “residual time” method means that “if the work is completed within the contracted time, the remaining time is considered a bonus.” For example, if $12$ hours are contracted and the work is completed in $10$ hours, $2$ hours are paid as a bonus (pp.
 396-397).
-Wages were paid by dividing one month into two periods (first and second semester).
-The average work hours are reported to be roughly 10 hours, with occasional all-night work (p.
+Wages are paid by dividing one month into two periods (the first and second half of the year).
+The average work hours are reported to be approximately ten hours, with occasional all-night work (p.
 402).
-Overtime hours were reportedly around 2 hours per day (Department of Health, Ministry of the Interior 1923a, p.
+Overtime hours are approximately two hours per day (Department of Health, Ministry of the Interior 1923a, p.
 402).
-Next, I summarize labor in the canning factory, which accounts for 24% of all Tsukishima factories.
+Next, I summarize labor in the canning factory, which accounts for $24$% of all the factories.
 The canning factory is engaged in shipbuilding, bridges, steel frames for construction, and other works, including fabrication and repair of gas reservoirs (p.
 406).
-Workers in a canning factory are called “boiler makers” and are typically involved in bending steel plates and joining them with rivets.
+Workers in a canning factory are called “boiler makers” and are typically involved in bending steel plates and jointing them with rivets.
 Workers are divided into rivetter, holder-up, plater, angle-ironsmith, and caulking.
-The work is done in groups, with the skilled angle-ironsmith and caulking.
+The work is usually done in groups, with the skilled angle-ironsmith and caulking.
 The “bōsin” of the rivetter gives the instructions as the head of the group.
@@
 Smithing is the second most common industry, accounting for $28$% of the total.
-Panel A of Table A.6 reveals that most of the smithing industry is relatively small scale.
+Panel A of Table A.6 reveals that most of the smithing industry is relatively small in scale.
 They make boulders, nuts, rivets, lathe stops, jigs, and spanners for can manufacturing and construction.
-They were subcontractors who manufactured metal parts to supply the larger factories in Tsukishima (Department of Health, Ministry of the Interior 1923a, p.
+They are subcontractors who manufacture metal parts for the larger factories in Tsukishima (Department of Health, Ministry of the Interior 1923a, p.
 412).
 There are three types of jobs: yokoza, sakite, and tatara-fuki (bellows-blower).
-The yokoza is self-employed skilled worker (oyakata), while the other two are young unskilled laborers.
-In small operations, wages were based on a percentage based on voluntary business discussions with the proprietor, and working hours were not regular (pp.
+The yokoza is a self-employed skilled worker (called oyakata), while the other two are young unskilled laborers.
+In small operations, wages are based on a percentage based on voluntary business discussions with the proprietor, and working hours are not regular (pp.
 411–413).
@@
 B.1 Tsukishima Survey: The First Social Survey in Japan
-Kawai (1980) summarizes the history of social research methodology.
-The methodology of social research in Japan was formed in the 1930s.
-Social research methodologies had been developed in the 1910s and 1920s (Kawai 1980, pp.
-47–48).
-During this period, “numerous social surveys were conducted by various survey entities, including government agencies, research groups, local administrative bodies, private organizations, survey research institutes, and researchers” (p.48).
-This is because “around WWI, there were many problems in labor, lifestyle, urban, health and sanitation, public safety, and civic due to the heavy and chemical industrialization, the monopolization of capitalism, and urban expansion.” The Report on Field Survey in Tsukishima, Kyobashi Ward, Tokyo (called Tsukishima Survey) [^92] Takano himself, who was responsible for overseeing the survey, refers to this field survey as the “Tsukishima Survey” (Department of Health, Ministry of the Interior 1923a, p.
+According to Kawai (1980), the methodology of social research in Japan was formed in the 1930s.
+During the 1910s and 1920s, “numerous social surveys were conducted by various survey entities, including government agencies, research groups, local administrative bodies, private organizations, survey research institutes, and researchers.” He saw that circa WWI, there were many problems in labor, lifestyle, urban, health and sanitation, public safety, and citizens due to the heavy and chemical industrialization, the monopolization of capitalism, and urban expansion (p.48).
+The Report on Field Survey in Tsukishima, Kyobashi Ward, Tokyo (called Tsukishima Survey) [^88] Takano himself, who was responsible for overseeing the survey, refers to this field survey as the “Tsukishima Survey” (Department of Health, Ministry of the Interior 1923a, p.
 6).
-Considering this, I call the Tsukishima Survey throughout this paper.
-used in this study is the most pioneering of these surveys of urban communities.
-Sekiya (1970, p.
-43) regards the Tsukishima Survey as a pioneering household survey in Japan.
+Considering this, I use the Tsukishima Survey as the name of this survey throughout this paper.
+used in this study is the most pioneering survey of urban communities (Sekiya 1970, p.
+43).
 First, I summarize the investigators, research questions, and subjects of the Tsukishima Survey.
@@
 33).
-He had the experience of conducting the first household survey in Japan in 1916.939393This is called the “Survey of Household Income and Expenditure of 20 Artisans and Workers in Tokyo” and was inspired by Charles Booth (Miyoshi 1980, p.
+He had the experience of conducting the first household survey in Japan in 1916898989This is called the “Survey of Household Income and Expenditure of 20 Artisans and Workers in Tokyo,” which was inspired by Charles Booth (Miyoshi 1980, p.
 33).
-Although the original documents of this survey were preserved at Tokyo Imperial University, those were destroyed in the Great Kantō Earthquake (Takano 1933, pp.
+The original documents of this survey, preserved at Tokyo Imperial University, were destroyed in the Great Kantō Earthquake (Takano 1933, pp.
 727–728).
-More precisely, before this household survey, Takano conducted a field survey in a poor area of Tokyo when he was a student at Tokyo Imperial University.
+Before this household survey, Takano conducted a field survey in a poor area of Tokyo when he was a student at Tokyo Imperial University.
 The findings were reported in April 1894 as an exercise entitled “East London in Tokyo” (Kawai 1981b, p.
 3).
-Takano has been a member of the seventh subcommittee (Rural Sanitation Subcommittee) of the Ministry of Home Affairs’ Health and Sanitation Investigation Committee since 1916.949494The Health and Sanitation Investigation Committee of the Ministry of Home Affairs was established on June 27, 1916, under the second Shigenobu Okuma Cabinet by issuing Imperial Ordinance No.
+He was a member of the seventh subcommittee (Rural Sanitation Subcommittee) of the Ministry of Home Affairs’ Health and Sanitation Investigation Committee since 1916.909090The Health and Sanitation Investigation Committee of the Ministry of Home Affairs was established on June 27, 1916, under the second Shigenobu Okuma Cabinet by issuing Imperial Ordinance No.
 172.
 Initially, 34 members were appointed, and the first meeting was held on July 8.
-The reason for the establishment of this committee was based on the need for a policy response regarding the establishment and institutionalization of various investigation organizations and the penetration of the capitalist economy during the Meiji period.
-Another reason is the requirement for a public health response to the rising mortality rate, exemplified by the tuberculosis epidemic (Kawai 1981a, pp.
+The reason for the establishment of this committee was the need for a policy response regarding the establishment and institutionalization of various investigation organizations and the penetration of the capitalist economy during the Meiji period.
+Another reason was the requirement for a public health response to the rising mortality rate, exemplified by the tuberculosis epidemic (Kawai 1981a, pp.
 10–13).
-In 1917, the Seventh Section evolved into the Section on Urban and Rural Sanitary Conditions (Kawai 1981a, pp.
+In 1917, the seventh subcommittee evolved into the Section on Urban and Rural Sanitary Conditions (Kawai 1981a, pp.
 12–13).
-On October 22, 1918, Takano submitted a proposal to the subcommittee for a field survey of urban sanitary conditions, which was approved.
+On October 22, 1918, Takano submitted a proposal to the subcommittee for a field survey of urban sanitary conditions.
 The survey aimed to determine the housing conditions, household budget status, and health status (child health, production, stillbirths, deaths, and diseases) in the areas where skilled-worker families congregated (Sekiya 1970, p.
@@
 Yoshino approached his former student Kotora Tanahashi, who had become a member of the fraternity club, a mutual aid organization for workers, to accept the investigator position.
-The diary written by Kotora Tanahashi (“Tanahashi Kotora Diary”) is in the possession of the OISR, and Takeda used the diary as the basis for her analysis of the circumstances surrounding the change of survey location.
-Tanahashi eventually recommended Yoshitsuru Yamana, who was his classmate in high school and Tokyo Imperial University, as a researcher and considered Tsukishima to be the best research area because of the ease of selecting survey households using the fraternity’s network (Takeda 2015, pp.
-76–77).
-On October 26, Tanahashi and Yamana, together with the Kyobashi Branch of the fraternity club, conducted an on-site inspection of Tsukishima.
+The diary written by Kotora Tanahashi (“Tanahashi Kotora Diary”) is in the possession of the OISR, and Takeda (2015, pp.
+76–77) used the diary to analyze the circumstances surrounding the change of survey location.
+Tanahashi eventually recommended Yoshitsuru Yamana, who was his classmate in high school and Tokyo Imperial University, as a researcher and considered Tsukishima to be the best research area because of the ease of selecting survey households using the fraternity’s network.
+On October 26, Tanahashi and Yamana—together with the Kyobashi Branch of the fraternity club—conducted an on-site inspection of Tsukishima.
 Based on this inspection, Tanahashi and Yamana approached Takano about locating the survey station in Tsukishima.
-Finally, Takano and Yamana conducted an on-site inspection of Tsukishima on October 30 and decided on Tsukishima as the survey site.
-Although the statistical method of sampling was not yet common at that time, Takano selected the survey sites based on his visionary thinking.
-In fact, “There can be no dispute that we have chosen Tokyo as the target of our metropolitan survey.
+Finally, Takano and Yamana conducted an on-site inspection there on October 30 and decided on Tsukishima as the survey site.
+Although the statistical sampling method was not yet common at that time, Takano selected the survey sites based on his visionary thinking.
+He stated that “there can be no dispute that we have chosen Tokyo as the target of our metropolitan survey.
 However, it is impossible to survey the entire Tokyo area.
@@
 2–3).
-Takano further explained the need to focus on the working-class, which makes up most of the citizens, rather than to see the political and commercial center of Tokyo: “Most of Tokyo’s citizens are wage earners who subsist on labor.
+Takano further explained the need to focus on the working-class, which makes up most of the citizens, rather than to see the political and commercial center: “Most of Tokyo citizens are wage earners who subsist on labor.
 As is usual in large cities, a large number of people are engaged in manual labor in Tokyo.
@@
 Therefore, the area where many families of skilled workers congregate can be regarded as a representative area of Tokyo.
-And it seemed to me that a study of the social and sanitary conditions in that area could explain a large part of the social and sanitary conditions in Tokyo.” (Department of Health, Ministry of the Interior 1923a, pp.
+Further, it seemed to me that a study of the social and sanitary conditions in that area could explain a large part of the social and sanitary conditions in Tokyo.” (Department of Health, Ministry of the Interior 1923a, pp.
 3).
-It is necessary to delve into the point where Takano states that “the most genuine type of worker is the skilled workers” and the fact that the main target of the Tsukishima Survey was skilled workers.
+It is necessary to investigate the point where Takano states that “the most genuine type of worker is the skilled workers” and the fact that the main target of the Tsukishima Survey was skilled workers.
 [^92] In other words, the Tsukishima Survey was not designed for poor households.
-Therefore, unlike the surveys conducted in the UK by C.
-Booth and B.S.
-Rowntree, no standard of living was set in the survey (Sekiya 1970, p.
+Therefore, unlike the surveys conducted in the UK by Booth and Rowntree, investigating the standard of living was not the main objective of the survey (Sekiya 1970, p.
 42).
@@
 13–18) is helpful in this regard.
-As summarized in Section 2, Japan’s heavy and chemical industrialization throughout WWI led to a large concentration of city workers.
-As a result, the ratio of small-scale conventional factories based on “master employees” as is often observed in the textile industry decreased.
-On the other hand, the number of factories with a direct “factory owner-employee” employment relationship increased, particularly in the machinery and equipment, metal, and chemical industries.
+As summarized in Section 2, Japan’s mechanical and chemical industrialization throughout WWI led to a large concentration of city workers.
+As a result, the ratio of small-scale conventional factories based on “master employees” (as is often observed in the textile industry) decreased.
+Conversely, the number of factories with a direct “factory owner-employee” employment relationship increased, particularly in the machinery and equipment, metal, and chemical industries.
 These workers in the modern industrial sector accumulated skills, formed households, and settled in the cities.
 It can be said that a new urban worker lifestyle was being formed.
-In short, although the city had a class of workers engaged in traditional industries, a new type of workers in the modern industrial sector, and a lower class, Takano believed that the skilled workers in the modern industrial sector would play a central role in terms of the labor force, labor movement, and urban life.
+In short, although the city had a class of workers engaged in traditional industries, Takano believed that the skilled workers in the modern industrial sector would play a central role in terms of the labor force, labor movement, and urban life.
 In addition, Sekiya notes the rise of the labor movement.
-As noted in Section 2, while wages rose throughout WWI, the price surge caused relative wages to decline.
+As explained in Section 2, while wages rose throughout WWI, inflation caused relative wages to decline.
 In 1916, Japan’s first labor protection legislation, the Factory Law, was enacted.
-However, the number of working days and hours did not decrease, and the number of labor disputes steadily increased (Sekiya 1970, pp.
+However, the number of working days and hours did not decrease.
+Consequently, the number of labor disputes steadily increased (Sekiya 1970, pp.
 24–26).
@@
 26–40) speculates that Takano’s intention in conducting the Tsukishima Survey was to obtain basic data for solving labor problems, specifically to identify the scope of matters to be solved by workers on their own initiative (or covered by labor protection legislation).
-His insights suggest that this historical background should be addressed as a reason for Takano’s focus on skilled workers.
 Kawai (1981b, pp.
-15; 17) assessed the 40 households covered by the household survey included in the Tsukishima Survey as “the representativeness of the sample has remained ambiguous.” There is some mention but no clear description of how the sample households were characterized among the worker households in Tsukishima.
-It is reported that “to find suitable persons to fill in the form, we must ask for the assistance of school principals, police officers, physicians, and laborers, and we have requested a meeting of these persons at the survey office to explain the purpose of the survey, and to solicit applicants for the distribution of the household account book” (Department of Health, Ministry of the Interior 1923a, pp.
+15; 17) assessed the $40$ households covered by the household survey included in the Tsukishima Survey as “the representativeness of the sample has remained ambiguous.” There are no systematic descriptions of how the sample households were characterized among the worker households in Tsukishima.
+The official report stated that “to find suitable persons to fill in the form, we must ask for the assistance of school principals, police officers, physicians, and laborers, and we have requested a meeting of these persons at the survey office to explain the purpose of the survey, and to solicit applicants for the distribution of the household account book” (Department of Health, Ministry of the Interior 1923a, pp.
 8–9).
 As explained, the target households were not randomly selected because the statistical sampling method was not prevalent.
-However, it is plausible that Takano designed this household budget survey in the sampling framework to capture the representative households in Tsukishima.
-As described in Section 3, the THBS households can be regareded as the standard skilled factory worker households in Tsukishima.
+However, as described in Section 3, the THBS households can be regarded as standard skilled factory worker households in Tsukishima.
+Therefore, it is plausible that Takano designed this household budget survey in the sampling framework to capture representative households in Tsukishima.
 Table B.1: Survey Categories in the Field Study of Tsukishima Survey
@@
 10).
-This was a pioneering attempt at the time and had influenced subsequent urban studies (Kawai 1980, pp.
-61-62).
+This was a pioneering attempt at the time and influenced subsequent urban studies (Kawai 1980, pp.
+61–62).
 The latter was a field survey including $14$ survey items (Table B.1).
@@
 social mapping and photography (appendix).
-The survey was conducted over two years, generally from November 1918 to November 1920, although the timing varied depending on the survey items (Table B.1).
-A unique feature of the survey methodology is using young researchers as commissioned researchers.
-The household budget survey and leisure life survey were conducted by Yasunosuke Gonda, a faculty member of the Doitsugaku kyōkai gakkō (Association for German Sciences), who had researched activity photography and entertainment.
+The survey was conducted over two years, from November 1918 to November 1920, although the timing varied depending on the survey items (Table B.1).
+A unique feature of the survey methodology is the use of young researchers as commissioned researchers.
+The household budget survey and leisure life survey were conducted by Yasunosuke Gonda, a faculty member of the Doitsugaku kyōkai gakkō (Association for German Sciences), who had researched cinematograph and entertainment.
 Tetsuo Hoshino, a medical doctor, was responsible for the sanitary survey.
-The labor survey was handled by Yoshitsuru Yamana, a graduate of Tokyo Imperial University who contributed to the organization of labor unions.
-In addition, Toyotaro Miyoshi, a sociology student at the Tokyo Imperial University, and others became temporary employees.
-In other words, although the Tsukishima Survey was initially planned as an insurance and sanitation survey by the Ministry of Home Affairs, in reality, it was a long-term survey conducted as a joint study by young researchers in different fields, led by Takano (Kawai 1980, p.
-63).
-At the time, the Health Bureau of the Ministry of the Interior was conducting several sanitary surveys of rural areas.
+The labor survey was managed by Yoshitsuru Yamana, a graduate of Tokyo Imperial University who contributed to the organization of labor unions.
+In addition, Toyotaro Miyoshi, a sociology student at the Tokyo Imperial University, and others became temporary employees (see footnote 55).
+At the time, the Health Bureau of the Ministry of the Interior conducted several sanitary surveys of rural areas.
 Takano may have designed the urban sanitary survey from this background (Kawai 1981a, pp.
@@
 Note: Figures B.1a and B.1b show the outside and inside of the survey office in Tsukishima, respectively. Source: Department of Health, Ministry of the Interior 1923a, figures 1 and 2.
-Takano stated, “To be as complete as possible, it is best to continue the survey over as long a period as possible and to obtain accurate facts by recording the workers’ household income and expenditures over a long period.” For this reason, he believed that “the best method of investigation would be to establish a survey station directly in the survey area and to have a full-time investigator reside there as much as possible, familiarize himself with the area and the people, and conduct the field survey” (Department of Health, Ministry of the Interior 1923a, p.
+Takano stated, “To be as complete as possible, it is best to continue the survey over as long period as possible and to obtain accurate facts by recording the workers’ household income and expenditures.” For this reason, he believed that “the best method of investigation would be to establish a survey station directly in the survey area and to have a full-time investigator reside there as much as possible, familiarize himself with the area and the people, and conduct the field survey” (Department of Health, Ministry of the Interior 1923a, p.
 4).
-In fact, through the mediation of the Tsukishima police station, a survey office was rented at 9-3 Higashinakadori, Tsukishima, where Yoshitsuru Yamana moved in from November 1918 and contacted worker households (Department of Health, Ministry of the Interior 1923a, p.
+In fact, through the mediation of the Tsukishima police station, a survey office was rented at 9-3 Higashinakadōri, Tsukishima, where Yoshitsuru Yamana moved in from November 1918 and contacted worker households (Department of Health, Ministry of the Interior 1923a, p.
 7; Takeda 2015, p.
@@
 Figure B.1 shows the photographs of the survey office.
-It is also important to note that Takano and his team held weekly meetings at the research center during the survey period.
-Toyotarō Miyoshi, who was an assistant in the survey, recalls the situation at that time as follows:
+It is also important to note that Takano and his team held weekly meetings at the office during the survey period.
+Toyotaro Miyoshi, who was an assistant in the survey, recalls the situation at that time as follows:
 After the start of the survey office, all the members gathered once a week with Iwasaburo Takano present to discuss the implementation and progress of the survey and administrative matters.
@@
 Even now, I can recall the warm face of sensei (i.e., Professor Takano) of those days.
-It was always a friendly, laboratory-like atmosphere in which the C.
-Booth and S.
-Rowntree investigations were occasionally discussed.
+It was always a friendly, laboratory-like atmosphere in which the Booth and Rowntree investigations were occasionally discussed.
 Miyoshi (1980, p. 38)
@@
 Kawai (1981b, p.
-28) assessed that there were no interconnections among the survey items mentioned above and that the survey “ended up being several fragmentary surveys rather than a comprehensive and systematic survey of workers’ lives.” This difficulty may be why previous studies have not frequently used the Tsukishima Survey.
+28) assessed that there were no interconnections among the survey items mentioned above and that the survey “ended up being several fragmentary surveys rather than a comprehensive and systematic survey of workers’ lives.” This difficulty may explain the reason why previous academic studies have rarely used the Tsukishima Survey.
 However, as already explained, Takano did not intend to conduct a large-scale social survey in Tsukishima in the first place.
-Instead, as he noted, the best method is to investigate a representative area “narrowly but deeply,” the goal was to improve the quality of the investigation rather than the scale of it.
-In this light, the fact that the researcher lived in the survey area and was constantly organizing and revising the survey results should be appreciated.
-Even though the Tsukishima Survey is the early social survey in Japan, it is of high quality as a prewar household survey.
+Instead, as Takano noted, the best method is to investigate a representative area “narrowly but deeply.” The goal was to improve the quality of the investigation rather than the scale of it.
+In this light, the fact that the researchers lived in the survey area and were constantly organizing and revising the survey results should be appreciated.
+Although the Tsukishima Survey is an early social survey in Japan, it is undoubtedly of high quality.
 Figure B.2: Official Published Report of the Tsukishima Survey
@@
 The first section is a comprehensive review by Takano and outlines the survey.
-The second section summarizes the living conditions of the workers and was written by Gonda (partly by Yamana).
-Hoshino wrote the third section and summarized the sanitary conditions of workers.
+The second section summarizes the living conditions of the workers, and was written by Gonda (partly by Yamana).
+Hoshino wrote the third section.
+It summarizes the sanitary conditions of workers.
 In the fourth section, Yamana summarizes the labor situation in Tsukishima.
-As for the two appendices, one contains statistical tables related to the above report, and the other organizes social maps and photographs.
+There are two appendices.
+One contains statistical tables related to the above report, and the other organizes social maps and photographs.
 The correspondence between the survey items and the report is summarized in Table B.1.
-The Ministry of Home Affairs has already published these materials, which are available at the Gonda Yasunosuke Library of the OISR and the University of Tokyo Library.979797 In the OISR, these are included in the Archives of Gonda Yasunosuke (7-2; 7-3; 7-4).
-In addition, a second manuscript of the Tsukishima Survey report exists.
+The Ministry of Home Affairs has already published these materials, which are available at the Gonda Yasunosuke Library of the OISR and the University of Tokyo Library.
+[^93] In the OISR, these are included in the Archives of Gonda Yasunosuke (7-2; 7-3; 7-4).
+In addition, there is a second manuscript of the Tsukishima Survey report.
 This is the manuscript of the second series, which was preserved in the home of Gonda Hayao, the second son of Gonda Yasunosuke, and was discovered by Terade Koji.
@@
 B.2 Original Micro-spreadsheets from the Household Budget Survey in Tsukishima Survey
-To analyze the consumption smoothing behavior of THBS households, it is necessary to use the original micro-spreadsheets.
-The official report does not include household-level information, and a large part of the original forms used to prepare the report have little survived.
+The official report does not include household-level information, and only a small quantity of the original forms used to prepare the report have survived.
 Fortunately, however, some of the original micro-spreadsheets are kept in the archives of the OISR because Takano was its first director.
 This study uses the original micro-spreadsheets from the household budget survey (No.
-9 in Table B.1), which was used in writing the second section of the official report by Gonda.
+9 in Table B.1), which was used by Gonda to write the second section of the official report.
 While they have yet to be publicly available, I received permission from the OISR to use the original forms stored in the archives: the OISR, Archives of the Tsukishima Survey (THBS, unreleased).
@@
 Mika Nakamura, an archivist at the institute, for her careful assistance during the research.
-Figure B.3: An Example Page of the Budget Book in the Tsukishima Household Budget Survey
-Notes: This photograph depicts an example page from the household expenses account of the THBS.
+Figure B.3: An Example Page of the Budget Book
+in the Tsukishima Household Budget Survey
+Notes: This photograph depicts a sample page from the household expenses account of the THBS.
 The right page documents the details of costs.
 The left page includes the details of revenue and receiving and making payments on the goods (actual things, including gifts).
-Source: The THBS, account #6, 15th February 1919. The tone was adjusted by the author using Adobe Photoshop 24.7.0.
+Source: The THBS, account #6, 15th February, 1919. The tone was adjusted by the author using Adobe Photoshop 24.7.0.
 The budget book used for the household budget survey is named the Kinsendeiri hikaechō (money receipts and disbursements account book).
-This budget book is designed based on the form used in Takano’s first household survey (Online Appendix B.1).
+The design of this budget book was based on the form used in Takano’s first household survey (Online Appendix B.1).
 [^95] Department of Health, Ministry of the Interior (1923a, p.
 152).
-The budget book is to be filled out daily and includes four sections: income, expenses, receipts from credit purchases, and receipts and disbursements of actual goods.
-Most of the entries are in the income and expenses sections, indicating that living was basically conducted through wage receipt and consumption in cash.
+The budget book was to be filled out daily and included four sections: income, expenses, receipts from credit purchases, and receipts and disbursements of actual goods.
+Most of the entries were in the income and expenses sections, indicating that their daily living was conducted through wage receipt and consumption in cash.
 Figure B.3 shows an example of the budget book.
-As described, the method used to select households for the survey was not recorded in detail (Section B.1).
-According to the report, they initially recruited just over 90 households through the factory, police station, and elementary school in Tsukishima.
+As described, the method used to select households for the survey was not recorded in detail (Online Appendix B.1).
+According to the report, they initially recruited just over $90$ households through the factories, police station, and elementary schools in Tsukishima.
 Of these, $50$ households completed the book for at least one month, but $10$ were deemed incomplete and not included in the report.
-In addition to the 40 budget books used in preparing the report, therefore, the OISR has a set of other budget books marked as “discarded”.
-According to my survey, about six incomplete books of the worker households have been determined to be in this “discarded” category.
-Therefore, based on the remaining conditions of the books, it is plausible to believe that roughly 50 households participated in the survey.
-Of the 40 households included in the report tally, two households had recorded for more than one year, 11 households had recorded for 6–12 months, and the remaining 27 households had registered for less than six months (Department of Health, Ministry of the Interior 1923a, pp.
+In addition to the $40$ budget books used to prepare the report, therefore, the OISR has another set of budget books marked as “discarded.” According to my survey, about six incomplete books of the worker households have been determined to be in this “discarded” category.
+Therefore, based on the condition of the remaining books, it is plausible to believe that approximately $50$ households participated in the survey.
+Of the $40$ households included in the report tally, two households maintained records for more than one year, $11$ households for 6–12 months, and the remaining $27$ households registered for less than six months (Department of Health, Ministry of the Interior 1923a, pp.
 152–153).
-Initially, the survey period was targeted to be six months, but the average bookkeeping period for the 40 households was four months.
+Initially, the survey period was intended to be six months, but the average bookkeeping period for the $40$ households was four months.
 B.3 Trimming
@@
 Strictly, semi-monthly, and monthly datasets include both panel and cross-sectional units because aggregation shall trim fractions in the edges of the survey period.
-Figure B.4b shows 35-panel units and 5 cross-sectional units (#29; #31-33; #35).
-Figure B.4c shows 28-panel units and 6 cross-sectional units (#20; #25; #26; #28; #34; #40).
-5 cross-sectional units in the semi-monthly panel do not appear in this figure (#29; #31-33; #35).
+Figure B.4b shows 35-panel units and five cross-sectional units (#29; #31-33; #35).
+Figure B.4c shows 28-panel units and six cross-sectional units (#20; #25; #26; #28; #34; #40).
+Five cross-sectional units in the semi-monthly panel do not appear in this figure (#29; #31-33; #35).
 Note that unit #30 having two semi-monthly cells is not included because those cells are separated into different adjusted months (February and March, respectively).
@@
 Although the official report by Gonda reported the average monthly income for this household (Department of Health, Ministry of the Interior 1923a, p.
-183), no documentation would allow me to know how this average income was calculated.
+183), no documentation enabled me to determine how this average income was calculated.
 Hence, I excluded this household’s budget book.
 Similarly, a household does not include information on the payments to the credit purchases (household #39).
-Although this household’s budget book documented the amount purchased in credit purchases, I excluded this household’s budget book because the expenditure series becomes systematically smaller due to the omission of the payment information.
+Although this household’s budget book documented the amount purchased in credit purchases, I excluded this household’s budget book because the expenditure series became systematically smaller due to the omission of the payment information.
 Second, the final pages of the budget books are sometimes omitted.
 This may be because the vendor who bound the budget books accidentally truncated the last page.
-All these incomplete sheets can not be included in the analysis.
+All these incomplete sheets could not be included in the analysis.
 Third, several households did not complete their budget books in the first and/or last month of the survey period.
-For example, the budget book of household #1 has statistics from 1st January 1919 to 12th January 1920.
-Similarly, the book of household #29 contains statistics from 26th January to 23rd February 1919.
+For example, the budget book of household #1 has statistics from January 1, 1919 to January 12, 1920.
+Similarly, the book of household #29 contains statistics from January 26 to February 23, 1919.
 In semi-monthly and monthly datasets, these censored budget books may not be included in the aggregation.
-In the latter case, for instance, the budget sheets of 26th January and after 11th February 1919 are not included in a semi-month cell because the semi-month cell is defined from the 27th to 11th of the next month in this case.
-Thus, household #29 only has a cross-sectional observation and can not be included in my analytical sample.
+In the latter case, for instance, the budget sheets of January 26 and after February 11, 1919, are not included in a semi-month cell because the semi-month cell is defined from the 27th to the 11th of the next month, in this case.
+Thus, household #29 only has a cross-sectional observation and could not be included in my analytical sample.
 Figure B.4 summarizes the structure of the $40$ THBS households.
@@
 Therefore, there are 35 units with a panel structure.
-After excluding two units without complete information on income and credit purchases (#37; #39), I get the 33 THBS households as the analytical sample for my panel data analysis.
-Regarding the time dimension, I trimmed all the observations before 12th January 1919 and after 11th November 1919 because the number of surveyed households is substantially small.
+After excluding two units without complete information on income and credit purchases (#37; #39), I obtained the 33 THBS households as the analytical sample for my panel data analysis.
+Regarding the time dimension, I trimmed all the observations before January 12, 1919, and after November 11, 1919, because the number of surveyed households is substantially small in the edge periods.
 For example, Figure B.4b shows that the number of cross-sectional units is three or less in both periods, causing substantial fluctuations in the net income.
-Consequently, my analytical sample includes 33 households measured between 12th January and 11th November 1919.
+Consequently, my analytical sample includes $33$ households measured between January 12 and November 11, 1919.
 The average number of semi-months per unit is $8.8$ (Std.
@@
 The characteristics of the $33$ THBS households are summarized in detail in Section 3.
-Regarding the adjusted monthly series, we further exclude several units that have insufficient observations for constructing the monthly panel: 26 THBS households are included in my adjusted monthly panel dataset.
+Regarding the adjusted monthly series, I further exclude several units that have insufficient observations for constructing the monthly panel: 26 THBS households are included in my adjusted monthly panel dataset.
 Figure B.4c shows the observations in the adjusted monthly dataset.
-B.4 Testing the Potential Influence of the Unbalancedness
+B.4 Testing the Potential Influence of the Lack of balance
 Table B.2: Results for the Balancing Tests
@@
 In this subsection, I test whether attrition leads to selection bias in the THBS sample.
-If the difference in the preference for consumptions predicted the large attritions, there must be statistically significant correlations between the family size variables and attritions.
+If the difference in the preference for consumption predicted large attritions, there must be statistically significant correlations between the family size variables and attrition.
 In other words, the units with shorter (longer) time-series observations have different preferences from those with longer (shorter) time-series observations.
@@
 Therefore, the family characteristics are similar between the households with short-term observations and those with long-term observations.
-This result supports the evidence that the unbalancedness in the THBS dataset is unlikely to lead to selection bias.
+This result supports the evidence that the lack of balance in the THBS dataset is unlikely to lead to selection bias.
 B.5 Comparing Distributions of the Household Size
 Figure B.5 illustrates the household size distributions measured in the census and THBS sample.
-Pearson chi-squared test does not reject the null hypothesis of the equality of distributions with $p$-value of $0.948$.
-This supports the evidence that the THBS sample has a similar household size distribution to the population of Tsukishima (Section 3.1).
-As described in Section 3.2, while the THBS sample does not cover households with one person and with a very large number of people, it covers most of the classes in the household sizes with reasonable approximation to the population distribution.100100100The number of households with one people is $362$, accounting for only $6$% of the total number of households.
+The Pearson chi-squared test does not reject the null hypothesis of the equality of distributions with $p$-value of $0.948$.
+This supports the evidence that the THBS sample has a similar household size distribution to that of the entire Tsukishima population (Section 3.1).
+As described in Section 3.2, while the THBS sample does not cover households with one person and with a very large number of people, it covers most of the classes in the household sizes with reasonable approximation to the population distribution.
+[^96] The number of households with one people is $362$, accounting for only $6$% of the total number of households.
 Similarly, the number of households with nine or more people is $294$, accounting for only $5$% (Section 3.2).
@@
 Population Census v. the THBS sample
-Notes: This figure shows the share of the household in each household size category to the total number of households in Tsukishima measured in the 1920 Population Census and the THBS samples.
-Pearson chi-squared test does not reject the null hypothesis of the equality of distributions with $p$-value of $0.948$.
+Notes: This figure shows the distribution of household size measured in the 1920 Population Census (red) and the THBS sample (green).
+The Pearson chi-squared test does not reject the null hypothesis of the equality of distributions with $p$-value of $0.948$.
 The share in each household size measured in the census is used as the theoretical probability in calculating $\chi^{2}$ test statistic (Pearson 1900).
@@
 B.6 Estimating Average Head’s Monthly Income
-Comprehensive estimation effort on the time series figures of average wages in prewar Japan was done in the Long-Term Economic Statistics (hereafter LTES).
-Specifically, the LTES estimates the average national salary series of “general occupations,” “manufacturing sector,” and “manufacturing occupations (by medium classification).” For the manufacturing occupations category, the LTES reports the average wage per day of factory workers in Tokyo City and Osaka cities from 1917 to 1922 (Ohkawa et al.
+The time-series figures of average wages in prewar Japan were provided by the Long-Term Economic Statistics (hereafter LTES).
+Specifically, the LTES estimates the average national salary series of “general occupations,” “manufacturing sector,” and “manufacturing occupations (by medium classification).” For the manufacturing occupations category, the LTES reports the average wage per day of factory workers in Tokyo and Osaka cities from 1917 to 1922 (Ohkawa et al.
 1967, p.
@@
 In this section, I use available wage and labor statistics to estimate the average monthly income of male factory worker households in Tokyo and compare it to the THBS heads’ average monthly earnings.
-To clarify the problem, I first estimate average monthly earnings most simply.
-Specifically, I use the ASCT (volume 17), which provides wage statistics as of the end of 1918, as a case study.
+To clarify the problem in estimating the monthly earnings, I first estimate average monthly earnings in the simplest way.
+To do so, I use the ASCT (Volume 17), which provides wage statistics as of the end of 1918, as a case study.
 For each industrial sector, the ASCT (Volume 17) provides the “average daily wage of male factory workers aged 15 and over,” “the number of male factory workers aged 15 and over,” and “average number of working days per year” for four factory categories: factories with ten or fewer workers, and factories using or not using motive power.
-Using the “number of workers over 15 years old” in each factory category cell ($c$) as weights, I can get the average wage per month for a worker, for example in the machinery sector, as:
-$\textit{Wage}^{15+}_{Machine}=\frac{\sum_{c}\{(\textit{Daily Wage}^{15+}_{Machine,c}\times\textit{Working Days}_{Machine,c})\times\textit{Workers}^{15+}_{Machine,c}\}}{12\sum_{c}\textit{Workers}^{15+}_{Machine,c}}.$ (5)
-This value is estimated to be $27.8$ yen.
+Using the “number of workers over 15 years old” in each factory category cell ($c$) as weight, I can obtain the average wage per month for a worker in each industrial sector.
+For example, in the manufacturing sector, this can be calculated as:
+$\textit{Wage}^{15+}_{Machine}=\frac{\sum_{c}\{(\textit{Daily Wage}^{15+}_{Machine,c}\times\textit{Working Days}_{Machine,c})\times\textit{Workers}^{15+}_{Machine,c}\}}{12\sum_{c}\textit{Workers}^{15+}_{Machine,c}},$ (5)
+which is estimated to be $27.8$ yen.
 This is the most straightforward estimate using daily wages, the number of workers, and the number of working days.
@@
 The sample mean is $33.2$ years old, which is unchanged if I exclude a head aged more than $60$ (i.e., $32.1$ years old).
-This means that I have to refer to the average daily wage for the 30s to 40s for comparison.
-However, since many workers are in their 20s in the distribution of entire workers, the average wage reported in the ASCT shall be considerably lower than the average wage for factory workers around 30 years old.
-Thus, I need the wage profiles by age to correct for this downward bias.101101101 The ASCT (volume 18), which contains statistics for 1919, does not list such wages by age group but only average wages.
+This means that I should refer to the average daily wage for the 30s to 40s for comparison.
+However, since many workers are in their twenties in the distribution of entire workers, the average wages reported in the ASCT are considerably lower than the average wage for factory workers at approximately 30 years old.
+This means that I need the wage profiles by age to correct this downward bias.
+[^97] The ASCT (Volume 18), which contains statistics for 1919, no longer lists such wages by age group (average wages only).
 As is clear, the same problem arises in this case.
-Second, there is no consideration of ancillary pay such as bonuses and various allowances.
-As explained later, these wages shared a non-negligible proportion of the head’s earnings.
-Thus, estimation without considering ancillary wages would have a strong downward bias.
-In the following, I summarize the procedure for correcting these downward biases.
+Second, there is no consideration of ancillary pay such as bonus and allowance.
+As explained later, these wages share a non-negligible proportion of the head’s earnings.
+Thus, estimation excluding ancillary wages would have a strong downward bias.
+In the following subsections, I summarize the procedure for correcting these factors leading to the downward bias in the monthly income estimation.
 B.6.1 Daily Wage Distribution by Ages
-Since the early wage surveys focused on average wages, official statistics in the 1910s do not include average wage statistics by age (bin).
-By the “Statistical Data Field Survey Law” (Law No.
+As the early wage surveys focused only on the average wage for entire workers, official statistics in the 1910s never included average wage statistics by age (bin).
+In the “Statistical Data Field Survey Law” (Law No.
 52) in April 1922, average wages by age group were surveyed for the first time in the “Labor Statistics Field Survey” in October 1924 conducted by the Social Affairs Bureau of the Ministry of Home Affairs (Committee for the History of the Labor Movement 1964, p.
@@
 Figures B.7a and B.7b show the distributions under 1918 and 1919 prices, respectively.
-The weighted average daily wage for a 30-39-year-old bin is shown in red.
+The weighted average daily wage for a 30–39-year-old bin is shown in red.
 The weighted averages for all and 14+ bins are shown in gray and blue lines, respectively.
@@
 For the 1919 price data, the fitted model is $\hat{y}^{\textit{Daily Wage}}=-13.18+5.47a^{1/2}-0.43a^{2}$ with $R^{2}=0.99$.
-Sources: Created by the author using the Tokyo City Statistics Division (1926a, pp. 16; 244-245). Data on the consumer price index for the cities are obtained from Ohkawa et al. (1967, p. 135).
+Sources: Created by the author using the Tokyo City Statistics Division (1926a, pp. 16; 244–245). Data on the consumer price index for the cities are obtained from Ohkawa et al. (1967, p. 135).
 Based on the LST, Figure B.7 shows the average daily wages of male factory workers in Tokyo City’s machinery and equipment manufacturing industry by age group.
@@
 The weighted average for ages $14$ and older is $2.58$ yen, which is about the same as the weighted average for daily wages of $2.56$ yen.
-The $2.58$ yen is between the average values for the 20-24 age bin and the 25-29 age bin.
-Thus, it likely to represent the daily wage of workers around 25.
-This is consistent with the fact that workers are concentrated in the 20s.
-However, the weighted average for the 30-39 age bin is $3.02$ yen, which deviates significantly from $2.58$ yen.
+The $2.58$ yen is between the average values for the 20–24 age bin and the 25–29 age bin.
+Thus, it is likely to represent the daily wage of workers at approximately $25$.
+This is consistent with the fact that workers are concentrated in the twenties.
+However, the weighted average for the 30–39 age bin is $3.02$ yen, which deviates significantly from $2.58$ yen.
 The distribution of wages at 1919 prices shows a similar result (Figure B.7b).
-Using weighted averages for earnings comparisons leads to the erroneous conclusion that the sample households are upwardly biased.
-To address this issue, I use the following steps to estimate the distribution of wages of the factory workers in Tokyo City in 1918-1919, the observation period of the THBS sample.
-Since the same estimation method is applied to both years, I use 1918 as an example below.
-First, the distribution of daily wages by age group was projected from statistics obtained from the LST.
+This means that the unweighted average leads to the erroneous conclusion that the sample households are upwardly biased.
+To address this issue, I use the following steps to estimate the distribution of wages of the factory workers in Tokyo City in 1918–1919, the observation period of the THBS sample.
+I use 1918 as an example below.
+First, the distribution of daily wages by age group was predicted from statistics obtained from the LST.
 Following Royston and Altman (1994), I consider a fractional polynomial of degree $m$ as follows:
@@
 $\phi_{2}(a;\varphi_{0},\varphi_{1},\varphi_{2},p_{1},p_{2})=\varphi_{0}+\varphi_{1}a^{(p_{1})}+\varphi_{2}a^{(p_{2})}.$ (7)
-To model the function, I consider a set of powers $\mathscr{P}=\{-2,-1,-0.5,0,0.5,1,2,3\}$ in the spirit of Royston and Altman (1994, p. 434).
+To model the function, I consider a set of powers $\mathscr{P}=\{-2,-1,-0.5,0,0.5,1,2,3\}$ following Royston and Altman (1994, p. 434).
 The maximum likelihood estimation is used to search for the most suitable candidate of the coefficients for a given $m$.
@@
 By adding the $8$ simple candidates under the nested models (i.e., the models with degree-1 fractional polynomial) to be conservative, I run $44$ regressions in total to find out the model with the highest maximized log-likelihood.
-The best power vector suggested is $(p_{1},p_{2})=(0.5,1)$.102102102The deviance, defined as the twice negative (maximized) log-likelihood, is calculated to be $-31.5$.
+The best power vector suggested is $(p_{1},p_{2})=(0.5,1)$.
+[^98] The deviance, defined as the twice negative (maximized) log-likelihood, is calculated to be $-31.5$.
 Royston and Altman (1994) proposed the deviance difference, $D(m,\mathbf{p})-D(m,\mathbf{\tilde{p}})$, where $\mathbf{\tilde{p}}$ is the best power vector to compare the different models.
@@
 $\hat{y}^{\textit{Daily Wage}}=-10.01+4.15a^{\frac{1}{2}}-0.33a^{1},\quad R^{2}=0.99.$ (8)
-The dashed curve in Figure B.7 is the predicted distribution in equation 8. The fitting suggests that the predicted distribution works properly, especially below the 50-54 bin.
+The dashed curve in Figure B.7 is the predicted distribution in equation 8. The fitting suggests that the predicted distribution works properly, especially below the 50–54 bin.
 Next, I shift the predicted distribution based on the difference in the average daily wages in the ASCT and LST.
-The weighted average of the daily wage is $2.56$ yen for the LST, whereas that for the ASTC is $1.05$ yen, meaning that there is roughly $1.5$ yen difference.103103103For the LST, I use the number of factory workers in each age bin as weight.
-The ASCT reports the average daily wage for four different categories of factories in each manufacturing sector: the factory scale (10+ or below ten workers) and engines (use or not).
-I calculated the weighted average of the daily wage using the number of factory workers in each age bin (12–14 and 15+) and in each category as weights.
-Thus, I shift the functional polynomial function by this amount of difference to yield the proxy distribution for 1918.104104104This operation, therefore, assumes that the wage-by-age distributions in 1918 and 1924 are reasonably similar.
-In the LST (i.e., the statistic on which the projected distribution is based), the weighted average for those aged 14 and older was 3.25 yen.
-In comparison, the overall weighted average was 3.23 yen, a difference of only 0.02 yen.
+The weighted average of the daily wage is $2.56$ yen for the LST, whereas that for the ASCT is $1.05$ yen, meaning that there is approximately $1.5$ yen difference.
+[^99] For the LST, I use the number of factory workers in each age bin as weight.
+The ASCT reports the average daily wage for four different categories of factories in each manufacturing sector: the factory scale (ten+ or below ten workers) and engines (use or not).
+I calculate the weighted average of the daily wage using the number of factory workers in each age bin (12–14 and 15+) and in each category as weights.
+Thus, I shift the functional polynomial function by this amount of difference to yield the proxy distribution for 1918.100100100This operation, therefore, assumes that the wage-by-age distributions in 1918 and 1924 are reasonably similar.
+In the LST (i.e., the statistic on which the predicted distribution is based), the weighted average for those aged $14$ and older was $3.25$ yen.
+In comparison, the overall weighted average is $3.23$ yen, a difference of only $0.02$ yen.
 A similar trend can be observed in the statistics obtained from the ASCT: the weighted average for those aged $15$ and older is $1.14$, while the overall average is $1.13$, a difference of only $0.01$ yen.
-At least, in the light of the first moment, this confirms that the wage distribution in the ASCT is similar to that in the LST.
+This suggests that the wage distribution in the ASCT is similar to that in the LST.
 Figure B.8a illustrates this operation: dotted curve is a fractional polynomial shifted from the original fractional polynomial shown as a short dashed curve.
@@
 In other words, the shifted curve shall get closer to the actual (unobservable) wage distribution as age increases.
-The long dashed curve in Figure B.8a examines this relationship by fitting another fractional polynomial model using the 12-14-year old bin figure of the ASCT and the predicted wages for 25–65 from the shifted fractional polynomial.105105105 The fitted model is as follows: $\hat{y}^{\textit{Daily Wage}}=-1.19+0.14a^{1}-0.002a^{2},\quad R^{2}=0.98.$ The deviance is reported to be $-149$.
+The long dashed curve in Figure B.8a examines this relationship by fitting another fractional polynomial model using the 12–14-year old bin figure of the ASCT and the predicted wages for 25–65 from the shifted fractional polynomial.101101101The fitted model is as follows: $\hat{y}^{\textit{Daily Wage}}=-1.19+0.14a^{1}-0.002a^{2},\quad R^{2}=0.98.$ The deviance is reported to be $-149$.
 I used the 25–65 age range for the prediction based on the age bins used in the LST.
@@
 However, the wage gaps for the younger ages show more significant differences.
-As long as our interest is in the late 20s to the early 40s, therefore, it is plausible to employ the predicted wages for the wage comparison in Section 3.
+If our interest is in the late 20s to the early 40s, it is plausible to employ the predicted wages for the wage comparison in Section 3.
 Figure B.8b illustrates the result of the same analysis from the wage data of 1919 price, reflecting the increasing trend in price between 1918 and 1919.
@@
 I calculated the weighted average of the daily wage using the number of factory workers in each age bin and the category as weights.
-The long-dashed curve shows the FP refitted using the average daily wage for 12-14-year-old bin ($0.42$ yen) measured in the ASCT and the daily wages for 25-65-year-old suggested in the shifted FP.
+The long-dashed curve shows the FP refitted using the average daily wage for 12–14-year-old bin ($0.42$ yen) measured in the ASCT and the daily wages for 25–65-year-old suggested in the shifted FP.
 The average daily wage for a 12–14-year-old bin is only available for the ASCT of 1918.
@@
 The blue dashed line is the average daily wage for 30–40-year-old in the refitted FP.
-Sources: Created by the author using the Tokyo City Office (1921, pp. 726–757); Tokyo City Statistics Division (1926a, pp. 16; 244-245).
+Sources: Created by the author using the Tokyo City Office (1921, pp. 726–757); Tokyo City Statistics Division (1926a, pp. 16; 244–245).
 B.6.2 Ancillary Wage
-Figure B.9: Log-Relationship between the Number of Workers and Share of Ancillary Wage: Original Data from the Factory Survey by Kitazawa (1924)
+Figure B.9: Log-Relationship between the Number of Workers and Share of Ancillary Wage: Original Data from the Factory Survey
+by Kitazawa (1924)
 Notes: This scatter plot shows the relationship between the log-transformed number of factory workers and the log-transformed share of ancillary wage.
@@
 In the case of fitting based on the raw data, the fitted fractional polynomial model is $\hat{y}^{\textit{Daily Wage}}=53.6-399.3a^{-1/2}+(2.84\times 10^{-10})a^{3}$ with $R^{2}=0.87$.
-I prefer to use linear regression herein because it can provide more precise fitted values for factories with a small number of workers.
+I prefer to use linear regression herein because it can provide more precise fitted values for small-and medium-scale factories.
 Source: Kitazawa (1924, pp. 20; 58–62).
 The daily wage estimated in Section B.6.1 is a fixed salary and thus, does not include additional salaries such as bonuses and allowances (hereafter, wages other than regular salaries are collectively referred to as ancillary wages).
-However, there are no systematic statistics on these ancillary wages throughout the prewar period (Committee for the History of the Labor Movement 1964, pp.
+However, there are no systematic statistics on the ancillary wages throughout the prewar period (Committee for the History of the Labor Movement 1964, pp.
 14–15).
@@
 Fixed wages are salaries added from a fixed hourly wage rate, while ancillary wages include bonuses and overtime pay, night shift pay, and labor premiums (profit sharing).
-In the case of a combination of hourly and piece-rate pay systems, the accessory wage includes a share of the piece-rate profit.
+In the case of a combination of hourly and piece-rate pay systems, the ancillary wage includes a share of the piece-rate profit.
 Panel A in Table B.3 summarizes the survey results.
@@
 One of the most frequently used statistics on factories in Tokyo is the Handbook of Factories (kōjyō tsūran), which covers all the factories with ten or more workers as of January 1, 1920.
-This handbook is the most comprehensive source for ascertaining the number of factories and their sizes around 1919.
+This handbook is the most comprehensive source for ascertaining the number of factories and their sizes circa 1919.
 This document does not cover the number of tiny working places with less than nine workers.
 Importantly, however, the purpose of Panel C in Table 2 is to compare the average monthly wage of the skilled factory workers in Tsukishima and the THBS heads’ average monthly earnings.
-Since the THBS heads did not work in the tiny workplaces, the censoring at the left tail of the factory size distribution is unlikely to influence this comparison.
+Since the THBS heads did not work in the small workplaces, the censoring at the left tail of the factory size distribution is unlikely to influence this purpose.
 Panel B of Table B.3 summarizes the distributions of the number of factories, the total number of workers, and the estimated average share of ancillary wage in Tokyo City (Column 1) and Tsukishima (Column 2).
@@
 As shown, Tsukishima factories tended to have larger factories than those of the entire city, which is reflected in the slightly larger estimate in the average ancillary wage for the Tsukishima factories.
-In Column (3) of Table B.3, the number of THBS heads whose occupation, factory name, and factory size are available are listed.
+In Column (3) of Panel C in Table B.3, the number of THBS heads whose occupation, factory name, and factory size are available are listed.
 While the percentage shares of the heads in each bin show a similar distribution to that of the Tsukishima factories, the share of middle (large) sized factories is larger (smaller) than the population.
 If interpreted literally, this tendency implies that the (unobservable) true average ancillary wage for the THBS heads may be slightly lower than the population mean.
-Note again that, however, the purpose of Panel C in Table 2 is to compare the average monthly wage of the skilled factory workers employed in the Tsukishima factories to the THBS heads’ average monthly earnings.
-Therefore, the purpose herein is to estimate the average ancillary wage for the target population, say the Tsukishima factories but not to calculate the average ancillary wage for the THBS sample itself.
+However, note that the purpose of Panel C in Table 2 is to compare the average monthly wage of the skilled factory workers employed in the Tsukishima factories to the THBS heads’ average monthly earnings.
+Therefore, the purpose herein is to estimate the average ancillary wage for the target population, for example, the Tsukishima factories, but not to calculate the average ancillary wage for the THBS sample itself.
 Table B.3: Estimating the Average Share of Monthly Ancillary Wage
@@
 D 623 77 68 32
-E 2837 NA 47 53
+E 2,837 NA 47 53
 F 928 41 60 40
@@
 There are 28 heads whose factory names and size are available.
-For a blacksmith included in these heads, I assume the number of workers to be 10, based on the report (Department of Health, Ministry of the Interior 1923a, p.
+For a blacksmith included in these heads, I assume the number of workers to be ten, based on the report (Department of Health, Ministry of the Interior 1923a, p.
 412).
@@
 Panel B: The number of factory workers is taken from the Ministry of Agriculture and Commerce (1921, pp.
-613-629; 674).
+613–629; 674).
 The average shares of ancillary wage are calculated using Kitazawa (1924, pp.
@@
 Panel C: Data on the number of factory workers and factories are from the Ministry of Agriculture and Commerce (1921, pp.
-613-629; 674).
+613–629; 674).
 Data on the number of heads are from the THBS sample.
@@
 1.
-Average daily wage ($w^{Daily}_{j}$): The average daily wage of male factory workers in the skilled age range (aged 30-40) in the machinery sector is calculated using two official reports of the manufacturing censuses documented in the ASCT and LST.
+Average daily wage ($w^{Daily}_{j}$): The average daily wage of male factory workers in the skilled age range (aged 30–40) in the machinery sector is calculated using two official reports of the manufacturing censuses documented in the ASCT and LST.
 Details in the calculation steps are summarized in Online Appendix B.6.1.
 2.
-Average annual working days ($l_{j}$): The ASCT documents the number of factories and average annual working days for each machinery sub-sector in 2 factory-size bins.
-I used the number of factories in each machinery sub-sector and factory size bin as the weights to calculate the weighted average of the annual working days in the entire machinery sector.
+Average annual working days ($l_{j}$): The ASCT documents the number of factories and average annual working days for each machinery sub-sector in two factory-size bins.
+I use the number of factories in each machinery sub-sector and factory size bin as the weights to calculate the weighted average of the annual working days in the entire machinery sector.
 3.
@@
 732-733); Tokyo City Statistics Division (1926a, pp.
-16; 244-245).
+16; 244–245).
 Data used to calculate the average annual working days are from the Tokyo City Office (1921, pp.
-726-741) and Tokyo City Office (1922, pp.
-716-725).
+726–741) and Tokyo City Office (1922, pp.
+716–725).
 Data used to calculate the average share of monthly ancillary wage are from Kitazawa (1924, pp.
-20; 58-62) and the Ministry of Agriculture and Commerce (1921, pp.
-613-629; 674).
-Following the institution described in Online Appendix B.6.2, the monthly wage of the male workers in the skilled age range ($w^{\textit{Total}}$) can be defined as follows:
+20; 58–62) and the Ministry of Agriculture and Commerce (1921, pp.
+613–629; 674).
+Following the institution described in Online Appendix B.6.2, the monthly wage of the male workers in the skilled age range ($w^{\textit{Total}}$) is defined as follows:
 $w^{\textit{Total}}=w^{\textit{Regular}}+w^{\textit{Ancillary}},$ (9)
 where $w^{\textit{Regular}}>0$ is the monthly regular wage and $w^{\textit{Ancillary}}\geq 0$ indicates the monthly ancillary wage.
-Online Appendix B.6.2 shows that $w^{\textit{Ancillary}}$ can be calculated as the product of the monthly regular wage and weight for the ancillary wage, $\vartheta\in(0,1)$.106106106Strictly, this weight is not required to be in this range.
-However, all the cases reported in Kitazawa (1924) take the weights within this range.
+Online Appendix B.6.2 shows that $w^{\textit{Ancillary}}$ can be calculated as the product of the monthly regular wage and weight for the ancillary wage, $\vartheta\in(0,1)$.102102102Strictly, this weight may not be required to be in this range.
+I use this range because all the cases reported in Kitazawa (1924) take the weights between zero and one.
 The monthly wage equation is then specified as follows:
@@
 To calculate equation 10, I have to conduct a few final tunings in my wage data.
-First, I fixed the deviations in the timing of the manufacturing censuses and the THBS observations.
+First, I fix the deviations in the timing of the manufacturing censuses and the THBS observations.
 Let $w^{\textit{Regular}}_{j}$ be the monthly regular wage calculated using wage statistics measured in year $j\in\{1918,1919\}$ as follows:
@@
 However, $w^{\textit{Regular}}_{j}$ is calculated using the wage statistics measured in the manufacturing census conducted in December of 1918 and 1919.
-This means that if I use $w^{\textit{Regular}}_{1919}$ alone for equation 10, it shall provide upward-biased wage due to the deviation between the timing of the THBS observations and of the manufacturing census.
-To cope with this issue, I use the weighted average between $w^{\textit{Regular}}_{1918}$ and $w^{\textit{Regular}}_{1919}$ to yield monthly wage series as follows:
+This means that if I used $w^{\textit{Regular}}_{1919}$ alone for equation 10, it shall provide upward-biased wage due to the deviation between the timing of the THBS observations and of the manufacturing census.
+To manage this issue, I use the weighted average between $w^{\textit{Regular}}_{1918}$ and $w^{\textit{Regular}}_{1919}$ to yield monthly wage series as follows:
 $w^{\textit{Regular}}_{m}=\{(12-m)w^{\textit{Regular}}_{1918}+mw^{\textit{Regular}}_{1919}\}/12,$ (12)
@@
 This way of weighting assumes that the average daily wage of machinery workers had increased during this period.
-This assumption is plausible given the historical fact that the average wage in the manufacturing industry indeed showed a secular increasing trend around 1920 in Tokyo City (Ohkawa et al.
+This assumption is plausible given the historical fact that the average wage in the manufacturing industry indeed showed a secular increasing trend circa 1920 in Tokyo City (Ohkawa et al.
 1967, p.
 255).
-Secondly, I use the number of observations in each year-month cell as the frequency weight.
+Second, I use the number of observations in each year-month cell as the frequency weight.
 Let $n_{m}$ be the number of households measured in year-month cell $m$.
@@
 The estimate is $56$ yen, which is similar to the average and median of monthly earnings in the THBS heads ($59$ and $56$ yen, respectively; see Panel C of Table 2).
-Therefore, I can conclude that the deviation from the population mean ($56$ yen) is not too large to infer the mean household behaviors of the skilled workers in the machinery factories in Tsukishima.
-Given that the average daily wage of the male factory workers in the skilled age range is $2.5$ yen (Figure B.8), $3$ yen difference from the average monthly income is roughly one day’s earnings.
+Therefore, I can conclude that the deviation from the population mean ($56$ yen) is not too large to infer the mean household strategies of the skilled workers in the machinery factories in Tsukishima.
+Given that the average daily wage of the male factory workers in the skilled age range is $2.5$ yen (Figure B.8), $3$ yen difference from the average monthly income is approximately one day’s earnings.
 This is consistent with the findings in my descriptive analyses that show the similarities in the distribution of family structure, average household size, and share of ancillary wage between the THBS households and Tsukishima households (Section 3; Online Appendix B.6.2).
@@
 The estimate is $54.5$ yen, which is slightly smaller than that in Column 1, reflecting the difference in the weights for the ancillary wages.
-If I see this figure as an unbiased estimate for the entire city, the difference in the average factory size between the entire Tokyo City and Tsukishima may be reflected in this deviation (Panel B in Table B.3).
+If I view this figure as an unbiased estimate for the entire Tokyo City, the difference in the average factory size between the entire Tokyo City and Tsukishima may be reflected in this deviation (Panel B in Table B.3).
 B.7 Measurement Errors in the Unadjusted Panel Dataset
@@
 The shifted series defines the days from the 2nd to the 16th of each month as the first half, whereas the days from the 17th to the 1st of the following month are the second half.
-Adjusted series is the baseline definition in the main text, which uses the days from 12th to 26th in each month as the “first-half” and those from 27th to 11th in the next month as the “second-half”.
+Adjusted series is the baseline definition in the main text, which uses the days from 12th to 26th in each month as the “first-half” and those from 27th to 11th in the next month as the “second-half.”
 As explained in Section 3, the raw series contains considerable measurement errors due to the timing of payday.
-The shifted series can deal with the measurement errors from the timing of payday but can not adjust the difference in the payday timing and expenses.
+The shifted series can manage the measurement errors from the timing of payday but cannot adjust the difference in the payday timing and expenses.
 In other words, the shifted series ignores the households’ decision on the consumption schedule in each semi-month.
-The adjusted series can deal with measurement errors and take the timing of the household expenses into account.
+The adjusted series reflecting the exact timing of the household expenses can manage measurement errors.
 Figures B.10bB.10dB.10f illustrate the average semi-monthly income minus expenditure calculated from unadjusted, shifted, and adjusted panel datasets, respectively.
 Source: Created by the author using the THBS sample.
-Figure B.10 explains how different ways of aggregation influence the measured time-series in income and expenditure.
-Figures B.10aB.10cB.10e show the average semi-monthly income and expenditure calculated from raw, shifted, and adjusted panel datasets, respectively.
+Figure B.10 explains how different methods of aggregation influence the time-series in income and expenditure.
+Figures B.10aB.10cB.10e show the average semi-monthly income and expenditure calculated from unadjasted, shifted, and adjusted panel datasets, respectively.
 The unadjusted (calender-month) series defines the days from the 1st to the 15th of each month as the first half and the days from the 16th to the end of each month as the second half.
 This series shall contain substantial measurement errors due to the mismatches in the income and expenditure data explained in Section 3.3.
-The time-series plots of average income in Figure B.10a have substantially large variances and fluctuate over time.
+The time-series plots of average income in Figure B.10a have a substantially large variance and fluctuate over time.
 This is because some semi-monthly income is assigned to the wrong semi-month cells.
-In these wrong cells, the measured incomes shall take unnaturally greater values than usual.
-Moreover, the time-series plots of average expenditure do not chase those of the income.
+In these wrong cells, the measured incomes take unnaturally greater values than usual.
+Moreover, the time-series plots of average expenditure do not chase those of average income.
 For example, the average expenditure in the second half of June greatly differs from the average income in the same cell.
-Consequently, the average net income (i.e., income minus expenditure) in each semi-month is relatively flat and takes a few odd cells with large positive values (Figure B.10b).
+Consequently, the average net income (i.e., income minus expenditure) in each semi-month takes unnaturally large positive values in a few cells (Figure B.10b).
 The shifted series defines the days from the 2nd to the 16th of each month as the first half and the days from the 17th to the 1st of the following month as the second half.
-Although the shifted series can partly deal with the measurement errors in income, it does not adjust the difference in the payday timing and expenses.
-Figure B.10c indeed gets a much smaller variance in the average incomes over time because all the paydays are now fixed to assign the suitable semi-month cells.
+Although the shifted series can partly manage the measurement errors in income, it does not adjust the difference in the timing of paydays and expenses.
+Figure B.10c obtains a much smaller variance in the average incomes over time because all the paydays are now fixed to assign suitable semi-month cells.
 However, the expenditure series still does not chase the income series well because the payday is placed on the latter edge in each semi-month cell.
@@
 However, it still has a few cells with large negative and large values due to the deviations of expenditure from income.
-The adjusted series uses the days from the 12th to 26th of each month as the “first half” and those from the 27th to 11th of the following month as the “second half”.
-By shifting the entire semi-month cells, this adjusted series can deal with the measurement errors from the payday shifts and fix the miss-matching in the timing of income and expenditure.
-In other words, each payday is assigned to a suitable semi-month cell, and each semi-month cell reflects the household’s decision on consumption given the latest income.
+The adjusted series uses the days from the 12th to 26th of each month as the “first half” and those from the 27th to 11th of the following month as the “second half.” By shifting the entire semi-month cells, this adjusted series can manage the measurement errors from the payday shifts and can fix the miss-matching in the timing of income and expenditure.
+In other words, each payday is assigned to a suitable semi-month cell, and each semi-month cell reflects the household’s decision on consumption in response to the latest income.
 Figure B.10e shows similar time-series plots of average income in Figure B.10c, in which has a corrected semi-monthly average income series only.
 Notably, the expenditure series are now chasing the income series in Figure B.10e because each semi-month cell’s expenditure is based on the income in the same cell.
-Accordingly, Figure B.10f now shows a smoother series in the net income.
-It shows a clear increasing trend in the net income and has a smaller variance over time.
-These illustrations support the evidence that the adjusted semi-monthly datasets can offer cleaner within variations in income and expenditure because they can fix the measurement errors and can reflect the households’ consumption behaviors based on their latest incomes measured in the same semi-month cell.
+Accordingly, Figure B.10f now shows a smoother series with a smaller variance in the net income.
+These illustrations support the evidence that the adjusted semi-monthly datasets can offer cleaner within variations in income and expenditure because they can fix the measurement errors and reflect the households’ consumption strategies based on their latest incomes measured in the same semi-month cell.
 B.8 Distribution of the Idiosyncratic Shocks
@@
 Source: Created by the author using the THBS sample.
-Figure B.11 illustrates the distribution of the idiosyncratic shocks predicted using the semi-monthly panel dataset.107107107I used the semi-monthly panel dataset to overview the entire distribution of the idiosyncratic shocks rather than the daily panel datasets.
-This is because the denser panels include a set of consecutive household-date cells with no income.
-These cells can not be used to predict the idiosyncratic shocks as the residuals economically meaningfully, and thus it is practically impossible to illustrate the idiosyncratic shocks based on the daily dataset.
-The idiosyncratic shocks are calculated as the residuals from the regression in the first difference in semi-monthly income on the first difference in aggregate semi-monthly income.
+Figure B.11 illustrates the distribution of the idiosyncratic shocks predicted using the semi-monthly panel dataset.103103103I use the semi-monthly panel dataset to overview the entire distribution of the idiosyncratic shocks rather than the daily panel datasets.
+This is because the daily panels include a set of consecutive household-date cells with no income.
+These cells cannot be used to predict the idiosyncratic shocks as the residuals are economically meaningfully, and thus it is practically impossible to illustrate the idiosyncratic shocks based on the daily dataset.
+The idiosyncratic shocks are calculated as residuals from the regression in the first difference in semi-monthly income on the first difference in aggregate semi-monthly income.
 Let $n_{t}$ be the number of households in semi-month $t$ in the year 1919.
@@
 $P_{m}$-statistic $p$-value 0.0000 0.0000
-The results of the Fisher-type panel unit root tests based on augmented Dickey-Fuller (ADF) tests are reported in this table.
+Notes: The results of the Fisher-type panel unit root tests based on augmented Dickey-Fuller (ADF) tests are reported in this table.
 The null hypothesis is that all the panels contain unit roots, whereas the alternative hypothesis is that at least one panel is stationary.
@@
 The demeaned data are used to address the effect of cross-sectional dependence.
-The number of lagged differences in the ADF regression equation reported is set as one because including two or more two lags reduces a large part of the panel units available for computing the test statistics due to the unbalancedness of the panels.
-I confirmed that including two lags in the semi-monthly panels does not change the results.
+The number of lagged differences in the ADF regression equation reported is set as one because including two or more two lags reduces a large part of the panel units available for computing the test statistics due to the lack of balance of the panels.
+I confirm that including two lags in the semi-monthly panels does not change the results.
 Including two lags in the adjusted monthly panels and three lags in the semi-monthly panels is impossible.
@@
 In my empirical setting, the heterogeneity in the risk preference is assumed to be removed by using the household fixed-effect.
-When the risk preferences are difference over households, say $\sigma=\sigma_{i}$, equation 26 becomes:
+When the risk preferences are differenced over households like $\sigma=\sigma_{i}$, equation 26 becomes:
 $\log c_{i,t}=\log c_{t}^{a}+\frac{1}{\sigma_{i}}(\log\omega_{i}-\omega^{a})+(\theta_{i,t}-\theta_{t}^{a}).$ (16)
 The first differencing removes the second term with the coefficient of relative risk aversion in the right-hand side of this equation, even though it varies over the households.
-Schulhofer-Wohl (2011) shows that the heterogeneous risk preference may lead to bias in the estimation of the income elasticity under the standard risk-sharing regressions and proposes a parametric test for that heterogeneity.
+Schulhofer-Wohl (2011) shows a specific case (utility function) that the heterogeneous risk preference may lead to bias in the estimation of the income elasticity under the standard risk-sharing regressions and proposes a parametric test for that heterogeneity.
 This test is valid under a particular utility function.
-Specifically, the bias could be raised under the case that the coefficient of relative risk aversion interacts with the aggregate shock in the first-order condition for consumption because the interaction means that the aggregate shock can not be removed using the time-fixed effect in the risk-sharing regressions.108108108Schulhofer-Wohl (2011, pp.
+Specifically, the bias could be raised in the case that the coefficient of relative risk aversion interacts with the aggregate shock in the first-order condition for consumption because the interaction means that the aggregate shock cannot be removed using the time-fixed effect in the risk-sharing regressions.104104104Specifically, Schulhofer-Wohl (2011, pp.
 928–929) considers the utility maximization problem under the following weighted sum of discounted expected utilities: $\sum_{i}\dot{\alpha}_{i}E\left[\sum_{t}\dot{\beta}^{t}\frac{[c_{i,t}(s_{t})]^{1-\dot{\gamma}_{i}}}{1-\dot{\gamma}_{i}}\right],$ (17) where $\dot{\alpha}_{i}$ is Parate weight and $\dot{\gamma}_{i}$ is the coefficient of relative risk aversion for household $i$.
@@
 Mozzoco and Saini (2012) consider a more general case for testing the homogeneity of risk preferences.
-Their test requires a panel dataset with a very long time dimension that can not be applicable to my THBS dataset.109109109The data used in their study covers the household-level monthly panel dataset surveyed in rural India between 1975 and 1985.
+Their test requires a panel dataset with a very long time dimension that is not applicable to my THBS dataset.105105105The data used in their study covers the household-level monthly panel dataset surveyed in rural India between 1975 and 1985.
 The households with fewer than 80 data points are not included in the homogeneity tests (Mozzoco and Saini 2012, p.
@@
 An alternative way to test potential heterogeneity in the risk preference is to use direct measures on the risk preferences of workers.
-Schulhofer-Wohl (2011) uses the results of experimental risk-tolerance questions on the workers’ risk preferences in the Health and Retirement Study to test whether the correlations between earnings and aggregate shocks differ across high- and low-risk-tolerance workers.110110110This question is asking if the respondents would be willing to take a risk to get a new job with a higher income, even if there’s a 50-50 chance of a reduction in income by a third (Schulhofer-Wohl 2011, P.
+Schulhofer-Wohl (2011) uses the results of experimental risk-tolerance questions on the workers’ risk preferences in the Health and Retirement Study to test whether the correlations between earnings and aggregate shocks differ across high- and low-risk-tolerance workers.106106106This question is asking if the respondents would be willing to take a risk to get a new job with a higher income, even if there was a 50-50 chance of a reduction in income by a third (Schulhofer-Wohl 2011, P.
 931).
@@
 Since the THBS has no experimental questions, it is impossible to apply the same test.
-Given this limitation that the direct test using experimental risk-tolerance question is unavailable, I compare the correlations between the heads’ income and aggregate consumption over sub-groups to get insights into the potential difference in the heads’ attitudes toward risks.
+Given this limitation—that the direct test using experimental risk-tolerance question is unavailable—I compare the correlations between the heads’ income and aggregate consumption over sub-groups to obtain insights into the potential difference in the heads’ attitudes toward risks.
 For household $i$ and semi-month $t$, the specification is characterized as follows:
@@
 The aggregate consumption is calculated as the cross-sectional average of the total expenditure among the THBS sample, given that the population mean of consumption expenditures among all the skilled factory workers’ households in Tsukishima is unavailable.
-$d_{i}$ is an indicator variable that takes one if household $i$ belongs to a specific group that might have had different preferences on their jobs than the other heads in the sample.
-Firstly, I consider the case that the heads who worked in the large-scale factory had different risk preferences.
-Although there are no specific anecdotes behind this, the workers in the large-scale factory might have had a lower risk tolerance than those in the small- and medium-scale enterprises.
-13 heads who worked in the Ishikawajima shipyard and governmental factory are classified into this group.
-Secondly, I assume that the heads who worked in the smithing and the non-machinery factories had different risk preferences.
-As described in Online Appendix A.5, skilled workers in the smithing sector could be small business owners.111111111Workers in the canning factories might have had a different preference on the risk preferences (Online Appendix A.6).
+$d_{i}$ is an indicator variable that takes one if a household $i$ belongs to a specific group that might have had different preferences on their jobs than the other heads in the sample.
+Firstly, I test the hypothesis that the heads who work in the large-scale factories have different risk preferences.
+Although there are no specific anecdotes behind this, the workers in the large-scale factories might have had a lower risk tolerance than those in the small- and medium-scale enterprises.
+Thirteen heads who worked in the Ishikawajima shipyard and governmental factory are classified into this group.
+Secondly, I test the hypothesis that the heads who work in the smithing and the non-machinery factories have different risk preferences.
+As described in Online Appendix A.5, skilled workers in the smithing sector could be small business owners.107107107Workers in the canning factories might have had a different preference on the risk preferences (Online Appendix A.6).
 However, my THBS sample does not include the workers in canning factories.
@@
 In Column (1), the estimated coefficient on the interaction term is close to zero and statistically insignificant.
-Thus, the sensitivities to aggregate consumption are similar across different scales of the factories.
+Thus, the sensitivities to aggregate consumption are similar across different scales of factories.
 Column (2) further shows that the estimated coefficient on the interaction term is not statistically significant under an alternative proxy of risk tolerance.
-It suggests that the heads worked in the smithing and non-machinery factories show similar responses to the aggregate shock.
+It suggests that the heads who work in the smithing and non-machinery factories show similar responses to the aggregate shock.
 Since the aggregate consumption is the cross-sectional sample average, the aggregate variable may contain noise in the period with the small number of cross-sectional observations.
-To deal with this potential issue, I exclude October and November 1919 in Columns (3) and (4).
+To manage this potential issue, I exclude October and November 1919 in Columns (3) and (4).
 As shown, the estimated coefficients on the interaction terms are largely unchanged in both columns.
-This finding is consistent with the fact that these heads have similar characteristics in nature because the THBS focused on skilled factory workers in a representative manufacturing area, Tsukishima.
-Therefore, it is plausible to see that the THBS households had more or less similar attitudes toward the risks.
+This finding is consistent with the fact that these heads have similar characteristics because the THBS focused on skilled factory workers in a representative manufacturing area, Tsukishima.
+Therefore, it is plausible that the THBS households had similar attitudes toward the risks.
 Table B.6: Results for Testing Heterogeneous Risk Preferences
@@
 Analytical Period
-(1) Full (2) Full Jan. to Sept. Jan. to Sept.
+(1) Full (2) Full (3) Jan. to Sept. (4) Jan. to Sept.
 Aggregate Consumption 0.277* 0.273 0.361** 0.362*
@@
 ‘Smithing/Non-machinery’ indicates an indicator variable that takes one if the head worked in the smithing, food, or paper-making factories.
-All the regressions include the family size variable and household fixed-effect.
+All the regressions include the family size variable and household fixed effect.
 Columns 1–2 show the results for the entire sample period.
-Columns 2–3 show the results for the period between January and September 1919 to remove year-month cells with the smaller (cross-sectional) number of households (Online Appendix Figure B.4).
+Columns 3–4 show the results for the period between January and September 1919 to remove year-month cells with the smaller (cross-sectional) number of households (Online Appendix Figure B.4).
 Appendix Appendix C Empirical Analysis Appendix
@@
 Following Mace (1991), I assume that the aggregate idiosyncratic shocks approach zero as $N$ gets large.
-Consider the situation that the household’s endowment can be decomposed into the permanent component ($e_{i,t}^{P}$), shock linked to macroeconomic trends ($\varrho_{i,t}^{a}(s_{t})$), and idiosyncratic shock ($\varsigma_{i,t}(s_{t})$) as:
+Consider the situation in which the household’s endowment can be decomposed into the permanent component ($e_{i,t}^{P}$), shock linked to macroeconomic trends ($\varrho_{i,t}^{a}(s_{t})$), and idiosyncratic shock ($\varsigma_{i,t}(s_{t})$) as:
 $e_{i,t}(s_{t})=e_{i,t}^{P}+\varrho_{i,t}^{a}(s_{t})+\varsigma_{i,t}(s_{t}).$ (22)
@@
 where $\sigma>0$ is the coefficient of relative risk aversion.
-Below, I consider a particular history of the state of the world so that the subscript $t$ can represent the dependence on the state (Hayashi et al.
+Below, I consider a particular history of the state of the world so that the subscript $t$ can represent dependence on the state (Hayashi et al.
 1996).
@@
 The aggregate consumption growth in equation(29) is assumed to reflect macroeconomic shocks in Tsukishima in the case of my empirical setting.
-As discussed in Section 3.4, the consumption of the THBS households indeed captures the overall macroeconomic trend.
-Despite this, interpreting $\alpha_{1}$ is not easy because the THBS households include the factory workers’ households in Tsukishima alone.
-To deal with this potential issue, I use a two-way fixed-effects model instead of the first-difference model in the spirit of Cochrane (1991).
-Another potential issue is that $\epsilon_{i,t}$ may include the household-specific time-varying preference shock, which might be correlated with the idiosyncratic income shock.
+As discussed in Section 3.4, the consumption of the THBS households captures the overall macroeconomic trend.
+Despite this, interpreting $\alpha_{1}$ is difficult because the THBS households include the factory workers’ households in Tsukishima alone.
+To manage this potential issue, I use a two-way fixed-effects model instead of the first-difference model, following Cochrane (1991).
+Another potential issue is that $\epsilon_{i,t}$ may include household-specific time-varying preference shock, which might be correlated with idiosyncratic income shock.
 Family size influences the household’s preference because the marginal utility of consumption increases with family size (Jappelli and Pistaferri 2017).
-As explained, the family size had been fairly stable during the sample period in the THBS households, meaning that the preference shifts had rarely occurred (Section 3.4).
+As explained, family size was stable during the sample period in the THBS households, meaning that the preference shifts rarely occurred (Section 3.4).
 To be conservative, however, I include the number of family members as a control variable for representing potential preference shock.
@@
 The estimate of $\gamma$ shall range from zero (i.e., idiosyncratic shocks are perfectly insured) to one (for the absence of insurance).
-In order to test the risk-coping mechanisms, I consider the following representation of individual consumption in the spirit of Fufchamps and Lund (2003).
+To test the risk-coping mechanisms, I consider the following representation of individual consumption following Fufchamps and Lund (2003).
 $c_{i,t}=y_{i}^{P}+y_{i,t}^{T}+r_{i,t},$ (31)
@@
 $r_{i,t}=\bm{z}^{\prime}_{i,t}\bm{1}$ is a linear combination of the net temporary incomes expressed using a $k\times 1$ net income vector ($\bm{z}$) and a vector of ones.
-Substituting equation 31 into equation 26 and taking the first-difference removes the time-constant components, including the permanent income.
+Substituting equation 31 into equation 26 and taking the first-difference removes the time-constant components, including permanent income.
 It follows that my empirical specification can be characterized as follows:
 $r_{i,t}=\kappa\tilde{y}_{i,t}+\eta x_{i,t}+\nu_{i}+\zeta_{t}+\epsilon_{i,t},$ (32)
-where $\tilde{y}_{i,t}$ is head’s earning, $x_{i,t}$ is the family size control, $\nu_{i}$ indicates the household-fixed effect, $\zeta_{t}$ indicates the time-fixed effect, and $\epsilon_{i,t}$ is a random error term.
-I herein use the head’s earnings ($\tilde{y}_{i,t}$) as the idiosyncratic shock variable instead of the disposable income.112112112In the analysis, I consider the labor-supply adjustment variable and sales of miscellaneous assets as the additional dependent variables in addition to the net income variables $r_{i,t}$ (Section 4.2).
-The former variable is the total earnings of the family members except for the head, meaning that the simultaneous issue rises if I use the disposable income in this equation.
+where $\tilde{y}_{i,t}$ is the head’s earning, $x_{i,t}$ is the family size control, $\nu_{i}$ indicates the household-fixed effect, $\zeta_{t}$ indicates the time-fixed effect, and $\epsilon_{i,t}$ is a random error term.
+I herein use the head’s earnings ($\tilde{y}_{i,t}$) as the idiosyncratic shock variable instead of the disposable income.108108108In the analysis, I consider the labor-supply adjustment variable and sales of miscellaneous assets as the additional dependent variables in addition to the net income variables $r_{i,t}$ (Section 4.2).
+The former variable is the total earnings of the family members except for the head, meaning that the simultaneous issue arises if I use the disposable income in this equation.
 The regression is separately run for each component in the net income vector ($\bm{z}$).
@@
 This means that a semi-month cell takes a minimal value when the payment of rent is not included in that cell.
-Figure C.1c indeed shows that the LD of rent generates some larger values in the semi-monthly panel dataset due to those water bill payments.
-These cells taking large LD values substantially increased the slope, as shown in Figure C.2c.
+Figure C.1c shows that the LD of rent generates some larger values in the semi-monthly panel dataset due to those water bill payments.
+These cells taking large LD values substantially increase the slope, as shown in Figure C.2c.
 Given this technical issue, the adjusted monthly panel dataset is suitable for estimating the income elasticity for the rent category.
@@
 (j) Transportation
-(k) Gift
+(k) Gifts
 (l) Miscellaneous
@@
 Figures C.1bC.1l show the distributions of the log-differences in semi-monthly expenditure for the 11 subcategories listed in panel A of Table 3.
-Sources: Created by the author using the THBS sample.
+Source: Created by the author using the THBS sample.
 (a) Total expenditure
@@
 (j) Transportation
-(k) Gift
+(k) Gifts
 (l) Miscellaneous
@@
 in Semi-monthly Income and Expenditure
-Notes: Figure C.2a illustrates the relationship between the log-differences in semi-monthly income and expenditure.
+Notes: Figure C.2a illustrates the relationship between the log-difference in semi-monthly disposable income and expenditure.
 Figures C.2bC.2l illustrate the relationships between the log-differences in semi-monthly disposable income and the 11 expenditure subcategories listed in Panel A in Table 3.
 The range of the y-axis is fixed across all the figures for comparability.
-Sources: Created by the author using the THBS sample.
+Source: Created by the author using the THBS sample.
 (a) Total expenditure
@@
 (j) Transportation
-(k) Gift
+(k) Gifts
 (l) Miscellaneous
@@
 Figures C.3bC.3l show the distributions of the log-differences in adjusted monthly expenditure for the 11 subcategories listed in panel A of Table 3.
-Sources: Created by the author using the THBS sample.
+Source: Created by the author using the THBS sample.
 (a) Total expenditure
@@
 (j) Transportation
-(k) Gift
+(k) Gifts
 (l) Miscellaneous
@@
 in Monthly Income and Expenditure
-Notes: Figure C.4a illustrates the relationship between the log-differences in semi-monthly income and expenditure.
+Notes: Figure C.4a illustrates the relationship between the log-difference in semi-monthly disposable income and expenditure.
 Figures C.4bC.4l illustrate the relationships between the log-differences in semi-monthly disposable income and the 11 expenditure subcategories listed in Panel A in Table 3.
@@
 C.3 Retail Sales in Tokyo City: A Retail Store Survey of 1935
-Table C.1 summarizes the contribution of regular customers to their total sales, social classes of the customers, and sales methods in the retailers selling white rice, fish, greengrocers, and fuels (firewood and charcoal) by scale of business.
-In the original document, the figures for “Total” in the firewood and charcoal category are missing in “Sales Method (%)”.
-For these three cells, I filled the averages over the three business-size categories in each sales method: 20.4% for “Cash”, 78.4% for “Credit purchase”, and 1.3% for “Monthly installment”.
+Table C.1 summarizes the contribution of regular customers to their total sales, social classes of the customers, and sales methods in the retailers selling white rice, fish, fruit and vegetable, and fuels (firewood and charcoal) by scale of business.
+In the original document, the figures for “Total” in the firewood and charcoal category are missing in “Sales Method (%).” For these three cells, I filled the averages over the three business-size categories in each sales method: 20.4% for “Cash,” 78.4% for “Credit purchase,” and 1.3% for “Monthly installment.”
 Table C.1: Customers, Sales Methods, and Competitiveness in the Retailers in Tokyo City
@@
 Notes:
-This table summarizes the statistics on the sales in the retail shops selling white rice, fish, fruit and vegetables, and firewood and charcoal in Tokyo surveyed in 1935.
+This table summarizes the statistics on the sales in the retailers selling white rice, fish, fruit and vegetables, and firewood and charcoal in Tokyo surveyed in 1935.
 1. Small, middle, and large-scale businesses include retailers with annual sales of less than 10,000, 10,000–30,000, and 30,000+ yen, respectively.
@@
 Table C.2 summarizes the use of credit purchase and the losses on the credit sales documented in the 1930 Survey Report.
-Following the report’s style, the item categories are divided into 6 categories: white rice, other grains, vegetables, fruit, fish, and fuel.
+Following the report’s style, the item categories are divided into six categories: white rice, other grains, vegetables, fruit, fish, and fuel.
 The total number of retailers in each category is $157$, $7$, $67$, $16$, $46$, and $64$, respectively.
@@
 Importantly, among the retailers using credit purchases, almost all of them suffered losses in sales.
-For example, 99% of white rice retailers experienced any losses.
+For example, $99$% of white rice retailers experienced losses.
 The minimum value of this share is still $97$% for the fuel retailers.
-This means that credit purchase was inherently a loss-making sales method for the retailers.
+This means that credit purchase was a loss-making sales method for the retailers.
 The statistics in Table C.2 also indicate that many cases have unknown loss amounts.
@@
 However, it is worth mentioning that a large part of losses was capped by $5$% of total sales in all these different retailers.
-In other words, this implies that while the retailers could not avoid losses, they might have set prices that anticipated predicted losses.
+This implies that while the retailers could not avoid losses, they might have set prices that anticipated losses.
 Table C.2: Sales Methods and Losses on Sales under Credit Purchases in Tokyo City
@@
 Share of retailers experienced losses (%) 99 100 100 100 100 97
-Notes: This table summarizes the statistics on the sales methods and losses on sales under credit purchases among the retailers measured in Tokyo City surveyed in 1930.
+Note: This table summarizes the statistics on the sales methods and losses on sales under credit purchases among the retailers measured in Tokyo City surveyed in 1930.
 Source: Tokyo City Office 1932, pp.
@@
   • (5)
-  • (6) Choi, In “Unit root tests for panel data”. Journal of International Money and Finance 20, no. 2 (2001): 249–272.
+  • (6) Choi, In “Unit root tests for panel data.” Journal of International Money and Finance 20, no. 2 (2001): 249–272.
   • (7)
-  • (8) James, John A., and Isao Suto. “Early Twentieth-century Japanese Worker Saving: Precautionary Behaviour before a Social Safety Net.” Cliometrica 5, no. 1 (2011): 1–25.
+  • (8) James, John A., and Isao Suto. “Early Twentieth-century Japanese Worker Saving: Precautionary Behavior before a Social Safety Net.” Cliometrica 5, no. 1 (2011): 1–25.
   • (9)
@@
   • (29)
-  • (30) Schulhofer-Wohl, Sam. “Heterogeneity and Tests of Risk Preference”. Journal of Political Economy 119, no.5, (2011): 925–958.
+  • (30) Schulhofer-Wohl, Sam. “Heterogeneity and Tests of Risk Preference.” Journal of Political Economy 119, no.5, (2011): 925–958.
   • (31)
@@
 Toshishominkinyū nikansuru chōsa, daiyonsatsu kousetsu shichiho (Survey of Financial Institutions among Common People, Volume.
-2, Pubic pawnshops) [in Japanese].
+2, Public pawnshops) [in Japanese].
 Tokyo Institute for Municipal Research, 1926.
@@
   • (3)
-  • (4) The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (7-2).
-  • (5) The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (7-3).
-  • (6) The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (7-4).
+  • (4) The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (7–2).
+  • (5) The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (7–3).
+  • (6) The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (7–4).
   • (7)
-  • (8) The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (8-12).
-  • (9) The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (8-13).
-  • (10) The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (8-14).
-  • (11) The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (8-15).
+  • (8) The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (8–12).
+  • (9) The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (8–13).
+  • (10) The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (8–14).
+  • (11) The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (8–15).
   • (12)

### v11

@@
 I found that while the households were so vulnerable that the shocks impacted their consumption levels, the estimated income elasticity for food consumption was relatively low in the short run.
-The event-study analysis using adverse health shock confirms the robustness of the results.
+The analysis using adverse health shock confirms the robustness of the results.
 The result of mechanism analysis suggests that credit purchases with local retailers smoothed short-run food consumption.
@@
 1 Introduction
-Economic theory illustrates the mechanism behind the smoothing of consumption in complete markets (Arrow and Debreu 1954).
-Additionally, a series of empirical studies have tested the smoothness of consumption among households in various economies under systematic empirical specification.
-Evidence shows that markets in contemporary developing countries are often imperfect.111 The early representative literature includes, for example, Townsend (1994), Townsend (1995), and Ravallion and Chaudhuri (1997).
+Economic studies that analyze the relationship between historical economic development and financial institutions have focused on the alleviation of constraints on corporate financing through official financial systems (Acemoglu et al.
+2005; Levine 2005).
+However, systematic research on the association between household financial institutions and consumption smoothing during industrialization remains limited.
+How did working-class households respond to idiosyncratic shocks and smooth their consumption before public insurance was established?
+Development economics research has confirmed that consumption smoothing is an essential mechanism for mitigating poverty and explaining human capital formation.
+[^1] Evidence shows that markets in contemporary developing countries are often imperfect.
+Thus, the households are vulnerable to risk, and idiosyncratic shocks impact the levels of their consumption.
+The early representative literature includes, for example, Townsend (1994), Townsend (1995), and Ravallion and Chaudhuri (1997).
 Morduch (1999), Townsend (1995), and Dercon (2004) provide reviews of the earlier studies.
 Attanasio and Pistaferri (2016) and Meyer and Sullivan (2023) provide recent discussions.
-The households are vulnerable to risk, and idiosyncratic shocks impact the levels of their consumption because shocks are not fully insured in the markets.
-Therefore, consumption smoothing is an essential mechanism for mitigating poverty, especially in developing economies.
-[^2] Smoothness of consumption is also an important factor in human capital formation.
-For example, Foster (1995) investigates the shock of the 1988 floods in Bangladesh and shows that children’s growth patterns are influenced by the inefficiency of the credit markets.
+Since shocks are not fully insured in the markets, smoothness of consumption is also an important factor in human capital formation.
+Foster (1995) investigates the shock of the 1988 floods in Bangladesh and shows that children’s growth patterns are influenced by the inefficiency of the credit markets.
 Rose (1999) investigates the relationship between consumption smoothing and excess losses of girls, revealing that the favorable rainfall events in early life increased the survival probabilities of girls in rural India.
 Gertler and Gruber (2002) also shows that consumption smoothing against health shocks is imperfect in Indonesia.
-Credit institutions in local society can mitigate idiosyncratic income shocks, especially when the availability of formal financial institutions is limited.
+Surprisingly, however, economic history studies have not yet provided an answer to this important question.
+We still know little about how historical households maintained their livelihoods during industrialization.
+This study assesses risk-coping behavior among the working-class households in early twentieth-century Tokyo.
+I digitized and analyzed a unique daily longitudinal budget survey on factory worker households conducted in a representative manufacturing area.
+The full insurance hypothesis is strongly rejected for overall consumption expenditure, indicating that factory worker households in the historical metropolis suffered certain consumption losses when they faced idiosyncratic adverse income shocks.
+However, expenses on indispensable consumption categories were partially smoothed in the short run.
+Credit purchases in local retailers played an important role in mitigating adverse idiosyncratic shocks, whereas the other risk-mitigation strategies were impractical.
+Credit purchases did not require an explicit contract and thus often led to a loss in sales for the retailers.
+However, to remain competitive, retailers allowed credit transactions so that their customers remained regular.
+This shows that informal credit institutions in urban local communities functioned as informal insurance networks in rural economies.
+This study’s primary contribution is to add analytical perspectives on consumption smoothing behavior to economic history studies.
+Studies of how working-class households compensated for income loss during the late 19th to the early 20th century have focused on Britain.
+These studies used cross-sectional variation in income to estimate the elasticities of the temporary income sources.
+Their detailed analyses have revealed the roles of financial institutions and labor supply adjustments in compensating for income loss in working-class households (Scott and Walker 2012).
+[^2] They found that interwar British working-class households had access to various institutions such as clothing clubs, hire-purchase, contractual savings, and insurance.
+The households with less earnings were likely to use credit-based institutions like clubs and hire-purchase, whereas those with higher incomes could use savings institutions.
+Horrell and Oxley (2000) highlight the important role of labor supply adjustments among family members to manage the income losses of the head in working-class households in late 19th-century Britain.
+Kiesling (1996) and Boyer (1997) focused on the risk-mitigation strategies among the cotton workers households in Lancashire during the recession caused by the Civil War.
+However, there is little evidence of how those risk-coping strategies smoothed their consumption.
+[^3] Other strands of literature provide supply-side (lender-side) evidence on the functions of the micro-financial institutions for the workers (O’Connell and Reid 2005), building image on the historical consumers in the long-run (Francks 2009), and describe the relationship between consumption and production (Gordon 2013).
+Hence, this study provides a rare window into urban household financial behavior during industrialization by utilizing unique household longitudinal budget data from prewar Japan.
+It also implements a systematic panel data analysis to offer comparable estimates of the income elasticity for future studies.
+The second contribution is revealing the roles of informal credit institutions in historical metropolis.
+Evidence from development economics studies indicates that the availability of informal credit institutions relates to the stability of consumption in rural economies, especially when the availability of formal financial institutions is limited.
 Udry (1994) showed that state-contingent loans among relatives and neighbors could partially insure against idiosyncratic shocks on rural households in northern Nigeria.
 A detailed field survey revealed that asymmetric information was no longer important in those private credit contracts (Udry 1990).
-Fafchamps and Lund (2003) also presented evidence that the households in the rural Philippines used informal networks with relatives and friends as a risk-mitigation strategy for idiosyncratic shocks.
+[^4] Fafchamps and Lund (2003) also presented evidence that the households in the rural Philippines used informal networks with relatives and friends as a risk-mitigation strategy for idiosyncratic shocks.
 Related studies consider risk-sharing among households with similar cultural and institutional backgrounds.
@@
 Munshi and Rosenzweig (2016) presented the importance of rural insurance networks among workers in the same caste in the context of spatial wage inequality in India.
-The evidence suggests that the availability of informal credit institutions relates to the stability of consumption in rural economies.
-This study assesses risk-coping behavior in informal credit transactions in the urban community in early twentieth-century Tokyo.
-I digitized and analyzed a unique daily longitudinal budget survey on factory worker households conducted in a representative manufacturing area.
-I found that the full insurance hypothesis is strongly rejected for overall consumption expenditure, meaning that the factory worker households in the historical metropolis suffered certain consumption losses when they faced idiosyncratic adverse income shocks.
-However, expenses on indispensable consumption categories were partially smoothed in the short run.
-Credit purchases in local retailers played an important role in mitigating adverse idiosyncratic shocks, whereas the other risk-mitigation strategies were impractical.
-Credit purchases did not require an explicit contract and thus often led to a loss in sales for the retailers.
-However, to remain competitive, retailers allowed credit transactions for their customers to remain regular customers.
-This shows that informal credit institutions in urban local communities functioned as informal insurance networks in rural economies.
-The first contribution of this study is to offer a case study of a metropolis in a developing economy.
-Urbanization is now a worldwide phenomenon.
-Moreover, the proportions of urban population in many developing countries, especially in East Asia, are reaching a similar level to the developed countries (Henderson and Turner 2020).
+Nonetheless, there is little evidence on the function of informal credit institutions in the historical context.
+Ogasawara (2024) observed the role of official financial institutions, such as postal savings and pawnshops, in compensating temporary income losses in the early 20th-century Osaka.
+[^5] Note that Osaka was middle scale city of approximately half the size of Tokyo City circa 1920 (Online Appendix Table A.1).
+However, the function of informal credit institutions in consumption smoothing has not been considered due to limited data availability.
+By considering the roles of both formal and informal financial institutions simultaneously, this current study documents the contribution of informal credit transactions by local retailers.
+Relatedly, an examination of how, and the extent to which, the local society could ensure household consumption provides valuable insights into existing urban history literature.
+After World War I, working-class households in Tokyo could establish and sustain their families (Nakagawa 1985).
+The results of this study suggest that the expansion of urbanization and credit transactions through the war underpins this traditional historical view.
+Another suggestion relates to the findings by Munshi and Rosenzweig (2016), who linked spatial wage inequality and men’s mobility in India with the availability of risk-mitigation networks among workers in the same caste.
+From this perspective, local credit institutions with lower barriers may motivate rural workers to urbanize, which may partially explain why Tokyo grew so fast circa WWI.
+This hypothesis may be worthy of testing in future research.
+The third contribution is to offer a case study of a metropolis in a developing economy.
+Extant development economics studies on consumption smoothing behavior have focused particularly on the rural economies (e.g., Townsend 1994; Ravallion and Chaudhuri 1997; Gertler and Gruber 2002).
+[^6] Recent field experimental studies also focus on village economies (Comola and Prina 2023; Macours et al.
+2022).
+While Townsend (1995), Skoufias and Quisumbing (2005), and Aguila et al.
+(2017) provided case studies on middle-sized contemporary economies, there is little evidence on agglomeration economies like prewar Tokyo.
+Urbanization is, however, now a worldwide phenomenon.
+The proportions of urban population in many developing countries, especially in East Asia, are reaching a similar level to the developed countries (Henderson and Turner 2020).
 This requires a deeper understanding of consumption behavior among urban working households in the economic development process.
-However, risk-mitigation strategies among urban worker households have been understudied.
-[^3] Extant studies have focused particularly on the rural economies (e.g., Townsend 1994; Ravallion and Chaudhuri 1997; Gertler and Gruber 2002).
-Recent field experimental studies also focus on village economies (Comola and Prina 2023; Macours et al.
-2022).
-Although several studies considered consumption smoothing among urban households in postwar Bangkok and prewar Osaka, the risk-mitigation mechanisms, particularly via credit institutions, are not fully evaluated.
-[^4] See Townsend (1995) and Ogasawara (2024), respectively.
-Osaka was middle scale city of approximately half the size of Tokyo City circa 1920 (Online Appendix Table A.1).
-(Skoufias and Quisumbing 2005; Aguila et al.
-2017) also provided case studies on middle-sized contemporary economies.
-However, they did not consider the roles of local credit institutions.
+Moreover, there is little research on the role of financial institutions among the household sector in agglomeration economies.
 This study considers a set of homogeneous households in a representative manufacturing area in prewar Tokyo, which provides a valuable empirical setting for testing risk-sharing strategies in the historical metropolis.
 [^7] Focusing on a sample that shares similar occupational and regional characteristics systematically mitigates bias in estimating income elasticity under the standard risk-sharing regressions due to the heterogeneity in the risk preferences (Schulhofer-Wohl 2011; Mazzocco and Saini 2012).
-My findings show that informal credit institutions in the local urban community could mitigate vulnerability among households similar to rural developing economies (Udry 1994; Fafchamps and Lund 2003).
-The second contribution is to add analytical perspectives to economic history studies.
-First, studies of how working-class households overcame their economic hardships during the late 19th to the early 20th century have focused primarily on Britain.
-The weight of evidence supports the roles of historical financial institutions and labor supply adjustments in compensating for income loss in working-class households.
-[^6] Horrell and Oxley (2000) highlight the important role of labor supply adjustments among family members to manage the income losses of the head in working-class households in late 19th-century Britain.
-Scott and Walker (2012) also found that interwar British working-class households had access to various institutions such as clothing clubs, hire-purchase, contractual savings, and insurance.
-The households with less earnings were likely to use credit-based institutions like clubs and hire-purchase, whereas those with higher incomes could use savings institutions.
-Kiesling (1996) and Boyer (1997) focused on the risk-mitigation strategies among the cotton workers households in Lancashire during the recession caused by the Civil War.
-However, there is little evidence of how those risk-coping strategies smoothed their consumption.
-Similarly, previous studies have used cross-sectional variations in income to estimate the elasticities of the temporary income sources.
-[^7] For example, Horrell and Oxley (2000) and James and Suto (2011) consider the deviations of actual earnings from the predicted earnings as the shocks.
-Scott and Walker (2012) employed the cross-sectional variations in the pooled cross-sectional sample to analyze the smoothness of expenditure (deviation from mean expenditure).
-An exceptional study is Saaritsa (2011), which shows the estimated elasticity of temporary income from a fixed effect one-way error component model using a panel household budget survey in 1928 Helsinki.
-The estimated elasticity is, therefore, potentially disturbed by household-specific unobservable confounding factors such as preference and learning ability.
-As risk-mitigation is a concept in the dynamic behavior of a specific household, the estimator derived from within variations of the units is more suitable for evaluating the roles of risk-mitigation strategies.
-This study implements a stylized econometric method for the panel dataset to offer comparable estimates for future studies.
-Second, this study adds micro-level evidence on consumer strategies during industrialization.
-Francks (2009) provides a comprehensive view of the long-run consumption history in Japan from the Edo era to the twentieth century.
-[^8] Several business history studies analyze the relationship between consumption and production.
-A representative work is Gordon (2013), which investigated the expansion of the sewing machine from many perspectives, including production, distribution, and gender social norms in Japan.
-However, these studies have not focused on the households’ risk-mitigation strategies.
-Some studies investigated the relationships between household chores and consumption strategies in interwar households (Tanimoto 2011, 2012).
-This study employs an alternative analytical view by focusing on short-run consumption strategies.
-An examination of how, and the extent to which, the local society could ensure household consumption provides valuable insights into consumption history literature.
-Third, several financial historical studies provide supply-side (lender-side) evidence on the functions of the micro-financial institutions for the workers (O’Connell and Reid 2005).
-This study is the first to document the function of informal credit transactions by local retailers in the context of consumption smoothing.
-Munshi and Rosenzweig (2016) linked spatial wage inequality and men’s mobility in India with the availability of risk-mitigation networks among workers in the same caste.
-From this perspective, lower barriers of local credit institutions may motivate rural workers to urbanize, which may partially explain why Tokyo grew so fast circa WWI.
-This hypothesis requires testing in future research.
+Caution is needed when comparing results, as the modern financial system differs in many ways from the past.
+However, my findings show that informal credit institutions in the local urban community could mitigate vulnerability among households similar to rural developing economies (Udry 1994; Fafchamps and Lund 2003).
 2 Historical Background
@@
 The population in Tokyo City has been growing since the end of the 19th century, and industrialization due to WWI spurred its growth.
-Tokyo City’s population share increased from $2.5$% in 1889 to approximately $4$% by 1920.141414In 1920, there were more than $10,000$ municipalities in Japan, suggesting how large Tokyo was.
+Tokyo City’s population share increased from $2.5$% in 1889 to approximately $4$% by 1920.131313In 1920, there were more than $10,000$ municipalities in Japan, suggesting how large Tokyo was.
 There was a difference of about 1 million between Tokyo and Osaka, the second-largest city by population.
@@
 While the sex ratio by age group in Tsukishima shows a similar trend to that of Tokyo and Kyobashi, the ratios for the 1414141419191919 and 2020202029292929 age bins are higher than those.
-Overall, these statistics reflect the fact that Tsukishima was a typical industrial area and many young male workers were employed there.
+Overall, these statistics reflect the fact that the typical household structure is a couple with two children.
 The slightly smaller household size in Tsukishima can be attributed to the relatively large number of young factory worker households.
@@
 The estimation strategy used for testing consumption-smoothing and risk-coping strategies requires panel structures for aggregate income and expenditure (Section 4).
-I use the semi-monthly series for baseline analysis, given that the head’s income was paid twice monthly.
+I use the semi-monthly series for the baseline analysis, as the head’s income was paid twice a month.
 [^30] Section 3.3 provides details on the definitions of aggregate income and expenditure.
@@
 Finally, I excluded a household for which no information on payments to the credit purchases was recorded.
-Consequently, my THBS sample for the semi-monthly series included $33$ households with panel structures from January 12 to November 11, 1919.323232I also used the adjusted monthly panel dataset for sub-analysis in Section 4.
+Consequently, my THBS sample for the semi-monthly series included $33$ households with panel structures from January 12 to November 11, 1919.313131I also used the adjusted monthly panel dataset for sub-analysis in Section 4.
 Following the same trimming steps, my adjusted monthly panel dataset included $26$ households for the same period.
@@
 Given the improvements in both income and expenditure measures, the net income series shows a much smoother trend in Figure 4d, which is unavailable under the unadjusted series (Figure 4f).
-In Section 4.1, I will demonstrate how these measurement errors in the unadjusted dataset attenuate the estimates.
+Online Appendix B.8 shows that these measurement errors in the unadjusted dataset strongly attenuate the estimates.
 3.4 Data Description
@@
 [^45] The family size variable was omitted in several regressions due to the multicollinearity in the fixed-effects models given its small within variations.
-Although the number of family members in different age bins is available from the THBS, those family composition variables are also time-constant and were absorbed by the household fixed effect.
+I then used the share of children below primary school age range as an alternative time-varying preference shock variable.
+Including the child share control variable does not influence the results, and the estimates are materially similar to those from my baseline specification (Online Appendix Table B.6).
 Panel C of Table 3 shows the summary statistics on the family size variable.
@@
 Table 4: Results of Estimating Income Elasticities
-(1) Semi-monthly Panels (2) Adj. Monthly Panels (3) Unadj. Monthly Panels
-Disposable income Obs. Disposable income Obs. Disposable income Obs.
-Coef. Std. error Coef. Std. error Coef. Std. error
-Total consumption 0.358 [0.033]*** 278 0.433 [0.059]*** 123 0.261 [0.054]*** 141
-      Food 0.136 [0.030]*** 259 0.293 [0.066]*** 114 0.100 [0.072] 132
-      Housing 0.099 [0.415] 134 0.013 [0.084] 105 -0.263 [0.167] 124
-      Utilities -0.003 [0.138] 223 0.343 [0.236] 110 0.197 [0.334] 130
-      Furniture 0.426 [0.223]* 169 -0.114 [0.706] 98 0.250 [0.339] 112
-      Clothes 0.378 [0.190]* 238 0.347 [0.424] 112 0.806 [0.262]*** 132
-      Education 0.101 [0.051]* 236 0.162 [0.128] 110 0.069 [0.146] 128
-      Medical expenses 0.039 [0.066] 258 0.194 [0.120] 114 -0.128 [0.094] 132
-      Entertainment expenses 0.307 [0.084]*** 250 0.716 [0.181]*** 113 0.194 [0.153] 132
-      Transportation 0.278 [0.110]** 187 0.605 [0.264]** 97 0.362 [0.291] 115
-      Gifts 0.523 [0.106]*** 234 0.469 [0.197]** 113 0.622 [0.306]* 130
-      Miscellaneous -0.076 [0.147] 214 0.522 [0.252]** 110 0.246 [0.267] 126
+(1) Semi-monthly Panels (2) Adjusted Monthly Panels
+Disposable income Obs. Disposable income Obs.
+Coef. Std. error Coef. Std. error
+Total consumption 0.358 [0.033]*** 278 0.433 [0.059]*** 123
+      Food 0.136 [0.030]*** 259 0.293 [0.066]*** 114
+      Housing 0.099 [0.415] 134 0.013 [0.084] 105
+      Utilities -0.003 [0.138] 223 0.343 [0.236] 110
+      Furniture 0.426 [0.223]* 169 -0.114 [0.706] 98
+      Clothes 0.378 [0.190]* 238 0.347 [0.424] 112
+      Education 0.101 [0.051]* 236 0.162 [0.128] 110
+      Medical expenses 0.039 [0.066] 258 0.194 [0.120] 114
+      Entertainment expenses 0.307 [0.084]*** 250 0.716 [0.181]*** 113
+      Transportation 0.278 [0.110]** 187 0.605 [0.264]** 97
+      Gifts 0.523 [0.106]*** 234 0.469 [0.197]** 113
+      Miscellaneous -0.076 [0.147] 214 0.522 [0.252]** 110
 ***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors in brackets are clustered at the household level.
@@
 The regressions of the 11 measures of log-transformed consumption on log-transformed disposable income, the family size control, household-fixed effect, and time-fixed effect are listed.
-Columns 1, 2, and 3 show the results for the semi-monthly, adjusted monthly, and unadjusted monthly panel datasets, respectively.
+Columns 1 and 2 show the results for the semi-monthly and adjusted monthly panel datasets, respectively.
 The estimated coefficients on log-transformed disposable income are listed in the columns named “Coef.”.
@@
 In my result, however, the estimates for the adjusted monthly panels are greater than those for the semi-monthly panels in a large part of subcategories, suggesting that households faced more difficulties with smoothing their consumption in the long run.
-Finally, I assess the validity of the aggregation method used in this study.
-Adjusting the aggregate data to account for the timing of paydays reveals more transparent relationships between income and expenditure than the unadjusted data (Section 3.3).
-To demonstrate this mechanism, I listed the estimates from the unadjusted monthly panel dataset in Column (3).
-The estimate for total consumption is $0.26$, which is much lower than the estimate for the adjusted series ($0.43$).
-In addition, most subcategories show systematically smaller estimates than those from the adjusted data listed in Column (2).
-For example, the estimate for the food category is now $0.10$ and statistically insignificant.
-This means that measurement errors in the unadjusted aggregation lead to strong attenuation bias, which results in false negatives (i.e., false rejections of the null of full risk-sharing hypothesis) in most of the consumption categories.
-I, therefore, specified a generating process of systematic measurement errors in the aggregate survey data, which has attracted attention in the macroeconomics literature (Cochrane 1991; Nelson 1994; Dynarski et al.
-1997; Gervais and Klein 2010).
-Moreover, several subcategories select values close to one or large negative values, which are no longer economically interpretable.
-[^59] For example, the estimate for housing was negative.
-The households paid a fixed rent amount monthly.
-Given that the rent payday was fixed, the shifts in wage paydays might create cells with less income with a fixed amount of rent.
-In the case of seasonal consumption categories, the shifts may coincide with the timing of seasonal spending.
-The THBS households tended to purchase work clothes in March to prepare for the start of the new fiscal year in April.
-They also spent more on clothing in June, likely to buy summer clothes.
-The shift of paydays to these months may have created a strong but false positive correlation.
-This means that the miss-assignment of paydays in the unadjusted data may cause critical statistical inference issues in a few subcategories.
 4.2 Risk-coping Mechanisms
@@
 An interesting finding—which is not explicitly described in the historical sources—is that the credit purchases in the local society and the informal gifts from personal networks might have provided some aid to the factory workers.
-4.3 Event-Study Analysis: Response to Adverse Health Shocks
+4.3 Response to Adverse Health Shocks
 The foregoing analyses used the within variations in the earnings to estimate income elasticity, meaning that the identification depends both on the adverse and favorable idiosyncratic shocks.
@@
 Results: To What Extent do Illnesses Affect Consumption?
-I first consider a reduced-form specification using the head’s illness as the exposure variable in equation 1 following Cochrane (1991).
-Column (1) in Table 7 shows the results.
-Overall, the coefficients on the exposure variable are negative, supporting evidence that adverse health shocks on the heads reduce consumption levels.
-Next, I explicitly model the structure underlying the relationship between the health shocks and earnings using the reduced-form equation of (3).
-If the negative impacts of the health shocks on consumption were derived from the income losses, the results from the structural estimation approach correspond to those in Column (1).
-The estimates from the structural equation are listed in Column (2) of Table 7.
-As shown, the results are congruent with the reduced-form results in Column (1).
-Importantly, the estimates are systematically larger than those reported in Table 4.
-The estimates for total and food consumption categories are $0.543$ and $0.355$, respectively, whereas the baseline estimates are $0.358$ and $0.136$.
-While this result partly reflects the bias reduction under the instrumental variable estimation, it is also noteworthy that this system is designed to reduce the bias by specifying the channel from adverse (instead of favorable) income shocks.
-For example, expenses on the allowance for children could be easily cut during unfavorable times but may not respond to favorable shocks.
-Similarly, the expenses for furniture are likely to be cut, but the households might not buy those frequently, meaning that there are no sufficient cases for reacting to rare health shock events.
-[^68] In other words, furniture may have only been purchased when the households had extra income.
-This may partly be reflected in the smaller value of the first-stage $F$-statistics in the furniture subcategory.
-[^69] Note that the logarithmic transformation leaves the cells with expenses on the furniture.
-Thus, the reduced-form equation captures the relationship between the health shocks and earnings in the cells with any expenses on furniture.
-While the first-stage $F$-statitics is $35.7$ in the linear-linear model, the second-stage result is robust against the functional form assumption.
-This may reflect the asymptotic theoretical advantage in the panel data model under nearly weak instruments case (Staiger and Stock 1997; Cai et al.
-2012; Baltagi et al.
-2012).
-The same argument is applicable to the housing category.
-A similar interpretation can be applied to the transportation category because family travel was not a frequent event and thus, unlikely to coincide with the timing of illness.
-By contrast, the entertainment expenses were still likely to be cut in case of illness, which might reflect the curtailment of tobacco, theater-going, and the allowance for heads.
-This may be another explanation for the difference in the results from my baseline results in Table 4.
-While the estimates for the clothes and gifts categories now exceed one, the null of equality to the population mean of one is not statistically rejected for both cases.
-Although the large estimates may be partly due to the operation under the instrument based on rare health shock events, the findings are materially like my baseline results—both categories are sensitive to the shocks as are those of the dispensable categories.
 Table 7: Results of Estimating Income Elasticities:
@@
 The first-stage $F$-statistics on the health shock variable are listed in the final column.
+I first consider a reduced-form specification using the head’s illness as the exposure variable in equation 1 following Cochrane (1991).
+Column (1) in Table 7 shows the results.
+Overall, the coefficients on the exposure variable are negative, supporting evidence that adverse health shocks on the heads reduce consumption levels.
+Next, I explicitly model the structure underlying the relationship between the health shocks and earnings using the reduced-form equation of (3).
+If the negative impacts of the health shocks on consumption were derived from the income losses, the results from the structural estimation approach correspond to those in Column (1).
+The estimates from the structural equation are listed in Column (2) of Table 7.
+As shown, the results are congruent with the reduced-form results in Column (1).
+Importantly, the estimates are systematically larger than those reported in Table 4.
+The estimates for total and food consumption categories are $0.543$ and $0.355$, respectively, whereas the baseline estimates are $0.358$ and $0.136$.
+While this result partly reflects the bias reduction under the instrumental variable estimation, it is also noteworthy that this system is designed to reduce the bias by specifying the channel from adverse (instead of favorable) income shocks.
+For example, expenses on the allowance for children could be easily cut during unfavorable times but may not respond to favorable shocks.
+Similarly, the expenses for furniture are likely to be cut, but the households might not buy those frequently, meaning that there are no sufficient cases for reacting to rare health shock events.
+[^66] In other words, furniture may have only been purchased when the households had extra income.
+This may partly be reflected in the smaller value of the first-stage $F$-statistics in the furniture subcategory.
+[^67] Note that the logarithmic transformation leaves the cells with expenses on the furniture.
+Thus, the reduced-form equation captures the relationship between the health shocks and earnings in the cells with any expenses on furniture.
+While the first-stage $F$-statitics is $35.7$ in the linear-linear model, the second-stage result is robust against the functional form assumption.
+This may reflect the asymptotic theoretical advantage in the panel data model under nearly weak instruments case (Staiger and Stock 1997; Cai et al.
+2012; Baltagi et al.
+2012).
+The same argument is applicable to the housing category.
+A similar interpretation can be applied to the transportation category because family travel was not a frequent event and thus, unlikely to coincide with the timing of illness.
+By contrast, the entertainment expenses were still likely to be cut in case of illness, which might reflect the curtailment of tobacco, theater-going, and the allowance for heads.
+This may be another explanation for the difference in the results from my baseline results in Table 4.
+While the estimates for the clothes and gifts categories now exceed one, the null of equality to the population mean of one is not statistically rejected for both cases.
+Although the large estimates may be partly due to the operation under the instrument based on rare health shock events, the findings are materially like my baseline results—both categories are sensitive to the shocks as are those of the dispensable categories.
 Results: How did the Households Compensate for Losses from illness?
-I tested the risk-coping strategies for mitigating adverse health shocks in Table 8.
-Column 1 lists the reduced-form results in the same manner as Table 5.
-The estimated coefficients on the net income variables are positive in most cases, indicating the households’ compensating strategies against the shocks.
-The estimate for the credit purchase category is weakly statistically significant, with the largest magnitude.
-It suggests that the one-day increase in the head’s illness increased the net credit purchases by $1.29$ yen, which accounts for approximately half of their average daily wages ($1.29/2.56$).
-Column 2 presents the estimates from the structural equation using the head’s earnings as the endogenous variable.
-A similar but clearer result is obtained—the estimate for the credit purchase category is approximately four times greater than that reported in Column (1) of Table 5.
-It indicates that a one yen decrease in the head’s earnings increases the net income from credit purchases by $0.55$ yen.
-This magnitude is consistent with the direct effect of the health shocks on the net credit purchase estimated in Column (1) of Table 8.707070This means that $2.56\times 0.55=1.4$ yen, which is close to $1.3$ yen in Column (1).
-The estimates for the savings and gifts categories are still negative but now statistically insignificant, supporting the evidence that credit purchase is the primary risk-mitigation strategy for dealing with health shocks.
-Since withdrawing and receiving gifts to use the money at the retailers takes more time (steps), the households likely prefer to use credit purchases.
-Overall, the event-study analysis using adverse health shocks confirms the baseline findings and provides a sharper illustration of consumption-smoothing strategies in the case of adverse idiosyncratic shock events.
-My results show that the unique available response to idiosyncratic income shocks among the THBS households was to use credit purchases.
-The available compensation via credit purchases accounts for approximately half of their temporary income losses, which partially contributes to smoothing their food consumption in the short term.
 Table 8: Results of Testing the Risk-coping Mechanisms:
@@
 All dependent variables are in yen.
+I tested the risk-coping strategies for mitigating adverse health shocks in Table 8.
+Column 1 lists the reduced-form results in the same manner as Table 5.
+The estimated coefficients on the net income variables are positive in most cases, indicating the households’ compensating strategies against the shocks.
+The estimate for the credit purchase category is weakly statistically significant, with the largest magnitude.
+It suggests that the one-day increase in the head’s illness increased the net credit purchases by $1.29$ yen, which accounts for approximately half of their average daily wages ($1.29/2.56$).
+Column 2 presents the estimates from the structural equation using the head’s earnings as the endogenous variable.
+A similar but clearer result is obtained—the estimate for the credit purchase category is approximately four times greater than that reported in Column (1) of Table 5.
+It indicates that a one yen decrease in the head’s earnings increases the net income from credit purchases by $0.55$ yen.
+This magnitude is consistent with the direct effect of the health shocks on the net credit purchase estimated in Column (1) of Table 8.686868This means that $2.56\times 0.55=1.4$ yen, which is close to $1.3$ yen in Column (1).
+The estimates for the savings and gifts categories are still negative but now statistically insignificant, supporting the evidence that credit purchase is the primary risk-mitigation strategy for dealing with health shocks.
+Since withdrawing and receiving gifts to use the money at the retailers takes more time (steps), the households likely prefer to use credit purchases.
+Overall, the event-study analysis using adverse health shocks confirms the baseline findings and provides a sharper illustration of consumption-smoothing strategies in the case of adverse idiosyncratic shock events.
+My results show that the unique available response to idiosyncratic income shocks among the THBS households was to use credit purchases.
+The available compensation via credit purchases accounts for approximately half of their temporary income losses, which partially contributes to smoothing their food consumption in the short term.
 5 Credit System in Local Economy
@@
 The information on the credit purchases is personal, as in the case of state-contingent loans in the rural credit market (Udry 1994) and thus, not well documented.
-Therefore, systematic information on credit purchases is unavailable.
 Fortunately, however, I found that some official survey reports documented information on this credit institution.
@@
 260; 263).
+Thus, credit purchases in Tokyo were neither region- nor religion-specific customs.
 A survey of retailers in 1935 published by the Tokyo City Chamber of Commerce and Industry provides more detailed information.
@@
 By contrast, if retailers did not allow households with high-income volatility to trade on credit, the estimate should be positive and reasonably large.
-Table 10: Volatility of Earnings and Use of Credit Purchases
-Dependent Variable: Indicator Variable for
-Non-CP Users Non-CP + Low-CP Users
-(1) (2) (3) (4) (5) (6)
-CV of Head’s Earnings $0.003$ $-0.002$ $-0.027$ $0.019$ $0.008$ $-0.005$
-$[0.012]$ $[0.014]$ $[0.050]$ $[0.013]$ $[0.017]$ $[0.049]$
-Sample mean of the DV $0.12$ $0.12$ $0.12$ $0.24$ $0.24$ $0.24$
-Model LPM LPM Probit LPM LPM Probit
-Number of THBS households 33 33 33 33 33 33
-Family Size Yes Yes Yes Yes Yes Yes
-Children aged 6–12 (%) Yes Yes Yes Yes Yes Yes
-Children aged 13–16 (%) Yes Yes Yes Yes Yes Yes
-Adults aged 17+ (%) Yes Yes Yes Yes Yes Yes
-Head’s age No Yes Yes No Yes Yes
-Head’s earnings No Yes Yes No Yes Yes
-Zero slope ($p$-value) 0.467 0.718 0.336 0.515 0.519 0.451
-***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. Robust standard errors are reported in brackets.
-Notes: The dependent variable in Columns (1)–(3) is the indicator variable for the households who did not use any credit purchases during the sample period (non-CP users).
-The dependent variable in Columns (4)–(6) is the indicator variable for the non-CP users and households less than the first quartile in the sample average of credit purchases.
-Columns (1)–(2) and (4)–(5) use the linear probability model (LPM), whereas Columns (3) and (6) use the Probit model.
-“CV of Head’s Earnings” indicates the coefficient of variations in the head’s semi-monthly earnings.
-The sample mean of the CV is $5.76$ (Std.
-Dev.
-= $5.72$).
-All the regressions include the family composition variables: average family size, share of children aged 6–12, share of children aged 13–16, and share of adults aged 17+ (share of children aged 0–5 is used as a reference group).
-The average head’s age and average head’s earnings are added to the control variables in Columns 3 and 6.
-$F$-statistics $p$-values for the null of the zero-slope hypothesis are reported in Columns (1)–(2) and (4)–(5).
-$\chi^{2}$-statistics $p$-values for the null of the zero-slope hypothesis are listed in Columns (3) and (6).
-The standard errors for the LPM are based on the HC2 estimator proposed by Horn et al.
-(1975).
-Column (1) of Table 10 presents the result for Equation 4.
-The estimated coefficient is close to zero and statistically insignificant.
-A potential concern may be the correlations between the volatility of earnings and the head’s age and earnings level.
-Column (2) shows the result from the extended model, including the head’s age and earnings in the controls, which shows the robustness of the baseline estimate in Column (1).
-Another potential issue is the functional form assumption in the linear probability model.
-Column (3) then lists an alternative non-linear model, including the full set of control variables used in Column (2).
-It confirms that the estimate is not statistically significantly different from zero.
-Next, I changed the threshold for the dependent indicator variable to consider a wider range of households including the households with less than the first quartile of the average credit purchases.
-Columns (4)–(6) summarize the results in the same column layouts as Columns (1)–(3), confirming that the estimates are still close to zero in all the specifications.
-Finally, the last row in this table shows that the null of the joint zero slope hypothesis is strongly rejected in all the regressions.
-Overall, there is no apparent trend in the characteristics of households that use credit for their purchases.
-This suggests that retailers may not have considered the volatility of earnings when providing credit to factory workers.
-This is consistent with the retailers’ motivation to secure their local customers discussed earlier, and with the fact that factory workers commonly used credit for purchases.
-[^78] Despite this, low-income households might have had more difficulty using credit for purchases.
-A famous essay drawn by a primary school student in the 1930s depicts an example of a rice retailer who refused to buy rice under credit purchase (Toyoda 1969, pp.
-346–355).
-The results so far indicate that retailers were generous in considering the earnings volatility risk.
+Online Appendix C.5 summarizes the results.
+Shortly, the estimated coefficient ($\hat{\varkappa}_{1}$) is close to zero and statistically insignificant in different specifications.
+This result remains unchanged when alternative thresholds for the dependent indicator variable are used to include the households with less than the first quartile of the average credit purchases.
+This suggests that retailers may be generous in considering the volatility of earnings when providing credit to factory workers, which is consistent with the retailers’ incentive to secure their local customers.
 This, in turn, implies that retailers may factor risk into retail prices or by charging interest on their loans.
@@
 They bought various goods in greengrocers, fish stores, and general stores between February 1 and February 14, 1919, and paid exactly the same amounts on the 15th.
-[^80] The repayment period in credit purchases among retailers and wholesalers (footnote 74) was often set at approximately one month (Tokyo City Office 1932, pp.
+[^77] The repayment period in credit purchases among retailers and wholesalers (footnote 72) was often set at approximately one month (Tokyo City Office 1932, pp.
 157-159).
@@
 33).
-He had the experience of conducting the first household survey in Japan in 1916898989This is called the “Survey of Household Income and Expenditure of 20 Artisans and Workers in Tokyo,” which was inspired by Charles Booth (Miyoshi 1980, p.
+He had the experience of conducting the first household survey in Japan in 1916868686This is called the “Survey of Household Income and Expenditure of 20 Artisans and Workers in Tokyo,” which was inspired by Charles Booth (Miyoshi 1980, p.
 33).
@@
 3).
-He was a member of the seventh subcommittee (Rural Sanitation Subcommittee) of the Ministry of Home Affairs’ Health and Sanitation Investigation Committee since 1916.909090The Health and Sanitation Investigation Committee of the Ministry of Home Affairs was established on June 27, 1916, under the second Shigenobu Okuma Cabinet by issuing Imperial Ordinance No.
+He was a member of the seventh subcommittee (Rural Sanitation Subcommittee) of the Ministry of Home Affairs’ Health and Sanitation Investigation Committee since 1916.878787The Health and Sanitation Investigation Committee of the Ministry of Home Affairs was established on June 27, 1916, under the second Shigenobu Okuma Cabinet by issuing Imperial Ordinance No.
 172.
@@
 The labor survey was managed by Yoshitsuru Yamana, a graduate of Tokyo Imperial University who contributed to the organization of labor unions.
-In addition, Toyotaro Miyoshi, a sociology student at the Tokyo Imperial University, and others became temporary employees (see footnote 55).
+In addition, Toyotaro Miyoshi, a sociology student at the Tokyo Imperial University, and others became temporary employees (see footnote 54).
 At the time, the Health Bureau of the Ministry of the Interior conducted several sanitary surveys of rural areas.
@@
 I calculate the weighted average of the daily wage using the number of factory workers in each age bin (12–14 and 15+) and in each category as weights.
-Thus, I shift the functional polynomial function by this amount of difference to yield the proxy distribution for 1918.100100100This operation, therefore, assumes that the wage-by-age distributions in 1918 and 1924 are reasonably similar.
+Thus, I shift the functional polynomial function by this amount of difference to yield the proxy distribution for 1918.979797This operation, therefore, assumes that the wage-by-age distributions in 1918 and 1924 are reasonably similar.
 In the LST (i.e., the statistic on which the predicted distribution is based), the weighted average for those aged $14$ and older was $3.25$ yen.
@@
 In other words, the shifted curve shall get closer to the actual (unobservable) wage distribution as age increases.
-The long dashed curve in Figure B.8a examines this relationship by fitting another fractional polynomial model using the 12–14-year old bin figure of the ASCT and the predicted wages for 25–65 from the shifted fractional polynomial.101101101The fitted model is as follows: $\hat{y}^{\textit{Daily Wage}}=-1.19+0.14a^{1}-0.002a^{2},\quad R^{2}=0.98.$ The deviance is reported to be $-149$.
+The long dashed curve in Figure B.8a examines this relationship by fitting another fractional polynomial model using the 12–14-year old bin figure of the ASCT and the predicted wages for 25–65 from the shifted fractional polynomial.
+[^98] The fitted model is as follows: $\hat{y}^{\textit{Daily Wage}}=-1.19+0.14a^{1}-0.002a^{2},\quad R^{2}=0.98.$ The deviance is reported to be $-149$.
 I used the 25–65 age range for the prediction based on the age bins used in the LST.
@@
 where $w^{\textit{Regular}}>0$ is the monthly regular wage and $w^{\textit{Ancillary}}\geq 0$ indicates the monthly ancillary wage.
-Online Appendix B.6.2 shows that $w^{\textit{Ancillary}}$ can be calculated as the product of the monthly regular wage and weight for the ancillary wage, $\vartheta\in(0,1)$.102102102Strictly, this weight may not be required to be in this range.
+Online Appendix B.6.2 shows that $w^{\textit{Ancillary}}$ can be calculated as the product of the monthly regular wage and weight for the ancillary wage, $\vartheta\in(0,1)$.
+[^99] Strictly, this weight may not be required to be in this range.
 I use this range because all the cases reported in Kitazawa (1924) take the weights between zero and one.
@@
 These illustrations support the evidence that the adjusted semi-monthly datasets can offer cleaner within variations in income and expenditure because they can fix the measurement errors and reflect the households’ consumption strategies based on their latest incomes measured in the same semi-month cell.
-B.8 Distribution of the Idiosyncratic Shocks
+B.8 Testing Attenuation Bias in the Unadjusted Panel Dataset
+In this subsection, I assess the validity of the aggregation method used in this study.
+Adjusting the aggregate data to account for the timing of paydays reveals more transparent relationships between income and expenditure than the unadjusted data (Section 3.3).
+To demonstrate this mechanism, I listed the estimates from the unadjusted monthly panel dataset in Column (2) of Table B.5.
+The estimate for total consumption is $0.26$, which is much lower than the estimate for the adjusted series ($0.43$) in Column (1).
+In addition, most subcategories show systematically smaller estimates than those from the adjusted data listed in Column (2).
+For example, the estimate for the food category is now $0.10$ and statistically insignificant.
+This means that measurement errors in the unadjusted aggregation lead to strong attenuation bias, which results in false negatives (i.e., false rejections of the null of full risk-sharing hypothesis) in most of the consumption categories.
+I, therefore, specified a generating process of systematic measurement errors in the aggregate survey data, which has attracted attention in the macroeconomics literature (Cochrane 1991; Dynarski et al.
+1997; Gervais and Klein 2010; Nelson 1994).
+Moreover, several subcategories select values close to one or large negative values, which are no longer economically interpretable.100100100For example, the estimate for housing was negative.
+The households paid a fixed rent amount monthly.
+Given that the rent payday was fixed, the shifts in wage paydays might create cells with less income with a fixed amount of rent.
+In the case of seasonal consumption categories, the shifts may coincide with the timing of seasonal spending.
+The THBS households tended to purchase work clothes in March to prepare for the start of the new fiscal year in April.
+They also spent more on clothing in June, likely to buy summer clothes.
+The shift of paydays to these months may have created a strong but false positive correlation.
+This means that the miss-assignment of paydays in the unadjusted data may cause critical statistical inference issues in a few subcategories.
+Table B.5: Testing Attenuation Bias in the Unadjusted Panel Dataset
+(1) Adjusted Monthly Panels (2) Unadjusted Monthly Panels
+Disposable income Obs. Disposable income Obs.
+Coef. Std. error Coef. Std. error
+Total consumption 0.433 [0.059]*** 123 0.261 [0.054]*** 141
+      Food 0.293 [0.066]*** 114 0.100 [0.072] 132
+      Housing 0.013 [0.084] 105 -0.263 [0.167] 124
+      Utilities 0.343 [0.236] 110 0.197 [0.334] 130
+      Furniture -0.114 [0.706] 98 0.250 [0.339] 112
+      Clothes 0.347 [0.424] 112 0.806 [0.262]*** 132
+      Education 0.162 [0.128] 110 0.069 [0.146] 128
+      Medical expenses 0.194 [0.120] 114 -0.128 [0.094] 132
+      Entertainment expenses 0.716 [0.181]*** 113 0.194 [0.153] 132
+      Transportation 0.605 [0.264]** 97 0.362 [0.291] 115
+      Gifts 0.469 [0.197]** 113 0.622 [0.306]* 130
+      Miscellaneous 0.522 [0.252]** 110 0.246 [0.267] 126
+***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors in brackets are clustered at the household level.
+Notes: This table shows the results of Equation 1.
+The regressions of the 11 measures of log-transformed consumption on log-transformed disposable income, the family size control, household-fixed effect, and time-fixed effect are listed.
+Columns 1 and 2 show the results for the adjusted monthly and unadjusted monthly panel datasets, respectively.
+The estimates in column (1) are the same as those reported in Table 4.
+The estimated coefficients on log-transformed disposable income are listed in the columns named “Coef.”.
+B.9 Considering Child Share as a Preference Shock Variable
+As explained in the main text, the child share variable exhibits significantly more within-variations than that of the family size.
+Table B.6 presents the results from specifications that include the share of children under the primary school age range, specifically those aged 0–5, as an alternative preference shock variable.
+Panels A and B show the results for the consumption smoothing and risk-coping strategy regressions, respectively.
+The results are materially similar to the main results presented in the main text (Tables 4 and 5).
+Note that the results remain robust to different definitions of age bins (not reported).
+Table B.6: Results for the Alternative Specification:
+Conditioning on the Share of Children Aged 0–5
+(1) Semi-monthly Panels (2) Adjusted Monthly Panels
+Disposable income Obs. Disposable income Obs.
+Panel A: Consumption Smoothing Coef. Std. error Coef. Std. error
+Total consumption 0.361 [0.034]*** 278 0.440 [0.058]*** 123
+      Food 0.130 [0.028]*** 259 0.277 [0.062]*** 114
+      Housing 0.097 [0.413] 134 0.015 [0.079] 105
+      Utilities -0.005 [0.140] 223 0.339 [0.232] 110
+      Furniture 0.431 [0.234]* 169 0.172 [0.665] 98
+      Clothes 0.381 [0.189]* 238 0.319 [0.412] 112
+      Education 0.103 [0.050]** 236 0.166 [0.127] 110
+      Medical expenses 0.037 [0.066] 258 0.182 [0.122] 114
+      Entertainment expenses 0.304 [0.085]*** 250 0.695 [0.176]*** 113
+      Transportation 0.281 [0.110]** 187 0.613 [0.267]** 97
+      Gifts 0.524 [0.107]*** 234 0.476 [0.192]** 113
+      Miscellaneous -0.079 [0.148] 214 0.506 [0.253]* 110
+(1) Semi-monthly Panels (2) Adj. Monthly Panels
+Head’s earning Obs. Head’s earning Obs.
+Panel B: Risk-Coping Strategy Coef. Std. error Coef. Std. error
+Panel B-1: Savings, insurance, borrowing, and gifts (yen)
+Net savings -0.043 [0.018]** 289 -0.035 [0.028] 124
+Net insurance 0.001 [0.013] 289 0.002 [0.032] 124
+Net borrowing -0.021 [0.027] 289 -0.026 [0.017] 124
+Net credit purchase -0.129 [0.039]*** 289 -0.021 [0.042] 124
+Net gifts -0.052 [0.020]** 289 -0.006 [0.029] 124
+Panel B-2: Labor supply adjustments & sales of miscellaneous assets (yen)
+Other member’s earnings 0.001 [0.010] 289 0.036 [0.033] 124
+Sales of miscellaneous assets -0.000 [0.000] 289 0.000 [0.001] 124
+***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors in brackets are clustered at the household level.
+Notes: Panel A: This table shows the results of Equation 1.
+The regressions of the 11 measures of log-transformed consumption on log-transformed disposable income, the proportion of children aged 0–5, household-fixed effect, and time-fixed effect are listed.
+Columns 1 and 2 show the results for the semi-monthly and adjusted monthly panel datasets, respectively.
+The estimated coefficients on log-transformed disposable income are listed in the columns named “Coef.”.
+Panel B: This table shows the results of equation 2.
+The proportion of children aged 0–5, household fixed effects, and time-fixed effects are included in all the regressions.
+Columns 1 and 2 show the estimates for the semi-monthly and adjusted monthly panel datasets, respectively.
+The estimated coefficients on the head’s earnings are listed in the “Coef.” columns.
+Panel B-1: Each net income variable is defined as the difference between income and expenses.
+“Net savings” refers to the difference between the amount withdrawn and the total deposits.
+“Net insurance” refers to the amount received from insurance minus the expenses paid for insurance.
+“Net borrowing” refers to the amount of borrowing minus the debt payments.
+“Net credit purchase” refers to the amount of credit purchases minus credit redeemed.
+“Net gifts” refers to the total received amounts, including both pecuniary and non-pecuniary gifts, minus the payments for gifts.
+Panel B-2: “Other member’s earnings” refers to the total income earned by all family members except for the head of the household.
+“Sales of miscellaneous assets” includes the sale of daily miscellaneous goods such as newspapers and empty bottles.
+All dependent variables are in yen.
+B.10 Distribution of the Idiosyncratic Shocks
 Figure B.11: Distribution of the Idiosyncratic Shocks
@@
 Source: Created by the author using the THBS sample.
-Figure B.11 illustrates the distribution of the idiosyncratic shocks predicted using the semi-monthly panel dataset.103103103I use the semi-monthly panel dataset to overview the entire distribution of the idiosyncratic shocks rather than the daily panel datasets.
+Figure B.11 illustrates the distribution of the idiosyncratic shocks predicted using the semi-monthly panel dataset.101101101I use the semi-monthly panel dataset to overview the entire distribution of the idiosyncratic shocks rather than the daily panel datasets.
 This is because the daily panels include a set of consecutive household-date cells with no income.
@@
 However, this figure also indicates that the residuals have rich variations, supporting that the THBS households had experienced both favorable and adverse idiosyncratic shocks during the survey periods.
-B.9 Unit Root and Autocorrelations
+B.11 Unit Root and Autocorrelations
 Table B.7 presents the results of the Fisher-type panel unit root tests for the total expenditure in the semi-monthly (Column 1) and adjusted monthly (Column 2) panel datasets.
@@
 Source: Calculated by the author using the THBS sample.
-B.10 Risk Preferences
+B.12 Risk Preferences
 The potential influences of heterogeneous risk preference on standard risk-sharing regressions are discussed in a few works (Schulhofer-Wohl 2011; Mozzoco and Saini 2012).
@@
 This test is valid under a particular utility function.
-Specifically, the bias could be raised in the case that the coefficient of relative risk aversion interacts with the aggregate shock in the first-order condition for consumption because the interaction means that the aggregate shock cannot be removed using the time-fixed effect in the risk-sharing regressions.104104104Specifically, Schulhofer-Wohl (2011, pp.
+Specifically, the bias could be raised in the case that the coefficient of relative risk aversion interacts with the aggregate shock in the first-order condition for consumption because the interaction means that the aggregate shock cannot be removed using the time-fixed effect in the risk-sharing regressions.102102102Specifically, Schulhofer-Wohl (2011, pp.
 928–929) considers the utility maximization problem under the following weighted sum of discounted expected utilities: $\sum_{i}\dot{\alpha}_{i}E\left[\sum_{t}\dot{\beta}^{t}\frac{[c_{i,t}(s_{t})]^{1-\dot{\gamma}_{i}}}{1-\dot{\gamma}_{i}}\right],$ (17) where $\dot{\alpha}_{i}$ is Parate weight and $\dot{\gamma}_{i}$ is the coefficient of relative risk aversion for household $i$.
@@
 Mozzoco and Saini (2012) consider a more general case for testing the homogeneity of risk preferences.
-Their test requires a panel dataset with a very long time dimension that is not applicable to my THBS dataset.105105105The data used in their study covers the household-level monthly panel dataset surveyed in rural India between 1975 and 1985.
+Their test requires a panel dataset with a very long time dimension that is not applicable to my THBS dataset.103103103The data used in their study covers the household-level monthly panel dataset surveyed in rural India between 1975 and 1985.
 The households with fewer than 80 data points are not included in the homogeneity tests (Mozzoco and Saini 2012, p.
@@
 An alternative way to test potential heterogeneity in the risk preference is to use direct measures on the risk preferences of workers.
-Schulhofer-Wohl (2011) uses the results of experimental risk-tolerance questions on the workers’ risk preferences in the Health and Retirement Study to test whether the correlations between earnings and aggregate shocks differ across high- and low-risk-tolerance workers.106106106This question is asking if the respondents would be willing to take a risk to get a new job with a higher income, even if there was a 50-50 chance of a reduction in income by a third (Schulhofer-Wohl 2011, P.
+Schulhofer-Wohl (2011) uses the results of experimental risk-tolerance questions on the workers’ risk preferences in the Health and Retirement Study to test whether the correlations between earnings and aggregate shocks differ across high- and low-risk-tolerance workers.104104104This question is asking if the respondents would be willing to take a risk to get a new job with a higher income, even if there was a 50-50 chance of a reduction in income by a third (Schulhofer-Wohl 2011, P.
 931).
@@
 Secondly, I test the hypothesis that the heads who work in the smithing and the non-machinery factories have different risk preferences.
-As described in Online Appendix A.5, skilled workers in the smithing sector could be small business owners.107107107Workers in the canning factories might have had a different preference on the risk preferences (Online Appendix A.6).
+As described in Online Appendix A.5, skilled workers in the smithing sector could be small business owners.105105105Workers in the canning factories might have had a different preference on the risk preferences (Online Appendix A.6).
 However, my THBS sample does not include the workers in canning factories.
@@
 where $\tilde{y}_{i,t}$ is the head’s earning, $x_{i,t}$ is the family size control, $\nu_{i}$ indicates the household-fixed effect, $\zeta_{t}$ indicates the time-fixed effect, and $\epsilon_{i,t}$ is a random error term.
-I herein use the head’s earnings ($\tilde{y}_{i,t}$) as the idiosyncratic shock variable instead of the disposable income.108108108In the analysis, I consider the labor-supply adjustment variable and sales of miscellaneous assets as the additional dependent variables in addition to the net income variables $r_{i,t}$ (Section 4.2).
+I herein use the head’s earnings ($\tilde{y}_{i,t}$) as the idiosyncratic shock variable instead of the disposable income.106106106In the analysis, I consider the labor-supply adjustment variable and sales of miscellaneous assets as the additional dependent variables in addition to the net income variables $r_{i,t}$ (Section 4.2).
 The former variable is the total earnings of the family members except for the head, meaning that the simultaneous issue arises if I use the disposable income in this equation.
@@
 454–469.
+C.5 Testing Relationship between Volatility of Earnings and Use of Credit Purchases
+Table C.3: Volatility of Earnings and Use of Credit Purchases
+Dependent Variable: Indicator Variable for
+Non-CP Users Non-CP + Low-CP Users
+(1) (2) (3) (4) (5) (6)
+CV of Head’s Earnings $0.003$ $-0.002$ $-0.027$ $0.019$ $0.008$ $-0.005$
+$[0.012]$ $[0.014]$ $[0.050]$ $[0.013]$ $[0.017]$ $[0.049]$
+Sample mean of the DV $0.12$ $0.12$ $0.12$ $0.24$ $0.24$ $0.24$
+Model LPM LPM Probit LPM LPM Probit
+Number of THBS households 33 33 33 33 33 33
+Family Size Yes Yes Yes Yes Yes Yes
+Children aged 6–12 (%) Yes Yes Yes Yes Yes Yes
+Children aged 13–16 (%) Yes Yes Yes Yes Yes Yes
+Adults aged 17+ (%) Yes Yes Yes Yes Yes Yes
+Head’s age No Yes Yes No Yes Yes
+Head’s earnings No Yes Yes No Yes Yes
+Zero slope ($p$-value) 0.467 0.718 0.336 0.515 0.519 0.451
+***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. Robust standard errors are reported in brackets.
+Notes: The dependent variable in Columns (1)–(3) is the indicator variable for the households who did not use any credit purchases during the sample period (non-CP users).
+The dependent variable in Columns (4)–(6) is the indicator variable for the non-CP users and households less than the first quartile in the sample average of credit purchases.
+Columns (1)–(2) and (4)–(5) use the linear probability model (LPM), whereas Columns (3) and (6) use the Probit model.
+“CV of Head’s Earnings” indicates the coefficient of variations in the head’s semi-monthly earnings.
+The sample mean of the CV is $5.76$ (Std.
+Dev.
+= $5.72$).
+All the regressions include the family composition variables: average family size, share of children aged 6–12, share of children aged 13–16, and share of adults aged 17+ (share of children aged 0–5 is used as a reference group).
+The average head’s age and average head’s earnings are added to the control variables in Columns 3 and 6.
+$F$-statistics $p$-values for the null of the zero-slope hypothesis are reported in Columns (1)–(2) and (4)–(5).
+$\chi^{2}$-statistics $p$-values for the null of the zero-slope hypothesis are listed in Columns (3) and (6).
+The standard errors for the LPM are based on the HC2 estimator proposed by Horn et al.
+(1975).
+Column (1) of Table C.3 presents the result for Equation 4.107107107 The estimated coefficient is close to zero and statistically insignificant.
+A potential concern may be the correlations between the volatility of earnings and the head’s age and earnings level.
+Column (2) shows the result from the extended model, including the head’s age and earnings in the controls, which shows the robustness of the baseline estimate in Column (1).
+Another potential issue is the functional form assumption in the linear probability model.
+Column (3) then lists an alternative non-linear model, including the full set of control variables used in Column (2).
+It confirms that the estimate is not statistically significantly different from zero.
+Next, I changed the threshold for the dependent indicator variable to consider a wider range of households including the households with less than the first quartile of the average credit purchases.
+Columns (4)–(6) summarize the results in the same column layouts as Columns (1)–(3), confirming that the estimates are still close to zero in all the specifications.
+Finally, the last row in this table shows that the null of the joint zero slope hypothesis is strongly rejected in all the regressions.
+Overall, there is no apparent trend in the characteristics of households that use credit for their purchases.
+This suggests that retailers may not have considered the volatility of earnings when providing credit to factory workers.
+This is consistent with the retailers’ motivation to secure their local customers discussed earlier, and with the fact that factory workers commonly used credit for purchases.108108108Despite this, low-income households might have had more difficulty using credit for purchases.
+A famous essay drawn by a primary school student in the 1930s depicts an example of a rice retailer who refused to buy rice under credit purchase (Toyoda 1969, pp.
+346–355).
 References used in the Appendices
@@
   • (11)
+  • (12) Julie A Nelson. On testing for full insurance using consumer expenditure survey data. Journal of Political Economy 102, no.2 (1994): 384–394.
+  • (13)
   • (14) Kawai, Takao. “Shakai chōsa houhoushi ni tsuite (On history of social research methods)” [in Japanese]. Hōgaku Kenkyū 53, no. 9 (1980): 41–67.
@@
   • (21)
+  • (22) Martin Gervais and Paul Klein. Measuring consumption smoothing in CEX data. Journal of Monetary Economics 57(8) (2010): 988–999.
+  • (23)
   • (24) Mazzocco, Maurizio and Shiv Saini. “Testing Efficient Risk Sharing with Heterogeneous Risk Preferences.” The American Economic Review 102, no. 1 (2012): 428–468.
@@
   • (30) Takano, Iwasaburo. Honpō shakaitōkei ron (Social Statistics in Japan) [in Japanese]. Tokyo: Kaizō sha.
+  • (31)
+  • (32) Toyoda, Masako. Tsuzurikata kyōshitsu (Composition Class) [in Japanese]. Rironsha, 1969.
   • (33)
@@
   • (44) Shibuya, Ryuichi. Kourigashikinyū no tenkai kōzō (Development and Structure of Usuries) [in Japanese]. Tokyo: Nihon Tosho Center, 2000.
+  • (45)
+  • (46) Susan Dynarski, Jonathan Gruber, Robert A Moffitt, and Gary Burtless. Can families smooth variable earnings? Brookings Papers on Economic Activity 1997, no. 1 (1997): 229–303.
+  • (47)
+  • (48) Susan D Horn, Roger A Horn, and David B Duncan. Estimating Heteroscedastic Variances in Linear Models. Journal of the American Statistical Association 70, no. 350 (1975): 380–385.
   • (49)

### v12

@@
 These studies used cross-sectional variation in income to estimate the elasticities of the temporary income sources.
-Their detailed analyses have revealed the roles of financial institutions and labor supply adjustments in compensating for income loss in working-class households (Scott and Walker 2012).
+Their detailed analyses have revealed the roles of financial institutions and labor supply adjustments in compensating for income loss in working-class households (e.g., Scott and Walker 2012).
 [^2] They found that interwar British working-class households had access to various institutions such as clothing clubs, hire-purchase, contractual savings, and insurance.
@@
 Source: Created by the author using the official shapefile (Ministry of Land, Infrastructure, Transport and Tourism, database).
-Figure 3: Factories in Tsukishima
-Notes: Blocks in green and light green indicate the shipyards.
-Blocks in red and pink indicate the factories in the machinery sector.
-Blocks in blue and light blue indicate the factories in the other manufacturing sectors.
-Blocks in brown are the warehouses.
-Small circles colored in black and white show the workplaces in the machinery and other manufacturing sectors, respectively.
-Source: Department of Health, Ministry of the Interior 1923c, second map. The tone was adjusted by the author using Adobe Photoshop 24.7.0.
 The subject of this paper is Tsukishima, an industrial area located southeast of Kyobashi Ward in Tokyo (Figure 2).
@@
 The focus of this study is the consumption behavior of these worker households in the industrial sector.
+Figure 3: Factories in Tsukishima
+Notes: Blocks in green and light green indicate the shipyards.
+Blocks in red and pink indicate the factories in the machinery sector.
+Blocks in blue and light blue indicate the factories in the other manufacturing sectors.
+Blocks in brown are the warehouses.
+Small circles colored in black and white show the workplaces in the machinery and other manufacturing sectors, respectively.
+Source: Department of Health, Ministry of the Interior 1923c, second map. The tone was adjusted by the author using Adobe Photoshop 24.7.0.
 Table 1: Industrial and Family Structures in Tsukishima
@@
 Hence, the details of this credit institution are not yet clear.
-Table 9: Customers, Sales Methods, and Competitiveness in the Retailers in Tokyo City
-Selling Item Number of Share of sales Social Class of Sales Method (%) Number of
-stores for regular the Regular Customer (%) Cash Credit Monthly competitors
-reported customer (%) Upper Middle Other purchase installment reported
-White Rice 101 73.1 16 45.7 38.2 26.9 71.5 1.7 8.5
-Fish 59 64.6 6.8 57.6 35.6 43.9 56.1 0 5.0
-Greengrocer 66 45.4 9.4 40.4 50.2 65.8 34.2 0 6.2
-Firewood and charcoal 100 51.1 15.7 50.9 33.4 20.4 78.4 1.3 8.6
-Notes: This table summarizes the statistics on the sales in the $326$ retailers selling white rice, fish, fruit and vegetables, and firewood and charcoal in Tokyo surveyed in 1935.
-The number of stores reported the share of regular customers (column 2) is $90$, $56$, $61$, and $93$ in the white rice, fish, greengrocer, and fuel categories, respectively.
-The “share of sales for regular customers” is defined as the percentage of sales for regular customers in the annual sales.
-The definition of the social classes (“Upper”; “Middle”; “Other”) is not defined in the report.
-The share of sales method is based on the percentage of sales using each method in the annual sales.
-The number of competitors indicates the number of peers within a three-cho, approximately $327$ meters, radius.
-The department stores and retail markets are not included in this survey.
-Source: Tokyo City Chamber of Commerce and Industry 1937, pp. legend; 20–21; questionnaire.
-The information on the credit purchases is personal, as in the case of state-contingent loans in the rural credit market (Udry 1994) and thus, not well documented.
-Fortunately, however, I found that some official survey reports documented information on this credit institution.
-A survey report on retailers in 1930 (hereafter 1930 Survey Report) revealed that credit purchase was a popular payment method in transactions with consumers.
-[^69] This survey was conducted in February and March 1930 by Tokyo City and covered $3,892$ small and medium enterprises (Tokyo City Office 1932a, p.
-1).
-Those included retailers as well as wholesale and other small businesses.
-The number of respondents varied depending on the question items.
-The number of respondents in the statistical table on the payment methods was $2,654$, of which $1,197$ were retailers (Tokyo City Office 1932b, p.
-273).
-To the best of my knowledge, this is the most comprehensive survey of the retailers in Tokyo circa 1930.
-In fact, $99$%($155/157$) of grain retailers, $81$%($35/43$) of fish retailers, $78$%($64/82$) of greengrocers, and $94$%($45/48$) of fuel retailers used credit purchases for their selling method (Tokyo City Office 1932b, pp.
-260; 263).
-Thus, credit purchases in Tokyo were neither region- nor religion-specific customs.
-A survey of retailers in 1935 published by the Tokyo City Chamber of Commerce and Industry provides more detailed information.
-Table 9 summarizes the information on the retailers’ sales measured in this survey.
-[^70] This survey covered $939$ retailers, and its report was edited on December 31, 1935.
-I digitized the statistics for all the retailers selling the representative items (rice, fish, fruit and vegetables, and fuels).
-The survey instrument included detailed questions regarding sales and purchases (Tokyo City Chamber of Commerce and Industry 1937, pp.
-legend, questionnaire).
-Online Appendix Table C.1 decomposes the statistics listed in Table 9 by scale of businesses.
-I confirm that the main findings are materially similar.
-First, many annual sales were from their regular customers, most of whom were from the non-upper classes.
-For example, $73$% of annual sales of rice retailers were from their regular customers, and more than $80$% of them were from non-upper classes, including factory workers.
-Second, credit purchase was a popular sales method, especially among rice and fuel retailers—more than $70$% of the sales were conducted with credit in both shops.
-[^71] These figures are systematically lower than those from the 1930 Survey Report.
-Given that both surveys show a systematic tendency toward credit purchases, this may be due to the differences in the samples, the definition of the retailers in each category, and the definition of the credit purchase itself.
-Although cash was also used among greengrocers, $34$% of the sales were settled using credit.
-This suggests that credit purchase was a popular payment method and that regular customers settled a large part of it.
-[^72] The 1930 Survey reveals that the retailers also commonly used credit purchases to buy their source goods from wholesalers (Tokyo City Office 1932a, pp.
-154–157).
-This means that credit purchase was a familiar payment method for retailers as well as consumers.
-Although these statistics are from the mid-1930s, this confirms the possibility that the retailers were engaged in credit transactions with long-term relationships with their regular customers.
-[^73] A textbook for retailers published in 1924 explains the psychology behind this long-term relationship—credit purchase was provided for the convenience of the customer’s budget but also for the “vanity” (status) of being regular customers (Shimizu 1924, pp.
-115–116).
-The final column of this table further provides suggestive evidence for this.
-The survey asked retailers whether the sales area was competitive or not, and if yes, they were asked to report the number of competitors within a three-cho (approximately $330$ meters) radius.
-As shown, rice and fuel stores have two to three more competitors than the retailers selling fish, fruit, and vegetables.
-Given the relatively high utilization of credit in rice and fuel retailers, they may have responded to competition by allowing their customers to trade on credit.
+This may mainly be because the information on credit purchases is personal, as in the case of state-contingent loans in the rural credit market (Udry 1994) and thus, not well documented.
+Accordingly, I analyze the institution using both the THBS dataset and available official documents.
+5.1 Availability and Premiums: Evidence from the THBS Dataset
 (a) Daily panel dataset:
@@
 Source: Created by the author using the THBS sample.
-Next, I examined the credit purchases included in the THBS dataset.
+I begin this descriptive analysis by examining the credit purchases included in the THBS dataset.
 Most THBS households ($29/33$) used any credit purchases during their sample period, confirming that credit purchase was a common practice among factory workers.
@@
 It is, therefore, reasonable that the average payment is greater than the average purchase in most semi-month cells.
-Thus, this is not a systematic trend but may be due to the smaller available number of households in the later sample period (Online Appendix Figure B.4).
+Note again that the wider variances in the latter sample period are based on the smaller available number of households (Online Appendix Figure B.4).
 I then analyzed whether the volatility of earnings restricted credit purchases.
 Since credit of worker was unobservable, I used the volatility of earnings as a proxy of credit risk in the following empirical specification:
-$I(\overline{CP}_{i}=0)=\varkappa_{0}+\varkappa_{1}CV_{i}+\bar{\mathbf{x}}_{i}^{\prime}\mathbf{\chi}+v_{i},$ (4)
+$I(\overline{CP}_{i}=0)=\varkappa_{0}+\varkappa_{1}CV_{i}+\bar{\mathbf{x}}_{i}^{\prime}\bm{\chi}+v_{i},$ (4)
 where $I(\cdot)$ is an indicator function, $\overline{CP}$ is the (unit-)average credit purchases, $CV$ is the coefficient of variation for head’s earnings, $\bar{\mathbf{x}}$ is a vector of the (unit-)average family composition variables, and $v$ is a random error term.
@@
 They bought various goods in greengrocers, fish stores, and general stores between February 1 and February 14, 1919, and paid exactly the same amounts on the 15th.
-[^77] The repayment period in credit purchases among retailers and wholesalers (footnote 72) was often set at approximately one month (Tokyo City Office 1932, pp.
+[^72] The repayment period in credit purchases among retailers and wholesalers (footnote 76) was often set at approximately one month (Tokyo City Office 1932, pp.
 157-159).
@@
 However, this case suggests that the risk premiums may have been included in the retail prices but not in the form of interest.
+5.2 Local Environment and Retailers’ Incentives: Evidence from the Official Surveys
+Next, I assess circumstances among the retailers using available survey reports on the credit institution.
+A survey report on retailers in 1930 (hereafter 1930 Survey Report) revealed that credit purchase was a popular payment method in transactions with consumers.
+[^73] This survey was conducted in February and March 1930 by Tokyo City and covered $3,892$ small and medium enterprises (Tokyo City Office 1932a, p.
+1).
+Those included retailers as well as wholesale and other small businesses.
+The number of respondents varied depending on the question items.
+The number of respondents in the statistical table on the payment methods was $2,654$, of which $1,197$ were retailers (Tokyo City Office 1932b, p.
+273).
+To the best of my knowledge, this is the most comprehensive survey of the retailers in Tokyo circa 1930.
+In fact, $99$%($155/157$) of grain retailers, $81$%($35/43$) of fish retailers, $78$%($64/82$) of greengrocers, and $94$%($45/48$) of fuel retailers used credit purchases for their selling method (Tokyo City Office 1932b, pp.
+260; 263).
+Thus, credit purchases in Tokyo were neither region- nor religion-specific customs.
+Table 9: Customers, Sales Methods, and Competitiveness
+in the Retailers in Tokyo City
+Selling Item Number of Share of sales Social Class of Sales Method (%) Number of
+stores for regular the Regular Customer (%) Cash Credit Monthly competitors
+reported customer (%) Upper Middle Other purchase installment reported
+White Rice 101 73.1 16 45.7 38.2 26.9 71.5 1.7 8.5
+Fish 59 64.6 6.8 57.6 35.6 43.9 56.1 0 5.0
+Greengrocer 66 45.4 9.4 40.4 50.2 65.8 34.2 0 6.2
+Firewood and charcoal 100 51.1 15.7 50.9 33.4 20.4 78.4 1.3 8.6
+Notes: This table summarizes the statistics on the sales in the $326$ retailers selling white rice, fish, fruit and vegetables, and firewood and charcoal in Tokyo surveyed in 1935.
+The number of stores reported the share of regular customers (column 2) is $90$, $56$, $61$, and $93$ in the white rice, fish, greengrocer, and fuel categories, respectively.
+The “share of sales for regular customers” is defined as the percentage of sales for regular customers in the annual sales.
+The definition of the social classes (“Upper”; “Middle”; “Other”) is not defined in the report.
+The share of sales method is based on the percentage of sales using each method in the annual sales.
+The number of competitors indicates the number of peers within a three-cho, approximately $327$ meters, radius.
+The department stores and retail markets are not included in this survey.
+Source: Tokyo City Chamber of Commerce and Industry 1937, pp. legend; 20–21; questionnaire.
+A survey of retailers in 1935 published by the Tokyo City Chamber of Commerce and Industry provides more detailed information.
+Table 9 summarizes the information on the retailers’ sales measured in this survey.
+[^74] This survey covered $939$ retailers, and its report was edited on December 31, 1935.
+I digitized the statistics for all the retailers selling the representative items (rice, fish, fruit and vegetables, and fuels).
+The survey instrument included detailed questions regarding sales and purchases (Tokyo City Chamber of Commerce and Industry 1937, pp.
+legend, questionnaire).
+Online Appendix Table C.1 decomposes the statistics listed in Table 9 by scale of businesses.
+I confirm that the main findings are materially similar.
+First, many annual sales were from their regular customers, most of whom were from the non-upper classes.
+$73$% of annual sales of rice retailers were from their regular customers, and more than $80$% of them were from non-upper classes, including factory workers.
+Second, credit purchase was a popular sales method, especially among rice and fuel retailers—more than $70$% of the sales were conducted with credit in both shops.
+[^75] These figures are systematically lower than those from the 1930 Survey Report.
+Given that both surveys show a systematic tendency toward credit purchases, this may be due to the differences in the samples, the definition of the retailers in each category, and the definition of the credit purchase itself.
+Although cash was also used among greengrocers, $34$% of the sales were settled using credit.
+This suggests that credit purchase was a popular payment method and that regular customers settled a large part of it.
+[^76] The 1930 Survey reveals that the retailers also commonly used credit purchases to buy their source goods from wholesalers (Tokyo City Office 1932a, pp.
+154–157).
+This means that credit purchase was a familiar payment method for retailers as well as consumers.
+This confirms that the retailers were engaged in credit transactions with long-term relationships with their regular customers.
+[^77] A textbook for retailers published in 1924 explains the psychology behind this long-term relationship—credit purchase was provided for the convenience of the customer’s budget but also for the “vanity” (status) of being regular customers (Shimizu 1924, pp.
+115–116).
+The final column of this table further provides suggestive evidence for this.
+The survey asked retailers whether the sales area was competitive or not, and if yes, they were asked to report the number of competitors within a three-cho (approximately $330$ meters) radius.
+As shown, rice and fuel stores have two to three more competitors than the retailers selling fish, fruit, and vegetables.
+Given the relatively high utilization of credit in rice and fuel retailers, they may have responded to competition by allowing their customers to trade on credit.
+Online Appendix C.3 provides a set of regression-based evidence on the positive relationship between the use of credit purchases and the number of local competitors.
 Finally, I assessed the losses suffered by retailers for credit purchases.
-According to the 1930 Survey Report, almost all the retailers suffered losses for credit purchases.
-For example, $99$% of rice and $97$% of fuel retailers lost their sales.
-This means that credit purchase was inherently a loss-making sales method for the retailers.
-As systematic statistics are unavailable to clarify the percentage of total profits corresponding to these losses, it is impossible to assess how severe these losses were for the retailers.
-However, an important fact is that these losses were usually capped at 5% of total sales.
+According to the 1930 Survey Report, $99$% of rice and $97$% of fuel retailers lost their sales.
+[^78] Online Appendix C.4 summarizes the statistics on the losses in credit purchases available from the 1930 Survey Report in detail.
+This implies that credit purchase was inherently a loss-making sales method for the retailers.
+However, these losses were usually marginal and capped at 5% of the total sales.
 For example, the available statistics suggest that $87$% of rice and $65$% of fuel retailers suffered losses amounting to $5$% of sales.
-This supports the evidence that while the retailers could not avoid losses in credit sales, they might have set prices that anticipated predicted losses to some extent.
+While the retailers could not avoid losses in credit sales, they might have set prices that anticipated predicted losses to some extent.
 This seems to be a plausible sales strategy for retailers to manage adverse selection issues under imperfect consumer information.
-Online Appendix C.4 summarizes the statistics on the losses in credit purchases available from the 1930 Survey Report in detail.
 To summarize, the available documents show that retailers competed with other retailers in the same sales area.
@@
 I digitized the daily longitudinal budget survey on factory worker households conducted in Tsukishima in 1919.
-I carefully assessed the potential measurement error due to the miss-assignments of paydays and proposed a way to avoid the assignment errors.
+I assessed the potential measurement error due to the miss-assignments of paydays and proposed a way to avoid the assignment errors.
 The estimated elasticity for total consumption indicates that the factory worker households suffered certain consumption losses when they faced idiosyncratic income shocks.
 The estimated elasticity for total expenditure ($0.43$) is in a similar range to those in rural economies in developing countries.
-According to Ravallion and Chaudhuri (1997), the elasticity for the rural households in Indian villages between 1975 and 1984 was approximately $0.5$.
-As for the urban households, the estimated elasticity was $0.4$ for Bangkok between 1975 and 1990, which had a similar GDP per capita to 1920 Japan (Townsend 1995).
+The elasticity for the rural households in Indian villages between 1975 and 1984 was reported to be approximately $0.5$.
+As for the urban households, the estimated elasticity was $0.4$ for Bangkok between 1975 and 1990, which had a similar GDP per capita to 1920 Japan.
 Since my estimate is from a specific occupation, skilled factory workers, we need to be careful when making comparisons with these estimates based on urban samples which include a variety of occupations.
 Nevertheless, the factory workers in the historical metropolis of Japan circa WWI were vulnerable to idiosyncratic income shocks.
-A recent study found an income elasticity of $0.39$ for factory workers’ households in Osaka city circa 1920 (Ogasawara 2024).
+A recent study found an income elasticity of $0.39$ for factory workers’ households in Osaka city circa 1920.808080The estimates for India, Bangkok, and Osaka are obtained from Ravallion and Chaudhuri (1997), Townsend (1995), and Ogasawara (2024), respectively.
 My estimate from the adjusted monthly panels is slightly larger than this estimate.
@@
 Therefore, inter-city and cross-national comparisons of income elasticities and analysis of the institutional characteristics behind consumption are subjects for future work.
+This would add a new analytical perspective to the study of economic history.
 Despite the vulnerability, the results from mechanism analyses show that credit purchases were commonly used among workers and played a role in mitigating adverse idiosyncratic shocks in several consumption subcategories, particularly food.
@@
 33).
-He had the experience of conducting the first household survey in Japan in 1916868686This is called the “Survey of Household Income and Expenditure of 20 Artisans and Workers in Tokyo,” which was inspired by Charles Booth (Miyoshi 1980, p.
+He had the experience of conducting the first household survey in Japan in 1916888888This is called the “Survey of Household Income and Expenditure of 20 Artisans and Workers in Tokyo,” which was inspired by Charles Booth (Miyoshi 1980, p.
 33).
@@
 3).
-He was a member of the seventh subcommittee (Rural Sanitation Subcommittee) of the Ministry of Home Affairs’ Health and Sanitation Investigation Committee since 1916.878787The Health and Sanitation Investigation Committee of the Ministry of Home Affairs was established on June 27, 1916, under the second Shigenobu Okuma Cabinet by issuing Imperial Ordinance No.
+He was a member of the seventh subcommittee (Rural Sanitation Subcommittee) of the Ministry of Home Affairs’ Health and Sanitation Investigation Committee since 1916.898989The Health and Sanitation Investigation Committee of the Ministry of Home Affairs was established on June 27, 1916, under the second Shigenobu Okuma Cabinet by issuing Imperial Ordinance No.
 172.
@@
 I calculate the weighted average of the daily wage using the number of factory workers in each age bin (12–14 and 15+) and in each category as weights.
-Thus, I shift the functional polynomial function by this amount of difference to yield the proxy distribution for 1918.979797This operation, therefore, assumes that the wage-by-age distributions in 1918 and 1924 are reasonably similar.
+Thus, I shift the functional polynomial function by this amount of difference to yield the proxy distribution for 1918.999999This operation, therefore, assumes that the wage-by-age distributions in 1918 and 1924 are reasonably similar.
 In the LST (i.e., the statistic on which the predicted distribution is based), the weighted average for those aged $14$ and older was $3.25$ yen.
@@
 In other words, the shifted curve shall get closer to the actual (unobservable) wage distribution as age increases.
-The long dashed curve in Figure B.8a examines this relationship by fitting another fractional polynomial model using the 12–14-year old bin figure of the ASCT and the predicted wages for 25–65 from the shifted fractional polynomial.
-[^98] The fitted model is as follows: $\hat{y}^{\textit{Daily Wage}}=-1.19+0.14a^{1}-0.002a^{2},\quad R^{2}=0.98.$ The deviance is reported to be $-149$.
+The long dashed curve in Figure B.8a examines this relationship by fitting another fractional polynomial model using the 12–14-year old bin figure of the ASCT and the predicted wages for 25–65 from the shifted fractional polynomial.100100100The fitted model is as follows: $\hat{y}^{\textit{Daily Wage}}=-1.19+0.14a^{1}-0.002a^{2},\quad R^{2}=0.98.$ The deviance is reported to be $-149$.
 I used the 25–65 age range for the prediction based on the age bins used in the LST.
@@
 where $w^{\textit{Regular}}>0$ is the monthly regular wage and $w^{\textit{Ancillary}}\geq 0$ indicates the monthly ancillary wage.
-Online Appendix B.6.2 shows that $w^{\textit{Ancillary}}$ can be calculated as the product of the monthly regular wage and weight for the ancillary wage, $\vartheta\in(0,1)$.
-[^99] Strictly, this weight may not be required to be in this range.
+Online Appendix B.6.2 shows that $w^{\textit{Ancillary}}$ can be calculated as the product of the monthly regular wage and weight for the ancillary wage, $\vartheta\in(0,1)$.101101101Strictly, this weight may not be required to be in this range.
 I use this range because all the cases reported in Kitazawa (1924) take the weights between zero and one.
@@
 Source: Created by the author using the THBS sample.
-Figure B.11 illustrates the distribution of the idiosyncratic shocks predicted using the semi-monthly panel dataset.101101101I use the semi-monthly panel dataset to overview the entire distribution of the idiosyncratic shocks rather than the daily panel datasets.
+Figure B.11 illustrates the distribution of the idiosyncratic shocks predicted using the semi-monthly panel dataset.103103103I use the semi-monthly panel dataset to overview the entire distribution of the idiosyncratic shocks rather than the daily panel datasets.
 This is because the daily panels include a set of consecutive household-date cells with no income.
@@
 This test is valid under a particular utility function.
-Specifically, the bias could be raised in the case that the coefficient of relative risk aversion interacts with the aggregate shock in the first-order condition for consumption because the interaction means that the aggregate shock cannot be removed using the time-fixed effect in the risk-sharing regressions.102102102Specifically, Schulhofer-Wohl (2011, pp.
+Specifically, the bias could be raised in the case that the coefficient of relative risk aversion interacts with the aggregate shock in the first-order condition for consumption because the interaction means that the aggregate shock cannot be removed using the time-fixed effect in the risk-sharing regressions.104104104Specifically, Schulhofer-Wohl (2011, pp.
 928–929) considers the utility maximization problem under the following weighted sum of discounted expected utilities: $\sum_{i}\dot{\alpha}_{i}E\left[\sum_{t}\dot{\beta}^{t}\frac{[c_{i,t}(s_{t})]^{1-\dot{\gamma}_{i}}}{1-\dot{\gamma}_{i}}\right],$ (17) where $\dot{\alpha}_{i}$ is Parate weight and $\dot{\gamma}_{i}$ is the coefficient of relative risk aversion for household $i$.
@@
 Mozzoco and Saini (2012) consider a more general case for testing the homogeneity of risk preferences.
-Their test requires a panel dataset with a very long time dimension that is not applicable to my THBS dataset.103103103The data used in their study covers the household-level monthly panel dataset surveyed in rural India between 1975 and 1985.
+Their test requires a panel dataset with a very long time dimension that is not applicable to my THBS dataset.105105105The data used in their study covers the household-level monthly panel dataset surveyed in rural India between 1975 and 1985.
 The households with fewer than 80 data points are not included in the homogeneity tests (Mozzoco and Saini 2012, p.
@@
 An alternative way to test potential heterogeneity in the risk preference is to use direct measures on the risk preferences of workers.
-Schulhofer-Wohl (2011) uses the results of experimental risk-tolerance questions on the workers’ risk preferences in the Health and Retirement Study to test whether the correlations between earnings and aggregate shocks differ across high- and low-risk-tolerance workers.104104104This question is asking if the respondents would be willing to take a risk to get a new job with a higher income, even if there was a 50-50 chance of a reduction in income by a third (Schulhofer-Wohl 2011, P.
+Schulhofer-Wohl (2011) uses the results of experimental risk-tolerance questions on the workers’ risk preferences in the Health and Retirement Study to test whether the correlations between earnings and aggregate shocks differ across high- and low-risk-tolerance workers.106106106This question is asking if the respondents would be willing to take a risk to get a new job with a higher income, even if there was a 50-50 chance of a reduction in income by a third (Schulhofer-Wohl 2011, P.
 931).
@@
 Secondly, I test the hypothesis that the heads who work in the smithing and the non-machinery factories have different risk preferences.
-As described in Online Appendix A.5, skilled workers in the smithing sector could be small business owners.105105105Workers in the canning factories might have had a different preference on the risk preferences (Online Appendix A.6).
+As described in Online Appendix A.5, skilled workers in the smithing sector could be small business owners.107107107Workers in the canning factories might have had a different preference on the risk preferences (Online Appendix A.6).
 However, my THBS sample does not include the workers in canning factories.
@@
 where $\tilde{y}_{i,t}$ is the head’s earning, $x_{i,t}$ is the family size control, $\nu_{i}$ indicates the household-fixed effect, $\zeta_{t}$ indicates the time-fixed effect, and $\epsilon_{i,t}$ is a random error term.
-I herein use the head’s earnings ($\tilde{y}_{i,t}$) as the idiosyncratic shock variable instead of the disposable income.106106106In the analysis, I consider the labor-supply adjustment variable and sales of miscellaneous assets as the additional dependent variables in addition to the net income variables $r_{i,t}$ (Section 4.2).
+I herein use the head’s earnings ($\tilde{y}_{i,t}$) as the idiosyncratic shock variable instead of the disposable income.108108108In the analysis, I consider the labor-supply adjustment variable and sales of miscellaneous assets as the additional dependent variables in addition to the net income variables $r_{i,t}$ (Section 4.2).
 The former variable is the total earnings of the family members except for the head, meaning that the simultaneous issue arises if I use the disposable income in this equation.
@@
 Source: Tokyo City Chamber of Commerce and Industry 1937, pp. legend; 20–21; questionnaire.
+In Table C.2, I analyze the relationship between credit purchases and competitiveness among local retailers using the selling item by scale of business cells reported in Table C.1.
+In Column (1), I regressed the sales in credit purchases (in percentage points) on the number of local competitors.
+I then conditioned the business scales in Column (2).
+The $F$-statistic $p$-value for joint zero-slope hypotheses for these scale dummies is estimated to be $0.990$.
+In Column (3), I included the selling item dummies instead of the scale dummies.
+The $F$-statistic $p$-value for joint zero-slope hypotheses is rejected at the conventional level, and the measure of fit is relatively high compared with the other two models reported in Columns (2) and (3).
+The estimate is $3.420$, suggesting that a one unit increase in the local competitors increases the use of credit purchase by roughly $3.4$ percentage points.
+Given that the standard deviation of the local competitors is approximately $2.3$, the estimated magnitude corresponding to the one standard deviation increase would be $7.9$ ($3.420\times 2.3$).
+This is roughly $41$% of the standard deviation of the dependent variable ($19.2$) and thus, is an economically meaningful magnitude.
+Table C.2: Use of Credit Purchases and Competitiveness among Local Retailers
+Dependent Variable: Sales in Credit Purchases (%)
+(1) (2) (3)
+Number of Competitors 7.1907.190***7.190 ∗ ∗ ∗ 7.2547.254***7.254 ∗ ∗ ∗ 3.4203.420**3.420 ∗ ∗
+$[2.041]$ $[2.145]$ $[1.310]$
+Sample mean of the DV $59.4$ $59.4$ $59.4$
+Model OLSE OLSE OLSE
+Number of observations 12 12 12
+Number of retail shops (weight) 326 326 326
+Scale FEs No Yes No
+Selling item FEs No No Yes
+$F$-statistic $p$-value for the FEs 0.990 0.000
+R-squared 0.577 0.578 0.986
+***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. Robust standard errors are reported in brackets.
+Notes: The observations used in this table are from the selling item by scale of business cells reported in Table C.1.
+The dependent variable in Columns (1)–(3) is the percentage of sales under credit purchases.
+“Number of competitors” indicates the number of peers within a three-cho (approximately 327 meters) radius.
+The sample mean and standard deviation of the dependent variable are $59.4$ and $19.2$, respectively.
+The sample mean of the number of competitors is $6.89$ (Std.
+Dev.
+= $2.26$).
+All the regressions are weighted by the number of stores in each selling item by scale of business cell (see third column of Table C.1).
+The scale of business dummies are added in Column 2, whereas the selling item dummies are included in Column 3.
+$F$-statistics $p$-values for these dummies are reported in Columns 2 and 3.
+The standard errors are based on the HC2 estimator proposed by Horn et al.
+(1975).
 C.4 Retail Sales in Tokyo City: A Small and Medium Scale Enterprises Survey of 1930
@@
 This suggests that retailers may not have considered the volatility of earnings when providing credit to factory workers.
-This is consistent with the retailers’ motivation to secure their local customers discussed earlier, and with the fact that factory workers commonly used credit for purchases.108108108Despite this, low-income households might have had more difficulty using credit for purchases.
+This is consistent with the retailers’ motivation to secure their local customers discussed earlier, and with the fact that factory workers commonly used credit for purchases.110110110Despite this, low-income households might have had more difficulty using credit for purchases.
 A famous essay drawn by a primary school student in the 1930s depicts an example of a rice retailer who refused to buy rice under credit purchase (Toyoda 1969, pp.

### v14

@@
-Consumption Smoothing in Metropolis: Evidence from Working-class Households in Prewar Tokyo
+Risk-coping Behaviors in Metropolis: Evidence from Working-class Households in Prewar Tokyo
 Abstract
-I analyzed the risk-coping strategies among factory worker households in early 20th-century Tokyo.
-I digitized and analyzed a unique daily longitudinal household budget survey to determine how consumption was impacted by idiosyncratic shocks.
-I found that while the households were so vulnerable that the shocks impacted their consumption levels, the estimated income elasticity for food consumption was relatively low in the short run.
-The analysis using adverse health shock confirms the robustness of the results.
-The result of mechanism analysis suggests that credit purchases with local retailers smoothed short-run food consumption.
-Despite the potential loss of profit, retailers in a competitive situation allowed consumers to trade on credit.
-This shows the roles of informal credit institutions in mitigating vulnerability among urban worker households.
+I analyze the risk-coping strategies of factory-worker households in early twentieth-century Tokyo.
+I digitized a unique daily longitudinal household budget survey conducted in Tsukishima, a representative manufacturing area, to examine how consumption was affected by idiosyncratic shocks.
+I find that although the households were vulnerable and their consumption levels were impacted by these shocks, the estimated income elasticity of indispensable consumption was relatively low in the short run.
+The results of the mechanism analysis suggest that credit purchases from local retailers helped smooth short-run consumption, highlighting the role of informal credit institutions in mitigating vulnerability among urban worker households.
 Keywords: Consumption smoothing; credit institution; urban local community; measurement error; risk-coping strategy; risk-sharing
@@
 1 Introduction
-Economic studies that analyze the relationship between historical economic development and financial institutions have focused on the alleviation of constraints on corporate financing through official financial systems (Acemoglu et al.
-2005; Levine 2005).
-However, systematic research on the association between household financial institutions and consumption smoothing during industrialization remains limited.
-How did working-class households respond to idiosyncratic shocks and smooth their consumption before public insurance was established?
-Development economics research has confirmed that consumption smoothing is an essential mechanism for mitigating poverty and explaining human capital formation.
+During the early stages of industrialization, workers who migrated to large cities in search of employment faced various uncertainties.
+Lacking the communal ties found in rural areas, they were forced to manage on their own.
+The vulnerability to idiosyncratic income shocks makes it challenging to smooth consumption and increases the risk of falling into poverty.
+Therefore, understanding how these workers managed their livelihoods helps clarify the mechanisms behind improvements in living standards in industrializing cities.
+This study examines risk-coping behaviors among working-class households in Tsukishima, a representative manufacturing area in early twentieth-century Tokyo City.
+Against the backdrop of industrialization during World War I (WWI), many workers migrated to the city, forming a notable portion of the population.
+Despite the absence of public insurance systems, evidence shows that these workers were able to form and maintain families during this period (Nakagawa 1985).
+This historical context offers a rare empirical opportunity to understand the risk-coping strategies of urban workers during industrialization.
+To provide micro-level evidence on risk-coping behaviors, I digitize a unique daily longitudinal budget survey of factory-worker households.
+Using a systematic consumption-smoothing regression, I find that the full insurance hypothesis is strongly rejected for overall consumption expenditure.
+However, I also find evidence that expenses on indispensable consumption categories are partially smoothed in the short run: the estimated income elasticity for the food category is less than half of that for overall consumption.
+I then use several temporary income categories within the same regression framework to assess the risk-coping behaviors.
+Considering both favorable and adverse income shocks, I find that savings, credit purchases, and gifts responded to these shocks.
+To focus specifically on responses to adverse income shocks, I use the household head’s absence due to the influenza pandemic as an instrument.
+Results from the instrumental variable approach show that credit purchases from local retailers played a significant role in mitigating adverse idiosyncratic shocks, whereas other risk-mitigation strategies were less practical.
+Given these results, I analyze the institutional aspects of credit purchases descriptively.
+First, I find that credit purchases were common among workers and that volatility in household heads’ earnings did not correlate with average credit purchases.
+Detailed cases suggest that risk premiums were not charged as explicit interest but were instead embedded in retail prices owing to imperfect customer information.
+Second, I examine conditions among retailers using several official survey reports.
+These reports provide suggestive evidence that retailers competed intensely within the same sales areas.
+Because credit purchases did not require explicit contracts, they often led to losses from unpaid sales.
+However, to retain regular customers and build loyalty, local retailers had strong incentives to offer credit transactions.
+The primary contribution of this study is the revelation of the role of informal credit institutions in a historical metropolis.
+By considering both formal and informal financial institutions simultaneously, this study documents how informal credit transactions with local retailers helped mitigate idiosyncratic income losses.
+This finding adds to the literature on consumption-smoothing behavior across various economies, [^1] The early representative literature includes Townsend (1994), Townsend (1995), and Ravallion and Chaudhuri (1997).
+Morduch (1999), Townsend (1995), and Dercon (2004) review earlier studies.
+Attanasio and Pistaferri (2016) and Meyer and Sullivan (2023) provide recent discussions.
+as well as the studies investigating the mechanisms behind human capital formation through consumption smoothing.
 [^2] Evidence shows that markets in contemporary developing countries are often imperfect.
-Thus, the households are vulnerable to risk, and idiosyncratic shocks impact the levels of their consumption.
-The early representative literature includes, for example, Townsend (1994), Townsend (1995), and Ravallion and Chaudhuri (1997).
-Morduch (1999), Townsend (1995), and Dercon (2004) provide reviews of the earlier studies.
-Attanasio and Pistaferri (2016) and Meyer and Sullivan (2023) provide recent discussions.
-Since shocks are not fully insured in the markets, smoothness of consumption is also an important factor in human capital formation.
-Foster (1995) investigates the shock of the 1988 floods in Bangladesh and shows that children’s growth patterns are influenced by the inefficiency of the credit markets.
-Rose (1999) investigates the relationship between consumption smoothing and excess losses of girls, revealing that the favorable rainfall events in early life increased the survival probabilities of girls in rural India.
-Gertler and Gruber (2002) also shows that consumption smoothing against health shocks is imperfect in Indonesia.
-Surprisingly, however, economic history studies have not yet provided an answer to this important question.
-We still know little about how historical households maintained their livelihoods during industrialization.
-This study assesses risk-coping behavior among the working-class households in early twentieth-century Tokyo.
-I digitized and analyzed a unique daily longitudinal budget survey on factory worker households conducted in a representative manufacturing area.
-The full insurance hypothesis is strongly rejected for overall consumption expenditure, indicating that factory worker households in the historical metropolis suffered certain consumption losses when they faced idiosyncratic adverse income shocks.
-However, expenses on indispensable consumption categories were partially smoothed in the short run.
-Credit purchases in local retailers played an important role in mitigating adverse idiosyncratic shocks, whereas the other risk-mitigation strategies were impractical.
-Credit purchases did not require an explicit contract and thus often led to a loss in sales for the retailers.
-However, to remain competitive, retailers allowed credit transactions so that their customers remained regular.
-This shows that informal credit institutions in urban local communities functioned as informal insurance networks in rural economies.
-This study’s primary contribution is to add analytical perspectives on consumption smoothing behavior to economic history studies.
-Studies of how working-class households compensated for income loss during the late 19th to the early 20th century have focused on Britain.
-These studies used cross-sectional variation in income to estimate the elasticities of the temporary income sources.
-Their detailed analyses have revealed the roles of financial institutions and labor supply adjustments in compensating for income loss in working-class households (e.g., Scott and Walker 2012).
-[^2] They found that interwar British working-class households had access to various institutions such as clothing clubs, hire-purchase, contractual savings, and insurance.
-The households with less earnings were likely to use credit-based institutions like clubs and hire-purchase, whereas those with higher incomes could use savings institutions.
-Horrell and Oxley (2000) highlight the important role of labor supply adjustments among family members to manage the income losses of the head in working-class households in late 19th-century Britain.
-Kiesling (1996) and Boyer (1997) focused on the risk-mitigation strategies among the cotton workers households in Lancashire during the recession caused by the Civil War.
-However, there is little evidence of how those risk-coping strategies smoothed their consumption.
-[^3] Other strands of literature provide supply-side (lender-side) evidence on the functions of the micro-financial institutions for the workers (O’Connell and Reid 2005), building image on the historical consumers in the long-run (Francks 2009), and describe the relationship between consumption and production (Gordon 2013).
-Hence, this study provides a rare window into urban household financial behavior during industrialization by utilizing unique household longitudinal budget data from prewar Japan.
-It also implements a systematic panel data analysis to offer comparable estimates of the income elasticity for future studies.
-The second contribution is revealing the roles of informal credit institutions in historical metropolis.
-Evidence from development economics studies indicates that the availability of informal credit institutions relates to the stability of consumption in rural economies, especially when the availability of formal financial institutions is limited.
-Udry (1994) showed that state-contingent loans among relatives and neighbors could partially insure against idiosyncratic shocks on rural households in northern Nigeria.
-A detailed field survey revealed that asymmetric information was no longer important in those private credit contracts (Udry 1990).
-[^4] Fafchamps and Lund (2003) also presented evidence that the households in the rural Philippines used informal networks with relatives and friends as a risk-mitigation strategy for idiosyncratic shocks.
-Related studies consider risk-sharing among households with similar cultural and institutional backgrounds.
-Grimard (1997) showed that insurance among households in the same ethnic group could partially mitigate idiosyncratic shocks in the multiethnic economy of Côte d’Ivoire.
-Munshi and Rosenzweig (2016) presented the importance of rural insurance networks among workers in the same caste in the context of spatial wage inequality in India.
-Nonetheless, there is little evidence on the function of informal credit institutions in the historical context.
-Ogasawara (2024) observed the role of official financial institutions, such as postal savings and pawnshops, in compensating temporary income losses in the early 20th-century Osaka.
-[^5] Note that Osaka was middle scale city of approximately half the size of Tokyo City circa 1920 (Online Appendix Table A.1).
-However, the function of informal credit institutions in consumption smoothing has not been considered due to limited data availability.
-By considering the roles of both formal and informal financial institutions simultaneously, this current study documents the contribution of informal credit transactions by local retailers.
-Relatedly, an examination of how, and the extent to which, the local society could ensure household consumption provides valuable insights into existing urban history literature.
-After World War I, working-class households in Tokyo could establish and sustain their families (Nakagawa 1985).
-The results of this study suggest that the expansion of urbanization and credit transactions through the war underpins this traditional historical view.
-Another suggestion relates to the findings by Munshi and Rosenzweig (2016), who linked spatial wage inequality and men’s mobility in India with the availability of risk-mitigation networks among workers in the same caste.
-From this perspective, local credit institutions with lower barriers may motivate rural workers to urbanize, which may partially explain why Tokyo grew so fast circa WWI.
-This hypothesis may be worthy of testing in future research.
-The third contribution is to offer a case study of a metropolis in a developing economy.
-Extant development economics studies on consumption smoothing behavior have focused particularly on the rural economies (e.g., Townsend 1994; Ravallion and Chaudhuri 1997; Gertler and Gruber 2002).
-[^6] Recent field experimental studies also focus on village economies (Comola and Prina 2023; Macours et al.
+Consequently, households remain vulnerable to risk, and idiosyncratic shocks affect their consumption levels, influencing human capital formation (Foster 1995; Rose 1999; Gertler and Gruber 2002).
+Specifically, I present a case study of an urban setting characterized by weak communal ties, which has often been overlooked in the literature.
+For example, development economics research shows that informal credit institutions based on relatives and neighbors help stabilize consumption in rural economies, especially where formal financial institutions are scarce (Udry 1994, 1990; Fafchamps and Lund 2003).
+[^3] Related studies examine risk-sharing among households with similar cultural and institutional backgrounds (Grimard 1997; Munshi and Rosenzweig 2016).
+Recent field experiments also focus on village economies (Comola and Prina 2023; Macours et al.
 2022).
-While Townsend (1995), Skoufias and Quisumbing (2005), and Aguila et al.
-(2017) provided case studies on middle-sized contemporary economies, there is little evidence on agglomeration economies like prewar Tokyo.
-Urbanization is, however, now a worldwide phenomenon.
-The proportions of urban population in many developing countries, especially in East Asia, are reaching a similar level to the developed countries (Henderson and Turner 2020).
-This requires a deeper understanding of consumption behavior among urban working households in the economic development process.
-Moreover, there is little research on the role of financial institutions among the household sector in agglomeration economies.
-This study considers a set of homogeneous households in a representative manufacturing area in prewar Tokyo, which provides a valuable empirical setting for testing risk-sharing strategies in the historical metropolis.
-[^7] Focusing on a sample that shares similar occupational and regional characteristics systematically mitigates bias in estimating income elasticity under the standard risk-sharing regressions due to the heterogeneity in the risk preferences (Schulhofer-Wohl 2011; Mazzocco and Saini 2012).
-Caution is needed when comparing results, as the modern financial system differs in many ways from the past.
-However, my findings show that informal credit institutions in the local urban community could mitigate vulnerability among households similar to rural developing economies (Udry 1994; Fafchamps and Lund 2003).
+Townsend (1995), Skoufias and Quisumbing (2005), Aguila et al.
+(2017), and Ogasawara (2024) provide case studies on middle-sized contemporary and historical economies, but evidence on agglomeration economies such as Tokyo City remains scarce.
+Another important strand of the literature highlights the role of social insurance provided by religious communities in rural contexts (Ager and Ciccone 2018; Ferrara and Testa 2023; Chen 2010).
+My findings suggest that, in the case of a metropolis, workers primarily relied on informal credit institutions arising from local commercial practices, as social institutions based on communal ties and religious organizations were largely absent.
+This result may also offer empirical lessons for rapidly urbanizing cities in developing countries.
+[^4] The real GDP per capita in 2011 dollars was $3,217$ in 1919 Japan (Bolt and van Zanden 2025).
+This figure is close to the $3,537$ reported for South and Southeast Asian countries in 2010, where rapid urbanization has been observed (Henderson and Turner 2020).
+The second contribution is the incorporation of analytical perspectives on consumption-smoothing behavior into economic history studies.
+I examine a set of homogeneous households in a representative manufacturing area, which provides a valuable empirical setting for testing consumption smoothing.
+[^5] Focusing on a sample that shares similar occupational and regional characteristics systematically reduces bias when estimating income elasticity (Schulhofer-Wohl 2011; Mazzocco and Saini 2012).
+This study advances economic history research that documents compensatory responses to income loss among working-class households but lacks direct evidence on consumption smoothing (Scott and Walker 2012; Horrell and Oxley 2000; Kiesling 1996; Boyer 1997).
+I implement a systematic panel data analysis to provide comparable estimates of income elasticity.
+This comparability improves understanding of the availability of risk-coping institutions at different stages of economic development and strengthens the potential for cross-period and cross-economy comparisons.
+It also adds an analytical perspective to other strands of literature on the supply-side evidence of financial institutions (O’Connell and Reid 2005) and historical patterns of consumption and production (Francks 2009; Gordon 2013).
+Lastly, examining how, and to what extent, local society could secure household consumption offers a new perspective within the urban history literature.
+A representative historical study on urban households shows that workers in Tokyo were able to establish and sustain their families around WWI (Nakagawa 1985).
+The findings of this study suggest that the expansion of urbanization and credit transactions supports this traditional view.
+Another implication relates to the findings of Munshi and Rosenzweig (2016), who link spatial wage inequality and men’s mobility in India to the availability of risk-mitigation networks among workers within the same caste.
+From this perspective, the existence of local credit institutions with lower entry barriers may have encouraged rural workers to migrate to urban areas, helping to explain Tokyo’s rapid growth around WWI.
 2 Historical Background
 2.1 Japanese Economy circa the First World War
-Between 1885 and 1915, the Japanese real GNP grew annually by 2.6%, whereas the population increased by 1.1%.
-[^8] Faster economic than population growth implies that modern economic growth had started in this period as defined by Simon Kuznets (Miyamoto 2008, pp.
+Between 1885 and 1915, Japan’s real gross national product (GNP) grew at an annual rate of 2.6%, while the population increased by 1.1%.
+[^6] Faster economic growth relative to population growth indicates that modern economic growth had begun during this period, as defined by Kuznets (Miyamoto 2008, pp.
 60; 64).
-In 1915, the percentage of the employed population in the primary, secondary, and tertiary industries were 63%, 20%, and 18%, respectively.
-Based on domestic genuine production, these industries accounted for just under 30%, over 30%, and approximately 40%, respectively.
-Soon after the outbreak of WWI, exports from European countries declined.
-Japan increased exports to the Asian market on behalf of those European countries and exported munitions and foodstuffs to the European Allies, leading to a current account surplus.
-The import substitution expanded the domestic market and led to a rise in the machinery, metal, and chemical sectors.
-A worldwide shortage of ships also stimulated the development of shipbuilding, which proliferated.
-The development of the shipbuilding industry further stimulated demand for the machinery and steel industries.
-Through this process of heavy and chemical industrialization, the total number of enterprises increased.
-[^9] Descriptions about the process of heavy chemical industrialization are based on Nakamura (1971, p.
-128), Nakamura and Odaka (1989, p.
-10), and Sawai and Tanimoto (2016, p.
-250).
-Year-on-year growth in the number of incorporated firms was 17.8%, 22.3%, and 13.5% in 1918, 1919, and 1920, respectively.
-These are the top three of the 34 years from after the Russo-Japanese War to 1940 (Nakamura and Odaka 1989, p.
-22).
-Figure 1 shows that the GNP growth from 1915–19 was 11.5% in nominal terms (4.4% in real terms), the highest growth throughout the 1910s and 1920s.
-This rapid industrialization impacted the Japanese economy in many ways.
-Due to export-led economic growth, large inflows of gold increased domestic currency balances, causing prices to soar.
-Although wages began to rise circa 1916 as a response to rising prices and labor demand due to the increased number of firms, wage increases failed to keep pace with soaring prices.
-By the end of the war, the number of labor disputes grew sharply, and the labor movement flourished.
-Furthermore, the rise of business accelerated urbanization as the rural population moved to cities in search of job opportunities.
-The share of large cities (with a population of $50,000$ or more) as a percentage of the total population was less than $10$% in 1889 but increased to nearly $16$% in 1920 (Online Appendix Figure A.1a).
-The growth rate was exceptionally high during WWI, indicating that structural changes through the war accelerated the rise of the urban population.
-Urbanization has increased public investment in large cities, especially in construction investment in social capital such as roads and water supply systems.
-[^10] Descriptions of the export-led growths, urbanization, and investments are from Nakamura and Odaka (1989, pp.
-21–23) and Nakamura (1971, pp.
-142; 147).
-Exports in the 1920s were lower because of the appreciation of the yen against the real exchange rate due to the accumulation of domestic and foreign currencies during WWI.
-As the current account deficit was settled through the disbursement of foreign currencies owned by the government and the Bank of Japan, there was no reduction in the domestic money supply despite the sizeable current account deficit.
-This resulted in the domestic prices (especially prices of non-tradable goods) remaining high.
-The first half of the 1920s saw a temporary decline in industrial investment as the capital investment boom of the WWI period came to a halt under declining demand.
-[^11] Conversely, construction investment by the private sector, including electric power and private railways, increased (Nakamura 1971, p.
-145).
-Descriptions about the post-war economy in this paragraph are from Sawai and Tanimoto (2016, pp.
-253–254).
+In 1915, the employed population was distributed across the primary, secondary, and tertiary sectors at 63%, 20%, and 18%, respectively.
+Based on domestic genuine production, these sectors accounted for just under 30%, over 30%, and approximately 40%, respectively.
+Descriptions in this subsection are based on Nakamura (1971, pp.
+128; 142; 145; 147), Nakamura and Odaka (1989, pp.
+10; 21–23), and Sawai and Tanimoto (2016, pp.
+250; 253–254).
+Soon after the outbreak of WWI, European exports declined.
+Japan increased its exports to Asian markets in place of European suppliers and exported munitions and foodstuffs to European allies, resulting in a current account surplus.
+Import substitution further expanded the domestic market, driving growth in the machinery, metal, and chemical sectors.
+A global shortage of ships also spurred rapid development in shipbuilding, which expanded significantly during this period.
+The growth of the shipbuilding industry, in turn, stimulated demand in the machinery and steel industries.
+Figure 1 shows that Japan’s nominal GNP grew by 11.511.5% from 1915 to 1919 (4.44.4% in real terms), marking the highest growth rate observed throughout the 1910s and 1920s.
 Figure 1: Gross National Product per Capita and Personal Consumption Expenditure (in yen)
-Note: This figure shows the gross national product per capita (solid line) and personal consumption expenditure (dashed line) between 1910 and 1926.
-The prices are deflated using the CPI in 1934–1935 price.
-Source: Created by the author using Ohkawa et. al. (1974, p. 237).
-As described, there was a short-term business cycle characterized by the boom caused by WWI and the subsequent recession.
-However, as a whole, Japan achieved a high economic growth rate relative to other countries in the early 20th century.
-As shown in Figure 1, the growth rates of personal consumption expenditure were particularly high during and immediately after WWI, with average growth rates reaching 112% in 1915–19 and 118% in 1920–24.
-This was a high level by international standards.
+Note: This figure shows Japan’s GNP per capita (solid line) and personal consumption expenditure (dashed line) from 1910 to 1926.
+Prices are deflated using the 1934–1935 consumer price index (CPI).
+Source: Created by the author based on Ohkawa et al.
+(1974, p.
+237).
+The rise of business accelerated urbanization as the rural population moved to cities in search of employment.
+The share of large cities (with a population of $50,000$ or more) in the total population was less than 1010% in 1889 but rose to nearly 1616% in 1920 (Online Appendix Figure A.1a).
+Growth was particularly rapid during WWI, indicating that structural changes during the war accelerated urban population growth.
+Urbanization also increased public investment in large cities, especially in social infrastructure such as roads and water supply systems.
+Exports in the 1920s were lower owing to the appreciation of the yen in real terms, caused by the accumulation of domestic and foreign currencies during WWI.
+The current account deficit was settled through the disbursement of foreign currencies held by the government and the Bank of Japan, so the domestic money supply remained unchanged despite the sizeable deficit.
+As a result, domestic prices stayed high.
+The first half of the 1920s saw a temporary decline in industrial investment as the WWI-era capital investment boom ended under falling demand.
+Overall, while a short-term business cycle emerged–marked by the WWI-induced boom and subsequent recession–Japan achieved a high economic growth rate relative to other countries in the early twentieth century.
+The average growth rates of personal consumption expenditure reached 112% in 1915–1919 and 118% in 1920–1924 (Figure 1), reflecting a high level by international standards.
 [^7] Total consumption as a percentage of the GNP remained above 80%.
-Although this level is not apparent among other countries (Kuznets 1962), the growth rate of real personal consumption per capita was relatively high: from 1875 to 1939, the average growth rate of personal consumption was 1.36%, and the rates never became negative except during the Russo-Japanese War period.
-Descriptions on the personal consumption expenditure are from Nakamura (1971, pp.
+Although this level is not internationally high-standard (Kuznets 1962), the growth rate of real personal consumption per capita was relatively high; from 1875 to 1939, the average growth rate of personal consumption was 1.36%, and it was never negative except during the Russo–Japanese War period (Nakamura 1971, pp.
 2–5; 27; 30–31).
-This highlights the importance of revealing the consumption behavior of urban households to understand the mechanisms underpinning the rapid Japanese economic growth in this period.
 2.2 Tsukishima: A Representative Manufacturing Area
-The population in Tokyo City has been growing since the end of the 19th century, and industrialization due to WWI spurred its growth.
-Tokyo City’s population share increased from $2.5$% in 1889 to approximately $4$% by 1920.131313In 1920, there were more than $10,000$ municipalities in Japan, suggesting how large Tokyo was.
-There was a difference of about 1 million between Tokyo and Osaka, the second-largest city by population.
-Online Appendix Figure A.1a summarizes the trends in the number of people in the top five cities between 1889 and 1920.
-Throughout the 1910s, the percentage of industrial and commercial workers increased in Tokyo City, building the city as a gigantic commercial and industrial place (Tanimoto 2002, p.
-8).
-Column 1 of Panel A in Table 1 shows that nearly half of the male workers in Tokyo worked in the industrial sector as of 1920.
-Figure 2: Administrative Wards in Tokyo City
-Notes: The name in each lattice indicates the administrative ward in Tokyo.
-The gray ward indicates the Kyobashi Ward and an island in the Kyobashi Ward indicates Tsukishima.
-The border surrounded by Asakusa, Honjyo, Nihonbashi, and Fukagawa wards includes the Sumida River.
-The borders of each ward are based on the administrative districts in 1920.
-Geographical coordinate system is based on JGD2000/(B, L).
-Source: Created by the author using the official shapefile (Ministry of Land, Infrastructure, Transport and Tourism, database).
-The subject of this paper is Tsukishima, an industrial area located southeast of Kyobashi Ward in Tokyo (Figure 2).
-[^14] The base of the island was reclaimed during the Edo and Meiji periods from a delta at the mouth of the Sumida River.
-Online Appendix A.3 summarizes the development history of Tsukishima in detail.
-Column (2) of Panel A in Table 1 shows that Kyobashi had an industrial composition like that of Tokyo City in 1920.
-Column (3) confirms that Tsukishima is a representative industrial area in the city; approximately $68$% of male workers were employed in the manufacturing sector.
-In the same year, there were $214$ factories on this island, and approximately $80$% of them belonged to the machinery and related sectors (Department of Health, Ministry of the Interior 1923a, pp 385–389).
-[^15] According to the census of 1920, the number of workers in the metal and machinery equipment manufacturing industry ($27.1$%) was the highest compared to other industries in the entire city of Tokyo.
-The same tendency was observed in Kyobashi Ward ($35.5$%).
-Although similar statistics are not available for Tsukishima, this supports the evidence that Kyobashi was not an area with a unique industrial pattern (Statistics Bureau of the Cabinet 1929c, pp.
-85–86; 108–109)).
-Figure 3 shows the spatial distribution of factories in Tsukishima as of June 30, 1918.
-Tsukishima had many machine factories and shipyards, and the workers’ residences were clustered there.
-[^16] For example, a representative shipbuilding plant was located on Ishikawajima, an island in the north of Tsukishima (Figure 3).
-Ships were strongly influenced by the import substitution that was underway at the beginning of the 20th century, and domestic production of ships progressed (Section 2.1).
-The production of a hydraulic 80-ton hardened crane by Ishikawajima Shipyard in 1917 is an example of the domestic output of a product that had previously relied on imports (Nakamura and Odaka 1989, pp.
+Figure 2: Factories in Tsukishima
+Notes: Blocks in green and light green represent shipyards.
+Blocks in red and pink represent factories in the machinery sector.
+Blocks in blue and light blue represent factories in other manufacturing sectors.
+Blocks in brown indicate warehouses.
+Small circles colored black and white show workplaces in the machinery and other manufacturing sectors, respectively.
+The upper right corner indicates north.
+Source: Department of Health, Ministry of the Interior 1923c, second map.
+The tone was adjusted by the author using Adobe Photoshop 24.7.0.
+The population of Tokyo City had been growing since the end of the 19th century, and industrialization during WWI further spurred its growth.
+[^8] Tokyo City’s population share increased from 2.52.5% in 1889 to approximately 44% by 1920.
+Its population exceeded Osaka’s, the second-largest city, by about 1 million.
+In 1920, Japan had more than $10,000$ municipalities, underscoring Tokyo’s relative size.
+Online Appendix Figure A.1a summarizes the population trends of the top five cities between 1889 and 1920.
+Throughout the 1910s, the share of industrial and commercial workers in Tokyo City increased.
+According to the 1920 census, nearly half of the male workers in Tokyo were employed in the industrial sector (Column 1 of Panel A in Table 1).
+The subject of this study is Tsukishima, an island located southeast of Kyobashi Ward.
+[^9] Online Appendix Figure A.3 shows the location of Tsukishima.
+The development history of Tsukishima is summarized in Online Appendix A.3.
+Kyobashi has an industrial composition similar to that of Tokyo City (Column 2), and Tsukishima represents a typical industrial area within the ward, with approximately 6868% of male workers employed in the manufacturing sector (Column 3).
+In 1920, Tsukishima had $214$ factories, of which approximately 8080% belonged to the machinery and related sectors.
+[^10] Department of Health, Ministry of the Interior (1923a, pp.
+385–389).
+According to the 1920 census, the metal and machinery equipment manufacturing industry (27.127.1%) employed the largest share of workers in Tokyo.
+A similar pattern was observed in Kyobashi Ward (35.535.5%) (Statistics Bureau of the Cabinet 1929c, pp.
+85–86; 108–109).
+Figure 2 shows the spatial distribution of factories in Tsukishima as of June 30, 1918, demonstrating the concentration of machine factories and shipyards.
+[^11] A representative shipbuilding plant, Ishikawajima Shipyard, was located on Ishikawajima, an island north of Tsukishima.
+Import substitution during this period strongly accelerated domestic ship production (Section 2.1).
+The production of a hydraulic 80-ton hardened crane by Ishikawajima Shipyard in 1917 exemplifies domestic output of a product that had previously relied on imports (Nakamura and Odaka 1989, pp.
 10; 32).
-The focus of this study is the consumption behavior of these worker households in the industrial sector.
-Figure 3: Factories in Tsukishima
-Notes: Blocks in green and light green indicate the shipyards.
-Blocks in red and pink indicate the factories in the machinery sector.
-Blocks in blue and light blue indicate the factories in the other manufacturing sectors.
-Blocks in brown are the warehouses.
-Small circles colored in black and white show the workplaces in the machinery and other manufacturing sectors, respectively.
-Source: Department of Health, Ministry of the Interior 1923c, second map. The tone was adjusted by the author using Adobe Photoshop 24.7.0.
 Table 1: Industrial and Family Structures in Tsukishima
@@
 Average age of males 25.3 25.3 25.1
-Notes:
-Panel A: This panel summarizes the industrial structures based on the occupations of male workers measured in the 1920 population census.
-Each share is calculated as the number of males who worked as regular workers in each sector divided by the total number of male workers (in percentage).
-All the figures do not include the number of unemployed (mugyō sha) males.
-Panel B: The average household size is the number of people divided by the number of households.
-The share of married males is the number of married males living with/without spouses divided by the total number of males (in percentage).
-The average age of males is calculated using the tables for population by age group reported in the census.
-As the tables are open-ended, the class for over 60 years old is rounded as $64.5$ following the range of the former category (i.e., 50–59).
-All the figures in this panel do not include a small number of quasi-households (jyun setai), which include the person, such as the students in dormitories and the patients in hospitals.
-The figures for Tsukishima listed in column 3 of both panels are based on the total sum of statistics of all the blocks (chōme) in Tsukishima (i.e., from Tsukudajima to Tsukishima dōri jyūni chōme).
-Sources: Columns 1 and 2 of Panel A use the Statistics Bureau of the Cabinet (1929c, pp.
+Notes: Panel A: This panel summarizes the industrial structure based on the occupations of male workers recorded in the 1920 Population Census.
+Each share is calculated as the number of males employed as regular workers in a given sector divided by the total number of male workers (in percentage).
+All figures exclude unemployed (mugyō sha) males.
+Panel B: The average household size is calculated as the total number of people divided by the number of households.
+The share of married males is the number of married males, living with/without spouses, divided by the total number of males (in percentage).
+The average age of males is calculated using the population tables by age group reported in the census.
+For open-ended age categories, the class for those over 60 years old is rounded to $64.5$, following the range of the previous category (50–59).
+All figures in this panel exclude a small number of quasi-households (jyun setai), which include individuals such as students in dormitories and patients in hospitals.
+The figures for Tsukishima listed in Column 3 of both panels are based on the total sum of statistics across all blocks (chōme) in Tsukishima (i.e., from Tsukudajima to Tsukishima dōri jyūni chōme).
+Sources: Columns 1 and 2 of Panel A: Statistics Bureau of the Cabinet (1929c, pp.
 18–19).
-Columns 1 and 2 of Panel B use the Statistics Bureau of the Cabinet (1929c, pp.
+Columns 1 and 2 of Panel B: Statistics Bureau of the Cabinet (1929c, pp.
 38–43); Tokyo City Office (1922b, pp.
 2–3; 42–46).
-Column 3 of Panel A uses the Tokyo City Office (1922c, pp.
+Column 3 of Panel A: Tokyo City Office (1922c, pp.
 42–47).
-Column 3 of Panel B uses the Tokyo City Office (1922a, pp.
+Column 3 of Panel B: Tokyo City Office (1922a, pp.
 26–31); Tokyo City Office (1922b, pp.
 2–3; 42–46).
-Panel B in Table 1 summarizes the population and occupational statistics from the census to examine the characteristics of working-class households in Tsukishima.
-First, the average household size in Tsukishima is $4.3$—similar to but slightly smaller than the values in Tokyo City ($4.6$) and Kyobashi Ward ($4.7$).
-The percentage of the married male population is $40$%, which is similar to the range of the figures for Tokyo ($38$%) and Kyobashi ($36$%).
-The average age of men in Tsukishima is $25.1$, slightly lower than the figures for Tokyo and Kyobashi.
-While the sex ratio by age group in Tsukishima shows a similar trend to that of Tokyo and Kyobashi, the ratios for the 1414141419191919 and 2020202029292929 age bins are higher than those.
-Overall, these statistics reflect the fact that the typical household structure is a couple with two children.
-The slightly smaller household size in Tsukishima can be attributed to the relatively large number of young factory worker households.
-The percentage of married men in Tsukishima suggests that a certain proportion of the households in their early 20s might have been couples before they started raising children.
-Finally, I provide a brief overview of the lives of the factory workers’ households in Tsukishima.
-Factories and workers’ residences were scattered throughout the island.
-The factory workers lived in similar tenements and used the commercial stores and restaurants in the central lot.
-[^17] Tsukishima is composed of Ishikawajima island at the northern end, the adjacent Tsukudajima and Shin-tsukudajima islands, and the rectangular Lots 1–3 in the south (Online Appendix Figure A.3).
-Specifically, factories and worker residences were scattered throughout Shin-tsukudajima and Lots 1 and 2, whereas Lot 3, a relatively new reclaimed site completed in 1913, was undeveloped.
-Online Appendix A.4 summarizes the housing in Tsukishima in detail using a housing survey.
-There were $209$, $109$, $74$, and $120$ stores selling daily necessities, clothing, tools, and food, drink, and entertainment, respectively.
-Regarding medical personnel, there were eight doctors, four dentists, eight obstetricians, and $14$ drug stores.
-There were also ten bathhouses (Department of Health, Ministry of the Interior 1923a, pp.
-41–47).
-There were two postal offices, $13$ pawnshops, and five moneylenders, which were scattered throughout the island (Online Appendix Figure A.2).
+Panel B in Table 1 summarizes population and occupational statistics from the census to examine the characteristics of households in Tsukishima.
+The average household size ($4.3$) is similar to, but slightly smaller than, the values in Tokyo City and Kyobashi Ward.
+The percentage of married males (4040%) is comparable to the figures in Columns (1) and (2), while the average age of men ($25.1$) is slightly lower.
+Sex ratios follow a similar trend to those of Tokyo and Kyobashi, although the ratios for the 14141919 and 20202929 age bins are higher.
+Overall, these statistics suggest that the typical household consists of a couple with two children.
+The slightly smaller household size in Tsukishima can be attributed to the relatively large number of young factory-worker households, and the percentage of married men indicates that a portion of households in their early 20s may have been couples before starting a family.
+Finally, I provide a brief overview of life in factory workers’ households.
+Factories and workers’ residences were scattered throughout Tsukishima.
+Factory workers lived in similar tenements and made use of a wide variety of commercial stores in the three lots: Tsukudajima and Shin-tsukudajima in the north, Lot 1 in the middle, and Lot 2 in the south (Figure 2).
+[^12] The northernmost Ishikawajima island contains only the Ishikawajima Shipyard, and the southernmost rectangular Lot 3 is uninhabited.
+Online Appendix Figures A.4 and C.5 provide detailed maps of the lots and retail shops, respectively.
+Online Appendix A.4 summarizes housing in Tsukishima.
 The main routes to the mainland were the wooden Aioi Bridge (Aioi bashi) and three ferries (watashi bune).
-The average daily total number of people using the ferries in 1918 was reported to be $53,318$, accounting for more than twice as many people as the Tsukishima population.
-[^18] See Department of Health, Ministry of the Interior (1923a, p.
+The average daily number of ferry users in 1918 was $53,318$, more than twice the population of Tsukishima.
+[^13] Department of Health, Ministry of the Interior (1923a, p.
 30).
-The number of people and households in Tsukishima at the end of December 1918 was reported to be $24,399$ and $5,562$, respectively (National Police Agency 1920a, pp.
+The number of people and households in Tsukishima at the end of December 1918 was $24,399$ and $5,562$, respectively (National Police Agency 1920a, pp.
 50–51).
-Online Appendix Figure A.3 shows the routes of the ferries.
-These descriptions illustrate Tsukishima as a busy machine manufacturing area during industrialization.
+Online Appendix Figure A.4 shows the ferry routes.
+These figures illustrate Tsukishima as a busy machine manufacturing area during industrialization.
 2.3 Risk-coping Institutions
-Working-class households were exposed to several risks that led to losses of income at that time.
-Generally, social insurance, such as health, accident, and unemployment insurance, can mitigate idiosyncratic income shocks on workers.
-[^19] Gertler and Gruber (2002) shows the impact of illness on consumption-smoothing behavior in the early 1990s in Indonesia and argues for the importance of public compensation for illness.
-Kantor and Fishback (1996) show that the installation of workers’ compensation reduced the precautionary savings in the interwar US.
-Gruber (1997) provides evidence that unemployment insurance smooths individual consumption as was the case in the US from 1968–1987.
-However, comprehensive social insurance systems did not exist in the early 1920s in Japan.
-The previous health insurance was enacted in 1927, and unemployment insurance was not established throughout the prewar period.
-Additionally, the Factory Act enacted in 1916 did not specify compensation for the sickness of workers.
-Therefore, they had no public support for their illness, injuries, and unemployment.
-Instead, they might have used private financial institutions to manage idiosyncratic risks.
-Several financial institutions were available for working-class households.
-A representative savings institution was the postal savings (yūbin chokin).
-In Tokyo City, approximately $1.4$ million people had postal savings accounts, covering approximately $67$% of the citizens in 1920 (Online Appendix Table A.2).
-The average savings amount per capita in Kyobashi Ward was $44.5$ yen (Tokyo City Office 1921, pp.
-892–893), which was less than one month’s earnings for the skilled-factory workers (Section 3.2).
-Savings bank (chochiku ginkō) was another famous savings institution among the working class in the 1910s (Ito and Saito 2019, pp.
-79–81).
-The number of people with savings accounts was approximately $1.4$ million, accounting for approximately $58$% of total people in Tokyo (Online Appendix Table A.2).
-However, $84.8$% of all the depositors were commercial and miscellaneous workers, and the manufacturing sector shared only $11.5$% in Kyobashi Ward (Online Appendix Table A.3).
-This reflects the fact that savings banks were institutions designed for lower-income workers rather than factory workers (Tokyo Institute for Municipal Research 1925b, p.
+Generally, social insurance such as health, accident, and unemployment insurance can mitigate idiosyncratic income shocks (Gertler and Gruber 2002; Kantor and Fishback 1996; Gruber 1997).
+However, comprehensive social insurance systems did not exist in Japan in the early 1920s.
+Health insurance was not enacted until 1927, and unemployment insurance was absent throughout the prewar period.
+The Factory Act enacted in 1916 did not specify compensation for workers’ sickness.
+As a result, factory workers had no public support for illness, injuries, or unemployment.
+Instead, several financial institutions were available for factory workers.
+[^14] Finer details of these institutions and the statistics used in this subsection are summarized in Online Appendix A.2.
+A representative savings institution was postal savings (yūbin chokin).
+In Tokyo City, approximately $1.4$ million people held postal savings accounts, covering approximately 6767% of the citizens in 1920.
+The average savings per capita in Kyobashi Ward was $44.5$ yen, less than one month’s earnings for skilled factory workers (Section 3.2).
+Savings banks (chochiku ginkō) were another well-known savings institution.
+About $1.4$ million people held accounts, representing approximately 6767% of Tokyo’s population.
+However, 84.884.8% of all depositors were commercial and miscellaneous workers, while only 11.511.5% belonged to the manufacturing sector in Kyobashi Ward.
+This reflects that savings banks were designed for lower-income workers such as those in miscellaneous industries and day laborers, rather than factory workers (Tokyo Institute for Municipal Research 1925b, p.
 81).
-This is consistent with the fact that Tsukishima did not have any savings bank branches, while there were two postal offices (Section 2.2).
-Although another savings institution was the mutual loan association (mujin), it was used among small business owners and merchants for running their businesses (Online Appendix A.2).
+This is also consistent with the fact that Tsukishima had no savings bank branches but did have a few post offices.
+Another savings institution, the mutual loan association (mujin), was mainly used by small business owners and merchants to run their businesses.
 [^15] Ordinary banks in prewar Japan provided large loans to businesses (Teranishi 2011).
-There was no bank in Tsukishima, as workers did not borrow money from banks to make ends meet.
-The official report states that “a commercial bank is not necessary in a laboring area such as Tsukishima, and even for large entrepreneurs such as factory owners, there is no particular need to establish a bank for their business transactions; transactions with banks in other areas are sufficient” (Department of Health, Ministry of the Interior 1923a, p.
+Tsukishima had no bank, since workers did not borrow money from banks to make ends meet (Department of Health, Ministry of the Interior 1923a, p.
 48).
-There were also few lending institutions.
-Credit purchases (kakegai) with retailers were the most popular institution.
-Although systematic statistics of credit purchases are unavailable, the THBS households often used credit purchases in various retailers for rice, fish, vegetables, firewood, and charcoal.
-[^21] Cooperative societies were also available for worker households to buy daily commodities, although they were not lending institutions.
-There were $26$ cooperative societies in Tokyo prefecture in 1924, of which 23 were organized by the citizens and workers (Central Federation of Industrial Associations 1925, p.
+Credit purchases (kakegai) with retailers were the most common informal credit institution.
+People often used credit purchases at various retailers for rice, fish, vegetables, firewood, and charcoal in Tokyo City.
+This institution is described in detail in Section 5.
+Pawnshops (shichiya) were widely distributed lending institutions in the city.
+[^16] The number of pawnshops in Tokyo was $1,334$ and $1,261$ in 1918 and 1919, with total transactions of $8,226,883$ and $7,573,406$ each year, respectively–more than three times the population of Tokyo City.
+However, evidence shows that the most frequent users were day laborers and workers classified as “miscellaneous,” accounting for 63.463.4% of all users.
+By contrast, the shares for agriculture, commerce, and manufacturing sectors were only $1.7$, $19.5$, and 15.615.6%, respectively (Tokyo Institute for Municipal Research 1926, p.
+25).
+In practice, pawnshop lenders did not need to screen borrowers’ credit, and borrowers did not face the risk of heavy debt, as the main items pawned were inexpensive clothes (Shibuya et al.
+1982).
+Interest rates were regulated by the Pawnbroker Regulation Act of 1895, and the average redemption rate was high under these lower rates.
+[^17] The average redemption rate was approximately 9494% in 1920 among $88$ pawnshops in Kyobashi (401,222401,222/$427,265$ cases) (Tokyo City Office 1922b, pp.
+888–889).
+In short, the primary users of pawnshops were low-income workers who needed credit for daily necessities, rather than factory workers (Tokyo City Social Affairs Bureau 1921, p.
+9).
+Other lending institutions included moneylenders (kinsen kashitsuke gyō) and credit unions (shinyō kumiai).
+Moneylenders were generally used by business owners and charged very high interest rates (Shibuya 2000, pp.
+184; 248).
+Credit unions had very limited coverage, likely serving only a small share of male workers in the manufacturing sector.
+Informal gifts may have served as another type of risk-mitigation institution.
+While systematic statistics are unavailable, a household survey of $185$ factory-worker households in Tokyo found that the average monthly gifts accounted for only 33% of total income in November 1922 (Social Affairs Division 1925, pp.
+58–59).
+Therefore, whether these small informal transfers functioned as effective mutual aid remains unclear.
+Life insurance was not designed to compensate for temporary income losses.
+Life insurance companies generally provided death insurance, which did not cover short-term income reductions, and they did not target working-class customers (Usami 1984, pp.
+87–88).
+Although postal life insurance (kani seimei hoken) operated by the government began in 1916, it offered only life and endowment insurance and did not cover illness or unemployment (Usami 1984, pp.
+102–107).
+[^18] Several employment agencies existed in Tokyo to introduce jobs to low-income job seekers.
+However, many were unscrupulous and not commonly used at the time (Machida 2016).
+Finally, cooperative societies were available to help worker households purchase daily commodities.
+In 1924, Tokyo prefecture had $26$ cooperative societies, of which 23 were organized by citizens and workers (Central Federation of Industrial Associations 1925, p.
 57–59).
-However, while the two largest societies had approximately $4,000$ members, the others were generally small societies with less than 100 members.
-This figure accounts for only a few percentage of the workers in the city, given that the number of male workers was more than $768$ thousand.
-In fact, the THBS households rarely used the cooperatives.
-I will describe this institution in detail in Section 5.
-The pawnshop (shichiya) was another lending institution.
-The accessibility was higher than for the other lending institutions.
-[^22] The number of pawnshops in Tokyo was $1,334$ and $1,261$ in 1918 and 1919, respectively.
-The total number of cases was $8,226,883$ and $7,573,406$ each year.
-These figures are more than three times greater than the entire population in Tokyo City.
-There were $13$ pawnshops in Tsukishima, more than twice the number of moneylenders (Section 2.2).
-However, the primary users of the pawnshops were the low-income working-class households rather than the factory workers (Tokyo City Social Affairs Bureau 1921, p.
-9).
-According to the statistics on the pawnshop users in Tokyo prefecture in 1923, the most representative users were the day laborers and workers classified as “miscellaneous,” accounting for $63.4$% of all users.
-By contrast, those for agriculture, commerce, and manufacturing sectors were only $1.7$, $19.5$, and $15.6$%, respectively (Tokyo Institute for Municipal Research 1926, p.
-25).
-The inexpensive articles pawned suggest that the pawnshops were only used to dealing with very short-run loans for necessities.
-Among the pawnshops of Kyobashi, for example, the average amount per case was $7.8$ yen, which accounts for $12$% of the monthly disposable income of the THBS households (Panel A of Table 3).
-Thus, the pawnshop was a small lending institution for low-income working-class households who needed credit.
-Indeed, lenders did not need to screen the borrower’s credit, and the borrowers did not worry about incurring heavy debt as the main articles pawned were inexpensive clothes (Shibuya et al.
-1982).
-Additionally, the interest rates were regulated by the Pawnbroker Regulation Act of 1895, and the average redemption rate was substantially high under the lower interest rates.
-[^23] The average redemption rate was approximately $94$% in 1920 among $88$ pawnshops in Kyobashi ($401,222$ cases/$427,265$ cases) (Tokyo City Office 1922b, pp.
-888–889).
-Other lending institutions were money lenders (kinsen kashitsuke gyō) and credit unions (shinyō kumiai).
-Generally, moneylenders were used by business owners with very high interest rates (Shibuya 2000, pp.
-184; 248).
-Similarly, the coverage of the credit unions was substantially low, possibly a few percentage points of the male workers in the manufacturing sector (Online Appendix A.2).
-Informal gifts might have been another type of risk-mitigation institution, although systematic statistics are not available.
-A household survey on the $185$ factory-worker households in Tokyo suggests that the share of average monthly gifts in total income was approximately $3$% in November 1922 (Social Affairs Division 1925, pp.
-58–59).
-While this means that the gifts comprised a few percentage points of the monthly income, it is unclear whether these informal income transfers functioned as mutual aid.
-Finally, life insurance was not designed to compensate for temporary income losses.
-Generally, life insurance companies provided death insurance, meaning that temporary income reductions were not covered.
-Moreover, they did not target the working class as customers (Usami 1984, pp.
-87–88).
-Although postal life insurance (kani seimei hoken) operated by the government started in 1916, it only provided life and endowment insurance.
-Thus, it never covered illness or unemployment (Usami 1984, pp.
-102–107).
-[^24] There were also several employment agencies that introduced jobs to low-income job-seekers in Tokyo.
-However, many of them were unscrupulous, and not commonly used at that time (Machida 2016).
-To summarize, while there were several lending institutions, few were available to the factory workers.
-Postal savings might have been the best available device for them.
-Credit purchase was the most plausible device regarding lending institutions.
-While pawnshops were accessible, they might not have provided sufficient money as the articles pawned were usually inexpensive.
-The factory workers in Tsukishima might have combined these devices to manage idiosyncratic income shocks.
+While the two largest societies had approximately $4,000$ members each, the others were generally small, with fewer than 100 members.
+Overall, this accounted for only a small share of workers in the city.
+To summarize, although several risk-coping institutions existed, few were accessible to factory workers.
+Postal savings were the main official financial institution available, while credit purchases were the most plausible informal mechanism.
+Factory workers in Tsukishima likely combined these devices to manage idiosyncratic income shocks.
 3 Data
@@
 The Tsukishima Survey was the first urban social survey conducted in Japan by the Ministry of Home Affairs circa 1919.
-The purpose was to investigate the status of the lives of urban worker households amid the rapid industrialization throughout WWI.
-It was designed as a field survey consisting of multiple survey items, including the budget survey (hereinafter THBS).
-[^25] The THBS was the first household survey in Japan to use the budget book method and is considered the prototype for subsequent household surveys in Japan (Sekiya 1970, p.
+Its purpose was to examine the living conditions of urban worker households amid rapid industrialization during WWI.
+The survey was designed as a field study with multiple components, including a budget survey (hereinafter THBS).
+The Ministry of Home Affairs published several official reports in 1921, which contained aggregated information on the target households.
+[^19] The THBS was the first household survey to use the budget book method and is considered the prototype for subsequent household surveys in Japan (Sekiya 1970, p.
 43).
-Online Appendix B.1 summarizes detailed information about the Tsukishima Survey.
-The Ministry of Home Affairs published a few official reports on the Tsukishima Survey in 1921, which contained some aggregated information on the target households.
-However, household-level budget information is required to analyze consumption-smoothing strategies.
-Fortunately, the household budget book (kinsendeiri hikaechō) gathered in the THBS is preserved in the archives of the Ohara Institute of Social Research (hereafter, OISR).
-[^26] The author was permitted to access those unreleased materials related to the Tsukishima Survey by the OISR: Archives of the Tsukishima Survey (THBS, unreleased).
-The official reports of the survey published in 1921 by the Ministry of Home Affairs are included in the archives of Gonda Yasunosuke at the OISR and are available to the public (The OISR, Archives of Gonda Yasunosuke (7-2; 7-3; 7-4)).
-These budget books were transferred to the OISR by Iwasaburo Takano, who was responsible for the Tsukishima Survey as a member of the 7th Section of the Health and Hygiene Research Committee of the Ministry of Home Affairs and became the first director of the OISR.
-[^27] Iwasaburo Takano was a professor at Tokyo Imperial University and a specialist in social statistics and social policy.
-He emphasized gathering statistics about workers through social surveys and is known as a researcher who significantly contributed to the formation of social survey methodology in Japan.
-His contribution to the Tsukishima Survey is summarized in detail in the Online Appendix B.1.
-I corrected and digitized all forty THBS budget books for the factory worker households that were included in the official report.
-[^28] The archives of the OISR also preserve several other household budget books of factory worker households.
-However, these budget books contain poor record-keeping.
-Therefore, I do not use these incomplete books in this study.
-The characteristics of the remaining household budget books are summarized in detail in the Online Appendix B.5.
-During the study period, each household recorded their income and expenses daily in a budget book.
-Notably, a survey office was set up in Tsukishima, where a few surveyors lived and were closely connected with the survey households.
-Furthermore, weekly administrative meetings were held to discuss the progress of the survey (Miyoshi 1980, p.
+The official reports published in 1921 by the Ministry of Home Affairs are included in the archives of Gonda Yasunosuke and are available to the public (The OISR, Archives of Gonda Yasunosuke (7-2; 7-3; 7-4)).
+However, household-level budget data are necessary to analyze consumption-smoothing strategies.
+Fortunately, the household budget books (kinsendeiri hikaechō) collected in the THBS are preserved in the archives of the Ohara Institute of Social Research (OISR).
+These budget books were transferred to the OISR by Professor Iwasaburo Takano of Tokyo Imperial University.
+He served on the research committee of the Ministry of Home Affairs responsible for the Tsukishima Survey and later became the first director of the OISR.
+[^20] Online Appendix B.1 provides further details on how this survey was conducted as part of Iwasaburo Takano’s ambitious social survey project.
+I collected and digitized all forty THBS budget books for the factory-worker households.
+[^21] The author was permitted access to the unreleased materials related to the Tsukishima Survey by the OISR: Archives of the Tsukishima Survey (THBS, unreleased).
+The characteristics of the budget books are summarized in detail in Online Appendix B.2.
+During the study period, each household recorded its daily income and expenses in a budget book.
+Notably, a survey office was established in Tsukishima, where a few surveyors lived and maintained close contact with the surveyed households.
+Weekly administrative meetings were held to monitor the survey’s progress (Miyoshi 1980, p.
 38).
-This confirms that the information in these budget books is reliable and that the THBS budget dataset ensures a quality that enables quantitative analysis.
-[^29] Online Appendix B.1 provides finder details on how the Tsukishima Survey was conducted carefully as a social survey project of Takano.
 3.2 Sample Characteristics
 Analytical Sample
-The estimation strategy used for testing consumption-smoothing and risk-coping strategies requires panel structures for aggregate income and expenditure (Section 4).
-I use the semi-monthly series for the baseline analysis, as the head’s income was paid twice a month.
+The estimation strategy requires panel structures for aggregate income and expenditure (Section 4).
+I use the semi-monthly series for the baseline analysis, as the household head’s income was paid twice a month.
 [^22] Section 3.3 provides details on the definitions of aggregate income and expenditure.
-Among the $40$ THBS households, there are five cross-sectional units in our semi-monthly data.
-I excluded these five cross-sectional units, leaving $35$ households.
-Next, I removed a household with incomplete income information.
-Finally, I excluded a household for which no information on payments to the credit purchases was recorded.
-Consequently, my THBS sample for the semi-monthly series included $33$ households with panel structures from January 12 to November 11, 1919.313131I also used the adjusted monthly panel dataset for sub-analysis in Section 4.
-Following the same trimming steps, my adjusted monthly panel dataset included $26$ households for the same period.
+Among the $40$ THBS households, five cross-sectional units appear in my semi-monthly data.
+I exclude these, leaving $35$ households.
+I then remove a household with incomplete income information and another for which no information on payments for credit purchases are recorded.
+Consequently, the THBS sample for the semi-monthly series includes $33$ households with panel structures from January 12 to November 11, 1919.232323I also use the adjusted monthly panel dataset for the sub-analysis in Section 4.
+Following the same trimming steps, the adjusted monthly panel dataset includes $26$ households for the same period.
 Note that several units have only three semi-month cells, for example, less than two adjusted months.
-These units have a panel structure in the semi-monthly panel dataset, but not in the adjusted monthly panel.
-Online Appendix B.3 summarizes the finer details of the original data structure and trimming.
-Before assessing the representatives of the THBS sample, I tested the potential selection bias due to attrition of the units.
-The panel units with shorter (longer) time-series observations might have different preferences on the consumption than those with longer (shorter) time-series observations.
-In other words, the households with shorter panels should have similar family size characteristics to those with more extended panels.
-To test this potential issue, I regressed an indicator variable for the shorter panel units on the family size variables.
+These units have a panel structure in the semi-monthly dataset but not in the adjusted monthly panel.
+Online Appendix B.3 summarizes the original data structure and trimming in detail.
+Before assessing the representativeness of the THBS sample, I evaluate potential selection bias due to unit attrition.
+Panels with shorter (or longer) time-series observations might have different consumption preferences than those with longer (or shorter) observations.
+In other words, households with shorter panels should have family size characteristics similar to those with longer panels.
+To test this, I regress an indicator variable for the shorter panel units on family size variables.
 Online Appendix Table B.2 summarizes the results.
-All the estimated coefficients on the covariates were close to zero and statistically insignificant, and the Wald statistics support the null results.
-This is robust to using different thresholds such as first quantile, median, and third quantile.
-This supports the evidence that the household’s preference was less likely to be correlated with a lack of balance.
+All estimated coefficients on the covariates are close to zero and statistically insignificant, and the Wald statistics support the null hypothesis.
+This result is robust to using different thresholds such as the first quantile, median, and third quantile.
+These findings suggest that household preferences are unlikely to be correlated with panel length.
 Representativeness
@@
 Survey subject All factories with All factories with # of 33 households
-# of workers $\geq 15$ workers $\geq 15$ workers $<15$
+# of workers $\geq 15$ workers $\geq 15$ workers <15<15
 Textile 0.2 2.9 0.0 0
@@
 Sex ratio 1.3 1.2 1.2
-Household size (share)
+Household size (% share)
      1 6.0 0
@@
 (Ave. age: 33.2) in machinery factories
-(Age range: 30-40)
+(Age range: 30–40)
 Average monthly earnings 59.1 (median = 56.2) 56.0
 95%CI [53.7, 64.5]
-Notes: Panel A: Column 1 shows the share of factory workers to the total factory workers in each sector (in percentage).
-The number of factory workers includes female and male factory workers in the 68 factories, with 15 workers and more.
-Column 2 indicates the share of factories to the total number of factories in each sector by factory size (in percentage).
-Column 3 summarizes the manufacturing sectors for the heads of the THBS.
-Columns 2 and 3 use the classification definition in the Statistics of the National Police Agency (column 1).
-Following this definition, several “woodworking” factories measured in the Tsukishima Factory Survey are included in the machinery sector in column 2, as it consists of shipbuilding and wooden pattern factories.
-The report also supports that the wooden patterns are used in the casting process and shall be included in the machinery sector (Department of Health, Ministry of the Interior 1923a, pp.
+Notes: Panel A: Column 1 shows the share of factory workers in each sector relative to the total number of factory workers (percentage).
+This includes both female and male workers in the 68 factories with 15 or more employees.
+Column 2 indicates the share of factories in each sector by factory size (percentage).
+Column 3 summarizes the manufacturing sectors of the THBS household heads.
+Columns 2 and 3 follow the classification used in the Statistics of the National Police Agency (Column 1).
+According to this classification, several “woodworking” factories recorded in the Tsukishima Factory Survey are included in the machinery sector in Column 2, as it covers shipbuilding and wooden pattern factories.
+The report also confirms that wooden patterns are used in the casting process and should therefore be included in the machinery sector (Department of Health, Ministry of the Interior 1923a, pp.
 388–389).
-Panel B: The first and second rows in columns 1 and 2 show the average household size and sex ratio measured in the 1920 Population Census and Statistics of the National Police Agency of 1919, respectively.
-The fourth to eighth rows in columns 1 and 3 show household size distributions in the census and the THBS sample, respectively.
-The figures in parentheses in Column (1) are the percentage share for family size bins between 2 and 8 people.
-The ninth row in column 3 shows the $p$-value from the Pearson $\chi^{2}$ test for the equality of the household size distributions between 2 and 8 people.
-The share in each household size bin measured in the census is used as the theoretical probability in calculating the test statistic (Online Appendix B.5).
-Panel C: Column 1 lists the average monthly earnings of the THBS heads calculated using the adjusted monthly panel dataset.
-Of the 33 THBS households, information from 26 households for which adjusted monthly income can be traced back is used in this panel to follow the unit in the manufacturing census.
-Comparison based on the average semi-monthly income from all 33 households shows a similar result.
-Gaussian-based 95% confidence interval (CI) using bootstrap standard error is reported in the brackets.
-Column 2 shows the average monthly wage of the male factory workers in the skilled age range (30–40 years old) in machinery factories in Tokyo City.
-This figure is calculated using the average monthly wage and average monthly ancillary wage measured in the manufacturing censuses.
-All the wage statistics are deflated using the CPI for cities provided by Ohkawa et al.
+Panel B: The first and second rows in Columns 1 and 2 show the average household size and sex ratio from the 1920 Population Census and the 1919 Statistics of the National Police Agency, respectively.
+Rows four to eight in Columns 1 and 3 present household size distributions in the census and the THBS sample, respectively.
+Figures in parentheses in Column 1 indicate the percentage share for family size bins from 2 to 8 people.
+Row nine in Column 3 reports the $p$-value from the Pearson $\chi^{2}$ test for equality of household size distributions between 2 and 8 people.
+The census shares in each household size bin are used as the theoretical probabilities in calculating the test statistic (Online Appendix B.5).
+Panel C: Column 1 lists the average monthly earnings of THBS household heads, calculated from the adjusted monthly panel dataset.
+Of the 33 THBS households, data from 26 households–where adjusted monthly income can be traced–are used in this panel to match units in the manufacturing census.
+Comparison using average semi-monthly income from all 33 households produces similar results.
+Gaussian-based 95% confidence intervals (CI) using bootstrap standard errors are reported in brackets.
+Column 2 shows the average monthly wage of male factory workers aged 30–40 in machinery factories in Tokyo City.
+This figure is calculated using average monthly wages and ancillary wages from the manufacturing censuses.
+All wage statistics are deflated using the CPI for cities from Ohkawa et al.
 (1967, p.
 255).
-Online Appendix B.6 summarizes the finer details of the process of this wage estimation.
-Sources: Panel A: Column 1 uses the National Police Agency (1920b, p.
-236).
-Column 2 uses the Department of Health, Ministry of the Interior (1923a, p.
-385–386).
-Column 3 uses the Department of Health, Ministry of the Interior (1923a, p.
+Online Appendix B.6 provides further details on the wage estimation process.
+Sources: Panel A: Column 1: National Police Agency (1920b, p.
+236); Column 2: Department of Health, Ministry of the Interior (1923a, p.
+385–386); Column 3: Department of Health, Ministry of the Interior (1923a, p.
 154).
-Panel B: Column 1 uses the Tokyo City Office (1922a, pp.
+Panel B: Column 1: Tokyo City Office (1922a, pp.
 26–31; 262–283); Tokyo City Office (1922b, pp.
-2–3; 42–46).
-Column 2 uses the National Police Agency (1920b, p.
-53).
-The figures in column 3 are calculated by the author using the THBS dataset.
-Panel C: Column 1 uses the THBS dataset.
-Column 2 uses the Tokyo City Statistics Division (1926a, pp.
+2–3; 42–46); Column 2: National Police Agency (1920b, p.
+53); Column 3: Calculated by the author from the THBS dataset.
+Panel C: Column 1: THBS dataset; Column 2: Tokyo City Statistics Division (1926a, pp.
 16; 244–245); Tokyo City Office (1921, pp.
 726–757).
-Next, I assess the representativeness of the THBS sample in Tsukishima using the manufacturing and population censuses.
 Panel A of Table 2 summarizes the share of male workers in the manufacturing industry by sector.
-Statistics of the National Police Agency (hereafter SNPA) covering factories with more than or equal to 15 workers show that approximately 90% of male workers were engaged in the machinery sector in 1919 (Column 1).
-Column 2 summarizes the statistics in the manufacturing census conducted in the Tsukishima Factory Survey.
-The proportion of machine factories is $85$% in larger factories and $92$% in smaller factories, indicating that most of the male workers worked in the machinery sector.
-[^32] I have confirmed that the paternal occupational statistics for the primary school students in Tsukishima provide materially similar results (Online Appendix A.5).
-In Column 3, I confirm that the THBS heads’ occupations show a similar proportion: $94$% worked in the machinery sector.
-In fact, the report states that “It is not too much to say that most of the industry in Tsukishima is machinery manufacturing.
+Statistics from the National Police Agency (SNPA) for factories with 15 or more workers show that approximately 9090% of male workers were employed in the machinery sector in 1919 (Column 1).
+Column 2 presents statistics from the Tsukishima Factory Survey.
+The proportion of machinery factories is 8585% among larger factories and 9292% among smaller factories, indicating that most male workers were in the machinery sector.
+[^24] Paternal occupational statistics for primary school students in Tsukishima yield materially similar results (Online Appendix A.5).
+Column 3 confirms that THBS household heads show a similar pattern, with 9494% working in the machinery sector.
+The report states: “It is not too much to say that most of the industry in Tsukishima is machinery manufacturing.
 Therefore, to describe the labor situation in Tsukishima, it is sufficient to describe its machinery industry” (Department of Health, Ministry of the Interior 1923a, p.
 389).
-Figure 3 illustrating a number of machinery factories supports this statement.
-As for the other sectors, the THBS includes a few heads who worked in food and miscellaneous sectors, and the proportions are similar to those listed in Column 2.
-Panel B of Table 2 summarizes the family structure.
-The average household sizes are identical in the Population Census (Column 1), SNPA (Column 2), and THBS (Column 3).
-The sex ratio is also similar across the surveys (1.21.21.21.21.31.31.31.3).
-The number of people in the THBS households ranges from 2 to 8, meaning the THBS sample does not include single or very large households.
-[^33] Those family households with 2–8 people cover most of the households in Tsukishima because the share of single and very large (9+ people) households was approximately $10$% (Column 1 of Panel B).
-In other words, the THBS mainly focuses on the consumption behavior in couples and households with a few children.
-For this household size range, the Pearson $\chi^{2}$ test does not reject the null of equality to the household size distribution measured in the census ($p$-value = $0.948$), meaning that the THBS sample reasonably approximates the distribution of family households in Tsukishima.
-Online Appendix B.5 illustrates the household size distributions in finer detail.
-This confirms the validity of the statement in the report that the THBS households can be “regarded as representative of the family form in Tsukishima” (Department of Health, Ministry of the Interior 1923a, p.
+The rest of the THBS heads work in food or miscellaneous industries, with proportions similar to those in Column 2.
+Panel B of Table 2 summarizes family structure.
+Average household sizes are identical across the Population Census (Column 1), SNPA (Column 2), and THBS (Column 3).
+Sex ratios are also similar across the surveys (1.21.21.31.3).
+THBS households range from 2 to 8 members, excluding single-person and very large households.
+In other words, the THBS mainly captures consumption behavior in couples and households with a few children.
+For this range, the Pearson $\chi^{2}$ test does not reject the null hypothesis of equality with the census household size distribution ($p$-value = $0.948$), indicating that the THBS sample reasonably approximates the distribution of family households in Tsukishima.
+Online Appendix B.5 provides a more detailed view of household size distributions.
+This supports the report’s statement that THBS households can be “regarded as representative of the family form in Tsukishima” (Department of Health, Ministry of the Interior 1923a, p.
 145).
-The official report of the Tsukishima Survey suggests that approximately nine out of ten factory workers in the skilled-age range were classified as skilled workers.
-Approximately six of those skilled workers were employed in the middle- to large-scale factories, and three worked in small-scale factories.
-[^34] Online Appendix A.5 summarizes this point in detail using the complete survey of the primary school students in the 4th to 6th grades in Tsukishima.
-Given this, the THBS was designed to investigate the budgets in skilled workers’ households.
-[^35] Shortly, Iwasaburo Takano, a survey director, regarded that the skilled workers in the modern industrial sector constituted the core of the labor force, labor movement, and citizens.
-The finer details in the background of this survey design are summarized in Online Appendix B.1.
-The average THBS head’s age is $33$, which is indeed in the skilled workers’ representative age range in the machinery factories (Kitazawa 1924).
-[^36] Specific occupations are not available for most heads because they are usually listed as “factory workers” in the occupation section.
-However, sub-categories available for some heads indicate that they were the typical skilled workers in machinery factories such as lathe operators and finishers.
-Online Appendix A.6 summarizes the work in the machinery factories in Tsukishima.
-Column 1 of Panel C in Table 2 indicates that the average monthly earning of the THBS heads is $59$ yen (median = $56$ yen).
-In Column 2, I calculated the average monthly earning of male machinery factory workers in the skilled-age range using the manufacturing censuses, which shows similar monthly earnings ($56$ yen).
-[^37] Shortly, this figure is calculated using the average daily wage of male factory workers aged 30–40 and average annual working days measured in the manufacturing censuses.
-Online Appendix B.6 provides the finer details of the calculation steps.
-The average wages for the male factory workers had an inverted-U-shaped distribution with respect to age, taking the greatest figure in the 30s (Online Appendix Figure B.7).
-This is a similar life-cycle pattern to the factory workers observed in late 19th century England (Horrell and Oxley 2000, p.
-42).
-This provides evidence that the monthly earnings of the THBS heads and skilled workers’ households in Tsukishima were materially similar.
-To summarize, this evidence indicates that the THBS sample is useful for inferring the mean tendency of those households.
+The official Tsukishima Survey report suggests that approximately nine out of ten factory workers in the skilled-age range are classified as skilled workers.
+Approximately six of these skilled workers are employed in medium- to large-scale factories, and three in small-scale factories.
+[^25] Online Appendix A.5 provides detailed evidence using the complete survey of primary school students in Tsukishima.
+Accordingly, the THBS is designed to examine household budgets of skilled workers.
+The average age of THBS household heads is $33$, which falls within the representative age range of skilled workers in machinery factories (Kitazawa 1924).
+[^26] Specific occupations are not available for most heads, as they are generally listed as “factory workers.” However, sub-categories available for some heads indicate typical skilled machinery factory roles such as lathe operators and finishers.
+Online Appendix A.6 summarizes work in the machinery factories in Tsukishima.
+The background and rationale for focusing on skilled factory workers in this survey are detailed in Online Appendix B.1.
+Column 1 of Panel C in Table 2 shows that the average monthly earnings of THBS heads is $59$ yen (median = $56$ yen).
+Column 2 reports the average monthly earnings of male machinery factory workers in the skilled-age range, calculated using the manufacturing censuses, which is similar at $56$ yen.
+[^27] This figure is calculated using the average daily wage of male factory workers aged 30–40 and the average annual working days measured in the manufacturing censuses.
+Online Appendix B.6 provides detailed calculation steps.
+Average wages for male factory workers follow an inverted-U-shaped distribution with respect to age, peaking in the 30s (Online Appendix Figure B.7).
+These results indicate that the monthly earnings of THBS household heads closely match those of skilled workers’ households in Tsukishima.
+Overall, this confirms that the THBS sample is suitable for analyzing the mean tendencies of skilled factory workers’ households in Tsukishima.
 3.3 Aggregation
@@
 by Different Time-series Frequencies
-Notes: Figure 4a4c, and 4e illustrate the time-series plots of the average daily, semi-monthly, and unadjusted semi-monthly income and expenditure, respectively.
-Figure 4b4d, and 4f illustrate the time-series plots of the average daily, semi-monthly, and unadjusted semi-monthly net income (i.e., income minus expenditure).
-Figures 4a and 4b show the daily data between January 12 and November 11, 1919.
-Figures 4c and 4d show the semi-monthly series calculated using the daily data between January 12 and November 11, 1919.
-The (adjusted) semi-monthly series defines the 12–26th for the first and 27–11th for the second half.
-Figures 4e and 4f show the unadjusted semi-monthly series calculated using the daily data between January 15 and October 31, 1919.
-The unadjusted semi-monthly series uses the 15th of each month as a threshold.
+Notes: Figures 3a3c, and 3e present time-series plots of average daily, semi-monthly, and unadjusted semi-monthly income and expenditure, respectively.
+Figures 3b3d, and 3f show the corresponding time-series plots of average daily, semi-monthly, and unadjusted semi-monthly net income (income minus expenditure).
+Figures 3a and 3b display daily data from January 12 to November 11, 1919.
+Figures 3c and 3d show the semi-monthly series calculated from the daily data over the same period.
+In the adjusted semi-monthly series, the first half covers the 12–26th and the second half the 27–11th.
+Figures 3e and 3f present the unadjusted semi-monthly series calculated from daily data between January 15 and October 31, 1919, using the 15th of each month as the threshold.
 Source: Created by the author using the THBS sample.
-It is necessary to summarize the daily budget information into a more extended time range because households need a set of time to smooth their consumption given a realized amount of income.
-To explore a reasonable method of aggregation, I first overviewed the daily time series on the total income and expenditure.
-Figure 4a shows that the head’s income was paid twice monthly: the semi-monthly wages were paid on the middle (14–16th) and last day (30th or 31st).
-[^38] The head’s income was the dominant method for earnings because the breadwinner households were typical at that time (Panel B in Table 3).
-Importantly, Figure 4a indicates that the households expensed most of their income on the payday and spent the rest of the payment little by little until the next payday.
-In other words, the expenditure never clearly increased just before the paydays.
-This means that the households made their consumption decisions (schedules until the next payday) based on the income they obtained.
-Note that calendar (semi-)monthly aggregation cannot reflect this household behavior because calendar (semi-)month does not place the payday at the beginning of its cell.
-[^39] For example, a calendar semi-month cell in January, for example, the cell from 1st to 14th, starts on the 1st but its payday is set for the 14th.
-Accordingly, this assumes that the aggregate consumption in this semi-month cell is based on the earnings at the end of its cell.
-This is clearly an unrealistic assumption.
-Another source of measurement error is the shift of the payday: the payday for the latter half of a month was sometimes set on the first date of the following month.
-Although this shift did not occur frequently, the calendar month cell might contain substantial interference from the shift as it leads to the cells including more (three paydays) and fewer (one payday) payments than normal.
-Nelson (1994) suggests that this sort of mismatching in the timing of income and expenditure induces type II errors when testing consumption smoothing.
+Summarizing daily budget information over a longer time span is necessary because households need time to smooth their consumption given their realized income.
+Figure 3a shows that the head’s income is paid twice a month, typically in the middle (14th–16th) and at the end of the month (30th or 31st).
+[^28] The head’s income is the primary source of earnings, as most households are typical breadwinner households (Panel B in Table 3).
+Importantly, Figure 3b indicates that households spend most of their income immediately on payday and gradually use the remainder until the next payday.
+In other words, expenditure does not increase sharply just before paydays.
+This suggests that households plan their consumption schedules based on the income they received.
+Calendar-based (semi-)monthly aggregation cannot capture this behavior because it does not align the payday with the start of the aggregation period.
+[^29] For example, a calendar semi-month cell in January (1st–14th) starts on the 1st, but the payday falls on the 14th.
+Aggregating in this way assumes that the consumption within the cell is based on income at the end of the cell, which is unrealistic.
+Another source of measurement error arises when the payday shifts: occasionally, the payday for the latter half of a month falls on the first day of the next month.
+While infrequent, this can distort calendar-month aggregation, which may then include three paydays in some months and only one in others.
+Such mismatches between income and expenditure timing can induce Type II errors when testing for consumption smoothing (Nelson 1994).
 Adjustment
-To manage the mismatch, I aggregated the daily observations between the 12th and 26th for the first half and those between the 27th and 11th of the next month for the second half.
-[^40] This means that the number of days included in a semi-month ranges from 13 (i.e., in February) to 16 (e.g., January).
-Note that this heterogeneity in the number of days over different semi-months is a cross-sectional constant.
-Thus, these differences are entirely captured using the semi-month fixed effects in the regression.
-This range is based on the fact that paydays in the second half of a month measured in the THBS dataset ranged from 27th to 31st.
-[^41] Note that I do not use the definition of semi-month starting from the 2nd (to the 16th) in each month because this method of stratification places the payday at the end of each semi-month cell.
-Online Appendix B.7 provides the evidence that while the definition using 2nd in each month as the threshold provides an improved series in income, the expenditure cannot chase the income correctly because the expenses measured under this definition are based on the income in the previous semi-month cell.
-My preferred definition was using the 12th of each month as this threshold offers a clean series of income and expenditure.
-Figure 4c shows the time-series plots of the income and expenditure from this adjusted semi-monthly panel dataset, whereas Figure 4e illustrates the unadjusted (calender) semi-monthly series using the 15th of each month as the threshold of the semi-month.
-Clearly, Figure 4e provides very rough measures of income and expenditure due to the measurement errors from the maladjustment of paydays.
-The most significant improvement in Figure 4c is that the expenditure can chase the income in each month, reflecting the households’ consumption strategies based on their income measured in the approximate exact timing.
-Given the improvements in both income and expenditure measures, the net income series shows a much smoother trend in Figure 4d, which is unavailable under the unadjusted series (Figure 4f).
-Online Appendix B.8 shows that these measurement errors in the unadjusted dataset strongly attenuate the estimates.
+To address the timing mismatch, I aggregate daily observations from the 12th to the 26th for the first half of each month and from the 27th to the 11th of the following month for the second half.
+[^30] This means that the number of days in a semi-month ranges from 13 (February) to 16 (January).
+This heterogeneity is constant across households, and differences are fully captured by semi-month fixed effects in the regression.
+This range is based on the observed paydays in the THBS dataset, which fall between the 27th and 31st for the second half of each month.
+[^31] I do not use a semi-month definition starting on the 2nd (2nd–16th) because it places the payday at the end of the semi-month cell.
+Online Appendix B.7 shows that although this alternative improves the income series, expenditure cannot correctly follow the income, as expenses measured under this definition are based on income from the previous semi-month.
+Figure 3c presents the time-series plots of income and expenditure using this adjusted semi-monthly panel, whereas Figure 3e shows the unadjusted (calendar-based) semi-monthly series, using the 15th of each month as the threshold.
+The unadjusted series clearly provides a rough measure of income and expenditure owing to measurement errors from misaligned paydays.
+The main improvement in Figure 3c is that expenditure now aligns with income in each month, capturing households’ consumption strategies based on the timing of actual income.
+As a result, the net income series shows a much smoother trend in Figure 3d, which is not evident in the unadjusted series (Figure 3f).
+Online Appendix B.8 demonstrates that measurement errors in the unadjusted dataset strongly attenuate the estimates.
 3.4 Data Description
 Variables
-The THBS budget book has two primary sections: expenditure and income.
-I digitized records in each category following the categorization used in the official report.
-Panel A of Table 3 lists the $11$ subcategories for expenditures: food, housing, utilities, furniture, clothes, education, medical, entertainment, transportation, gifts, and miscellaneous.
-These are used to test the smoothness of the consumptions in each subcategory.
-To test the roles of risk-coping strategies, I considered five net income categories—savings, insurance, borrowing, credit purchase, and gifts.
-Net variables are defined as the income minus expenditure in each category; for example, net savings refers to the difference between the withdrawal amount and deposits to savings.
-I also considered the sales of miscellaneous goods and labor earnings from the family members (except for the head) to test the contribution of the sales of small assets and intrahousehold labor-supply adjustments.
+The THBS budget book has two main sections: expenditure and income.
+Panel A of Table 3 lists the $11$ expenditure subcategories: food, housing, utilities, furniture, clothes, education, medical, entertainment, transportation, gifts, and miscellaneous.
+These subcategories are used to test the smoothness of the consumptions.
+To examine the role of risk-coping strategies, I consider five net income categories: savings, insurance, borrowing, credit purchases, and gifts.
+Net variables are defined as income minus expenditure in each category; for example, net savings refer to withdrawals minus deposits to savings.
+I also include the sales of miscellaneous goods and family members’ (excluding the household head’s) labor earnings to assess the contribution of small asset sales and intrahousehold labor supply adjustments.
 Panel B of Table 3 summarizes these variables.
-[^42] Online Appendix B.11 presents the results for the panel unit root tests.
-The null hypothesis of unit roots in all the panels is rejected at the conventional level for all these variables.
-Alterations in family size may shift the household’s preferences (Jappelli and Pistaferri 2017).
-[^43] In village economies, landholding and asset stocks such as livestock and grain may also be correlated with the household’s risk-coping strategies (e.g., Udry 1994, 1995).
-As this study analyzes urban factory workers’ households, the roles of these assets can be discarded.
-Fortunately, the archives of the Tsukishima Survey include a set of sheets documenting the changes in the composition of the THBS households (Kazoku idō hyō).
-[^44] The Ohara Institute for Social Research, Archives of the Tsukishima Survey (THBS, unreleased).
-I found that the size of households was stable over time: one household experienced changes in the number of family members due to the birth of a son during the sample periods.
-This supports the evidence that the preference shifts rarely occurred among the THBS households.
-Although family size was included in all regressions, to be conservative, it had little effect on the results.
-[^45] The family size variable was omitted in several regressions due to the multicollinearity in the fixed-effects models given its small within variations.
-I then used the share of children below primary school age range as an alternative time-varying preference shock variable.
-Including the child share control variable does not influence the results, and the estimates are materially similar to those from my baseline specification (Online Appendix Table B.6).
-Panel C of Table 3 shows the summary statistics on the family size variable.
+[^32] Online Appendix B.11 presents the results of panel unit root tests.
+The null hypothesis of unit roots is rejected at conventional levels for all these variables.
+Changes in family size may affect household preferences (Jappelli and Pistaferri 2017).
+[^33] In village economies, landholding and assets such as livestock and grain may also correlate with households’ risk-coping strategies (e.g., Udry 1994, 1995).
+As this study focuses on urban factory-worker households, the roles of these assets are not relevant.
+Fortunately, the archives include sheets documenting changes in the composition of THBS households (Kazoku idō hyō).
+[^34] OISR, Archives of the Tsukishima Survey (THBS, unreleased).
+I find that household size remains stable over time.
+During the sample period, only one household experienced a change in family size due to the birth of a son.
+This indicates that preference shifts rarely occurred among THBS households.
+Although family size is included in all regressions, it has minimal effect on the results.
+[^35] The family size variable is omitted in some regressions owing to multicollinearity in fixed-effects models, given its small within-household variation.
+As an alternative time-varying measure of preference shocks, I use the share of children below primary school age.
+Including this child-share variable does not affect the results; estimates are materially similar to the baseline specification (Online Appendix Table B.6).
+Panel C of Table 3 presents the summary statistics for family size.
 Table 3: Summary Statistics
 Panel A: Variables for testing consumption smoothing (in yen)
-Semi-monthly Panel Dataset Adj. Monthly Panel Dataset
+Semi-monthly panel dataset Monthly panel dataset
 Level Log Level Log
@@
 Panel B: Variables for testing risk-coping mechanisms (in yen)
-Semi-monthly Panel Dataset Adj. Monthly Panel Dataset
+Semi-monthly panel dataset Monthly panel dataset
 Mean Std. Dev. Obs. Mean Std. Dev. Obs.
-Panel B-1: Savings, insurance, borrowing, and gifts
+Panel B-1: Net savings, insurance, borrowing, and gifts (income - expenditure)
     Net savings -0.01 3.97 289 0.08 5.78 124
@@
 Panel B-2: Labor supply adjustments & sales of miscellaneous assets
-      Other family member’s earning 3.74 6.11 289 8.21 16.44 124
+    Other family members’ earnings 3.74 6.11 289 8.21 16.44 124
     Sales of miscellaneous assets 0.02 0.12 289 0.06 0.17 124
@@
 (other family members)
-Notes:
-Panel A: This panel shows the summary statistics of the variables used to test the consumption smoothing for the 33- and 26-panel units in the semi-monthly and adjusted monthly panel datasets, respectively.
-Disposable income is the total income minus tax payments.
-Panel B: This panel shows the summary statistics of the variables used to test the risk-coping strategies for the 33- and 26-panel units in the semi-monthly and adjusted monthly panel datasets, respectively.
-Each net income variable is defined as the difference between income and expenses.
-“Net savings” refers to the difference between the amount withdrawn and the total deposits.
-“Net insurance” refers to the amount received from insurance minus the expenses paid for insurance.
-“Net borrowing” refers to the amount of borrowing minus the debt payments.
-“Net credit purchase” refers to the amount of credit purchases minus credit redeemed.
-“Net gifts” refers to the total received amounts, including both pecuniary and non-pecuniary gifts, minus the payments for gifts.
-“Other family member’s earnings” refers to the total income earned by all family members except for the head of the household.
-“Sales of miscellaneous assets” includes the sale of miscellaneous goods such as old newspapers and empty bottles.
-Panel C: The summary statistics of the family size variable for the 33-panel units in the semi-monthly panel dataset are listed.
-The statistics for the adjusted monthly panel dataset are not reported because they are materially similar.
-Panel D: The summary statistics of the number of days with illness for family members among the 33-panel units in the semi-monthly panel dataset are listed.
-As those are lagged variables, the initial period is omitted in all the units.
-Accordingly, a household with two semi-month observations is excluded.
+Notes: Panel A: This panel presents summary statistics for the variables used to test consumption smoothing for the 33- and 26-panel units in the adjusted semi-monthly and adjusted monthly panel datasets, respectively.
+Disposable income is defined as total income minus tax payments.
+Panel B: This panel presents summary statistics for the variables used to examine risk-coping strategies for the 33- and 26-panel units in the adjusted semi-monthly and adjusted monthly panel datasets, respectively.
+Each net income variable is defined as income minus expenditure.
+“Net savings”: withdrawals minus deposits to savings.
+“Net insurance”: insurance payments received minus expenses paid for insurance.
+“Net borrowing”: the amount borrowed minus debt repayments.
+This category includes net income from both pawnshops and other lending institutions.
+“Net credit purchase”: credit purchases minus credit repayments.
+“Net gifts”: total gifts received, including both pecuniary and non-pecuniary, minus any payments for gifts.
+“Other family members’ earnings”: total income earned by all family members except the household head.
+“Sales of miscellaneous assets”: include the sale of goods such as old newspapers and empty bottles.
+Panel C: This panel shows summary statistics for the family size variable for the 33-panel units in the adjusted semi-monthly panel dataset.
+Statistics for the adjusted monthly panel dataset are not reported because they are materially similar.
+Panel D: This panel shows summary statistics for the number of days family members were ill among the 33-panel units in the adjusted semi-monthly panel dataset.
+Since these are lagged variables, the initial period is omitted for all units.
+Accordingly, households with only two semi-month observations are excluded.
 Source: Created by the author from the THBS sample.
 Trend
-Next, I illustrate the time-series figures of the THBS dataset to overview the trend in income and expenditure.
-Figure 4c shows a weak increasing trend in the average income, which is consistent with the overall macroeconomic trend in nominal income during this period.
-This trend in earnings was, however, offset by steep inflation caused by expansions after the war (Section 2.1).
-In Figure 4d, the deviation of expenditure from income in January captures the expenses on the goods for the New Year events.
-The gaps in June to the first half of July may reflect the payments of bonuses.
-The peaks in March may reflect the preparation for the new fiscal and academic year from April, suggesting a higher proportion of temporary income from withdrawals and loans.
-These show that the budget data from the THBS households can capture a substantial aspect of the macroeconomic trends.
+Figure 3c shows a slight upward trend in average income, consistent with the overall macroeconomic increase in nominal income during this period.
+However, this growth in earnings is offset by steep inflation following postwar expansions (Section 2.1).
+In Figure 3d, the deviation of expenditure from income in January reflects spending on goods for New Year events.
+The gaps from June to the first half of July likely correspond to bonus payments.
+Peaks in March may indicate preparation for the new fiscal and academic year starting in April, suggesting a higher share of temporary income from withdrawals and loans.
+These patterns demonstrate that the THBS household budget data capture key aspects of macroeconomic trends.
 Idiosyncratic Shocks
-Online Appendix Figure B.11 illustrates the residuals from the regression in the first difference in semi-monthly income on the first difference in aggregate semi-monthly income.
-[^46] Let $n_{t}$ be the number of households in semi-month $t$ in the year 1919.
-For household $i$ in semi-month $t$ cell, the residual is then defined as $\Delta\hat{\upsilon}_{i,t}=\Delta\textit{Income}_{i,t}-\hat{\alpha}\Delta\overline{\textit{Income}}_{.,t}$, where $\overline{\textit{Income}}_{.,t}=\sum_{i=1}^{n_{t}}\frac{\textit{Income}_{i,t}}{n_{t}}$ and $\hat{\alpha}$ is the estimated coefficient.
-While the residuals satisfy the zero-average property, they occasionally take large negative and positive values, suggesting that the workers experienced idiosyncratic income shocks in some cells.
-Regardless of favorable or adverse shocks, these cells that deviated from the mean tendency help investigate how those households coped with the idiosyncratic shocks.
-Despite this, it is useful to consider illness as it is the least predictable shock (Gertler and Gruber 2002).
-Fortunately, the illness among the THBS families is documented systematically and thus, available for analysis.
-As explained in Section 2.3, neither a comprehensive health insurance system nor compensation for illness from the factories was available.
-Moreover, the sanitary conditions in the cities were low, leading to higher mortality rates.
-[^47] See Schneider and Ogasawara (2018) for the overviews on the sanitary conditions in interwar Japan.
-The official report mentioned that the workers in the machinery factories in Tsukishima faced the risk of diseases under severe sanitary conditions (Department of Health, Ministry of the Interior 1923a, p.
-14).
-Although unemployment might be another shock, the THBS heads rarely changed their occupation during the survey period as they were skilled workers who were less likely to move to other jobs.
-[^48] Another possibility is that the THBS heads who changed their jobs might have been omitted from the survey.
-However, I confirmed that the households with longer and shorter time-series observations had similar family characteristics (Online Appendix B.3).
-Thus, this omission does not lead to selection bias in terms of the potential vulnerabilities to the idiosyncratic risk.
-Moreover, the number of dismissals at the Tsukishima factories would be small.
-In fact, the average monthly dismissal rate (i.e., the number of dismissals/workers) from June 1919 to May 1920 at the one factory available from the official report of the Tsukishima Survey was only $1.5$% (Department of Health, Ministry of the Interior 1923a, p.
+Online Appendix Figure B.11 illustrates the residuals from the regression of the first difference in semi-monthly income on the first difference in aggregate semi-monthly income.
+[^36] Let $n_{t}$ be the number of households in semi-month $t$ in 1919.
+For household $i$ in semi-month $t$, the residual is defined as $\Delta\hat{\upsilon}_{i,t}=\Delta\textit{Income}_{i,t}-\hat{\alpha}\Delta\overline{\textit{Income}}_{.,t}$, where $\overline{\textit{Income}}_{.,t}=\sum_{i=1}^{n_{t}}\frac{\textit{Income}_{i,t}}{n_{t}}$ and $\hat{\alpha}$ is the estimated coefficient.
+Although the residuals satisfy the zero-average property, they occasionally take large positive or negative values, indicating that workers experienced idiosyncratic income shocks in some cells.
+These deviations from the mean help examine household responses to idiosyncratic shocks.
+Fortunately, the THBS archives (Kazoku idō hyō) document the exact timing and duration of family members’ illnesses.
+To understand the mechanisms behind consumption smoothing, I treat illness as the least predictable adverse shock (Gertler and Gruber 2002).
+While unemployment could also act as a shock, THBS household heads rarely changed occupations during the survey period, as they were skilled workers unlikely to move to other jobs.
+[^37] Household heads who changed jobs may have been omitted from the survey.
+However, households with longer and shorter time-series observations exhibit similar family characteristics (Online Appendix B.3), so this omission does not create selection bias regarding vulnerability to idiosyncratic risk.
+Moreover, dismissals at the Tsukishima factories were rare.
+For example, the average monthly dismissal rate from June 1919 to May 1920 at the one factory with available data was 1.51.5% (Department of Health, Ministry of the Interior 1923a, p.
 415).
-Panel D of Table 3 shows the summary statistics on the health shock variables.
+Panel D of Table 3 reports summary statistics for the health shock variables.
 Risk Preferences
-The different attitudes toward risks among households with different occupations and regions may lead to bias in the standard risk-sharing regressions (Schulhofer-Wohl 2011; Mazzocco and Saini 2012).
-In contrast, this study investigates consumption-smoothing strategies among homogeneous households — skilled factory workers’ households in the machinery manufacturing sector in Tsukishima.
-To assess the homogeneity of the risk preference among THBS households, I tested whether the earnings of the heads who worked in a few specific factories or sectors showed different responses to the aggregate shocks, following Schulhofer-Wohl (2011).
-Specifically, I consider the case that the heads who worked in large-scale or governmental factories might have had different risk preferences because those workers might have had lower risk tolerance than those who worked in smaller enterprises.
-I confirmed that the sensitivities to the aggregate shocks are statistically similar.
-[^49] Online Appendix B.12 summarizes the finer details of the analysis.
-In short, the heads’ earnings are regressed on the aggregate consumption, its interaction term with respect to the indicator variable for the large-scale and governmental factories, family size variable, and household fixed effects.
-The results are unchanged if I trim a few semi-months with a smaller number of cross-sectional observations.
-This result is unchanged if I assume that the heads who worked in the smithing factory or in the non-machinery manufacturing sectors had different risk preferences.
-[^50] The skilled workers in the smithing sector could be business owners (Online Appendix A.6).
-Similarly, the workers in the non-machinery manufacturing sectors in Tsukishima—a representative machinery production area—could have a different, presumably higher, risk tolerance.
-This shows that it is plausible to assume that the THBS households shared similar risk preferences, conditional on the household fixed effects.
+Different attitudes toward risk across households with varying occupations and in diverse regions can bias standard risk-sharing regressions (Schulhofer-Wohl 2011; Mazzocco and Saini 2012).
+However, this study focuses on consumption-smoothing strategies among homogeneous households—skilled factory workers in the machinery manufacturing sector in Tsukishima.
+To assess the homogeneity of risk preferences among THBS households, I test whether the earnings of household heads in specific factories or sectors responded differently to aggregate shocks, following Schulhofer-Wohl (2011).
+The results show that the sensitivities to aggregate shocks were statistically similar.
+[^38] The heads’ earnings are regressed on aggregate consumption, its interaction with an indicator for large-scale or governmental factories (or for smithing and non-machinery factories), the family size variable, and household fixed effects.
+The results remain unchanged when a few semi-months with smaller cross-sectional samples are trimmed.
+THBS households shared similar risk preferences, conditional on household fixed effects.
+Online Appendix B.12 provides additional details of the analysis.
 4 Empirical Analysis
@@
 Estimation Strategy
-I consider a linear fixed-effects model to test consumption smoothing.
-[^51] Another empirical specification used to test consumption smoothing is the first-difference model, including the aggregate consumption to capture macroeconomic shocks in the economy (Mace 1991).
-This alternative model requires the interpretation of the estimated coefficient on the aggregate measure of consumption.
-However, the THBS households do not include entire households in the target economy—Tsukishima.
-This may disturb the interpretation of the coefficient on aggregate consumption.
-To be conservative, therefore, I use the two-way fixed-effects model to control for macroeconomic shocks (Cochrane 1991; Ravallion and Chaudhuri 1997).
-Online Appendix C.1 provides the conceptual framework and the derivation of my empirical specification.
-For household $i$ in time $t$, the reduced form equation is characterized as follows:
+I use a linear fixed-effects model to test consumption smoothing.
+Online Appendix C.1 presents the conceptual framework and the derivation of my empirical specification.
+For household $i$ at time $t$, the reduced form equation is
 $\log c_{i,t}=\gamma\log y_{i,t}+\delta x_{i,t}+\mu_{i}+\phi_{t}+u_{i,t},$ (1)
-where $c_{i,t}$ is consumption, $y_{i,t}$ is disposable income, $x_{i,t}$ is the family size control, $\mu_{i}$ is the household fixed effect, $\phi_{t}$ is the semi-month fixed effect, and $u_{i,t}$ is a random error term.
-The household fixed effect captures the time-constant unobservable factors such as permanent income and consumption preference.
-The time-fixed effect effectively removes the unobservable macroeconomic shocks and trends described in Section 3.4.
-The estimate of $\gamma$ shows income elasticity, ranging from zero for perfect insurance to one for the absence of insurance.
-[^52] This is comparable to the estimate from the same conceptual framework: this comparability is important in the consumption-smoothing literature.
-Note that relevant previous studies have not integrated the recent theoretical discussion on the fixed effect model with two-way error components under the difference-in-differences (TWFE-DID) setting yet.
-In fact, the bias-corrected estimator for the TWFE-DID model with continuous treatment for multiple periods is underdeveloped (see de Chaisemartin and D’Haultfœuille (2023) for a comprehensive review).
-The cluster-robust variance-covariance matrix estimator is used to manage heteroskedasticity and serial dependency (Arellano 1987).
+where $c_{i,t}$ is consumption, $y_{i,t}$ is disposable income, $x_{i,t}$ is the family size control, $\mu_{i}$ is the household fixed effect, $\phi_{t}$ is the time fixed effect, and $u_{i,t}$ is a random error term.
+The household fixed effect captures time-invariant unobservables such as permanent income and consumption preferences.
+The time fixed effect accounts for unobservable macroeconomic shocks and trends described in Section 3.4.
+The estimate of $\gamma$ represents income elasticity, ranging from zero under perfect insurance to one in the absence of insurance.
+Since identification relies on both adverse and favorable idiosyncratic income shocks, the estimated elasticity reflects the overall response to these shocks.
+A cluster-robust variance-covariance matrix estimator is used to address heteroskedasticity and serial correlation (Arellano 1987).
+Another specification commonly used in the literature is the first-difference model, which uses aggregate consumption to capture macroeconomic shocks (Mace 1991).
+However, the aggregate measure may not accurately reflect macroeconomic shocks, particularly when the number of households is relatively small.
+To be conservative, I therefore employ the two-way fixed-effects model, which systematically controls for macroeconomic shocks through the time fixed effect (Cochrane 1991; Ravallion and Chaudhuri 1997).
 Results
 Table 4: Results of Estimating Income Elasticities
-(1) Semi-monthly Panels (2) Adjusted Monthly Panels
+(1) Semi-monthly panels (2) Monthly panels
 Disposable income Obs. Disposable income Obs.
@@
 ***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors in brackets are clustered at the household level.
-Notes: This table shows the results of equation 1.
-The regressions of the 11 measures of log-transformed consumption on log-transformed disposable income, the family size control, household-fixed effect, and time-fixed effect are listed.
-Columns 1 and 2 show the results for the semi-monthly and adjusted monthly panel datasets, respectively.
-The estimated coefficients on log-transformed disposable income are listed in the columns named “Coef.”.
-Column (1) of Table 4 presents the results from the semi-monthly panel dataset.
-In the total consumption, the estimated coefficient is $0.358$.
-This means that a one percentage point decrease in disposable income decreases the total expenditure by $0.36$ percentage point, suggesting that the factory worker households could not thoroughly smooth their consumption.
+Notes: This table presents the results of equation 1.
+It reports regressions of 11 measures of log-transformed consumption on log-transformed disposable income, the family size control, the household fixed effect, and the time fixed effect.
+Columns 1 and 2 show the results for the adjusted semi-monthly and adjusted monthly panel datasets, respectively.
+The estimated coefficients on log-transformed disposable income are shown in the columns labeled “Coef.” The estimated elasticities reflect the overall responses of households to both adverse and favorable idiosyncratic income shocks.
+Column 1 of Table 4 presents the results from the semi-monthly panel dataset.
+The estimated coefficient for total consumption indicates that a one percentage point decrease in disposable income reduces total expenditure by $0.358$ percentage points, suggesting that factory-worker households could not fully smooth their consumption.
 An important finding is that the estimate for food categories is smaller ($0.136$) but highly statistically significant.
-This implies that while the households tended to reduce their food expenses when they faced adverse income shocks, they might have had some means to manage those shocks to smooth food consumption.
-The estimates for the entertainment and transportation expenses ($0.307$ and $0.278$) are also highly statistically significant.
-This may be understandable because those expenses include family travel and theatergoing.
-[^53] The workers could walk to the factories in Tsukishima, so the costs of commuting were not usually documented in the budget books.
-Similarly, the estimate for gifts shows a large amount ($0.523$), reflecting that the expenses on the others were optional.
-The results for the other subcategories are statistically insignificant or unclear.
-The estimates for furniture and clothes are weakly statistically significant.
-Although not conclusive, this may reflect that both categories were luxury categories.
-The estimate for the education category is small but also weakly statistically significant ($0.101$).
-This may be because it includes an allowance for children, which could be easily decreased.
-[^54] Toyotaro Miyoshi served as a research assistant at the Tsukishima Survey while he was a student at the Tokyo Imperial University.
-He left important retrospectives on the survey in his works (Online Appendix B.1).
-According to his book, since there were no nurseries, children in these working-class households regularly received a small allowance to play outside while their parents were working (Miyoshi 1989, pp.
+This suggests that while households reduced food expenses in response to adverse income shocks, they still had some means to smooth food consumption.
+The estimates for entertainment and transportation expenses ($0.307$ and $0.278$) are also statistically significant.
+This is understandable because these expenses include family travel and theater visits.
+[^39] On this small island, workers could walk to the factories, so commuting costs were typically not recorded in the budget books.
+Similarly, the larger estimate for gifts ($0.523$) reflects that these expenses were discretionary.
+The results for other subcategories are statistically insignificant or ambiguous.
+Estimates for furniture and clothing are weakly statistically significant, which may reflect that both are luxury categories.
+The estimate for the education category is small but also weakly statistically significant ($0.101$), as it includes an allowance for children that can be easily reduced.
+[^40] Toyotaro Miyoshi, who served as a research assistant at the Tsukishima Survey, provided important retrospective accounts (Online Appendix B.1).
+According to his book, children in these factory-worker households regularly received a small allowance to play outside while their parents worked, since no nurseries were available (Miyoshi 1989, pp.
 383–384; 386; 387).
-I present the results for the adjusted monthly panel dataset in Column (2) to see whether the income elasticities changed over a longer span.
-Not surprisingly, the adjusted monthly series yields greater estimates than the semi-monthly series.
-The estimate for total consumption was $0.433$, which is approximately $1.2$ times greater than that from the semi-monthly series.
-One potential explanation for this difference is that the available sources of temporary income might have been limited, and those sources were not useful for compensating for adverse income shocks that lasted for at least one month.
-The most noticeable example is the food category, which comprised the largest portion of the expenditure (Panel A of Table 3).
-The estimate is $0.293$, more than twice the estimate for the semi-monthly series ($0.136$).
-This implies that the households’ food expenditures were likely to be insured for a few weeks but not over one month.
-I will investigate the mechanism underlying this behavior in the following subsections, and credit purchases may be used to smooth food consumption in the short run.
-Similarly, the expenses for entertainment and transportation were much more sensitive ($0.716$ and $0.605$) to the shocks in the adjusted monthly series.
-The estimate for gifts is slightly smaller but in a similar range.
-In short, while the household could compensate for their income losses for a few weeks, they could not do so for several weeks, presumably because of borrowing constraints and/or a limited amount of liquidity assets.
-The expenses on those unnecessary categories might particularly be cut in response to relatively persistent shocks.
-In contrast, a few subcategories were not sensitive to the idiosyncratic shocks over different time-series frequencies.
-For example, the estimates for the housing, utilities, and medical categories are not statistically significant in both semi-monthly and adjusted monthly series.
-This may reflect the fact that the rent and utility payments were relatively fixed.
-[^55] For example, rent payment was usually fixed and paid once per month.
-Online Appendix C.2 shows that the housing category should be analyzed based on the adjusted monthly series rather than the semi-monthly series.
-The medical expenses included the fee for bathing (yuya).
-People took baths every few days at that time, which was less likely to be cut, although hairdressing for women (kamiyui) in the same category was likely to be cut.
-The estimates for the furniture, clothes, and education categories are statistically insignificant in the adjusted monthly series.
-While this could partly be the efficiency issue under a smaller sample size, those expenses may be sensitive to unexpected shocks in consumption, such as unanticipated losses of household goods, clothes, and stationery, including writing paper.
-[^56] An alternative explanation for the clothes category relates to seasonality.
-The workers had bought their work clothes for the new season (e.g., the OISR, Archives of the Tsukishima Survey (THBS #40, March 31, 1919)).
-As those clothes were necessary items, those seasonal expenses may attenuate the estimate.
-This may systematically attenuate the estimates for the specific categories in the adjusted monthly series.
-[^57] Generally, the panels with short-run time bins were more likely to be influenced by unexpected consumption shocks, which attenuated income elasticity (Nelson 1994).
-If this mechanism worked, subcategories that were sensitive (insensitive) to unexpected consumption shocks should be more (less) attenuated.
-In my result, however, the estimates for the adjusted monthly panels are greater than those for the semi-monthly panels in a large part of subcategories, suggesting that households faced more difficulties with smoothing their consumption in the long run.
+Column 2 reports the results for the monthly panel dataset examining whether income elasticities change over a longer period.
+As expected, the monthly series yields larger estimates than the semi-monthly series.
+The estimate for total consumption is $0.433$, approximately $1.2$ times higher than that from the semi-monthly series.
+One possible explanation is that the sources of temporary income are limited and insufficient to offset adverse income shocks lasting at least one month.
+The most notable example is the food category, which shows more than twice the estimate ($0.293$) compared with the semi-monthly series.
+This implies that households’ food expenditures could be insured for a few weeks but not over a full month.
+Similarly, expenses for entertainment and transportation are much more sensitive ($0.716$ and $0.605$, respectively) to shocks in the monthly series.
+The estimate for gifts is slightly lower but remains in a similar range.
+Expenses on these discretionary categories are particularly responsive to relatively persistent shocks.
+By contrast, some subcategories are not sensitive to idiosyncratic shocks across different time-series frequencies.
+For example, estimates for housing, utilities, and medical expenses are not statistically significant in either the semi-monthly or monthly series.
+This may reflect the fact that rent and utility payments are relatively fixed.
+[^41] Rent is usually fixed and paid once per month.
+Online Appendix C.2 provides quantitative evidence that the housing category should be analyzed using the monthly series rather than the semi-monthly series.
+Medical expenses include fees for bathing (yuya).
+People bathe every few days, a habit unlikely to change, although the frequency with which women opt for hairdressing services(kamiyui) in the same category can reduce.
+The estimates for furniture, clothing, and education are statistically insignificant in the monthly series.
+While this may partly reflect efficiency issues, these expenses are generally insensitive to unexpected demand shocks such as unanticipated losses of household goods, clothing, or stationery, including writing paper.
+[^42] An alternative explanation for the clothing category relates to seasonality.
+Workers bought new work clothes for the season (OISR, Archives of the Tsukishima Survey (THBS #40, March 31, 1919)).
+As these clothes were necessary, seasonal purchases attenuate the estimates.
+This systematically reduces the estimates for specific categories in the monthly series.
+[^43] Panels with short-run time bins are more likely to be influenced by unexpected consumption shocks, which attenuate income elasticity (Nelson 1994).
+When this mechanism operates, subcategories sensitive (insensitive) to unexpected consumption shocks are more (less) attenuated.
+In my results, however, estimates for the monthly panels are higher than those for the semi-monthly panels for many subcategories, suggesting that households faced greater difficulties smoothing consumption over the longer term.
 4.2 Risk-coping Mechanisms
-The preceding results show that the skilled workers’ households in Tsukishima were less likely to smooth their consumption.
-The estimated elasticities, however, became smaller when I focused on the semi-monthly frequency.
-Specifically, food expenditure was likely to be insured within at least half a month.
-The workers might have coped with their shorter-run shocks relatively well but could not manage longer-run shocks.
-In this section, I test the mechanisms behind their smoothing behavior using several variables in the risk-coping strategies.
+The preceding results suggest that workers were able to cope with short-term shocks to some extent.
+In this section, I examine the mechanisms underlying their consumption-smoothing behavior using information on risk-coping strategies.
 Estimation Strategy
-To test the roles of risk-coping strategies, I employed the empirical specification suggested by Fafchamps and Lund (2003).
-For household $i$ at time $t$, the specification is characterized as follows:
+To test the roles of risk-coping strategies, I employ the empirical specification proposed by Fafchamps and Lund (2003). For household $i$ at time $t$, the specification is
 $r_{i,t}=\kappa\tilde{y}_{i,t}+\eta x_{i,t}+\nu_{i}+\zeta_{t}+\epsilon_{i,t},$ (2)
-where $\tilde{y}_{i,t}$ is head’s earning, $x_{i,t}$ is the family size control, $\nu_{i}$ is the household fixed effect, $\zeta_{t}$ is the time fixed effect, and $\epsilon_{i,t}$ is a random error term.
+where $\tilde{y}_{i,t}$ is the head’s earnings, $x_{i,t}$ is the family size control, $\nu_{i}$ is the household fixed effect, $\zeta_{t}$ is the time fixed effect, and $\epsilon_{i,t}$ is a random error term.
 Online Appendix C.1 summarizes the derivation of this specification.
-I used the head’s income instead of the disposable income in Equation 2 to avoid the potential endogeneity.
-[^58] For instance, the household’s total income is the outcome of the intra-household responses to the shocks, such as the labor-supply adjustments.
-I confirmed that the results were largely unchanged if I used the disposable income as the idiosyncratic income shock variable instead of the head’s income (not reported).
-This is consistent with the fact that the labor-supply adjustments did not occur among the THBS households (Table 5).
-Despite this, I preferred to use the conservative setting to manage potential biases and to set a comparable setting to the structural equation used in Section 4.3.
-See Asdrubali et al.
-(2020) for a recent study using the head’s income as the shock variable in testing the full risk-sharing.
-Regarding $r_{i,t}$, I considered several variables to overview various risk-coping strategies.
-These included the net income on savings, insurance, borrowing, credit purchases, and gifts.
-Each net variable was defined as the difference between income and expenses.
-The estimates of $\kappa$ were therefore expected to be negative since net income becomes negative in good times but positive in unfavorable times.
-I also considered the earnings of the family members (except for the head) and the sales of miscellaneous goods because both intra-household labor supply adjustments and asset sales might alleviate idiosyncratic shocks as alternative risk-mitigation strategies (Udry 1995; Blundell et al.
+I use the head’s income instead of total disposable income in Equation 2 to avoid potential endogeneity.
+[^44] For example, total household income is the outcome of intra-household responses to shocks, such as labor supply adjustments (Asdrubali et al.
+2020).
+The results are largely unchanged if disposable income is used instead of the head’s income (not reported).
+This is consistent with the absence of labor supply adjustments among THBS households (Table 5).
+Nevertheless, I use the conservative approach to reduce potential biases and maintain comparability with the structural equation in Section 4.3.
+Regarding $r_{i,t}$, I consider several variables to capture different risk-coping strategies, including net income from savings, insurance, borrowing, credit purchases, and gifts.
+Each net variable is defined as income minus expenses.
+The estimates of $\kappa$ are expected to be negative, since net income is typically negative in good times and positive in unfavorable times.
+I also consider earnings of other family members (excluding the head) and income from sales of miscellaneous goods (Udry 1995; Blundell et al.
 2016).
-[^59] Rosenzweig and Stark (1989) show that marriage cum migration in South Indian villages contributed to reducing the variability of household food consumption.
-Since the THBS focused on the households living in Tsukishima throughout the sample period, I did not need to address the ex-ante risk-mitigation strategy through migration.
-The estimated coefficients for the sales and labor supply adjustment variables may be negative, reflecting additional earnings when the head’s income is reduced.
+[^45] Rosenzweig and Stark (1989) show that marriage combined with migration in South Indian villages reduced variability in household food consumption.
+Since the THBS tracked households living in Tsukishima throughout the sample period, ex-ante risk mitigation via migration is not relevant here.
+Estimated coefficients for sales and labor supply adjustment variables may be negative, reflecting additional earnings when the head’s income declines.
 Results
 Table 5: Results of Testing the Risk-coping Mechanisms
-(1) Semi-monthly Panels (2) Adj. Monthly Panels
-Head’s earning Obs. Head’s earning Obs.
+(1) Semi-monthly panels (2) Monthly panels
+Head’s earnings Obs. Head’s earnings Obs.
 Coef. Std. error Coef. Std. error
-Panel A: Savings, insurance, borrowing, and gifts (yen)
+Panel A: Net savings, insurance, borrowing, and gifts (income - expenditure, yen)
 Net savings -0.043 [0.018]** 289 -0.035 [0.028] 124
@@
 Panel B: Labor supply adjustments & sales of miscellaneous assets (yen)
-Other member’s earnings 0.001 [0.010] 289 0.036 [0.033] 124
+Other members’ earnings 0.001 [0.010] 289 0.036 [0.033] 124
 Sales of miscellaneous assets -0.000 [0.000] 289 0.000 [0.001] 124
 ***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors in brackets are clustered at the household level.
-Notes:
-This table shows the results of equation 2.
-The family size control, household fixed effects, and time-fixed effects are included in all the regressions.
-Columns 1 and 2 show the estimates for the semi-monthly and adjusted monthly panel datasets, respectively.
-The estimated coefficients on the head’s earnings are listed in the “Coef.” columns.
-Panel A: Each net income variable is defined as the difference between income and expenses.
-“Net savings” refers to the difference between the amount withdrawn and the total deposits.
-“Net insurance” refers to the amount received from insurance minus the expenses paid for insurance.
-“Net borrowing” refers to the amount of borrowing minus the debt payments.
-“Net credit purchase” refers to the amount of credit purchases minus credit redeemed.
-“Net gifts” refers to the total received amounts, including both pecuniary and non-pecuniary gifts, minus the payments for gifts.
-Panel B: “Other member’s earnings” refers to the total income earned by all family members except for the head of the household.
-“Sales of miscellaneous assets” includes the sale of daily miscellaneous goods such as newspapers and empty bottles.
-All dependent variables are in yen.
+Notes: This table reports the results of equation 2.
+All regressions include the family size control, household fixed effects, and time fixed effects.
+Columns 1 and 2 present estimates for the adjusted semi-monthly and adjusted monthly panel datasets, respectively.
+The estimated coefficients on the head’s earnings are shown in the “Coef.” columns.
+Panel A: Each net income variable is defined as income minus expenditure.
+“Net savings”: withdrawals minus deposits to savings.
+“Net insurance”: insurance payouts minus insurance expenses.
+“Net borrowing”: borrowing minus debt repayments.
+This category includes net income from both pawnshops and other lending institutions.
+“Net credit purchase”: credit purchases minus credit repayments.
+“Net gifts”: the total received gifts, both pecuniary and non-pecuniary, minus payments for gifts.
+Panel B: “Other members’ earnings”: total income earned by all household members except the head.
+“Sales of miscellaneous assets”: include the sale of daily items such as newspapers and empty bottles.
+All dependent variables are measured in yen.
 Table 5 presents the results.
-Columns (1) and (2) summarize the results for the semi-monthly and adjusted monthly panel datasets, respectively.
-Panel A shows the results for the net income variables, whereas Panel B shows those for the labor supply adjustments and sales of small assets.
-The estimated coefficient for the net savings is negative and statistically significant in Column (1).
-The estimate suggests that a one-yen decrease (increase) in the head’s earnings increases (decreases) the net income from savings by $0.043$ yen.
-The estimate from the adjusted monthly series is similar ($0.035$), albeit statistically insignificant (Column (2)).
-This is consistent with the historical fact that many workers had savings accounts in postal savings and savings banks at that time (Section 2.3).
-However, the estimated magnitude was small, suggesting that the withdrawals from savings did not compensate for the earnings losses.
-The estimates for the insurance category are close to zero and statistically insignificant in both columns.
-The budget books indicate that the primary source in this category was the mutual loan associations (mujin).
-As explained in Section 2.2, these associations’ rules were unsuitable for compensating for the idiosyncratic shocks.
-For example, a THBS household won a lottery and received $100$ yen in a monthly cell, which is independent of the worker’s earnings.
-[^60] The OISR, Archives of the Tsukishima Survey (THBS #5, May 17, 1919).
-The bidding rule may be used to support a specific member in the association if members have a common belief in the aid.
-However, most of the income from mujin measured in the budget books was less than one yen, which is not the case on average.
-For the net borrowing category, the estimated coefficients are close to zero and statistically insignificant in both columns.
-Section 2.2 shows that pawnshops were a popular lending institution for low-income households.
-Although THBS households occasionally used pawnshops, this was not frequent.
-[^61] The OISR, Archives of the Tsukishima Survey (THBS #2; 19; 22).
-This is understandable because skilled factory workers were not classified as low-income working class at that time.
-The estimated coefficient for net income from credit purchases is negative and strongly statistically significant in Column (1).
-The estimate suggests that a one-yen decrease (increase) in the head’s earnings increases (decreases) the net income from credit purchases by $0.13$ yen.
-This magnitude is the largest in all the categories.
-According to the THBS budget books, the credit purchases were mainly used to buy rice, fish, vegetables, and fuels (firewood and charcoal).
-As rice—a staple food—is usually an expensive item, credit purchases might have played an important role among these households.
-In addition, credit purchase was occasionally used for the call doctor and pharmacy.
-This is consistent with my baseline result for medical expenses in the semi-monthly series (Column (1) in Table 4).
-While these results suggest that credit purchase was the primary risk-mitigation strategy, Column (2) shows that credit was no longer effective in the long term.
-This is reasonable because borrowers could make monthly or semi-monthly payments but did not borrow for more than a month (Section 5).
-Given the estimated magnitude ($0.13$), however, it is noteworthy that the credit could only partially compensate for the idiosyncratic shocks.
-The estimates for the gifts category are negative and statistically significant in Column (1).
-The budget books show that the gifts included seasonal gifts, celebration money, and return gifts.
-Among them, the return gifts can be regarded as mutual aid from personal networks, which may be reflected in this result.
-According to an official report of Tokyo, however, the share of average monthly gifts in total income was approximately $3$% (Social Affairs Division 1925, pp.
+Columns 1 and 2 summarize the findings for the semi-monthly and monthly panel datasets, respectively.
+Panel A reports the results for the net income variables, while Panel B reports those for labor supply adjustments and sales of small assets.
+The estimated coefficient for net savings is negative and statistically significant in Column 1.
+This suggests that a one yen decrease (increase) in the head’s earnings increased (decreased) net savings by $0.043$ yen.
+The estimate from the monthly series is similar but statistically insignificant (Column 2).
+While this reflects that many workers had savings accounts in postal savings and savings banks (Section 2.3), the magnitude was small, indicating that withdrawals from savings did not fully compensate for earnings losses.
+Estimates for the insurance category are near zero and statistically insignificant in both columns.
+The budget books indicate that the primary source was mutual loan associations (mujin).
+However, these associations’ rules were unsuitable for offsetting idiosyncratic shocks.
+For example, a THBS household won a lottery and received $100$ yen in a monthly cell, which was independent of the worker’s earnings.
+[^46] OISR, Archives of the Tsukishima Survey (THBS #5, May 17, 1919).
+The bidding rule could support a specific member if all members have a common belief in aid, but most income from mujin recorded in the budget books was less than one yen, which is not representative on average.
+For net borrowing, estimated coefficients are close to zero and statistically insignificant in both columns.
+Although THBS households occasionally used pawnshops, these instances were infrequent.
+[^47] OISR, Archives of the Tsukishima Survey (THBS #2; 19; 22).
+This is understandable because skilled factory workers were not classified as low-income, who were the primary users of pawnshops (Section 2.3).
+The estimated coefficient for net income from credit purchases is negative and strongly statistically significant in Column 1.
+This suggests that a one yen decrease (increase) in the head’s earnings increased (decreased) net income from credit purchases by $0.13$ yen.
+This magnitude is the largest among all categories.
+According to the budget books, credit purchases were primarily used to buy rice, fish, vegetables, and fuel (firewood and charcoal).
+As rice—a staple food—was usually expensive, credit purchases played an important role for these households.
+Credit purchases were also occasionally used for medical expenses such as doctor visits and pharmacy purchases, consistent with the baseline results for medical expenses (Column 1 in Table 4).
+While these results indicate that credit purchases were the primary risk-mitigation strategy in the short term, Column 2 shows that credit ceases to be effective over the longer term.
+This is reasonable, as borrowers typically made monthly or semi-monthly payments and did not borrow for periods longer than a month (Section 5).
+Estimates for the gifts category are negative and statistically significant in Column 1.
+The budget books show that the gifts include seasonal gifts, celebration money, and return gifts.
+Most of these gifts were not aid from kinship networks, including rural relatives, but rather monetary and material exchanges with acquaintances based on a culture of gift-giving.
+According to an official report from Tokyo City, average monthly gifts accounted for approximately 33% of total income (Social Affairs Division 1925, pp.
 58-59).
-In my adjusted monthly panels, this share is calculated to be approximately $4$%.
-In this light, while the gifts comprised a few percentage points of the monthly income, they may not have provided sufficient aid for the households facing economic hardship.
-In fact, the estimate is no longer statistically significant in the adjusted monthly series (Column (2)).
-Finally, Panel B tests the roles of the intra-household labor supply adjustment and sales of miscellaneous assets.
+Similarly, this share is approximately 44% in my monthly panel dataset.
+Even if mutual aid from personal networks is included in this small figure, it likely provided insufficient support for households facing economic hardship.
+Accordingly, the estimate ceases to be statistically significant in the monthly series (Column 2) and becomes negligible when focusing on the impact of adverse income shocks (Section 4.3).
+Finally, Panel B tests the roles of intra-household labor supply adjustments and sales of miscellaneous assets.
 The estimated coefficients in both columns of the first row are close to zero and statistically insignificant.
 Miyoshi (1989, pp.
-383–386) documented that the wives regularly engaged in side jobs rather than working temporarily to compensate for the income loss.
-In fact, although sewing was the most popular side job among THBS wives, their earnings were too small to replace the lost income from their breadwinning husbands.
-[^62] The OISR, Archives of the Tsukishima Survey (THBS#1, January 13, 1919).
-In the second row, I show that selling miscellaneous assets such as newspapers, bottles, and clocks was not an effective risk-mitigation strategy.
-To summarize, the evidence suggests that credit purchases mitigated idiosyncratic shocks.
-The results also indicate that the savings institutions and personal networks might have been used as risk-coping institutions among the workers.
-However, these were available for the short term —at most half a month.
-The estimated magnitudes indicate that the workers could only partially compensate for the temporary losses of income by using these institutions.
-This supports the findings on partial risk-sharing in Section 4.1.
-Overall, the results are aligned with the descriptives from the historical documents on the financial institutions available in Tokyo City (Section 2.3).
-An interesting finding—which is not explicitly described in the historical sources—is that the credit purchases in the local society and the informal gifts from personal networks might have provided some aid to the factory workers.
+383–386) documented that wives regularly engaged in side jobs such as sewing, rather than taking temporary work to compensate for income losses.
+However, their earnings were too small to replace the lost income from their breadwinning husbands.
+[^48] OISR, Archives of the Tsukishima Survey (THBS#1, January 13, 1919).
+In the second row, sales of miscellaneous assets such as newspapers, bottles, and clocks are shown to be ineffective as a risk-mitigation strategy.
+Heterogeneity
+The residential area of Tsukishima spanned less than $1.5$ kilometers, with a width of $500$ meters (Chūō Ward 1994).
+Within this narrow area, workers had easy access to shops in different lots (Department of Health, Ministry of the Interior 1923a, pp.
+40–46).
+Nevertheless, I conduct several additional tests to explore heterogeneous responses using maps of retailers and pawnshops documented in the report.
+First, I calculate the density of retail shops for daily necessities in each lot and include it in the regression for credit purchases.
+The estimate from this expanded regression is largely consistent with my baseline estimate (Column 1 in Panel A of Table 5).
+As an alternative test, I add an interaction term between the head’s earnings and the density of retailers into Equation 2.
+The interaction effect is approximately zero and remains unchanged when using an alternative definition of density.
+These results indicate that shop accessibility was homogeneous for the workers.
+Online Appendix C.3 provides detailed summaries of these analyses.
+Second, I examine the role of local pawnshops separately from the net borrowing regression in Table 5.
+In European countries and the U.S., religious communities played an important role in providing social insurance (Ferrara and Testa 2023; Ager and Ciccone 2018).
+Although some philanthropists assisted the poor before government-managed social welfare became available, few social facilities run by religious communities existed in Japan (Yoshida 2004).
+Unlike in European countries (Murhem 2016), pawnshops in Japan were private, profit-oriented enterprises until public pawnshops with social functions were established in 1927 under the Public Pawnbroking Act.
+Despite this, local pawnshops may have had informal ties with workers, as they served the same residents of Tsukishima.
+When I use temporary income from pawnshops as the dependent variable in Equation 2, the estimated coefficient on the head’s earnings is close to zero and statistically insignificant.
+[^49] The cooperative society is another local risk-coping institution, though it is rarely used (Section 2.3).
+Using total purchases from the cooperatives as a dependent variable in Equation 2 yields a coefficient on the head’s earnings that is nearly zero and statistically insignificant (Online Appendix C.4).
+This result remains unchanged after including the density of pawnshops in each lot in the same regression.
+I then estimate an expanded regression of Equation 2 for temporary income from pawnshops, which includes an interaction term between the head’s earnings and the density of pawnshops.
+Although the interaction effect is negative, it is small and statistically insignificant.
+A similar result is obtained when using an alternative definition of the density variable.
+These findings suggest that the use of pawnshops was independent of idiosyncratic shocks.
+Online Appendix C.4 summarizes the results.
 4.3 Response to Adverse Health Shocks
-The foregoing analyses used the within variations in the earnings to estimate income elasticity, meaning that the identification depends both on the adverse and favorable idiosyncratic shocks.
-While it is useful to illustrate the overall responses of the households to the shocks, it does not reveal the pathways behind the risk-coping strategies.
-Next, I focus on temporary illness of household heads that could cause temporary income losses.
+The identification in the analyses above relies on both adverse and favorable idiosyncratic shocks.
+To focus on the mechanisms behind consumption smoothing in response to adverse shocks, I use short-term income losses due to the temporary illnesses of household heads for the identification in this section.
 Estimation Strategy
 Figure 4: Number of Influenza Deaths in Tokyo between 1918 and 1920
-Notes: This figure shows the time-series plots of the monthly death counts from influenza (ryūkōsei kanbō) in Tokyo prefecture between January 1918 and December 1920.
-The death statistics used in this figure were collected based on the comprehensive registration system.
-Source: Created by the author using the 1918, 1919, and 1920 editions of the Statistics of Causes of Deaths of the Empire Japan (Statistics Bureau of the Cabinet 1921; 1922; 1923).
-I began my analysis by testing whether the illness of the head impacted the earnings.
-To do so, I decided the exact timing and duration of illnesses of the family members using the Kazoku idō hyō (Section 3.4).
-Figure 5 shows the number of monthly deaths from influenza (ryūkōsei kanbō) in Tokyo between 1918 and 1920.
-There appears to have been a moderate spike in early 1919, which is partially captured in the THBS period.
-In fact, a few THBS budget books indicate that the family members contracted influenza during this period.
-Five households experienced the illness during the sample period, and illness affecting the heads was concentrated between February and May 1919, when influenza exposure was most prevalent.
-[^63] Among four cells with health shocks on the heads, two are observed in February, during the peak of the epidemics (Figure 5), and the others are in March and May.
-In addition, the average duration for the heads is $5.8$ days (Std.
+Notes: This figure shows time-series plots of the monthly death counts from influenza (ryūkōsei kanbō) in Tokyo prefecture from January 1918 to December 1920.
+The death statistics are collected through the comprehensive registration system.
+Source: Created by the author using the 1918, 1919, and 1920 editions of the Statistics of Causes of Deaths of the Empire of Japan (Statistics Bureau of the Cabinet 1921; 1922; 1923).
+I begin my analysis by testing whether the head’s illness affected earnings.
+Figure 4 shows the monthly death counts from influenza (ryūkōsei kanbō) in Tokyo between 1918 and 1920.
+Early 1919 saw a moderate spike, partially captured in the THBS period.
+Indeed, a few THBS budget books indicate that family members contracted influenza during this time.
+According to the Kazoku idō hyō (Section 3.4), five households experienced illness during the sample period, with illness affecting the heads concentrated between February and May 1919, when influenza exposure was most prevalent.
+The average duration of illness for the heads was $5.8$ days (Std.
 Dev.
-$=3.6$), which is consistent with the fact that Type A influenza caused fever for less than six days (Uyeki et al.
+=3.6=3.6), consistent with Type A influenza, which causes fever lasting less than six days (Uyeki et al.
 2022).
-Therefore, it is plausible that influenza was responsible for illnesses during this period, providing a quasi-experimental setting for analyzing to what extent the adverse idiosyncratic shock of illness influences the households’ consumption strategies.
-Earnings in a semi-month are based on the work done during the previous semi-month.
-This means that the head’s earnings were influenced by the number of illness days in the previous semi-month cell.
-Table 6 presents the results from the regression of the head’s earnings on the number of illness days in the previous semi-monthly cell.
-While the instantaneous health shock has little impact on the head’s earnings (Column 1), the lagged health shock has a statistically significant effect on their earnings (Column 2).
-This result is unchanged if I include both variables in the same regression in Column 3, supporting the evidence that the health shocks in the previous cell reduced the earnings of the head.
-The estimate in Column (2) suggests that a one-day increase in illness decreases the head’s semi-monthly income by approximately $2.4$ yen.
-Notably, this is close to the average daily wage rate of skilled factory workers (Online Appendix B.6).
-This reflects the reductions in their salary due to absence: Miyoshi (1989, pp.
-383; 388) highlighted that income was reduced due to absence when the head was infected with a disease.
-Finally, Column (4) adds the health shocks of the other family members into the specification of Column (3).
-The estimated coefficients on these placebo variables are statistically insignificant, supporting the robustness of the result in Column (3).
-These results confirm that the head who suffered illness lost earnings.
-[^64] Given that the average number of illness days was approximately $2.0$, the temporary health shock losses were more than five yen in that month, accounting for approximately $10$% of the average monthly income.
-Although this is a rough estimate and the losses depend on the severity of the shocks, my result shows that the household income responded to the adverse health shock of illness among the factory workers’ households.
+Therefore, influenza likely caused the illnesses during this period.
+Earnings in a semi-month are based on work performed during the previous semi-month.
+Table 6 presents the regression results of the head’s earnings on the number of illness days in the previous semi-monthly cell.
+While instantaneous health shocks have little impact on earnings (Column 1), lagged health shocks have a statistically significant effect (Column 2).
+This result remains unchanged when both variables are included in the same regression (Column 3), supporting the evidence that health shocks in the previous cell reduce the head’s earnings.
+The estimate in Column 2 suggests that a one day increase in illness reduces the head’s semi-monthly income by approximately $2.4$ yen.
+Notably, this is close to the average daily wage of skilled factory workers, reflecting salary reductions due to absence (Online Appendix B.6).
+This aligns with anecdotal evidence that the head’s illness directly caused income loss due to absence (Miyoshi 1989).
+Finally, Column 4 adds the health shocks of other family members to the specification in Column 3.
+The estimated coefficients on these placebo variables are statistically insignificant, supporting the robustness of the results in Column 3.
 Table 6: Results of Testing the Impacts of Adverse Health Shocks
-on the Head’s Earning
+on the Head’s Earnings
 Dependent variable: semi-monthly earnings of the head
@@
 Health shocks on:
-The head [level] $-0.061$ $-0.972$ 1.319-1.319*- 1.319 ∗
+The head [level] 0.061-0.061 0.972-0.972 1.319-1.319*
 $[0.679]$ $[0.699]$ $[0.666]$
-The head [lag] 2.371-2.371***- 2.371 ∗ ∗ ∗ 2.868-2.868***- 2.868 ∗ ∗ ∗ 2.994-2.994***- 2.994 ∗ ∗ ∗
+The head [lag] 2.371-2.371*** 2.868-2.868*** 2.994-2.994***
 $[0.644]$ $[0.734]$ $[0.765]$
-The others [level] $0.889$
+Others [level] $0.889$
 $[1.275]$
-The others [lag] $3.675$
+Others [lag] $3.675$
 $[3.312]$
@@
 ***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors in brackets are clustered at the household level.
-Notes: This table shows the results from the regressions of the head’s earnings on the health shock variables using the semi-monthly panel dataset.
-A household with two semi-month observations is excluded from the sample in all the regressions because the lagged health shock variables are included.
-The family size variable, household fixed effects, and time-fixed effects are included in all the regressions.
-“Health shocks on the head” refers to the number of days that the head suffered illness.
-“Health shocks on the others” refers to the number of days that family members, excluding the head, were ill.
-“[level]” indicates the variable of the head’s earnings in the same semi-month cell.
-“[lag]” indicates the lagged variable (i.e., the shocks measured in the previous semi-month cell).
-Although the correlation coefficient between the level and lag of the health shock variables for the other family members is substantially high (Pearson’s coefficient $=0.99$), including either of the two variables does not change the finding on the negligible impacts of the health shocks on the other family members.
-Following the evidence, I characterize Equation 1 as a structural consumption-smoothing equation by using the first-stage reduced form equation as follows:
+Notes: This table presents results from regressions of the head’s earnings on health shock variables using the adjusted semi-monthly panel dataset.
+Households with only two semi-month observations are excluded from all regressions because lagged health shock variables are included.
+All regressions include the family size control, household fixed effects, and time fixed effects.
+“Health shocks on the head”: the number of days the head was ill.
+“Health shocks on others”: the number of days other family members (excluding the head) were ill.
+“[level]”: the variable measured in the same semi-month cell.
+“[lag]”: the variable from the previous semi-month cell.
+Although the correlation between the level and lag of health shocks for other family members is high (Pearson’s coefficient =0.99=0.99), including either variable does not change the finding that health shocks for other family members have negligible effects.
+I then characterize equation 1 as a structural equation by using the first-stage reduced form equation as follows:
 $\log y_{i,t}=\beta h_{i,t-1}+\psi x_{i,t}+\varpi_{i}+\tau_{t}+e_{i,t},$ (3)
-where $h_{i,t-1}$ is the number of days of the head’s illness in the previous semi-month cell, $\varpi_{i}$ is the household fixed effect, $\tau_{t}$ is the semi-month fixed effect, and $e_{i,t}$ is a random error term.
-For the risk-sharing mechanisms analysis under equation 2, the log-transformed disposable income in equation 3 ($y_{i,t}$) is replaced with the head’s earnings ($\tilde{y}_{i,t}$).
-The health shock variable is plausibly exogenous because the illnesses were from the influenza epidemics, as explained above.
-[^65] As noted in Section 2.3, the factories did not offer compensation for sick leave because the Factory Act did not specify benefits for workers’ illness (Study Group of the Factory Act 1916).
-Even in the largest factory in Tsukishima, the Ishikawajima Shipyard, any descriptions of the influenza epidemics are documented in the company’s history (Arai 1930).
-In fact, Miyoshi (1989) documented the issue that the head’s disease directly caused income loss due to the absence at that time.
-In addition, there is a correlation between the health shock variable and the head’s earnings at a statistically valid level (Table 6).
-Both support the evidence that the exogeneity of the instrument and relevance condition hold under this just-identified system.
+where $h_{i,t-1}$ is the number of days the head was ill in the previous semi-month cell, $\varpi_{i}$ is the household fixed effect, $\tau_{t}$ is the semi-month fixed effect, and $e_{i,t}$ is a random error term.
+For the analysis using equation 2, the log-transformed disposable income in equation 3 ($y_{i,t}$) is replaced with the head’s earnings ($\tilde{y}_{i,t}$).
+The health shock variable is plausibly exogenous because the illnesses resulted from influenza epidemics.
+[^50] As noted in Section 2.3, factories did not provide sick leave compensation because the Factory Act did not specify benefits for workers’ illness (Study Group of the Factory Act 1916).
+Even at the largest factory in Tsukishima, the Ishikawajima Shipyard, the company did not record any compensation for influenza epidemics (Arai 1930).
+In addition, the health shock variable is statistically correlated with the head’s earnings (Table 6).
+Together, this evidence supports the exogeneity of the instrument and satisfies the relevance condition in this just-identified system.
 Results: To What Extent do Illnesses Affect Consumption?
@@
 ***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors in brackets are clustered at the household level.
-Notes:
-Column 1 shows the results of equation 1, in which the log-transformed disposable income is replaced with the head’s illness (i.e., the number of days with illness in the previous semi-month cell).
-The dependent variables are the 11 measures of log-transformed consumption, and all the regressions include the family size control, household-fixed effects, and time-fixed effects.
-The estimated coefficient on the health shock variable is listed in the sub-column named “Coef.”.
-Column 2 shows the results of equation 1, in which the log-transformed disposable income is regarded as the endogenous variable.
-The head’s illness is an instrumental variable (i.e., excluded exogenous regressor) for the first-stage reduced-form equation (3).
-The estimated coefficient of the disposable income is listed in the sub-column named “Coef.”.
-The first-stage $F$-statistics on the health shock variable are listed in the final column.
-I first consider a reduced-form specification using the head’s illness as the exposure variable in equation 1 following Cochrane (1991).
-Column (1) in Table 7 shows the results.
-Overall, the coefficients on the exposure variable are negative, supporting evidence that adverse health shocks on the heads reduce consumption levels.
-Next, I explicitly model the structure underlying the relationship between the health shocks and earnings using the reduced-form equation of (3).
-If the negative impacts of the health shocks on consumption were derived from the income losses, the results from the structural estimation approach correspond to those in Column (1).
-The estimates from the structural equation are listed in Column (2) of Table 7.
-As shown, the results are congruent with the reduced-form results in Column (1).
-Importantly, the estimates are systematically larger than those reported in Table 4.
-The estimates for total and food consumption categories are $0.543$ and $0.355$, respectively, whereas the baseline estimates are $0.358$ and $0.136$.
-While this result partly reflects the bias reduction under the instrumental variable estimation, it is also noteworthy that this system is designed to reduce the bias by specifying the channel from adverse (instead of favorable) income shocks.
-For example, expenses on the allowance for children could be easily cut during unfavorable times but may not respond to favorable shocks.
-Similarly, the expenses for furniture are likely to be cut, but the households might not buy those frequently, meaning that there are no sufficient cases for reacting to rare health shock events.
-[^66] In other words, furniture may have only been purchased when the households had extra income.
-This may partly be reflected in the smaller value of the first-stage $F$-statistics in the furniture subcategory.
-[^67] Note that the logarithmic transformation leaves the cells with expenses on the furniture.
-Thus, the reduced-form equation captures the relationship between the health shocks and earnings in the cells with any expenses on furniture.
-While the first-stage $F$-statitics is $35.7$ in the linear-linear model, the second-stage result is robust against the functional form assumption.
-This may reflect the asymptotic theoretical advantage in the panel data model under nearly weak instruments case (Staiger and Stock 1997; Cai et al.
+Notes: Column 1 reports results of equation 1, replacing log-transformed disposable income with the head’s illness (number of days ill in the previous semi-month).
+Dependent variables are the 11 measures of log-transformed consumption.
+All regressions include the family size control, household fixed effects, and time fixed effects.
+The coefficient on the health shock variable is shown in the “Coef.” sub-column.
+Column 2 treats log-transformed disposable income as endogenous, using the head’s illness as an instrument in the first-stage equation (3).
+The estimated coefficient on disposable income appears in the “Coef.” sub-column, and first-stage $F$-statistics for the instrument are listed in the final column.
+I first consider a reduced-form specification using the head’s illness as the exposure variable in equation 1, following Cochrane (1991).
+Column 1 in Table 7 presents the results, providing evidence that adverse health shocks to the heads reduced consumption levels.
+Next, I explicitly model the structure underlying the relationship between health shocks and earnings using the reduced-form equation in (3).
+If the negative impacts of health shocks on consumption derive from income losses, the results from the instrumental variable approach should correspond to those in Column 1.
+The estimates are listed in Column 2 of Table 7.
+As shown, these results align with the reduced-form findings in Column 1.
+Importantly, the estimates are systematically larger than those reported in Table 4, with total and food consumption estimates of $0.543$ and $0.355$, respectively.
+While this partly reflects bias reduction under instrumental variable estimation, it also reflects the system’s design to reduce bias by specifying the channel from adverse (rather than favorable) income shocks.
+For example, allowances for children can be cut during unfavorable periods but may not increase during favorable ones.
+Similarly, furniture expenses are more likely to be reduced during shocks, but households purchase furniture infrequently, providing few observations for rare health shocks.
+[^51] In other words, furniture may only have been purchased when households had extra income.
+This pattern may partly explain the smaller first-stage $F$-statistics for the furniture subcategory.
+[^52] Note that the logarithmic transformation retains cells with furniture expenses.
+Thus, the reduced-form equation captures the relationship between health shocks and earnings in cells with any furniture expenses.
+While the first-stage $F$-statistic is $35.7$ in the linear-linear model, the second-stage result is robust to the functional form assumption.
+This may reflect the asymptotic advantage of panel data models under nearly weak instruments (Staiger and Stock 1997; Cai et al.
 2012; Baltagi et al.
 2012).
-The same argument is applicable to the housing category.
-A similar interpretation can be applied to the transportation category because family travel was not a frequent event and thus, unlikely to coincide with the timing of illness.
-By contrast, the entertainment expenses were still likely to be cut in case of illness, which might reflect the curtailment of tobacco, theater-going, and the allowance for heads.
-This may be another explanation for the difference in the results from my baseline results in Table 4.
-While the estimates for the clothes and gifts categories now exceed one, the null of equality to the population mean of one is not statistically rejected for both cases.
-Although the large estimates may be partly due to the operation under the instrument based on rare health shock events, the findings are materially like my baseline results—both categories are sensitive to the shocks as are those of the dispensable categories.
+The same argument applies to the housing category.
+A similar interpretation applies to the transportation category, as family travel is infrequent and unlikely to coincide with illness events.
+By contrast, entertainment expenses are still likely to be cut in case of illness, reflecting reductions in tobacco use, theater visits, and allowances for heads.
+This may explain the differences from the baseline results in Table 4.
+Although the estimates for clothes and gifts now exceed one, the null hypothesis of equality to the population mean of one is not statistically rejected for either category.
+While the large estimates may partly result from the instrument based on rare health shock events, the findings are broadly consistent with the baseline results—both categories remain sensitive to shocks, similar to other dispensable categories.
 Results: How did the Households Compensate for Losses from Illness?
 Table 8: Results of Testing the Risk-coping Mechanisms:
-Adverse Health Shock as the Idiosyncratic Shock
+Adverse Health as the Idiosyncratic Shock
 Semi-monthly panels
 (1) Reduced-form approach (2) Instrumental variable approach
-Head’s illness Obs. Head’s earning Obs.
+Head’s illness Obs. Head’s earnings Obs.
 Coef. Std. error Coef. Std. error
-Panel A: Savings, insurance, borrowing, and gifts (yen)
+Panel A: Net savings, insurance, borrowing, and gifts (income - expenditure, yen)
 Net savings 0.397 [0.318] 254 -0.167 [0.117] 254
@@
 Panel B: Labor supply adjustments & sales of miscellaneous assets (yen)
-Other member’s earnings 0.001 [0.081] 254 -0.001 [0.034] 254
+Other members’ earnings 0.001 [0.081] 254 -0.001 [0.034] 254
 Sales of miscellaneous assets 0.002 [0.002] 254 -0.001 [0.001] 254
 ***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors in brackets are clustered at the household level.
-Notes:
-Column 1 shows the results of equation 2, in which the head’s earning is replaced with the head’s illness (i.e., the number of days with illness in the previous semi-month cell).
-The family size control, household fixed effects, and time-fixed effects are included in all the regressions.
-The estimated coefficients on the head’s illness are listed in the sub-column named “Coef.”.
-Column 2 shows the results of equation 2, in which the head’s earning is treated as the endogenous variable.
-The head’s illness is an instrumental variable (i.e., excluded exogenous regressor) in the first-stage reduced-form equation.
-The first-stage result is identical to the result listed in Column (2) of Table 6 with the first-stage $F$-statistics of $13.6$.
-The estimated coefficients on the head’s earnings are listed in the sub-column named “Coef.”.
-Panel A: Each net income variable is defined as the difference between income and expenses.
-“Net savings” refers to the difference between the amount withdrawn and the total deposits.
-“Net insurance” refers to the amount received from insurance minus the expenses paid for insurance.
-“Net borrowing” refers to the amount of borrowing minus the debt payments.
-“Net credit purchase” refers to the amount of credit purchases minus credit redeemed.
-“Net gifts” refers to the total received amounts, including both pecuniary and non-pecuniary gifts, minus the payments for gifts.
-Panel B: “Other member’s earnings” refers to the total income earned by all family members except for the head of the household.
-“Sales of miscellaneous assets” includes the sale of daily miscellaneous goods such as newspapers and empty bottles.
+Notes: Column 1 reports the results of equation 2, where the head’s earnings are replaced by the head’s illness (i.e., the number of days ill in the previous semi-month cell).
+Family size, household fixed effects, and time fixed effects are included in all regressions.
+The estimated coefficients on the head’s illness are listed in the sub-column labeled “Coef.” Column 2 reports the results of equation 2, treating the head’s earnings as the endogenous variable.
+The head’s illness serves as an instrumental variable in the first-stage reduced-form equation.
+The first-stage results match those in Column 2of Table 6, with a first-stage $F$-statistic of $13.6$.
+The estimated coefficients on the head’s earnings are listed in the sub-column labeled “Coef.” Panel A: Each net income variable is defined as income minus expenditure.
+“Net savings”: withdrawals minus deposits.
+“Net insurance”: insurance receipts minus premiums paid.
+“Net borrowing”: borrowing minus debt repayments.
+This category considers net income from both pawnshops and other lending institutions.
+“Net credit purchase”: total credit purchases minus repaid credit.
+“Net gifts”: total received gifts, including pecuniary and non-pecuniary, minus any gift payments.
+Panel B: “Other members’ earnings”: the total income earned by all household members except the head.
+“Sales of miscellaneous assets”: include sales of daily items such as newspapers and empty bottles.
 All dependent variables are in yen.
-I tested the risk-coping strategies for mitigating adverse health shocks in Table 8.
-Column 1 lists the reduced-form results in the same manner as Table 5.
-The estimated coefficients on the net income variables are positive in most cases, indicating the households’ compensating strategies against the shocks.
-The estimate for the credit purchase category is weakly statistically significant, with the largest magnitude.
-It suggests that the one-day increase in the head’s illness increased the net credit purchases by $1.29$ yen, which accounts for approximately half of their average daily wages ($1.29/2.56$).
-Column 2 presents the estimates from the structural equation using the head’s earnings as the endogenous variable.
-A similar but clearer result is obtained—the estimate for the credit purchase category is approximately four times greater than that reported in Column (1) of Table 5.
-It indicates that a one yen decrease in the head’s earnings increases the net income from credit purchases by $0.55$ yen.
-This magnitude is consistent with the direct effect of the health shocks on the net credit purchase estimated in Column (1) of Table 8.686868This means that $2.56\times 0.55=1.4$ yen, which is close to $1.3$ yen in Column (1).
-The estimates for the savings and gifts categories are still negative but now statistically insignificant, supporting the evidence that credit purchase is the primary risk-mitigation strategy for dealing with health shocks.
-Since withdrawing and receiving gifts to use the money at the retailers takes more time (steps), the households likely prefer to use credit purchases.
-Overall, the event-study analysis using adverse health shocks confirms the baseline findings and provides a sharper illustration of consumption-smoothing strategies in the case of adverse idiosyncratic shock events.
-My results show that the unique available response to idiosyncratic income shocks among the THBS households was to use credit purchases.
-The available compensation via credit purchases accounts for approximately half of their temporary income losses, which partially contributes to smoothing their food consumption in the short term.
-5 Credit System in Local Economy
-The foregoing results enlighten the roles of credit institutions among skilled workers’ households during industrialization.
-As the credit for purchases was provided by local retailers, this suggests the importance of the relationships between the workers and retailers.
-To the best of my knowledge, however, there is little historical research on credit purchases in prewar Japan.
-Hence, the details of this credit institution are not yet clear.
-This may mainly be because the information on credit purchases is personal, as in the case of state-contingent loans in the rural credit market (Udry 1994) and thus, not well documented.
-Accordingly, I analyze the institution using both the THBS dataset and available official documents.
+Table 8 presents the analysis of the risk-coping strategies for mitigating adverse health shocks.
+Column 1 presents the reduced-form results.
+The estimated coefficients on the net income variables are positive in most cases, indicating households’ responses to the shocks.
+The estimate for the credit purchase category is weakly statistically significant and has the largest magnitude.
+This suggests that a one standard deviation increase in the duration of the head’s flu exposure increases net credit purchases by $4.6$ yen ($1.29\times 3.6~\text{days}$), which represents approximately 1616% of their average semi-monthly earnings of the head (Table 3).
+Column 2 reports the estimates from the instrumental variable approach, treating the head’s earnings as endogenous.
+The estimate for the credit purchase category is approximately four times larger than that reported in Column 1 of Table 5.
+It implies that a one yen decrease in the head’s earnings increases net income from credit purchases by $0.55$ yen.
+This magnitude aligns with the direct effect of health shocks on net credit purchases estimated in Column 1 of Table 8.535353Given their average daily wage of $2.56$ (Online Appendix B.6), $2.56\times 0.55=1.4$ yen, which is close to $1.3$ yen in Column 1.
+The estimates for the savings and gifts categories remain negative but are now statistically insignificant.
+Overall, these results suggest that the primary response to adverse idiosyncratic income shocks among THBS households was to use credit purchases.
+Compensation via credit purchases accounted for roughly half of their temporary income losses, helping to smooth consumption in the short term.
+5 Credit System in the Local Economy
+The results highlight the importance of the relationships between workers and local retailers.
+To the best of my knowledge, however, little historical research exists on credit purchases in prewar Japan, and the details of this credit institution remain unclear.
+As with state-contingent loans in the rural credit market (Udry 1994), information on credit purchases is personal and poorly documented.
+Accordingly, I analyze the institution using the THBS dataset along with a few available official documents.
 5.1 Availability and Premiums: Evidence from the THBS Dataset
@@
 by Different Time-series Frequencies
-Notes: Figure 6a, and 6c illustrate the time-series plots of the average daily and semi-monthly credit purchase and payment, respectively.
+Notes: Figures 5a and 5c plot average daily and semi-monthly credit purchases and payments, respectively.
 The blue and red dashed lines indicate the average credit purchases and payments, respectively.
-Figure 6b and 6d illustrate the time-series plots of the average daily and semi-monthly credit purchase minus payment.
-The red solid line highlights the zero value.
-Figures 6a and 6b show the daily data between January 12 and November 11, 1919.
-Figures 6c and 6d show the semi-monthly series calculated using the daily data between January and November 11, 1919.
-All figures are based on the 29 THBS households that used any credit purchases.
+Figures 5b and 5d plot credit purchases minus payments.
+The red solid line marks zero.
+Daily data cover January 12–November 11, 1919.
+Figures 5c and 5d show the semi-monthly series derived from the same daily data.
+All figures use the 29 THBS households that made credit purchases.
 Source: Created by the author using the THBS sample.
-I begin this descriptive analysis by examining the credit purchases included in the THBS dataset.
-Most THBS households ($29/33$) used any credit purchases during their sample period, confirming that credit purchase was a common practice among factory workers.
+I begin this descriptive analysis by examining credit purchases in the THBS dataset.
+Most THBS households ($29/33$) made credit purchases during the sample period, confirming that this practice was common among factory workers.
 Figure 5a illustrates the average daily credit purchases and payments.
-Generally, there is a peak per semi-month in payments, which reflects the practice of making the payments for the credit purchase on or soon after their paydays.
-The purchase series shows a similar tendency but with lower spikes and larger variance.
-This indicates that the households first purchased goods for a certain amount on or soon after their paydays, and then used purchased perishable foods such as fish, fruit, and vegetables when they needed them.
-Figure 6b illustrating the net purchase confirms this behavior: the households used the credit on a daily basis and redeemed it near paydays.
-Figure 6c shows that the payment slightly exceeded purchase, suggesting that many households could approximately repay their loans.
-Although the difference between payment and purchase is relatively obvious after August 1919, Figure 6d indicates that this difference is not statistically significant in all the semi-months.
-[^69] Strictly, payment can include accrued liabilities that should have been paid prior to the previous period.
-It is, therefore, reasonable that the average payment is greater than the average purchase in most semi-month cells.
-Note again that the wider variances in the latter sample period are based on the smaller available number of households (Online Appendix Figure B.4).
-I then analyzed whether the volatility of earnings restricted credit purchases.
-Since credit of worker was unobservable, I used the volatility of earnings as a proxy of credit risk in the following empirical specification:
+Payments generally peak semi-monthly, reflecting the practice of settling credit purchases on or shortly after paydays.
+The purchase series shows a similar pattern but with smaller peaks and greater variance.
+This indicates that households first make purchases on or soon after paydays and then consume perishable goods such as fish, fruit, and vegetables.
+Figure 5b, which illustrates net purchases, confirms this behavior: households used credit daily and repaid it near paydays.
+Figure 5c shows that payments slightly exceed purchases, implying near-full loan repayment for most households.
+Although the difference between payment and purchase becomes more noticeable after August 1919, it is not statistically significant in all semi-months (Figure 5d).
+[^54] Strictly, payments can include accrued liabilities from prior periods.
+Therefore, average payments reasonably exceed average purchases in most semi-month cells.
+Note that the wider variances in the later sample period are owing to the smaller number of available households (Online Appendix Figure B.4).
+Next, I analyze whether earnings volatility restricts credit purchases.
+Because household credit is unobservable, I use earnings volatility as a proxy for credit risk in the following empirical specification:
 $I(\overline{CP}_{i}=0)=\varkappa_{0}+\varkappa_{1}CV_{i}+\bar{\mathbf{x}}_{i}^{\prime}\bm{\chi}+v_{i},$ (4)
-where $I(\cdot)$ is an indicator function, $\overline{CP}$ is the (unit-)average credit purchases, $CV$ is the coefficient of variation for head’s earnings, $\bar{\mathbf{x}}$ is a vector of the (unit-)average family composition variables, and $v$ is a random error term.
+where $I(\cdot)$ is an indicator function, $\overline{CP}$ is the (unit-)average credit purchases, $CV$ is the coefficient of variation for the head’s earnings, $\bar{\mathbf{x}}$ is a vector of (unit-)average family composition variables, and $v$ is a random error term.
 All variables are calculated using the semi-monthly series.
-[^70] For example, the coefficient of variation for household $i$ is defined as $CV_{i}=\{\sum_{t}^{T_{i}}(\tilde{y}_{i,t}-\bar{\tilde{y}}_{i,.})/(T_{i}-1)\}^{\frac{1}{2}}/\bar{\tilde{y}}_{i,.}$, where $T_{i}$ indicates the number of time-series observations, and $\bar{\tilde{y}}_{i,.}$ indicates the time-series average of the head’s income.
-In this cross-sectional model, I include a set of family composition variables to control for unobservable preferences because the household-fixed effect cannot be modeled as an error component.
-If retailers allowed households to use credit purchases regardless of the variability of earnings, the estimate of $\varkappa_{1}$ becomes close to zero.
-By contrast, if retailers did not allow households with high-income volatility to trade on credit, the estimate should be positive and reasonably large.
+[^55] For example, the coefficient of variation for household $i$ is defined as $CV_{i}=\{\sum_{t}^{T_{i}}(\tilde{y}_{i,t}-\bar{\tilde{y}}_{i,.})^{2}/(T_{i}-1)\}^{\frac{1}{2}}/\bar{\tilde{y}}_{i,.}$, where $T_{i}$ indicates the number of time-series observations, and $\bar{\tilde{y}}_{i,.}$ is the time-series average of the head’s income.
+In this cross-sectional model, I include family composition variables to control for unobserved preferences, since household fixed effects cannot be included as an error component.
+If retailers allow households to use credit regardless of earnings variability, the estimate of $\varkappa_{1}$ should be close to zero.
+By contrast, if households with high earnings volatility are restricted from trading on credit, the estimate should be positive and sizable.
 Online Appendix C.7 summarizes the results.
-Shortly, the estimated coefficient ($\hat{\varkappa}_{1}$) is close to zero and statistically insignificant in different specifications.
-This result remains unchanged when alternative thresholds for the dependent indicator variable are used to include the households with less than the first quartile of the average credit purchases.
-This suggests that retailers may be generous in considering the volatility of earnings when providing credit to factory workers, which is consistent with the retailers’ incentive to secure their local customers.
-This, in turn, implies that retailers may factor risk into retail prices or by charging interest on their loans.
-Unfortunately, there are no systematic statistics on the interest charged on credit purchases.
-It may be possible to reverse calculate the risk premiums by differencing payments and purchases.
-This reverse calculation technique requires information on both the price and type of retailers.
-However, budget books usually document price-by-item information but not price-by-retailer information.
-This is problematic because the same items were often sold in different types of retailers: for example, eggs were frequently sold in greengrocers (yaoya) but sometimes also in general stores (aramonoya).
-Fortunately, I found a specific month in a budget book that has documented the price-by-retailer information.
+In short, the estimated coefficient ($\hat{\varkappa}_{1}$) is close to zero and statistically insignificant.
+This finding remains unchanged when alternative thresholds for the dependent indicator variable are applied, including households below the first quartile of average credit purchases.
+This suggests that retailers were generous in considering earnings volatility when extending credit to factory workers, consistent with their incentive to retain local customers (Section 5.2).
+This, in turn, suggests that retailers may factor risk into retail prices or by charging interest on loans.
+Unfortunately, no systematic statistics exist on interest charged for credit purchases.
+One might attempt to reverse-calculate risk premiums by differencing payments and purchases.
+However, this approach requires information on both the price and type of retailer.
+Budget books usually record prices by item, but not by retailer, which is problematic because the same items are often sold by different types of retailers–for example, eggs are commonly sold in greengrocers (yaoya) but occasionally also in general stores (aramonoya).
+Fortunately, a specific month in a budget book provides price-by-retailer information.
 [^56] The OISR, Archives of the Tsukishima Survey (THBS #4, February 1919).
-According to this budget book, the reimbursements were made semi-monthly, as seen in Figure 6b.
-They bought various goods in greengrocers, fish stores, and general stores between February 1 and February 14, 1919, and paid exactly the same amounts on the 15th.
-[^72] The repayment period in credit purchases among retailers and wholesalers (footnote 76) was often set at approximately one month (Tokyo City Office 1932, pp.
+According to this record, reimbursements were made semi-monthly, as shown in Figure 5b.
+Between February 1 and February 14, 1919, the household purchased various goods from greengrocers, fish stores, and general stores, and paid exactly the same amounts on the 15th.
+[^57] The repayment period for credit purchases among retailers and wholesalers (footnote 61) was often set at approximately one month (Tokyo City Office 1932, pp.
 157-159).
 Therefore, customers likely made repayments at semi-monthly or monthly intervals.
-In the latter half of February, the household paid the same amounts to greengrocers and general stores on the 28th, but paid less to the fish stores.
-As the budget book for March does not have information on the type of retailers, it is difficult to determine exactly how the reimbursement shortfall was managed.
-However, this case suggests that the risk premiums may have been included in the retail prices but not in the form of interest.
+In the latter half of February, the household paid the same amounts to greengrocers and general stores on the 28th, but less to fish stores.
+Because the March budget book lacks information on retailer type, determining exactly how the reimbursement shortfall was handled is difficult.
+Nevertheless, this case suggests that risk premiums were incorporated into retail prices rather than charged as explicit interest.
 5.2 Local Environment and Retailers’ Incentives: Evidence from the Official Surveys
-Next, I assess circumstances among the retailers using available survey reports on the credit institution.
-A survey report on retailers in 1930 (hereafter 1930 Survey Report) revealed that credit purchase was a popular payment method in transactions with consumers.
+Next, I examine the circumstances among retailers using a survey report published by the Tokyo City Office in 1930 (hereafter, the 1930 Survey Report).
 [^58] This survey was conducted in February and March 1930 by Tokyo City and covered $3,892$ small and medium enterprises (Tokyo City Office 1932a, p.
-1).
-Those included retailers as well as wholesale and other small businesses.
-The number of respondents varied depending on the question items.
-The number of respondents in the statistical table on the payment methods was $2,654$, of which $1,197$ were retailers (Tokyo City Office 1932b, p.
+1), including retailers, wholesalers, and other small businesses.
+The number of respondents varies by question.
+The statistical table on payment methods includes $2,654$ respondents, of whom $1,197$ were retailers (Tokyo City Office 1932b, p.
 273).
-To the best of my knowledge, this is the most comprehensive survey of the retailers in Tokyo circa 1930.
-In fact, $99$%($155/157$) of grain retailers, $81$%($35/43$) of fish retailers, $78$%($64/82$) of greengrocers, and $94$%($45/48$) of fuel retailers used credit purchases for their selling method (Tokyo City Office 1932b, pp.
+To the best of my knowledge, this is the most comprehensive survey of retailers in Tokyo circa 1930.
+Although a similar survey around 1920 is unavailable, the industrial and household structures in 1920 and 1930 were generally comparable (Online Appendix Table C.6).
+Thus, the 1930 report can serve as a reference to illustrate how credit purchases functioned in the city.
+According to the report, credit purchases were a popular payment method across consumer transactions: 9999%($155/157$) of grain retailers, 8181%($35/43$) of fish retailers, 7878%($64/82$) of greengrocers, and 9494%($45/48$) of fuel retailers used credit purchases as part of their selling methods (Tokyo City Office 1932b, pp.
 260; 263).
-Thus, credit purchases in Tokyo were neither region- nor religion-specific customs.
+This widespread use indicates that credit purchases were neither region- nor religion-specific customs.
 Table 9: Customers, Sales Methods, and Competitiveness
-in the Retailers in Tokyo City
+among Retailers in Tokyo City
 Selling item Number of Share of sales Social class of Sales method (%) Number of
-stores for regular the Regular Customer (%) Cash Credit Monthly competitors
-reported customer (%) Upper Middle Other purchase installment reported
+stores for regular regular customers (%) Cash Credit Monthly competitors
+customers (%) Upper Middle Other purchase installment
 White Rice 101 73.1 16 45.7 38.2 26.9 71.5 1.7 8.5
 Fish 59 64.6 6.8 57.6 35.6 43.9 56.1 0 5.0
-Greengrocer 66 45.4 9.4 40.4 50.2 65.8 34.2 0 6.2
+Fruits and vegetables 66 45.4 9.4 40.4 50.2 65.8 34.2 0 6.2
 Firewood and charcoal 100 51.1 15.7 50.9 33.4 20.4 78.4 1.3 8.6
-Notes: This table summarizes the statistics on the sales in the $326$ retailers selling white rice, fish, fruit and vegetables, and firewood and charcoal in Tokyo surveyed in 1935.
-The number of stores reported the share of regular customers (column 2) is $90$, $56$, $61$, and $93$ in the white rice, fish, greengrocer, and fuel categories, respectively.
-The “share of sales for regular customers” is defined as the percentage of sales for regular customers in the annual sales.
-The definition of the social classes (“Upper”; “Middle”; “Other”) is not defined in the report.
-The share of sales method is based on the percentage of sales using each method in the annual sales.
-The number of competitors indicates the number of peers within a three-cho, approximately $327$ meters, radius.
-The department stores and retail markets are not included in this survey.
-Source: Tokyo City Chamber of Commerce and Industry 1937, pp. legend; 20–21; questionnaire.
-A survey of retailers in 1935 published by the Tokyo City Chamber of Commerce and Industry provides more detailed information.
-Table 9 summarizes the information on the retailers’ sales measured in this survey.
-[^74] This survey covered $939$ retailers, and its report was edited on December 31, 1935.
-I digitized the statistics for all the retailers selling the representative items (rice, fish, fruit and vegetables, and fuels).
-The survey instrument included detailed questions regarding sales and purchases (Tokyo City Chamber of Commerce and Industry 1937, pp.
+Notes: This table summarizes sales statistics for $326$ retailers in Tokyo surveyed in 1935, covering white rice, fish, fruits and vegetables, and firewood and charcoal.
+The number of stores reporting the share of regular customers (Column 2) is $90$, $56$, $61$, and $93$ for the white rice, fish, fruits and vegetables, and fuel categories, respectively.
+The “share of sales for regular customers” refers to the percentage of annual sales made to regular customers.
+Social classes (“Upper,” “Middle,” and “Other”) are not defined in the report.
+The share of sales by method is based on the percentage of annual sales conducted using each method.
+The number of competitors indicates peers within a three-cho radius (approximately $327$ meters).
+Department stores and retail markets are excluded from this survey.
+Source: Tokyo Chamber of Commerce and Industry 1937, pp.
+legend; 20–21; questionnaire.
+A survey of retailers in 1935, published by the Tokyo Chamber of Commerce and Industry, provides detailed information on sales practices.
+Table 9 summarizes these statistics.
+[^59] The survey covered $939$ retailers, and the report was finalized on December 31, 1935.
+It includes detailed questions regarding sales and purchases (Tokyo Chamber of Commerce and Industry 1937, pp.
 legend, questionnaire).
-Online Appendix Table C.1 decomposes the statistics listed in Table 9 by scale of businesses.
-I confirm that the main findings are materially similar.
-First, many annual sales were from their regular customers, most of whom were from the non-upper classes.
-$73$% of annual sales of rice retailers were from their regular customers, and more than $80$% of them were from non-upper classes, including factory workers.
-Second, credit purchase was a popular sales method, especially among rice and fuel retailers—more than $70$% of the sales were conducted with credit in both shops.
-[^75] These figures are systematically lower than those from the 1930 Survey Report.
-Given that both surveys show a systematic tendency toward credit purchases, this may be due to the differences in the samples, the definition of the retailers in each category, and the definition of the credit purchase itself.
-Although cash was also used among greengrocers, $34$% of the sales were settled using credit.
-This suggests that credit purchase was a popular payment method and that regular customers settled a large part of it.
-[^76] The 1930 Survey reveals that the retailers also commonly used credit purchases to buy their source goods from wholesalers (Tokyo City Office 1932a, pp.
-154–157).
-This means that credit purchase was a familiar payment method for retailers as well as consumers.
-This confirms that the retailers were engaged in credit transactions with long-term relationships with their regular customers.
-[^77] A textbook for retailers published in 1924 explains the psychology behind this long-term relationship—credit purchase was provided for the convenience of the customer’s budget but also for the “vanity” (status) of being regular customers (Shimizu 1924, pp.
+Online Appendix Table C.3 decomposes the statistics in Table 9 by business scale, confirming that the main findings are materially similar.
+First, a large share of annual sales come from regular customers, most of whom belonged to non-upper classes.
+7373% of annual sales by rice retailers come from regular customers, and more than 8080% of these customers were from non-upper classes, including factory workers.
+Second, credit purchase was a popular sales method, particularly among rice and fuel retailers: more than 7070% of sales in both categories were conducted on credit.
+[^60] These figures are systematically lower than those in the 1930 Survey Report.
+Both surveys, however, show the same general tendency toward credit purchases; differences may reflect variations in the samples, retailer category definitions, or the definition of credit purchase itself.
+Cash was also accepted by greengrocers, but 3434% of their sales were on credit.
+This confirms that retailers engaged in credit transactions, often maintaining long-term relationships with regular customers.
+[^61] A 1924 textbook for retailers explains the psychology behind these long-term relationships: credit purchases are convenient for customers from a budget perspective and confer “vanity” or status as regular customers (Shimizu 1924, pp.
 115–116).
-The final column of this table further provides suggestive evidence for this.
-The survey asked retailers whether the sales area was competitive or not, and if yes, they were asked to report the number of competitors within a three-cho (approximately $330$ meters) radius.
-As shown, rice and fuel stores have two to three more competitors than the retailers selling fish, fruit, and vegetables.
-Given the relatively high utilization of credit in rice and fuel retailers, they may have responded to competition by allowing their customers to trade on credit.
-Online Appendix C.3 provides a set of regression-based evidence on the positive relationship between the use of credit purchases and the number of local competitors.
-Finally, I assessed the losses suffered by retailers for credit purchases.
-According to the 1930 Survey Report, $99$% of rice and $97$% of fuel retailers lost their sales.
-[^78] Online Appendix C.4 summarizes the statistics on the losses in credit purchases available from the 1930 Survey Report in detail.
-This implies that credit purchase was inherently a loss-making sales method for the retailers.
-However, these losses were usually marginal and capped at 5% of the total sales.
-For example, the available statistics suggest that $87$% of rice and $65$% of fuel retailers suffered losses amounting to $5$% of sales.
-While the retailers could not avoid losses in credit sales, they might have set prices that anticipated predicted losses to some extent.
-This seems to be a plausible sales strategy for retailers to manage adverse selection issues under imperfect consumer information.
-To summarize, the available documents show that retailers competed with other retailers in the same sales area.
-Under this circumstance, they provided regular consumers with the opportunity to trade on credit without interest for at least one month.
-The 1930 Survey Report indeed stated that “intense competition among competitors has led to a high frequency of credit purchases” (Tokyo City Office 1932b, p.
+Retailers also commonly used credit purchases to buy source goods from wholesalers (Tokyo City Office 1932a, pp.
+154–157), making credit a familiar payment method for both retailers and consumers.
+The final column of Table 9 provides suggestive evidence on competition.
+The survey asks whether the sales area is competitive and, if so, how many competitors are located within a three-cho (approximately $327$ meters) radius.
+Rice and fuel stores have two to three more competitors than retailers selling fish or fruits and vegetables.
+Given the relatively high use of credit among rice and fuel retailers, this suggests that they may have responded to competition by extending credit to their customers.
+Online Appendix C.5 presents regression-based evidence of a positive relationship between credit use and the number of local competitors.
+Finally, I assess the losses suffered by retailers from unpaid credit purchases.
+According to the 1930 Survey Report, 9999% of rice retailers and 9797% of fuel retailers experienced losses on credit sales.
+[^62] Online Appendix C.6 summarizes detailed statistics on losses from credit purchases in the 1930 Survey Report.
+This indicates that credit purchases were inherently loss-making for retailers.
+However, these losses were generally small, typically capped at 5% of total sales.
+For example, available statistics show that 8787% of rice retailers and 6565% of fuel retailers suffered losses amounting to 55% of sales.
+Although retailers could not avoid losses from credit sales, they may have set prices to anticipate these expected losses.
+This appears to be a plausible strategy for managing adverse selection under imperfect consumer information (Section 5.1).
+In summary, the available documents show that retailers competed with others in the same sales area and provided regular customers the opportunity to trade on credit without interest for at least one month.
+The 1930 Survey Report noted that “intense competition among competitors has led to a high frequency of credit purchases” (Tokyo City Office 1932b, p.
 415).
-This accessibility to credit purchases could help workers manage consumption after unexpected individual setbacks in the short term.
+This informal credit system could help workers manage consumption following unexpected individual setbacks in the short term.
 [^63] Another potential explanation for the popularity of credit purchases might be theft.
-However, the number of thefts in $100$ people in 1919 was $0.5$ in Tsukishima, which is close to or even slightly lower than the city average of $1.3$ (National Police Agency 1920b, pp 53–54; 244–245).
-In fact, Tsukishima Police Department does not mention that Tsukshima was an area with a particularly high rate of theft (Tsukishima Police Department 1976, pp.
+However, the number of thefts per $100$ people in 1919 was $0.5$ in Tsukishima, close to or slightly lower than the city average of $1.3$ (National Police Agency 1920b, pp 53–54; 244–245).
+The Tsukishima Police Department does not report that the area had an unusually high rate of theft (Tsukishima Police Department 1976, pp.
 244–245).
-Therefore, the crime channel is not relevant to explain the use of credit purchases.
+Therefore, the crime channel does not explain the use of credit purchases.
 6 Conclusion
-This study investigates the consumption-smoothing strategies used by working-class households in prewar Tokyo.
-I digitized the daily longitudinal budget survey on factory worker households conducted in Tsukishima in 1919.
-I assessed the potential measurement error due to the miss-assignments of paydays and proposed a way to avoid the assignment errors.
-The estimated elasticity for total consumption indicates that the factory worker households suffered certain consumption losses when they faced idiosyncratic income shocks.
-The estimated elasticity for total expenditure ($0.43$) is in a similar range to those in rural economies in developing countries.
-The elasticity for the rural households in Indian villages between 1975 and 1984 was reported to be approximately $0.5$.
-As for the urban households, the estimated elasticity was $0.4$ for Bangkok between 1975 and 1990, which had a similar GDP per capita to 1920 Japan.
-Since my estimate is from a specific occupation, skilled factory workers, we need to be careful when making comparisons with these estimates based on urban samples which include a variety of occupations.
-Nevertheless, the factory workers in the historical metropolis of Japan circa WWI were vulnerable to idiosyncratic income shocks.
-A recent study found an income elasticity of $0.39$ for factory workers’ households in Osaka city circa 1920.808080The estimates for India, Bangkok, and Osaka are obtained from Ravallion and Chaudhuri (1997), Townsend (1995), and Ogasawara (2024), respectively.
-My estimate from the adjusted monthly panels is slightly larger than this estimate.
-If taken literally, this implies that the Tsukishima factory workers were slightly more vulnerable than those in Osaka.
-This indicates that higher levels of urbanization do not necessarily correspond to a greater availability of risk-mitigation institutions.
-Therefore, inter-city and cross-national comparisons of income elasticities and analysis of the institutional characteristics behind consumption are subjects for future work.
-This would add a new analytical perspective to the study of economic history.
-Despite the vulnerability, the results from mechanism analyses show that credit purchases were commonly used among workers and played a role in mitigating adverse idiosyncratic shocks in several consumption subcategories, particularly food.
-My findings offer a new historical view of the interdependence between consumers and local retailers.
-As researchers build an understanding of local credit institutions we can comprehend regional consumption behavior and may understand how living standards improved in prewar Japan.
-Appendices
+This study investigates the consumption-smoothing strategies of skilled factory-worker households by digitizing the longitudinal budget survey conducted in Tsukishima in 1919.
+Despite this narrow geographical focus, the analytical sample represents average skilled factory-worker households in this representative manufacturing area of Tokyo.
+The estimated elasticity for total consumption indicates that households experienced measurable consumption losses when facing idiosyncratic income shocks.
+The estimated elasticity for total expenditure ($0.43$) is similar to estimates found in developing countries.
+The elasticity for rural households in Indian villages between 1975 and 1984 is reported to be approximately $0.5$.
+In Bangkok, which had a GDP per capita similar to that of Japan in 1920, the estimated elasticity for urban households is $0.4$ between 1975 and 1990.646464The estimates for India, Bangkok, and Osaka are from Ravallion and Chaudhuri (1997), Townsend (1995), and Ogasawara (2024), respectively.
+A recent study reports an income elasticity of $0.39$ for factory-worker households in Osaka around 1920.
+My estimate from the monthly panels is slightly higher.
+Taken literally, this suggests that Tsukishima factory workers were slightly more vulnerable than those in Osaka.
+This indicates that higher levels of urbanization do not necessarily correspond to greater availability of risk-coping institutions.
+A limitation of this study is that it does not conduct inter-city and cross-national comparisons of income elasticities, nor does it analyze the institutional characteristics affecting consumption; these are left for future research.
+Despite this vulnerability, results from the mechanism analyses show that credit purchases were commonly used by workers and helped mitigate adverse idiosyncratic shocks in several consumption subcategories, particularly food.
+These findings offer a new historical perspective on the interdependence between workers and local retailers.
+As researchers develop a better understanding of local credit institutions, we may gain insight into regional consumption behavior and the factors that improved living standards in prewar Japan.
 Appendix Appendix A Background Appendix
@@
 Figure A.1b shows the share of people in Tokyo City (diamond) and the cities with more than $50,000$ people (circle) in total population in Japan.
-Sources: Umemura et al. (1983, pp. 303–305); Umemura et al. (1988, pp. 166; 168).
-Table A.1: City Characteristics:
-Tokyo, Osaka, Kyoto, and Nagoya
-Name of cities
-Tokyo Osaka Kyoto Nagoya
-Panel A: Population and Workers
-Population 2,173,201 1,252,983 591,323 429,997
-Male workers 768,633 476,229 189,922 138,275
-Panel B: Industrial Structure (share in male workers)
-Agriculture 1.0 0.8 2.2 1.8
-Fisheries 0.1 0.1 0.0 0.4
-Mining 0.4 0.3 0.2 0.2
-Manufacturing 44.5 45.6 49.0 45.8
-Commerce 32.4 34.0 32.3 29.6
-Transport 7.6 10.6 5.5 8.0
-Public service and professions 11.3 6.7 9.0 11.3
-Housework 0.3 1.8 0.1 0.0
-Other industry 2.4 0.1 1.5 2.8
-Panel C: Local Infrastructures
-Number of electrical lights per people in 1921 0.9 0.8 0.9 0.9
-Accessibility to modern water supply (%) 78.9 80.5 46.5 39.4
-Panel D: Meteorological Features
-January 1918 – December 1923
-Average daily temperature 14.1 15.4 14.1 14.6
-Average daily precipitation 4.7 3.9 4.6 4.8
-February – April only
-Average daily temperature 8.0 8.8 7.8 8.3
-Average daily precipitation 4.2 3.4 4.1 4.3
-Notes:
-Panel A: The number of people and male workers measured in the Population Census conducted in October 1920 are listed.
-Panel B: Occupations are classified based on the industrial classification of the 1920 Population Census. Unemployed people are not included in the calculation.
-Panel C: The number of electrical lights measured at the end of 1921 is divided by the number of people in the 1920 Population Census.
-The number of houses accessible to water supply measured at the end of September 1920 is divided by the number of total households measured in the 1920 Population Census (%).
-Panel D: The average daily temperature and precipitation between January 1, 1918, and December 31, 1923, and between February 1, 1920, and April 30, 1920 are listed.
-Sources: The number of people in 1920 is from the 1920 Population Census (Statistics Bureau of the Cabinet 1928, pp.
-186; 214; 222; 225).
-The number of male workers and their industrial structures are from the Statistics Bureau of the Cabinet (1929a, pp.
-20–21; 26–29).
-The number of houses accessible to the water supply is from the Water Works Association (1922, pp.
-69–70).
-The total number of households is from the 1920 Population Census (Statistics Bureau of the Cabinet 1928, pp.
-186; 214; 222; 225).
-The number of electrical lights is from the Ministry of Posts and Telecommunications (1922, p.
-558) The daily temperature and precipitation are from the Japan Meteorological Agency (database).
-Figure A.1a illustrates the developments of the large cities in Japan between 1889 and 1920.
-Table A.1 summarizes the scales, industrial structures, infrastructures, and meteorological features of the large cities.
+Sources: Umemura et al.
+(1983, pp.
+303–305); Umemura et al.
+(1988, pp.
+166; 168).
+Figure A.1a shows the developments of the large cities in Japan between 1889 and 1920.
 A.2 Financial Institutions in the Cities
-Several financial institutions were available for working-class households in the cities circa 1920.
-A representative savings institution was the postal savings (yūbin chokin).
-Postal savings was initially designed as a small savings institution for low-income households, and the share of postal savings to total savings was approximately $10$% in the 1910s (Tanaka 2018, p.
-29).
-However, the number of postal savings accounts per total population in the early 1910s exceeded $20$% (Tanaka 2018, p.
-33).
-James and Suto (2011) indeed show that the saving rate in prewar Japan was higher than that of the US.
-As shown in Panel A of Table A.2, approximately $1.4$ million people had accounts in Tokyo in 1920.
-This covers approximately $67$% of the entire population in Tokyo City and the average savings amount per capita was $64.5$ yen.
-In Kyobashi Ward, the average figure for the postal savings accounts was $44.5$ yen at the end of 1919 (Tokyo City Office 1921, pp.
-892–893).
-This amount was less than the average monthly earnings of the skilled-factory workers (Section 3.2).
-Unfortunately, the statistics on postal savings by depositors’ occupation are unavailable.
-[^81] There are some statistics for the whole of Japan.
-According to the Postal Savings Bureau (1924, p.
-72), the share of depositors in the manufacturing sector was $4.8$% (the share based on the deposit amount was $5.5$%).
-However, this implies that, while a large proportion of the workers might have had accounts, postal savings may not have provided sufficient temporary income.
-Savings bank (chochiku ginkō) was another popular savings institution for working-class households in the 1910s (Ito and Saito 2019, pp 79–81).
-Panel A of Table A.2 shows that the number of people who had savings accounts was approximately $1.46$ million, accounting for approximately $67.3$% of the total people in Tokyo City at that time.
-Although this figure is similar to the postal savings, the workers in the manufacturing industry were less likely to rely on the savings banks than the workers in the other industrial sectors.
-Table A.3 summarizes the number of depositors and the amounts of deposits in savings banks by occupation in Tokyo City (Panel A) and Kyobashi Ward (Panel B).
-In Kyobashi, the commerce and miscellaneous sectors shared approximately $84.8$% of all the depositors, whereas the manufacturing sector had only $11.5$%.
-This is unchanged if I use the amounts of deposits to calculate the shares.
-The Tokyo Institute for Municipal Research confirmed that savings banks were designed as financial institutions for low-income households (Tokyo Institute for Municipal Research 1925b, p.
-81).
-Thus, skilled factory worker households—like the THBS households—were less likely to use the savings banks.
-This is consistent with the fact that Tsukishima did not have any savings bank branches, while there were two postal offices.
-[^82] There were 123 savings banks in Tokyo City in 1922, and ten of them were in Kyobashi Ward (Tokyo Institute for Municipal Research 1925, p.
-76).
-However, Tsukishima—a representative manufacturing area—did not have savings banks at that time.
-Mutual loan association (mujin) was another type of savings institution.
-The members of this association paid a fixed amount of money several times, and each member withdrew the total amount of payments each time under specific rules.
-The lotteries or bidding system decided which members could withdraw money each time.
-This means that the workers could not use this institution to manage acute adverse income shocks.
-[^83] This is the most representative mutual loan system called tsūjyō (i.e., common) mujin.
-There was a different type called sueoki chokin (or tsumitate kai) from which people could receive their stakes with interests after a fixed amount had been accumulated for a certain period (Tokyo Legislative Research Association 1915, pp.
-3–4).
-However, these were also not flexible for compensating for income shocks.
-Instead, these were used by small business owners and merchants to generate money.
-A report suggests that the workers in the manufacturing sector shared only $10$% of all the mutual loan associations in Tokyo City (Tokyo City Office 1935, p.
-25).
-Moreover, the Tokyo Chamber of Commerce stated that the mutual loan association was used for business and rarely for household consumption (Tokyo Chamber of Commerce 1918, p.
-121).
-Cooperative societies were also available for worker households.
-Cooperative societies in Tokyo used credit as their payment method—households could buy goods in advance and pay at the end of the month or twice (in the middle and at the end of) a month.
-The commodity prices were similar to or slightly lower than those in the retailers (Central Federation of Industrial Associations 1925, pp.
-18–19).
-However, the cooperatives covered a small proportion of citizens.
-There were $26$ cooperative societies in Tokyo prefecture in 1924, and 23 of them were organized by the citizens and workers (Central Federation of Industrial Associations 1925, p.
-57–59).
-Among these $23$ cooperatives, two were large societies with approximately $4,000$ members, and the others were small societies with less than $100$ members.
-Given that the number of male workers was more than $768$ thousand (Table A.1), these account for only a few percentage points of the total workers in the city.
-In fact, the THBS households rarely used the cooperatives.
-An alternative borrowing institution was credit purchase in retailers.
-The details of the credit purchase are little documented, and the systematic statistics on credit purchases are also unavailable.
-I surmise that this was because the information on the transactions was personal.
-Despite this, a few available documents show that credit purchases were used in various retailers at that time (Section 5).
-The THBS households frequently used credit for purchases in various retailers for rice, fish, vegetables, alcohol, firewood, charcoal, and medicines.
-Pawnshops were the most popular lending institution among low-income working-class households.
-Lenders did not need to screen the borrower’s credit, and the borrowers did not worry about incurring heavy debt as the primary articles pawned were inexpensive clothes (Shibuya et al.
-1982).
-In addition, the interest rates were regulated by the Pawnbroker Regulation Act of 1895.
-Thus, the average redemption rate was substantially high under the lower interest rates.
-Among $88$ pawnshops in Kyobashi, the average redemption rate was approximately $94$% ($401,222$ cases/$427,265$ cases) in 1920 (Tokyo City Office 1922b, pp.
-888–889).
-The number of pawnshops in Tokyo was $1,334$ and $1,261$ in 1918 and 1919, respectively.
-The total number of cases was $8,226,883$ and $7,573,406$ each year.
-These figures were more than three times greater than the entire population in Tokyo City.
-This means that pawnshops were more accessible to workers than other institutions.
-As described in the main text, there were $13$ pawnshops in Tsukishima, whereas there were five moneylenders.
-However, the inexpensive articles pawned meant that pawnshops were available only for short-run necessities.
-Among the pawnshops of Kyobashi, for example, the average amount per case was $7.8$ yen, which was approximately $10$% of the monthly earnings of the THBS households.
-While the pawnshop was a popular lending institution, it could not offer enough temporary income for relatively large idiosyncratic shocks.
-According to the statistics on the pawnshop users in Tokyo prefecture in 1923, the share of users in the agriculture, commerce, and manufacturing sectors was $1.7$%, $19.5$%, and $15.6$%, respectively.
-In turn, the most representative users were the day laborers and workers classified as “miscellaneous,” accounting for $63.4$% (Tokyo Institute for Municipal Research 1926, p.
-25).
-In fact, the Tokyo City Office stated that the pawnshops were mainly used among low-income (poor) working-class households (Tokyo City Social Affairs Bureau 1921, p.
-9).
-This indicates the evidence that the primary pawnshop users were not the skilled-worker households like the THBS households.
-Another lending institution was the money lenders (kinsen kashitsuke gyō).
-Generally, business owners used money lenders.
-The interest rates of the money lenders were very high compared with those in the pawnshops (Shibuya 2000, pp.
-184; 248).
-Tokyo Institute for Municipal Research was indeed concerned about the increasing number of unscrupulous companies arrested by the National Police Agency (Tokyo Institute for Municipal Research 1925b, pp.
-106–108).
-Given this fact, while there were two money lenders in Tsukishima, they were less likely to be used by the factory worker households.
-Credit union (shinyō kumiai) was another institution available for working-class households.
-In Tokyo City, there were $215$ unions in 1923.
-According to the survey of the $184$ unions among them, however, the number of union members in the manufacturing sector was only $5,662$ (Tokyo Institute for Municipal Research 1925c, p.
-20).
-This accounts for only a few percentage points of the male workers in the manufacturing sector in Tokyo measured in the population census of 1920 (Online Appendix Table A.1).
-Table A.2: Financial Institutions in Tokyo, Osaka, Kyoto, and Nagoya Cities in 1920
+Table A.1: Financial Institutions in Tokyo, Osaka,
+Kyoto, and Nagoya Cities in 1920
 Name of cities
@@
 Amount per loan (yen) 6.40 1.87 3.54 9.33
-Notes:
-1. Panel A: The statistics on postal savings were measured at the end of March 1921. The statistics on savings banks were measured at the end of 1920.
-[…]
+Notes: Panel A: The statistics on postal savings were measured at the end of March 1921.
+The statistics on savings banks were measured at the end of 1920.
+Panel B: Pawnshops in Tokyo, Osaka, and Kyoto were measured in 1920.
+Those for Nagoya (suggested in $\dagger$}) were measured in 1923 because the statistics in the former years are unavailable.
 Sources: Panel A: Tokyo City Office (1923, pp.
@@
 12).
+A representative savings institution was postal savings (yūbin chokin).
+Panel A of Table A.1 shows that approximately $1.46$ million people held postal savings accounts in Tokyo in 1920, covering about 6767% of the city’s population.
+However, the average savings per capita was insufficient.
+In Kyobashi Ward, the average balance of postal savings accounts was only $44.5$ yen in 1919 (Tokyo City Office 1921, pp.
+892–893).
+This amount is lower than the average monthly earnings of skilled factory workers (Section 3.2), suggesting that postal savings may not have provided adequate temporary income.
+[^65] Although statistics on postal savings by depositors’ occupation are unavailable, some nationwide data exist.
+According to the Postal Savings Bureau (1924, p.
+72), depositors in the manufacturing sector accounted for 4.84.8% of the total, and their share based on deposit amounts was 5.55.5%).
+Savings banks (chochiku ginkō) were another common savings institution (Ito and Saito 2019, pp 79–81).
+Panel A of Table A.1 shows that the number of people with savings accounts was approximately $1.46$ million, representing approximately 6767% of Tokyo’s total population.
+Although this figure is similar to that for postal savings, savings banks were designed as financial institutions for low-income workers (Tokyo Institute for Municipal Research 1925b, p.
+81).
+[^66] There were 123 savings banks in Tokyo City in 1922, and 10 of them were in Kyobashi Ward (Tokyo Institute for Municipal Research 1925, p.
+76).
+However, Tsukishima—a representative manufacturing area—did not have savings banks at the time.
+Table A.2 summarizes the number of depositors and the amounts of deposits in savings banks by occupation in Tokyo City (Panel A) and Kyobashi Ward (Panel B).
+In Kyobashi, the commerce and miscellaneous sectors together accounted for about 84.884.8% of all depositors, while the manufacturing sector accounted for only 11.511.5%.
+These shares remain unchanged when calculated based on deposit amounts.
+Mutual loan associations (mujin) represent another type of savings institution.
+Members contribute a fixed amount of money over several periods, and, under specific rules, each member withdraws the total contributions once during the cycle.
+A lottery or bidding system determines withdrawal order, which means that workers could not rely on these associations to manage sudden adverse income shocks.
+[^67] This is the most representative mutual loan system called tsūjyō (i.e., common) mujin.
+There was a different type called sueoki chokin (or tsumitate kai) from which people could receive their stakes with interests after a fixed amount had been accumulated for a certain period (Tokyo Legislative Research Association 1915, pp.
+3–4).
+However, these were also not flexible for compensating for income shocks.
+Instead, they were mainly used by small business owners and merchants to generate revenue (Tokyo Chamber of Commerce 1918, p.
+121).
+Workers in the manufacturing sector accounted for only 1010% of all mutual loan associations in Tokyo City (Tokyo City Office 1935, p.
+25).
+Cooperative societies were also available for workers.
+In Tokyo, these societies used credit as their payment method; households could purchase goods in advance and settle their payments at the end of the month or in two installments, mid-month and month-end.
+Commodity prices in these cooperatives were similar to, or slightly lower than, those in retail shops (Central Federation of Industrial Associations 1925, pp.
+18–19).
+In 1924, Tokyo Prefecture had $26$ cooperative societies, of which 23 were organized by citizens and workers (Central Federation of Industrial Associations 1925, p.
+57–59).
+Among these $23$ cooperatives, two were large, with approximately $4,000$ members each, while the rest were small, with fewer than $100$ members.
+Given that the number of male workers exceeded $768$ thousand (Statistics Bureau of the Cabinet 1929a, pp.
+20–21; 26–29), these cooperatives accounted for only a small share of all workers in the city.
+Another borrowing option was purchasing on credit from retailers.
+Details on credit purchases are poorly documented, and systematic statistics are unavailable because transaction records are personal.
+Nevertheless, a few existing sources indicate that credit purchases were widely used across various retailers (Section 5).
+The THBS households frequently used credit to buy goods, including rice, fish, vegetables, alcohol, firewood, charcoal, and medicine, from different retailers.
 Table A.2: Deposits in the Savings Banks in Tokyo City by Occupations
@@
 Agriculture Commerce Manufacturing Miscellaneous Total
-Pane A: Entire Tokyo City
+Panel A: Entire Tokyo City
     Number of depositors 76,210 550,784 166,498 619,963 1,413,455
@@
     Amount of deposits (%) 4.2 40.2 14.0 41.6 100.0
-Pane B: Kyobahi Ward
+Panel B: Kyobashi Ward
     Number of depositors 3,809 34,665 11,664 51,481 101,619
@@
 Panel A and B list the statistics for Tokyo City and Kyobashi Word, respectively.
-Source: Tokyo City Office (1922a, pp. 864–865).
+Source: Tokyo City Office (1922a, pp.
+864–865).
+According to the Tokyo City Office, pawnshops were primarily used by low-income workers (Tokyo City Social Affairs Bureau 1921, p.
+9).
+Lenders did not need to assess borrowers’ creditworthiness, and borrowers faced limited risk of heavy debt because most pawned items were inexpensive clothes (Shibuya et al.
+1982).
+The average redemption rate was high, supported by relatively low interest rates regulated under the Pawnbroker Regulation Act of 1895.
+Among the $88$ pawnshops in Kyobashi, the average redemption rate was approximately 9494% (401,222401,222/$427,265$ cases) in 1920 (Tokyo City Office 1922b, pp.
+888–889).
+In Tokyo, overall, the number of pawnshops was $1,334$ in 1918 and $1,261$ in 1919, with total transactions of $8,226,883$ and $7,573,406$, respectively.
+These figures exceed three times the total population of Tokyo City, indicating that pawnshops were more accessible to workers than other financial institutions.
+Figure A.2 shows that Tsukishima had $13$ pawnshops compared to five moneylenders.
+However, because the items pawned were inexpensive, pawnshops primarily served short-term needs.
+For example, in Kyobashi, the average loan per transaction was $7.8$ yen, approximately 1010% of the monthly earnings of the THBS households.
+Statistics on pawnshop users in Tokyo Prefecture in 1923 show that 1.71.7% worked in agriculture, 19.519.5% in commerce, and 15.615.6% in manufacturing.
+The largest group of users consisted of day laborers and workers classified as “miscellaneous,” accounting for 63.463.4% (Tokyo Institute for Municipal Research 1926, p.
+25).
 Figure A.2: Pawnshops and Money Lenders
@@
 The white circle shows the rickshaw station (kuruma yado).
-Source: Department of Health, Ministry of the Interior 1923c, tenth map. The tone was adjusted by the author using Adobe Photoshop 24.7.0.
+The upper right corner indicates north.
+Source: Department of Health, Ministry of the Interior 1923c, tenth map.
+The tone was adjusted by the author using Adobe Photoshop 24.7.0.
+Moneylenders were another lending institution(kinsen kashitsuke gyō).
+They were generally used by business owners, and their interest rates were significantly higher than those charged by pawnshops (Shibuya 2000, pp.
+184; 248).
+The Tokyo Institute for Municipal Research expressed concern over the growing number of unscrupulous moneylenders, many of whom were arrested by the National Police Agency (Tokyo Institute for Municipal Research 1925b, pp.
+106–108).
+Credit unions (shinyō kumiai) were another option available to working-class households.
+In Tokyo City, there were $215$ unions in 1923.
+However, according to a survey of $184$ of these unions, the number of members from the manufacturing sector was only $5,662$ (Tokyo Institute for Municipal Research 1925c, p.
+20).
+This represented a small share of male workers in the manufacturing sector, as measured in the 1920 Population Census.
 A.3 Development of Tsukishima
-Figure A.3: Tsukishima: Ishikawajima, Tsukudajima, Shin-tsukudajima, and Lots 1–3
+Figure A.3: Administrative Wards in Tokyo City
+Notes: The name in each lattice indicates the administrative ward in Tokyo.
+The gray ward indicates the Kyobashi Ward and an island in the Kyobashi Ward indicates Tsukishima.
+The border surrounded by Asakusa, Honjyo, Nihonbashi, and Fukagawa wards includes the Sumida River.
+The borders of each ward are based on the administrative districts in 1920.
+Geographical coordinate system is based on JGD2000/(B, L).
+Source: Created by the author using the official shapefile (Ministry of Land, Infrastructure, Transport and Tourism, database).
+Figure A.4: Map of Tsukishima: Ishikawajima, Tsukudajima, Shin-tsukudajima, and Lots 1–3
 Notes: Tsukishima contains Ishikawajima, Tsukudajima, Shin-tsukudajima, and Tsukishima Lots 1–3.
@@
 “Aioi bridge” links Shin tsukudajima and Fukagawa ward.
-Source: Department of Health, Ministry of the Interior 1923c, 15th map. The author modified and adjusted the tone using Adobe Photoshop 24.7.0.
-Tsukishima (transliterated as “moon island”) is a reclaimed land on a sandbar of the Sumida River.
-Ishikawajima and Tsukudajima islands at the northern end of Tsukishima (Figure A.3) are the starting point of its history.
-The following description in this section is based on Chūō Ward (1994, pp.
+Source: Department of Health, Ministry of the Interior 1923c, 15th map.
+The author modified and adjusted the tone using Adobe Photoshop 24.7.0.
+Tsukishima (translated as “moon island”) is a reclaimed area on a sandbar of the Sumida River (Figure A.3).
+[^68] The following description in this section is based on Chūō Ward (1994, pp.
 122–124).
-Ishikawajima was built in 1626 by a boatman, Hachizaemon Ishikawa, who received a fief from the Shogunate.
-In 1790, the Ishikawajima Prison (ninsoku yoseba) was built on Ishikawajima.
-This was a facility for the reintegration of minor offenders and those without lodgings.
-After Matthew Perry arrived in Japan in 1853, the Shogunate ordered the Mito Clan to build a shipyard on Ishikawajima.
-A month later, a groundbreaking ceremony was held for a Western-style wooden sailing ship delivered to the Shogunate in 1856.
-In 1866, the Ishikawajima Beacon was built on the west bank of the Ishikawajima Prison and served to signal the position of ships entering the Nihonbashi direction.
-In 1870, the Ishikawajima Prison became a place of imprisonment (toba) and was renamed the Ishikawajima Jail Station of the Metropolitan Police Department in 1877.
-The prison was finally moved to Sugamo in Kitatoshima County in Tokyo prefecture, and the penitentiary on Ishikawajima was closed in 1895.
-The area adjacent to Ishikawajima is called Tsukudajima, where fishermen from “Tsukuda” in Osaka prefecture settled in 1644 after receiving land from the Shogunate.
-In 1645, the “Tsukuda ferry” began, and throughout the Edo period, Tsukudajima was known as a fishing island for whitefish for the Shogun’s household.
-In 1863, the Tsukudajima Battery was built on the south side of Tsukudajima—one of $13$ batteries used to defend Edo Bay.
-This battery later became the starting point for the reclamation of Lot 1 of Tsukishima.
-In 1884, the Port of Tokyo Waterway Dredging Project was initiated.
-This was to secure the passage of ships by sweeping away the sediment in the shipping channel, and the Tsukishima landfill was created using the deposit.
-The first Tsukishima site was completed in 1891, and the “Tsukishima ferry” began the same year.
-The second site was created in 1894; the Shin Tsukudajima (namely, new Tsukudajima) was built in 1896.
-From the west shore of Shin Tsukudajima, the island overlooks the Boso Peninsula in Chiba Prefecture, making it a favorite vacation spot for writers.
-Cultural figures such as Toson Shimazaki and Kaoru Osanai stayed here from 1907 to the early Taisho period.
-In 1903, the Aioi Bridge was completed, bringing water and electricity to Tsukishima and improving transportation convenience, which contributed to development of industry.
-[^84] Until then, a retailer named Suzukiiya had been selling water to supply the entire island using a ferryboat.
-The “Kachidoki Ferry” began to deliver to Lot 2 in 1905.
-Lot 3 was finally completed in 1913.
+The islands of Ishikawajima and Tsukudajima at the northern end mark the starting point of Tsukishima’s territory (Figure A.4).
+Ishikawajima was established in 1626 by Hachizaemon Ishikawa, a boatman who received a fief from the shogunate.
+In 1790, Ishikawajima Prison (ninsoku yoseba) was constructed to accommodate minor offenders and homeless individuals, providing a facility for reintegration into society.
+After Matthew Perry’s arrival in Japan in 1853, the shogunate instructed the Mito Clan to construct a shipyard on Ishikawajima.
+A groundbreaking ceremony was held a month later, and a Western-style wooden sailing ship was delivered to the shogunate in 1856.
+In 1866, the Ishikawajima Beacon was built on the west bank of the prison to signal the position of ships entering the Nihonbashi area.
+In 1870, the prison was converted into an official place of imprisonment (toba) and was renamed the Ishikawajima Jail Station of the Metropolitan Police Department in 1877.
+The facility was eventually relocated to Sugamo in Kitatoshima County, Tokyo Prefecture, and the penitentiary on Ishikawajima was closed in 1895.
+Adjacent to Ishikawajima lies Tsukudajima, where fishermen from Tsukuda in Osaka Prefecture settled in 1644 after receiving land from the shogunate.
+In 1645, the Tsukuda Ferry began operations.
+Throughout the Edo period, Tsukudajima became known as a fishing hub, supplying whitefish for the shogun’s household.
+In 1863, the Tsukudajima Battery was constructed on the southern side of the island as one of $13$ batteries built to defend Edo Bay.
+This site later became the starting point for the reclamation of Lot 1 of Tsukishima.
+In 1884, the Port of Tokyo Waterway Dredging Project began to secure navigable passages by clearing sediment from the shipping channel, and the reclaimed land of Tsukishima was created using the dredged deposits.
+Lot 1 was completed in 1891, and the Tsukishima Ferry began operations the same year.
+Lot 2 was reclaimed in 1894, and Shin-tsukudajima (literally, new Tsukudajima) was established in 1896.
+From the western shore of Shin-tsukudajima, the island offers views of the Bōsō Peninsula in Chiba Prefecture, making it a favored retreat for writers.
+Cultural figures such as Tōson Shimazaki and Kaoru Osanai stayed there from 1907 through the early Taishō period.
+In 1903, the completion of the Aioi Bridge improved transportation access and brought water and electricity to Tsukishima, contributing to the growth of the local industry.
+The Kachidoki Ferry began serving Lot 2 in 1905, and Lot 3 was finally completed in 1913.
 A.4 Housing in Tsukishima
-The Tsukishima Survey included a housing survey (Table B.1).
-Table A.4 summarizes the housing types measured in this survey.
-As shown, $88$ ($46$%) of the detached homes are in Tsukudajima.
+Table A.3 summarizes the housing types recorded in the housing survey (Table B.1).
+As shown, $88$ detached homes (4646%) were located in Tsukudajima.
 Tsukudajima is reported to be “superior to other areas in terms of housing quality, and the percentage of inferior housing is meager.” [^69] The description in this section is based on the Department of Health, Ministry of the Interior (1923a, pp.
 69–76).
-Shin-tsukudajima had a low percentage of detached homes.
-The report states that it was indeed “inhabited by unskilled laborers” and “occupied the most inconvenient housing status of all the islands.” Lot 1 was the center of life in Tsukishima and was a cluster of numerous residences, mainly tenements.
-Lot 2 had a relatively high percentage of single-family and two-family houses and appears to be characterized as a “workers’ residential area” than Lot 1.
-These indicate that there might have been some geographic differences in income distribution regarding housing type.
-Overall, however, of the $989$ dwellings in Tsukishima, $799$ ($81\%$) were row houses, and only $190$ were single-family homes.
-This confirms that Tsukishima was an industrial area with many working families.
-The workers lived in similar tenements and with similar standards of living.
-Figure A.4 is a photo showing these tenement houses.
+By contrast, Shin-tsukudajima had a low percentage of detached homes.
+According to the report, it was “inhabited by unskilled laborers” and “occupied the most inconvenient housing status of all the islands.” Lot 1, considered the center of life in Tsukishima, consisted mainly of numerous tenement-style residences.
+Lot 2, by comparison, had a relatively high percentage of single-family and two-family houses and appears to have been characterized more as a “workers’ residential area” than Lot 1.
+These findings suggest possible geographic differences in income distribution reflected in housing type.
+Despite this, of the $989$ dwellings in Tsukishima, $799$ ($81\%$) were row houses, while only $190$ were single-family homes.
+This indicates that most workers lived in similar tenements and experienced comparable living standards.
+Figure A.5 shows a photo of these tenement houses.
 Table A.3: Housing Types in Tsukishima
@@
 Figure A.5: Photographs of the Tenements in Tsukishima
-Note: Figures A.4a and A.4b show the two- and one-story tenements in Tsukishima, respectively.
-Source: Department of Health, Ministry of the Interior 1923c, figures 62–63; 66.
+Note: Figures A.5a and A.5b show the two- and one-story tenements in Tsukishima, respectively. Source: Department of Health, Ministry of the Interior 1923c, figures 62–63; 66.
 A.5 Machinery Factories in Tsukishima
-First, I used the results of the Primary School Children Survey included in the Tsukishima Survey to analyze the occupation (industry) of the household heads in Tsukishima.
-This is a survey of 4th-6th graders at Tsukishima 1st and 2nd Elementary Schools in July 1919.
-There were $1,173$ children in these grades, and the survey included all $1,100$ children who resided in Tsukishima.
-At that time, there were no other elementary schools.
-Thus, this survey is a complete survey for all children in grades 4–6 in Tsukishima.
-Although it does not include households without children in grades 4–6, it is undoubtedly a beneficial survey that provides an overview of the households with children—the main target of this study.
-Panel A of Table A.5 shows the industrial structures.
-The composition ratios are similar to the census-based statistics reported in Panel A of Table 1.
-It does not specify whether brothers and sisters in the same household were counted twice.
-Despite this, the similarity confirms that the statistics obtained from this survey are useful to identify overall trends in family households in Tsukishima.
-Panel B of Table A.5 divides the “social classes” of the household heads working in the manufacturing sector into six categories.
-In the report, large business owners were classified into the capitalist class; small business owners, free enterprise owners, and directors were regarded as the middle class; workers were defined as the working class (Department of Health, Ministry of the Interior 1923a, p.
+First, I use the results of the Primary School Children Survey, included in the Tsukishima Survey, to analyze the occupations of household heads.
+This survey was conducted among 4th–6th graders at Tsukishima 1st and 2nd Elementary Schools in July 1919.
+A total of $1,173$ children were in these grades, and the survey covered all $1,100$ children residing in Tsukishima.
+Since no other elementary schools existed, this represents a complete survey of all 4th–6th graders in Tsukishima.
+Although it excludes households without children in these grades, it remains a valuable source for understanding households with children.
+Panel A of Table A.4 presents the industrial structures.
+The composition ratios are similar to the census-based statistics shown in Panel A of Table 1.
+The report does not clarify whether siblings in the same household are counted more than once.
+Nevertheless, the similarity between the results confirms that this survey provides reliable statistics for identifying overall trends in family households in Tsukishima.
+Panel B of Table A.4 classifies the “social classes” of household heads working in the manufacturing sector into six categories.
+According to the report, large business owners are classified as the capitalist class; small business owners, free enterprise owners, and directors are grouped as the middle class; and workers are defined as the working class (Department of Health, Ministry of the Interior 1923a, p.
 50).
-Most heads of households were workers, with only 1% of the capitalist class and less than 20% of the middle class.
-While the definition of “executive” is unclear, “small business owners” must be small factory owners who could have had similar earnings to the skilled worker households (Department of Health, Ministry of the Interior 1923a, p.
+Most household heads were workers, with only 1% in the capitalist class and less than 20% in the middle class.
+Although the definition of “executive” is unclear, “small business owners” appear to refer to small factory owners whose earnings are similar to those of skilled worker households (Department of Health, Ministry of the Interior 1923a, p.
 64).
@@
 Panel A summarizes the heads’ industries, and Panel B sorts the “worker type” of the heads in the manufacturing sector, respectively.
-Source: Department of Health, Ministry of the Interior 1923a, pp. 386–387.
-In November 1920, there were $214$ factories in Tsukishima, $168$ of them (approximately $80$%) were classified as machinery factories.
-The official report documented that the wooden pattern factories could be included in the machinery sector because they were incorporated into the casting processes (Department of Health, Ministry of the Interior 1923a, p.
+Source: Department of Health, Ministry of the Interior 1923a, pp.
+386–387.
+In November 1920, $214$ factories existed in Tsukishima, of which $168$ (approximately 8080%) were classified as machinery factories.
+The official report notes that wooden pattern factories could be included in the machinery sector because they were part of the casting processes (Department of Health, Ministry of the Interior 1923a, p.
 388).
-The share of machinery factories becomes approximately $90$% if I include those $24$ wooden pattern factories.
-Panel A of Table A.6 uses the manufacturing survey in the Tsukishima Survey to classify the $168$ machinery factories by size and industrial sector.
-First, more than $90$% of the plants are in the three sectors of machinery and equipment manufacturing, can manufacturing, and blacksmithing, with only $15$ plants in the casting sector.
-$113$ ($67$%) of the machinery factories have less than $15$ workers.
-Second, many smithing factories have fewer than $15$ workers.
-Hence, almost all of Tsukishima’s smithing industry was classified as small businesses.
-[^86] The report suggests that $43$ of the $47$ smithing factories were “town factories” or “workplaces.” The definition of “town factory” is not clear in the report (p.
-392).
-Since they were always counted more than the number of factories with 15 or fewer workers, they are likely to include smaller factories operated as small businesses.
-Panel B of Table A.6 classifies the types of the $620$ workers based on the results of the primary school survey used in Table A.5.
-There are $455$ skilled and $165$ unskilled workers.
-The same report shows that $490$ out of $620$ workers are industrial workers.
-The day laborers are classified as non-industrial workers.
-In fact, $620$ workers minus $120$ day laborers become $490$, which is identical to the number of industrial workers documented.
-Therefore, approximately $90$% of factory workers ($455/490$) can be regarded as skilled workers in the industrial sector: $65$% work in factories, $17$% in cottage industries, and $18$% are craftsmen.
-Since workers in cottage and handicraft industries are classified as working in small factories, $65$% of these workers might have worked in the middle- to large-scale factories (with $15+$ workers).
-Notably, this survey only includes household heads with primary school children aged 9–11 (i.e., 4th to 6th grades), suggesting that most of these heads are in their thirties, in the skilled workers age range.
-Therefore, the proportion of skilled workers to the entire worker group could be smaller than the $90$%.
-Nevertheless, this confirms that most of the Tsukishima workers in their 30s were skilled workers at that time.
+Including the $24$ wooden pattern factories raises the share of machinery factories to approximately 9090%.
+Panel A of Table A.5 uses data from the manufacturing survey in the Tsukishima Survey to classify the $168$ machinery factories by size and industrial sector.
+First, more than 9090% of the plants were concentrated in three sectors: machinery and equipment manufacturing, can manufacturing, and blacksmithing, with only $15$ plants in the casting sector.
+Among the machinery factories, $113$ (6767%) employed fewer than $15$ workers.
+Second, many smithing factories also had fewer than $15$ workers.
+Hence, almost the entire smithing industry in Tsukishima was classified as small businesses.
+[^70] The official report indicates that $43$ of the $47$ smithing factories were small factories and workplaces called (machi kōba).
+The report does not clearly define machi kōba.
+However, since their number consistently exceeds the count of factories with 15 or fewer workers, they likely include smaller factories operated as small businesses (Department of Health, Ministry of the Interior 1923a, pp.
+392; 410–411).
+Panel B of Table A.5 classifies $620$ workers based on the primary school survey data used in Table A.4.
+Of these, $455$ are skilled workers and $165$ are unskilled.
+Among the unskilled workers, 45 worked in factories and cottage industries, while 120 were day laborers considered low-income workers.
+According to the same report, $490$ of the $620$ workers were industrial workers.
+Day laborers are classified as non-industrial workers.
+This is consistent with the calculation: $620$ workers minus $120$ day laborers equals $490$ industrial workers.
+Therefore, approximately 9090% of industrial workers ($455/490$) can be regarded as skilled workers: 6565% worked in factories, 1717% in cottage industries, and 1818% were craftsmen.
+Since workers in cottage and handicraft industries are classified as working in small factories, roughly 6565% of these skilled workers likely worked in medium- to large-scale factories (with $15+$ workers).
+Notably, this survey includes only household heads with primary school children aged 9–11 (i.e., 4th to 6th grades).
+This suggests that most household heads were in their thirties, an age range typically associated with skilled workers.
+Therefore, the share of skilled workers among the entire working population may be lower than 9090%.
+Nevertheless, the data confirm that most Tsukishima workers in their 30s were skilled.
 Table A.5: Scale of Machinery Factories and Type of Workers in Tsukishima
@@
 # of workers $\geq 15$ 22 15 10 8 55
-# of workers $<15$ 44 25 5 39 113
+# of workers <15<15 44 25 5 39 113
 Total 66 40 15 47 168
@@
 The threshold of the scale (i.e., 15 workers) is based on the application of the Factory Act enacted in 1916.
-Panel B sorted the heads by their skill level and job type. Statistics are obtained from the Complete Survey on the 4th–6th graders in Tsukishima1st and 2nd Primary Schools in July 1919.
-Sources: Panel A: Department of Health, Ministry of the Interior 1923a, pp. 386–387; 410–414. Panel B: Department of Health, Ministry of the Interior 1923a, p. 149.
+Panel B sorted the heads by their skill level and job type.
+Statistics are obtained from the Complete Survey on the 4th–6th graders in Tsukishima 1st and 2nd Primary Schools in July 1919.
+Sources: Panel A: Department of Health, Ministry of the Interior 1923a, pp.
+386–387; 410–414.
+Panel B: Department of Health, Ministry of the Interior 1923a, p.
+149.
 A.6 Works in the Machinery Factories
-Tsukishima Survey includes the factory survey that counts the number of machinery factories in Tsukishima.
-Panel A of Table A.6 summarizes the number of machinery factories by type of plant.
-The machinery and equipment manufacturing factories shared $39$% of all the machinery factories.
-Typical jobs for workers in machinery and equipment manufacturing plants are finishers (or fitters) and turners.
-A finisher is “a skilled worker responsible for finishing, assembling, and commissioning machine components” and is “one of the finest skilled workers because it requires skills in all aspects to correct defects in machine making” (Department of Health, Ministry of the Interior 1923a, p.
+Panel A of Table A.5 summarizes the number of machinery factories by plant type.
+Machinery and equipment manufacturing factories accounted for 3939% of all machinery factories.
+Typical jobs in these factories included finishers (or fitters) and turners.
+A finisher is defined as “a skilled worker responsible for finishing, assembling, and commissioning machine components” and is considered “one of the finest skilled workers because the role requires proficiency in all aspects to correct defects in machine making” (Department of Health, Ministry of the Interior 1923a, p.
 393).
-The turner is the “middle axis of machinists,” to which most semi-skilled and skilled workers belong.
-Becoming a full-fledged finisher and lathe operator takes at least five years.
-Generally, apprentices enter the workforce at approximately 14–18, but they do not work in the same factory for many years within the apprenticeship system.
-After learning the skills for 1–2 years, they often move to another factory to improve their skills (Department of Health, Ministry of the Interior 1923a, p.
+The turner is described as the “middle axis of machinists,” a category that includes most semi-skilled and skilled workers.
+Becoming a fully qualified finisher or lathe operator typically took at least five years.
+Apprentices usually entered the workforce between ages $14$ and $18$, but did not remain in the same factory throughout their apprenticeship.
+After acquiring basic skills for 1–2 years, they often moved to another factory to further develop their expertise (Department of Health, Ministry of the Interior 1923a, p.
 394).
-Wages are usually paid on a piece-work basis or on a time basis.
-In Tsukishima, a time contract is broadly adopted, and payment is based on the “residual time” method.
-The “residual time” method means that “if the work is completed within the contracted time, the remaining time is considered a bonus.” For example, if $12$ hours are contracted and the work is completed in $10$ hours, $2$ hours are paid as a bonus (pp.
+Wages were generally paid either on a piece-work or time basis.
+In Tsukishima, time contracts were widely adopted and payment followed the “residual time” method, whereby “if the work is completed within the contracted time, the remaining time is treated as a bonus.” For example, if $12$ hours are contracted but the work is completed in $10$ hours, the remaining $2$ hours are paid as a bonus (Department of Health, Ministry of the Interior 1923a, pp.
 396-397).
-Wages are paid by dividing one month into two periods (the first and second half of the year).
-The average work hours are reported to be approximately ten hours, with occasional all-night work (p.
+Wages were distributed twice a month, divided into two payment periods.
+Average working hours were reported to be approximately $10$ hours per day, with occasional all-night shifts (Department of Health, Ministry of the Interior 1923a, p.
 402).
-Overtime hours are approximately two hours per day (Department of Health, Ministry of the Interior 1923a, p.
+Overtime averaged approximately two hours per day (Department of Health, Ministry of the Interior 1923a, p.
 402).
-Next, I summarize labor in the canning factory, which accounts for $24$% of all the factories.
-The canning factory is engaged in shipbuilding, bridges, steel frames for construction, and other works, including fabrication and repair of gas reservoirs (p.
+Next, I provide a summary of the workforce in the canning factories, which accounted for 2424% of all factories.
+Canning factories were engaged in shipbuilding, constructing bridges, manufacturing steel frames for construction, and other work, including the fabrication and repair of gas reservoirs (Department of Health, Ministry of the Interior 1923a, p.
 406).
-Workers in a canning factory are called “boiler makers” and are typically involved in bending steel plates and jointing them with rivets.
-Workers are divided into rivetter, holder-up, plater, angle-ironsmith, and caulking.
-The work is usually done in groups, with the skilled angle-ironsmith and caulking.
-The “bōsin” of the rivetter gives the instructions as the head of the group.
-Wages are paid on a piece-rate contract basis.
-Because of the contracted wage, the cannery workers’ income is unstable; thus, they are paid a fixed wage (daily wage amount) during the off-season.
-Smithing is the second most common industry, accounting for $28$% of the total.
-Panel A of Table A.6 reveals that most of the smithing industry is relatively small in scale.
-They make boulders, nuts, rivets, lathe stops, jigs, and spanners for can manufacturing and construction.
-They are subcontractors who manufacture metal parts for the larger factories in Tsukishima (Department of Health, Ministry of the Interior 1923a, p.
+Workers in these factories were called “boiler makers” and typically bent steel plates and joined them with rivets.
+They were divided into rivetter, holder-up, plater, angle-ironsmith, and caulker.
+Work was usually performed in groups, led by skilled angle-ironsmiths and caulkers.
+The “bōsin” of the rivetter served as the group leader and provided instructions.
+Wages were paid on a piece-rate basis, with a fixed daily wage during the off-season.
+Smithing was the second most common industry, accounting for 2828% of the total.
+Panel A of Table A.5 shows that most smithing operations were relatively small.
+Workers produced boulders, nuts, rivets, lathe stops, jigs, and spanners for can manufacturing and construction.
+These small shops often acted as subcontractors supplying metal parts to larger factories (Department of Health, Ministry of the Interior 1923a, p.
 412).
-There are three types of jobs: yokoza, sakite, and tatara-fuki (bellows-blower).
-The yokoza is a self-employed skilled worker (called oyakata), while the other two are young unskilled laborers.
-In small operations, wages are based on a percentage based on voluntary business discussions with the proprietor, and working hours are not regular (pp.
+Three main jobs existed: yokoza, sakite, and tatara-fuki (bellows-blower).
+Yokoza referred to self-employed skilled workers (oyakata), while the other two referred to young, unskilled laborers.
+In small operations, wages were usually determined as a percentage through voluntary agreement with the proprietor, and working hours were irregular (Department of Health, Ministry of the Interior 1923a, pp.
 411–413).
@@
 B.1 Tsukishima Survey: The First Social Survey in Japan
-According to Kawai (1980), the methodology of social research in Japan was formed in the 1930s.
-During the 1910s and 1920s, “numerous social surveys were conducted by various survey entities, including government agencies, research groups, local administrative bodies, private organizations, survey research institutes, and researchers.” He saw that circa WWI, there were many problems in labor, lifestyle, urban, health and sanitation, public safety, and citizens due to the heavy and chemical industrialization, the monopolization of capitalism, and urban expansion (p.48).
-The Report on Field Survey in Tsukishima, Kyobashi Ward, Tokyo (called Tsukishima Survey) [^87] Takano himself, who was responsible for overseeing the survey, refers to this field survey as the “Tsukishima Survey” (Department of Health, Ministry of the Interior 1923a, p.
+According to Kawai (1980), the methodology of social research in Japan was established in the 1930s.
+During the 1910s and 1920s, “numerous social surveys were conducted by various entities, including government agencies, research groups, local administrative bodies, private organizations, survey research institutes, and individual researchers.” The Report on Field Survey in Tsukishima, Kyobashi Ward, Tokyo (called the Tsukishima Survey) represents a pioneering survey of urban communities (Sekiya 1970, p.
+43).
+First, I summarize the investigators, research questions, and subjects of the Tsukishima Survey.
+[^71] Takano himself refers to this field survey as the “Tsukishima Survey” (Department of Health, Ministry of the Interior 1923a, p.
 6).
 Considering this, I use the Tsukishima Survey as the name of this survey throughout this paper.
-used in this study is the most pioneering survey of urban communities (Sekiya 1970, p.
-43).
-First, I summarize the investigators, research questions, and subjects of the Tsukishima Survey.
-The person responsible for the Tsukishima Survey was Iwasaburō Takano (hereafter “Takano”), a professor of social statistics and social policy at the Law School of Tokyo Imperial University.
-Takano became a professor of economics at Tokyo Imperial University in 1919, and became the director of the OISR in 1920.
-Takano placed great importance on gathering objective statistics about workers through social surveys (Sekiya 1970, p.
+The person responsible for the survey was Iwasaburō Takano (hereafter “Takano”), a professor of social statistics and social policy at the Law School of Tokyo Imperial University.
+Circa WWI, Takano observed widespread problems in labor, lifestyle, urban infrastructure, health and sanitation, public safety, and citizen welfare, driven by heavy and chemical industrialization, capitalist monopolization, and urban expansion.
+In this context, Takano emphasized the importance of gathering objective statistics about workers through social surveys (Sekiya 1970, p.
 33).
-He had the experience of conducting the first household survey in Japan in 1916888888This is called the “Survey of Household Income and Expenditure of 20 Artisans and Workers in Tokyo,” which was inspired by Charles Booth (Miyoshi 1980, p.
+He conducted Japan’s first household survey [^72] This is called the “Survey of Household Income and Expenditure of 20 Artisans and Workers in Tokyo,” which was inspired by Charles Booth (Miyoshi 1980, p.
 33).
@@
 3).
-He was a member of the seventh subcommittee (Rural Sanitation Subcommittee) of the Ministry of Home Affairs’ Health and Sanitation Investigation Committee since 1916.898989The Health and Sanitation Investigation Committee of the Ministry of Home Affairs was established on June 27, 1916, under the second Shigenobu Okuma Cabinet by issuing Imperial Ordinance No.
+and had served on the seventh subcommittee (Rural Sanitation Subcommittee) of the Ministry of Home Affairs’ Health and Sanitation Investigation Committee since 1916.737373The Health and Sanitation Investigation Committee of the Ministry of Home Affairs was established on June 27, 1916, under the second Shigenobu Okuma Cabinet by issuing Imperial Ordinance No.
 172.
@@
 On October 22, 1918, Takano submitted a proposal to the subcommittee for a field survey of urban sanitary conditions.
-The survey aimed to determine the housing conditions, household budget status, and health status (child health, production, stillbirths, deaths, and diseases) in the areas where skilled-worker families congregated (Sekiya 1970, p.
+The survey aimed to determine the housing conditions, household budgets, and health status (including child health, live births, stillbirths, deaths, and diseases) in areas where skilled-worker families resided (Sekiya 1970, p.
 6).
-Initially, the survey area was planned to be Yanagibashi Yokogawa-chō in Honjo Ward.
-However, at a committee meeting on November 23, the survey site was changed to Tsukishima in Kyobashi Ward, which had more skilled worker households than in Yokogawa-chō (Kawai 1980, pp.
+Initially, the survey area was planned as Yanagibashi Yokogawa-chō in Honjo Ward.
+However, at a committee meeting on November 23, the survey site was changed to Tsukishima in Kyobashi Ward, which had more skilled-worker households than Yokogawa-chō (Kawai 1980, pp.
 55–59).
@@
 Finally, Takano and Yamana conducted an on-site inspection there on October 30 and decided on Tsukishima as the survey site.
-Although the statistical sampling method was not yet common at that time, Takano selected the survey sites based on his visionary thinking.
-He stated that “there can be no dispute that we have chosen Tokyo as the target of our metropolitan survey.
+Although statistical sampling was not yet common at the time, Takano selected the survey site based on his judgment.
+He stated: “There can be no dispute that we have chosen Tokyo as the target of our metropolitan survey.
 However, it is impossible to survey the entire Tokyo area.
 We have no choice but to conduct a partial survey.
-For this purpose, it would be the best idea to select a representative area in Tokyo and survey that area narrowly but deeply.” (Department of Health, Ministry of the Interior 1923a, pp.
+For this purpose, it would be the best idea to select a representative area in Tokyo and survey that area narrowly but deeply” (Department of Health, Ministry of the Interior 1923a, pp.
 2–3).
-Takano further explained the need to focus on the working-class, which makes up most of the citizens, rather than to see the political and commercial center: “Most of Tokyo citizens are wage earners who subsist on labor.
+Takano further emphasized the importance of focusing on the working class, which constitutes the majority of citizens, rather than the political and commercial center: “Most of Tokyo citizens are wage earners who subsist on labor.
 As is usual in large cities, a large number of people are engaged in manual labor in Tokyo.
@@
 Therefore, the area where many families of skilled workers congregate can be regarded as a representative area of Tokyo.
-Further, it seemed to me that a study of the social and sanitary conditions in that area could explain a large part of the social and sanitary conditions in Tokyo.” (Department of Health, Ministry of the Interior 1923a, pp.
+Further, it seemed to me that a study of the social and sanitary conditions in that area could explain a large part of the social and sanitary conditions in Tokyo” (Department of Health, Ministry of the Interior 1923a, pp.
 3).
-It is necessary to investigate the point where Takano states that “the most genuine type of worker is the skilled workers” and the fact that the main target of the Tsukishima Survey was skilled workers.
+Examining Takano’s statement that “the most genuine type of worker is the skilled worker” and the fact that the main target of the Tsukishima Survey was skilled workers is important.
 [^75] In other words, the Tsukishima Survey was not designed for poor households.
@@
 42).
-The description by Kawai (1981a, pp.
+Kawai’s description (1981a, pp.
 13–18) is helpful in this regard.
-As summarized in Section 2, Japan’s mechanical and chemical industrialization throughout WWI led to a large concentration of city workers.
-As a result, the ratio of small-scale conventional factories based on “master employees” (as is often observed in the textile industry) decreased.
-Conversely, the number of factories with a direct “factory owner-employee” employment relationship increased, particularly in the machinery and equipment, metal, and chemical industries.
-These workers in the modern industrial sector accumulated skills, formed households, and settled in the cities.
-It can be said that a new urban worker lifestyle was being formed.
-In short, although the city had a class of workers engaged in traditional industries, Takano believed that the skilled workers in the modern industrial sector would play a central role in terms of the labor force, labor movement, and urban life.
-In addition, Sekiya notes the rise of the labor movement.
-As explained in Section 2, while wages rose throughout WWI, inflation caused relative wages to decline.
-In 1916, Japan’s first labor protection legislation, the Factory Law, was enacted.
-However, the number of working days and hours did not decrease.
-Consequently, the number of labor disputes steadily increased (Sekiya 1970, pp.
+As summarized in Section 2, Japan’s mechanical and chemical industrialization during WWI led to a large concentration of city workers.
+Consequently, the ratio of small-scale conventional factories based on “master employees” (as commonly seen in the textile industry) declined.
+By contrast, the number of factories with a direct “factory owner–employee” relationship increased, particularly in the machinery and equipment, metal, and chemical industries.
+Workers in the modern industrial sector acquired skills, formed households, and settled in cities, giving rise to a new urban worker lifestyle.
+Although the city still had a class of workers engaged in traditional industries, Takano considered skilled workers in the modern industrial sector to be central in terms of labor force, labor movements, and urban life.
+Sekiya also notes the rise of the labor movement.
+As explained in Section 2, although wages rose during WWI, inflation caused relative wages to decline.
+In 1916, Japan enacted its first labor protection legislation, the Factory Law.
+However, the number of working days and hours remained unchanged, leading to a steady increase in labor disputes (Sekiya 1970, pp.
 24–26).
-Given this, Takano believed that workers should organize labor unions to help themselves with labor problems that could not be resolved through labor protection legislation.
+Takano therefore believed that workers should organize labor unions to address problems not resolved through labor protection legislation.
 Sekiya (1970, pp.
-26–40) speculates that Takano’s intention in conducting the Tsukishima Survey was to obtain basic data for solving labor problems, specifically to identify the scope of matters to be solved by workers on their own initiative (or covered by labor protection legislation).
+26–40) speculates that Takano’s purpose in conducting the Tsukishima Survey was to obtain basic data for solving labor issues, specifically to identify problems that workers could address independently or through existing labor legislation.
 Kawai (1981b, pp.
-15; 17) assessed the $40$ households covered by the household survey included in the Tsukishima Survey as “the representativeness of the sample has remained ambiguous.” There are no systematic descriptions of how the sample households were characterized among the worker households in Tsukishima.
-The official report stated that “to find suitable persons to fill in the form, we must ask for the assistance of school principals, police officers, physicians, and laborers, and we have requested a meeting of these persons at the survey office to explain the purpose of the survey, and to solicit applicants for the distribution of the household account book” (Department of Health, Ministry of the Interior 1923a, pp.
+15; 17) assesses the $40$ households covered by the Tsukishima Survey as having “ambiguous representativeness.” No systematic description exists of how sample households were characterized among worker households in Tsukishima.
+The official report noted: “To find suitable persons to fill in the form, we must ask for the assistance of school principals, police officers, physicians, and laborers, and we have requested a meeting of these persons at the survey office to explain the purpose of the survey, and to solicit applicants for the distribution of the household account book” (Department of Health, Ministry of the Interior 1923a, pp.
 8–9).
-As explained, the target households were not randomly selected because the statistical sampling method was not prevalent.
-However, as described in Section 3, the THBS households can be regarded as standard skilled factory worker households in Tsukishima.
-Therefore, it is plausible that Takano designed this household budget survey in the sampling framework to capture representative households in Tsukishima.
+As noted, the target households were not randomly selected because statistical sampling was not yet prevalent.
+However, as described in Section 3, the THBS households can be regarded as standard skilled factory-worker households in Tsukishima.
+Therefore, Takano may have designed this household budget survey to capture representative households within the sampling framework of Tsukishima.
 Table B.1: Survey Categories in the Field Study of Tsukishima Survey
@@
 Sources: Department of Health, Ministry of the Interior (1923a, 1923b, 1923c).
-Next, I overview the survey items and methodology.
+Next, I provide an overview of the survey items and methodology.
 The investigation was primarily divided into “documentary” and “field” surveys.
-The former was an analysis of the characteristics of Tsukishima, using published statistics such as the Police Department Statistics Book (Department of Health, Ministry of the Interior 1923a, p.
+The former analyzed the characteristics of Tsukishima using published statistics, such as the Police Department Statistics Book (Department of Health, Ministry of the Interior 1923a, p.
 10).
-This was a pioneering attempt at the time and influenced subsequent urban studies (Kawai 1980, pp.
+This was a pioneering approach that influenced subsequent urban studies (Kawai 1980, pp.
 61–62).
-The latter was a field survey including $14$ survey items (Table B.1).
-It is not possible to describe all the survey items in detail, but following the report, they can be divided into four large categories: 1.
-survey on workers’ livelihood (chapter 2); 2.
-survey on sanitary conditions (chapter 3); 3.
-survey on labor conditions (chapter 4); and 4.
+The field survey included $14$ items (Table B.1).
+According to the report, these can be grouped into four main categories: 1.
+workers’ livelihood (chapter 2); 2.
+sanitary conditions (chapter 3); 3.
+labor conditions (chapter 4); and 4.
 social mapping and photography (appendix).
-The survey was conducted over two years, from November 1918 to November 1920, although the timing varied depending on the survey items (Table B.1).
-A unique feature of the survey methodology is the use of young researchers as commissioned researchers.
-The household budget survey and leisure life survey were conducted by Yasunosuke Gonda, a faculty member of the Doitsugaku kyōkai gakkō (Association for German Sciences), who had researched cinematograph and entertainment.
-Tetsuo Hoshino, a medical doctor, was responsible for the sanitary survey.
-The labor survey was managed by Yoshitsuru Yamana, a graduate of Tokyo Imperial University who contributed to the organization of labor unions.
-In addition, Toyotaro Miyoshi, a sociology student at the Tokyo Imperial University, and others became temporary employees (see footnote 54).
-At the time, the Health Bureau of the Ministry of the Interior conducted several sanitary surveys of rural areas.
-Takano may have designed the urban sanitary survey from this background (Kawai 1981a, pp.
+The survey was conducted over two years, from November 1918 to November 1920, with timing varying depending on the survey items (Table B.1).
+A distinctive feature of the methodology was the use of young commissioned researchers.
+The household budget and leisure life surveys were conducted by Yasunosuke Gonda, a faculty member of the Doitsugaku Kyōkai Gakkō (Association for German Sciences) who had conducted research on cinematography and entertainment.
+Tetsuo Hoshino, a medical doctor, oversaw the sanitary survey.
+Yoshitsuru Yamana, a Tokyo Imperial University graduate who contributed to labor union organization, managed the labor survey.
+Additionally, Toyotaro Miyoshi, a sociology student at Tokyo Imperial University, and others served as temporary employees (footnote 40).
+At the time, the Health Bureau of the Ministry of the Interior conducted several rural sanitary surveys, and Takano may have drawn on this experience in designing the urban sanitary survey (Kawai 1981a, pp.
 8–9).
-However, based on the idea that “health and sanitation surveys can at least encompass economic and social surveys,” Takano included a wide range of survey items, not limited to sanitation surveys (Department of Health, Ministry of the Interior 1923a, p.
+However, operating on the principle that “health and sanitation surveys can at least encompass economic and social surveys,” Takano included a broad range of survey items beyond sanitation (Department of Health, Ministry of the Interior 1923a, p.
 56).
 Kawai (1980, p.
-55) credits the Tsukishima Survey as a noteworthy pioneering concept: “going beyond a mere insurance and sanitation survey, it attempted to accurately depict the state of urban workers’ lives as they became established in the city during modernization, with skilled workers as its core.”
+55) credits the Tsukishima Survey as a notable pioneering effort: “Going beyond a mere insurance and sanitation survey, it attempted to accurately depict the state of urban workers’ lives as they became established in the city during modernization, with skilled workers as its core.”
 (a) Outside of the office
@@
 Note: Figures B.1a and B.1b show the outside and inside of the survey office in Tsukishima, respectively. Source: Department of Health, Ministry of the Interior 1923a, figures 1 and 2.
-Takano stated, “To be as complete as possible, it is best to continue the survey over as long period as possible and to obtain accurate facts by recording the workers’ household income and expenditures.” For this reason, he believed that “the best method of investigation would be to establish a survey station directly in the survey area and to have a full-time investigator reside there as much as possible, familiarize himself with the area and the people, and conduct the field survey” (Department of Health, Ministry of the Interior 1923a, p.
+Takano stated, “To be as complete as possible, it is best to continue the survey over as long a period as possible and to obtain accurate facts by recording the workers’ household income and expenditures.” For this reason, he believed that “the best method of investigation would be to establish a survey station directly in the survey area and to have a full-time investigator reside there as much as possible, familiarize himself with the area and the people, and conduct the field survey” (Department of Health, Ministry of the Interior 1923a, p.
 4).
-In fact, through the mediation of the Tsukishima police station, a survey office was rented at 9-3 Higashinakadōri, Tsukishima, where Yoshitsuru Yamana moved in from November 1918 and contacted worker households (Department of Health, Ministry of the Interior 1923a, p.
+With the mediation of the Tsukishima police station, a survey office was rented at 9-3 Higashinakadōri, Tsukishima.
+Yoshitsuru Yamana moved in from November 1918 and began contacting worker households (Department of Health, Ministry of the Interior 1923a, p.
 7; Takeda 2015, p.
 77).
-The survey office was open from November 1918 to December 1920, including the entire survey period.
-Figure B.1 shows the photographs of the survey office.
-It is also important to note that Takano and his team held weekly meetings at the office during the survey period.
-Toyotaro Miyoshi, who was an assistant in the survey, recalls the situation at that time as follows:
+The office remained open for the entire survey period, from November 1918 to December 1920.
+Figure B.1 shows photographs of the survey office.
+Importantly, Takano and his team held weekly meetings at the office throughout the survey period. Toyotaro Miyoshi, who worked as an assistant on the survey, recalls:
 After the start of the survey office, all the members gathered once a week with Iwasaburo Takano present to discuss the implementation and progress of the survey and administrative matters.
@@
 Miyoshi (1980, p. 38)
-He also states the following about Takano:
+He also notes:
 When the Tsukishima Survey was launched, the Faculty of Economics (in the Tokyo Imperial University) was becoming independent, and he was swamped both inside and outside the university.
@@
 Kawai (1981b, p.
-28) assessed that there were no interconnections among the survey items mentioned above and that the survey “ended up being several fragmentary surveys rather than a comprehensive and systematic survey of workers’ lives.” This difficulty may explain the reason why previous academic studies have rarely used the Tsukishima Survey.
-However, as already explained, Takano did not intend to conduct a large-scale social survey in Tsukishima in the first place.
-Instead, as Takano noted, the best method is to investigate a representative area “narrowly but deeply.” The goal was to improve the quality of the investigation rather than the scale of it.
-In this light, the fact that the researchers lived in the survey area and were constantly organizing and revising the survey results should be appreciated.
+28) observes no clear interconnections among the survey items, and that the survey “ended up being several fragmentary surveys rather than a comprehensive and systematic survey of workers’ lives.” This may explain why previous academic studies have rarely used the Tsukishima Survey.
+However, as explained earlier, Takano did not aim to conduct a large-scale social survey in Tsukishima.
+Instead, he argued that the best method is to investigate a representative area “narrowly but deeply,” prioritizing the quality of the investigation over its scale.
+In this light, the fact that researchers lived in the survey area and continuously organized and revised the results is significant.
 Although the Tsukishima Survey is an early social survey in Japan, it is undoubtedly of high quality.
@@
 Notes: This photograph shows an example of the official report of the Tsukishima Survey published by the Department of Health, Ministry of the Interior, in 1923.
-Source: Department of Health, Ministry of the Interior 1923b, title page. The tone was adjusted by the author using Adobe Photoshop 24.7.0.
+Source: Department of Health, Ministry of the Interior 1923b, title page.
+The tone was adjusted by the author using Adobe Photoshop 24.7.0.
 Finally, I summarize the format of the published report and the unpublished materials.
-The results of the Tsukishima Survey were officially summarized in one report published in December 1921 by the Sanitary Bureau of the Ministry of the Interior and in two appendices (Figure B.2 is an example of the front cover).
-The report was published under the name “Report on Field Survey in Tsukishima, Kyobashi, Tokyo,” and consists of four sections.
-The first section is a comprehensive review by Takano and outlines the survey.
-The second section summarizes the living conditions of the workers, and was written by Gonda (partly by Yamana).
-Hoshino wrote the third section.
-It summarizes the sanitary conditions of workers.
+The results of the Tsukishima Survey are officially documented in a report published in December 1921 by the Sanitary Bureau of the Ministry of the Interior, along with two appendices (Figure B.2 shows an example of the front cover).
+The report, published under the title “Report on Field Survey in Tsukishima, Kyobashi, Tokyo,” consists of four sections.
+The first section, a comprehensive review by Takano, outlines the survey.
+The second section, written by Gonda (partly by Yamana), summarizes the living conditions of workers.
+Hoshino authored the third section, which addresses workers’ sanitary conditions.
 In the fourth section, Yamana summarizes the labor situation in Tsukishima.
-There are two appendices.
-One contains statistical tables related to the above report, and the other organizes social maps and photographs.
-The correspondence between the survey items and the report is summarized in Table B.1.
-The Ministry of Home Affairs has already published these materials, which are available at the Gonda Yasunosuke Library of the OISR and the University of Tokyo Library.
-[^92] In the OISR, these are included in the Archives of Gonda Yasunosuke (7-2; 7-3; 7-4).
-In addition, there is a second manuscript of the Tsukishima Survey report.
-This is the manuscript of the second series, which was preserved in the home of Gonda Hayao, the second son of Gonda Yasunosuke, and was discovered by Terade Koji.
-The second series includes a summary of the results of a household survey of elementary school teachers’ households in Tokyo, conducted as a supplementary survey to the Tsukishima Survey (see Terade (1982) for details).
-The manuscript of this second series was later donated to the OISR and is now available as part of the Gonda Yasunosuke Collection at the Institute (The OISR, Archives of Gonda Yasunosuke (8-12; 8-13; 8-14; 8-15)).
+The two appendices complement the main report.
+One contains statistical tables corresponding to the report, while the other organizes social maps and photographs.
+The correspondence between survey items and the report is summarized in Table B.1.
+These materials have been published by the Ministry of Home Affairs and are available at the Gonda Yasunosuke Library of the OISR and the University of Tokyo Library.
+At the OISR, they are included in the Archives of Gonda Yasunosuke (7-2; 7-3; 7-4).
 B.2 Original Micro-spreadsheets from the Household Budget Survey in Tsukishima Survey
-The official report does not include household-level information, and only a small quantity of the original forms used to prepare the report have survived.
-Fortunately, however, some of the original micro-spreadsheets are kept in the archives of the OISR because Takano was its first director.
+The official report does not provide household-level information.
+Fortunately, some of the original micro-spreadsheets are preserved in the OISR archives, as Takano was its first director.
 This study uses the original micro-spreadsheets from the household budget survey (No.
-9 in Table B.1), which was used by Gonda to write the second section of the official report.
-While they have yet to be publicly available, I received permission from the OISR to use the original forms stored in the archives: the OISR, Archives of the Tsukishima Survey (THBS, unreleased).
-[^93] I would like to express my sincere thanks to Professor Kazue Enoki, Deputy Director of the OISR, for allowing me access to the documents.
+9 in Table B.1), which Gonda used to write the second section of the official report.
+Although these spreadsheets are not yet publicly available, I obtained permission from the OISR to use the original forms stored in the archives: the OISR, Archives of the Tsukishima Survey (THBS, unreleased).
+[^76] I would like to express my sincere thanks to Professor Kazue Enoki at the OISR, for allowing me access to the documents.
 I wish to thank Ms.
@@
 The left page includes the details of revenue and receiving and making payments on the goods (actual things, including gifts).
-Source: The THBS, account #6, 15th February, 1919. The tone was adjusted by the author using Adobe Photoshop 24.7.0.
-The budget book used for the household budget survey is named the Kinsendeiri hikaechō (money receipts and disbursements account book).
-The design of this budget book was based on the form used in Takano’s first household survey (Online Appendix B.1).
+Source: The THBS, account #6, 15th February, 1919.
+The tone was adjusted by the author using Adobe Photoshop 24.7.0.
+The budget book is titled the Kinsendeiri hikaechō (money receipts and disbursements account book).
+Its design is based on the form used in Takano’s first household survey (Online Appendix B.1).
 [^77] Department of Health, Ministry of the Interior (1923a, p.
 152).
-The budget book was to be filled out daily and included four sections: income, expenses, receipts from credit purchases, and receipts and disbursements of actual goods.
-Most of the entries were in the income and expenses sections, indicating that their daily living was conducted through wage receipt and consumption in cash.
-Figure B.3 shows an example of the budget book.
-As described, the method used to select households for the survey was not recorded in detail (Online Appendix B.1).
-According to the report, they initially recruited just over $90$ households through the factories, police station, and elementary schools in Tsukishima.
-Of these, $50$ households completed the book for at least one month, but $10$ were deemed incomplete and not included in the report.
-In addition to the $40$ budget books used to prepare the report, therefore, the OISR has another set of budget books marked as “discarded.” According to my survey, about six incomplete books of the worker households have been determined to be in this “discarded” category.
-Therefore, based on the condition of the remaining books, it is plausible to believe that approximately $50$ households participated in the survey.
-Of the $40$ households included in the report tally, two households maintained records for more than one year, $11$ households for 6–12 months, and the remaining $27$ households registered for less than six months (Department of Health, Ministry of the Interior 1923a, pp.
+The book is completed daily and comprises four sections: income, expenses, receipts from credit purchases, and receipts and disbursements of actual goods.
+Most entries appear in the income and expenses sections, indicating that daily life was largely sustained through wages and cash expenses.
+Figure B.3 provides an example of the budget book.
+As noted, the method for selecting households for the survey was not recorded in detail (Online Appendix B.1).
+According to the report, just over $90$ households were initially recruited through factories, the police station, and elementary schools in Tsukishima.
+Of these, $50$ households completed the book for at least one month, but $10$ household’s entries were considered incomplete and excluded from the report.
+In addition to the $40$ budget books used for the report, the OISR holds another set marked as “discarded.” My survey indicates that about six incomplete books from worker households fall into this category.
+Based on the condition of the remaining books, approximately $50$ households potentially participated in the survey.
+Among the $40$ households included in the report, two maintained records for more than one year, $11$ households for 6–12 months, and the remaining $27$ households recorded data for less than six months (Department of Health, Ministry of the Interior 1923a, pp.
 152–153).
-Initially, the survey period was intended to be six months, but the average bookkeeping period for the $40$ households was four months.
+Although the survey was initially planned for six months, the average bookkeeping period for these $40$ households was four months.
 B.3 Trimming
@@
 (b) Semi-monthly panel dataset
-(c) Adj. monthly panel dataset
+(c) Monthly panel dataset
 Figure B.4: Structures of the THBS Datasets by Different Time Frequencies
-Note: Figure B.4aB.4b, and B.4c show the observations in the daily, semi-monthly, and adjusted monthly panel datasets, respectively.
+Notes: Figure B.4aB.4b, and B.4c show the observations in the daily, adjusted semi-monthly, and adjusted monthly panel datasets, respectively.
 For example, a marker in Figure B.4a indicates an observation in a specific year-month-date cell.
@@
 Source: Created by the author using the THBS sample.
-To check whether the statistics from THBS budget books contained any missing information, I digitized all $40$ household budget books.
-Throughout this digitization process, I found that there are a few misreporting issues in the budget books.
+I digitized all $40$ household budget books. Throughout this digitization process, I found that there are a few misreporting issues in the budget books.
 First, there is a household that does not include income information (household #37).
@@
 Similarly, a household does not include information on the payments to the credit purchases (household #39).
-Although this household’s budget book documented the amount purchased in credit purchases, I excluded this household’s budget book because the expenditure series became systematically smaller due to the omission of the payment information.
+I excluded this household’s budget book because the expenditure series became systematically smaller due to the omission of the payment information.
 Second, the final pages of the budget books are sometimes omitted.
@@
 Dev.
-$=5.0$).
+=5.0=5.0).
 The characteristics of the $33$ THBS households are summarized in detail in Section 3.
@@
 The null of the zero-slope hypothesis is not rejected in all the specifications at the conventional level (not reported).
-In this subsection, I test whether attrition leads to selection bias in the THBS sample.
 If the difference in the preference for consumption predicted large attritions, there must be statistically significant correlations between the family size variables and attrition.
 In other words, the units with shorter (longer) time-series observations have different preferences from those with longer (shorter) time-series observations.
-Table B.2 summarizes the results of the balancing tests (Section 3.2).
+Table B.2 summarizes the results of the balancing tests.
 The dependent variable is an indicator variable that takes one if the household has time-series observations less than the median (column 1), 75 percentile (column 2), and 25 percentile (column 3).
@@
 As described in Section 3.2, while the THBS sample does not cover households with one person and with a very large number of people, it covers most of the classes in the household sizes with reasonable approximation to the population distribution.
-[^95] The number of households with one people is $362$, accounting for only $6$% of the total number of households.
-Similarly, the number of households with nine or more people is $294$, accounting for only $5$% (Section 3.2).
+[^78] The number of households with one people is $362$, accounting for only 66% of the total number of households.
+Similarly, the number of households with nine or more people is $294$, accounting for only 55% (Section 3.2).
 Figure B.5: Distribution of the Household Size:
@@
 The share in each household size measured in the census is used as the theoretical probability in calculating $\chi^{2}$ test statistic (Pearson 1900).
-Sources: Created by the author using the Tokyo City Office (1922a, pp. 262–283) and the THBS dataset.
+Sources: Created by the author using the Tokyo City Office (1922a, pp.
+262–283) and the THBS dataset.
 B.6 Estimating Average Head’s Monthly Income
@@
 As the early wage surveys focused only on the average wage for entire workers, official statistics in the 1910s never included average wage statistics by age (bin).
-In the “Statistical Data Field Survey Law” (Law No.
-52) in April 1922, average wages by age group were surveyed for the first time in the “Labor Statistics Field Survey” in October 1924 conducted by the Social Affairs Bureau of the Ministry of Home Affairs (Committee for the History of the Labor Movement 1964, p.
+In the Statistical Data Field Survey Law (Law No.
+52) in April 1922, average wages by age group were surveyed for the first time in the Labor Statistics Field Survey in October 1924 conducted by the Social Affairs Bureau of the Ministry of Home Affairs (Committee for the History of the Labor Movement 1964, p.
 14).
-The result of this survey in Tokyo was published as the “Labor Statistics Table for Tokyo City and Suburbs” (hereafter, LST) by the Tokyo City Statistics Division in 1926.
+The result of this survey in Tokyo was published as the Labor Statistics Table for Tokyo City and Suburbs (hereafter, LST) by the Tokyo City Statistics Division in 1926.
 (a) 1918
@@
 For the 1919 price data, the fitted model is $\hat{y}^{\textit{Daily Wage}}=-13.18+5.47a^{1/2}-0.43a^{2}$ with $R^{2}=0.99$.
-Sources: Created by the author using the Tokyo City Statistics Division (1926a, pp. 16; 244–245). Data on the consumer price index for the cities are obtained from Ohkawa et al. (1967, p. 135).
+Sources: Created by the author using the Tokyo City Statistics Division (1926a, pp.
+16; 244–245).
+Data on the consumer price index for the cities are obtained from Ohkawa et al.
+(1967, p.
+135).
 Based on the LST, Figure B.7 shows the average daily wages of male factory workers in Tokyo City’s machinery and equipment manufacturing industry by age group.
 In the LST, there are $13$ age bins, and the number of workers and average wages are listed for each bin.
+The average daily wages have an inverted-U-shape, which is a similar life-cycle pattern to the factory workers observed in late 19th century England (Horrell and Oxley 2000, p.
+42).
 Figure B.7a shows the distribution of wages adjusted to 1918 prices using the consumer price index (CPI) for urban areas (Ohkawa et al.
@@
 $\phi_{m}(a;\bm{\varphi},\mathbf{p})=\varphi_{0}+\sum_{j=1}^{m}\varphi_{j}a^{(p_{j})},$ (6)
-where $a$ ($>0$) denotes class value of age bin, $\bm{\varphi}$ is a vector of coefficients and $\mathbf{p}$ is a vector of powers, satisfying $p_{1}<p_{2}<...<p_{m}$.
+where $a$ (>0>0) denotes class value of age bin, $\bm{\varphi}$ is a vector of coefficients and $\mathbf{p}$ is a vector of powers, satisfying $p_{1}<p_{2}<...<p_{m}$.
 $a^{(p_{j})}$ indicates the Box-Tidwell transformation, taking $\text{ln}(a)$ if $p_{j}=0$ and $a^{p_{j}}$ if $p_{j}\neq 0$.
@@
 To model the function, I consider a set of powers $\mathscr{P}=\{-2,-1,-0.5,0,0.5,1,2,3\}$ following Royston and Altman (1994, p. 434).
-The maximum likelihood estimation is used to search for the most suitable candidate of the coefficients for a given $m$.
-Since $m=2$, there are ${}_{8}\mathrm{H}_{2}$ ($=36$) repeated combinations for the candidates of the parameters.
-By adding the $8$ simple candidates under the nested models (i.e., the models with degree-1 fractional polynomial) to be conservative, I run $44$ regressions in total to find out the model with the highest maximized log-likelihood.
-The best power vector suggested is $(p_{1},p_{2})=(0.5,1)$.
-[^97] The deviance, defined as the twice negative (maximized) log-likelihood, is calculated to be $-31.5$.
-Royston and Altman (1994) proposed the deviance difference, $D(m,\mathbf{p})-D(m,\mathbf{\tilde{p}})$, where $\mathbf{\tilde{p}}$ is the best power vector to compare the different models.
-This statistic has an asymptotic chi-squared distribution with $m$ degree of freedom.
-The estimated deviance difference with the simplest linear model (i.e., $m=1$ and $p_{1}=1$) is $54.6$ with $p$-value = $0.000$.
-Similarly, the difference with an another non-linear model with $m=1$ and $p_{1}=-2$ is $34.2$ with $p$-value = $0.000$.
-The projection model fitted using the best power vector is as follows:
+[…]
 $\hat{y}^{\textit{Daily Wage}}=-10.01+4.15a^{\frac{1}{2}}-0.33a^{1},\quad R^{2}=0.99.$ (8)
@@
 I calculate the weighted average of the daily wage using the number of factory workers in each age bin (12–14 and 15+) and in each category as weights.
-Thus, I shift the functional polynomial function by this amount of difference to yield the proxy distribution for 1918.999999This operation, therefore, assumes that the wage-by-age distributions in 1918 and 1924 are reasonably similar.
+Thus, I shift the functional polynomial function by this amount of difference to yield the proxy distribution for 1918.828282This operation, therefore, assumes that the wage-by-age distributions in 1918 and 1924 are reasonably similar.
 In the LST (i.e., the statistic on which the predicted distribution is based), the weighted average for those aged $14$ and older was $3.25$ yen.
@@
 In other words, the shifted curve shall get closer to the actual (unobservable) wage distribution as age increases.
-The long dashed curve in Figure B.8a examines this relationship by fitting another fractional polynomial model using the 12–14-year old bin figure of the ASCT and the predicted wages for 25–65 from the shifted fractional polynomial.100100100The fitted model is as follows: $\hat{y}^{\textit{Daily Wage}}=-1.19+0.14a^{1}-0.002a^{2},\quad R^{2}=0.98.$ The deviance is reported to be $-149$.
+The long dashed curve in Figure B.8a examines this relationship by fitting another fractional polynomial model using the 12–14-year old bin figure of the ASCT and the predicted wages for 25–65 from the shifted fractional polynomial.
+[^83] The fitted model is as follows: $\hat{y}^{\textit{Daily Wage}}=-1.19+0.14a^{1}-0.002a^{2},\quad R^{2}=0.98.$ The deviance is reported to be 149-149.
 I used the 25–65 age range for the prediction based on the age bins used in the LST.
@@
 The blue dashed line is the average daily wage for 30–40-year-old in the refitted FP.
-Sources: Created by the author using the Tokyo City Office (1921, pp. 726–757); Tokyo City Statistics Division (1926a, pp. 16; 244–245).
+Sources: Created by the author using the Tokyo City Office (1921, pp.
+726–757); Tokyo City Statistics Division (1926a, pp.
+16; 244–245).
 B.6.2 Ancillary Wage
@@
 I prefer to use linear regression herein because it can provide more precise fitted values for small-and medium-scale factories.
-Source: Kitazawa (1924, pp. 20; 58–62).
+Source: Kitazawa (1924, pp.
+20; 58–62).
 The daily wage estimated in Section B.6.1 is a fixed salary and thus, does not include additional salaries such as bonuses and allowances (hereafter, wages other than regular salaries are collectively referred to as ancillary wages).
@@
 Moreover, since the Tsukishima Survey targets skilled worker households, information on the ancillary wages paid to skilled workers is necessary.
-Fortunately, I found a survey report named “Wage Survey Report” edited by Kitazawa (1924).
+Fortunately, I found a survey report named Wage Survey Report edited by Kitazawa (1924).
 This report includes the result of a survey on the wages for machine finishers (fitters) and turners at six machine factories in Tokyo from May–October 1921.
@@
 The surveyed factories included $86$, $157$, $376$, $623$, $928$, and $2,837$ workers, ranging from small to large plants.
-The percentage of skilled workers (finishers and turners) in the surveyed factories ranged from $27$% to $77$%, and the portion of monthly ancillary wages was 12–53%.
+The percentage of skilled workers (finishers and turners) in the surveyed factories ranged from 2727% to 7777%, and the portion of monthly ancillary wages was 12–53%.
 Figure B.9 is a linear regression fit using log-transformed data on the percentage of ancillary wages and the number of workers.
@@
 Panel B of Table B.3 summarizes the distributions of the number of factories, the total number of workers, and the estimated average share of ancillary wage in Tokyo City (Column 1) and Tsukishima (Column 2).
-The average ancillary wage at the Tsukishima factories is $12$%, which is slightly larger than the value for the entire city of Tokyo ($9$%).
+The average ancillary wage at the Tsukishima factories is 1212%, which is slightly larger than the value for the entire city of Tokyo (99%).
 The percentage shares of the number of factory workers by each bin are listed in Columns (1) and (2) of Panel C in Table B.3.
@@
 Total 47,679 100 5,816 100 28 100
-Notes:
-Panel A: This panel shows the size of factories and the share of fixed and ancillary monthly wages in six factories surveyed in Kitazawa (1924).
+Notes: Panel A: This panel shows the size of factories and the share of fixed and ancillary monthly wages in six factories surveyed in Kitazawa (1924).
 The number of factory workers surveyed in factories A–F are 14, 23, 9, 11, 11, and 16, respectively.
@@
 $n_{10}$ 4 4
-Panel B: Estimated monthly earning: $w^{\textit{Total}}$
+Panel B: Estimated Monthly Earning: $w^{Total}$
 $w^{\textit{Total}}$ 56.0 54.5
-Notes:
-1.
+Notes: 1.
 Average daily wage ($w^{Daily}_{j}$): The average daily wage of male factory workers in the skilled age range (aged 30–40) in the machinery sector is calculated using two official reports of the manufacturing censuses documented in the ASCT and LST.
@@
 where $w^{\textit{Regular}}>0$ is the monthly regular wage and $w^{\textit{Ancillary}}\geq 0$ indicates the monthly ancillary wage.
-Online Appendix B.6.2 shows that $w^{\textit{Ancillary}}$ can be calculated as the product of the monthly regular wage and weight for the ancillary wage, $\vartheta\in(0,1)$.101101101Strictly, this weight may not be required to be in this range.
+Online Appendix B.6.2 shows that $w^{\textit{Ancillary}}$ can be calculated as the product of the monthly regular wage and weight for the ancillary wage, $\vartheta\in(0,1)$.
+[^84] Strictly, this weight may not be required to be in this range.
 I use this range because all the cases reported in Kitazawa (1924) take the weights between zero and one.
@@
 The shifted series defines the days from the 2nd to the 16th of each month as the first half, whereas the days from the 17th to the 1st of the following month are the second half.
-Adjusted series is the baseline definition in the main text, which uses the days from 12th to 26th in each month as the “first-half” and those from 27th to 11th in the next month as the “second-half.”
-As explained in Section 3, the raw series contains considerable measurement errors due to the timing of payday.
+Adjusted series is the baseline definition in the main text, which uses the days from 12th to 26th in each month as the “first-half” and those from 27th to 11th in the next month as the “second-half.” As explained in Section 3, the raw series contains considerable measurement errors due to the timing of payday.
 The shifted series can manage the measurement errors from the timing of payday but cannot adjust the difference in the payday timing and expenses.
@@
 B.8 Testing Attenuation Bias in the Unadjusted Panel Dataset
-In this subsection, I assess the validity of the aggregation method used in this study.
+I assess the validity of the aggregation method used in this study.
 Adjusting the aggregate data to account for the timing of paydays reveals more transparent relationships between income and expenditure than the unadjusted data (Section 3.3).
@@
 1997; Gervais and Klein 2010; Nelson 1994).
-Moreover, several subcategories select values close to one or large negative values, which are no longer economically interpretable.102102102For example, the estimate for housing was negative.
+Moreover, several subcategories select values close to one or large negative values, which are no longer economically interpretable.
+[^85] For example, the estimate for housing was negative.
 The households paid a fixed rent amount monthly.
@@
 B.9 Considering Child Share as a Preference Shock Variable
-As explained in the main text, the child share variable exhibits significantly more within-variations than that of the family size.
+As explained in the main text, the child share variable exhibits more within-variations than that of the family size.
 Table B.6 presents the results from specifications that include the share of children under the primary school age range, specifically those aged 0–5, as an alternative preference shock variable.
 Panels A and B show the results for the consumption smoothing and risk-coping strategy regressions, respectively.
-The results are materially similar to the main results presented in the main text (Tables 4 and 5).
+The results are materially similar to the main results (Tables 4 and 5).
 Note that the results remain robust to different definitions of age bins (not reported).
@@
 Conditioning on the Share of Children Aged 0–5
-(1) Semi-monthly Panels (2) Adjusted Monthly Panels
+(1) Semi-monthly Panels (2) Monthly Panels
 Disposable income Obs. Disposable income Obs.
@@
      Miscellaneous -0.079 [0.148] 214 0.506 [0.253]* 110
-(1) Semi-monthly Panels (2) Adj. Monthly Panels
+(1) Semi-monthly Panels (2) Monthly Panels
 Head’s earning Obs. Head’s earning Obs.
@@
 Panel B-2: Labor supply adjustments & sales of miscellaneous assets (yen)
-Other member’s earnings 0.001 [0.010] 289 0.036 [0.033] 124
+Other members’ earnings 0.001 [0.010] 289 0.036 [0.033] 124
 Sales of miscellaneous assets -0.000 [0.000] 289 0.000 [0.001] 124
 ***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors in brackets are clustered at the household level.
-Notes: Panel A: This table shows the results of Equation 1.
-The regressions of the 11 measures of log-transformed consumption on log-transformed disposable income, the proportion of children aged 0–5, household-fixed effect, and time-fixed effect are listed.
-Columns 1 and 2 show the results for the semi-monthly and adjusted monthly panel datasets, respectively.
-The estimated coefficients on log-transformed disposable income are listed in the columns named “Coef.”.
-Panel B: This table shows the results of equation 2.
+Notes: Panel A: This table presents the results of equation 1.
+It reports regressions of 11 measures of log-transformed consumption on log-transformed disposable income, the family size control, the household fixed effect, and the time fixed effect.
+Columns 1 and 2 show the results for the adjusted semi-monthly and adjusted monthly panel datasets, respectively.
+The estimated coefficients on log-transformed disposable income are shown in the columns labeled “Coef.” Panel B: This table shows the results of equation 2.
 The proportion of children aged 0–5, household fixed effects, and time-fixed effects are included in all the regressions.
-Columns 1 and 2 show the estimates for the semi-monthly and adjusted monthly panel datasets, respectively.
+Columns 1 and 2 show the estimates for the adjusted semi-monthly and adjusted monthly panel datasets, respectively.
 The estimated coefficients on the head’s earnings are listed in the “Coef.” columns.
-Panel B-1: Each net income variable is defined as the difference between income and expenses.
-“Net savings” refers to the difference between the amount withdrawn and the total deposits.
-“Net insurance” refers to the amount received from insurance minus the expenses paid for insurance.
-“Net borrowing” refers to the amount of borrowing minus the debt payments.
+Panel B-1: Each net income variable is defined as income minus expenditure.
+“Net savings”: withdrawal amount minus deposits to savings.
+“Net insurance”: insurance payments received minus the expenses paid for insurance.
+“Net borrowing”: the amount borrowed minus debt payments.
+This category includes net income from both pawnshop and other lending institutions.
 “Net credit purchase” refers to the amount of credit purchases minus credit redeemed.
-“Net gifts” refers to the total received amounts, including both pecuniary and non-pecuniary gifts, minus the payments for gifts.
-Panel B-2: “Other member’s earnings” refers to the total income earned by all family members except for the head of the household.
-“Sales of miscellaneous assets” includes the sale of daily miscellaneous goods such as newspapers and empty bottles.
+“Net gifts”: total gifts received, including both pecuniary and non-pecuniary, minus any payments for gifts.
+Panel B-2: “Other family members’ earnings”: total income earned by all family members except the household head.
+“Sales of miscellaneous assets”: include the sale of goods such as old newspapers and empty bottles.
 All dependent variables are in yen.
@@
 Figure B.11: Distribution of the Idiosyncratic Shocks
-Notes: This figure shows the distribution of idiosyncratic shocks in the semi-monthly THBS panel dataset.
+Notes: This figure shows the distribution of idiosyncratic shocks in the adjusted semi-monthly THBS panel dataset.
 The idiosyncratic shocks are estimated as the residuals from the regression of the first difference in income on the first difference in aggregate income.
-The red line indicates the mean of the idiosyncratic shocks ($-0.235$).
+The red line indicates the mean of the idiosyncratic shocks (0.235-0.235).
 Source: Created by the author using the THBS sample.
-Figure B.11 illustrates the distribution of the idiosyncratic shocks predicted using the semi-monthly panel dataset.103103103I use the semi-monthly panel dataset to overview the entire distribution of the idiosyncratic shocks rather than the daily panel datasets.
+Figure B.11 illustrates the distribution of the idiosyncratic shocks predicted using the semi-monthly panel dataset.
+[^86] I use the semi-monthly panel dataset to overview the entire distribution of the idiosyncratic shocks rather than the daily panel datasets.
 This is because the daily panels include a set of consecutive household-date cells with no income.
@@
 B.11 Unit Root and Autocorrelations
-Table B.7 presents the results of the Fisher-type panel unit root tests for the total expenditure in the semi-monthly (Column 1) and adjusted monthly (Column 2) panel datasets.
+Table B.7 presents the results of the Fisher-type panel unit root tests for the total expenditure in the semi-monthly (Column 1) and monthly (Column 2) panel datasets.
 Both columns show that the null hypothesis of unit roots in all the panels is rejected at the conventional level.
@@
 Table B.7: Results of the Unit Root Tests
-Test statistics (1) Semi-Monthly (2) Adjusted Monthly
-$P$-statistic $p$-value 0.0000 0.0001
-$Z$-statistic $p$-value 0.0000 0.0000
+Test statistics (1) Semi-Monthly Panels (2) Monthly Panels
+$P$-statistic pp-value 0.0000 0.0001
+$Z$-statistic pp-value 0.0000 0.0000
 $L^{*}$-statistic $p$-value 0.0000 0.0000
@@
 The number of lagged differences in the ADF regression equation reported is set as one because including two or more two lags reduces a large part of the panel units available for computing the test statistics due to the lack of balance of the panels.
-I confirm that including two lags in the semi-monthly panels does not change the results.
-Including two lags in the adjusted monthly panels and three lags in the semi-monthly panels is impossible.
+I confirm that including two lags in the adjusted semi-monthly panels does not change the results.
+Including two lags in the adjusted monthly panels and three lags in the adjusted semi-monthly panels is impossible.
 See Choi (2001) for the details of the tests.
@@
 This test is valid under a particular utility function.
-Specifically, the bias could be raised in the case that the coefficient of relative risk aversion interacts with the aggregate shock in the first-order condition for consumption because the interaction means that the aggregate shock cannot be removed using the time-fixed effect in the risk-sharing regressions.104104104Specifically, Schulhofer-Wohl (2011, pp.
+Specifically, the bias could be raised in the case that the coefficient of relative risk aversion interacts with the aggregate shock in the first-order condition for consumption because the interaction means that the aggregate shock cannot be removed using the time-fixed effect in the risk-sharing regressions.
+[^87] Specifically, Schulhofer-Wohl (2011, pp.
 928–929) considers the utility maximization problem under the following weighted sum of discounted expected utilities: $\sum_{i}\dot{\alpha}_{i}E\left[\sum_{t}\dot{\beta}^{t}\frac{[c_{i,t}(s_{t})]^{1-\dot{\gamma}_{i}}}{1-\dot{\gamma}_{i}}\right],$ (17) where $\dot{\alpha}_{i}$ is Parate weight and $\dot{\gamma}_{i}$ is the coefficient of relative risk aversion for household $i$.
@@
 Mozzoco and Saini (2012) consider a more general case for testing the homogeneity of risk preferences.
-Their test requires a panel dataset with a very long time dimension that is not applicable to my THBS dataset.105105105The data used in their study covers the household-level monthly panel dataset surveyed in rural India between 1975 and 1985.
+Their test requires a panel dataset with a very long time dimension that is not applicable to my THBS dataset.
+[^88] The data used in their study covers the household-level monthly panel dataset surveyed in rural India between 1975 and 1985.
 The households with fewer than 80 data points are not included in the homogeneity tests (Mozzoco and Saini 2012, p.
@@
 An alternative way to test potential heterogeneity in the risk preference is to use direct measures on the risk preferences of workers.
-Schulhofer-Wohl (2011) uses the results of experimental risk-tolerance questions on the workers’ risk preferences in the Health and Retirement Study to test whether the correlations between earnings and aggregate shocks differ across high- and low-risk-tolerance workers.106106106This question is asking if the respondents would be willing to take a risk to get a new job with a higher income, even if there was a 50-50 chance of a reduction in income by a third (Schulhofer-Wohl 2011, P.
+Schulhofer-Wohl (2011) uses the results of experimental risk-tolerance questions on the workers’ risk preferences in the Health and Retirement Study to test whether the correlations between earnings and aggregate shocks differ across high- and low-risk-tolerance workers.
+[^89] This question is asking if the respondents would be willing to take a risk to get a new job with a higher income, even if there was a 50-50 chance of a reduction in income by a third (Schulhofer-Wohl 2011, P.
 931).
@@
 Since the THBS has no experimental questions, it is impossible to apply the same test.
-Given this limitation—that the direct test using experimental risk-tolerance question is unavailable—I compare the correlations between the heads’ income and aggregate consumption over sub-groups to obtain insights into the potential difference in the heads’ attitudes toward risks.
+Given this limitation, I compare the correlations between the heads’ income and aggregate consumption over sub-groups to obtain insights into the potential difference in the heads’ attitudes toward risks.
 For household $i$ and semi-month $t$, the specification is characterized as follows:
@@
 Secondly, I test the hypothesis that the heads who work in the smithing and the non-machinery factories have different risk preferences.
-As described in Online Appendix A.5, skilled workers in the smithing sector could be small business owners.107107107Workers in the canning factories might have had a different preference on the risk preferences (Online Appendix A.6).
+As described in Online Appendix A.5, skilled workers in the smithing sector could be small business owners.
+[^90] Workers in the canning factories might have had a different preference on the risk preferences (Online Appendix A.6).
 However, my THBS sample does not include the workers in canning factories.
@@
 This group has only three heads: one worked in a smithing factory, and two worked in food and paper-making factories.
-Table B.8 presents the results.
-In Column (1), the estimated coefficient on the interaction term is close to zero and statistically insignificant.
-Thus, the sensitivities to aggregate consumption are similar across different scales of factories.
-Column (2) further shows that the estimated coefficient on the interaction term is not statistically significant under an alternative proxy of risk tolerance.
-It suggests that the heads who work in the smithing and non-machinery factories show similar responses to the aggregate shock.
-Since the aggregate consumption is the cross-sectional sample average, the aggregate variable may contain noise in the period with the small number of cross-sectional observations.
-To manage this potential issue, I exclude October and November 1919 in Columns (3) and (4).
-As shown, the estimated coefficients on the interaction terms are largely unchanged in both columns.
-This finding is consistent with the fact that these heads have similar characteristics because the THBS focused on skilled factory workers in a representative manufacturing area, Tsukishima.
-Therefore, it is plausible that the THBS households had similar attitudes toward the risks.
 Table B.8: Results for Testing Heterogeneous Risk Preferences
@@
 ***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors in brackets are clustered at the household level.
-Notes: This table shows the results of equation 19 for the semi-monthly panel dataset.
+Notes: This table shows the results of equation 19 for the adjusted semi-monthly panel dataset.
 ‘Aggregate Consumption’ indicates the cross-sectional sample average of the total expenditures.
@@
 Columns 3–4 show the results for the period between January and September 1919 to remove year-month cells with the smaller (cross-sectional) number of households (Online Appendix Figure B.4).
+Table B.8 presents the results.
+In Column (1), the estimated coefficient on the interaction term is close to zero and statistically insignificant.
+Thus, the sensitivities to aggregate consumption are similar across different scales of factories.
+Column (2) further shows that the estimated coefficient on the interaction term is not statistically significant under an alternative proxy of risk tolerance.
+It suggests that the heads who work in the smithing and non-machinery factories show similar responses to the aggregate shock.
+Since the aggregate consumption is the cross-sectional sample average, the aggregate variable may contain noise in the period with the small number of cross-sectional observations.
+To manage this potential issue, I exclude October and November 1919 in Columns (3) and (4).
+As shown, the estimated coefficients on the interaction terms are largely unchanged in both columns.
+This finding aligns with the similarity of these household heads, as the THBS targeted skilled factory workers within a geographically narrow area.
+Therefore, it is plausible that the THBS households had similar attitudes toward the risks.
 Appendix Appendix C Empirical Analysis Appendix
 C.1 Conceptual Framework
-Consider the economy where household $i=1,...,N$ receives an uncertain income at time $t$ depending on the state of the world $s_{t}\in S_{t}$ as $y_{i,t}(s_{t})$.
+Consider the economy where household i=1,,Ni=1,...,N receives an uncertain income at time $t$ depending on the state of the world $s_{t}\in S_{t}$ as $y_{i,t}(s_{t})$.
 For infinite time horizon, the state $s_{t}$ realizes with probability $\pi(s_{t})\in[0,1]$.
@@
 $\sum_{i=1}^{N}\omega_{i}\sum_{t=0}^{\infty}\rho^{t}\sum_{s_{t}\in S_{t}}\pi(s_{t})u(c_{i,t}(s_{t}),\theta_{i,t}(s_{t})),$ (20)
-where $\omega_{i}$ is the social planner’s weight representing the reciprocal of the marginal utility of each household ($10<\omega_{i}<$ 1 and $\sum_{i=1}^{N}\omega_{i}=1$), $10<\rho^{t}<$ 1 is the discount factor, $c_{i,t}(s_{t})$ is consumption, and $\theta_{i,t}(s_{t})$ is a preference shock.
+where $\omega_{i}$ is the social planner’s weight representing the reciprocal of the marginal utility of each household ($0<\omega_{i}<1$ and $\sum_{i=1}^{N}\omega_{i}=1$), $0<\rho^{t}<1$ is the discount factor, $c_{i,t}(s_{t})$ is consumption, and $\theta_{i,t}(s_{t})$ is a preference shock.
 For all states $s_{t}$, the aggregate resource constraint is given by:
@@
 The estimate of $\gamma$ shall range from zero (i.e., idiosyncratic shocks are perfectly insured) to one (for the absence of insurance).
+[^91] I use the conventional within estimator.
+The bias-corrected estimator for the two-way fixed-effects model in a difference-in-differences (TWFE-DID) design with continuous treatment over multiple periods remains underdeveloped (de Chaisemartin and D’Haultfœuille 2023).
+Therefore, previous studies have not yet incorporated recent theoretical developments in the TWFE-DID framework.
 To test the risk-coping mechanisms, I consider the following representation of individual consumption following Fufchamps and Lund (2003).
@@
 where $\tilde{y}_{i,t}$ is the head’s earning, $x_{i,t}$ is the family size control, $\nu_{i}$ indicates the household-fixed effect, $\zeta_{t}$ indicates the time-fixed effect, and $\epsilon_{i,t}$ is a random error term.
-I herein use the head’s earnings ($\tilde{y}_{i,t}$) as the idiosyncratic shock variable instead of the disposable income.108108108In the analysis, I consider the labor-supply adjustment variable and sales of miscellaneous assets as the additional dependent variables in addition to the net income variables $r_{i,t}$ (Section 4.2).
+I herein use the head’s earnings ($\tilde{y}_{i,t}$) as the idiosyncratic shock variable instead of the disposable income.
+[^92] I consider the labor-supply adjustment variable and sales of miscellaneous assets as the additional dependent variables in addition to the net income variables $r_{i,t}$ (Section 4.2).
 The former variable is the total earnings of the family members except for the head, meaning that the simultaneous issue arises if I use the disposable income in this equation.
@@
 C.2 Raw Relationship between Income and Expenditures
-Figure C.1 summarizes the densities of the log-differences (LD) of the semi-monthly expenditures.
-Figure C.1a illustrates the LD distribution of the total expenditure, whereas Figures C.1bC.1l summarize the LD distributions of the 11 subcategories of the expenditure.
-These figures confirm no specific outliers in all these expenditure categories.
-Figure C.2 summarizes the scatter plots between the LD of the semi-monthly disposable income and expenditures.
-Figure C.2a illustrates the relationship between the LD of the semi-monthly disposable income and total expenditure.
-Figures C.2bC.2l illustrate the correlations between the LD of the semi-monthly disposable income and 11 expenditure subcategories.
-Generally, rent was paid once per month.
-This means that the semi-monthly dataset fails to capture the variations in rent payment because it includes some semi-monthly cells with zero values.
-In addition, the water bills are included in the rent subcategory but not in the utility subcategory.
-This means that a semi-month cell takes a minimal value when the payment of rent is not included in that cell.
-Figure C.1c shows that the LD of rent generates some larger values in the semi-monthly panel dataset due to those water bill payments.
-These cells taking large LD values substantially increase the slope, as shown in Figure C.2c.
-Given this technical issue, the adjusted monthly panel dataset is suitable for estimating the income elasticity for the rent category.
-In fact, Figure C.4c confirms that the scatter plot from the adjusted monthly variation shows a flatter relationship.
-(a) Total expenditure
+(a) Total
 (b) Food
@@
 Figure C.1: Distributions of the Log-differences of the Semi-monthly Expenditures
-Notes: The distribution of the log-difference in semi-monthly total expenditure is shown in Figure C.1a.
-Figures C.1bC.1l show the distributions of the log-differences in semi-monthly expenditure for the 11 subcategories listed in panel A of Table 3.
+Notes: The distribution of the log-difference in adjusted semi-monthly total expenditure is shown in Figure C.1a.
+Figures C.1bC.1l show the distributions of the log-differences in adjusted semi-monthly expenditure for the 11 subcategories listed in panel A of Table 3.
 Source: Created by the author using the THBS sample.
-(a) Total expenditure
+(a) Total
 (b) Food
@@
 in Semi-monthly Income and Expenditure
-Notes: Figure C.2a illustrates the relationship between the log-difference in semi-monthly disposable income and expenditure.
-Figures C.2bC.2l illustrate the relationships between the log-differences in semi-monthly disposable income and the 11 expenditure subcategories listed in Panel A in Table 3.
+Notes: Figure C.2a illustrates the relationship between the log-difference in adjusted semi-monthly disposable income and expenditure.
+Figures C.2bC.2l illustrate the relationships between the log-differences in adjusted semi-monthly disposable income and the 11 expenditure subcategories listed in Panel A in Table 3.
 The range of the y-axis is fixed across all the figures for comparability.
 Source: Created by the author using the THBS sample.
-(a) Total expenditure
+(a) Total
 (b) Food
@@
 Source: Created by the author using the THBS sample.
-(a) Total expenditure
+(a) Total
 (b) Food
@@
 in Monthly Income and Expenditure
-Notes: Figure C.4a illustrates the relationship between the log-difference in semi-monthly disposable income and expenditure.
-Figures C.4bC.4l illustrate the relationships between the log-differences in semi-monthly disposable income and the 11 expenditure subcategories listed in Panel A in Table 3.
+Notes: Figure C.4a illustrates the relationship between the log-difference in monthly disposable income and expenditure.
+Figures C.4bC.4l illustrate the relationships between the log-differences in monthly disposable income and the 11 expenditure subcategories listed in Panel A in Table 3.
 The range of the y-axis is fixed across all the figures for comparability.
 Sources: Created by the author using the THBS sample.
-C.3 Retail Sales in Tokyo City: A Retail Store Survey of 1935
+Figure C.1 summarizes the densities of the log-differences (LD) of the semi-monthly expenditures.
+Figure C.1a illustrates the LD distribution of the total expenditure, whereas Figures C.1bC.1l summarize the LD distributions of the 11 subcategories of the expenditure.
+These figures confirm no specific outliers in all these expenditure categories.
+Figure C.2 summarizes the scatter plots between the LD of the adjusted semi-monthly disposable income and expenditures.
+Figure C.2a illustrates the relationship between the LD of the adjusted semi-monthly disposable income and total expenditure.
+Figures C.2bC.2l illustrate the correlations between the LD of the adjusted semi-monthly disposable income and 11 expenditure subcategories.
+Generally, rent was paid once per month.
+This means that the semi-monthly dataset fails to capture the variations in rent payment because it includes some semi-monthly cells with zero values.
+In addition, the water bills are included in the rent subcategory but not in the utility subcategory.
+This means that a semi-month cell takes a minimal value when the payment of rent is not included in that cell.
+Figure C.1c shows that the LD of rent generates some larger values in the semi-monthly panel dataset due to those water bill payments.
+These cells taking large LD values substantially increase the slope, as shown in Figure C.2c.
+Given this technical issue, the monthly panel dataset is suitable for estimating the income elasticity for the rent category.
+In fact, Figure C.4c confirms that the scatter plot from the monthly variation shows a flatter relationship.
+C.3 Testing Heterogeneity: Accessibility to Retailers
+The approximate locations of retail shops for daily necessities are reported in the official document and shown in Figure C.5.
+Density varies across lots: the number of retail shops per $1,000$ population is $8.8$ in the Tsukuda area, $7.8$ in Lot 1, and $4.6$ in Lot 2.
+Lot 2 has a lower density because it is relatively newer than the others (Online Appendix A.3).
+First, I include the density of retail shops in each lot as a control variable in Equation 2.
+Because this variable is time-constant, I interact it with the time dummies.
+Column 1 of Table C.1 shows an estimate of 0.132-0.132, which is close to the baseline estimate of 0.128-0.128 (Table C.7).
+Second, I expand Equation 2 by adding an interaction term between the head’s earnings and retail shop density.
+In this expanded model, a statistically significantly negative coefficient on the interaction term would indicate that households in denser areas rely more on credit purchases.
+However, Column 2 shows that the estimated coefficient on the interaction term is 0.004-0.004 and not statistically significant.
+In Column 3, I use an alternative definition of density, adopting the number of households as the denominator to account for variation in family size.
+The estimated coefficient on the interaction term remains near zero (0.002-0.002).
+In both columns, the marginal effect of the head’s earnings is estimated at 0.137-0.137, similar to the estimate in Column 1.
+The total length of the residential area in Tsukishima is roughly $1.4$ kilometers with a width of $500$ meters (Chūō Ward 1994, p.
+38).
+Thus, although the living areas are divided into three lots, residents can easily access shops in different lots on foot (Department of Health, Ministry of the Interior 1923a, pp.
+40–46).
+Figure C.5: Retail Shops for Daily Necessities: Food and Fuel Retailers
+Notes: This map shows the locations of retail shops for daily necessities (seikatsu hitsujyuhin hanbaiten) in Tsukishima.
+Cross mark: rice shop (kome ya).
+Circle with a cross: liquor shop (saka ya).
+Black triangle: greengrocer (yao ya) Circle with a minus: fish store (sakana ya).
+Black square: tōfu store.
+Half-black circle: firewood and charcoal store (shintan ya).
+Square with a cross: butcher store (niku ya).
+White circle: sugar, pickles, and dried food store (satō, tsukemono, kanbutsu ya).
+Black circle: tsukudani shop.
+tsukudani is a food simmered in sweet soy sauce.
+Half circle: tea shop (cha ya).
+White triangle: goshikiage shop.
+Note: Lot 3, the southernmost lot, had no households in 1919.
+Goshikiage ya is a fried fish paste containing vegetables.
+Source: Department of Health, Ministry of the Interior 1923c, fourth map.
+The tone was adjusted by the author using Adobe Photoshop 26.6.1.
+Table C.1: Testing Heterogeneity: Density of the Retailers in Each Lot
+Dependent Variable: Net Credit Purchase (income minus expenditure, yen)
+Semi-monthly Panel Adjusted Monthly Panel
+(1) (2) (3) (4) (5) (6)
+Head’s Earnings 0.132-0.132*** 0.107-0.107 0.099-0.099 0.026-0.026 $0.010$ $0.013$
+$[0.044]$ $[0.117]$ $[0.135]$ $[0.061]$ $[0.160]$ $[0.190]$
+Head’s Earnings $\times$ Density 0.004-0.004 0.002-0.002 0.003-0.003 0.001-0.001
+$[0.020]$ $[0.007]$ $[0.026]$ $[0.009]$
+Marginal Effect ($\partial r_{it}/\partial\tilde{y}_{it}$) 0.137-0.137*** 0.137-0.137*** 0.008-0.008 0.008-0.008
+$[0.045]$ $[0.044]$ $[0.051]$ $[0.051]$
+Density $\times$ Time FEs Yes No No Yes No No
+Family Size Yes Yes Yes Yes Yes Yes
+Denominator of the Density Population Population Households Population Population Households
+Observations 279 279 279 120 120 120
+***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. Robust standard errors are reported in brackets.
+Notes: Columns 1 and 4 report the estimated coefficient on the head’s earnings in the expanded regression of Equation 2, which includes interaction terms between retailer density and time dummies as additional controls.
+Columns 2, 3, 5, and 6 present results from the expanded regression of Equation 2, which includes the interaction between the head’s earnings and retailer density.
+The density variable in Columns 2 and 5 is measured per population, whereas it is measured per household in Columns 3 and 6.
+Columns 1–3 show results for the semi-monthly panel, while Columns 4–6 show results for the adjusted monthly panel.
+“Marginal Effect ($\partial r_{it}/\partial\tilde{y}_{it}$)” indicates the marginal effect of the head’s earnings at the mean of the density variable.
+All regressions control for family size.
+Households with missing residence information are excluded from the analytical sample.
+Results remain unchanged if these households are assumed to reside in any of the three lots.
+A few households living on the mainland in Kyobashi Ward are assigned to the nearest lot of Tsukishima.
+Sources: See the main text for the THBS sample.
+Data on the number of people and households in the three lots are from the Department of Health, Ministry of the Interior (1923b, p.
+2).
+The number of retail shops in each lot is from the Department of Health, Ministry of the Interior (1923c, fourth map).
+C.4 Testing the Roles of Cooperatives and Pawnshops
+Table C.2: Testing Roles of the Cooperatives and Pawnshops
+Panel A: Estimates from the Baseline Specification
+Dependent Variable:
+Prices of Purchased Items
+from Cooperatives (yen) Income from Pawnshops (yen)
+(1) (2) (3) (4)
+Head’s Earnings 0.001-0.001 0.006-0.006 $0.000$ $0.005$
+$[0.002]$ $[0.004]$ $[0.002]$ $[0.006]$
+Type of Panels Semi-monthly Monthly Semi-monthly Monthly
+Family Size Yes Yes Yes Yes
+Observations 289 124 289 120
+Panel B: Testing Heterogeneity using Density of the Pawnshop in Each Lot
+Dependent Variable: Income from Pawnshops (yen)
+Semi-monthly Panel Adjusted Monthly Panel
+(1) (2) (3) (4) (5) (6)
+Head’s Earnings 0.003-0.003 $0.013$ $0.018$ $0.001$ $0.113$ $0.101$
+$[0.003]$ $[0.022]$ $[0.025]$ $[0.004]$ $[0.126]$ $[0.115]$
+Head’s Earnings $\times$ Density 0.079-0.079 0.032-0.032 0.686-0.686 0.182-0.182
+$[0.129]$ $[0.044]$ $[0.753]$ $[0.203]$
+Marginal Effect ($\partial r_{it}/\partial\tilde{y}_{it}$) $0.001$ $0.001$ $0.006$ $0.007$
+$[0.002]$ $[0.003]$ $[0.009]$ $[0.010]$
+Density $\times$ Time FEs Yes No No Yes No No
+Family Size Yes Yes Yes Yes Yes Yes
+Denominator of the Density Population Population Households Population Population Households
+Observations 279 279 279 120 120 120
+***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. Robust standard errors are reported in brackets.
+Notes: Panel A: This panel reports results from Equation 2, using prices of purchased items from cooperatives (Columns 1–2) or temporary income from pawnshops (Columns 3–4) as the dependent variable.
+Columns 1 and 3 use the semi-monthly panel, while Columns 2 and 4 use the adjusted monthly panel.
+All regressions control for family size.
+Panel B: Columns 1 and 4 present the estimated coefficient on the head’s earnings in the expanded regression of Equation 2, which includes interaction terms between pawnshop density and time dummies as additional controls.
+Columns 2, 3, 5, and 6 present results from the expanded regression, accounting for the interaction between the head’s earnings and pawnshop density.
+The density variable in Columns 2 and 5 is measured per population, while in Columns 3 and 6, it is measured per household.
+Columns 1–3 show results for the semi-monthly panel, and Columns 4–6 show results for the adjusted monthly panel.
+“Marginal Effect ($\partial r_{it}/\partial\tilde{y}_{it}$)” indicates the marginal effect of the head’s earnings at the mean of the density variable.
+Households with no address are excluded from the analytical sample.
+Results are unchanged if these households are assumed to reside in any of the three lots.
+A few households living on the mainland in Kyobashi Ward are assigned to the nearest lot of Tsukishima.
+Sources: See the main text for the THBS sample.
+Data on the number of people and households in the three lots are from the Department of Health, Ministry of the Interior (1923b, p.
+2).
+The number of pawnshops in each lot is from the Department of Health, Ministry of the Interior (1923c, 15th map).
+In the THBS sample, five households used cooperatives to buy daily commodities in a few semi-monthly cells.
+Although this is a small portion, cooperatives may have offered some compensation to their members.
+To examine this possibility, I use the total prices of items purchased from cooperatives as the dependent variable in Equation 2.
+If such purchases increase following an idiosyncratic shock, the estimated coefficient on the head’s income would turn negative and become statistically significant.
+Columns 1 and 2 in Panel A of Table C.2 show that the estimates are negative but close to zero and statistically insignificant in both the semi-monthly and adjusted monthly panels.
+Similarly, a few THBS households used pawnshops in a couple of semi-monthly cells.
+I examine the role of pawnshops using the nominal income from pawnshops in Equation 2.939393Nominal income is used instead of net income because the THBS budget book summarizes payments in a column labeled payment (bensai), which includes amounts paid not only to pawnshops but also to other lending institutions.
+If workers rely on pawnshops to mitigate hardships, the estimated coefficient on the head’s income would be negative.
+Columns 3 and 4 in Panel B show that the estimates are near zero and statistically insignificant in both the semi-monthly and adjusted monthly panels.
+This result aligns with the main finding for the borrowing category (Table 5).
+In Panel B of the same table, I examine potential heterogeneity in pawnshop usage following the approach in Table C.1.
+The official report provides the locations of pawnshops, allowing me to calculate their density, measured as the number of pawnshops per $1,000$ people, in each lot: $0.13$, $0.18$, and $0.16$ in the Tsukuda area, Lot 1, and Lot 2, respectively.
+Column 1 includes density as a control variable in Equation 2.
+Column 2 adds an interaction term between the head’s income and the density variable.
+Column 3 tests sensitivity by using the number of households as an alternative denominator for density.
+In all columns, the estimated coefficients are near zero and not statistically significant.
+Columns 4–6, using the adjusted monthly panel, produce the same results.
+This pattern aligns with the findings in Table C.1, consistent with Tsukishima’s geographically narrow layout.
+C.5 Retail Sales in Tokyo City: A Retail Store Survey of 1935
 Table C.3 summarizes the contribution of regular customers to their total sales, social classes of the customers, and sales methods in the retailers selling white rice, fish, fruit and vegetable, and fuels (firewood and charcoal) by scale of business.
@@
 Total 100 51.1 15.7 50.9 33.4 20.4 78.4 1.3 8.6
-Notes:
-This table summarizes the statistics on the sales in the retailers selling white rice, fish, fruit and vegetables, and firewood and charcoal in Tokyo surveyed in 1935.
+Notes: This table summarizes the statistics on the sales in the retailers selling white rice, fish, fruit and vegetables, and firewood and charcoal in Tokyo surveyed in 1935.
 1. Small, middle, and large-scale businesses include retailers with annual sales of less than 10,000, 10,000–30,000, and 30,000+ yen, respectively.
-2. The number of stores reported the share of regular customers (%) are 90, 56, 61, and 93 in the white rice, fish, greengrocers, and firewood and charcoal categories, respectively.
+2. The number of stores reporting the share of regular customers (%) is 90, 56, 61, and 93 in the white rice, fish, greengrocers, and firewood and charcoal categories, respectively.
 3. The share of regular customers is defined as the percentage of sales for regular customers in the annual total sales in 1935.
@@
 6. The number of competitors indicates the number of peers within a three-cho (approximately 327 meters) radius.
-Source: Tokyo City Chamber of Commerce and Industry 1937, pp. legend; 20–21; questionnaire.
+Source: Tokyo Chamber of Commerce and Industry 1937, pp. legend; 20–21; questionnaire.
 In Table C.4, I analyze the relationship between credit purchases and competitiveness among local retailers using the selling item by scale of business cells reported in Table C.3.
@@
 Given that the standard deviation of the local competitors is approximately $2.3$, the estimated magnitude corresponding to the one standard deviation increase would be $7.9$ ($3.420\times 2.3$).
-This is roughly $41$% of the standard deviation of the dependent variable ($19.2$) and thus, is an economically meaningful magnitude.
+This is roughly 4141% of the standard deviation of the dependent variable ($19.2$) and thus, is an economically meaningful magnitude.
 Table C.4: Use of Credit Purchases and Competitiveness among Local Retailers
@@
 (1) (2) (3)
-Number of Competitors 7.1907.190***7.190 ∗ ∗ ∗ 7.2547.254***7.254 ∗ ∗ ∗ 3.4203.420**3.420 ∗ ∗
+Number of Competitors 7.1907.190*** 7.2547.254*** 3.4203.420**
 $[2.041]$ $[2.145]$ $[1.310]$
@@
 (1975).
-C.4 Retail Sales in Tokyo City: A Small and Medium Scale Enterprises Survey of 1930
-Table C.3 summarizes the use of credit purchase and the losses on the credit sales documented in the 1930 Survey Report.
-Following the report’s style, the item categories are divided into six categories: white rice, other grains, vegetables, fruit, fish, and fuel.
-The total number of retailers in each category is $157$, $7$, $67$, $16$, $46$, and $64$, respectively.
-As explained in Section 5, most of them used credit purchases as their selling method.
-Importantly, among the retailers using credit purchases, almost all of them suffered losses in sales.
-For example, $99$% of white rice retailers experienced losses.
-The minimum value of this share is still $97$% for the fuel retailers.
-This means that credit purchase was a loss-making sales method for the retailers.
-The statistics in Table C.3 also indicate that many cases have unknown loss amounts.
-Overall, however, most of the cases seem to be grouped into bins of less than $5$% of losses.
-For example, $87$% ($97$ retailers) of the $111$ available rice retailers suffered losses amounting to $5$% of sales.
-Similarly, $65$% ($32$ retailers) of the $49$ available fuel retailers are grouped into the same loss category.
-Systematic statistics are unavailable to indicate the percentage of total profits corresponding to these losses.
-However, it is worth mentioning that a large part of losses was capped by $5$% of total sales in all these different retailers.
-This implies that while the retailers could not avoid losses, they might have set prices that anticipated losses.
+C.6 Retail Sales in Tokyo City: A Small and Medium Scale Enterprises Survey of 1930
+Table C.5 summarizes the use of credit purchases and the associated losses documented in the 1930 Survey Report.
+Following the report’s classification, items are divided into six categories: white rice, other grains, vegetables, fruits, fish, and fuel.
+The total number of retailers in each category is $157$, 77, $67$, $16$, $46$, and $64$, respectively.
+These retailers used credit purchases as their primary selling method.
+Notably, almost all retailers using credit purchases experienced sales losses.
+For example, 9999% of white rice retailers reported losses, while the minimum share was 9797% for fuel retailers.
+This indicates that credit purchase was generally a loss-making sales method.
+Table C.5 also shows that many cases have unknown loss amounts.
+Overall, most losses appear to be concentrated in bins of less than 55% of sales.
+For instance, 8787% ($97$ retailers) of the $111$ rice retailers with available data suffered losses amounting to 55% of sales.
+Similarly, 6565% ($32$ retailers) of the $49$ fuel retailers fell into the same loss category.
+Systematic statistics on the percentage of total profits corresponding to these losses are unavailable.
+However, a large portion of the losses is capped at 55% of total sales across all retailer types.
+This suggests that, although retailers could not entirely avoid losses, they may have set prices to account for anticipated losses.
 Table C.5: Sales Methods and Losses on Sales under Credit Purchases in Tokyo City
@@
        21–30% 2 0 1 0 0 0
-              30+0 0 0 2 0 0
+       30+% 0 0 0 2 0 0
        Unknown 44 4 19 0 16 12
@@
 454–469.
-C.5 Testing Relationship between Volatility of Earnings and Use of Credit Purchases
+Lastly, Table C.6 summarizes the industrial and family structures in 1930 Tokyo using the official report of the Population Census.
+Column 1 summarizes the share of male workers in each industry, showing a similar distribution as that in 1920 (Column 1 of Panel A in Table 1).
+While the development of the city throughout the 1920s slightly increased the share of commercial industries, manufacturing and commercial industries covered approximately 75% (77% in 1920).
+Column 2 confirms that the average household size, share of married males, sex ratio, and average age of males are in a similar range to those measured in the 1920 Population Census (Column 1 of Panel B in Table 1).
+These show that not only the industrial structure but also the average household characteristics had not changed essentially between 1920 and 1930.
+While Tokyo experienced the Great Kanto Earthquake in 1923, the population structure did not change dramatically, except for a temporary evacuation soon after the earthquake.
+Table C.6: Industrial and Family Structures in Tokyo: 1930 Population Census
+(1) Industrial structure (2) Family Structure
+Name of survey 1930 Population census 1930 Population census
+Survey area Tokyo City Tokyo City
+Survey subject Complete survey Complete survey
+Survey month and year October 1930 October 1930
+Agriculture 0.5 Average household size (in person) 5.1
+Fisheries 0.1 Share of married males (%) 34.2
+Mining 0.0 Sex ratio (males/females) 1.2
+Manufacturing 37.2 Average age of males 25.2
+Commerce 37.3
+Transport 7.2
+Public service and professions 13.5
+Housework 0.6
+Other industry 3.6
+Notes: Panel A: This panel summarizes industrial structures based on the occupations of male workers recorded in the 1930 Population Census.
+Each share is calculated as the number of males working as regular employees in each sector divided by the total number of male workers (percentage).
+Unemployed males (mushoku) are excluded from these figures.
+Panel B: The average household size is calculated as the total number of people divided by the number of households.
+The share of married males is calculated as the number of married males living with/without spouses divided by the total number of males (percentage).
+The average age of males is calculated from the population tables by age group reported in the census.
+For the open-ended age category over 65, the value is rounded to $67$, following the range of the previous category (i.e., 60–64).
+All figures in this panel exclude a small number of quasi-households (jyun setai) such as students in dormitories and patients in hospitals.
+Sources: Column 1: Statistics Bureau of the Cabinet (1933, pp.
+52–53).
+Column 2: Statistics Bureau of the Cabinet (1933, pp.
+14; 6–7; 357).
+C.7 Testing Relationship between Volatility of Earnings and Use of Credit Purchases
 Table C.7: Volatility of Earnings and Use of Credit Purchases
@@
 (1) (2) (3) (4) (5) (6)
-CV of Head’s Earnings $0.003$ $-0.002$ $-0.027$ $0.019$ $0.008$ $-0.005$
+CV of Head’s Earnings $0.003$ 0.002-0.002 0.027-0.027 $0.019$ $0.008$ 0.005-0.005
 $[0.012]$ $[0.014]$ $[0.050]$ $[0.013]$ $[0.017]$ $[0.049]$
@@
 (1975).
-Column (1) of Table C.4 presents the result for Equation 4.109109109 The estimated coefficient is close to zero and statistically insignificant.
+Column (1) of Table C.7 presents the result for Equation 4.
+The estimated coefficient is close to zero and statistically insignificant.
 A potential concern may be the correlations between the volatility of earnings and the head’s age and earnings level.
@@
 This suggests that retailers may not have considered the volatility of earnings when providing credit to factory workers.
-This is consistent with the retailers’ motivation to secure their local customers discussed earlier, and with the fact that factory workers commonly used credit for purchases.110110110Despite this, low-income households might have had more difficulty using credit for purchases.
+This is consistent with the retailers’ motivation to secure their local customers discussed earlier, and with the fact that factory workers commonly used credit for purchases.
+[^94] Despite this, low-income workers might have had more difficulty using credit for purchases.
 A famous essay drawn by a primary school student in the 1930s depicts an example of a rice retailer who refused to buy rice under credit purchase (Toyoda 1969, pp.
@@
   • (3)
+  • (4) de Chaisemartin, Clement., and Xavier D’Haultfœuille.
+“Two-way fixed effects and difference-in-differences with heterogeneous treatment effects: a survey.” The Econometrics Journal 26, issue.
+3 (2023): C1–C30.
+  • (5)
   • (6) Cochrane, John H. “A simple test of consumption insurance.” Journal of Political Economy 99, no. 5 (1991): 957–976.
@@
   • (8) Choi, In “Unit root tests for panel data.” Journal of International Money and Finance 20, no. 2 (2001): 249–272.
+  • (9)
+  • (10) Horrell, Sara., and Deborah Oxley. Work and prudence: Household responses to income variation in nineteenth-century Britain. European Review of Economic History 4, no.1 (2000): 27–57.
   • (11)
@@
   • (49)
+  • (50) Susan D Horn, Roger A Horn, and David B Duncan. Estimating heteroscedastic variances in linear models. Journal of the American Statistical Association 70, no. 350 (1975): 380–385.
+  • (51)
   • (52) Susan Dynarski, Jonathan Gruber, Robert A Moffitt, and Gary Burtless. Can families smooth variable earnings? Brookings Papers on Economic Activity 1997, no. 1 (1997): 229–303.
-  • (47)
-  • (48) Susan D Horn, Roger A Horn, and David B Duncan. Estimating Heteroscedastic Variances in Linear Models. Journal of the American Statistical Association 70, no. 350 (1975): 380–385.
   • (53)
@@
 12 (2015): 70–80.
-  • (51)
-  • (52) Tanaka, Hikaru. Mouhitotsu no kinyū system (Another Financial System) [in Japanese]. Nagoya: Nagoya University Press, 2018.
   • (55)
@@
 Altman.
-“Regression using Fractional Polynomials of Continuous Covariates: Parsimonious Parametric Modeling.” Journal of the Royal Statistical Society.
+“Regression using factional polynomials of continuous covariates: parsimonious parametric modeling.” Journal of the Royal Statistical Society.
 Series C (Applied Statistics) 43 no.
@@
   • (25)
+  • (26) Ministry of Land, Infrastructure, Transport and Tourism (database).
+Gyōsei kuiki data (Administrative Area Data).
+[in Japanese] (https://nlftp.mlit.go.jp/ksj/gml/datalist/KsjTmplt-N03-v2˙2.html, accessed on August 15, 2023).
+  • (27)
   • (28) Ministry of Posts and Telecommunications. Denki jigyō yōran, dai jyūyon kai (Handbook of Electricity Industry, vol. 14). [in Japanese] Tokyo: Electric Association, 1922.
@@
 [in Japanese] Tokyo: Statistics Bureau of the Cabinet, 1929a.
+  • (41)
+  • (42) Statistics Bureau of the Cabinet.
+Shōwa gonen kokuseichōsahōkoku, fuken hen, dai yonkan (Population Census of Japan, Tokyo Prefecture, 1930 Edition).
+[in Japanese] Tokyo: Statistics Bureau of the Cabinet, 1933.
   • (43)
@@
   • (7)
-  • (8) The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (8–12).
-  • (9) The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (8–13).
-  • (10) The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (8–14).
-  • (11) The Ohara Institute for Social Research, Archives of Gonda Yasunosuke (8–15).
-  • (12)